A Complete Bibliography of Publications in Communications in Statistics: Simulation

and Computation: 2010–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

15 June 2022 Version 1.13

Title word cross-reference

100(1 − p) [DB18]. 12 [WLWT16]. ^{2} [Sei19]. ^{n} [KP10]. ^{p−1} [Aok15]. A [AJ17, HB10, YZZ17a]. α [KTA17, SN13b, XYL19]. AR(1) [UA17]. AR(m) [Tew17]. AR(p) [DK17b]. X̄ [CGST17, Kao17, KM19b, LK19, PMNF18]. β [Z.18]. C_{p} [US10]. C_{P M} [DZC10]. C_{pyk} [DSMJ18]. W [GS11]. χ^{2} [CDGML12, Liu13]. D [ATV^{+}14, AJ17, HB10, Mie12, PM12, VM11, WPLL15, YM15]. δ [WCSL17]. δ = P (X < Y) [CJTL18]. E [YZZ17a]. ϵ [ZM13]. F [ASSH10, AK19, EY17, KPIY16, Liu13, RSW12]. G [AK19, Ery15, FTE17, SL14, TN16]. GARCH(1, 1) [BB17d]. I(0) [Góm13]. I(1) [Góm13]. J [OWY18]. K [ZLH^{+}19, RLTJ16, AB13a, AY18a, ADLS10, BGU15, BS15, BSL15, BKJ10, BJ14a, CM15, DK17b, DK17c, DN16, EY17, EK10, Ery10, Ery15, FTE17, HSP^{+}17, KFO14, LKTW12, Naj17, Naj19, Ng17, PSHP17, PRA14, QJ14, REM19, ZWCC12, ZW14, ZYS19]. k − 1 [Wu18b]. L [SP17a]. L^{p} [SJT19]. L^{p}(1 < p ≤ 2) [WLF19]. L_{1} [ET16, HW14]. L_{1/2} [AA18b]. L_{2} [BSK19, Zha15c]. l_{p} [LNT17]. Λ [CBDD16]. LQ [SP17a]. M [DZ17, EKG17, GTPS14, HM12b, JK16, PKK12, She13b, Sun17, Ert18, EK10, KFO14, LKTW12]. M^{[X]}/G/1 [SSOTB19]. MA(q) [DK17c]. m: G [ZW14]. n [BD17a, Bay11, EY17, Ery10, Ery15, FTE17, HS12, HN15, KM19a, Sab18, ZYS19]. n: F [BGU15, EK10, KFO14, QJ14]. ω_{t} [SN13b]. P [Koc17, AE17, AL15, Aok15, HN15, Mar15, TGSA18, WY15b, XL18, ZX17]. P (X < Y) [Gre17]. p ≥ 3 [Liu17a]. Φ_{p} [FT16]. Pr(X < Y) [Bak16]. Pr(Y < X) [AVR17]. p → 0 [DB18]. P {X < Y } [Jov17]. Q [LDK17]. R [AMP10, Ban17, Kao17, BD17a]. R^{2} [CXC18]. RCINAR(p) [WZ10]. s − K [HWX16]. t [AOK13, AMJ17, CS19, CMM12, CLW12, HN13, Hod13, KPV17, Liu13, MN18, Ndu18, RNJ16, SBNG18, SMMA19, SMY12, Wu14, WZTX16]. T^{2} [NA12, PKC^{+}11, SSA18, SKBA18, YSH19, YKLR14, YKSC16]. t_{2} [AN12]. t_{3} [AN12]. θ [SN13b]. T L [SP17a]. U [KK10b]. VARFIMA(0, d, 0) [LPZ13]. X [NGT14, YC10]. y [WLF19]. z [WLF19]. |S| [LK17a, LK18a].

-Bar [NGT14, YC10]. -best [Ng17]. -chart [AMJ17, SL14]. -charts [CGST17]. -control [MKFP18, KM19b, PMNF18]. -copulas [KPV17]. -Distributed [HN13, MN18, Ndu18]. -estimates [ET16]. -estimation [DA18, DZ17, PKK12, Sun17]. -estimator [Ert18, EKG17, JK16]. -estimators [She13b]. -Factor [SMY12, DN16]. -FCvM [HW14]. -inflated [Z.18]. -Means [ADLS10, BJ14a, REM19, ZLH^{+}19]. -mixing [XYL19]. -moment [SP17a]. -norm [LNT17, BSK19]. -Normal [Wu14]. -of- [LKTW12]. -Optimal [ATV^{+}14, Mie12, VM11, AJ17, YM15, YZZ17a]. -optimality [FT16, WPLL15]. -Optimum [PM12]. -Out-of- [BGU15, EK10, KFO14, BD17a, EY17, Ery10, Ery15, FTE17, ZYS19, QJ14, ZW14]. -player [Sab18]. -Procedures [HM12b]. -prototypes [AY18a]. -Quantile [GTPS14]. -Records [AB13a]. -Rotatability [HB10]. -series [KTA17]. -shock [WCSL17]. -Spatial [BKJ10]. -Squares [ASSH10]. -Statistic [AOK13]. -Statistics [KK10b]. -SVR [ZM13]. -Test [CMM12, SMMA19]. -Tests [RSW12]. -type [WZTX16, Ban17]. -unit [PRA14]. -Value [AL15, XL18]. -Values [WY15b, AE17, Mar15, ZX17, Koc17]. -Within- [EK10, KFO14]. -year [WLWT16].

1} [Ban17, IOTBO17, MSR12, OTZA17, ZX10].

2 [ABAJ14].

5th [CD14].

65 [Grc17].

A-Optimal [SH12]. A20 [CZL16]. A20-year [CZL16]. Aalen [She14]. AB/BA [Lui17a]. abilities [Yao18]. Ability [BAPA15, Yam16]. Absence [cCyYnH13]. Absolute [Hay14, Van15]. Abundance [LA13]. accelerate [CTXW17]. Accelerated [AHAH15, BZL15, CSY17, CW12a, GTFX14, Jin11, LLZ16, PB10, PBS11, PS14, SHM12, TY16, WCL12, XTG14, ZHL14, BC17, CXC18, DOZ17, FS17, Lin17, NP17, SS16, WGX18, WSZ17]. Acceptance [ABAJ14, DS13, GA17, Bal17, ML19b, MNBB17]. acceptance-rejection [ML19b, MNBB17]. accident [DARMM^{+}17]. accompanied [SGNS17]. Account [GBO14, MS14]. Accounted [BH16]. Accounting [BL15, Roy12, SM17, YLM17]. Accumulated [HdL10]. Accuracy [JMCT16, SB17, TS17, TS18, ZK18]. accurate [Dem17b]. achieved [CTGR17]. Achievements [RM16c]. Activity [RR13, ZKL^{+}19]. Adaboost [LqQlX11]. Adaptation [De 11]. adaption [STLH17]. Adaptive [AM14, AZCS10, ANA14, BM10, BD14a, BJ14b, DCM16, FZDF14, KJ13, KYP11, KK10b, Lee10b, LZZW13, LGM12, ML19a, ND10, NCF15, PLW^{+}14, Ste12, Wan17, YLL17, ZLZW16, AM17a, AHM17, CBB^{+}17a, HK17, KL17b, Mor19, TTC19, XT19, YCS17]. Additive [ASM15, CW12b, LLX16, MMM16, Rod16, SP18a, She13a, She14, She15a, UHP11, YGW16, ZX13, dJvBS16, FCC18, LYG17, MC19, UA17, WB18, WX19]. address [DRAM18]. adjacency [GPM18]. Adjacent [OBD16, GPM18]. Adjacent-category [OBD16]. adjust [BLC17]. Adjusted [HM12a, Hua14, Hua17, LqQlX11, ZGC16, AK17, AAC17, HKNC17, KA19a, LDK17]. Adjusting [HHH16, PB17b]. Adjustive [Ron10]. Adjustment [Kna19, MLZ15, Mak11, PL13, YKY10, CB17, WY19, WB18]. adjustments [MMV18, MO17]. adopted [ASKT19]. Advanced [CD14]. Advances [FP14, Tri18a]. Advantages [GBB^{+}12]. Aeppli [BYAJ17, CM15, CM19, QLAJB17]. affect [Wes17]. Affected [MSS10a]. affecting [Bah18]. affine [MO18]. AFT [BCGV13]. After [BA13, BA14, GE17]. AG [Bia17a]. Against [BAM11, Che15, LS10, OK15, ZTT10, KPG^{+}17, LXPL17, SFS12]. Age [BSV13, LZZ16]. Aggregate [KR12]. Aggregated [AS12c]. aggregation [TS17, TS18]. Agreement [TKX13]. agriculture [EA17]. AIC [HWD18]. Aided [YSWM13]. AIDS [HD13b]. Air [BNKM15, ABH17, BBA17]. airborne [Mor19]. Alarm [BGM13]. Alcohol [McI14, HBC17]. alert [AD19]. Algorithm [AM14, ADLS10, ART^{+}12, CCL10, CK14, Ell11, FT16, FWY14, Far16, HN15, JZ14, KK14b, Kom15, Koz14, MTL10, Mey13, MM15, MNH10, NGT14, Pap14, PK16, Sat12, SID13, Wan14, XC15, XS11, XZ16, YBW^{+}12, ZM13, hZxNL14, BHF17, CHJL17, CDE19, CCH18, CX17, DI17, DGMC18, GRM17, LL17a, LTZ17, LL17b, MZ17, ÖA19, PB18a, PGCT17, SF19, SKLK19, TRM18, Wal17, WSCW18, YY17a, YY17b, ZLH^{+}19, Zha18, ZHKR17]. Algorithmic [DGH19, GQYY18, MX17a]. Algorithms [FS12, FG15, HN13, KWW14, LDS10, Pic13, VM11, VKB10, AB19, AY18a, MN18, MNBB17, Ndu18, PLT18]. aligned [Lue17]. All-to-Next [SH12]. Allocation [AN11, KGA13, WL15a]. allowed [Sab18]. Allowing [LLZ16]. Almost [WY14, Wu16, Das17, WA18]. Alo [Hou16]. Alo-group [Hou16]. alpha [AHB19]. alpha-divergence [AHB19]. Alternation [SOSB14a, SOSB14b]. Alternative [AK11, BAM11, CY12, EB14, GAPSÁ10, HHH16, LYY14, MH14, OK15, RR13, YA17, ZTT10, BDL17, CB17, CS17b, DGMC18, GGK17, GSP^{+}17, KR17, MR18, NWR17, ÖA19, YL17, ZSY17, dSBHM19]. Alternatives [AR12, JKBR16, RB16, XD15]. Always [GBO14]. American [NMS18]. Among [HTN14, TNS12, YHL12, GCR12, LLH17, RAAK18]. Among-Group [YHL12, LLH17]. amount [HZNN17]. Analyses [KP10, WBG16a, YKSC16, AMKS19, LDK17, WMH17]. Analysis [AT16, ADLS10, AKV16a, AC15a, AU17, Ati16, BGP14, BCV15, BG16, CP16, Che14, CCLT14, CDGML12, CJV13, DD16, Ema13, ES16, GGNB10, HM12a, HK13, HA12, HHHZ10, JMCT16, JH14, Jin11, KK10a, KMSK15, KSJ^{+}12, Kim14, KS16, KSGM13, lKW13, Li12, LDW14, LLZ18, LRVF10, LZ16, MNZ13, McI14, MPEGMGMR11, MS14, ONS16, PGGF14, PWT11, PLLF16, PSS10, PAM15, PSO13, Pic13, RR13, SNM16, SKZA16, SW11a, SZFM11, SWL14, SL17, TK15a, TZ14, UA17, WTCZ10, WL11, jWzM13, WP13, WMK14, WTLWL16, WMK16, XTG14, YSWM13, Z.18, ZP14, ZW14, ZW15b, dWD16, AK17, AKZK17, AO17, Ban17, BB17e, BVC^{+}19, BSK19, Car19, CKP17, CK19, CTH17, DL17, DKA18, DWW18, DGMC18, EGO^{+}17, Ery17, FN17, FT17, FNA17, FA17, GGP19, GLGG17, Grc17, HK18a, HT17, HBC17]. analysis [HZNN17, HW17b, KJK17, Koc17, KC17, LL18, LNUT17, LN17, Lan17, Lue18, MCM16, Maz17, MP17b, MP19b, MCC18, MMB19, NK17, NFM19, PTVP19, PFOC17, PUN17, PR19a, PR19b, RG18, SSOTB19, SU18, She17b, SS17b, SDS17, VRN18, Vil17, WGX18, WLX18, WC17, WSZ17, WMK19, XYZ17, ZLA17, dSBHM19]. Analytic [GG19]. Analytical [BR16b, PAS19, SA17a]. analyze [Fan18]. Analyzers [SMY12, YXY17]. Analyzing [BG12, FL13, KMS11, KP12, LW11b, XLC14, DKL18, DK19, FKG19, GS17, NC19, SAD17, ZLW17]. and/or [BFB17, dODD19]. Anderson [Ras17]. Animation [ZW14]. ANOVA [CS19, DI17, FZ17, LZ16, RSW12, WTCZ10, WTZ11, XYCY15, Zha15a]. ANOVA-type [FZ17]. Answering [WTLWL16]. Anticipated [Fen16, HWZ18, SJT19]. Any [DH16, MMN12, BK17]. anything [BBK17]. APARM [CD14]. Appication [ZWZC12]. Application [Bon16, BSFG10, CLO13, CCM^{+}16, CC10, Che11, CK14, CCD^{+}12, DCM16, ES16, EA17, GVC10, GDGG16, HB15, Hou16, HD13b, Isl11, KK10a, KMSK15, KPIY16, LMS13, Pra15, Rut15, RB16, SN13a, SSK16, TN19, TL15, XZZ15, XS11, vZ11, Abd18, CL22, CCH18, CX17, DARMM^{+}17, EGO^{+}17, FNA17, Gun17, HA18, HC17b, KPG^{+}17, LL19, LL18, MKS18, PSHP17, SMS17b, Tas17, Usa17, WDC18, ZKL^{+}19, vNBA17]. Applications [ABB12, BZ15, DKV15, EK12, HXX15, KJR13, Kri16, Li12, Li13, Mey13, SNM16, SI13, TCA^{+}16, TM14, WLWT16, Xu14, ABS^{+}18, ARM^{+}19, BS17a, Ben17, CC19a, CBH19, CMN17, CMBS18, Ery18, GDSCO19, KKH17, LNT17, LLB19, LWT17, MJ17, RTJ17, SNT17, SF19, SBS17, TB17, Vin17, YAR^{+}19, ZLW17]. Applied [PC13, dMBSFB17, CR17, MAD17, MHAD19].

Applying [Van15]. Appraisal [GGNB10]. Approach [ANA14, ALH14, Asi14, BHI12, BG12, Bon16, BR16b, CCM^{+}16, CIB10, CY16, CRS13, DAM15, FL13, GKA12, HD13b, JA15, KGK14, LC12, LNL11, LW16, MYS^{+}16, MSS10b, MG12, Man15b, Mar13b, MH14, MS10, NQ15, Rai16, RSC16, RM16c, RAN10, RAN15, SMB12, SKS16, SG12, SCH14, lTHT13, TA13, TY16, YSWM13, YZ11, YC16, ZHL14, Zha15a, ZAK14, Baz17, BR18, dMBSFB17, CTXW17, CR17, Che17, CN19, DZT19, Fan18, GGB17, GG17b, GP17, GB19, HZHK17, HW17b, JK17a, LSY17, LXPL17, MV19, MRF17, MMB17, NÖ18, NBOM17, PGH^{+}18, RG18, RH17, Sei19, SMS17a, SAD17, WLZL19, XY17b, XL18, XZ19, YE17, YA17, YSS17, ZSZ17, Zuo19]. Approach-based [RSC16]. Approaches [ARW16, CMM14, HHH16, KO16, Mah12, MMF14, MSB11, RLF11, Wan12, YT12, KPG^{+}17, MH17b, Sta17]. Approximate [EFA17, Mar13b, RSW12, VL13, Vol14, WA14, AR17a, Bah18, RSR17]. Approximating [AFM12, Liu17a, Ouy13, RCM17]. Approximation [AOK13, AL15, AC15b, Bia17b, Bia17a, DPCJ16, GP16, GKA19, Kri16, LSY16, RCR14, SV13, VHV^{+}16, WSW16, XX16, Bar17c, EL18, JU19, NFM19, Pic19, RK17]. Approximations [AL16, BS16a, Ima11, LYY14, RV16, SR12, Sta12, SW13, Bee17]. arbitrarily [Lin17]. Arbitrary [Ery14]. ARCH [Iqb11, JTY10, Lak11, LAG17]. ArcTan [GDSCO19]. Area [CS12, CSG15, GTPS14, LQW^{+}16, PC13, KKGK19, Rah17, SDSM18]. area-biased [SDSM18]. Area-Level [PC13]. Areas [ARW16, cCyYnH13, ABH17, MAD17]. ARFIMA [PS18b]. ARIMA [DK17b, DK17c, KH19a]. ARIMA/MLP [KH19a]. ARL} [PAS19]. ARMA [CW14, LXPL17, MAD17, TS18, YLL17]. ARMA-GARCH [YLL17]. array [EZ17]. Arrays [AAK13, WTZ11]. Arrival [AC15a, BR18]. Artificial [CCC16b, CX17, MERN18, YDA19]. Asbestos [YBW^{+}12]. Asbestos-Related [YBW^{+}12]. Asia [CD14]. Asian [WSQ15]. Aspect [BSFG10]. Aspects [AJB11, FP14, KR17, ROC^{+}18]. Assay [vdH10]. assess [ZKL^{+}19, dLHT19]. Assessing [Alk17, CMC^{+}17, HM17, HM13, TKX13, ZK18, GE17, LLWY17, PGH^{+}18]. Assessment [CCM^{+}16, CN11, FMMV16, RYW16, CTXW17, JS19b]. assignable [MKFP18]. Associated [CDL12, JS14, Kom15, LX13, XY17a, ZKL^{+}19]. Association [PLLF16, ZF16, DL17, KA19a, ZLGS18]. Assuming [OY16]. Assumption [AOK13, CDS15, KU14, LZZ16, TVS13, OAB17, SU18]. Assumptions [MSK10, MN15, HKNC17]. assurance [CH17]. assuring [BJU18]. Asymmetric [CDL12, HS12, PWT11, RYW16, RR12, RB16, Xu14, AM17b, BPJ17, KH17, Sab18, SMS17a, YSY19]. Asymmetry [Hol10, Tas17]. Asymptote [DM14]. Asymptotic [Dar12, KX17, MP17a, SS10, SR12, Wil12, EA17, HA17b, Zim17]. Attributable [RT12]. Attribute [BAA16, ABJM17, Asl19, SK17c]. Attribute-variable [BAA16]. Attributes [LC16, LS14, BB17a, GA19, RA17]. AUC [LLJ^{+}12]. Augmented [MMB19]. Australia [Rah17, ZLW17]. Autocorrelated [AS15, HXB13, KJS10, PB16, RMS10]. autocorrelation [CA17, MAT17, Res19, SGNS17, Sun17]. automatic [CCH18]. Autoregressive [AB13b, Asi14, ART^{+}12, BM10, BD14a, BD14b, CW12b, DH12, EMH15, ET16, HN13, HHKM16, LTQ15, LM16b, AM17b, Ami19a, Ami19b, BM17a, Elh17, JT19, JU19, KL18, LN17, MHAD19, MN18, MC19, MWSB18, Ndu18, Ott17, PB18b, SS19b, SS17b, SSY18, TSL^{+}19, YA17]. Auxiliary [AS14b, CSG15, CWYF12, GK14b, GSSCR16, KGS14, SS14, SG14, SSA15, ASS18, KAAJ19, LSD17, Mar19, MSS17, PS18a, SAAJ19, SG17, Wes17]. Availability [LKHH11, GLGG17, LKYH17]. Available [DAI14]. AVC [VKB10]. Average [AS12b, HXB13, Lee12, MW10, SFLC14, WSQ15, Yan13, AHM17, AGKJ17, AAMJ17, Car19, CHB17, JU19, MHAD19, NH18, PB18b, PW17, Roo17, Tas17, Wu18a]. average-EWMA [AGKJ17]. averaged [Liu17c]. Averaging [FWY14, AM17b, FLW19, XZ19].

B [Din18]. B-splines [Din18]. BA [Lui17a]. Back [PV17]. backfitted [WX19]. Backfitting [XY12]. Backoff [LGM12]. Backward [Fen16, HWZ18]. Bag [ZZZ10]. Bagging [WZG15]. Baker [ZNA16]. Balakrishnan [Ate17]. Balanced [Man15b, SS12a, YAA15, GPM18, JPV^{+}18, KAS^{+}19, RRA^{+}19b]. balanced-Repeated [KAS^{+}19]. balancing [TK17]. Band [GVC10, Chv17, Das17]. Bands [ZGC16]. Bandwidth [MWH10, Ste12, ZH19, MRN17]. Banerjee [KM19b, YC10]. Bar [NGT14, YC10]. Bartlett [LBP17, MLZ15]. Based [AHAH15, ASSH10, AJM10, AY15, AB13b, ALH14, AZSA11, ARAM15, BHI12, BDKM11, BAA16, BR16a, BDS10, CP16, Che14, CZL16, CMB11, CYW13, Chi15, CCD^{+}12, De 11, DAA13, DC15, EK16, FG15, FK15, GCR12, GAPSÁ10, Góm13, HN13, HJST15, HLLV14, Hou16, Hua14, IMS14, Ima11, JS14, KHH12, KU14, KM13, KJS10, KNJ16, KK10b, KP12, KGXR16, LDW14, LZZ16, LHB08, LHB10, LBW11, LZS10, LqQlX11, LZZW13, LH13, LZZ15, LKTW12, LZ16, Mah15, MA16, Mur15, NB15, NA13, PGGF14, PB10, PS14, Par11, Par15, PKC^{+}11, PRA14, QMLS16, SKST14, Sat12, SB12b, SVWH10, SNO16, SA16, SW11a, SS10, SO10, Sim11, SG12, SMY12, SW13, TK15a, TKCL16, TKKC14, TK15b, TT15, VG15, VG10, VAGE15, WTCZ10, WL11, WTZ11, WBG16b, WGS10, WLL10, WL15a, Xu14, XZ16]. Based [YC16, yYjHwK14, Zgh10, ZA10, ZJNG13, ZDX14, ZW15b, Zha15c, ZWZ14b, ZLGZ15, ZTT10, AB18, ARY18, AHB19, AO17, AS17, ABH17, AGAB17, Ate17, AMKS19, AY18b, BHF17, BG17, BSLF19, Bal17, BH17, CHJL17, CC19b, CN19, CX17, CQH17, DARMM^{+}17, DBM17, DAY19, DDD16, DKV15, DB17, EDAAAM17, ELS17, EY17, EA12, EKG17, EG16, FN17, GK18, GRM17, GS17, GG17b, GLGG17, Gre17, GE17, GA17, HPS17, Hua16, HZ19, Hwa17, Jov17, KAN17, Kao17, Kao18, KE17, KLK19, KEB17, Kiz19, KL17b, KR18, Kri16, KOC16, LL19, LTZ17, MO18, MO19, MAT17, MMP16, Maz17, MH17b, MWK16, MM16, MN18, MM17, NAJ18, NP17, NÖ18, Ndu18, NS18, NWB^{+}19, ND17, Nou17b, Nou17c, Nou17d, ÖEG17, OWY18, PB18a, Pal17, PB18b, Par17, PS19b, PW17, PR19b, RSC16, RSR17]. based [RRR19, RBX^{+}15, RSBZ18, RPRPMG^{+}18, RM16c, RH17, RGP17, SBH18, SAAJ19, SMMA19, SFC19, SA17b, SXYF17, SC18b, Sun17, Tas17, TNO19, Tri18a, Ul 17, VAR19, WBG16a, WLX18, WMH17, Wu17b, Wu17a, Wu18a, Wu19, YLA17, YCL17, YWZ18, YSY19, ZSZ17, ZSY17, ZHLW17, ZLHW19, ZHKR17, ZM17, Zim17, dMdCQSH17, vZ18]. Baseline [MR14, ZLS14, YW17b]. Baselines [ZCR10]. Basis [Aok15, DPCJ16, KK10a, BVC^{+}19, CDE19]. bathtub [SKK17]. bathtub-shaped [SKK17]. Bayes [BBF13, Bay11, CKC^{+}14, DAA13, GP17, GR17, HKP18, KK17a, KU14, MOS18, OWD14, RK17, SU18, TTC19]. Bayesian [LHB10, AT16, AARQ19, ATV^{+}14, AGM19, AM17a, Agh18, AM17b, Alh16, AA18a, AA18b, AA19, Alk17, AC18, Ami19a, AJ17, AR17a, ALH14, ABH17, ASM15, AHGJRC16, BS10, CIB10, CCC16a, CK19, CSL17, CTH17, CL18, CNK17, CQH17, DSR15, Din18, Don10, DWW18, DZT19, DGMC18, EZ15, ESSH18, FWY14, FK14, FGMW10, GLGG17, Gut15, HK18a, HZHK17, Han19, HD13a, HE17, HNW17, HFKM16, HD13b, HSR10, JA11, KSS17, KGK14, KK17b, KJK17, KRK18, Kiz19, KBAD11, KL17b, KR18, KSGM13, Kyu16, LC12, Lak11, LAG17, Lan10, Li10b, LL18, LXPL17, LHB08, LW11b, Lin17, Liu17c, LN17, Lan17, MYS^{+}16, MV19, MMM16, Mar13b, Maz17, MP17b, MSR12, MC19, MMB19, ML17, MS14, NP17, OG14, PLW^{+}14, PLT18, PLA16, QEK17, RZZ10, RF10, RW17, RLF11]. Bayesian [RV16, SGKM17, SS17a, SKK17, SK17b, SB12b, SSK16, SS17b, SK17c, SXYF17, SSY15, TK15a, TZ14, TA17, TSL^{+}19, TTC19, Usa17, VND19, VB16, WL11, jWzM13, WA14, XTG14, XZ17, XT19, YW16, YE17, YCL17, Yao13, YZ11, YZW19, ZAK14, dSBHM19, dlHMRB17, vNBA17]. BBPSO [Wan14]. beat [BBK17]. bedload [CR17]. Behavior [AFS14, BLBK16, KX17, MP17a]. Behrens [NL17, SOY17]. belief [AKJ17]. BEM [VKB10]. Benchmark [CL15]. Benefits [RBX^{+}15]. Bennett [ZSÁ17]. Bera [HESK18]. Berkson [WF17]. Bernoulli [DDD16, GAK16, MHAD19]. Bernstein [DZ17]. Berry [XYL19]. Bertholon [CSA17]. Best [AMM17, AMM18, CM16, GP17, Ng17]. Beta [Bid12, CZ19, OCH11, PCN14, SPCNL16, ZW15a, AK17, AP17, BRFD18, BCN17, KSS17, KH17, KL17a, LBP17, MNBB17, MMB19, PB18b, Per19]. beta-binomial [KSS17, KL17a]. Beta-Generated [CZ19]. Beta-Weibull [OCH11]. Better [HHH16]. Between [Góm13, KM13, LTPT11, McI14, OG14, RAN15, TA13, Yos15, ARK17, AdBPdBP19, BS16b, BLC17, CHB17, HM17, HBC17, JZ18, SBH18, SOBH10, Sin14, SR12, Wes17, XHD18, YWY17, dSBHM19]. between-within [BLC17]. Bhattacharyya [NRB13]. BHEP [Ten17]. Bi [RG18]. Bi-directional [RG18]. Bias [GHRM11, Hay14, LÁJGAF11, MMD19, Ouy13, WW17, ZL17, BH17, CNK17, GG19, GBP16, HNW17, JZ19, LASX19]. Bias-corrected [MMD19, WW17, ZL17]. Bias-Correction [Hay14]. Biased [Özk14, TP17, GGK17, PB18a, QZF19, SDSM18]. Biasing [KMS15]. BIBDs [RTV16]. Bifurcating [ET16]. Big [CCC10, SSYS19]. Bilinear [TW16, BM19]. Billor [OM17]. Binary [AA15, CK14, DAD14, NHLL10, RZZ10, SG15b, SID13, lTHT13, WPLL15, XLPC13, XLC14, ZLS14, AG17, BSH17, CBB^{+}17b, CTH17, CNK17, Dem17a, Fan18, GS17, LSTM17, MAT17, WHK17, YMS17]. Bingham [ARC13]. BINMA [SKJ17]. BinNor [DAD14]. Binomial [ABV16, Bay11, FL13, HAZSK16, Li10a, ON13, SSK16, ZAK14, AP17, EL18, EFA17, HA18, HW18, KD18, KSS17, KL17a, KÖ19, PGH^{+}18, PSHP17, TÖ18]. Biochemical [CKPR13]. Biomarker [VVHP^{+}16, HBC17]. biomarkers [YH19]. Biometric [BSFG10]. biometry [BKSM18]. Biparametric [LC12]. Birds [AC15a]. Birnbaum [BZ15, BA19, SCN17, SS16, WGX18]. biserial [DH16]. bit [Els19]. Bivariate [ALH14, AC15b, BB14, CMB11, GBPDL16, KLLM12a, KPIY16, Kim14, KNJ16, ON13, PBS11, PK13a, SA16, She12, SKBA18, Xia15, ZNA16, AR17b, BSLF19, BYAJ17, BPJ17, DWW18, Ery19, FL19, HPG17, HS17, JZ17, Kim17, MRF17, MP19b, MMB17, MAMM17, QLAJB17, RRR19, SKJ17, dODD19]. black [HWD18]. black-box [HWD18]. Blackwellized [DH12]. Blaker [LP16]. Blanket [LY16]. Blinded [PK13b]. Block [AA15, BRT^{+}16, CU19, FT16, GCR12, HB15, Man15b, YAA15, BFB17, DGH19, Lui17b, Lui19, NK17, XSA^{+}17, ZSY17, ZF17]. Blockage [JH14]. Blocked [ASSH10]. Blocks [ASA13, SZSA17]. BLUP [YZCW15]. BMAP [Ban17, KD12]. BMAP/PH/N [KD12]. BMAP/R/1} [Ban17]. Board [Ano11b, Ano12, Ano17a, Ano18b]. Bond [SW11a, NMS18]. Boosting [WZG15, Zha15c]. Bootstrap [AG19, AZ10, BNKM15, Baz17, BA13, BA14, CD12, DSMJ18, DZC10, Dut14, GN15, GL16, HB15, HHKM16, HNV12, IMS18, Li12, LBP17, MMV18, MWH10, MS10, Nam17, NA12, Ouy13, PB18b, PKC^{+}11, RAN15, Sma17, SGB16, WMK14, WMK16, XYCY15, XY16, XQW^{+}16, Zha15a, BFB17, EGO^{+}17, JSKV18, Kao18, NÖ18, SJC18, Sta17, Usa17, WMK19, XW17, XP19, YSS17, ZLW17, ZK18]. Bootstrap-Based [PKC^{+}11, PB18b]. Bootstrapping [BR13, GH15, LKHH11]. border [GPM18]. both [XP19]. bound [CHJL17, XYL19]. Boundaries [KRAA15]. boundary [GKA19]. Bounded [GL15]. Bounding [TC14]. Bounds [NRB13, PR15b, TC14, AND^{+}17, VND19, WPTL17, YAT19]. box [HWD18, AID17, BSZ16, DAI14, DI17]. Boxplot [DC15]. Boxplot-Based [DC15]. Bradley [Fed12]. Bradley-Terry [Fed12]. Branching [Yar12]. Break [OK15, Nam17, Sen18]. Breaking [CNPW15, ZGR^{+}17]. Breaks [RR12, JZ18]. Breast [PR19a, PR19b, ZKL^{+}19]. Bridge [CCY17, SEG15, AA19, CHJL17]. Brief [Olk16]. Brier [VL16]. Brownian [CSL17, HWZ18, KX17, SEG15, XZZ15, XY17b, ZL10, Zha11a]. BSDEs [SJT19, WLF19]. Burr [AMS16, ABJM17, CW12a, CJTL18, HK18a, KY15, KTRA17, LTAJ14, PH17, WCC^{+}18]. bursts [MCC18]. business [Tas17].

Calculate [SF12]. Calculating [Rah17]. Calculation [Lon14, PK13b, Sin14, SC18b]. Calibration [CIB10, HB15, KK16, Lan10, PA18b, RSC16, SHM12, SEG12, ZX17, dlHMRB17]. Cancer [MSB11, PR19a, PR19b, SSK16, ZKL^{+}19]. Canonical [SZFM11, SKS16, HBC17]. can’t [SSYS19]. Capability [BDS10, CT12, DZC10, HA12, JL12, RYW16, AMJ17, Bal17, DSMJ18, JG17a, XP19]. Capacitated [LH14]. Capacitor [FSJ12]. Capacity [GH16, ZYS19]. Capillary [Har12]. Cardiovascular [SKZA16]. cards [BSH17]. care [LWT17]. Carlo [Asl14a, AMKS19, FGL10, GNV15, HESK18, HE17, HC17b, JA11, KÖ19, LL14, MKS18, MBF17, MMN12, Nou17b, Özk14, PLT18, RK17, TJ17, WMH17, WMK19, Xu10]. carry [Lui17b]. carry-over [Lui17b]. carryover [Lui17a]. cascading [AR17b]. Case [AA17, Bar11, Dei12, Kor12, PLLF16, RBX^{+}15, SOSB14a, SOSB14b, ASKT19, BBA17, Els19, PDL19, PV17, SKLK19, WLZL19]. case-cohort [PDL19]. Case-Control [PLLF16]. catch [KRK18]. Categorical [KK14b, RAN15, cIC17, KAN17, SKZD18, WSX18]. Categories [LqQlX11]. Categorization [LLS16, WC17]. categorizing [CCH18]. category [OBD16]. Cauchy [TN17]. Causal [CZL16, SWL14, YBW^{+}12, GMR^{+}18, NH18, PW17]. causality [LL19, Vin17]. Cause [AZSA11, She12, PMNF18]. Cause-Selecting [AZSA11]. Cause-Specific [She12]. causes [MKFP18]. causing [YDA19]. CCA [Pol17]. CDF [BAN16, BAND16, MM18a]. cDNA [DGH19]. cement [TB17]. Censored [ALH14, ARAM15, BM14, BCV15, BB14, BZL15, CL15, CCL10, CW12a, DTL12, DDK16, DAA13, FK15, HAT14, KNJ16, KBAD11, KRK15, LM10, Li14, LHB08, LHB10, LBW11, MN15, ONa13, Par15, SB12b, Sha16, sS12c, She13a, She13b, She14, She15a, She15b, TK15a, WCL12, Wan14, WY12, WCC16, YXC10, YXS14, ZDX14, ZTT10, AGM19, AYR17, ARY18, Ate17, AY18b, BK18, BSLF19, BS17a, CY17, CTTT16, CSA17, DZX18, EDAAAM17, ES16, FA17, FL19, GRY17, HPS17, HK17, JG17b, Jov17, KSS17, Kim17, KEB17, KC17, LJ17, Lin17, MWL17, MAMM17, Nou17c, Par17, POCdSB19, RSR17, RGP17, SP18a, She17b, She17c, SKLK19, SL17, SC18b, Tri18b, VAR19, WODY17, Wu17b, Wu18a, Wu19, YCL17]. Censoring [AHAH15, AY15, BMN12, BLN12, BS15, BSL15, BS10, BR16a, CLJ13, CCLT14, HK13, Kor12, SSK16, SS16, dFR13, dR11, AEQEH17, BW19, CSY17, CD18, DOZ17, GP17, JZ17, JJB17, KTRA17, NKKC17, PN17, SK17a, SKK17, SK17b, WSZ17, WCC^{+}18, ZS17]. Censorship [AD12, AHGJRC16]. Central [Wan10b, CLW12]. Centrality [CLW12]. Certain [CLLH16, CU19]. CEV [MBF17]. Chain [HXB13, JA11, MYS^{+}16, MMN12, Ngu17, YCSD10, ZLGS18, Bal17, HST17, JU19, PLT18, Sei19]. Chains [VG15, Mar19]. Change [Ade12, BB11, Che15, DK11, DK13, FEP16, GV10, HLZK15, HZP15, JTY10, KNA15, LDW14, LTQ15, VG10, ZQL14, ZZGL15, AA17, CH17, CN19, DPS17, EB18, GGP19, GVTGF^{+}17, JJB17, MP17b, ML17, ND17, QDT18, SZMY17, SS17b, SA17b, SGNS17, XHD18]. Change-Point [Ade12, JTY10, FEP16, ZZGL15, CN19, DPS17, GVTGF^{+}17, MP17b, ND17]. change-points [GGP19]. Changed [BJ14b, XY17b]. changepoint [SGNS17]. Changes [HHKM16, HM12b, PLZ15, QDTL16, JZZZ17, MT18, SNT17, ZSGW17, ZSZW18]. Changing [ND10, RTKB18]. channel [Vil17]. channels [ML19b]. Characteristic [DBM17, JS14, OBD16, Van15, ZLTX19, vZ18]. Characteristics [Ani16, BS16a, KFO14, MMP^{+}13, PS14, ZL10, AAMJ17, YZZ17b, Za̧d17]. Characterization [BB14, VAGE15, ZA10, CU19, KH18, PH17]. Characterizations [AN12, YAR^{+}19]. characterizing [ZZS17]. Charlier [LW17b]. Chart [AMP10, AS12b, ANA14, AZSA11, AS15, CYW13, DK11, Gad14, Lee10b, Lee12, LK14, LK15a, LK15b, LDW14, LZZW13, LZZ15, MMM13, MW10, MM10, NA12, PS10, PB16, SL14, TKCL16, TGRGCF^{+}12, Yan13, Yan16, YC10, YKLR14, YKSC16, ZZW11, ZLGZ15, AHM17, AAAJ18, AKJ17, AGKJ17, AAMJ17, AMJ17, Asl19, CHB17, CYY18, GA19, HC19, Haq19, JU19, Kao18, KMV19, KKYW17, LK17a, LK17b, LK18a, LK18b, LTZ17, ML17, PAS19, RW17, SL15, SAAJ19, SKLC18, SM17, Sei19, SSA18, SKBA18, ZHLW17, LT15, LK19]. Charts [AMW15, BGM13, CP11, Chi15, HLH14, KJS10, Lee10a, LTAJ14, MGC13, NT10, NGT14, PKC^{+}11, TKKC14, TK15b, YKY10, AARQ19, BW19, CGST17, CJB^{+}17, DKL18, GK18, GAK19, HKNC17, KAN17, Kao17, KM19b, KA19b, MKFP18, PMNF18, RRA19a, Rak17, SWGX18, Tri18a, XP19, ZSGW17, ZSZW18]. Chatterjee [OM17]. Checking [SJP12]. Chen [DKRL17]. Chi [Hwa17, ZF16, Bar17a, CSA17, MO19]. Chi-Square [Hwa17, ZF16, Bar17a, CSA17]. Children [DCM16, YCSP19]. China [BZX11, ZB16]. Chinese [XZZ15]. chirp [Maz17]. Choice [ATV^{+}14, CJV13, GE13, KHH12, MSS10b, ZCR10, AdBPdBP19, HSP^{+}17, STLH17]. Choosing [EF14, dSBHM19]. chosen [CTGR17]. Chromatography [Har12]. Churn [LqQlX11]. CIC [XYZ17]. CIR [NMS18]. circle [GY19]. Circular [AHRM12, IRHM13, KFO14, KSG18, MRKI16, RRA^{+}19b, AK19, JPV^{+}18, KAS^{+}19, RRR19, Whi18]. claim [ZZ17]. Class [AJ13, AFV16, CJV13, GH16, GK14b, HWX16, HXB13, KJS10, Li14, Lin15, MSS10a, NCF15, Pap14, Pra15, RNJ16, SLLW11, ZZZJ14, ABS^{+}18, BRFD18, BPJ17, BR17, cIC17, CY17, CBH19, KA19a, KG17, LC16, LNUT17, MJ17, SPH19, dODD19, vNBA17]. Classes [AA15, SZSA17]. Classical [AGM19, Dei12, FL13, RF10, RLF11, Wil12, Wil15, AM17b, WF17]. Classification [AN11, AMA15, AAP16, Cha12, CDGML12, GL14, Gór18, JK17b, KO16, KJS10, KOC16, MBL12, RSW12, SG15a, SLLW11, SMY12, SLB11, BVC^{+}19, cIC17, HWW18, SA17a]. Classification-Based [KJS10]. classifications [dMdCQSH17]. Classifier [NHLL10, SXYF17]. Classifiers [GK15]. clinic [ABH17]. Clinical [Ani16, PK13b, ZCR10, Zha11a, ZLS14, CBB^{+}17a, CTH17, CH17, Llo17, YLM17]. cloning [dlHMRB17]. Closed [GBB^{+}12, HNM^{+}18]. closed-form [HNM^{+}18]. Closeness [AB13a, BDKM11]. Closer [YGL12, RTKB18]. Cluster [MV10, YYY16, YLK11, AM19, AKZK17, AO17, BLC17, HNM^{+}18, ISS17, KP17, LDK17, LNUT17, XYZ17]. cluster-based [AO17]. cluster-level [BLC17]. Clustered [LLJ^{+}12, LW11b, MH17b, She13a, She15a, Spa12, YLK11, dGEB17, GB17a, MB17, SAD17]. Clustering [AY18a, AAT15, AKS13a, AK14, BKJ10, BJ14a, CCD^{+}12, FG15, GZ10, MMP16, MM16, MCC18, RH11, VL16, Whi18, BB17a, CCH18, KK17b, REM19, WY17, ZLH^{+}19]. Clustering-based [MMP16]. Clusters [LLZ16, XLC14]. co [SF19]. co-sufficient [SF19]. Cobb [KABM17]. Coefficient [AB13b, AZ10, BK11, CSS13, Che14, CWYF12, CY16, DS14, FP16, Hua14, LZ10, LM10, LTQ15, Lu14, SSmK14, Wei11, WWL15, Xia15, ZM12, ZW15b, AG19, ASS18, KL18, KKYW17, MO17, XC18, YLA17, YY17c, YY17d, YKLL17, ZLW17]. Coefficients [AFvRvR15, BBF13, TZD^{+}13, VL13, BH17, Elh17, EA17, HJ18, JK17a, SJC18, SaN17, ZSÁ17]. Cognitive [CZL16]. Cohort [BL15, PDL19]. Coiflets [PR14]. Cointegrated [GD15]. Cointegrating [Kur13, Mak11]. Cointegration [Mak11, PD16, SMS17a, SMS17b]. Cojocaru [BB17b]. cold [FTE17]. Collapsed [PM16]. Collecting [Ang13]. Collinear [FX12]. Collinearity [AHGJRC16, BHI12, MSS10a]. Collinearity-Influential [BHI12]. Color [SH12, CCH18]. colour [DGH19]. Column [LZS10, GQYY18]. Column-Based [LZS10]. column-orthogonal [GQYY18]. COM [KJ13, MP19a]. COM-Poisson [KJ13]. Combination [AH13a, GCR12, HLH14, Kri16, SS12b, WL15a, EFA17, Gór18, HA17b]. Combination-Based [GCR12]. Combinations [Ger16, LLB19]. Combinatorial [Man15a, ZSZ17]. Combinatorics [BG15]. Combined [DKV15, EF14, KIM16a, LK17a, NFM19, UH17]. Combining [CDkP14, GK15, MC19, SG12, VVHP^{+}16, WY15b, ZWZ^{+}14a, GSP^{+}17, Ten17]. Comments [BB17b, CL22]. Common [BH16, PUN16, SN13b, dWD16, GE17, JK17a, JK17b, PUN17, PUN18, Tri18b, Vil17, YW17b]. Commonly [MMP^{+}13]. community [YC16]. Comparative [Bon16, Che14, HHS16, Kim16b, KP10, MSB11, Önd16, Rod16, lTHT13, TKX13, VPF14, WWJ14, ARK17, AY18b, BA19, KH19a, KNLK19]. Comparing [AGJ15, ANB19, BRT^{+}16, CPL10, CSK16, CJV13, DC17, DMMA15, GMR^{+}18, GMG13, JS14, JMCT16, KJ13, LSY16, Lon14, MH17b, MMF14, OWY13, OWY18, PFOC17, SaJMF16, SG15a, WANB13, Yao18, NÖ18, WHK17, XL18, ZK18]. Comparison [AMW15, AZ10, AAK13, ASM15, Bak16, BK10, BSV13, BW19, BAPA15, CTTT16, CDS15, CS17b, DL17, DTZG16, FS12, FG15, GS17, GCR12, HJST15, HLZK15, HD13a, HFKM16, HD13b, HAAQ11, ISS17, JKKR16, KO16, Koz14, KWW14, LYY14, LCD18, Lue18, MCM16, Mar13a, MMP^{+}13, MWSB18, Mur15, OG14, PGHA15, PD16, PUN16, PC13, RF10, SKZD18, SOBH10, SOY17, SDD17, SR12, Sta17, TNS12, TW11, TK15b, VL13, WK17, Wes14, WL15b, Yos15, YT12, YZCW15, YR11, Zha11b, ZSY17, dWD16, vdH10, Agh18, AdBPdBP19, BKSM18, CBB^{+}17b, DARMM^{+}17, DD17, EL18, FKG19, HESK18, HWD18, JK17c, MERN18, NWR17, Nou17b, SWM17, SP17a, TA17, TK17, Usa17, Vil17, YY17e, Zen17, Zha17a, dSBHM19].

Comparisons [AAK^{+}17, GBB^{+}12, GGK17, HST17, HTN14, MGG11, MGL17, OAB17, Sin14, SI14, Wan12, XY12, ZKL^{+}19, HS17, MSK^{+}17, PS18b, Wil17, Wu17b, Wu18a, Wu18b, Wu19, ZVHC17]. Competing [HLLV14, MN17b, She12, AYR17, ARY18, BT17, HC19, HK17, WSZ17, YH19]. Competitive [SFLC14]. complementary [HA17a]. Complete [GCR12, Jov17, KC17, XSA^{+}17, ZSY17]. Complex [AL16, ARC13, FMMV16, KS16, ZZZJ14, ZGC16, Elh17, Pic19]. Complexity [MM16]. Compliance [VB16]. Component [Che14, DB16, KK10a, LM16b, PRA14, TPS15, WY15a, YHL12, dWD16, BB17e, BLB19, Cha18, FTE17, MB17, MCC18, Roo17, WF17, WA18, YSS17]. Components [ASSH10, ARKB13, BGU15, Ery10, HWX16, LW11a, TNS12, UB14, WL10, ZC11, ZX13, BD17a, CYY18, Ery17, Ery19, FTE17, MO18, Man17, PA18a, PUN17, PUN18]. Componentwise [CCC16a]. Composite [BEJC17, Fed12, KMSK15, SN13b, CX17, LY17]. compositional [WWWS19]. Compound [AZSA11, Bee17, MJ17]. Compound-Estimator [AZSA11]. Compounded [CC19a, CMN17]. compounds [PM18]. Compromise [KGA13]. Compulsory [WLWT16]. Computation [Gui18, LKHH11, PR15b, Sim11, WA14, WTLWL16, AR17a, Bee17, BD17b, LL17a, LLB19, Zuo19]. Computational [EA12, FP14, JA15, JK17a, KMSK15, Kyu16, GG17b, ROC^{+}18, Zen17]. Computations [HXB13]. Compute [Liu13]. Computer [AA12, DD16, LH14, MDX17]. Computer-Generated [AA12]. Computing [DH16, HWD18, LP16, LZ14, Llo17, Phi12, Wal14, YZZ17a, IMS18]. Concerning [BM14]. Concurrent [DAD14, Dem17a]. Condition [EY17, Far16, HB10]. Condition-based [EY17]. Conditional [BR14, BM10, CW12b, DTZG16, Hod13, MAM11, Sta12, Wan12, CS17a, SF19]. conditionally [SS17a]. Conditions [LS14, MR14, RA17, VA17]. Confidence [AK15, BK10, BK11, Bar11, BS16b, CMM12, cCyYnH13, CLJ13, CTTT16, Chv17, FP16, HNV12, JW18, JL12, KBJ14, LP16, LTPT11, Naj19, ONa13, SaN17, SFC19, SD16, SZS10, SGB16, lTHT13, WN15a, WN15b, WWJ14, WC16, XX16, XD17, YR11, ZX10, Zha11a, ZGC16, vdH10, AND^{+}17, AG19, BLB19, CLB^{+}18, CCZ19, DSMJ18, EFA17, Gui17, LLWY17, LLH17, Llo17, MX17b, PFOC17, Rah17, SJC18, SOY17, WN17, WPTL17, YW17b]. Confidence-stabilizing [FP16]. Conforming [LK15b, CA17]. confounding [BLC17]. Conjecture [RTV16]. Consecutive [EK10, Ery10, KFO14, KM13, EY17]. Consecutive- [EK10, KFO14, EY17]. Consideration [RYW16]. Considerations [LH14, ZP14]. Considering [PLZ15]. Consistency [Kor12, PS13, SW16, XY17a, XYL19]. Consistent [Asl14b, Góm13]. Consisting [Ery14, FTE17]. Consolidation [SMB12]. Constant [Ade12, CW12a, FS17, GTFX14, HRS14, PBS11, UHP11, Wal11, ZQL14, GG19, SSY18]. Constant-partially [FS17]. Constant-Stress [CW12a, GTFX14, PBS11]. Constrained [MNH10, TZD^{+}13]. constrains [DZ17]. constraint [BA19]. constraints [PDL19, WB18]. construct [KPG^{+}17]. Constructing [AM14, AFL^{+}12, CP11, Els19, SP16, VM11, EZ17]. Construction [AK15, AGJ15, AKS13b, KAS^{+}19, KMSG13, LW14, Man15b, YAA15, ZGR^{+}17, DGH19, GQYY18, MX17a, PGCT17]. Consumer [GBTV16]. Consumption [McI14, HBC17]. Contagion [SKS16]. contaminated [RRA19a, Xue17]. Contamination [AFS14, RH11]. contemporaneous [Tas17]. Contention [LGM12]. Context [Sim11]. Contingency [Lin15]. continuity [EL18]. Continuous [DH16, DZ16, HMH^{+}10, YXS14, ZWJ10, ZLS14, BM19, Dem17a, HWZ18, Kna19, Lui17b, YH19, Z.18, ZZ18, dODD19]. continuous-time [BM19]. Contrast [PM12]. Contrasts [SH12]. Control [AGJ15, AS12b, AMW15, ANA14, AZSA11, AS15, BGM13, BG16, CSK16, CP11, CJB^{+}17, Chi15, DK11, Gad14, GMG13, HC19, Kan16, KJS10, Lee10a, Lee10b, Lee12, LK14, LK15a, LT15, LDW14, LTAJ14, LZZW13, LZZ15, LGM12, MMM13, MW10, MM10, MGG11, MGC13, NGT14, NA12, PLLF16, PS10, PKC^{+}11, PB16, SH12, TGRGCF^{+}12, TKKC14, TK15b, VPF14, YKY10, Yan13, YC10, ZZGL15, ZLGZ15, AARQ19, AHM17, AAAJ18, AKJ17, AGKJ17, Asl19, BW19, CGST17, DKL18, ELS17, GK18, GA19, GAK19, Haq19, KLS17, KAN17, Kao17, Kao18, KMV19, KM19b, KKYW17, KA19b, LK18a, LK18b, LTZ17, MKFP18, ML17, PMNF18, RRA19a, RW17, SAAJ19, SKLC18, SM17, Sei19, SSA18, SXYF17, SKBA18, SWGX18, Tri18a, Wan17, Wu17b, Wu18b, Wu19, XP19, ZSGW17, ZSZW18, ZF17, ZM17]. control-limit [CGST17]. controlled [JLGD17]. controller [CTXW17]. Controlling [LJ12]. conventional [KuAH19, PN17]. Convergence [Góm13, Alk17, SY17b]. Convertible [SW11a]. Convex [Ell11, SS12b]. Convolution [GNV15, Bar17c]. Cooling [MPG16]. Coordinate [PK16]. Coordinates [KWW14, MPEGMGMR11]. Coplot [Ati16]. Copula [BB14, CP16, DPCJ16, JWW14, NQ15, SN13a, SW11a, TA13, BR18, CDE19, ELS17, KH17, KH19b, RG18, SC18b, WSZ17]. copula-based [ELS17]. copula-coverage [RG18]. Copulas [AK14, Boz13, BPJ17, KPV17, SKBA18]. Coronary [McI14]. Correct [SG15a, JZ19]. corrected [MMD19, WW17, ZL17]. Correcting [NA13, FP18]. Correction [Hay14, LÁJGAF11, LA13, XLPC13, EL18, GG19, LBP17, MV19, SS19b, SMS17a, SMS17b, WY19, ZLTX19]. Corrections [De 11, LHB10]. Correlated [AD15, ASLFLR12, Asl14b, ART^{+}12, CWSW16, FZDF14, HPG17, KIM16a, KPIY16, Lee10a, MMM16, NGT14, Özk14, WW10, XT11, XLPC13, BF17, PM17, ST17, TÖ17, Z.18, ZLC^{+}17, ZYX18]. Correlation [AFvRvR15, AAK13, BDKM11, BBF13, BH16, BZX11, CY16, DS14, DH16, DKV15, FP16, GHY14, LM16a, LWH^{+}15, Önd16, PW17, SSmK14, SG12, VAGE15, WN15a, WLWT16, Xia15, YGW16, ZB16, Zha15c, Car19, FLW19, GNG17, HBC17, ISS17, JK17a, LK18b, MNBB17, PH17, Vin17, Wes17, XYZ17]. Correlation-Based [Zha15c]. Correlation-Type [BDKM11]. correlations [UMR17]. Correspondence [CDGML12, HZNN17]. Corrigendum [Ano11a, Ano14a, Ano18a]. cosine [CBH19]. Cost [GKA12, NMS18]. Costs [ARKB13]. Count [ASM15, CK16, HHS16, LK15b, MR14, XT11, ZAK14, NFM19, PGH^{+}18, RRR19, TK17]. count-based [RRR19]. counterexample [OM17]. Counting [GLS12]. Counts [KIM16a, Wei10, AA17, LL19]. Coupon [Ang13, NMS18]. Course [CCD^{+}12]. Covariance [DK13, Ima13, KS15, PUN16, SKZA16, Wes14, Baz17, GAK19, GNG17, HNM^{+}18, MT18, Mai17, MP19b, MN17a, NK17, PUN18, SS17c]. Covariate [DAI14, HHH16, Roy12, YXS14, ZLS14, PR19a, PR19b, XHD18]. Covariates [CWYF12, Hua16, SLX14, TZ14, ZHL14, BPG17, BLC17, HD17, HZ19, Lin17, SDS17, YLM17, ZZ17]. Coverage [RM16a, vdH10, RG18]. Covering [RM16a]. COVRATIO [IRHM13]. Cox [Alk17, ATC17, AID17, AdBPdBP19, DAI14, DI17, ES17a, KPG^{+}17, LLZ18, Rut15, WODY17, ZLL17]. Cox-Stuart [Rut15]. CPI [ZLW17]. credit [Zen17]. crime [LL19]. crisis [DC17]. Criteria [AAK14, AH13b, JM13, Lin15, SOBH10, AK19, BCN17, CTGR17, DGMC18, HCW17, Mai17, XYZ17, dSBHM19]. Criterion [GL11, Góm13, KP12, MMK10, AMAA17, Del17, RL17]. Critical [CN11, EL18, RR13, Sin14]. Cronbach [SN13b]. crops [Das17]. Crosier [CYY18]. Cross [AGJ15, Dut16, HdL10, HMS16, RSW12, Ste12, HWD18, Pol17, Pol18]. cross-language [Pol17, Pol18]. Cross-Over [AGJ15]. Cross-sectional [HMS16]. Cross-Validation [HdL10, Ste12, Dut16, HWD18]. Crossing [BLN12, HC17a]. Crossover [SCpL12, Lui17a, Lui17b, Lui19]. CUB [MZ17]. CUBE [Ian16]. Cubic [LZZ16, KX17]. Cucconi [RB16]. Cum [SG14, JW12]. Cumulative [BR16a, KLLM12a, KLLM12b, Lee10a, LK15b, Liu13, RCR14, YS12, BD17b, Gui18, HKNC17, KM19a, Par17, RBC17]. Cure [CLO13, GBPDL16, Kim17, dFR13, dR11, GRM17, MP19a, PB18a, She17c]. cured [CY17]. Current [Kim14, LW11b, LLX16, BKSM18]. curvatures [HK18b]. Curve [LM16a, VL13, KKGK19, SKZD18, SS17c, WB18]. Curves [cCyYnH13, JS12, JS14, KOC16, HC17a, ZZS17]. Customer [AAT15, Bon16, LqQlX11, LLS16, GB17a]. Customers [KD12, BR18]. CUSUM [AARQ19, AK11, CYY18, HLH14, LZZ15, MGC13, NT10, PAS19, RRA19a, ZZ13]. Cut [LX13, JK17c]. cut-off [JK17c]. Cut-Points [LX13]. Cutaneous [CLO13]. cutoff [HCW17]. cutoffs [YH19]. Cycle [CCM^{+}16]. cycles [Tas17]. Cyclic [YAA15, KAS^{+}19]. cylindrical [SIMS19].

D [HB16]. D-minimax [HB16]. daily [HYZ19]. Darling [Ras17]. Data [AHRM12, AOJ12, AZCS10, AL16, ALP16, AD15, ARC13, AS12c, AKV16a, AC15a, ASM15, Aus10, ARAM15, ART^{+}12, BG12, BM14, BCV15, BSZ16, BS10, Ben14, BZW13, BR13, BZL15, BSFG10, CLO13, CS12, CSG15, CL15, CCL10, CW12a, CTTT16, CK14, CCD^{+}12, CDGML12, CJV13, CK16, DAD14, DTL12, Don10, DMMA15, DAA13, EMH15, ES16, EE10, FL13, FEP16, FX12, FZDF14, GGNB10, GHY14, GP15, HAT14, HHS16, HFKM16, HM13, HD13b, HMH^{+}10, Isl11, JS12, Jin11, JKKR16, KO16, KR12, KK14b, KSJ^{+}12, Kim14, Li14, LDW14, LQW^{+}16, LW11b, LBW11, Lin15, LLX16, LMS13, LW16, MYS^{+}16, MR14, MRKI16, MSB11, MDD13, MS14, ONS16, ONa13, PWT11, PLLF16, PKK12, Par15, PRA14, QWM^{+}16, RZZ10, RT12, SB12a, Sat12, SKZA16, SB12b, Sha16, Sha15, She12, sS12c, She13a]. Data [She13b, She14, She15a, She15b, SKM^{+}10, SSK16, SLLW11, Ste12, SC13, SI13, SF12, SID13, Sun16, TK15a, TZ14, TX15, TY16, VVHP^{+}16, WCL12, Wan14, WPLL15, WY12, WLWT16, XT11, XLC14, Xu14, XTG14, YXC10, YYY16, YC10, YCSD10, YLK11, YGL12, ZDX14, Zha15b, Zha15a, ZGC16, ZWZC12, ZTT10, vdH10, Abd18, AK17, AYR17, ARY18, ATC17, AD17, AA17, AY18b, AAK^{+}17, Bah18, BSLF19, BB17a, BC17, BS18, BB17e, BDK17, CSY17, CTXW17, CMC^{+}17, CL22, CR17, CY17, CTH17, CNK17, CSA17, CCZ19, CS17b, Dem17a, DÖ17, EDAAAM17, EGO^{+}17, Fan18, FCC18, FS17, GB17a, GS17, GGM17, GGB17, HPG17, HPS17, HT17, HK17, HW17b, Iva18, JG17b, JCL18, KSS17, KKH17, KK17b, Kim17, KL17a, KC17, LJ17, LL19, Lin17, LNUT17, LSTM17, LLZ18, Lan17, Lui17b, MB17, Man17, MR18]. data [MH17b, MMB19, NK17, NFM19, NC19, ND17, Nou17c, OAC^{+}17, Par17, PS19b, PGH^{+}18, POCdSB19, PR19a, PR19b, QG19, REM19, RGP17, SIMS19, SKZD18, SMMA19, SP18a, SS17a, SSYS19, She17b, She17c, sS19a, SA17a, SKLK19, SDS17, SAD17, SL17, SC18b, TA17, TK17, UH17, Usa17, WSX18, WDC18, WWWS19, WC17, WODY17, WMK19, XW17, XCJ^{+}18, XT19, XLX^{+}19, YSH19, YW17a, YH19, YDA19, ZSZ17, ZLL17, ZLC^{+}17, ZLZ17, ZLW17, Zha18, ZZS17, dlHMRB17]. Data-Adaptive [Ste12]. Data-Generation [Aus10]. data-parametric [HW17b]. data-segmentation [ND17]. Dataset [XZ16, AY18a, BK17, TB17]. Datasets [WMK14, WMK16, Hir17, NCF18, QZF19, ZCY^{+}19]. Day [MPG16]. Dead [Pic16]. Dead-time [Pic16]. Dealing [LO16, XZ16]. Death [MMF14]. debugging [DSD19]. Decay [YCH16]. Decide [Góm13]. Decision [GMG13, Hou16, RDV12, AAK^{+}17]. Deck [PL13]. decks [BSH17]. Decomposition [BGP14, EG16, GP16, QMLS16, SB12a, Pal17]. decremental [Ros19]. DEE [KP17]. Deferred [TC14]. definite [EGO^{+}17]. Deformation [MWK16]. Degradation [FSJ12, PB10, PBS11, PS14, RPRPMG^{+}18, WBG16a, fYyP14, AE18, BD17b, CTXW17, DWW18]. Degree [MPG16, ZF16, KCC16]. Degrees [BAPA15, CLW12, GK14a, Hod13, HWD18]. Delay [KS16, LKYH17]. delayed [BHF17]. delays [SSOTB19]. Delete [GK14a]. Delete-a-Group [GK14a]. Deletion [BHI12, WLZL19]. Deliberate [KS16]. delta [MSK^{+}17]. demand [GDSCO19]. DEMATEL [SW16]. denoising [KL17b]. Densities [NCF15]. Density [Bee17, CP16, CDL12, CDS15, Kan16, MWH10, Nic10, OY16, Pic16, RCR14, VG10, Xu14, CCH18, HNW17, MH17a, WB18]. Density-Based [VG10]. Dependence [AJB11, HMS16, TA13, WZCJ16, AR17a, AA17, BR18, BPJ17, KH17, KH19b, KPV17, ZF17]. Dependencies [AK14]. Dependency [LW16, AK19, DD17, WMK19, Zha17a]. Dependent [BS10, HM12b, JWW14, AAAJ18, Asl19, BT17, CSY17, DOZ17, Ery17, Ery19, FL19, FS17, GA19, KM19a, KAAJ19, SEG15, WSZ17, ZM17]. dependently [LLZ18, She17b, ZLL17]. Depth [KOC16, LDW14, LZ14, BG17, Liu17a, PS19b, Zuo19]. derivative [MDX17]. Derivatives [GL14]. Descent [PK16]. Descriptive [Sha15, BOOTD17, IOTBO17, OTZA17]. Design [AHAH15, AM14, Chi15, DS13, FL13, GCR12, KCC16, Lee10a, Lee12, LW14, LC16, LZ16, MP13b, MMP^{+}13, NGT14, PS14, PLW^{+}14, SOSB14b, TKCL16, Tom12, VB16, WTCZ10, YHL12, YKLR14, YPH14, ZCJ12, AKZK17, AHM17, CBB^{+}17a, CH17, Del17, KM19b, LK18b, Lui17b, Lui19, MH17b, MDX17, NP17, PDL19, PMNF18, Sei19, SC18a, SSA18, Wil17]. design-based [MH17b]. Designing [AKJ17, BU15, LP10, NS18, SL15, YLA17, fYyP14, BJU18]. Designs [ATV^{+}14, ASSH10, AGJ15, AA12, ASA13, AA15, AFL^{+}12, ASLFLR12, AAK14, AKV16a, AKS13b, Aok15, BRT^{+}16, DAM15, DCM16, EB14, FT16, GP16, GH16, GKA12, HB16, KMS11, KP12, KMSG13, LRVF10, MTCP10, MP13a, Man15a, Man15b, Mie12, NHLL10, OD16, PM12, Rot12, SP16, SH12, VM11, WL15b, XQW^{+}16, YM15, YAA15, YKSC16, AJ17, CKP17, CU19, DM19, DGH19, DKL18, DK19, Els19, EZ17, GQYY18, JPV^{+}18, KAS^{+}19, LK19, MKFP18, MX17a, PM17, PM18, PGCT17, QQYZ17, RRA^{+}19b, SZSA17, XGTM18, XSA^{+}17, YZZ17a, ZSY17]. destructive [MP19a, PB18a]. Detect [Don10, SR12]. Detectability [LA13]. detected [KLS17]. Detecting [ARC13, KCM17, LL17b, PCN14, SI14, WL10, ZZW11, Che17, SKLC18, WLZL19, ZSZW18]. Detection [Ade12, BA13, CW12b, DK13, DC15, GV10, HM12b, IRHM13, MP13b, Spa12, SP14, SGNS17, VG10, AU17, DPS17, MT18, MV19, MC19, UA17, YWZ18, ZCY^{+}19, ZGR^{+}17]. determinacy [BH17]. Determination [Bus15, CH10, LS14, MMK10, EA17, MO17]. Determine [TZD^{+}13, PGH^{+}18, RW17]. Determining [CP11, KRAA15]. Deterministic [KMSK15]. Detrended [JH14, SSS14]. Developing [BM17b]. Development [ND10, RBX^{+}15, Vin17]. Developments [ZYM13]. deviance [AAC17, SFC19]. deviances [RM17]. Deviation [DS13, Hay14, KL14, Mac11a, SI13, Van15]. Device [HAZSK16, BC17, BS17b, SU18]. Devices [Bla14, Lan10]. DF [Wu10]. DF-GLS [Wu10]. Diabetic [CCLT14]. diagnosing [KAN17]. Diagnosis [BL15, VPF14, ZWZ14b]. Diagnostic [AM13, BHI12, BA14, LX13, MV10, SJP12, SSY15, TA13, Alk17, HYS18, MP19b]. diagnostics [AO17, AAC17]. Diagonal [WL14, RL17]. dichotomized [WSCW18]. Dichotomizing [ZLS14]. Dichotomous [RT12, TA17, YW17b, YWY17, YDA19, dMdCQSH17]. Dickey [Su11a]. dif [YDA19]. Differ [OWY13]. Difference [BS16b, CLW12, cCyYnH13, Che15, EG16, HHMW11, JW12, LTPT11, SG14, SR12, SSY15, TT15, Che17, Fan18, Fen17, HM17, HA17b, JW18, Pal17]. Difference-Based [TT15, EG16, Pal17]. Difference-cum-Ratio [JW12]. Difference-Cum-Ratio-Type [SG14]. Difference-Stationarity [Che15]. Differences [Aus10]. Different [AAK13, ARW16, BAPA15, Góm13, Gui17, KSGM13, MSK10, MAR11, Önd16, VPF14, Yos15, Ami19b, BH17, BR18, BW19, DÖ17, GMR^{+}18, HYZ19, MRA18, RRA^{+}19b, SKZD18]. Differential [BL15, Fen16, LW16, PGCT17, YS12, DGMC18, HWZ18, SDS17]. Difficulty [BAPA15]. Diffusion [AS14a, KL14, EGO^{+}17, QEK17]. diffusion-tensor-imaging [EGO^{+}17]. Diffusions [Nic10, RMM15]. Dimension [HTN14, LSY17, LY16, Wan10b, SSA18, SS17c, Zha17b]. Dimensional [ART^{+}12, HS12, HM13, Su11b, TX15, WTCZ10, WTZ11, BB17e, CLLH16, CCY17, GPM18, HB16, Iva18, LZ18, LNUT17, LW16, LZ16, LL17b, Man17, Maz17, MN17a, NK17, NCF18, ONS16, PK16, SFG15, TTC19, WLZL19, WLF19, YE17, ZZ17, ZYX18]. Dimensionality [WL10, DRAM18]. dimensions [Liu17a, Zuo19]. Direct [LSLW11, YKLL17]. Direction [CSK16, CFSC17, KK17a]. Direction-mixed [CSK16]. Directional [KH17, KH19b, RG18]. Dirichlet [DSR15, HC17b, MS14, Roo17, Wan10a, ZYW17]. Discordance [AHRM12]. Discordancy [MRKI16, SIMS19]. Discovery [HAT15, Kan16, LJ12, LY16, YBW^{+}12, Zha17a]. discrepancy [MRN17]. Discrete [AKV16a, CIB10, JS19a, KM12, LP16, MYS^{+}16, Mar15, Pra15, SO10, SW13, TP17, Wan12, YCSD10, YCS12, BOOTD17, Bar17a, BF17, Lue17, ZLGS18, dODD19]. discrete-event [BOOTD17]. Discrete-Time [YCSD10, YCS12, MYS^{+}16, ZLGS18]. Discretization [DZ16, JU19]. Discriminant [AJM10, AMS16, BGP14, KWW14, LNUT17, MRM10, MPEGMGMR11, PUN17, SZFM11, SS10, ZP14, BVC^{+}19, HT17, NK17, RCM17]. Discriminating [ARK17, RAAK18]. Discrimination [AAK14, MA16, KKGK19]. Discussion [Olk16]. Disease [CCD^{+}12, HAAQ11, McI14, ZWZ14b, YCSP19]. Diseases [YBW^{+}12]. disparities [LNUT17]. Dispersion [DAM15, Gad14, Gut15, TZ14, AP17, ASS18, BCN17, DZT19, GK18, LK17a, WSX18, XLW17, XP19]. dissimilarity [Gór18]. Distance [CDGML12, KM13, LPZ13, AY18a, JK17c, MX17a, SBH18]. Distances [Mur15, MO18]. Distortion [LS10]. Distributed [HN13, KP10, YC10, AGKJ17, AAMJ17, BB17a, GGK17, KÖ19, MN18, Ndu18]. Distribution [AT16, AHAH15, AFM12, AY15, AOK13, AN12, ART^{+}12, BAN16, Bak16, BR14, BDKM11, BK10, Bay11, Bee15, Ben14, Bid12, BAN15, CH10, CLJ13, CLLH16, CM16, DPCJ16, DH16, DDK16, Dut15, FK14, FG15, GDCO15, GCD13, GTFX14, GV10, HXX15, HK13, HW14, HM12c, HAZSK16, KLLM12a, KLLM12b, KBJ14, KY15, KRAA15, KPIY16, KMF15, KNJ16, KBAD11, KRK15, Kom15, KJR13, KGXR16, LÁJGAF11, LHB08, LHB10, LBW11, Liu13, LS14, Mar12, dBMdSF16, MLC16, MMN12, MAM11, MPG16, MGC13, NB15, NRB13, OWD14, ON13, OCH11, PLZ15, PSO13, PRA14, Ras13, RR13, SKST14, Sha16, Sha15, She12, SS10, SKJH11, SW11b, SS16, TCA^{+}16, TKCL16, TTZ14, TM14, TP15, Van12, Van15, VPSV13, WL11, WCL12, Wu14, WC16, WCC16, Xia15, Yan16, fYyP14, ZXJ15, ZZGL15, Abd18, ABFAM17]. distribution [AKZK17, ARM^{+}19, AR17b, AKJ17, Asl19, Ate17, BK18, BSLF19, BYAJ17, BS17a, BDL17, BA19, BJU18, BMJ17, Ben17, CCRX18, CL22, CS19, CCZZ17, CC19b, CN19, CJB^{+}17, Dem18, DKRL17, DD17, DPS17, EDAAAM17, EL18, FKG19, FNA17, FA17, Gen17, GDSCO19, GLGG17, GA17, Gui17, Gui18, HA18, Han19, HA17a, Hly17, HC17b, HSP^{+}17, JS19a, JG17b, JZ19, KTA17, KSS17, KTRA17, KKH17, KEB17, Kor17, KG17, KR18, KR17, LNT17, LCN19, LL17a, MM18a, MGL17, MMD19, MAMM17, NL17, NAJ18, Nou17a, Nou17d, PSHP17, QLAJB17, RSR17, RA17, RRR19, RCM17, RBC17, RSBZ18, RW17, RTJ17, RGP17, SBNG18, SY17a, SBZB19, SK17a, SKK17, SK17b, SBS17, sS19a, SP17a, TN19, VRN18, VA17, WK17, WK18, WW17, WGX18, WLX18, WDC18, WN17, WCC^{+}18, YMS17, YSY19, YT17]. distribution [ZW17, TP17]. Distribution-Free [ART^{+}12, MGC13, Yan16, ZZGL15]. Distributional [MSK10, MN15, OAB17, PB17b]. Distributions [AK15, Afi14, AJM10, AFM12, AMS16, AL16, AJB11, ALH14, ABV16, AFV16, BLBK16, BK16, BKB^{+}15, CZ19, CLW12, EK16, EE10, GAMA15, GL15, GEVA10, HSR10, JA15, KK14a, LC12, LDS10, LP16, MWK16, Olk16, OWY13, RNJ16, RV16, SL15, SO10, STHS16, SW13, TN16, Wil12, YCH16, ZNA16, ZWJ10, vZS12, AGM19, ARK17, ARY18, ABS^{+}18, AMM18, ABJM17, BK17, Bar17c, Bee17, CC19a, CK19, CL18, CBH19, CJTL18, CMN17, CFSC17, DSMJ18, GGZG17, GRSN19, HK18a, HE17, HK18b, HC17b, JA17, Kao17, KE17, KG17, LLB19, MJ17, MMB19, PN17, PH17, PA18b, RAAK18, SBH18, SaN17, SDSM18, SDD17, Sko19, Wu18b, YCL17, ZN18, vNBA17]. Disturbances [Asi14]. Divergence [BMPZ13, MMK10, SNO16, SO10, Zha11b, AHB19, Ami19b, TNO19]. Diverging [WSW16]. Divisible [KSS13]. dMEWMA [AS12b]. Do [Fen17, LML11, SSY18]. Does [BBK17, Sen18]. Domain [HN15, ZW15b]. Domains [TRG13]. Doray [BB17b]. Dose [CL15, NYD16, Rot12, PM18]. dose-response [PM18]. Double [AS12b, AGKJ17, CCUR12, JW12, LK17b, LK18a, LC16, MW10, RV15, SMWR14, VB16, YA19, ZC11, Ami19a, LK19, Ros19, SKLC18]. Double-Time-Scale [SMWR14]. Doubly [Pic14, sS12c, She13a, She13b, She15a, Zha15b, BPG17, Wu17b, Wu18a, Wu19]. Douglas [KABM17]. Down [MP19b, TNS12, BSV13]. Downside [BZW13]. Draftnikology [CMM14]. Drift [KNA15, XZZ15, GKA19]. Driven [Mac11a, HWZ18, HHS16, JZZZ17]. Dropout [BG12, GGNB10, XLX^{+}19]. Drug [RBX^{+}15, SKHL15]. DRUR [SSS14]. DTW [GL14, Gór18]. Dual [LZZ15, Lu14, EDE18, RRA19a]. dual-CUSUM [RRA19a]. due [BLC17, JZ18]. Duration [CW12b, HW17b, She17a]. Durations [BR14]. during [DC17]. Dynamic [DSR15, EK10, LW16, WH17, BD17b, STLH17]. dynamical [ZCY^{+}19]. Dynamics [BZX11, YCH16, Elh17, YCSP19].

E-Bayesian [GLGG17, Han19, Kiz19]. E-MSE [Han19]. Early [CZL16]. ECM [BG12]. Ecological [Yam16]. Econometric [CCUR12].

Econometrics [Cal19]. Economic [Chi15, KM19b, Lee10a, LK18b, NGT14, YC10, YKSC16, LK19, MKFP18, PMNF18, Sei19]. Economic-Statistical [Chi15, NGT14, LK18b]. Economically [YKLR14, NS18]. economics [Vin17]. economy [MAD17]. ECR [Pap14]. Edgeworth [ZJNG13]. editor [BB17b]. Editorial [Ano11b, Ano12, Ano17a, Ano18b]. Education [WLWT16]. Educational [FMMV16]. Effect [Aok15, CC10, CB17, De 11, GE13, Hod13, JY10, MMM13, NYD16, PK13b, SD16, She13a, SLX14, VL16, WZG15, WWJ14, YC11, DÖ17, Fen17, GMR^{+}18, HZNN17, Llo17, MCAK19, MNBB17, NH18, TS17, TS18, YW17b]. effecting [KKGK19]. Effective [DZ16, GK14a]. effectiveness [SDD17]. Effects [AKZK17, AKS13a, DS14, DCM16, GBPDL16, GTPS14, Gut15, HD13b, Jin11, MV10, MP13b, Mie12, MKFP18, Phi12, Roy12, Sin14, SOSB14a, SOSB14b, ST13, VL13, Vol14, Wan10a, WBG16b, YT12, YZCW15, ABH17, AD19, CA17, DGH19, EH18, JLGD17, LASX19, LLH17, Lui17a, Lui17b, PM16, PW17, QG19, RPRPMG^{+}18, Sun17, XGTM18, YCH16]. efficacy [ZKL^{+}19]. Efficiency [KEB17, MSM12, CU19, EA17, LAG17, PW17, YA17]. Efficient [AH13b, ARA13, BAN16, BAND16, BDS10, HZP15, HAT15, KCC16, LWT17, MBF17, NB15, SP16, SV13, TS11, WWL15, Xu10, HZHK17, HLWH17, HNM^{+}18, ML19b, Ott17, WX19]. EGARCH [CC10, TT14]. eigenvalues [UMN17]. Eigenvector [PUN16]. Elastic [Alh16, FX12, FZDF14, AM17a, AA18a]. Electric [ES16]. Electrodynamics [MSS16]. elements [BB17a]. elicitation [CK18]. Eligible [Pit14]. elimination [Mar19]. EM-Based [HN13, MN18, Ndu18]. emanating [CU19]. emphasis [KG17]. Emphasizing [KS16]. Empirical [BCV15, Bay11, BZL15, HXX15, Hua14, Hua16, JZ17, JKKR16, LLX16, Liu17b, MDD13, QMLS16, SOBH10, SVWH10, sS12c, SSmK14, Sun16, TKX13, VG10, XS11, Yos15, Zgh10, ZWJ10, ZX13, CS17a, GP17, HKP18, MOS18, Nou17d, ÖA19, PGH^{+}18, RK17, XZF19, vZ18]. Empirical-Likelihood-Based [Hua14, Hua16]. endogenous [HZ19]. Endpoint [BL15, BB17c]. Endpoints [XLPC13, BB17c]. Energy [Che11]. Enforcing [WB18]. enrollment [SC18a]. Ensemble [ZZZ10]. Ensure [HMH^{+}10]. Entries [CDGML12]. Entropy [BR16a, CZ19, EA12, KFM16, MSS10b, MSS16, Mah15, NA13, Nou17b, SM10, Tou11, YS12, AM17c, KM19a, KEB17, NWB^{+}19, TB17, Mac17]. Entropy-Based [Mah15, Nou17b, NWB^{+}19]. Environments [Yar12]. EOV [Ano11b, Ano12, Ano17a, Ano18b]. Epidemic [FLB15, ZWZ14b, DN18, QEK17]. Epidemiological [CCC10]. Equality [BK16, BRT^{+}16, JK17a, Li14, Mar12, SKM^{+}10, ZWJ10, ZWCC12, AJ18, AP17, CPL17, EKK19, GG17b, GG17a, GNG17, JA17, Naj17, UH17]. Equation [BHF17, Fen16, GHY14, LW16, Rai16, CB17, TA17]. Equation-solving [BHF17]. Equations [Önd16, BG17, CCZZ17, FLW19, HWZ18, Wal17, XZF19]. Equilibrium [Lon14, SBH18]. Equivalence [CH11, CWSW16, Man15a, MC12, PZ18].

Ergodic [RMM15]. Erratum [Ano14b, Ano15a, Ano15b, Ano17b]. Error [AP18, GL11, HdL10, Hua16, HM12c, IMS14, LK14, LY13, LH14, LRVF10, Lu13, PLZ15, RMS10, Roy12, RH11, WLL10, WL15b, WL15a, ZHL14, CCRX18, CMC^{+}17, CS19, DD17, EH18, FP18, GRB17, Han19, HD17, ISS17, LASX19, LZ18, Lin17, LCD18, MWL17, MV19, RL17, SS19b, SDD17, SMS17a, SMS17b, TGSA18, Wan17, WF17, WX19, YSS17]. error-correction [SMS17b]. error-in-variable [HD17]. Error-prone [Hua16, HD17]. Errors [ARA13, Asl14b, HHH16, LML11, OR11, Özk14, RYW16, RM16b, Su11a, Sun16, Wei11, hWjJsH13, YGW16, YGL12, BRFD18, CKKP17, JG17a, KuAH19, NBOM17, RN17, ST17, TÖ17, TSL^{+}19, XC18, XZ17, YLL17]. Errors-in-Variables [Wei11, hWjJsH13, Sun16, YGW16, XC18]. Escalation [Rot12]. Esseen [XYL19]. establishment [BK17]. Estimate [SSmK14, Van12, Wan14, vZ11, HYS18, MOS18, WB18]. Estimated [AMW15, CSG15, MM10, Mah12, TKCL16, Rak17, SM17, ZSZW18]. Estimates [AS12c, JKKR16, Phi12, WN15b, AM17b, ET16, FK17, Pal17, SP17a, Wes17, ZLW17, ZYW17]. Estimating [AID17, AA17, Asi14, BM17a, BK10, BK11, BSV13, CW12a, CM16, DS14, DZC10, GNV15, GHY14, HdL10, HB16, MYS^{+}16, Mak11, Mar13b, NQ15, NZL13, Önd16, OBD16, Sha15, SKJH11, SKHL15, SKJ17, Tan16, YCSD10, YW17b, ZWZC12, Zha17a, vZ11, BG17, DI17, FLW19, GMR^{+}18, LCD18, TK17, XZF19, Yao18, ZLZ17]. Estimation [AD12, AMA15, AJM10, AY15, AFvRvR15, AMS16, AB13b, ASS18, AKM11, ALH14, AHGJRC16, AVR17, ARA13, Ate17, BAN16, BAND16, BG12, BBF13, BB11, Bay11, BB14, Bee15, BD14a, BVP16, BAPA15, Cah13, CS12, CSG15, cCyYnH13, CP16, CL15, CCL10, CW12b, CB14, Dar12, De 14, De 11, DZ17, DK11, DAA13, DTZG16, DZX18, Dut15, EZ15, EMH15, ES16, EK16, EO16, FFNM17, FEP16, GAMA15, GH16, GHRM11, GDCO15, GE13, HZP15, HD13a, HRS14, HHMW11, Hod13, HKC10, HSR10, IMS14, JA11, JTY10, JJB17, Jov17, KK14a, Kan16, KY15, KR12, KK14b, KNA15, KJ13, KGA11, KMS15, Kim16b, KNJ16, KRK15, KL14, Kur16, LS16, LWZH10, LY13, LY17, LG11, Lu14, Lui16, LLZ16, MCM16, MA12, MZ17, MAC11b, MMM16, MH14, MMP^{+}13, Mer18, MWH10]. Estimation [MNH10, MAMM17, NB15, NCF15, Nic10, NH18, OG14, OD16, PS18a, PR14, PR15a, PGHA15, PKK12, PK13a, PC13, PLA16, QJ14, RMM15, Ras13, Rod16, SN13a, SKST14, SS14, SS12a, SS12b, SG17, She12, She15b, She17a, SKS16, SO10, SSA15, SSK16, SP17b, SSY15, SLX14, SID13, SS16, TN16, TVS13, TPS15, TP15, TP17, TS11, TT15, VL10, Van15, Wan10b, WW10, WTZ11, WBG16b, Wei11, hWjJsH13, WWL15, Wes14, Wil15, WY12, XC18, Xu10, Xu14, Yam16, YMS17, YXS14, YS12, YT17, YT12, YZCW15, ZZZ10, ZDX14, ZYHZ16, vZS12, vdH10, AGM19, ASKT19, AP18, AO17, AC18, AM17c, BA19, Bar17a, Bar17b, BH17, BM19, BLB19, BBK17, Cal19, CSL17, CCZZ17, CGST17, CX17, CC19c, CQH17, DN18, Del17, Dem18, DOZ17, DKRL17, Din18, DA18, DB17, DB18]. estimation [ES17a, ELS17, EB18, FLW19, FL19, GRM17, GGZG17, GMR^{+}18, GLGG17, Gui17, GB17b, HLWH17, Han19, HNW17, HS19a, HC18, HW17a, HS17, HKP18, HD17, IMS18, JT19, JG17a, JK17c, KK17a, Kar18, KP17, KNLK19, KEB17, Kiz19, LSD17, LH17, LCN19, Lui17a, MM18a, MCAK19, MGL17, MN18, MMB17, ML17, MN17a, MSEN18, NÖ18, ÖA19, PA18a, PM17, PUN18, QG19, Rah17, RSR17, Res19, SKZD18, SS17a, SKK17, SK17b, SPH19, ST17, SA17b, SK17c, Sun17, TN19, Tas17, TNO19, TS17, TS18, Tri18b, TGSA18, WK17, WK18, WW17, WZT17, Wei19, YE17, YY17d, YCS17, YWY17, YZZ17b, ZLA17, ZGE17, ZSZ17, ZYX18, ZLHW19, ZHKR17, ZZ18]. Estimations [ABV16, KBAD11, CLY19, HJ18]. Estimator [AOJ12, AZSA11, AS14b, CZ19, CDL12, CXW16, Dei12, DTL12, DX16, GL11, GA12, IE13, Iqb11, JW12, Lu13, RSC16, RMS10, RM16b, RA14, RA16, SG14, STK15, TRG13, WL14, WY15a, Wu16, Zha15b, AAW17, AG17, BHF17, BBK17, Cha18, DK17a, DKA18, EKG17, Ert18, FP18, FK18, GKS17, JK16, KuAH19, KP17, KÖ19, Mac17, Ras17, RH17, sS19a, SKLK19, TÖ17, WA17, WF17, WA18, YXY17, Yil17, Yil19]. Estimators [ATB14, AG16, Bak16, Che14, CT12, CDS15, GBP16, GK14b, GSSCR16, HWX16, KGS14, MSK10, MSS10a, MSS16, MWH10, NAK13, NA13, ÖW12, Özk14, Ron10, TK16, Vrb15, WY14, XY12, ZXJ15, AB18, BR17, CS17a, CCY17, GVTGF^{+}17, GG19, GGK17, HNM^{+}18, ISS17, KE17, LCD18, MMD19, PW17, RL17, SAB19, SBZB19, SPH19, She13b, TB17, TÖ18, Wu17a, Xue17, YA17, ZL17]. Ethernet [SMWR14]. Euclidean [CMM14]. Evaluate [TK16, Ami19b, LDK17, MR18]. Evaluating [AMAA17, BB17e, BS16b, GK14a, JLGD17, SKZA16, MRA18]. Evaluation [DK17a, EK10, GBTV16, LL14, MW10, MAR11, NT10, RM16a, SMWR14, Sta12, VA17, WZCJ16, XS11, CLT17, DGMC18, GVTGF^{+}17, JSKV18, MO17, Nou17a, WCC^{+}18, ZYS19]. Evaluations [Bon16]. Event [BS10, SCpL12, SF12, BOOTD17, BDK17, sS19a, SL17, XHD18]. Events [FEP16, MCM16, CHB17, Whi18]. Evidence [Asl14a, GP15, JZZZ17, ZB16]. Evidential [Ema13]. Evolutionary [GE13, PGCT17]. EWMA [ANA14, AGKJ17, ABJM17, BAA16, Chi15, Haq19, Lee10b, LZZW13, SAAJ19, Spa12, SP14, Yan16]. EWMA-Ordered [SP14]. Exact [BS15, BSL15, HW14, JG17b, LP16, LHB08, LHB10, LSTM17, MCM16, NL17, RLTJ16, SJC18, Sha16, SVWH10, Sim11, Sol16, SW11b, TKCL16, WGS10, BBK17, GR17, LL17a, Llo17, ZVHC17]. Example [GAK16, Li10b, Han19]. exams [Cal19]. Exceedance [MGC13]. Exceeds [CM13]. excess [XLW17]. Exchange [VM11]. Exchangeable [BGU15]. exciting [CLY19]. excluding [GPM18]. Existing [SI14]. Exists [TK16]. Exit [AS14a]. exogeneity [MRF17]. Exogenous [JY10, PB16]. Expansion [SS10, VVHP^{+}16, ZJNG13, BVC^{+}19]. Expansions [KK10a]. EXPAR [Mer18]. Expected [CMM12, CNA16, PR15b, DB18]. expectile [ZZ18]. expensive [MM17]. Experiment [Kan16, LZ16, WTCZ10, PM17, SU18]. Experimental [AM14, RSW12, SP16, Els19, NKKC17, PGCT17]. Experiments [AKS13a, CRS13, HWB13, IMS14, KP10, MP13a, SH12, SOSB14a, SOSB14b, jWzM13, DGH19, MDX17]. Explicit [MPG16, PAS19, Rif19]. Exploitation [AM14]. Exploration [AM14]. Exploration-Exploitation [AM14]. exploratory [BH17]. Exploring [Kur13, VL16]. Exponential [ALH14, AHGJRC16, Bak16, BS15, BSL15, BK16, Bid12, CLJ13, DM14, EE10, FK14, GNV15, GK14b, GTFX14, GSSCR16, KMF15, KU14, KNJ16, KRK15, KGS14, LC12, LHB08, LHB10, tLB14, MGG11, Ng13, SL15, SKST14, Sha16, SW11b, TTZ14, WA14, fYyP14, ZWZC12, dFR13, dR11, ARY18, AND^{+}17, AVR17, AKJ17, CN19, CJB^{+}17, FS17, Gen17, GGM17, GP17, GLGG17, GA17, HA18, Han19, JG17b, Jov17, KKH17, MGL17, Ng17, Ros19, SGKM17, SaN17, SP18a, SK17b, SSK16, SWGX18, Tri18b, VRN18, Wu17b, Wu18a, Wu18b, Wu19, YCL17, YT17]. Exponential-Geometric [Bid12, EE10]. Exponential-Type [KGS14]. Exponentiality [AHB19, BR16a, Mah15, SBH18, Nou17c, PZ18, SU18]. Exponentially [AS12b, Lee12, MW10, AHM17, AGKJ17, AAMJ17]. Exponentiated [ABS^{+}18, BAN16, BAN15, DKRL17, Gui17, JA11, KBAD11, ABFAM17, FNA17, SP18a]. exponentiated-exponential [SP18a]. Expression [CCD^{+}12, Isl11, ZWZC12, KK17b]. expressions [Rif19]. Extended [AK17, AFM12, EE10, KY15, TS11, Abd18, Ben17, CL22, LCN19, MGL17, SBS17]. Extension [BAND16, BAN15, YC10, ARM^{+}19, AR17b, TJ17]. Extensions [SOSB14a]. extensive [Nou17a]. extinction [CJLV18]. extra [AP17]. extra-dispersion [AP17]. extracted [PTVP19]. extremal [NKKC17, Yeh17]. Extreme [BDKM11, CH10, DB17, GHRM11, GBP16, WGS10, vZ11, HE17, Ten17, WH17]. Extremely [ML19b]. Extremes [CXW16, GN15, RBC17].

Factor [BH16, CZL16, PSS10, Sol16, SMY12, dWD16, BH17, DN16, HS19a, MSK^{+}17, NP17, PTVP19, XQW^{+}16, YXY17]. Factorial [AKS13a, Aok15, GH16, KP10, Man15a, jWzM13, YM15, Fan18, XGTM18]. Factorization [Den11, YW16, Zha11b]. Factors [BH16, BR16b, Hay14, PSS10, KKGK19, SN13b, Sol16, YDA19, AKZK17, Bah18, PTVP19, WY19, ZKL^{+}19]. fading [ML19b]. Fails [SG15a]. Failure [BZL15, Ery15, HHHZ10, JWW14, Jin11, KS16, LLZ16, SA16, SHM12, TY16, WZCJ16, YC10, YLK11, ZHL14, AK19, BDK17, CXC18, DOZ17, HA17a, Lin17, LKYH17, MJ17]. failures [AR17b, VAR19, ZSZ17]. False [BGM13, HAT15, Kan16, LJ12, Zha17a]. Families [BLBK16, CSK16, EK16, AND^{+}17, AdBPdBP19, KE17]. Family [BBF13, BMPZ13, DC15, LC12, Xia15, Zha15c, AYR17, BK17, CK19, HK18b]. familywise [Wan17]. Farlie [AJB11, Boz13]. Farlie-Gumbel-Morgenstern [AJB11, Boz13]. Fast [AMP10, HNM^{+}18, Ouy13]. Fat [KO16]. Fat-tailed [KO16]. fatigue [WGX18]. Fault [RM16a, TC14, VPF14]. Fault-Tolerant [TC14]. Faure [VC15]. Fay [Yos15]. FBST [AdBPdBP19]. FCvM [HW14]. FDR [KLS17, ZF17]. Feasibility [KU14]. Feasible [RA14, RH17, Zha18]. Feature [LLS16, WZG15, Zha15c]. fetal [BKSM18]. fetal-growth [BKSM18]. Fibonacci [Hly17]. fiducial [XL18]. FIEGARCH [PLA16].

Fields [KSS13, NCF18]. fifth [Dem17a]. fifth-order [Dem17a]. Filter [AB13b, DH12, YSWM13, ZHKR17]. Filtering [KYP11, KLS17]. filters [WH17]. Finance [BR14]. Financial [CC10, DC17, DB17]. Finite [AK11, DKV15, EO16, HT17, Hay14, HW14, Mar13b, NT10, RSC16, SSmK14, SSA15, TPS15, XC15, yYjHwK14, CS17a, GKS17, JW18, PS18a, SG17, ZLGS18]. Finite-Horizon [NT10]. Finite-Sample [Hay14, DKV15, HT17, CS17a]. finite-state [ZLGS18]. Fire [HB15]. firm [HW17b]. First [AB13b, AS14a, BD14a, ET16, GL14, HHHZ10, KCC16, LSY16, SC13, BM19, KP17, YA17]. First-degree [KCC16]. First-Order [AB13b, BD14a, ET16, BM19, YA17]. first-phase [KP17]. fish [CX17]. Fisher [Bia17b, Bia17a, GR17, NL17, SOY17, SMS17b]. Fit [AD12, BDKM11, BCGV13, BCV15, CZ19, GEVA10, HXX15, Li11, NA13, PSW16, VPSV13, Zgh10, AID17, BPG17, BC17, BS18, BK17, CBB^{+}17b, DÖ17, EAG18, ES17b, FK15, HWXH18, KGXR16, MO19, MDX17, PN17, Par17, YWZ18, vZ18]. Fitted [MLC16]. Fitting [Ben14, CC19b, KCC16, RSS17, SC13, GY19, XLX^{+}19, YXY17]. Five [KMSG13, MV10]. Five-Level [KMSG13]. Fixed [Far16, SD16, VL13, Vol14, XY16, AD19, ML19a, QG19]. flavor [HKP18]. Flexible [XD15, HSP^{+}17, TN19]. Flow [JH14, KMSK15]. Fluctuation [ARW16, JH14]. Fold [YHL12]. Foldover [EB14]. Follow [BL15]. Follow-up [BL15]. Following [JZ14]. Food [AMA15]. Foods [GVC10]. football [ZLGS18]. force [MMB19]. Forecast [CCC16b, LSLW11, SDZL15, SS19b, SDD17]. forecasters [MRA18]. Forecasting [GAPSÁ10, MNZ13, PS18b, SDZL15, DC17, GSP^{+}17, KH19a, LPC17, Maz17, SSY18]. Forecasts [CDkP14, SDD17]. forest [Wes17]. Forgetting [DH12]. Form [PFST13, She13b, HNM^{+}18]. Formation [CK16]. Formula [Bia14, GKA19, GC17]. Formulae [BGM13]. formulas [GKA19]. forward [UMR17]. Four [ASSH10, LSY16, Yeh17, ARM^{+}19, ROC^{+}18]. four-parameter [ARM^{+}19]. Fourier [Gür19, KK17b]. Fourth [AE18]. Fourth-order [AE18]. FPRs [SSY15]. Fraction [RT12, CA17]. Fractional [Aok15, FGL10, GH16, LZZ16, Man15a, MA16, PL13, Rod16, jWzM13, XZZ15, ZL10, Zha11a, CSL17, KX17, NMS18, ÖSY17]. Frailty [AKV16b, ES15, HD13a, SCpL12, VL16, ES17a, HPG17, HPS17]. Frame [Lu14]. framework [ARM^{+}19, LLZ18, SMS17b]. frameworks [HCW17]. Fréchet [KJR13]. Free [ART^{+}12, GV10, Mar12, MGC13, Pic19, Yan16, YZW19, ZZGL15]. Freedom [CLW12, GK14a, Hod13, HWD18]. Frequencies [WGS10]. Frequency [BZW13, ZW15b, STHS16, XY17a, XYL19]. Fresh [lKW13]. Freund [AR17b]. Friedman [PAM15, ST13, XSA^{+}17]. Frontier [GR10, OG14]. Full [GH16]. Fuller [Su11a]. Fuller-Type [Su11a]. Function [AJM10, AMS16, AC15b, CDL12, CLLH16, Chi15, DAA13, Dut15, GAPSÁ10, GL16, HXX15, Iva18, KLLM12a, KLLM12b, Kpa15, Liu13, PR14, PR15b, RCR14, SNM16, SS12a, SNO16, SW11a, She12, SS10, SO10, TA13, TP15, UHP11, Van15, WW10, Zgh10, AP18, AJ17, BM17a, CX17, DBM17, Del17, GG19, Gür19, Han19, JZ17, KABM17, MGB17, MAMM17, Nou17d, RCM17, RK17, She17a, sS19a, TNO19, War17, WSZ17, vZ18]. function-based [DBM17, TNO19]. Function-on-function [Iva18]. Functional [AMA15, DB16, DD16, KK10a, KOC16, PFST13, Sha15, BBA17, Elh17, GKA19, Iva18, Mat19, PB17a, SDS17, ZZS17]. functional-coefficients [Elh17]. Functionals [Dar12]. Functions [ALH14, DPCJ16, Li14, MRM10, SS12b, SKJH11, SW13, Wes14, Zha11b, ZC11, ZQL14, ZTT10, CDE19, CCH18, CBH19, FK17, Gui18, Gun17, NS18, Par17, SDD17, SKLK19, VRN18, YL17, ZLZ17]. Funds [Pit14]. Further [BM14, VCDB14]. Fused [FX12, FZDF14]. Fusion [PLLF16]. Future [LOL16, VAR19, Za̧d17]. Futures [CCC16b]. Fuzzy [CY16, SZFM11, VG15, WLWT16, WTLWL16, CR17].

G [MSR12, OTZA17, ZX10]. gambler [Sab18, HS12]. game [Ros18]. game-theory [Ros18]. GAMLSS [ARM^{+}19]. Gamma [BKB^{+}15, Dei12, HD13a, MAC11b, NRB13, PB10, PS14, RR13, WL11, YC10, AHB19, AJ17, Bar17c, HPG17, Kor17, MP19a, MCC18, ML17, NC19, POCdSB19, Rif19, RPRPMG^{+}18, SBZB19, WDC18, vNBA17, AAC17, RBC17]. gamma-divergence [AHB19]. gamma-mixture [vNBA17]. gamma-normal [Kor17]. gamma-ray [MCC18]. gap [MAMM17, sS19a]. GARCH [DBM17, HC18, JM13, KCM17, KH17, Kur13, LS10, LPC17, Su11a, YLL17, ZGR^{+}17]. GARCH-Type [JM13, LPC17]. GARMA [DN16, dALA17]. Gas [Har12]. Gauge [RYW16]. Gaussian [ALP16, BR14, CW15, DWW18, EKK19, GKA19, GDGG16, HZHK17, JT19, KK14a, KTA17, KH17, LXS19, LH13, Liu17c, MP17a, MM18b, Nou17a, SEG12, SEG15, SJR19, SC13, TK15a, XZF19]. GEE [TX15, XYZ17]. GEL [LH17]. Gene [CCD^{+}12, HWB13, Isl11, LW16, YBW^{+}12, ZWZC12, CBB^{+}17a, KK17b]. gene-treatment [CBB^{+}17a]. General [HMS16, KMSG13, MSS10a, MG12, SJP12, Sim11, Zha15c, BPG17, KN17, LKYH17, NWR17, SJT19, SKLK19, Yeh17]. Generalization [ON13, TTZ14, AB18, KKH17, WDC18]. Generalized [AHAH15, AFM12, AY15, AA16, AJB11, ABV16, Ben14, CH10, CC10, CCUR12, cCyYnH13, CLJ13, CM16, DAM15, DK11, EA12, EO16, Ger16, GDCO15, GHY14, GCD13, GK14b, GTFX14, Gut15, HWZ18, Hua14, HSP^{+}17, IMS14, JW12, KuAH19, KSG15, KMF15, KYP11, KNJ16, Kyu16, LÁJGAF11, LOL16, LH13, Lu13, MAD17, NRB13, Önd16, PR14, RSS17, SP16, SG14, Sha16, SO10, SN13b, TRG13, Tor15, Van12, Vin17, VPSV13, WZ10, WBG16b, WZTX16, YM15, YZCW15, dJvBS16, ARK17, AMAA17, AVR17, BK18, BG17, BS17a, Ben17, BR17, CKP17, CK19, CCZZ17, CC19b, CJB^{+}17, CCY17, Dem18, Dem17b, DSMJ18, EDE18, FLW19, Gen17, HA17a, HE17, HWD18, HZ19, KAN17, Kar18, KA19a, KM19a, KKH17, Kor17, KG17, KR17, KÖ19, Liu17b, LLH17, LCD18, MP19a, MX17b, Naj17]. generalized [Naj19, NC19, PDL19, PMNF18, POCdSB19, RSR17, SY17a, SK17b, SZMY17, SZSA17, SA17b, TB17, TTC19, WK17, Wan17, WH17, XLW17, XZF19, YE17, Mac17]. generally [AAMJ17, CHB17]. Generated [AA12, CZ19, VCDB14]. Generating [AD15, ART^{+}12, MAM11, SNO16, SO10, Zgh10, RSS17, TNO19]. Generation [Aus10, DAD14, KLLM12a, KLLM12b, AD17, Dem17b, Dem17a]. generator [WLF19]. Generators [KM13, SJT19]. generic [BS17b]. Genetic [HN15, Koz14, NGT14, Sat12, BVC^{+}19, CDE19, DL17, KA19a]. Genome [PLLF16]. Genome-wide [PLLF16]. genomic [YLM17]. genotype [DL17]. Geometric [Bid12, EE10, GDCO15, GDSCO19, NN15, OR11, WSQ15, CMN17, FNA17, HA17a, HK18b, HSP^{+}17, Jov17]. geophysics [Mor19]. GEV [SP17a]. GFI [CB17]. Gibbs [AM17a, CCC16a, CFSC17, lKW13, NCF18, PM16, YSY19]. Gini [ANB19, BAM11, KMV19, SL15, WANB13]. Gini-chart [SL15]. give [SSYS19]. Given [PR15b, Xia15]. GLD5 [RSBZ18]. GLG [KK14b]. GLM [NQ15]. GLMMs [RK17]. Global [LZ16, WTCZ10, WTZ11, WC15, Wil17, dLHT19, DC17, DKA18, WZT17]. GLS [Wu10]. GMC [LH17]. GMC/GEL [LH17]. GME [Mac17]. GMM [KJ13, YA17]. Gold [cCyYnH13, CCC16b]. Gompertz [CL22, Abd18, EDAAAM17, RTJ17, WCC16]. Good [AR12, Ng13]. Goodness [AD12, BDKM11, BCGV13, BCV15, BC17, CZ19, FK15, GEVA10, HXX15, KGXR16, NA13, PSW16, VPSV13, YWZ18, Zgh10, AID17, BPG17, BS18, CBB^{+}17b, EAG18, ES17b, HWXH18, MO19, PN17, Par17, vZ18]. Goodness-of-Fit [AD12, BCV15, CZ19, GEVA10, HXX15, NA13, VPSV13, BC17, FK15, KGXR16, AID17, BPG17, CBB^{+}17b, EAG18, HWXH18, PN17, Par17, vZ18]. Goodness-of-fit-based [YWZ18]. Governed [Mac11a]. Gower [AY18a]. GPD [SP17a]. GQL [KJ13]. Graded [MMM16]. Gradient [BDS10]. Gradient-Based [BDS10]. Granger [PFST13]. Graph [WA14]. Graphical [ARC13, CK18, HK10, SDS17, Liu17c, MM18b, SP18b]. graphics [IMS18]. Graphs [DAM15]. Gravitation [AS12c]. grids [JK17c]. Groparu [BB17b]. Groparu-Cojocaru [BB17b]. gross [JG17a]. Group [ABAJ14, BHI12, DX16, FZDF14, Gad14, GK14a, KO16, KSJ^{+}12, LW14, LSLW11, ND10, Par11, PR19a, PR19b, WWJ14, YHL12, ZP14, ZL10, AE17, BVC^{+}19, CHJL17, GK18, HW17a, Hou16, HJ18, HW18, LLH17, Mat19, NS18, NC19, PB17a, SKLC18, XD17, YLA17]. group-Lasso [PB17a]. group-sampling [NS18]. Grouped [LQW^{+}16, LBW11, Xu14, BDK17, QG19]. Grouping [FX12, CBB^{+}17b]. Groups [LX13, PUN16, RM16c, WC15, LSTM17, NÖ18, OWY18]. Growth [Das17, DM14, VL13, BKSM18, CJLV18, SKZD18, SJT19, SS17c]. Gumbel [AJB11, Boz13, AJM10, ARM^{+}19, BAN16, Dem18, vZS12].

Haar [PR15a]. Hadamard [FK17]. Hadi [OM17]. Haibing [Sin14]. Half [AY15, Gui17, SK17a]. half-logistic [SK17a]. Half-normal [AY15]. Halfspace [LZ14, Zuo19]. Halton [VC13, VG15]. Hamiltonian [HE17, MKS18]. Handling [MMM13, Pap14, ATC17]. Hardware [Bla14]. Hastings [BHF17, MM15]. Hausman [BR13, JY10]. Having [Far16, AK19, Ery17, WCSL17]. Hazard [BLN12, UHP11, ZQL14, HPS17, JJB17, Kiz19, Par17, TN19]. Hazards [HM12a, LLX16, SHM12, SLX14, TS11, AYR17, PDL19, PMNF18, ZS17]. health [GDSCO19, LWT17]. Heart [McI14]. Heating [MPG16]. Heavy [JTY10, MWK16, CCRX18, DBM17, JZZZ17, REM19]. Heavy-Tailed [JTY10, CCRX18, DBM17, JZZZ17, REM19]. Hedging [KPG^{+}17, WSQ15]. Hellinger [SO10]. Hellinger-Type [SO10]. hemodynamic [Del17]. Henze [HESK18]. Heritability [RT12]. Herriot [Yos15]. Heterogeneity [MH14, OWY13, WGS10, EKK19, LDK17]. Heterogeneous [CLT17, EE10, BD17a, DZT19, LWT17, SS19b, SSY18]. Heteroscedastic [ARA13, Asl14a, CA10, Özk14, PR14, PR15a, RM16b, WWL15, WC15, YHL12, CCRX18, LW17a, LL18, LLH17, NC19, POCdSB19, XL18, ZLW17]. Heteroscedasticity [Asl14b, GMG13, HFKM16, MGG11, MN15, Res19, XYCY15, XY16, XQW^{+}16, DD17, GGM17, LSY17, Nam17, Wu17b, Wu18a, Wu18b, Wu19]. Heteroscedasticity-Consistent [Asl14b]. Heteroscedasticity-robust [Res19]. Heteroskedastic [BM10, Hod13]. heteroskedasticity [PV17]. heuristic [AB19]. Heyse [CBB^{+}17b]. Hidden [KGK14, MTL10, Rod11, YCS12, KJK17]. Hierarchic [Ani16]. Hierarchical [GH16, Hou16, KGK14, Lan10, YW16, CL18, CNK17, HNM^{+}18, NCF18, SU18]. High [Asi14, ARA13, ART^{+}12, BZW13, FSJ12, HM13, LW16, LZ16, ONS16, PK16, SFG15, TX15, WTCZ10, WTZ11, ZYX18, BB17e, CCY17, DRAM18, LZ18, LNUT17, LL17b, Man17, MN17a, NK17, NCF18, SP17a, SS17c, TTC19, WLZL19, YE17, ZZ17, Zuo19]. High-correlated [ZYX18]. High-Dimensional [ART^{+}12, HM13, TX15, WTCZ10, WTZ11, LW16, LZ16, ONS16, PK16, SFG15, BB17e, CCY17, LZ18, LNUT17, LL17b, Man17, MN17a, NK17, NCF18, TTC19, WLZL19, YE17, ZZ17, ZYX18]. High-Frequency [BZW13]. High-Order [Asi14]. High-Performance [FSJ12]. Higher [AOK13, RV16, MSS17]. Higher-order [RV16]. highly [SA17a]. HITS [XS11]. Hitting [SC13, MP17a]. HIV [YCH16]. HNBUE [Pra15]. hoc [PAM15]. Hölder [HRS14]. Homogeneity [BMN12, BAM11, ES15, LH13, LL14, SW11b, ZW15a, ZJNG13, ZTT10, Baz17, JNMR17, LSTM17, XLW17]. Homogeneous [HJST15, KFM16, Yar12, CM19, LL17b]. Homoscedasticity [SJP12]. Horizon [NT10]. Horizontal [DM14, Vil17]. Hosmer [CBB^{+}17b]. Hot [PL13]. Hotelling [SKBA18]. housing [Rah17]. Hurst [CSL17]. Hybrid [HK13, KRK15, KSGM13, LHB08, LHB10, MBL12, SB12b, Sha16, SDZL15, TKKC14, Ate17, CSY17, EDAAAM17, KTRA17, KH19a, PS18b, SK17b, VAR19, WSZ17, YCL17]. Hydrological [Rut15, RB16]. hydrology [MAD17]. Hyperbolic [RSS17, QEK17].

Hypergeometric [ZXJ15]. Hyperparameter [ZZZ10]. Hyperparameters [RF10, NCF18, SFC19]. Hypersphere [FG15]. Hypotheses [CSK16, RLF11, ZWZC12]. Hypothesis [AK11, CM13, DKV15, GB13, Kur13, LP10, NYD16, PFST13, Pra15, SCN17, CLT17, DAY19, dGEB17, HM17, SMS17b, Sma17, TJ17, WCC^{+}18, ZX17].

I/II [PLW^{+}14]. Identical [CLW12, WC15, PA18a]. Identifiability [PGHA15]. Identification [BR16b, CMB11, LW16, XZZ15, Ami19a, CJLV18, LZ18]. identify [RBC17]. Identifying [ARW16, DS14, JS12, NYD16, ND17, NZL13, SWL14, STHS16]. Idiosyncratic [Sol16]. Ignorable [MYS^{+}16, YCSD10, FKG19]. II [LHB10, AGM19, ARY18, AMW15, ARAM15, BK18, BS15, BSL15, BS17a, BR16a, DDK16, DAA13, EDAAAM17, GRY17, GP17, HK13, HK17, KAN17, KNJ16, KRK15, LHB08, Mah12, MAT17, MCAK19, NKKC17, NN15, Nou17c, PLW^{+}14, Par17, Pit14, QLAJB17, RSR17, SK17a, SKK17, SZMY17, SSK16, SS16, YCL17, YT17, ZS17]. iid [BS16a, MP17a]. III [AMS16, CTH17, KY15, PH17]. Illness [MMF14]. Illness-Death [MMF14]. illustration [LWT17]. images [CCH18]. Imaging [SKZA16, EGO^{+}17]. Imbalanced [LQW^{+}16, XZ16, ZWZ^{+}14a]. Impact [BT17, Kur13, RH11]. impacts [HT17]. Impairment [CZL16]. Imperfect [LKHH11, RM16a]. Implement [DAI14]. Implementation [BGM13]. Implementing [BOOTD17]. Implication [Che11]. implications [RTKB18]. Implied [MLC16, SS19b]. Importance [Ngu17, Su11b, SCH14, GGZG17, JS19b]. important [SZSA17]. Improved [AL16, AS14a, CLW12, CYW13, CNPW15, GSSCR16, HT14, HWB13, JNMR17, LSD17, MMK10, SAB19, SG14, Tri18b, WN15b, XZ16, YSH19, Yan16, AM19, BSH17, CX17, Mac17, MKS18, ZYX18]. Improvement [AZ10, LS10, OM17, KKGK19]. improvements [PB18b]. Improving [AMA15, BJ14b, NCF18, SB17, Zha15c, HYS18, SPH19, YA17]. Imputation [BS17c, DMMA15, HMH^{+}10, MSS17, PL13, dJvBS16, CS17b, PR19b, ZLC^{+}17, ZKL^{+}19]. Imputing [PR19a]. IMSE [GP16]. IMSE-optimal [GP16]. in-control [MKFP18]. INAR [AA17, PK13a, ZW15b]. INARCH [Wei10]. incapability [LLWY17]. income [Gun17]. Incomplete [DMMA15, FT16, GHY14, Man15b, YAA15, CU19, JCL18, Lui17b, Lui19, SS17a, ZLC^{+}17]. incomplete-Block [CU19]. incorporating [CNK17, DL17]. Incorrect [TVS13, YKY10]. increased [DB18]. increasing [HWZ18]. Incremental [Ros19, WWWS19]. Incremental-decremental [Ros19]. Independence [RAN15, TNS12, ZNA16, DK19, NWB^{+}19, Orz17]. Independent [CLW12, GA19, GAK16, HHMW11, Kri16, LLJ^{+}12, NHLL10, WC15, BD17a, CL18, EDE18, HM17, HA17b, MMB19, NÖ18, OWY18, PA18a]. Index [Bia14, CT12, DZC10, GHRM11, GBP16, GP15, HB15, LZ10, LWH^{+}15, LX13, SN13a, SI13, TZD^{+}13, Van15, WCC16, YC11, vZ11, AGAB17, AMJ17, Bal17, Bia17b, Bia17a, CB17, DSMJ18, DRAM18, DWKL19, Hua16, KMV19, Kim16b, Kur16, LWZH10, LY17, WCC^{+}18, XP19, YLG16, YY17c, YY17d, YZW19, ZYHZ16]. Indices [BAM11, JL12, WTZ11, AKZK17, Bia17a, CB17, WZT17]. indirect [Cal19]. Individual [WLL10, WL15a, Hwa17]. inducted [MNBB17]. Industry [AMA15]. Inequality [GP15, Gun17]. Inference [AHAH15, AZCS10, AA16, AFV16, ARA13, ARAM15, BK18, BM14, BS15, BSL15, BZ15, BJ14b, CLW12, CJTL18, CKPR13, DDD16, DW16, FK14, FLB15, FP14, Ger16, GP15, GCD13, HN13, HLLV14, KTRA17, LM16b, McI14, MWK16, MSR12, PBS11, PRA14, QLAJB17, RZZ10, RF10, RNJ16, RGP17, SMB12, SK17a, Sha16, SNO16, SA16, SPCNL16, Sun16, Tor15, WL11, Wes14, ZM12, ZW15b, Zha15b, ZLL17, BSLF19, BYAJ17, BS17a, CC19a, DBM17, EDAAAM17, Gre17, HE17, JK17a, JG17b, JZ17, KTA17, KSS17, KR18, LXS19, LKYH17, LN17, PB18a, PDL19, Pic19, QEK17, SGKM17, SC18a, Sta17, XC18, XZ17, YL17, dlHMRB17]. Inferences [AH13a, Gun17, JA15, Li10b, RSBZ18, WCC16, YHL12, dMdCQSH17, GGM17, HA17b, XW17]. inferential [KABM17, MP19a, PB18b]. Inferiority [ZCJ12]. Infinite [Che15, CW14, DN16, JKBR16, KM12, Rod11]. Infinitely [KSS13]. Inflated [CS12, HLH14, Li11, LS14, MR14, PCN14, SPCNL16, ZWJ10, AK17, DDD16, HHS16, KR17, LBP17, RA17, WK18, Z.18]. Inflation [XT11]. Influence [AAC17, Kpa15, MPEGMGMR11, Car19, GRB17]. Influences [EA17]. Influential [ARC13, BHI12, BA13, DS14, WLZL19]. Information [AH13b, CN19, GSSCR16, JM13, KP12, KGS14, Lin15, MMK10, SS14, SSA15, AMAA17, ASS18, CDE19, DGMC18, Mah17, MDX17, Nou17c, PS18a, Pol17, Pol18, SPH19]. information-theoretic [Mah17]. Informative [CTH17, YLK11, dFR13, dR11, ATC17]. Infusion [KMSK15]. INGARCH [LL19]. Inhomogeneous [ARW16, Nic10]. INID [ZC11]. Initial [AMP10, LO16, RSS17, ZSZ17]. INLA [EZ15]. Innovations [DN16, HN13, JTY10, Kur13, DBM17, HF18, MN18, Ndu18]. innovative [UA17]. input [WLX18]. Inputs [LSY16, PB16, SB12a]. insights [RSS17]. Inspection [BAA16, yYjHwK14, GK18, NBOM17]. inspection-based [GK18]. inspired [DKL18]. instance [SXYF17]. instance-based [SXYF17]. Instrumental [HS19a, HZ19, JY10, ZYHZ16, JLGD17, SSYS19]. insurance [ZZ17]. Integer [GKA12, Man15b, BB17d]. integer-valued [BB17d]. Integrals [SV13]. Integrated [MA16, NFM19, Sol16, MV19, MN17a, KKGK19, ZM17]. Integrating [BVC^{+}19]. Integration [CCC16b, SMB12, SW11a, ÖSY17]. integrators [MKS18]. Intensity [MAC11b, AC18, CQH17]. Interaction [XYCY15, XY16, CBB^{+}17a, YWY17]. interactions [XGTM18, YLM17]. Intercept [CZL16, AMKS19]. interdependence [HW17b]. Interim [HM12a]. Intermittent [MYS^{+}16, QWM^{+}16, YCSD10, YCS12]. International [CD14, PR19a, PR19b]. interpoint [MO18]. Interpolation [LZZ16, AMM17]. Interpreting [SXYF17, CCC10]. Interval [AY15, ABV16, Bak16, CMM12, cCyYnH13, CLJ13, CCLT14, CCL10, CYW13, CY16, De 14, DTL12, GVC10, HAT14, HHMW11, Kor12, LP16, LK15a, LK15b, Lui16, NGT14, SSY15, lTHT13, WY12, YSWM13, YXS14, YT12, ZXJ15, AYR17, BSLF19, BLB19, CY17, KSS17, KKYW17, Kim17, KC17, LK18a, LLH17, RGP17, She17c, SSA18, SKLK19, Sta17, WCC^{+}18]. Interval-Censored [DTL12, HAT14, WY12, YXS14, KC17, She17c, SKLK19]. Intervals [AK15, BK10, BK11, Bar11, BS16b, CTTT16, HNV12, JL12, KBJ14, LTPT11, LOL16, MS10, ONa13, SD16, SB12b, SZS10, SGB16, WN15a, WN15b, WWJ14, XX16, Yos15, You13, YR11, ZX10, Zha11a, vdH10, AG19, CCZ19, DSMJ18, EDE18, EFA17, JW18, LLWY17, MX17b, Naj19, PFOC17, Rah17, SJC18, SaN17, SFC19, SOY17, XD17]. intervened [VA17]. intra [ISS17, XYZ17]. intra-cluster [ISS17, XYZ17]. Intraclass [AFvRvR15, BBF13, DS14, GNG17, SS17c]. Intraday [BZX11, ZB16]. Invariant [GHRM11, NB15, MO18]. inventory [Wes17]. Inverse [BR14, CBDD16, Den11, DD16, DWW18, GDGG16, KK14a, LaM14, LW17a, LH13, MAM11, PSO13, SC13, TK16, ZWZ^{+}14a, Zha15b, ZXJ15, AJ17, Ate17, BMJ17, CC19a, EKK19, Hil17, JS19b, KTA17, KPG^{+}17, LXS19, MMD19, Nou17a]. Inverse-Probability-Weighted [Zha15b]. Inversion [Pic16]. Inversive [VC13]. inverted [CK19, Gen17]. Investigate [LW16]. Investigating [HWXH18, PTVP19, WZG15]. Investigation [WP13, Del17, RTKB18, SRP^{+}17]. Involving [lTHT13]. IP [AA17]. IPW [sS19a]. Irrelevant [Cha13]. IRT [BAPA15, dSBHM19]. IRUSRT [ZWZ^{+}14a]. island [GPM18]. Isotonic [Sun16]. isotope [SJR19]. Issue [FP14]. Issues [BM14, BG15, GP15, SC13, KABM17, MP19a]. item [FN17, McI18, Yao18]. Iterated [Ouy13]. iteration [ZLTX19]. Iterative [ZLHW19, GY19, Kur16, Wal17, YY17a, YY17b, ZLH^{+}19]. Iteratively [GA12]. IV [SKS16].

Jackknife [BA13, BA14, BZL15, GK14a, JZ19]. Jackknife-After-Bootstrap [BA13, BA14]. Jackknifed [EA12, TÖ18]. Jarque [HESK18]. Jeffrey [SNO16]. Job [LMS13]. Joinpoint [HKC10]. Joint [BS15, BSL15, DAA13, KJ13, PLLF16, TY16, Wu14, WZTX16, YLK11, YCSP19, CL18, FKG19, Gui17, She17a, sS19a, WY19, XZ17, YW17a]. Jump [NZL13, ZB16, HLWH17, PLT18, YZZ17b]. Jumps [YZ11]. Jurecková [Mur15]. Jurecková-Kalina [Mur15].

Kalina [Mur15]. Kalman [AB13b, KYP11, MC19]. kappa [MGB17, War17]. kappas [MGB17, War17]. Kendall [GS11]. Kenward [Sta17]. Kernel [CDL12, CDS15, Dut15, Kan16, KWW14, LZS10, MT18, MAC11b, MWH10, SLB11, TY16, Xu14, ZAK14, AE18, BEJC17, DKA18, LXS19, MCC18, Vin17, WX19]. Kernelized [YW16]. Kernels [ZM13]. Khalaf [UMN17]. Kibria [Asl14a, EF14, UMN17]. Kind [CLLH16]. King [BDL17]. knot [YZW19]. Knowing [SMF15]. knowledge [Sen18]. Known [FP16, MA12, SMF15, TPS15, ANB19, LSD17]. Kolassa [Tan16].

Kolmogorov [DTZ10, Zen17, vZ11]. kriging [MWSB18]. Kruskal [BG15]. Kruskal-Wallis [BG15]. Kshirsagar [NRB13]. Kullback [Ami19b, Nou17c]. Kumaraswamy [ABFAM17, CMBS18, SBNG18, TN16]. Kumaraswamy- [TN16]. Kumaraswamy-transmuted [ABFAM17]. Kurtosis [AOJ12, OWY13, PB17b, ZLA17].

Label [Pap14]. Labeling [Yao13, CKKP17, HT17]. labels [cIC17]. labor [MMB19]. Lack [Afi14, Li11]. Lack-of-Fit [Li11]. LAD [YLG16]. Lag [Wu10, GGK17]. Lambda [AFM12, CM16]. Langevin [KK17a, ZGE17]. language [Pol17, Pol18]. Laplace [GCD13, KYP11, LLB19, MLC16, NFM19, SL15, XY17b, YY17a, YSY19]. Large [AR12, HN15, Mac11a, Mai17, PWT11, Sat12, SGB16, WA14, WMK14, WMK16, Hir17, MO19, NCF18, SMMA19, XCJ^{+}18, Zha17b, vZ18]. Large-Scale [PWT11, SGB16, SMMA19, XCJ^{+}18]. Larger [HTN14]. Largest [WC16]. Lasso [KSJ^{+}12, PB17a, Pol18, QMLS16, TL15, XT19, Mat19, TTC19, TW16, TSL^{+}19, YLL17, XCJ^{+}18]. LASSO-Regularized [TSL^{+}19]. Lasso-Type [TL15]. Latency [LH14]. Latent [CJV13, FMMV16, Lin15, Pap14, QWM^{+}16, RM16c, TZD^{+}13, ZZZJ14, KA19a, NCF18]. Later [CZL16]. Latin [CU19]. Law [WL11, ASKT19]. Layer [MMN16]. Layout [MMP16]. Layouts [PLZ15]. LCSS [Gór18]. Learning [CKC^{+}14, LZS10, Rod11, SFLC14, SEG12, ZWZ^{+}14a, HZHK17, KPG^{+}17, MM18b]. Least [AKM11, Bar17a, Ben14, IMS14, Van15, vZS12, CTGR17, DA18, FKA17, WY19]. Least-Squares [vZS12, Bar17a, WY19]. Leffler [Cah13]. Left [BM14, CTTT16, ES16, She12, She14, She15b, BT17, HPS17, Kao17, She17b, She17c]. left-right [Kao17]. Left-Truncated [She12, She14, She15b, CTTT16, ES16, She17b, She17c]. Legendre [DPCJ16]. Lehmann [WY12]. Lehmer [KM13]. Leibler [Ami19b, Nou17c]. Lemeshow [CBB^{+}17b]. Length [CMM12, HXB13, LKTW12, STHS16, TKCL16, WP13, Wu10, CQH17, HNW17, PB18a, SDSM18]. length-biased [PB18a]. Length-frequency [STHS16]. Lengths [GL16]. Lenth [XGTM18]. Leone [FA17]. Letter [BB17b]. Leukemia [ZWZC12]. Level [AAK13, AKS13b, CSG15, EB14, KMSG13, OK15, PC13, Sta12, jWzM13, AU17, BLC17, CKP17, EZ17, NKKC17, UA17]. Levels [NHLL10, Önd16]. Leverage [ARA13, RN17]. Lévy [VAGE15]. Library [AFL^{+}12]. Life [AHAH15, CCM^{+}16, CW12a, FSJ12, GTFX14, KSGM13, LBW11, Olk16, SS16, WCL12, WBG16b, XTG14, BC17, BJU18, CSY17, Els19, FS17, GA17, NP17, PR19a, PR19b, SKHL15, WGX18, WSZ17]. Lifetime [AGAB17, BZ15, CL15, CTTT16, ES16, GRSN19, PRA14, SC13, WCC16, fYyP14, Abd18, BD17a, CL22, CN19, OAC^{+}17, PN17, PB18a, SP18a, VRN18, WCC^{+}18, YAR^{+}19]. lifetimes [GRY17, Wu18b]. like [BGM13]. Likelihood [BM14, BS15, BSL15, BYAJ17, Bee15, BVP16, BZL15, CLW12, CH11, CXW16, CCL10, DTZG16, HN13, HJST15, HLLV14, Hua14, Ima11, KGXR16, LH13, LLX16, MCM16, Mar13b, Nic10, OG14, PB18a, Phi12, Pic19, SNT17, SVWH10, sS12c, SSmK14, Sol16, SG12, SW11b, Sun16, Tor15, VG10, ZW15a, ZWCC12, ZYM13, ZHLW17, ZX13, BA19, CS17a, CCY17, ES17a, Fan18, FL19, GG19, Gre17, HKP18, Hua16, JZ17, Kar18, KLK19, LCN19, Liu17b, MMD19, MN18, ÖA19, Ras17, RK17, SP17a, SKLK19, Usa17, WW17, WZTX16, XZF19, YXY17, YW17a, YWY17, ZVHC17]. Likelihood-Based [HN13, HJST15, HLLV14, SG12]. Likelihood-free [Pic19]. likelihood-ratio [ZVHC17]. Likelihoods [Fed12, Wal14]. Likert [WC17]. Likert-type [WC17]. Limit [ABB12, GLS12, LW14, CGST17]. Limits [CLJ13, Llo17, RW17, Zim17]. Lindley [TP17, ABS^{+}18, BMJ17, Ben17, DSMJ18, GAMA15, JJB17, MM18a, NAJ18, OAC^{+}17, WW17]. Line [Rod11, Chv17, Mor19]. line-transect [Mor19]. Linear [ASA13, AL15, ALP16, AA16, AMW15, AM13, AH13a, ATB14, ARA13, Asl14b, BA13, CW15, DAM15, DTL12, DW16, DM14, FP14, GL11, Ger16, GK14b, HWX16, HA12, Hua14, Hua16, HM12c, Iqb11, JK16, KLLM12a, KLLM12b, KNA15, KHH12, Kim16b, Kri16, LaM14, LWZH10, LM10, Li13, LY13, LM16b, Lu13, MSS10a, Mah12, Man15b, MH14, OCH11, Özk14, PM16, PC13, RSR17, Ron10, RAN10, SP16, SD16, SJP12, sS12c, SS10, Sun16, TKX13, TTZ14, TX15, Tor15, TT15, US10, VVHP^{+}16, hWjJsH13, WWL15, WL14, WY14, WY15a, Wu16, WZTX16, XY12, YXC10, YXS14, YM15, YGW16, YAT19, YT12, ZP14, ZX13, ZYHZ16, AARQ19, AE17, AMM17, AE18, AHGJRC16, AK19, BPG17, Bah18, BVC^{+}19, Cha18, CLB^{+}18, CKP17, CY17, Chv17, CMBS18, CC19c, CCY17, DGH19, DZ17, DZX18, DGMC18]. linear [EY17, Ert18, EFA17, FT17, FCC18, GRB17, GAK19, HA17a, HT17, HA17b, Hir17, HJ18, HD17, HZ19, JG17a, KAN17, KL18, KA19a, KSG18, KN17, KÖ19, LJ17, LSY17, LZ18, LXS19, LXPL17, LLB19, Liu17b, LCD18, LYG17, MJ17, NK17, NC19, ÖA19, RL17, SZMY17, SZSA17, SA17b, Sta17, TÖ17, TA17, TSL^{+}19, TTC19, Wal17, WF17, WX19, YMS17, YY17a, YY17b, Yil19, YLL17, YY17e, YZW19, ZSGW17, ZSZW18, ZX17, Zha18, ZLHW19]. Linearized [GL11, JK16]. Linex [MRN17, ALH14, DAA13]. link [AJ17]. Linnik [JS19a]. Lipschitz [WLF19]. Liquidity [TT14]. Lists [SMB12]. Literature [XS11]. Liu [AG16, DK17a, EKG17, FK18, IE13, Ron10, STK15, TÖ17, WY15a, Wu16, WA17, Wu17a, Yil17, Yil19]. Liu-Type [IE13, Ron10, AG16, EKG17, WY15a, Wu17a, Yil17, Yil19]. Lloyd [Bia17b]. LM [LO16, Wu17a]. Load [WZCJ16]. Load-sharing [WZCJ16]. Loadings [dWD16, BLB19]. Local [Dut14, Dut15, SDS17, YBW^{+}12, ZZS17, BBK17, DWKL19, dLHT19]. Location [AS12b, DC15, EK16, GHRM11, GEVA10, Mar13a, MGG11, MGC13, MAR11, NB15, ÖW12, OWY13, RB16, SKST14, Wu14, ZLZW16, CCZ19, GGM17, QZF19, SAAJ19, Tas18, Tri18b, WSX18, Wu17b, Wu18a, Wu19, XZ17]. location-biased [QZF19]. Location-Invariant [GHRM11]. Location-Scale [DC15, GEVA10, EK16, RB16]. locations [PS19b]. Log [AT16, AHGJRC16, HK13, JA15, KRAA15, Kom15, OCH11, PM12, TVS13, BRFD18, CSY17, GG17b, JA17, NC19, OAC^{+}17, PM18, POCdSB19, RAAK18, RGP17, SCN17, YMS17]. log-Birnbaum [SCN17]. Log-Linear [OCH11, AHGJRC16, YMS17]. Log-logistic [AT16, OAC^{+}17, PM18, POCdSB19, RAAK18]. Log-Normal [HK13, JA15, KRAA15, Kom15, BRFD18, CSY17, GG17b, JA17, RAAK18, RGP17]. Log-Normality [TVS13]. Logarithmic [CW12b, CN19, KR17]. Logistic [ATV^{+}14, AG16, BA14, Bus15, CC19b, DB16, HHH16, IE13, JZ14, LÁJGAF11, NHLL10, Önd16, PSW16, SL15, STK15, TG10, TM14, AT16, AAW17, Asa17, AG17, BK18, BS17a, BDL17, CJLV18, CBB^{+}17b, GE17, Gui17, HWXH18, KA19b, Mat19, Nou17d, OAC^{+}17, PM18, PLT18, POCdSB19, RSR17, RAAK18, SB17, SK17a, WA17, WA18, Xue17]. Logistic-Uniform [TM14]. Logit [Roy12]. LogitBoost [SLLW11]. Logits [OBD16]. Lognormal [AK15, BM14, Bee15, CLLH16, Bee17]. Lognormal-Pareto [Bee15]. Lognormals [BS16a]. Logspline [YXS14]. Long [AR17a, FGL10, LG11, She15b, Zha11a, ZQL14, BBK17, JZ18, PAS19]. long-memory [PAS19]. Long-range [AR17a]. Long-Term [She15b, Zha11a, ZQL14]. Longitudinal [ASM15, CK14, CCD^{+}12, DMMA15, GGNB10, GHY14, HHS16, KJ13, MR14, MSB11, QWM^{+}16, SID13, TZ14, TX15, TY16, Bah18, BS17c, Fan18, SAD17, SL17, Usa17, XLX^{+}19, YW17a]. Look [lKW13, YGL12]. Loss [ALH14, Chi15, DAA13, SS12a, SS12b, Xu14, YKY10, Yan13, Han19, NS18, SSOTB19]. Lot [ABAJ14, GL15, SKHL15, BM17b, NBOM17]. Lot-to-lot [SKHL15]. Lots [BU15, BAA16, RA17]. Low [LZS10]. Low-Rank [LZS10]. Lowe [Bia17a]. Lower [CLJ13, WPTL17, CHJL17]. LR [Wu17a]. LRC [GNG17]. LRNN [PS18b]. LS [CMB11]. LS-SVM [CMB11]. LSTAR [LS10]. LTS [DA18, RH17, Sat12].

M [CQH17, IOTBO17, MMB17, CQH17, IOTBO17, MSR12, OTZA17, ZX10]. M-periodogram [MMB17]. M/G/1 [MSR12, OTZA17, ZX10]. M/M/1 [IOTBO17]. M/M/s [CQH17]. MA [DK17b, DK17c]. MA-ARIMA [DK17b, DK17c]. Machine [BKJ10, BJ14a, SW11a, SEG12, CKKP17]. Macugen [CTH17]. Magnitudes [CCC10]. Main [Aok15, SLX14]. Maintainability [YPH14]. Maintenance [ARKB13, CD14, EY17, FN17, ZM17]. make [Fen17]. Makeham [CL22, Abd18]. Making [RDV12]. MALDI [SJR19]. MALDI-TOF [SJR19]. Mallows [LPZ13]. Manhattan [KM13]. manifold [ZGE17]. Manifolds [SG15a]. Mann [Tan16]. MANOVA [XY16]. Many [PSS10, HS19a]. Maps [FK15]. Marginal [Fed12, KBJ14, Kar18, Li14, Mar13b, RAN15, Wal14, GMR^{+}18, KPG^{+}17]. Marginals [PR15b, Xia15, AD17, KH17]. Markers [BSV13, YLM17]. Market [BZX11, SI13, ZB16, CLT17, LPC17, YZZ17b]. Marketing [LSLW11]. Markets [Che11, SDZL15, XZZ15]. Markov [ABB12, Aok15, AK19, CMC^{+}17, GA19, GGNB10, HXB13, JA11, JU19, KGK14, KJK17, LY16, MYS^{+}16, Mac11a, MTL10, MAT17, Mar19, MMN12, Ngu17, ÖSY17, PLT18, Rod11, Sei19, VG15, YCSD10, YCS12, YZ11, ZW17, ZLGS18, dLHT19]. Markov-dependent [GA19]. Markovian [AC18, LGM12, Yeh17]. Marshall [BSLF19, Ben17, FS17, KJR13, MGL17]. Masked [XTG14, CSY17, FS17]. Mass [MDD13]. massive [Zha18]. Matched [LL14, Lui16, ZCJ12, KLS17]. Matched-Pairs [ZCJ12]. matches [ZLGS18]. Matching [CZL16, KLK19]. Mathematical [HEF13]. Matrices [KS15, LZS10, PUN16, SG12, Wes14, Baz17, EGO^{+}17, FK17, GNG17, MT18, YL17, dlHMRB17]. Matrix [Den11, Ima13, YW16, Zha11b, GK18, GAK19, MP19b, MN17a, NK17, PUN18]. Max [GEVA10, LK17b]. Max-Stable [GEVA10]. Maxima [SMF15]. maximin [MX17a]. Maximum [BA19, Bee15, BVP16, CBDD16, CXW16, CCL10, DTZG16, EA12, FL19, KFM16, MSS10b, MSS16, Mac17, MN18, OG14, Phi12, SMF15, YWY17, ZYM13, GG19, Kar18, LCN19, MMD19, Ras17, SP17a, SKLK19, TB17, WW17]. Maximum-Likelihood [Phi12, YWY17]. Maxwell [SBS17, SDSM18]. may [BK17, Wes17]. McDonald [RTJ17, SN13b]. MCMC [Hly17, LL17b, Maz17, NZL13, PLA16]. MCMC-based [Maz17]. MCP [JZ14]. MCS [LPC17]. Mean [ABV16, AS14b, AS15, BGU15, BK10, CWSW16, DK13, Ery15, GL11, GRB17, GK14b, GSSCR16, Gut15, HTN14, Ima11, Ima13, JW12, JTY10, KS15, KGS14, LK14, MA12, dBMdSF16, SMF15, SKM^{+}10, SF12, TVS13, Yan13, Yan16, ZA10, ZZW11, ZZ13, AB18, AHM17, AAMJ17, Baz17, GAK19, GKS17, Haq19, JK17b, KK17a, KuAH19, LSD17, Liu17b, NAJ18, ÖSY17, PS18a, SKLC18, SG17, SPH19, ST17, SSA18, SP17b, SGNS17, Wu18b, XP19, YSH19, Zha17b, Bia17a]. Mean-shift [dBMdSF16, SGNS17]. mean-variance [NAJ18]. Means [AK15, ADLS10, BJ14a, BRT^{+}16, CPL10, CH11, DW16, HN13, JA15, KBJ14, KGA11, KP12, SS12b, VL16, WC16, dWD16, EKK19, EGO^{+}17, GG17b, JA17, KLK19, NC19, ÖEG17, OWY18, REM19, SDD17, UH17, ZLH^{+}19]. Measure [BHI12, BZW13, Hol10, MNH10, RT12, RM16c, YC16, CXC18, Mah17, REM19, WPTL17]. Measured [PK13b]. Measurement [HHH16, LY13, OR11, RYW16, Roy12, ZHL14, ZF16, BRFD18, CCRX18, FP18, GRB17, KuAH19, KAS^{+}19, LZ18, Lin17, MWL17, MV19, RN17, RRA^{+}19b, ST17, WX19]. Measures [AL16, LRVF10, MA16, OWY13, RCR14, SW16, ZX10, DÖ17, GS17, GRB17, Gór18, Gun17, KuAH19, KKGK19, YW17b]. Measuring [BZW13]. mechanism [HZNN17]. mechanisms [HT17]. media [SRP^{+}17]. Median [CTTT16, Hay14, LKTW12, TKCL16, Van12, WANB13, ANB19, LSD17, Liu17a]. Medians [BKJ10, Agh18, Che17, Wil17]. Mediated [YT12]. Melanoma [CLO13]. Memory [FGL10, LG11, BBK17, JZ18, PAS19]. Merg [MSS16]. Meta [Koc17, MCM16, ZLA17, LDK17]. meta-analyses [LDK17]. Meta-analysis [Koc17, MCM16, ZLA17]. metabolic [HCW17]. Metabolomic [VVHP^{+}16]. metamodel [MM17]. metamodel-based [MM17]. Metamodels [AM14]. Method [AL15, AH13a, AKV16a, Asl14a, BNKM15, Ben14, BA13, CD12, CM16, Don10, EF14, EK16, EA12, GZ10, GBB^{+}12, GL10, GHY14, HAT14, HWB13, HMH^{+}10, KMS11, KMSG13, Koz14, Kyu16, LP10, LW11a, LLS16, LG11, LSLW11, Lu13, MBL12, MAM11, NB15, RDV12, SDZL15, SMWR14, SW16, SF12, TW16, Tan16, TL15, XZ16, ZZZ10, AM17a, AE18, BB17a, BM17a, BM19, CC19b, DOZ17, DA18, DKL18, GE17, GSP^{+}17, HLWH17, HYZ19, JG17a, KH18, LJ17, LCD18, MWL17, ML19b, MSK^{+}17, MM17, MC19, OWY18, PH17, QQYZ17, Ras17, RK17, SMMA19, SZMY17, WZT17, XGTM18, ZLTX19, ZLC^{+}17, ZLZ17, ZZ18, ZGR^{+}17]. Methodologies [GN15, VPF14]. Methodology [AKM11, BKSM18, DAI14, DARMM^{+}17, Sen18, Zim17]. Methods [AMA15, ARC13, BSV13, BAPA15, CJV13, DMMA15, EG16, GGNB10, GAPSÁ10, GL10, GK15, HJST15, HLZK15, Hua14, HYS18, HAAQ11, JMCT16, KG14, LSY16, LZS10, LA13, LLZ16, MCM16, MMM13, Mar13a, PGHA15, PUN16, SFG15, TKX13, VCDB14, WL15b, YGL12, vdH10, Asa17, AAK^{+}17, BT17, BH17, BFB17, BB17e, Cal19, CBB^{+}17b, CS17b, EL18, FKA17, GS17, GMR^{+}18, Gui17, HST17, Kar18, LSTM17, LCD18, Lue18, MT18, MGL17, PR19b, SKZD18, ST17, Sma17, SDS17, UMN17, WK17, Whi18, XHD18, YCS17, Yao18, YW17a, Zen17, ZGE17, ZKL^{+}19, ZK18]. Metrics [CMM14]. Metropolis [BHF17, SF19]. MEWMA [LTZ17, MMM13, MM10, RW17, SM17]. Microarray [HWB13, Kan16, SH12, DGH19]. microsimulation [Rah17]. microstructure [YZZ17b]. middle [AEQEH17, SP18a]. Midzuno [SKJH11]. Migration [YC11]. Migratory [AC15a]. Miller [ZSÁ17]. Mineral [ARW16]. Minima [BS16a]. Minimal [ASA13, Aok15, Mar19, RSW12, LM17]. Minimax [Dar12, HB16]. Minimization [Tou11, FFNM17]. minimizing [TNO19]. Minimum [CH10, Car19, HS17, LSS13, NYD16, SS12b, ZP14, Bar17a, CDE19, Ras17]. Mining [CCC16b, XS11]. Mis [WBG16a, JZ18]. Mis-specification [WBG16a]. mis-specified [JZ18]. Misaligned [YGL12]. misapplication [BDK17]. Misclassification [Roy12, WLL10, WL15a, KSHD19, XD17]. misclassifications [dMdCQSH17]. Misclassified [RZZ10, TP17]. Misconceptions [ST13]. MISE [MWH10]. Missing [ALP16, EMH15, KHH12, Lin15, MYS^{+}16, MMM13, SS10, SKM^{+}10, Sun16, TZ14, Wei11, YCSD10, YCS12, ATC17, BPG17, FT17, FKG19, Liu17b, PR19a, PR19b, XT19, YSH19, ZLZ17]. Missingness [QWM^{+}16, HZNN17, XLX^{+}19]. Misspecification [PLZ15, PCN14, SPCNL16, DPS17, EA17, KPG^{+}17]. misspecifications [PB17b]. Misspecified [Tom12, Cha18]. misuse [Cal19]. Mittag [Cah13]. Mittag-Leffler [Cah13]. Mixed [Afi14, DCM16, HD13b, JH14, Jin11, Kor12, LaM14, Lu13, MTCP10, MH14, Mie12, MRM10, PM16, PC13, SD16, She13a, TZ14, Tor15, Vol14, Wan10a, XY16, YCH16, YZCW15, AY18a, AMKS19, Bah18, Bal17, CSK16, DGH19, FFNM17, GB19, GDGG16, KA19a, KC17, LW17a, Lan17, LCD18, MMB19, PLT18, SP18b, SWM17, Sta17, Sun17, TTC19, XLW17, Yao18, Yil17, Z.18, Zha18]. Mixed-Effects [HD13b, Jin11, YZCW15, PM16, YCH16]. mixing [XYL19]. Mixture [ASSH10, AJM10, AMS16, ALP16, BD14b, BVP16, CK14, Ema13, EE10, EO16, FGMW10, HSR10, KCC16, LC12, MP13b, MLC16, PM12, PGHA15, TPS15, XC15, Yao13, ZW15a, ZJNG13, ZWZC12, ZTT10, DA18, FA17, HK18a, HYZ19, JSKV18, LL18, NAJ18, PM17, PM18, SJR19, TRM18, WF17, WY17, YSY19, vNBA17]. Mixtures [AK14, BJ14b, Boz13, JCL18, MP13a, Mar13b, PGGF14, SMY12, CA17, IMS18, PA18b, YXY17]. MixTVEM [YSC19]. ML [RSR17, YZCW15]. MLE [CLLH16, JZ19, Kor12]. MLP [KH19a]. MM [MNH10]. Mobile [AAT15]. modal [LYG17, WY17, YY17d, ZH19]. Mode [LS16, QMLS16, SS14, Pal17]. Model [AD12, ATV^{+}14, AS12c, AAK14, AHGJRC16, Asl14a, Asl14b, AdBPdBP19, ARAM15, BAND16, BZ15, BB11, BCN17, BH16, BD14a, BVP16, BZL15, CLO13, CMB11, CW12b, CW15, CM15, CCD^{+}12, CKPR13, CB14, DS14, DH12, DTL12, DDD16, DTZG16, DM14, EF14, ET16, ES16, FWY14, Fed12, FSJ12, FLB15, FZDF14, FGMW10, GBPDL16, GL11, GVC10, Ger16, GR10, GGNB10, GDGG16, GB13, HZP15, Har12, HWX16, Hod13, HKC10, Hou16, Hua14, HSR10, IRHM13, JK16, JA11, Jin11, JL12, KIM16a, KK14b, KMSK15, KJ13, KFM16, KGK14, Kim14, Kur16, LZ10, LS10, LW11a, LNL11, Li13, LLJ^{+}12, Lin15, LLX16, LLZ16, LZ16, MSS10a, MJ17, MAT17, MMM16, MH14, dBMdSF16, MN15, MMF14, MM16, MPG16, OCH11, PM12, PBS11, PLZ15, PUN16, PCN14, PSW16, Phi12, QWM^{+}16, RSW12, RBX^{+}15, Rod11]. Model [RM16c, Roy12, SOBH10, SJP12, Sha16, SA16, She13a, She14, SCpL12, SG15b, SSK16, STK15, Sol16, SEG12, SPCNL16, SMY12, SCH14, TA17, Tom12, TZD^{+}13, TS11, TT14, UHP11, VL10, VL16, Van12, WL10, WTCZ10, WTZ11, Wei10, WY12, WY14, Wu16, Xu14, XYCY15, YXS14, YC10, YLK11, Yos15, YZ11, ZJNG13, ZZZJ14, ZWZC12, ZZGL15, dFR13, dR11, AM19, ASKT19, ARM^{+}19, AAC17, AJ17, AE18, AAW17, AR17b, AMKS19, Bah18, BDL17, BA19, Bar17a, BSH17, BH17, BB17d, CJLV18, CCRX18, Cha18, CN19, CX17, CNK17, CSA17, CB17, CMBS18, DARMM^{+}17, DGH19, DOZ17, DRAM18, DZ17, DÖ17, DZT19, EH18, ES17b, Ert18, EB18, dGEB17, FN17, FLW19, GRM17, GB19, GDSCO19, GB17b, GKS17, HF18, HPS17, HWW18, IMS18, ISS17, JSKV18, JJB17, Jov17, KA19a, KPG^{+}17]. model [KM19b, Kim17, KH17, KJK17, KL17a, Kiz19, LSY17, LL18, LZ18, LXPL17, Lin17, Liu17c, LLH17, LN17, LL17b, MP19a, MH17b, MBF17, MAD17, MHAD19, MMB17, MDX17, NMS18, NFM19, NC19, Nou17d, ÖA19, OAC^{+}17, OAB17, PB18a, PM18, PB18b, PDL19, PB17a, PMNF18, PAS19, PUN17, PUN18, PA18b, KKGK19, POCdSB19, QEK17, Rah17, ROC^{+}18, RBC17, SNT17, SGKM17, SP18a, SB17, SP18b, She17c, SS19b, SA17b, STLH17, SSY18, SJR19, SS17c, SKJ17, TRM18, Tew17, TGSA18, TÖ18, Usa17, WSCW18, WA17, Wu17a, WSZ17, WA18, WX19, XZ19, XLX^{+}19, YMS17, YY17a, YY17b, Yil19, YAR^{+}19, YA17, YZW19, ZLL17, ZS17, Zha18, ZYX18, ZM17, ZZS17, ZZ18, ZLGS18, dSBHM19, dALA17]. model-averaged [Liu17c]. model-averaging [XZ19]. Model-Based [CCD^{+}12, SMY12, MM16, RBX^{+}15, RM16c, MH17b, SA17b]. Model-Robust [AAK14]. Modeling [Ani16, BR14, BS10, BNKM15, BD17a, BKSM18, CWYF12, CK14, CCD^{+}12, CD14, CK16, EE10, Ery19, FMMV16, GVC10, GBTV16, Gut15, HA18, HAAQ11, KMSK15, LM17, MH17a, MMF14, MNH10, PB10, PK16, Pra15, QWM^{+}16, TY16, BR18, BPJ17, BD17b, BDK17, CL18, DN18, DA18, FKG19, FL19, GB17a, KRK18, RPRPMG^{+}18, WH17]. Modelling [BLN12, WWWS19, ZLHW19]. Models [Ade12, ALP16, AB13b, AA16, AS14a, ASLFLR12, AKV16b, Aok15, ASM15, ARA13, BCGV13, Bar11, BD14b, BA13, BR13, BG16, CC10, CCUR12, CCC16a, CWYF12, CK14, CA10, DSR15, DAM15, DCM16, EZ15, EMH15, Ema13, EO16, FWY14, FS12, FP14, GD15, GH16, GTPS14, GH15, GL16, Gut15, HN13, HD13a, HHS16, HFKM16, HHKM16, HXB13, HD13b, Hua16, HM12c, Ian16, JM13, JMCT16, JTY10, KU14, Kim16b, Kyu16, LaM14, Lak11, LSS13, LWZH10, Li10a, LM10, LNL11, Li11, LY13, LWH^{+}15, LM16b, LW11b, Lin15, Lu13, MTL10, MA16, MWK16, MRM10, NQ15, NZL13, ONa13, OG14, Özk14, PR14, PR15a, PGHA15, Pap14, PM16, PFST13, PC13, PLA16, QDTL16, RF10, RR12, Ron10, RAN10, RA14, RA16, SaJMF16, SP16, SD16, sS12c, She15a, SKS16, SHM12]. Models [Sta12, SC13, Sun16, TK15a, TW16, TZ14, TPS15, TX15, Tor15, TT15, VVHP^{+}16, VL13, Vol14, Wan10a, Wan12, WA14, WBG16a, Wei11, hWjJsH13, WWL15, Wu14, WZTX16, Xu10, XY12, XC15, XY16, Yam16, YCH16, YXC10, YM15, YLG16, YGW16, YC10, YCS12, YGL12, YT12, YZCW15, ZW15a, ZM12, ZHL14, ZX13, ZYHZ16, dJvBS16, AK17, AEQEH17, AM17b, AP18, Ami19a, Ami19b, AU17, AD19, BOOTD17, BPG17, BM17a, Bar17b, BRFD18, BKSM18, BBK17, BLC17, CK18, CMC^{+}17, CXC18, CKP17, CY17, CC19c, CCY17, DARMM^{+}17, DBM17, DM19, DK17b, DK17c, DC17, Din18, DZX18, DWW18, ES17a, Elh17, ELS17, Ery18, Ery19, FT17, FCC18, FFNM17, GGP19, GRB17, GE17, GGK17, GSP^{+}17, HLWH17, HPG17, HWD18, HS19a, HC18, HNM^{+}18, HJ18, HD17, HZ19, JSKV18, JCL18, KCM17, KL18, KPG^{+}17]. models [KSHD19, KH19a, KH19b, KÖ19, LW17a, LAG17, LH17, LJ17, LL19, LXS19, LXPL17, Liu17c, Liu17b, Lan17, LCD18, LPC17, LYG17, MZ17, MRF17, MM18b, Mat19, McI18, Mer18, MN18, MMB19, MWSB18, MERN18, Ndu18, NCF18, PLT18, PGH^{+}18, Pic19, PS18b, QDT18, QG19, RPRPMG^{+}18, RH17, Ros18, SKZD18, SFC19, SP18b, She17b, SWM17, SMS17a, Sta17, SC18b, Sun17, TÖ17, Tas17, TS17, TS18, TA17, TSL^{+}19, TTC19, WK18, WZT17, WODY17, WF17, WY17, XC18, XLW17, XY17b, XZ17, XW17, XL18, XT19, Xue17, YY17c, YY17d, YCS17, Yao18, YZZ17a, YW17a, YLL17, YAR^{+}19, YCSP19, YSS17, ZLC^{+}17, ZLW17, ZLHW19, ZGR^{+}17, dSBHM19, dLHT19, dODD19]. Moderate [KL14]. Modern [FL13]. Modes [CDS15, STHS16]. Modification [Sat12]. Modified [AS12c, AY18b, CXW16, DTZ10, Hir17, Hol10, KM19b, KBAD11, Kri16, LOL16, MS10, NK17, She13b, TGRGCF^{+}12, VC15, WLL10, Wu16, ZW15a, ABFAM17, AMAA17, CC19b, CSA17, Ert18, KLK19, LCN19, LTZ17, NL17, NWB^{+}19, Rak17, SKLC18, SMMA19, Wu19, XSA^{+}17, XYZ17, YCL17, LT15].

Modules [Ery14]. Modulus [YS12]. Moment [BM19, Lu13, SVWH10, SW13, Zgh10, ZJNG13, CCZZ17, SP17a]. Moment-Based [SW13, ZJNG13]. moment-equations [CCZZ17]. Moments [Bee15, LSY16, NA13, Vrb15, BS17a, GKA19, LCD18, MSS17, Rif19]. Monitor [YKY10, Yan13]. Monitored [ZL10]. Monitoring [AS12b, AM13, AZSA11, AAMJ17, AS15, Che15, DK13, ESAD17, HJST15, HLH14, HHKM16, KKYW17, KJS10, Lee10a, LJ12, LTQ15, LTAJ14, NA12, NN15, PB16, QDTL16, ZC11, ZZ13, ZSGW17, AARQ19, AAAJ18, GAK19, GB19, HC19, Haq19, Hwa17, KAN17, KN17, KA19b, LK17a, MAT17, MCAK19, MP19b, QDT18, SAAJ19, SZMY17, SSA18, SRP^{+}17, WSX18, YWZ18, YKLL17, ZHLW17]. Monotone [SS10, SZS10, YSH19]. Monotonic [SA17b, SGB16, AA17, WLF19]. monotonicity [KRK18, SJT19]. Monte [HE17, Asl14a, AMKS19, FGL10, GNV15, HESK18, HC17b, JA11, KÖ19, LL14, MKS18, MBF17, MMN12, Nou17b, Özk14, PLT18, RK17, TJ17, WMH17, WMK19, Xu10]. Moran [DWKL19]. morbidity [ASKT19]. Morgenstern [AJB11, Boz13]. Morphing [BSFG10]. mortality [ASKT19]. MOSUM [QDT18]. mother [YCSP19]. Motion [ZL10, Zha11a, CSL17, HWZ18, KX17]. Motions [Mac11a, XZZ15]. motivated [GRSN19]. Moulton [Bia17b]. Moving [AS12b, CXW16, Lee12, MW10, AHM17, AGKJ17, AAMJ17, CHB17, JU19, MHAD19, PB18b, Tas17]. MPL [ZC11]. MRL [ZC11]. MS [SJR19]. MSC [Grc17]. MSE [CYW13, Han19]. Muller [BSZ16]. Multi [AZCS10, BPS13, BSFG10, BL15, Cha12, FSJ12, FX12, HAAQ11, KGA13, KGK14, MMF14, OK15, OD16, RM16a, Su11b, ZZ17, CQH17, KA19a, KSHD19, KJK17, MSEN18, NP17, NKKC17, Tas18, Vil17, ZYS19]. Multi-Aspect [BSFG10]. multi-channel [Vil17]. Multi-Cohort [BL15]. Multi-Collinear [FX12]. Multi-Dimensional [Su11b]. multi-factor [NP17]. multi-level [NKKC17]. Multi-Phase [FSJ12]. Multi-ranker [OD16]. Multi-Response [KGA13, MSEN18]. multi-sample [Tas18]. multi-server [CQH17]. Multi-Sided [BPS13]. Multi-Site [KGK14, KJK17]. Multi-Stage [AZCS10, KA19a]. Multi-State [HAAQ11, MMF14, RM16a, KSHD19, ZYS19]. Multi-Step [Cha12]. Multi-type [ZZ17]. Multi-Variance [OK15]. Multiclass [SG15a, Mat19]. multicollinearity [AMAA17, Asa17, Ert18, LASX19]. Multidimensional [MMM16, Ott17, Yao18]. Multilevel [MSM12, PSW16, TG10, ZZZJ14, BS18, ES17b, EB18]. Multinomial [JKBR16, TW11, VL10, VKB10, BS17c, DM19, HWXH18, WN17]. Multiobjective [GKA12, HAB16]. MultiOrd [AD15]. Multiple [AM13, BDS10, cIC17, CW12a, CB14, DMMA15, GBB^{+}12, GK15, GMG13, GTFX14, HMH^{+}10, HTN14, KAAJ19, tLB14, MGG11, MSK^{+}17, PB10, PS14, PB16, SV13, SS12a, Spa12, SP14, TNS12, US10, WN15a, Wu17b, Wu18a, XLPC13, YC16, ZQL14, dJvBS16, Agh18, AAAJ18, Asl19, BBA17, BA19, Ban17, Chv17, DD17, GGP19, GAK19, LSTM17, Maz17, MKFP18, NH18, PV17, STLH17, WLZL19, WCSL17, Wu18b, Wu19, YLA17, ZLC^{+}17, Zha17a]. multiple-case [WLZL19]. Multiple-class [cIC17]. Multiple-community [YC16]. multiple-step [Maz17]. Multiple-Steps [PB10]. multiple-variety [STLH17]. Multiplicative [ASM15, Hou16, BRFD18]. multiplicity [PMNF18]. multiplicity-cause [PMNF18]. Multiply [WCL12, JG17b]. multipower [CLT17]. Multiresolution [CP16, PGGF14, KL17b]. Multiresponse [MP13a]. multiscale [KLS17]. Multistage [AZSA11, RM16b, ZF16, AS17, KN17, MSEN18]. Multistate [Kim14, CMC^{+}17]. Multistratum [VM11]. Multivariate [AOJ12, ALP16, AS12b, AFV16, BG12, BMPZ13, Bon16, CCM^{+}16, CP11, CCLT14, Che11, DPCJ16, DK13, Gad14, GKA12, HAB16, HHHZ10, HMH^{+}10, Ima11, Ima13, Iqb11, KBJ14, KNA15, KYP11, KGA11, KJS10, Kur13, Lee10b, Lee12, LK15a, LDW14, LW11b, LRVF10, Mur15, NAJ18, PKC^{+}11, SOSB14a, SMY12, SGB16, TKKC14, TK15b, VPF14, WC16, YKLR14, YKSC16, ZZGL15, ZLGZ15, AO17, AD17, BBA17, CYY18, CFSC17, DKA18, GAK19, HESK18, HC17b, Hwa17, IMS18, KAN17, MO18, MO19, MN18, Naj17, Naj19, PTVP19, PS19b, SSA18, Sma17, SXYF17, Ten17, YMS17, Yeh17, ZLC^{+}17]. Multiwavelets [CP16, DPCJ16].

N [KD12]. Naive [CKC^{+}14, Sta17]. Nakagami [ML19b]. Near [GLS12]. Near-Records [GLS12]. Nearest [GL10]. Nearly [ASSH10, HW18, GQYY18]. Necessary [BMPZ13, GBO14, WC17]. need [SSY18]. Needed [Aus10]. NEF [ZJNG13]. NEF-QVF [ZJNG13]. Negative [ABV16, CDGML12, HAZSK16, KD12, Li10a, ON13, KD18, MNBB17, PGH^{+}18, PSHP17, TÖ18, Zha11b]. negatively [WDC18, XY17a]. Negatives [SSY15]. negligible [XGTM18]. Neighbor [AA12, ASA13, AA15, GL10, JPV^{+}18, SZSA17]. Nested [ASLFLR12, HFKM16, JMCT16, LW11a, XQW^{+}16, YHL12, EH18, ISS17, NFM19]. nested-error [EH18]. Nesterov [TL15]. Net [Alh16, FX12, FZDF14, AM17a, AA18a, KKGK19]. Network [CKPR13, LW16, PWT11, RR13, HZHK17, MV19]. Networks [CCC16b, LH14, WA14, MERN18, PV17]. Neumann [Rai16]. Neural [AS14a, CCC16b, MERN18, PV17]. Newton [CHJL17]. Next [SH12]. NFL [CMM14]. No [DAI14]. nodes [ML19a]. Noise [KYP11, KG14, WW10, YZZ17b]. Noisy [Dar12, JS12, YYY16, JCL18]. Non [CLW12, CH11, CM19, CDGML12, Dar12, HJST15, HHH16, HA12, KBAD11, Kur16, LML11, NGT14, SL14, SLX14, VHV^{+}16, jWzM13, WL14, WY15b, Yar12, YC10, Zha11b, ZCJ12, ZM13, CA17, DKA18, FKG19, GG19, JT19, KuAH19, MKFP18, PA18a, PS18a, PTVP19, RL17, SKLK19, SKJ17, WZT17, WLF19, YY17a, YY17b]. non-conforming [CA17]. non-constant [GG19]. non-conventional [KuAH19]. Non-Diagonal-Type [WL14, RL17]. non-Gaussian [JT19]. Non-Homogeneous [HJST15, CM19]. non-identical [PA18a]. non-ignorable [FKG19]. Non-iterative [Kur16, YY17a, YY17b]. non-Lipschitz [WLF19]. Non-Normal [NGT14, jWzM13]. Non-Normality [SL14, MKFP18]. Non-Normally [YC10]. non-orthogonal [PTVP19]. Non-parametric [HHH16, VHV^{+}16, DKA18, SKLK19, WZT17]. Non-Proportional [SLX14]. Non-PSD [ZM13]. non-response [PS18a]. Non-Stationary [WY15b, SKJ17]. Nonadditivity [MMP16]. Noncausality [PFST13]. Noncentral [AOK13, KPIY16, Liu13, ON13]. Nonconforming [DS13]. Nonconformities [GL15]. nonextreme [Ten17]. Nonhomogeneous [GGNB10, HLZK15]. Nonidentical [GAK16]. Nonignorable [TZ14]. nonignorably [XT19]. Noninformative [dR11, CTH17]. Nonlinear [Ben14, CA10, DCM16, EF14, GVC10, GKA12, Gut15, HF18, LML11, Mak11, MG12, MRM10, PFST13, Rai16, RMS10, RM16b, TZ14, Vol14, Wan10a, WBG16a, ABH17, Elh17, GB19, Gür19, MZ17, SMS17a, Sun17, TA17, XC18, XZ17]. Nonnormal [BDS10, Dem17a]. nonnormality [CGST17]. Nonparametric [AN11, AM17c, AZ10, BCV15, BK10, BRT^{+}16, CY12, CWYF12, CCZ19, DB18, EGO^{+}17, dGEB17, FGMW10, GGP19, GB17a, GP15, GB13, GB17b, HM12c, JKKR16, KLLM12a, KLLM12b, KJK17, KRK18, LZZW13, LZZ15, MTCP10, MTL10, Mar13a, NA12, Orz17, Par11, Par15, PS10, PS19b, SN13a, SB12a, Ste12, SZS10, SI13, Tri18a, TP15, UB14, WMK16, XD15, YZZ17b, ZDX14, Zha15b, ZWZC12, ZX13, ZAK14, AP18, Che17, DD17, HK18a, HLWH17, Haq19, HNW17, KNLK19, Lue18, MR18, QDT18, RRR19, Zar19, ZVHC17, ZH19]. Nonproportional [HM12a]. Nonregular [EB14]. Nonrespondents [RSC16]. Nonresponse [HAB16]. Nonsensitive [GSSCR16, KGS14, SPH19]. nonstationary [BBK17]. Nonzero [DW16]. norm [BSK19, LNT17]. Normal [AL15, BG12, BGP14, BSZ16, Bar11, CCM^{+}16, CLB^{+}18, CSS13, CWSW16, DAD14, Fed12, FP16, GCD13, HYZ19, HK13, HFKM16, HA12, HD13b, Ima11, Ima13, JA15, KRAA15, Kom15, Kri16, LML11, Lon14, dBMdSF16, MLC16, NGT14, jWzM13, WN15a, Wu14, WC16, ZWCC12, AJ18, AY15, AD17, BB17a, BRFD18, CSY17, CFSC17, CA17, CMBS18, EL18, GVTGF^{+}17, GG17b, GG17a, Gui18, HS17, HK18b, IMS18, JA17, JK17b, KLK19, Kor17, KG17, LL18, LLWY17, MAD17, MP19b, MMB19, NAJ18, Naj17, Naj19, ÖEG17, PS19a, PA18b, RAAK18, RGP17, SNT17, SSA18, Sko19, UH17, WBG16b, XZ17, ZL17]. Normal-based [Kri16]. Normal-Laplace [MLC16]. normal-Pareto [LL18]. normal/independent [CL18]. Normality [AOK13, BMPZ13, DTZ10, HT14, HM13, SL14, SJP12, SVWH10, SFS12, TVS13, VAGE15, Zgh10, ZA10, HESK18, MO18, MO19, Man17, MKFP18, PFOC17, SY17b, Ten17, Ul 17, vZ18]. Normalization [McI18, SLLW11, HST17]. Normalizing [Wal11]. Normally [YC10]. Note [AR10, CPL10, Den11, GE13, HRS14, LT15, Rod16, Su11b, TL15, TP15, XT11, ZW15a, ZWCC12, AC18, JPV^{+}18, Ott17]. Notes [AM13, LL14, Lui16, Lui17a, Lui17b, Lui19, Tan16]. Novel [AM14, Asi14, BHI12, GZ10, PLLF16, SMWR14, SF12, YSWM13, ZWZ^{+}14a, HKP18, ZLTX19]. Novelty [TKKC14, TK15b]. NRL [Orz17]. Null [Che15, ZWZC12]. Number [AR12, CP11, CM13, CDS15, Ery10, Far16, GL15, LASX19, WSW16, ML19a, MMB17, ZHKR17]. Numbers [Aus10, Bla14, KM13, Cal19]. Numerical [ARC13, BLBK16, RCR14, SG12, XY12, Yos15, EAG18, Grc17, MZ17].

Obesity [QWM^{+}16]. Objective [AT16]. oblique [Car19]. Observation [BHI12, HHS16, LO16]. Observation- [HHS16]. observational [TK17, YWY17]. Observations [ASLFLR12, ARC13, AAK13, BA13, Che15, Dar12, DS14, HM12b, KM12, Mie12, YCS12, GVTGF^{+}17, Hwa17, JZZZ17]. Observed [HHHZ10, WW10, BS17c]. Obtain [MWH10, SF19]. Obtained [Bla14, SDD17]. occasion [SP17b]. Occupancy [Wil12, Wil15]. occurrence [MHAD19, YDA19]. odd [OAC^{+}17, POCdSB19]. Odds [AD12, CCC10, LW11b, LL14, Lui16, RZZ10, ES17b, HYS18]. off [JK17c]. Olkin [KJR13, BSLF19, Ben17, FS17, MGL17]. Omega [De 14]. omnibus [EAG18]. On-Line [Rod11]. Oncology [Rot12]. One [BK10, CJV13, DS14, JL12, KJS10, LP16, MGG11, SB12b, She13b, WLL10, WL15a, Wu18b, YHL12, yYjHwK14, Zha15a, AK17, AND^{+}17, BC17, LLH17, Llo17, Maz17, SU18, WLF19, Wil17, ZX17]. One- [SB12b]. One-Class [KJS10]. one-dimensional [Maz17, WLF19]. One-Fold [YHL12]. one-parameter [AND^{+}17]. One-Shot [yYjHwK14, BC17, SU18]. One-Sided [WLL10, WL15a, Llo17, ZX17]. One-Stage [CJV13]. One-Way [DS14, JL12, Zha15a, LLH17, Wil17]. Online [HHKM16]. Only [SMF15, HHHZ10]. onto [AB19]. Operating [JS14, MMP^{+}13, OBD16, ZL10]. Operational [Ani16]. Operator [HZP15, WZ10, YA19]. Opportunities [Ras17]. Optimal [AHAH15, ATV^{+}14, ASSH10, AFL^{+}12, AHM17, ASLFLR12, AH13a, BU15, BJU18, BG16, CP11, CYW13, CFSC17, DS13, FLW19, GTFX14, GL16, Kan16, LLJ^{+}12, MNZ13, Mie12, NHLL10, NKKC17, PS14, RL17, SH12, SC18a, SSA18, Tew17, Tom12, VM11, VC13, WL15a, WC16, YCL17, YKLR14, YH19, YPH14, ZZZ10, AJ17, dMBSFB17, CDE19, CD18, DK17b, DK17c, DGH19, Els19, GP16, HW18, NP17, NS18, PGCT17, YM15, YZZ17a, ZLH^{+}19]. optimality [CU19, CTGR17, FT16, WPLL15]. Optimization [BBA17, BDS10, Ell11, SA16, dMBSFB17, LTZ17, MZ17, MM17, MSEN18]. Optimized [STHS16]. Optimizing [CDE19]. Optimum [AS14b, AAP16, DSD19, GKA12, KRAA15, MP13a, MP13b, PM12, PM17, PM18]. option [MBF17, NMS18, XY17b]. Optional [MSR12, SG15b, GKS17]. Options [WSQ15]. Oracally [WX19]. Oracle [FZDF14]. Order [AS12a, AOK13, AB13b, Asi14, BNKM15, BD14a, CM15, EMH15, EK16, FGL10, KMSG13, LOL16, MMK10, SGKM17, VCDB14, XC15, ARY18, AMM17, AMM18, AE18, BS17a, Baz17, BM19, Dem17a, EDE18, ET16, HB16, HSP^{+}17, KE17, KM19a, LLB19, MSS17, PSHP17, RV16, SY17a, Tri18a, Tri18b, YA17, Zar19]. Ordered [BAM11, KK14b, Sin14, SP14, XD15, Agh18, BFB17, GGM17, KR18, SKLK19, WSX18, ZSY17]. Ordinal [AD15, cIC17, Don10, DMMA15, GGNB10, LLJ^{+}12, LX13, MS14, CS17b, GGB17, ZLC^{+}17]. ordinary [EDE18]. Organs [lTHT13]. Original [Bia14].

Ornstein [BRF17, JT19, KM12, ZW17]. Orthogonal [AAK13, WTZ11, AB19, EZ17, GQYY18, PTVP19]. Orthogonally [ASSH10]. Orthonormal [DPCJ16]. Other [CMM14, GN15, SEG15, WANB13, ANB19, Wil17]. Out-of-Bag [ZZZ10]. out-of-control [SXYF17]. Outbreak [Spa12]. Outbreaks [SP14]. Outcome [XLC14, LJ17, YW17b, YWY17]. Outcomes [lTHT13, ZLS14, WHK17]. Outlier [CW12b, DC15, IRHM13, dBMdSF16, TT15, AO17, MC19, YWZ18, ZCY^{+}19]. Outliers [Dei12, Don10, GTPS14, tLB14, MSS10a, Rod16, CB17, Ert18, KCM17, RBC17, UA17]. Output [DD16]. Outputs [BDS10]. Ovarian [SSK16]. Overdispersed [Wei10, KL17a]. Overdispersion [ASM15, Ian16, KIM16a, PGH^{+}18]. Overlap [YC16]. Oversized [Su11a]. Overview [PAM15, UMR17]. Own [Pit14].

P [AMA15, KCM17]. P-spline [AMA15]. Pacific [CD14]. Package [AD15, CNA16, DAD14, AD17, BF17, MN17b, XLX^{+}19]. Pair [DPCJ16, CDE19]. pair-copula [CDE19]. Paired [BSFG10, cCyYnH13, Li14, MC12, Mie12, lTHT13, LSTM17, SC18b, UH17, Usa17]. Paired-Samples [MC12]. Pairs [KM13, LL14, Lui16, PWT11, ZCJ12]. Pairwise [DTZG16]. Panel [BR13, HMS16, AY18a, CMC^{+}17, LL17b, QG19, SMS17b, XW17, YA17]. Panels [WY15b]. paradigm [CH17]. Parallel [Ery14, GL16, JU19, PRA14, WZCJ16, WWJ14, ZC11, Car19, FN17, GRSN19, IMS18, PA18a, WC17]. Parallel-Group [WWJ14]. Parameter [AD12, AB13b, AKM11, AKV16b, AH13b, BLBK16, Bay11, CJLV18, CXW16, CCZZ17, CX17, EF14, EA12, HEF13, HHS16, KRK15, Kom15, KL14, LP16, LTZ17, LG11, MSS10b, NB15, NGT14, QJ14, Ras13, Rod16, SKST14, She15b, SO10, TNO19, Van12, Vrb15, XZZ15, fYyP14, vZS12, AO17, AND^{+}17, ARM^{+}19, AID17, AG17, AVR17, Bak16, Bar17b, Cha18, CSL17, CTGR17, CLY19, DI17, Dem18, DDK16, DPS17, HC19, JT19, Mac17, MCAK19, ÖA19, ÖSY17, PA18a, PM17, PDL19, QEK17, RSBZ18, SaN17, SBZB19, SK17a, SK17b, SWGX18, TGSA18, ZHKR17, ZYW17]. Parameter-driven [HHS16]. Parameter-Tuned [NGT14]. Parameterization [Kom15]. Parameters [AY15, AMW15, ASM15, AG16, BVP16, BKB^{+}15, Cah13, CLW12, CW12a, CM16, Dei12, DTZG16, Ger16, GR10, GDCO15, Gut15, HD13a, Hod13, HKC10, Iqb11, KY15, KU14, KMS15, KNJ16, LH13, LMS13, MM10, Mah12, MNZ13, MGG11, MAR11, SKST14, SD16, TN16, TKCL16, TZD^{+}13, Van12, Wan14, WSW16, WC15, AP17, AA17, BA19, GVTGF^{+}17, GGM17, Gun17, Kar18, KSS17, Lan17, MGL17, MMD19, MMB17, Rak17, SM17, SS17b, SKJ17, TN19, Tri18b, TP17, WK17, WW17, WLX18, Wu17b, Wu18a, Wu19, YT17, ZLA17, ZSZW18]. Parametric [BCGV13, BK10, CK16, DTL12, GP15, GV10, Kor12, LM16b, XYCY15, XY16, XQW^{+}16, XW17, XP19, Zha15a, AEQEH17, DKA18, HHH16, HW17b, JT19, JK17c, MERN18, SKLK19, Sta17, VHV^{+}16, WZT17, WF17, YSS17, ZVHC17]. Pareto [AHAH15, AGM19, AFV16, Bee15, CCZZ17, Gun17, JZ19, LL18, SY17a, SFS12, TCA^{+}16, Van12, YT17]. part [AK17]. Partial [CLY19, DH12, HZP15, KD12, Sun16, Car19, FKA17, LXS19, WF17, WX19, YZW19]. Partially [AHAH15, CW12a, CW15, Hua14, Hua16, Kim16b, LWZH10, Li13, LM16b, PM16, RAN10, TX15, TT15, WCL12, hWjJsH13, WWL15, XY12, YXC10, YGW16, ZX13, ZYHZ16, DZ17, DZX18, FT17, FCC18, FS17, HD17, KAS^{+}19, LZ18, LYG17, XC18, ZN18]. Particle [DH12, YSWM13, WH17, ZHKR17]. partition [GGP19]. Partitioning [PKK12, RMM15]. partly [TÖ17]. patches [MC19]. Path [PK16, SEG15]. Path-dependent [SEG15]. Paths [JZ14]. Pathways [YBW^{+}12]. Patient [MDD13]. Pattern [CMB11, MBL12, WP13]. PD [AFL^{+}12]. PDF [BAN16, BAND16, MM18a]. Penalized [AKM11, AKV16b, Ell11, EO16, JZ14, LZ10, LCN19, WZTX16, YLG16, CHJL17, CTGR17]. Penalty [WSW16, Zha15c, AA18b]. pentavariate [Gui18]. Percentage [AOK13, GNG17]. Percentile [HNV12, KK14a, BJU18, Chv17, NÖ18]. Percentiles [CLJ13, Kri16, LTAJ14, CC19b, CJB^{+}17, Hua17, KH18, PH17]. Perform [LML11]. Performability [SMWR14]. Performance [AOJ12, AMW15, Asl14a, BS16b, CCM13, Cha18, DD17, FSJ12, HKNC17, JM13, LMS13, MMM13, MM10, Mah12, MAR11, PS14, PR19b, RM16a, RMS10, RM16b, SG15a, TRG13, UMN17, Wes14, WCC16, WF17, WA18, ZX10, ZXJ15, AGAB17, BB17e, DK17a, GC17, KD18, MH17b, MRA18, OM17, PFOC17, KKGK19, Rak17, SM17, UA17, WCC^{+}18, YSC19, vZ18]. Performances [PD16, XYZ17]. performing [ELS17]. Performs [CM16]. Period [HNV12, Lui19]. Periodic [BD14a, BD14b, BB17d, GH15, DSD19, Mer18, ZSZ17]. Periodicity [BM10, Isl11, VG15]. Periodogram [HNV12, MMB17]. periods [HYZ19, RRA^{+}19b, Za̧d17]. Permutation [AL15, BPS13, CRS13, GCR12, GGB17, JMCT16, Li14, PS13, RM17, SOSB14a, SOSB14b, AS17, NWB^{+}19, SMMA19]. Permuted [SM10]. Persistence [Che15, FP18]. Person [DTZG16, Fen17]. Personality [MS14]. Perspective [ZCR10, DWW18, GR17]. Perspectives [RF10]. perturbing [ZLZ17]. PH [KD12, BCGV13]. Pharmacokinetic [ASLFLR12]. Pharmacokinetics [DCM16]. Phase [Ery18, OWD14, SG14, ASKT19, BOOTD17, CTH17, GB19, KP17, LS16, MCAK19, RSC16, SM17, SG17, AMW15, FSJ12, KAN17, KN17, LDW14, Mah12, MAT17, NN15, PLW^{+}14, SZMY17, YWZ18]. Phase-I [KN17]. Phase-II [KAN17]. Phase-Type [OWD14]. physical [ZKL^{+}19]. picked [WMH17]. picked-points [WMH17]. Piecewise [Ade12, KLLM12a, KLLM12b, OK15, UHP11, WODY17, ZQL14]. Piecewise-Linear [KLLM12a, KLLM12b]. Pigeon [CBB^{+}17b]. pitfall [SSYS19]. Pitfalls [Mak11]. Pitman [REM19]. pivotal [SK17a]. PK [AFL^{+}12]. PK/PD [AFL^{+}12]. PkStaMp [AFL^{+}12]. Plan [ABAJ14, BU15, LW14, Bal17, NS18, NKKC17, VA17, YLA17]. Planned [RBX^{+}15]. Planning [LBW11]. Plans [DS13, EB14, LHB08, LHB10, LC16, LS14, VB16, ABJM17, GPM18, GA17, KAAJ19, RA17, YCL17]. Plausible [HMH^{+}10]. player [Sab18]. Players [HS12]. Plot [CRS13, FL13, HK10, IMS14, VM11]. Plots [SJP12, CLB^{+}18, LJ17, MO19]. PMCMC [YCS17]. Point [Ade12, AOK13, DH16, DK11, DK13, EK12, GV10, HZP15, JTY10, KNA15, LDW14, MAC11b, PA18a, Pic13, Pic14, VG10, AB19, CN19, CLY19, DPS17, FEP16, GVTGF^{+}17, JJB17, MP17b, ML17, ND17, SZMY17, SA17b, ZZGL15]. Point-biserial [DH16]. Points [ARA13, LX13, ZQL14, GGP19, GNG17, HLWH17, HCW17, QZF19, WLZL19, WMH17, YMS17, ZGR^{+}17]. PoisNor [AD17]. Poisson [AD17, ABV16, AK11, Bee17, BS16b, CPL10, Dem17b, EK12, ES17a, FS12, GDGG16, HHS16, HLH14, KD18, KBJ14, KJ13, KFM16, KMS15, KP10, KÖ19, LL19, LTPT11, Li11, LS14, MP19a, MCAK19, MB17, MP17b, OAC^{+}17, OAB17, PB18a, PGH^{+}18, PK13a, Pic14, RA17, SWM17, SSK16, SKJ17, VA17, WK17, WK18, XLW17, Z.18, Zim17]. Poisson-exponential [SSK16]. Poisson-generalized [MP19a]. Poisson-lognormal [Bee17]. Poisson-mixed [GDGG16]. Poissonity [LW17b]. Policies [GV10, VG10, yYjHwK14]. Policy [BAA16, EY17, MAD17]. Pollution [BNKM15, ABH17]. Poly [RLTJ16]. Poly- [RLTJ16]. Pólya [BYAJ17, CM15, CM19, QLAJB17]. polygons [XY17a, XYL19]. polymerase [HST17]. Polynomial [DPCJ16, KCC16, LZZ16, QDTL16, BVC^{+}19, MH17a, ND17, Sko19]. polynomial-normal [Sko19]. polynomials [Dem17a, DZ17, LW17b]. polytomous [dSBHM19]. Pool [GAK16]. Population [AFL^{+}12, AS14b, BK11, BJ14b, CSG15, CSS13, DX16, GK14b, HJST15, JW12, KRAA15, KGA11, KG14, LA13, MA12, OD16, RT12, RSC16, SG14, SSmK14, SSA15, Wes17, Wil15, GP17, GKS17, KuAH19, LSD17, PS18a, SG17, ST17, SP17b, VND19, XD17, Xue17, YW17b, Za̧d17, dlHMRB17]. Populations [BS15, BSL15, CCM^{+}16, CWSW16, Ng13, ZP14, ZWJ10, ZWCC12, AJ18, CPL17, GG17a, HM17, JK17b, Naj17, Naj19, Ng17, Tri18b]. Portland [TB17]. Portmanteau [CW14]. Positive [GBP16, EGO^{+}17, GKA19]. positive-definite [EGO^{+}17]. Positively [BK10]. possible [Nam17]. Possibly [TVS13, HS19a, LNUT17]. Post [PAM15, DPS17]. post-change [DPS17]. Post-hoc [PAM15]. Posterior [Wal11, Wal14, Ami19b, FA17, RV16, YY17a, YY17b]. Posttest [WL15b]. potential [MO17]. Potentially [BDS10]. Power [BMPZ13, CLO13, CMM12, ES16, FK15, GAMA15, HA18, LDK17, Mur15, OK15, SR12, SC18b, Tan16, VPSV13, WL11, WL15b, ZCR10, ABS^{+}18, AID17, Bah18, BMJ17, CTXW17, CMC^{+}17, CH17, Dem17a, DSMJ18, FZ17, GRM17, HWXH18, Kar18, KH18, MJ17, Nou17b, Nou17a, TNO19, VRN18]. Power-Generalized [VPSV13]. Powerful [OY16, PLLF16, AS17, MO18]. powers [Naj19]. Practical [BGM13, BRF17, MRF17]. Practice [SZFM11]. precise [Dem17b]. Precision [vdH10, Wes17, WPTL17, ZYW17]. Predictability [DKV15]. Predicting [MSB11, MDD13]. Prediction [CCM13, EZ15, EDE18, FSJ12, GVC10, HdL10, JMCT16, LaM14, LOL16, LqQlX11, MSR12, MS10, SB12b, Ste12, TG10, VAR19, WL11, Yos15, AMM18, GRY17, KR18, LW17a, RL17, Za̧d17]. Predictions [CMM14, HJ18, MWSB18]. Predictive [Ani16, BJ14b, GVC10, Yam16, DK17a, RN17, RW17, SB17]. Predictor [ATV^{+}14, dJvBS16]. predictors [KD18, LASX19, LL17b]. Preface [CD14, FP14, GA15, Mel12a, Mel12b, MMS16, RL16, SZ16, ZS11]. preference [GGB17]. Preferences [GBTV16]. Pregnancy [BSV13]. Preliminary [KABM17, Wu17a, XY12]. Prequential [EAG18]. Presence [AAK13, ARA13, BLN12, Cha13, Dei12, GSSCR16, HAB16, JKBR16, KHH12, KGS14, OK15, SSA15, AMAA17, AP17, CBB^{+}17a, FL19, KuAH19, KSG18, Lui17a, MAT17, NBOM17, PS18a, Sen18, ST17, XD17]. presentation [YWY17]. prespecified [ZSGW17, ZSZW18]. Pretest [WL15b, ZLA17]. Pretest-Posttest [WL15b]. preventive [FN17]. Price [Bia14, Bia17b, Bia17a, GSP^{+}17, KH19a]. Prices [CCC16b]. Pricing [NMS18, MBF17, XY17b]. Principal [Che14, DB16, HWX16, KK10a, UB14, WL10, WY15a, dWD16, BB17e, BLB19, Cha18, CYY18, MO18, MB17, Man17, MCC18, PUN17, PUN18, SDS17, WA18, YMS17]. Prior [ATV^{+}14, CW15, DW16, HSR10, OY16, RV15, WL11, CK18, FA17, KLK19, LN17, ZYW17]. Priors [Lon14, ATC17, Ami19b, CTH17, SS17a]. Probabilistic [Rai16, YW16, YC16, BB17a]. Probabilities [AK11, KR12, MYS^{+}16, YCSD10, YR11, MP17a, WHK17]. Probability [AJ13, Bee15, SNO16, SO10, SW13, TK16, VPSV13, WL15b, Zha15b, vdH10, CLB^{+}18, CCZZ17, CCH18, CBH19, KPG^{+}17, MO19, MH17a, SDD17, TNO19, WB18, YAT19, Zar19, ZLL17]. Probit [CCC16a, LMS13, MRF17, ZLC^{+}17]. Problem [Ang13, CIB10, CY12, HS12, KGA13, KK10b, Pap14, Rai16, SKM^{+}10, Su11a, VKB10, Wil12, Wil15, ATC17, dMBSFB17, DRAM18, NBOM17, Sab18, SOY17]. Problems [DD16, LP10, MG12, McI14, SGB16, TL15, BBA17, DN18, MM17, MSEN18, Tas18]. Procedure [CB14, DZC10, DK13, HAT15, ND10, NZL13, OBD16, Ouy13, Par11, Par15, RBC17, Sin14, TNS12, WLL10, WL15a, ZLZW16, Agh18, DPS17, GRM17, GY19, Lui19, OM17, SNT17, Wan17, WSX18, WCC^{+}18, YKLL17, ZVHC17]. Procedures [AM13, BSFG10, GBB^{+}12, GMG13, HHKM16, HM12b, LRVF10, MMP^{+}13, Ng13, ZZ13, GE17, KNLK19, Ng17]. Process [AS12b, Aus10, AS15, BKJ10, BDS10, CT12, CMB11, CW15, DZC10, DN16, Gad14, GLS12, HLH14, HA12, JL12, KMSK15, KL14, Lee10a, LJ12, MMK10, MS14, PK13a, PB16, RYW16, WZ10, WL11, WBG16a, WBG16b, YKY10, Yan13, ZZW11, ZZ13, ZW15b, ZZGL15, AAAJ18, AA17, AAMJ17, AMJ17, Bal17, BW19, CM19, DSMJ18, DWW18, ELS17, HWZ18, Haq19, KTA17, LK17a, LLWY17, ML17, PGCT17, RPRPMG^{+}18, SAAJ19, SSA18, SXYF17, XP19, ZHLW17, ZW17, ZM17, ZYW17]. Process-based [WBG16a]. Processes [AZSA11, BM10, CW14, EK12, Far16, FGL10, Góm13, HXB13, HAAQ11, KM12, KFM16, KJS10, LPZ13, Mac11a, MAC11b, NT10, PB10, PS14, Pic13, Pic14, SEG12, SEG15, Wan10a, BRF17, BM19, CLY19, EAG18, ESAD17, GA19, JT19, KAN17, KN17, MP19b, Ott17, RRA19a, RSS17, UA17, WPTL17, Yeh17, ZZ17, ZM17]. processing [IMS18]. Product [RSC16, BM17a, BS17a, Bal17, GGP19, KLK19, Naj19, ST17]. Production [OG14, CX17, KABM17]. Products [LC16, WCC16, WGX18, WCC^{+}18]. Profile [AMW15, Mah12, MDD13, NA12, ONS16, SID13, ESAD17, GB19, KLK19, MAT17]. Profiles [AM13, HA12, KNA15, NN15, AARQ19, GAK19, KN17, KA19b, MAT17, MCAK19, ND17, SZMY17, SA17b, ZSGW17, ZSZW18]. Profust [AK19]. Programming [EF14, GKA12, HEF13, MG12, Man15b, Mey13, HWW18, YAT19, YZZ17a]. Progression [HAAQ11]. Progressive [AHAH15, AY15, BSL15, BR16a, CLJ13, CCL10, DAA13, HK13, KBAD11, KSGM13, SSK16, BSLF19, CD18, GP17, KSS17, KTRA17, MR18, NKKC17, RGP17, SKK17, SK17b, WGX18, WCC^{+}18, ZS17, CSY17]. Progressively [AYR17, ARY18, ARAM15, DDK16, LBW11, BK18, BS17a, EDAAAM17, GRY17, HK17, Nou17c, SK17a, WSZ17]. Projection [Sim11, Tou11, AB19, Hir17, Liu17a, dlHMRB17]. Promotion [GBPDL16]. prone [Hua16, HD17]. propagation [PV17]. Propensity [CZL16, TK16, GMR^{+}18, NH18, PW17]. Properties [ARM^{+}19, Ben17, CBB^{+}17a, CC19a, CT12, DKRL17, HNV12, ÖW12, RNJ16, RTJ17, Rut15, SBS17, ST13, TCA^{+}16, dALA17, CS17a, Gen17, MJ17, SBNG18, SJR19]. Property [FZDF14]. Proportion [DX16, MPG16, ZWZC12, CY17, Lui17a, VND19, XD17, Zha17a]. Proportional [AD12, AJ13, LW11b, SHM12, SLX14, AK17, AYR17, ES17b, Kiz19, Lan17, PDL19, PMNF18, ZS17]. Proportions [AH13a, DS13, MCM16, OY16, SR12, AP17, EFA17, HW17a, HA17b, HW18, LSTM17, Roo17]. proposal [CB17, OM17]. Prospective [CZL16]. Prostate [MSB11]. Protection [KD12]. Proteomic [MDD13]. Protocol [LGM12]. prototypes [AY18a]. Proved [Bla14]. PSD [ZM13]. Pseudo [JY10, RV16, WZTX16, Fan18, XZF19]. Pseudo-Exogenous [JY10]. Pseudo-likelihood [WZTX16, XZF19]. Pseudo-posterior [RV16]. PSO [LTZ17]. psychophysical [MSK^{+}17]. public [MAD17]. Pursuit [Sim11, Tou11]. put [NMS18].

q [KCM17]. Quadratic [ALH14, KCC16, Mey13, Mie12, PM12, WSW16, Wes14, ZP14, Del17, HWW18, RCM17, YL17]. Quadratically [Ell11]. Quadrature [GP16]. quadrivariate [Gui18]. Qualitative [NHLL10]. Quality [AMA15, VPF14, AGKJ17, BBA17, CB17, PR19a, PR19b]. quantification [SJR19]. Quantify [CBDD16]. Quantile [Alh16, DAM15, GTPS14, KCC16, Kur16, MA12, NB15, NÖ18, Par11, SNM16, YY17c, hZxNL14, AA18b, AA19, BEJC17, DZX18, DB17, ESSH18, Hwa17, LY17, MWL17, Per19, RN17, SP17a, TSL^{+}19, XCJ^{+}18, XT19, YY17b, YSY19, YZW19, ZLZ17]. quantile-based [Hwa17]. Quantiles [Par15, AJ18, Das17, HM17, KNLK19, Wil17]. quantitative [BS17b, HST17]. quantities [SK17a]. Quantization [SFLC14]. Quantum [MSS16]. quartile [AG19]. Quasi [CCY17, CHJL17, CU19, MOS18]. Quasi-likelihood [CCY17]. quasi-Newton [CHJL17]. quasi-Semi-Latin [CU19]. Question [Lin15]. Queue [KD12, Ban17, CQH17]. Queueing [ZX10]. Queues [ABB12, AC18, BR18, IOTBO17]. queuing [OTZA17, SSOTB19]. quick [BJU18]. Quintile [Kpa15]. QVF [ZJNG13].

R [AD15, Ban17, DAD14, AD17, BF17, CNA16, ELS17, MN17b, XLX^{+}19]. Rahim [KM19b, YC10]. rainfall [MHAD19]. Random [AD12, AM14, ARKB13, AB13b, AS14b, BB14, Bla14, CC10, CWYF12, CZL16, DS14, DZ16, GBPDL16, GGZG17, GGNB10, GTPS14, Gut15, HXX15, Hil17, JL12, KLLM12a, KLLM12b, KSS13, KM13, KGA11, KK16, Koz14, Kri16, LTQ15, Lon14, Mac11a, MBL12, MAM11, MRM10, Phi12, Ras13, Roy12, TNS12, VL13, WA14, WBG16b, Yar12, ZWZ^{+}14a, ZDX14, ZW15b, AD19, AMKS19, EH18, FT17, GKA19, HJ18, Liu17b, LLH17, MP17a, RPRPMG^{+}18, Sun17, TN17, VAR19, WC17, ZHKR17]. Random-Effects [Phi12, Sun17]. Randomization [HAZSK16, KHH12, TC14, BS17b, JLGD17]. randomize [HSP^{+}17]. Randomized [AB19, BRT^{+}16, GCR12, HSR10, LSS13, SG15b, WWJ14, ZLS14, AM19, BSH17, GKS17, JLGD17, LDK17, MOS18, WSCW18, XSA^{+}17, ZSY17]. Randomly [LM10, Roo17, WZG15, YXC10, AY18b]. Randomness [GR10, KM13, SI13, SI14]. Range [BGM13, SSS14, AHM17, AR17a]. Rank [AL15, CCM^{+}16, DAY19, GR10, GA12, HWB13, LZS10, LZZW13, LLZ16, vZS12, AE17, AMKS19, FZ17, Hua17, Lue17, WMH17, ZSY17]. Rank-Based [LZZW13, DAY19, AMKS19, WMH17]. Ranked [CXW16, MA12, Mah15, OD16, SV13, SKST14, SMB12, AM17c, AS17, AGAB17, KR18, ÖEG17, SAB19, SB17, VND19]. ranker [OD16]. Ranking [CMM14, ZDX14, Ul 17, YLM17]. Ranks [KHH12, LZZ15, Mur15, Tas18]. Rao [DH12]. Rao-Blackwellized [DH12]. Rare [MCM16, ZP14, SK17c]. Rasch [DTZG16]. Rate [CLO13, GBPDL16, HAT15, Kan16, LJ12, LH14, She15b, YCH16, dFR13, dR11, AA17, Fan18, FP18, GRM17, HPS17, HA17a, JJB17, Kim17, Kiz19, MJ17, MP19a, MNBB17, PB18a, TN19, TK17, Wan17]. rate-geometric [HA17a]. rate-power [MJ17]. Rates [BLN12, BS16b, Góm13, LTPT11, LRVF10, RH11, ASKT19]. Ratio [CBDD16, CCC10, CH11, De 14, GK14b, Ima11, JW12, KG14, Kpa15, LH13, LL14, Lui16, SG14, ST17, SVWH10, sS12c, Sol16, SW11b, VG10, ZW15a, ZWCC12, HA17b, Hua17, HYS18, Kao17, KRK18, Lui17a, Par17, SPH19, ZHLW17, ZVHC17]. Ratio- [CBDD16]. ratio-type [SPH19]. Rational [SW13]. Ratios [AK15, AS12a, CCC10, KBJ14, RZZ10, SD16, TK17]. ray [MCC18]. Rayleigh [ARK17, DDK16, KR18, KL14]. RCA [AB13b]. RCCA [Pol18]. Re [MMP^{+}13, XZ16, CTH17]. re-analysis [CTH17]. Re-Estimation [MMP^{+}13]. Re-sampling [XZ16]. reaction [HST17]. Real [JKKR16, HST17, LqQlX11]. real-time [HST17]. Realized [BZW13, SS19b, SSY18]. Realizing [AH13a]. Reasonable [SY17b]. Reboot [KS16, LKYH17]. Recalculation [PK13b]. Receiver [JS14, OBD16]. reclassification [KKGK19]. Recognizer [CMB11]. Reconsidered [GS11]. reconstructed [MO17]. Reconstruction [MNZ13]. Record [AN16, BSZ16, CJTL18, Kiz19, SAB19]. Records [AB13a, GLS12, AVR17, PS19a, Ste19]. Recovery [BKJ10]. Recurrence [BS17a]. Recurrent [BS10, FEP16, BDK17, sS19a]. recursive [HC18, NBOM17]. Reduced [GHRM11, GBP16, LA13]. Reduced-Bias [GHRM11, GBP16]. Reduced-Variance [LA13]. Reducing [VG15]. Reduction [LY16, Wan10b, BB17c, IOTBO17, OTZA17]. Reference [ABAJ14, KHH12, vZS12]. refined [BOOTD17, IOTBO17, LJ17, OTZA17]. Reflection [Fen16]. regard [PTVP19]. regarding [Sen18]. Regenerative [AR10, MMN12, Ngu17, TC14, WN15b]. Regime [HKC10, ÖSY17]. Region [WC16, WN17]. regions [Gui17, KLS17, YW17b]. Registered [OY16]. Regression [ATV^{+}14, ALP16, Alh16, AM13, ATB14, AHGJRC16, ARA13, Asl14a, Asl14b, AS14b, BZ15, BA13, BA14, Bus15, Cha13, CA10, DB16, DDD16, DW16, EF14, FWY14, FS12, GL11, GVC10, GR10, GDGG16, GA12, GK15, HAT14, HKC10, HK10, HHH16, HA12, HN15, HM12c, IRHM13, JK16, JR12, JZ14, JKKR16, KR12, KJ13, KMS15, Kim14, Kur16, Kyu16, Li10a, LM10, Li11, LLJ^{+}12, LZ14, LMS13, Lu14, MSK10, MSS10a, MRF17, MN15, Mie12, NAK13, NHLL10, ONa13, OCH11, Özk14, PL13, PR14, PWT11, PKK12, PK16, QMLS16, RMS10, RM16b, Ron10, RAN10, RA16, SaJMF16, SS12a, SFG15, She13b, She15b, Sim11, STK15, SC13, SZS10, Sun16, SGB16, TG10, TRG13, TZD^{+}13, UB14, US10, Van12, Van15, VVHP^{+}16, WC15, WY12, WL14, WY14, WY15a, Wu16, Xu10, YXS14, YLG16]. Regression [YGL12, You13, YT12, hZxNL14, ZAK14, vZS12, AK17, AA18a, AA18b, AA19, AP18, AO17, ARM^{+}19, Alk17, ATC17, AAC17, AJ17, ABH17, AAW17, Asa17, AG17, AY18b, BEJC17, BRFD18, BCN17, CBB^{+}17b, CHJL17, CR17, Cha18, CLB^{+}18, Chv17, CMBS18, DRAM18, DA18, DZX18, DZT19, DGMC18, EH18, ESSH18, EKG17, Ert18, FK18, FKA17, GE17, GB17b, GKS17, HWXH18, Hil17, ISS17, Iva18, JG17a, JZZZ17, JZ18, JS19b, JCL18, Kar18, KA19b, KH17, KP17, LASX19, LJ17, LY17, LBP17, LYG17, MP19a, MCAK19, MB17, MWL17, Mat19, MP17b, MMB19, NKKC17, OM17, ÖA19, OAB17, PB17a, Per19, POCdSB19, QDT18, ROC^{+}18, RN17, RL17, RH17, SJC18, SP18a, SB17, She17c, SA17b, TÖ17, TSL^{+}19, TGSA18, TÖ18, WLZL19, WA17, Wu17a, WF17, WA18, XL18, XCJ^{+}18, XT19, XZ19, YE17]. regression [YY17a, YY17b, YY17c, YSY19, YZZ17a, Yil19, YLL17, YAR^{+}19, YY17e, YZW19, YSS17, ZLZ17, ZLHW19, ZH19, Zim17, dODD19, Mac17]. Regression-type [Van15]. Regressions [Asi14, BJ14b, CK16, Önd16, PCN14, SPCNL16, PV17, SCN17]. Regressors [Asl14b]. Regularization [AH13b, MSS16, Liu17c]. Regularized [ESSH18, KK10a, PUN18, TSL^{+}19]. Regulatory [LW16]. reinspection [BM17b]. Rejection [LDS10, BHF17, ML19a, ML19b, MNBB17]. Relabel [WY17]. Related [BMN12, BG15, YBW^{+}12, HC17b, KABM17, SBNG18]. Relational [TKX13]. Relations [SVWH10, BS17a]. Relationship [Mak11, McI14, KL18]. Relationships [LM10, SWL14]. Relative [PK13b, Tou11, EA17, HST17]. release [DSD19]. Reliability [ARKB13, Bar11, BD17b, CD14, EK10, Ery17, FTE17, GAMA15, JWW14, KFO14, LH14, SNM16, SN13b, TC14, WZCJ16, WGX18, YPH14, ZW14, AGM19, AK19, Bar17b, CTXW17, Ery18, GLGG17, HS17, PA18a, SAB19, VRN18, ZSZ17]. Reliable [LLWY17]. Remaining [BGU15, Hod13]. REML [AMKS19, YZCW15]. Removal [SSK16]. removals [KSS17]. Renewal [AC15b, Pic16]. Renyi [SBH18]. Repair [TC14, AD19, LKYH17]. Repairable [KS16, LKHH11, LKYH17]. Repairing [ARKB13]. repairman [WCSL17]. repeatability [PTVP19]. Repeated [KAS^{+}19, LRVF10, RAN15, Zha11a, GS17, RRA^{+}19b]. repeatedly [BS17c]. Repetitive [LW14, KAAJ19, NS18, SAAJ19, YLA17, dMdCQSH17]. replacement [FN17]. Replacing [ARKB13, GL10]. Replicated [SOSB14a, CCRX18]. reported [HBC17]. Represent [Liu13]. Representation [HK10, LZ16, WTCZ10, WTZ11, LNT17]. Representative [QZF19]. Reproductive [TZ14]. require [Sen18]. Requirements [LSS13]. Resampling [AS12a, BAA16, DKV15, EMH15, GN15, KOC16, Sha15, Su11b, GGZG17, ZLZ17]. Resampling-based [DKV15, KOC16]. Research [CLLH16, Grc17, JWW14]. Residual [BSK19, CTTT16, IMS14, LJ17, WBG16b, Zar19]. Residual-Based [IMS14]. Residuals [BCGV13, AAC17, CLB^{+}18, JK17c, Per19, ZYX18]. Resin [KMSK15]. Resistance [GTPS14]. Resistant [TT15]. resisting [JG17a]. Resolution [Aok15]. respect [SDD17]. Respondent [MS14]. Response [AMP10, AKS13b, AK11, GGNB10, HB16, HSR10, KGA13, KP10, LSS13, MMM16, RCR14, Roy12, SCpL12, SG15b, SOSB14a, WPLL15, Wei11, AM19, BS17b, BSH17, Del17, FT17, Fan18, GKS17, HSP^{+}17, LW17a, Liu17b, McI18, MOS18, MSEN18, MDX17, PS18a, PM18, RBC17, WSCW18, XT19, BKSM18]. Responses [HD13b, KCC16, RAN15, SSA15, jWzM13, BBA17, FKG19, KÖ19, Liu17b, PM17, Z.18, dODD19]. restratified [KP17]. Restricted [AAW17, CD18, KÖ19, LY13, NAK13, STK15, hWjJsH13, WY14, Wu16, Baz17, Din18, Mer18, Pol17, RH17, SGKM17, Tri18b, WA17, Yil19]. restrictions [GRB17, Yil17]. Resubmitted [BU15, RA17]. Results [GZ10, LP16, LLJ^{+}12, Mac11a, ÖW12, OWY13, VCDB14, Els19, McI18]. retesting [HW17a]. Retinopathy [CCLT14]. retrial [IOTBO17]. retrieval [Pol17, Pol18]. Retrospective [GV10]. Returns [SN13a, HYZ19, SNT17]. reversed [HPS17]. Reversible [NZL13, PLT18]. Review [SI14, EL18, YY17e]. Reviews [LLS16]. Revised [LKTW12]. revision [CYY18]. Revisit [RR13, CPL17]. Revisited [BGM13, Dut16, PSS10]. Revisiting [HW17a]. Reweighted [GA12]. Ridge [ATB14, Asl14a, EF14, EA12, GL11, HEF13, JK16, JR12, KR12, KMS15, LMS13, MSK10, MSS10b, NAK13, RAN10, RA14, RA16, AAW17, AG17, BR17, Ert18, FKA17, KD18, Kar18, KÖ19, WA18, YE17, Mac17]. ridge-type [WA18]. Riemann [ZGE17]. Right [BM14, CL15, CTTT16, ES16, LM10, Par15, She14, She15b, AEQEH17, AY18b, FL19, JJB17, Kao17, She17b, SC18b, WODY17]. Right-Censored [CL15, LM10, She14, She15b, CTTT16, ES16, AY18b, FL19, She17b, SC18b, WODY17]. Rigorous [RBX^{+}15]. Risk [Aus10, BSV13, BZW13, BG16, CNA16, CZL16, Che11, CM15, LYY14, Pit14, PK13b, SS12b, She14, UHP11, HKNC17, YCS17]. risk-adjusted [HKNC17]. Risks [HLLV14, PR15b, She12, AYR17, ARY18, BT17, HC19, HK17, MN17b, SP18a, WSZ17, YH19]. RMM [BKSM18]. Road [MNH10, DARMM^{+}17]. Roberts [DPS17, ZLZW16]. Robust [AB18, ALP16, AO17, AA16, AAK14, ARA13, Ati16, AS15, BG17, BB11, BLB19, CCRX18, CK19, Che14, CC19c, DAM15, DM19, DB16, DA18, ES17a, EKG17, EB18, GTPS14, HS19b, JK16, KY15, LLZ16, MCM16, MN17a, MSEN18, MX17b, ÖW12, PB17a, PKK12, PK16, PLZ15, Pit14, RRA19a, RMS10, RM16b, SKK17, SK17b, SFS12, TVS13, Tom12, TGSA18, WZTX16, Xue17, YXY17, YY17d, YY17e, ZLW17, AMKS19, BPG17, CR17, CCH18, CKKP17, DL17, FK18, GY19, HLWH17, JG17a, MMB17, NC19, OM17, Res19, RH17, UMR17]. Robustness [De 11, RCR14, SL14, AEQEH17, DPS17]. ROC [JS14, cCyYnH13, Li12, LQW^{+}16, PFOC17, KKGK19, SG15a, WMK14, WMK16, WMK19, YLM17]. Roger [Sta17]. Role [CDGML12, PTVP19]. rolling [Ros19]. Root [AFS14, CYW13, CNPW15, Lak11, LML11, LO16, LS10, OK15, SSS14, Wu10, ZYM13, Gür19, LAG17, MMV18, Sen18]. Roots [HMS16, Sol16]. Rotatability [HB10]. Rotatable [KMSG13]. rotation [PTVP19]. round [Cal19]. router [Els19]. routines [ELS17]. RROS [AS12a]. Ruin [HS12, Sab18]. Rule [LKTW12, MMV18, SSYS19]. Rules [GP16, LK15b, LK18b, McI18, Rak17, ZSGW17]. Run [AMP10, HXB13, LKTW12, TKCL16]. Running [lKW13]. Runs [Gad14, LK15b, LKTW12, LK18b, Rak17, SKLC18, ZSGW17].

S [LK15a, CQH17]. Saddlepoint [AE17, AL16]. SAEM [KK14b]. Safety [MNH10]. Sample [AJM10, AL15, ANA14, BK10, BKB^{+}15, Bus15, CH10, CSS13, CYW13, CM13, CNPW15, De 11, Ger16, GBO14, Hay14, HW14, HTN14, KS15, KRAA15, KGA13, KBAD11, KK10b, Lee10b, LSS13, tLB14, Lon14, MLZ15, MMP^{+}13, ND10, Önd16, Ouy13, ÖW12, OWY13, PK13b, RMM15, SM10, SKST14, SB12b, SKM^{+}10, SFS12, Vol14, WSCW18, ZP14, AM17c, CS17a, DÖ17, DKV15, DB18, EGO^{+}17, FZ17, GRY17, HT17, KKYW17, KP17, KEB17, LK18a, LK18b, MO19, Mai17, MH17b, PZ18, SMMA19, SY17b, Sta17, SC18b, Tas18, VAR19, WMK16, Wu17b, Wu18a, Wu19, YKLL17, ZVHC17, Zim17]. Sampler [lKW13, CFSC17]. Samplers [CCC16a]. Samples [ALH14, BMN12, BS16a, CP11, ES15, FK15, KNJ16, Lee10a, LHB08, LHB10, Lon14, MA12, MC12, NGT14, SS10, WCC16, BK18, CJTL18, Jov17, NWR17, RSR17, SJC18, SB17, SF19, Tri18b, UH17, VAR19, XY17a, XYL19, YCL17, Zha17b, vZ18]. Sampling [AJ13, ABAJ14, AS14b, BU15, CXW16, CYW13, DZC10, DS13, GKA12, GL15, HAB16, JW12, KSG15, KGA11, KK16, LS16, LDS10, LK15a, LK15b, LHB08, LHB10, LW14, LC16, LS14, LL14, Lui16, Mah15, MWK16, MMN16, Ngu17, NGT14, NCF18, OD16, PM16, RSC16, SV13, SG14, SKJH11, STHS16, SCH14, VB16, VHV^{+}16, Wal11, Wal14, WTLWL16, XCJ^{+}18, ZWZ^{+}14a, ZDX14, AM17a, AS17, AAAJ18, AGAB17, ABJM17, Asl19, BOOTD17, Bal17, BJU18, GGZG17, GPM18, GA17, IOTBO17, JK17c, KSG18, KAAJ19, KKYW17, KR18, LK17b, LK18a, LK19, LWT17, ML19a, MOS18, Mor19, NS18, OTZA17, ÖEG17, RA17, SAB19, SAAJ19, SKLC18, SM17, SG17, SSA18, SP17b, VA17, VND19, XZ16, YLA17, YY17a, YY17b, YCL17, YSY19]. Sampling-based [MWK16]. SARFIMA [PAS19]. SARIMA [AU17]. Satisfaction [Bon16]. Satterthwaite [vdH10]. Saunders [BZ15, BA19, SCN17, SS16, WGX18]. SAVE [HS19b]. SCAD [JZ14, WSW16]. Scalar [SS12b, ZL17]. Scale [ALP16, CXW16, DC15, GEVA10, KK10b, LH13, Mar13a, PWT11, SKST14, SMWR14, Wu14, EK16, GGM17, KE17, PA18b, RM17, RB16, SMMA19, SGB16, Tri18b, XZ17, XCJ^{+}18]. Scaled [HN13, Han19, MN18, RSR17]. Scales [CCM13, Mar12]. scaling [BSK19]. Scan [Mar15, Spa12]. scenarios [BB17e]. schedule [DSD19]. Scheme [LJ12, Ate17, GP17, SAB19, YKLL17, ZCY^{+}19]. Schemes [AN16, HK13, LP10, SKJH11, AEQEH17, BM17b, GPM18, SK17a]. School [RM16c]. Score [CZL16, TK15b, VL16, XLW17, BH17, GMR^{+}18, NH18, PW17]. Score-Based [TK15b, PW17]. Scores [TK16, TKKC14, Fen17, GMR^{+}18, TK17]. Scoring [CMM14, MMV18, Zen17]. Scrambled [SSA15, VC15]. Scrambling [VC13]. Screening [GAK16, WLL10, WL15a, DK19]. Search [AAP16, Koz14, LMS13, BVC^{+}19, EZ17]. Searching [YPH14, ZHKR17]. Seasonal [CNPW15, Ami19a, AU17, Whi18]. Second [GL14, KMSG13, MSR12, ESAD17, SC18a]. Second-Order [KMSG13]. sectional [HMS16]. Seemingly [Asi14, RA14]. Segmentation [AAT15, CJV13, ND17]. select [CDE19]. Selected [HAAQ11, WZG15, SDD17]. Selecting [AR12, AZSA11, GP17, Ng13, XGTM18, Ng17]. Selection [AKV16b, AH13b, Cha13, CKC^{+}14, ES16, EO16, FX12, FZDF14, GBO14, GHY14, HEF13, HWB13, JKBR16, KMS11, Kur16, LNL11, LQW^{+}16, Lin15, LLS16, LSLW11, PUN16, SOBH10, SFG15, SLB11, Ste12, TX15, TW11, US10, VKB10, VVHP^{+}16, Wu10, Wu14, WZTX16, XC15, YXC10, ZZZJ14, Zha15c, ZWZ14b, BCN17, CNK17, DGMC18, FCC18, FA17, GE17, HWW18, HZ19, IMS18, JS19b, LY17, LYG17, MP17b, MRN17, PLT18, QQYZ17, SP18b, TTC19, UMR17, Wei19, XYZ17, XLX^{+}19, XZF19, YY17c, YY17d, YZW19, YA19, ZLL17, ZS17, ZH19]. Selections [CCC16a]. Selective [ARKB13, UMR17]. Self [HC18, KA19b, CLY19, HBC17, ZKL^{+}19]. self-exciting [CLY19]. self-reported [HBC17]. Self-Starting [KA19b]. Self-weighted [HC18]. SEM [FS12]. Semantic [SWL14]. Semi [ABB12, CU19, HLLV14, JT19, KL17b, Mac11a, MTL10, MM17, MERN18, Z.18, ZLW17]. Semi-Competing [HLLV14]. semi-continuous [Z.18]. semi-expensive [MM17]. Semi-Markov [ABB12, Mac11a, MTL10]. Semi-parametric [JT19, MERN18]. Semi-supervised [KL17b]. semi-varying [ZLW17]. semidefinite [YZZ17a]. Semiparametric [CIB10, CK16, DOZ17, EO16, FEP16, HAT14, Kim16b, Kyu16, Li10a, MB17, RAN10, RA14, RA16, She17b, SID13, TS11, TT15, ZHL14, ZQL14, BT17, Din18, DZT19, GB19, HF18, Lin17, RH17, She17c, ZS17]. Semisupervised [CKC^{+}14]. Sen [SKJH11]. Sensitive [GSSCR16, HAZSK16, KGS14, GKS17, PSHP17, SKLC18, SPH19, SK17c]. Sensitivity [BG16, Che14, CGST17, FT17, Lan17, LZ16, VRN18, WTCZ10, WTZ11, WPLL15, YKSC16, dSBHM19, Ami19b, DKA18, LN17, WZT17, WLX18]. sentencing [NBOM17]. separate [AdBPdBP19]. Separation [Afi14]. Sequence [MBL12, Ras13, VC13, VG15, WP13, GVTGF^{+}17]. Sequences [CDL12, VC15, EDE18]. Sequential [AM14, DZC10, LP10, LZZW13, MDX17, ND10, NYD16, QDTL16, Rot12, Ste12, WPLL15, WP13, WLL10, WL15a, ZL10, ZQL14, ARY18, BD17a, Els19, HW18, Mar19, MMB17, Ros18, ZCY^{+}19, dMdCQSH17]. Serial [BZX11, DKV15, LWH^{+}15, YGW16, ZB16, Orz17]. Series [CD12, CLO13, CC10, CNPW15, CB14, Ery14, GL14, Isl11, JH14, LM16a, LTQ15, MS10, Rut15, SB12a, TW16, Wei10, ABS^{+}18, Ami19a, Ami19b, CSY17, CCZ19, DAY19, DB17, Elh17, ELS17, FN17, FCC18, FS17, GRM17, GRSN19, Gór18, JZ18, KTA17, KH19a, KR17, LL19, LNUT17, MJ17, MC19, NWB^{+}19, SS17b, SGNS17, SKJ17, TS17, Tew17, VRN18]. Serum [BSV13]. server [AC18, CQH17]. Service [AAT15, KD12, MSR12]. Services [Bon16, GDSCO19]. Set [CXW16, KHH12, MA12, Mah15, OD16, SKST14, TP15, ZDX14, AB19, AM17c, AS17, AGAB17, KR18, ÖEG17, SAB19, SB17, VND19]. Set-up [TP15]. Sets [Sat12, MP17a]. setting [DWKL19, HCW17]. Setup [SS12a]. Several [AFvRvR15, CCM^{+}16, CPL10, CSK16, CTTT16, GMG13, Mar13a, SaJMF16, SR12, AJ18, CPL17, EKK19, GG17b, GG17a, JK17a, JA17, XYZ17, ZYW17]. Shanghai [SDZL15]. Shannon [CZ19]. Shape [ADLS10, BKB^{+}15, KU14, DZ17, Din18, HC19, RTKB18, WB18, ZZS17]. shaped [SKK17]. Shapiro [HT14]. Share [Kpa15]. Shared [HD13a, HPS17]. sharing [GPM18, WZCJ16]. Sharp [PR15b]. Shelf [SKHL15]. Shelf-life [SKHL15]. Shenzhen [SDZL15]. Shewhart [PS10, TKCL16]. Shewhart-Type [PS10]. shift [AU17, GRB17, KM19a, MR18, dBMdSF16, SGNS17, UA17]. shift-dependent [KM19a]. Shifts [CMB11, OK15, YAA15, ZLZW16, AHM17, KAS^{+}19, SKLC18]. Shiryaev [DPS17, ZLZW16]. shock [Ery19, PMNF18, Tas17, WCSL17]. shocks [BSLF19]. Short [Isl11, SGNS17]. Shortest [CMM12]. Shortfall [CNA16, DB18]. Shot [yYjHwK14, BC17, SU18]. Shrinkage [ATB14, AG16, RV15, TP15, Wan10b, KL17b, ZLA17]. shrunken [YA19]. Shurkur [UMN17]. Si [ZGR^{+}17]. Si-GARCH [ZGR^{+}17]. side [SKLC18]. side-sensitive [SKLC18]. Sided [AH13a, BPS13, Ima11, Ima13, WLL10, WL15a, KG17, Llo17, ZX17]. Siegmund [GKA19]. Sieve [MS10]. Signal [KG14, Maz17, MMB17]. Signal-to-Noise [KG14]. Signalling [ZLZW16]. signals [MMB17, SXYF17]. signature [CBB^{+}17a]. Signatures [Ery14]. Signed [WZ10, Hua17, MMV18]. significance [GE17]. Significant [BR16b, XGTM18]. SIMEX [MWL17]. similar [PM18, ZZS17]. Similarity [YC11, AKZK17, REM19]. Simon [SC18a]. Simple [AMW15, AM13, AS14b, ART^{+}12, BAM11, CD12, DZ16, HSR10, HMH^{+}10, LW11a, Mah12, Mey13, Sha16, SVWH10, SS16, Yao13, AE18, SGKM17, SBZB19, Sen18, SAD17]. Simple-Ordered [BAM11]. Simpler [BGM13]. Simplex [MAM11, DZT19, Lan17]. Simplified [KGXR16, GRM17, YSH19]. simply [Agh18]. Simulate [Far16]. Simulated [ART^{+}12, CN11, LSS13, Lu13, Nic10, SV13, LCD18]. Simulating [KIM16a, Sko19, ML19b, TN17]. Simulation [AM14, Alm10, AS14a, ASM15, AR10, BSZ16, Bia14, Bia17b, BR16b, Bus15, DARMM^{+}17, DDD16, DMMA15, DD16, FS12, GD15, GHRM11, GB13, HT14, HC17b, HAAQ11, JA11, JS14, KSS13, KMS15, LNL11, MSK10, McI18, MMF14, NAK13, NP17, Özk14, PLW^{+}14, RM16c, Rut15, SFG15, SEG15, Tas17, TA13, VG15, VL13, VCDB14, WN15b, XLC14, yYjHwK14, ZCR10, BOOTD17, BT17, BF17, CMC^{+}17, ES17a, GS17, GC17, GE17, HA18, KÖ19, LWT17, MM17, MN17b, NFM19, Ott17, OTZA17, PS19a, PTVP19, Ros18, SKZD18, SWM17, Ste19, WK17, Whi18, XLX^{+}19, YSC19, YDA19, ZSÁ17]. Simulation-Based [yYjHwK14, DARMM^{+}17, DDD16, NP17, Tas17, GS17]. Simulations [AL15, CDkP14, DWKL19, HS12, Pic14, dWD16, HC17b, KMV19, PR19b, dALA17]. Simultaneous [AK15, AB13a, BDKM11, BR18, BS16b, Ger16, GGM17, KBJ14, Mar13a, MMB17, SD16, TNS12, BLB19, GAK19]. Simultaneously [DK13]. sine [CBH19]. Single [ATV^{+}14, Hua16, Kim16b, Kur16, LWZH10, LZ10, LWH^{+}15, LS14, Maz17, RSC16, SOSB14a, SOSB14b, Yan13, YLG16, ZYHZ16, AC18, BS17a, DRAM18, FN17, LY17, RA17, VA17, YY17c, YY17d, YZW19]. Single-Case [SOSB14a, SOSB14b]. Single-Index [LZ10, Hua16, Kim16b, Kur16, LWZH10, YLG16, ZYHZ16, DRAM18, LY17, YY17c, YY17d, YZW19]. single-item [FN17]. single-server [AC18]. Single-step [Maz17]. Singular [EG16, MNZ13, MM18b, Vil17]. sinusoidal [MMB17]. SIR [QEK17]. Site [KGK14, KJK17]. Six [ZWJ10, BSH17, CU19]. six-Treatment [CU19]. Size [AJM10, AJ13, ANA14, AFS14, Bus15, CH10, CYW13, CM13, De 11, HJST15, HLZK15, HTN14, LSS13, Lon14, MLZ15, MMP^{+}13, ND10, PK13b, RSW12, TPS15, TP17, Wil15, WWJ14, YLK11, ZP14, DB18, FZ17, KKYW17, LK18a, SSA18, SC18b, WSCW18, YKLL17]. Sized [Spa12]. SiZer [ZM12]. Sizes [BBF13, KRAA15, Lee10b, Önd16, DÖ17, HNM^{+}18, MH17b, RRA^{+}19b, SY17b, VAR19]. Skeleton [JS12]. Skew [BG12, HD13b, dBMdSF16, RNJ16, TK15a, WBG16b, Wu14, BDL17, CS19, CL18, HF18, HK18b, KPV17, Kor17, MMB19, PA18b, SNT17, SBNG18, XZ17, ZL17]. Skew- [RNJ16, Wu14, SBNG18]. Skew-Normal [BG12, HD13b, WBG16b, CL18, HK18b, MMB19, XZ17]. skew-normal/independent [CL18]. skew-symmetric [HF18]. Skewed [BK10, KO16, KRAA15, YCH16, DD17, Kao17, MR18, NC19, SA17a, TN17, UMN17, WDC18]. Skewness [AR10, Hol10, OWY13, PB17b, SL17]. Skip [ABAJ14, BM17b]. Skip-Lot [ABAJ14, BM17b]. SkSP [ABAJ14]. SkSP-2 [ABAJ14]. Slack [Mar15]. slash [Kor17]. sliced [Hil17, JS19b]. slightly [Xue17]. Slope [CZL16]. Slopes [HB16]. SLS [AU17]. Small [AJM10, BLBK16, BKB^{+}15, CS12, CSG15, CNPW15, ES15, Ger16, GTPS14, HN15, MLZ15, Ouy13, ÖW12, OWY13, PZ18, PC13, Rah17, SFS12, TRG13, Vol14, XLC14, ABH17, MH17b, SJC18, Sta17, Zha17b, Zim17]. Small-Sample [CNPW15, Ouy13, ÖW12, OWY13, Vol14, Sta17, Zim17]. Smallest [WC16]. Smiles [MLC16]. Smirnov [DTZ10, Zen17, ZLGZ15, vZ11]. SMO [ZM13]. Smoking [CZL16]. Smooth [AN11, GL16, TX15, TY16, LN17]. Smooth-Threshold [TX15]. Smoothed [LLZ16, Tas18, CHJL17]. smoother [MC19]. Smoothing [Dut14, Dut15, LWZH10, Pal17, CTGR17, Ros19, WX19, XZ19]. smoothness [CTGR17]. SMOTE [XZ16]. Social [WA14, SRP^{+}17]. Soft [WTLWL16]. software [Bar17b, DSD19, LW17a, ZVHC17]. solid [CTXW17]. Solution [CLLH16, PK16, Rai16, SOSB14b, XY16, Yao13]. Solutions [SOSB14a, JW18, PAS19, SJT19, WLF19]. solve [Asa17]. Solvency [Pit14]. Solving [CIB10, TL15, ZM13, BHF17, MM17, Wal17]. Some [ARK17, AA15, AN16, AL16, AM13, Asa17, Asl14a, Bah18, BM14, BCV15, BK10, BLBK16, DN18, EK12, GL10, HB16, KFO14, KMS15, LDS10, LMS13, MSK10, NAK13, ÖW12, OWY13, PUN16, Pic14, Rai16, RRR19, SBZB19, SOBH10, SZSA17, Sim11, TKX13, Tri18a, AKZK17, CU19, HC17b, HKNC17, KE17, KR17, MP17a, MP19a, Nou17a, PN17]. Sortino [De 14]. Space [GH15, CC19c, SFC19]. Spacings [MRKI16, LL17a, Rif19]. Span [yYjHwK14]. spanning [LM17]. Spanos [Hod13]. Spare [LH14]. Sparse [HBC17, KK10a, LY16, Mat19, PK16, VVHP^{+}16, Wei19, Zha15c, BB17e, Wal17]. Sparsity [LZ18]. Spatial [ARW16, BKJ10, CCUR12, GB13, KK14b, LM16b, PLZ15, SKZA16, TK15a, Tor15, Hir17, KL18, MWSB18, Ott17, Rah17, RK17, SWM17]. Spatial-Temporal [GB13]. Spatially [Spa12, YGL12]. Spatio [AC15a, SP14, DWKL19]. Spatio-Temporal [SP14, AC15a, DWKL19]. Spatiotemporal [BB11, HJST15, HLZK15]. Spearman [PH17]. Special [FP14, VB16]. Specific [She12, vZS12]. Specification [GD15, LW14, WBG16a]. Specified [Aus10, SR12, Xia15, JZ18, SS17a].

Speckman [XY12]. Spectral [BGP14, GP16]. Spectrometry [MDD13]. Spectrum [Far16, GE13, MNZ13, Vil17]. sphericity [SS17c]. Spline [HZP15, LWZH10, LZ10, AMA15, AY18b, MRF17, WX19, XZ19, YZW19]. spline-backfitted [WX19]. splines [Din18]. Split [CRS13, FL13, IMS14, TC14]. Split-Plot [CRS13, FL13, IMS14]. Splitting [MSM12, RDV12]. sports [Usa17]. Spread [AL16]. Spring [AC15a]. Spurious [JZ18, JZZZ17]. Square [CYW13, Hwa17, LK14, Bar17a, CSA17, ZF16]. Squared [GL11, Han19, MO19]. Squares [ASSH10, AKM11, Ben14, IMS14, vZS12, Bar17a, CU19, CTGR17, DA18, FKA17, RL17, WY19]. SSMGRDS [SKLC18]. SSR [HK10]. Stabilizing [AOJ12, FP16]. Stable [GEVA10, KM12, Van15, AMM18, BRF17, SJR19, TRM18]. Staden [BDL17]. Stage [AZCS10, ABAJ14, CJV13, MMP^{+}13, PLW^{+}14, RMS10, AM19, DCM16, ESAD17, KA19a, LL17a, Lui19, PLLF16, QQYZ17, QG19, SC18a, SG15b, WPLL15, WZT17, Wu18b, ZM17]. Standard [Asl14b, cCyYnH13, DS13, Lu13, TW11, WSQ15, dFR13, dR11, LCD18, PR19b, SJR19]. standardization [GMR^{+}18]. Standardized [MLC16]. Standby [LKHH11, FTE17, WCSL17]. STAR [LN17]. Starting [KA19b]. State [GH15, HAAQ11, MMF14, SV13, AAAJ18, Asl19, CTXW17, CC19c, KSHD19, KAAJ19, RM16a, SFC19, YLA17, ZW17, ZYS19, ZLGS18]. State-Space [GH15]. states [Mar19]. Static [WSQ15]. Stationarity [Che15, dLHT19]. Stationary [Ban17, CDL12, OK15, PFST13, WY15b, OTZA17, SKJ17, Yeh17, ZW17]. Statistic [AOK13, BG15, HW14, IRHM13, Mar15, US10, AKJ17, LDK17, NL17, Orz17, ÖEG17, RN17, SFC19, Zen17]. Statistical [ADLS10, AS12c, AFV16, ARAM15, BSLF19, BG15, Chi15, CKPR13, EDAAAM17, FK14, FLB15, GBTV16, HK13, HCW17, HW17b, HD17, KTA17, KMSK15, LM16b, LXS19, LKYH17, NWR17, NGT14, PDL19, PSO13, RH11, SNO16, TKX13, WP13, WGS10, WL15b, WTLWL16, WCC16, WSZ17, YKSC16, ZCR10, ZZGL15, BG17, Bah18, CH17, ELS17, FN17, GE17, HK18b, KPG^{+}17, KM19b, LK18b, LK19, MKFP18, MSEN18, PMNF18, RTKB18, Sei19, SXYF17, WGX18, ZVHC17, ZM17]. Statistics [AHRM12, AS12a, BNKM15, EK16, GCR12, GN15, KK10b, LOL16, Mey13, NB15, ONS16, Sha15, Sim11, Spa12, SP14, Su11b, VCDB14, ARY18, AMM17, AMM18, BS17a, DAY19, EDE18, KE17, LLB19, MMV18, SY17a, Tri18a, UA17, Wu17a, YSH19]. Status [Kim14, LW11b, LLX16]. Steady [SV13]. Stein [TP15]. Step [AHAH15, Cha12, CM13, IMS14, MP19b, PB10, PS14, Sha16, SA16, SS10, SS16, TNS12, XTG14, BA19, FCC18, FFNM17, Maz17, SGKM17, YW17a]. Step-Down [MP19b, TNS12]. Step-Partially [AHAH15]. Step-Stress [PB10, PS14, Sha16, XTG14, BA19, SGKM17]. Steps [AAP16, PB10]. Stepwise [BS16b, Vil17]. Stochastic [AR12, EZ15, Fen16, FLB15, FK17, GR10, Har12, HAB16, KS16, LNL11, LY13, MTL10, MG12, OG14, Pic14, Ros18, SCH14, WY14, XX16, YZ11, CJLV18, EAG18, GRB17, HWZ18, KH19b, LH17, LNT17, LWT17, Pic19, QEK17, Xue17, YCS17, Yil17, Yil19, ZGE17]. stochastically [SKLK19]. Stock [BZX11, SN13a, SDZL15, SI13, XZZ15, ZB16, HYZ19, KH19a, LPC17, SNT17]. Stocks [CY16]. Storage [FSJ12, yYjHwK14, RTKB18, ZSZ17]. Strata [KRAA15, Kao18]. Strategies [De 11, STHS16]. Strategy [AAP16, CY16, HC17a]. Stratification [Koz14, dMBSFB17, NH18]. Stratified [GKA12, HAT15, HAB16, KGA11, KGA13, KK16, LL14, Lui16, SLLW11, SG17]. Streams [YYY16, WWWS19]. Strength [BGU15, Bar11, GAMA15, AGM19, BK18, BD17b, Ery18, HS17, PA18a, RSBZ18, SAB19]. Stress [Bar11, CW12a, GAMA15, GTFX14, PB10, PBS11, PS14, SA16, SS16, AGM19, BK18, BA19, BD17b, Ery18, HS17, NKKC17, PA18a, Rah17, RSBZ18, SAB19, SGKM17, Sha16, WGX18, XTG14, YCSP19]. Stress-accelerated [SS16]. Stress-Strength [Bar11, GAMA15, AGM19, BK18, BD17b, Ery18, HS17, PA18a, RSBZ18, SAB19]. Stringency [Ul 17]. Stringency-based [Ul 17]. strong [XY17a, XYL19]. Strongly [FZDF14, Góm13]. Structural [BB11, LTQ15, RF10, RR12, CB17, EB18, JZZZ17, JZ18, KPG^{+}17, Nam17, TA17]. Structure [ART^{+}12, GHY14, MM18b, SWL14, FLW19, GNG17, HZHK17, Mai17, Wei19, Wes17, XYZ17]. Structures [AAK13, Ery14, SKZA16, SN13b, Car19, DÖ17, GRSN19, ROC^{+}18, SS17c]. Stuart [Rut15]. Student [AN12]. Studentized [SS10]. Students [RM16c]. Studies [AFL^{+}12, BL15, CCC10, PLW^{+}14, PLLF16, Rot12, SCpL12, lTHT13, WMK14, WWJ14, ARK17, Fan18, KÖ19, Mor19, TK17, WMK19, YWY17]. Study [Alm10, ASM15, BLBK16, BNKM15, BGM13, Bia14, Bus15, Che14, CCLT14, CZL16, DMMA15, DD16, FS12, FGL10, FKA17, GHRM11, GB13, HT14, HHS16, HAAQ11, JS14, KU14, KMS15, Kim16b, KP10, MSK10, MMF14, MSB11, MDD13, NAK13, Önd16, Özk14, PR19a, PR19b, RBX^{+}15, RF10, RM16c, Rut15, SL14, SFG15, SDZL15, TKX13, TA13, VPF14, VKB10, Xu10, XLC14, XS11, YR11, ZCR10, ZWJ10, Zha11b, ZM13, AA17, AMKS19, AY18b, BBA17, BT17, BA19, Bia17b, BDK17, CMC^{+}17, CXC18, EAG18, ES17a, Els19, GS17, GC17, HK18b, JG17a, KH19a, KNLK19, NFM19, PTVP19, SKZD18, SA17a, SJR19, WK17, WMH17, Whi18, YL17, YSC19, YDA19, ZSÁ17, Zen17]. Studying [MM16]. sub [HYZ19, KX17]. sub-fractional [KX17]. sub-periods [HYZ19]. SubBag [ZZZ10]. subcomponents [Ery17]. Subject [LKHH11, ZW14, ZHL14, Lin17]. Subjects [RAN15]. Subpopulation [Afi14, Za̧d17]. Subsample [GL16]. Subsampling [RSC16]. subset [HZNN17]. Subsets [WZG15]. Subsoil [ARW16]. Subspace [GAPSÁ10, Wan10b]. Substitution [CY16]. Substudy [RZZ10]. Subsystems [ZW14]. Successive [Sin14, She17a, SP17b]. Sufficient [BFB17, HB10, LY16, Wan10b, SF19]. Suitable [Kom15, FA17]. Sum [AMP10, HWB13, Lee10a, HKNC17, RL17]. Summary [TZD^{+}13]. sums [LL17a, MP17a]. sup [KEB17]. sup-entropy [KEB17]. superimposed [MMB17]. Supermodular [PR15b]. Supersaturated [AKV16a, KMS11, KP12, CKP17, DKL18, DK19, QQYZ17]. supervised [KL17b]. supplementary [SPH19, ZSGW17]. Support [AAT15, BKJ10, BJ14a, SW11a, WP13, CKKP17, DRAM18]. Sure [DK19, Das17]. Surface [AKS13b, CBDD16, HB16, KCC16, OBD16, MSEN18, MDX17]. Surfaces [RCR14]. surrogate [BB17c]. Surveillance [HLZK15]. Survey [AZCS10, GP15, MS14, WTLWL16, ZGC16]. Surveys [HAZSK16, KGA13, Lu14, BS17c, PSHP17]. Survival [ALH14, CCM13, EE10, HM12a, JMCT16, KSJ^{+}12, Li14, MDD13, She15b, SF12, VL16, YR11, ZTT10, AEQEH17, BS18, Che17, dGEB17, FNA17, GB17a, HK18a, HPG17, HC17a, JZ17, KKH17, KNLK19, MERN18, PR19a, PR19b, She17a, SKLK19, SC18b, WDC18, YW17a, YH19]. Survivors [ZQL14, ZKL^{+}19]. survsim [MN17b]. Suspended [ZW14]. SVM [CMB11]. SVR [ZM13]. swarm [CX17]. Swedish [LMS13]. Switching [LKHH11, Pap14, TK16, Xu10, YZ11, BJU18, GB17b, JLGD17, LKYH17, ÖSY17, ZW17]. Symmetric [CGST17, EK16, HM12c, MLC16, War17, Yar12, HF18, LNT17, TRM18]. Symmetrical [LZ16]. Symmetrizing [CSS13]. Symmetry [WANB13, ANB19, RRR19]. symplectic [MKS18]. Symposium [CD14]. Syndrome [BSV13, HCW17]. syndromic [SRP^{+}17]. Synergistic [MP13b]. Synthetic [Lee12, LK14, LK15a, PS10, SWGX18, YKLR14, YKSC16, Haq19, LK17a, LK19]. System [ARKB13, AR12, BGU15, EK10, Ery10, GAMA15, KFO14, KYP11, KS16, LKHH11, Pit14, PRA14, QJ14, RM16a, WZCJ16, YPH14, yYjHwK14, ZX10, BM17b, BJU18, CSY17, EY17, FS17, GLGG17, LKYH17, OTZA17, PA18a, RG18, SSOTB19, WCSL17, ZYS19]. Systematic [KSG15, KSG18]. Systems [Ery14, Ery15, JWW14, SG15a, TC14, ZC11, ZW14, AK19, BD17a, Ery17, Ery19, FN17, FTE17].

T [Sei19, SL15]. Table [CDGML12, ZF16]. tables [Hua17]. Taguchi [Chi15]. Tail [KPV17, MWK16, Ras13, SN13a, Kao17]. tail-weighted [Kao17]. Tailed [JTY10, CCRX18, DBM17, JZZZ17, KO16, REM19]. Tails [SFS12]. Taiwan [ABH17, WLWT16]. Take [GBO14]. tanks [RTKB18]. TAR [NZL13]. target [Wes17]. Taylor [SCH14]. Technique [CBDD16, TP15, ZWZ^{+}14a]. Techniques [AS14a, FL13, Sha15]. Technology [GR10]. temperature [LL19]. Tempered [BRF17, KM12]. Temporal [GB13, MAC11b, SP14, AC15a, DWKL19, TS17, TS18]. ten [MGL17]. tensor [EGO^{+}17]. Term [BSV13, She15b, Zha11a, ZQL14, SSY18]. terminal [SL17]. Terms [WL15b, CS19, TGSA18, UMR17]. Terraces [AGJ15]. Terry [Fed12]. Tessellations [VHV^{+}16]. Test [AD12, AHRM12, AGJ15, AZ10, BDKM11, BR16a, BM10, BR13, CZ19, CMM12, CY12, CW14, DTZ10, ES15, GR10, GEVA10, HW14, HWB13, Hua16, Ima11, JA15, JY10, KS15, KHH12, KM13, KK10b, LO16, LP10, LLJ^{+}12, LL14, MLZ15, Mah15, MMP16, Mar12, MRKI16, OY16, ONS16, PS14, Par11, Par15, PLZ15, PAM15, PSW16, RLTJ16, Rut15, RB16, RH11, SSS14, Sol16, SOSB14a, SOSB14b, ST13, SI13, SS16, Tan16, VAGE15, WANB13, WPLL15, XD15, XY12, XYCY15, fYyP14, ZNA16, ZW15a, Zgh10, ZA10, ZWCC12, ZJNG13, ZCJ12, ZLGZ15, ZF16, vZ11, AHB19, AS17, AdBPdBP19, ANB19, BPG17, BS18, Baz17, CPL17, CSA17, ES17b, FZ17, GR17, Gür19, HESK18, HM17, HW18, Hua17, JK17a, KAN17, LAG17, LSY17, LL19, LPC17, MO18, MO19, Mai17, Man17, MSK^{+}17, Nam17]. test [NWB^{+}19, NC19, ÖEG17, Par17, SBH18, SIMS19, SMMA19, Ten17, UA17, WGX18, Wu17a, XLW17, XSA^{+}17, YSH19, Zar19, ZSÁ17, ZHLW17, Zha17b]. test-based [KAN17, ZHLW17]. Testing [AJ18, AP17, AFS14, BLN12, BCGV13, BAM11, BK16, CLW12, CWSW16, CM13, DX16, ES15, EKK19, FGL10, FK15, GR10, GBB^{+}12, GG17b, GG17a, GB13, HAT15, HXX15, Hol10, HMS16, Ian16, Isl11, JA17, KL18, KG14, Kur13, Lak11, LM10, LS10, LW11a, Li11, LWH^{+}15, LXPL17, LH13, Mar13a, Naj17, Nou17c, Nou17d, PN17, PFST13, Pra15, RLF11, RAN15, SKM^{+}10, SMS17b, SPCNL16, SS17c, SW11b, SFS12, Sun17, UH17, VB16, XLPC13, XTG14, YGW16, ZQL14, ZCJ12, ZTT10, BC17, DAY19, EH18, dGEB17, GNG17, HW17a, HC17a, KA19a, KABM17, LW17b, LSTM17, Lui17b, Mah17, MRF17, MSK^{+}17, NKKC17, Orz17, SNT17, SCN17, Sen18, SU18, Sma17, WSZ17, WCC^{+}18, XD17, YE17, ZSZ17, ZX17, SMF15]. Tests [AHAH15, AL15, Alm10, Aok15, AK11, BMN12, BCV15, BMPZ13, BPS13, BRT^{+}16, BKB^{+}15, BAPA15, CH11, CW12a, CNPW15, DKV15, GTFX14, HT14, HM12c, Ima11, Ima13, JS14, KGXR16, KSGM13, LML11, Li14, LBW11, tLB14, MV10, MTCP10, MC12, Mar13a, MAR11, Mur15, NYD16, NA13, OK15, Par11, PD16, PS13, RSW12, SVWH10, SR12, SSY15, SI14, Su11a, TNS12, TA13, VPSV13, VL13, Vol14, WCL12, WANB13, WC15, WGS10, Wu10, XT11, XQW^{+}16, ZWJ10, ZYM13, AE17, Alk17, AID17, ANB19, Bah18, BC17, CSY17, CBB^{+}17b, DL17, DD17, EAG18, FS17, GA17, HWXH18, HESK18, JNMR17, MR18, NP17, NWR17, Nou17b, Nou17a, OAB17, PZ18, PFOC17, PS19b, PB17b, RRR19, RM17, SF19, Ten17, Tew17, TJ17, Ul 17, Xue17, ZSY17, ZVHC17, dLHT19, vZ18]. Text [CCC16b, LLS16]. th [DK17b, DK17c]. Their [ABB12, CMN17, CLY19, HT17, LASX19, LLB19, Rif19, WLX18]. Theorems [ABB12, GLS12]. theoretic [Mah17]. Theoretical [Wes14]. Theory [ABS^{+}18, MRKI16, PGGF14, SZFM11, Ros18]. There [PSS10, Wil17]. Thinning [WZ10]. Third [HB16, Fen17]. Third-order [HB16]. third-person [Fen17]. Three [AAK13, AKS13b, AAP16, GMG13, HB16, Kom15, LP10, LqQlX11, LC16, LX13, NB15, PLW^{+}14, RSW12, Van12, Wan12, XQW^{+}16, AK17, BOOTD17, BSLF19, EZ17, FFNM17, Kna19, QQYZ17, RRA^{+}19b]. Three-class [LC16]. Three-Decision [GMG13]. Three-dimensional [HB16]. Three-factor [XQW^{+}16]. Three-Hypothesis [LP10]. Three-Level [AAK13, AKS13b, EZ17]. Three-Parameter [Kom15, NB15, Van12]. three-part [AK17]. three-phase [BOOTD17]. Three-Stage [PLW^{+}14, QQYZ17]. Threshold [BD14a, KU14, NYD16, SWL14, TX15, vZ11, LXPL17, PGH^{+}18, Tas17, ZZ18]. threshold-ARMA [LXPL17]. Thresholding [XC15]. thresholds [MSK^{+}17]. thumb [SSYS19]. tied [Wil17]. ties [Sab18, XHD18]. Time [AS14a, ABJM17, Asl19, BZL15, BL15, CD12, CC10, CCM13, CZL16, CB14, Ery15, GBPDL16, GL14, HHHZ10, JH14, Jin11, lKW13, LM16a, LTQ15, LG11, LLZ16, MWK16, MS10, Nic10, RR13, SB12a, SMWR14, SCpL12, SHM12, SC13, SF12, TW16, TY16, Wei10, YCH16, YCSD10, YCS12, YLK11, yYjHwK14, ZHL14, Ami19a, Ami19b, BA19, BM19, BDK17, CHB17, CXC18, CH17, CCZ19, DSD19, DAY19, DOZ17, DB17, Elh17, ELS17, FCC18, Gór18, HST17, LL19, Lin17, LNUT17, MYS^{+}16, MC19, NWB^{+}19, Pic16, SS17b, SGNS17, SKJ17, TS17, Tew17, XY17b, ZLGS18]. time-changed [XY17b]. time-constraint [BA19]. Time-Inhomogeneous [Nic10]. time-series [SGNS17]. Time-to-Event [SCpL12, SF12]. Time-truncated [ABJM17]. Time-Varying [LG11, CZL16, YCH16]. times [MKFP18, MAMM17, She17a, sS19a, VAR19, XHD18]. Tippett [BGM13]. Tippett-like [BGM13]. Tobit [Alh16, FL19, AA18b]. Toeplitz [Mai17]. TOF [SJR19]. Tolerable [LH14]. Tolerance [You13, Zim17]. Tolerances [RYW16]. Tolerant [TC14]. Tool [BA14, KIM16a, SP18b]. Topp [FA17]. Total [HTN14, RSC16, WY19]. TPRs [SSY15]. trace [FK17]. Traces [JR12]. traditional [WMH17]. Traffic [JH14, AC18, CQH17]. Training [RH11]. Traits [MMM16, MS14]. Trajectories [CCD^{+}12]. Trajectory [CK14]. transaction [NMS18]. transect [Mor19]. Transfer [GAPSÁ10]. Transform [PR15a, Lue17, XY17b]. Transformation [BSZ16, CSS13, CYW13, DAI14, GK14b, sS12c, She15a, AID17, CY17, DI17, NC19, She17b]. Transformations [FP16, KH18, MH17a]. Transformed [dALA17]. Transformers [ES16]. transient [UA17]. Transition [KR12, MYS^{+}16, Nic10, YCSD10, LN17, dLHT19]. Transmuted [KKH17, TTZ14, ABFAM17, FNA17]. Transport [Bon16, CR17]. Travel [GKA12]. Treat [Aus10]. Treatment [CU19, PK13b, SH12, Sin14, SOSB14a, SOSB14b, TK16, Agh18, CBB^{+}17a, JLGD17, Llo17, Wu18b, YW17b]. Treatment-Control [SH12]. Treatments [AGJ15, CWSW16, GMG13, YWY17]. Tree [ZWZ^{+}14a, AAK^{+}17, LM17]. Trees [AAP16, Cha12, Cha13]. Trend [Alm10, AC15a, CTGR17, QDTL16, Rut15, KSG18, SGNS17]. Trending [CNPW15]. trends [Grc17]. Trial [GAK16, CBB^{+}17a, CH17, JLGD17, Lui17a]. Trials [Ani16, AK11, PK13b, PR19a, PR19b, ZCR10, ZL10, Zha11a, ZLS14, CTH17, LDK17, Llo17]. trimmed [AB18, DA18, OWY18, YXY17]. Trinomial [PSO13]. Trisomies [RLF11]. True [BB17c, ZWZC12]. Truncated [BM14, Gen17, KK14b, LDS10, She12, She14, She15b, SKJH11, US10, ZA10, Zha15b, ABJM17, Asl19, CTTT16, CFSC17, ES16, GA17, HW18, JZ19, LLZ18, She17b, She17c, TN17, WPTL17, ZLL17, ZN18]. Truncating [EG16]. truncation [BT17]. Tsallis [TB17]. Tsiatis [CBB^{+}17b]. tuber [Das17].

Tukey [LT15, TGRGCF^{+}12]. Tuned [NGT14]. Tuning [AKM11, AKV16b, EA12, Kar18]. Turing [GC17]. Turkish [ASKT19, Tas17]. TV [AAT15]. Twin [RT12]. Two [AJM10, AJ13, AMS16, ASLFLR12, AH13a, AG17, ABAJ14, AS14b, Bak16, BK16, BG16, CS19, CLW12, CLLH16, CM13, CDGML12, CJV13, Dei12, DDK16, DCM16, EB14, FCC18, GK18, GRY17, GPM18, HS12, HHMW11, HKC10, Hua14, HB16, IMS14, Ima11, Ima13, JA15, KO16, KS15, KK10b, LS16, LTPT11, Li14, Lon14, LSLW11, MMP16, MMP^{+}13, OY16, PLLF16, PUN16, Phi12, RSC16, RMS10, RAN15, SKST14, SH12, SG14, SB12b, SS10, SKM^{+}10, SG15b, SLLW11, SR12, Sta12, SSY15, TA13, VCDB14, jWzM13, WPLL15, WC15, WMK16, XYCY15, XY16, YW17a, fYyP14, ZP14, ZTT10, ZF16, AM19, AL15, AP17, AVR17, AK19, BR18, BS16b, BVC^{+}19, CBB^{+}17b, Cha18, CKP17, DGH19, EDE18, Ery17, Ery19, ESAD17, FL19, FTE17, HC19, HM17, HA17b, HC17a, Iva18, JK17b, KG17]. two [LL17a, LCD18, Lui19, MSK^{+}17, NÖ18, NWR17, ÖEG17, PM17, PM18, QG19, SSOTB19, SMMA19, SaN17, SBZB19, SK17a, SK17b, SG17, SC18a, SP17b, Tas18, Tri18b, UH17, WZT17, WHK17, YSS17, ZW17, ZVHC17, ZM17]. Two- [HB16]. Two-Class [SLLW11]. Two-Color [SH12]. two-colour [DGH19]. Two-Dimensional [HS12, CLLH16, GPM18, Iva18]. two-factor [MSK^{+}17]. Two-Group [ZP14, BVC^{+}19]. Two-Group-Forecast-and-Selection [LSLW11]. Two-Level [EB14, Sta12, jWzM13, CKP17]. two-occasion [SP17b]. Two-Parameter [SKST14, fYyP14, AG17, Bak16, DDK16, AVR17, Cha18, SaN17, SBZB19, SK17a, SK17b]. two-period [Lui19]. Two-Phase [SG14, LS16, RSC16, SG17]. Two-Regime [HKC10]. Two-Sample [KK10b, SB12b, GRY17, WMK16, AL15, SMMA19, ZVHC17]. Two-Sided [AH13a, Ima11, Ima13, KG17]. Two-Stage [ABAJ14, CJV13, MMP^{+}13, RMS10, DCM16, PLLF16, SG15b, WPLL15, AM19, ESAD17, LL17a, Lui19, QG19, SC18a, WZT17, ZM17]. two-state [ZW17]. Two-Step [CM13, IMS14, SS10, FCC18, YW17a]. Two-Way [CDGML12, Phi12, XYCY15, CS19, MMP16, XY16, ZF16, YSS17]. Type [AMS16, ABAJ14, BDKM11, BK10, BMPZ13, GK14b, GSSCR16, HWX16, IE13, JM13, JW12, KMS15, KGS14, LHB10, OWD14, PS10, Ron10, RGP17, SG14, SO10, Su11a, TL15, VB16, WL14, AGM19, ASKT19, AG16, ABJM17, AK19, Ban17, BW19, CC19b, DK17a, EKG17, Ery18, FZ17, HESK18, JK17c, LPC17, Mur15, RL17, SPH19, Van15, WC17, WY15a, WZTX16, Wu17a, WF17, WCC^{+}18, WA18, YCL17, Yil17, Yil19, ZZ17, AHAH15, AY15, AYR17, ARY18, ARAM15, BK18, BS15, BSL15, BYAJ17, BS17a, BR16a, CSY17, CMC^{+}17, CLJ13, CCL10, DDK16, DAA13, EDAAAM17, GRY17, GP17, HK13, HK17, JG17b, KY15, KSS17, KTRA17, KNJ16, KEB17, KRK15, LÁJGAF11, LHB08, LBW11, LTAJ14, LRVF10, NKKC17, Nou17c, PN17, PH17, Par17, QLAJB17, RSR17, SK17a, SKK17, SK17b, SB12b, Sha16, SSK16, SS16]. Type [VAR19, WL15b, ZS17]. Type-I [LHB10, BW19, AHAH15, AY15, AYR17, CSY17, CLJ13, CCL10, JG17b, KSS17, KEB17, LHB08, LBW11, PN17, RSR17, SK17b, SB12b, Sha16, VAR19]. Type-II [LHB10, YCL17, ARY18, BK18, BS15, BSL15, BS17a, DDK16, EDAAAM17, GRY17, GP17, HK13, HK17, KNJ16, KRK15, LHB08, NKKC17, RSR17, SK17a, SKK17, SSK16, SS16, ZS17]. Type-X [LTAJ14]. types [FTE17, Yao18].

Uhlenbeck [BRF17, JT19, KM12, ZW17]. ultraspherical [MH17a]. Umbrella [CY12, XD15]. unary [JG17a]. Unbalanced [LW11a, Phi12, SLLW11, XQW^{+}16, YHL12, Zha15a, vdH10, CKP17, GGM17, XW17]. Unbiased [WY14, Wu16, YS12, AMM17, AMM18, HS17, WA18, Yil17]. uncertain [DL17]. Uncertainties [KMSK15, WLX18]. Uncertainty [DD16, LJ17]. Unconstrained [Ell11]. uncorrelated [ML19b]. Undecimated [PR15a]. Underground [RTKB18]. Underlying [DH16]. Understanding [MM15]. Unequal [BBF13, KS15, SaJMF16, WWJ14, Zha15a, AP17, HNM^{+}18, JK17b]. unidimensional [BSK19]. Unified [Liu13, Ate17]. Uniform [MAM11, TM14, dR11]. Uniform-Exponential [dR11]. Uniformity [CBDD16, Mah17, Nou17b]. Uniformly [XY17a, XYL19]. Unilateral [LW14]. union [YAT19]. Unit [AFS14, CNPW15, HMS16, Lak11, LML11, LO16, LS10, OK15, SSS14, Sol16, Wu10, ZYM13, Gür19, LAG17, PRA14, Sen18]. Units [AJ13, GPM18, IMS18]. Univariate [Che11, LW11b, LRVF10, NZL13, ÖW12, CK18, JZ17, Koc17, WSX18, ZX17, vZ18]. Unknown [AL16, DS13, Hod13, Ima13, NB15, PFST13, SMF15, SKHL15, Wal11, WW10, WANB13, Baz17, GVTGF^{+}17, SWGX18]. unlabeled [HT17]. unmeasured [BLC17]. Unobserved [Lak11, LAG17, MV10]. Unobserved-ARCH [Lak11, LAG17]. unpaired [UH17]. Unrelated [Asi14, RA14]. unreliable [SSOTB19]. Unrepairable [ZW14]. Unreplicated [AKS13a, MMP16]. Untested [Lan10]. Unverified [SSY15]. unweighted [MGB17, War17, ZSY17]. Use [Cal19, Gut15, LSY16, SKZA16, YGL12, BLC17]. Used [MMP^{+}13, PD16]. useful [ARM^{+}19]. Usefulness [Wil12]. Using [AGJ15, AM14, AL15, AL16, ATC17, Asl14b, AS14b, BAA16, BKJ10, BJ14a, BSV13, Bee15, BZW13, BAPA15, CSG15, CBDD16, CP16, CXW16, CY16, CW15, CK16, DPCJ16, DZC10, DX16, Dut14, EE10, EZ17, FSJ12, GGNB10, GL14, HEF13, HNV12, HSR10, HAZSK16, IRHM13, JWW14, JZ14, JZ19, KLLM12a, KLLM12b, KCC16, KK14b, Kim14, KOC16, LS16, Li14, LQW^{+}16, Lon14, LG11, MA12, Man17, MRKI16, Mor19, MS14, NZL13, NA12, OBD16, PKK12, Pol17, Pol18, QWM^{+}16, RW17, Roy12, SJP12, SS14, SS12a, SG14, SDZL15, SSA15, Spa12, SP14, SW13, TP15, TK16, UHP11, US10, VG15, Van15, Wan14, YKY10, Yan13, YS12, YAA15, YZW19, dJvBS16, AARQ19, ASKT19, AY18a, ARM^{+}19, ASS18, AM17c, AAAJ18, AKJ17, AAMJ17, AMJ17, ABJM17, Asl19, AAK^{+}17, BSH17, Bia17b, Bia17a]. using [CCRX18, CDE19, CL18, CYY18, CLT17, DGH19, DA18, DGMC18, FKG19, GKA19, Gór18, GKS17, HK18a, HWD18, HW17b, IOTBO17, IMS18, KAN17, KuAH19, KAS^{+}19, KKYW17, LW17a, LSD17, Mah17, Maz17, MN18, MSS17, ML17, MDD13, MDX17, NK17, Orz17, OTZA17, PS18a, Pal17, PS18b, RM17, SAB19, SV13, SB17, SS17a, SK17a, SG17, SPH19, SMS17a, SXYF17, TC14, Tri18b, UH17, Usa17, Whi18, WB18, WSZ17, XGTM18, XD17, YMS17, YXY17, YCS17, YKLL17, YBW^{+}12, YLM17, ZLL17, ZLC^{+}17, ZSGW17, dSBHM19]. UST [RTKB18, RTKB18]. Utilized [Wes14].

vacation [SSOTB19]. vacations [Ban17, WCSL17]. Validation [HdL10, RZZ10, Ste12, WMK16, Dut16, HWD18, JSKV18, KL17a, ZGR^{+}17]. validity [Nou17d]. Value [AL15, BDKM11, CH10, CNA16, Che11, EG16, GHRM11, GBP16, LYY14, Pit14, PR15b, SW11a, YCS17, vZ11, HE17, WH17, XL18, ZLTX19]. Value-At-Risk [LYY14, Che11]. valued [BB17d]. Values [CN11, KHH12, MMM13, SWL14, Sin14, WGS10, WY15b, AE17, Kiz19, Koc17, Mar15, RSS17, Wil17, ZX17]. VaR [Bee17, LSY16]. Variability [SKHL15, WMK14, Yan13, ZZW11, ZZ13, SM17, WMK19, ZHLW17]. Variable [ANA14, Cha13, CCC16a, CYW13, CKC^{+}14, DZ16, DGMC18, EO16, FMMV16, FX12, FZDF14, GK14b, GSSCR16, HWW18, JS19b, KMS11, KGS14, KP10, Kur16, LK15a, LK15b, LQW^{+}16, LHB08, LHB10, LYG17, NGT14, QWM^{+}16, SFG15, SCpL12, SLB11, TX15, US10, Wu14, WZTX16, XZF19, YXC10, ZS17, ZWZ14b, BAA16, FCC18, GE17, GKS17, HS19a, HD17, HZ19, JLGD17, KAAJ19, KKYW17, LSD17, LK18a, LY17, MP17b, MSS17, PLT18, QQYZ17, SAAJ19, SSA18, TTC19, YCL17, YY17c, YY17d, YKLL17, YZW19, ZLH^{+}19]. Variable-Selection-Based [ZWZ14b]. Variables [AS14b, BU15, BB14, Cha13, CM13, DS13, FG15, Gut15, JY10, Kri16, LW14, Mar15, RAN15, RB16, SG14, Wei11, hWjJsH13, YYY16, YGL12, ZYHZ16, dJvBS16, BF17, BS17b, BS17c, FL19, HS19a, Kna19, Lue17, SG17, SPH19, SSYS19, Sun16, TN17, WLX18, Wes17, XC18, YGW16]. Variance [Ade12, BH16, Che15, CMB11, CW14, DN16, DK11, FP16, IOTBO17, KP17, LW11a, LA13, MCM16, OK15, OTZA17, OD16, PR14, PR15a, Sab18, SG14, SSA15, VL10, WLX18, YHL12, AP18, HS17, Hwa17, Kao17, LLH17, Lan17, Lue18, NAJ18, Pal17, QDT18, ZLHW19]. Variance-based [WLX18]. Variance-stabilizing [FP16]. Variances [CSK16, FP16, HHMW11, IMS14, OWY13, SaJMF16, WWJ14, Zha15a, CPL17, GG17a, JK17b, JNMR17, Naj17, Naj19]. variant [Ras17, EFA17]. Variate [KLLM12a, KLLM12b]. variates [Dem17b]. Variation [Ang13, BK11, CSS13, KM12, AG19, AAAJ18, CLT17, KKYW17, KX17, SaN17, YLA17, YKLL17, ZSÁ17, ZZS17]. Variational [DW16, OWD14, Pal17, TTC19]. Variations [AZ10, JR12]. variety [STLH17]. VARIMA [GD15]. variogram [JK17c]. Various [BSV13, LSS13, TK16, ZK18]. Varying [CCM13, Hua14, LZ10, LM10, LG11, Wei11, WWL15, ZM12, ZLZW16, BCN17, CZL16, KL18, XC18, YCH16, YY17c, YY17d, ZLW17].

Varying-Coefficient [Hua14, LZ10, LM10, Wei11, YY17c, YY17d]. Vasicek [SEG12]. Vector [AAT15, BKJ10, BJ14a, BD14b, Mak11, SFLC14, SW11a, SS19b, TNS12, AM17b, CKKP17, DRAM18, GAK19, MN18, SSA18, YSH19, Zha17b]. Vectorial [SEG15]. Vectors [HTN14, Ima11, Ima13, KS15, Kur13, MAM11, SKM^{+}10, Baz17]. verification [SSOTB19]. Version [MTL10, KM19b, Mac17]. versions [KSG18]. Versus [GP15, XGTM18, dWD16]. VI [PR19a, PR19b]. Via [AK14, AH13b, LDS10, Pap14, AID17, BR16a, BR18, CDkP14, Cha13, CKP17, CBH19, EZ15, EO16, FK17, GGZG17, GP16, GP17, HE17, HWW18, JA11, KK10a, KH17, Kor17, MZ17, NFM19, NH18, SMY12, VC13, VHV^{+}16, Wal14, WSW16, WH17, WZTX16, WY17, XZF19, YZZ17a, hZxNL14]. View [Ema13, OR11, BRF17]. VII [PR19a]. VII* [PR19b]. visits [ABH17]. Visualization [GZ10]. visualizing [LJ17]. VMA [Far16]. volatilities [SSY18]. Volatility [EZ15, LNL11, MLC16, RR12, SN13a, SCH14, YZ11, ZB16, CLT17, DC17, GSP^{+}17, KH19b, LH17, LPC17, SS19b, YCS17, ZGE17]. Volume [SG15a]. Voronoi [VHV^{+}16]. vote [Lue18]. voting [Ros18]. vs [HdL10]. VSI [YKY10]. VSS [Sei19].

W [Wu17a]. Waerden [Lue18]. Wald [WF17]. Wald-type [WF17]. Walk [MRM10, GKA19]. Walks [Yar12]. Wallis [BG15]. warm [WCSL17]. Warner [HSR10]. warning [Lue17]. Wasserman [Li10b]. water [AAK^{+}17]. Watson [FG15]. Wave [Mac11a]. Wavelet [ABH17, GE13, HN15, LS10, LG11, NCF15, RV15, WW10, KL17b]. Wavelet-based [ABH17, KL17b]. Wavelets [BBK17, MA16, VVHP^{+}16]. Way [CDGML12, DS14, JL12, Liu13, Phi12, RSW12, XYCY15, Zha15a, AARQ19, CS19, LLH17, MMP16, Wil17, XY16, YSS17, ZF16]. WCF [CTXW17]. Weak [Li10b, PS13, SJT19, Ste19]. weakly [WLF19]. Weather [HB15]. Web [XS11]. Weibull [ARK17, ABFAM17, ABJM17, Asl19, Ate17, ARAM15, BAND16, BSLF19, BJU18, Bar17a, BAN15, BW19, CC19a, FNA17, GRY17, HC19, JA11, JS19a, KSS17, KSHD19, KU14, KEB17, KBAD11, KGXR16, LCN19, LBW11, LOL16, MMD19, MKFP18, NB15, NKKC17, OCH11, PB18a, PGHA15, RAAK18, RV15, TCA^{+}16, VPSV13, WCL12, Wan14, YS12, vZS12]. Weight [FX12, FZDF14, PL13, Sta12, TZD^{+}13, KPG^{+}17, WY19]. Weight-Fused [FZDF14]. weight-model [KPG^{+}17]. Weighted [AS12b, BJ14a, Bee15, ET16, Ery15, FK14, GHY14, Gre17, Iqb11, KMF15, KRK15, Lee12, LY13, MW10, MGB17, NCF15, TT14, WY19, XLPC13, Zha15b, vZS12, AHM17, AAMJ17, BG17, CHB17, CCZZ17, FKG19, FTE17, GA17, HC18, HKP18, Kao17, Kao18, KX17, Roo17, Usa17, WW17, Yil17, ZSY17, ZLHW19, ZYS19]. Weighting [HXX15, KK16, TK16, Wes14, XHD18, GGZG17, YL17, ZLH^{+}19]. Weights [LLJ^{+}12, KPG^{+}17]. Welch [NC19]. Where [PK13b]. Which [CM16]. While [Ery10]. whiteness [DAY19]. Whitney [Tan16]. Whittle [BBK17].

Whole [VM11]. Whole-Plot [VM11]. whose [WLF19]. wide [PLLF16]. Widely [PD16]. Wiener [FSJ12, PB10, SA16, WBG16a, WBG16b]. Wilcoxon [HWB13, Hua17, Mur15, Tan16]. Wilcoxon-type [Mur15]. Wilk [HT14]. Win [WHK17]. Win-probabilities [WHK17]. Window [LGM12, Ros19]. Windows [Spa12]. Within [EK10, KFO14, BLC17, DH12, LK18b, MAT17]. within-profile [MAT17]. Without [AOK13, XYCY15, KAAJ19, XY16]. Wood [CK16]. Working [Ery10, GHY14, Gut15, LLJ^{+}12, Ban17, XYZ17].

X [ABJM17, LTAJ14]. XGLD [AFM12]. XII [CW12a, CJTL18, HK18a, KTRA17, PH17, WCC^{+}18].

year [CZL16, WLWT16]. Youden [LX13]. Young [Bia17a]. Yuen [NC19].

Zenga [GP15]. Zero [CS12, DDD16, HHS16, HLH14, Li11, LS14, MR14, VL10, WK18, XT11, ZWJ10, AK17, KR17, NMS18, RA17]. zero-coupon [NMS18]. Zero-Inflated [CS12, HLH14, Li11, LS14, MR14, ZWJ10, DDD16, HHS16, WK18, KR17, RA17]. Zero-Inflation [XT11]. zero-one [AK17]. zeros [XLW17]. Zirkler [HESK18]. Zone [ZSZW18].

References

Ahmed:2012:CGN

[AA12] Rashid Ahmed and Munir Akhtar. Computer-generated neighbor designs. Communications in Statistics: Simulation and Computation, 41(10):1834–1839, 2012. CODEN CSSCDB. ISSN 0361-0918.

Ahmed:2015:SCB

[AA15] Rashid Ahmed and Munir Akhtar. Some classes of binary block neighbor designs. Communications in Statistics: Simulation and Computation, 44(5):1253–1256, 2015. CODEN CSSCDB. ISSN 0361-0918.

Alqallaf:2016:RIG

[AA16] Fatemah Alqallaf and Claudio Agostinelli. Robust inference in generalized linear models. Communications in Statistics: Simulation and Computation, 45(9):3053–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ashuri:2017:EMC

[AA17] Atefeh Ashuri and Amirhossein Amiri. Estimating monotonic change in the rate and dependence parameters of INAR(1) process (case study: IP counts data). Communications in Statistics: Simulation and Computation, 46(7): 5023–5053, 2017. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2018:BEN

[AA18a] Rahim Alhamzawi and Haithem Taha Mohammad Ali. The Bayesian elastic net regression. Communications in Statistics: Simulation and Computation, 47(4):1168–1178, 2018. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2018:BTQ

[AA18b] Rahim Alhamzawi and Haithem Taha Mohammad Ali. Bayesian tobit quantile regression with L_{1/2} penalty. Communications in Statistics: Simulation and Computation, 47 (6):1739–1750, 2018. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2019:BBQ

[AA19] Rahim Alhamzawi and Zakariya Yahya Algamal. Bayesian bridge quantile regression. Communications in Statistics: Simulation and Computation, 48(3):944–956, 2019. CODEN CSSCDB. ISSN 0361-0918.

Arshad:2018:CCM

[AAAJ18] Asma Arshad, Muhammad Azam, Muhammad Aslam, and Chi-Hyuck Jun. A control chart for monitoring process variation using multiple dependent state sampling. Communications in Statistics: Simulation and Computation, 47(8): 2216–2233, 2018. CODEN CSSCDB. ISSN 0361-0918.

Amin:2017:IDG

[AAC17] Muhammad Amin, Muhammad Amanullah, and Gauss M. Cordeiro. Influence diagnostics in the Gamma regression model with adjusted deviance residuals. Communications in Statistics: Simulation and Computation, 46(9):6959–6973, 2017. CODEN CSSCDB. ISSN 0361-0918.

Androulakis:2013:CTL

[AAK13] E. Androulakis, P. Angelopoulos, and C. Koukouvinos. A comparison of three-level orthogonal arrays in the presence of different correlation structures in observations. Communications in Statistics: Simulation and Computation, 42(3): 552–569, 2013. CODEN CSSCDB. ISSN 0361-0918.

Androulakis:2014:MDC

[AAK14] E. Androulakis, P. Angelopoulos, and C. Koukouvinos. Model discrimination criteria on model-robust designs. Communications in Statistics: Simulation and Computation, 43(7):1575–1582, 2014. CODEN CSSCDB. ISSN 0361-0918.

Azam:2017:CDT

[AAK^{+}17] Muhammad Azam, Muhammad Aslam, Khushnoor Khan, Anwar Mughal, and Awais Inayat. Comparisons of decision tree methods using water data. Communications in Statistics: Simulation and Computation, 46(4):2924–2934, 2017. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2017:MPM

[AAMJ17] Muhammad Aslam, Ali Hussein Al-Marshadi, and Chi-Hyuck Jun. Monitoring process mean using generally weighted moving average chart for exponentially distributed characteristics. Communications in Statistics: Simulation and Computation, 46(5):3712–3722, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Azam:2016:TSS

[AAP16] Muhammad Azam, Muhammad Aslam, and Karl Peter Pfeiffer. Three steps strategy to search for optimum classification trees. Communications in Statistics: Simulation and Computation, 45(2):548–565, 2016. CODEN CSSCDB. ISSN 0361-0918.

Abbas:2019:BWM

[AARQ19] Tahir Abbas, Shabbir Ahmad, Muhammad Riaz, and Zhengming Qian. A Bayesian way of monitoring the linear profiles using CUSUM control charts. Communications in Statistics: Simulation and Computation, 48(1):126–149, 2019. CODEN CSSCDB. ISSN 0361-0918.

Albuquerque:2015:SVC

[AAT15] Pedro Albuquerque, Solange Alfinito, and Claudio V. Torres. Support vector clustering for customer segmentation on mobile TV service. Communications in Statistics: Simulation and Computation, 44(6):1453–1464, 2015. CODEN CSSCDB. ISSN 0361-0918.

Asar:2017:RRE

[AAW17] Yasin Asar, Mohammad Arashi, and Jibo Wu. Restricted ridge estimator in the logistic regression model. Communications in Statistics: Simulation and Computation, 46(8): 6538–6544, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ahmadi:2013:SCR

[AB13a] Jafar Ahmadi and N. Balakrishnan. Simultaneous closeness of k-records. Communications in Statistics: Simulation and Computation, 42(1):107–120, 2013. CODEN CSSCDB. ISSN 0361-0918.

Allal:2013:PEF

[AB13b] Jelloul Allal and Mohammed Benmoumen. Parameter estimation for first-order random coefficient autoregressive (RCA) models based on Kalman filter. Communications in Statistics: Simulation and Computation, 42(8):1750–1762, 2013. CODEN CSSCDB. ISSN 0361-0918.

Adam:2018:REB

[AB18] Lukás Adam and Premysl Bejda. Robust estimators based on generalization of trimmed mean. Communications in Statistics: Simulation and Computation, 47(7):2139–2151, 2018. CODEN CSSCDB. ISSN 0361-0918.

Abbasov:2019:RHA

[AB19] Majid E. Abbasov and Vladimir M. Bure. Randomized heuristic algorithms for orthogonal projection of a point onto a set. Communications in Statistics: Simulation and Computation, 48(10):2866–2876, 2019. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2014:SLS

[ABAJ14] Muhammad Aslam, Saminathan Balamurali, Muhammad Azam, and Chi-Hyuck Jun. Skip-lot sampling plan of type SkSP-2 with two-stage group acceptance sampling plan as reference plan. Communications in Statistics: Simulation and Computation, 43(4):777–789, 2014. CODEN CSSCDB. ISSN 0361-0918.

Afanasyeva:2012:LTS

[ABB12] Larisa Afanasyeva, Elena Bashtova, and Ekaterina Bulinskaya. Limit theorems for semi-Markov queues and their applications. Communications in Statistics: Simulation and Computation, 41(6):688–709, 2012. CODEN CSSCDB. ISSN 0361-0918.

AbdEl-Bar:2018:EGM

[Abd18] Ahmed M. T. Abd El-Bar. An extended Gompertz– Makeham distribution with application to lifetime data. Communications in Statistics: Simulation and Computation, 47(8):2454–2475, 2018. CODEN CSSCDB. ISSN 0361-0918. See comments [CL22].

Al-Babtain:2017:KTE

[ABFAM17] Abdulhakim Al-Babtain, Ahmed A. Fattah, A-Hadi N. Ahmed, and Faton Merovci. The Kumaraswamy-transmuted exponentiated modified Weibull distribution. Communications in Statistics: Simulation and Computation, 46(5): 3812–3832, 2017. CODEN CSSCDB. ISSN 0361-0918.

Angers:2017:WBB

[ABH17] Jean-François Angers, Atanu Biswas, and Jing-Shiang Hwang. Wavelet-based Bayesian nonlinear regression for the air pollution effects on clinic visits in small areas of Taiwan. Communications in Statistics: Simulation and Computation, 46(7):5504–5515, 2017. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2017:TTA

[ABJM17] Muhammad Aslam, S. Balamurali, Chi-Hyuck Jun, and Aneela Meer. Time-truncated attribute sampling plans using EWMA for Weibull and Burr type X distributions. Communications in Statistics: Simulation and Computation, 46 (6):4173–4184, 2017. CODEN CSSCDB. ISSN 0361-0918.

Alizadeh:2018:EPL

[ABS^{+}18] M. Alizadeh, S. F. Bagheri, E. Bahrami Samani, S. Ghobadi, and S. Nadarajah. Exponentiated power Lindley power series class of distributions: Theory and applications. Communications in Statistics: Simulation and Computation, 47 (9):2499–2531, 2018. CODEN CSSCDB. ISSN 0361-0918.

Arefi:2016:IEM

[ABV16] M. Arefi, G. R. Mohtashami Borzadaran, and Y. Vaghei. Interval estimations for the mean of the generalized Poisson and generalized negative binomial distributions. Communications in Statistics: Simulation and Computation, 45(6): 1838–1864, 2016. CODEN CSSCDB. ISSN 0361-0918.

Arab:2015:STT

[AC15a] Ali Arab and Jason R. Courter. Spatio-temporal trend analysis of spring arrival data for migratory birds. Communications in Statistics: Simulation and Computation, 44(10): 2535–2547, 2015. CODEN CSSCDB. ISSN 0361-0918.

Arunachalam:2015:ABR

[AC15b] Viswanathan Arunachalam and Álvaro Calvache. Approximation of the bivariate renewal function. Communications in Statistics: Simulation and Computation, 44(1):154–167, 2015. CODEN CSSCDB. ISSN 0361-0918.

Almeida:2018:NBE

[AC18] Márcio A. C. Almeida and Frederico R. B. Cruz. A note on Bayesian estimation of traffic intensity in single-server Markovian queues. Communications in Statistics: Simulation and Computation, 47(9):2577–2586, 2018. CODEN CSSCDB. ISSN 0361-0918.

Abadi:2012:GFT

[AD12] R. Hasan Abadi and M. Doostparast. Goodness-of-fit test and parameter estimation for a proportional odds model of random censorship. Communications in Statistics: Simulation and Computation, 41(8):1430–1443, 2012. CODEN CSSCDB. ISSN 0361-0918.

Amatya:2015:MRP

[AD15] Anup Amatya and Hakan Demirtas. MultiOrd: An r package for generating correlated ordinal data. Communications in Statistics: Simulation and Computation, 44(7):1683–1691, 2015. CODEN CSSCDB. ISSN 0361-0918.

Amatya:2017:PRP

[AD17] Anup Amatya and Hakan Demirtas. PoisNor: An R package for generation of multivariate data with Poisson and normal marginals. Communications in Statistics: Simulation and Computation, 46(3):2241–2253, 2017. CODEN CSSCDB. ISSN 0361-0918.

Atlekhani:2019:RAM

[AD19] Mohammad Atlekhani and Mahdi Doostparast. The repair alert models with fixed and random effects. Communications in Statistics: Simulation and Computation, 48(2):385–395, 2019. CODEN CSSCDB. ISSN 0361-0918.

Assane:2019:MCS

[AdBPdBP19] Cachimo Combo Assane, Basilio de Bragança Pereira, and Carlos Alberto de Bragança Pereira. Model choice in separate families: A comparison between the FBST and the Cox test. Communications in Statistics: Simulation and Computation, 48(9):2641–2654, 2019. CODEN CSSCDB. ISSN 0361-0918.

Adelfio:2012:CPD

[Ade12] Giada Adelfio. Change-point detection for variance piecewise constant models. Communications in Statistics: Simulation and Computation, 41(4):437–448, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Amaral:2010:MAS

[ADLS10] Getulio J. A. Amaral, Luiz H. Dore, Rosangela P. Lessa, and Borko Stosic. k-means algorithm in statistical shape analysis. Communications in Statistics: Simulation and Computation, 39(5):1016–1026, 2010. CODEN CSSCDB. ISSN 0361-0918.

Abd-Elfattah:2017:SVG

[AE17] Ehab F. Abd-Elfattah. Saddlepoint p-values for group of linear rank tests. Communications in Statistics: Simulation and Computation, 46(6):4274–4280, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Arif:2018:FOK

[AE18] Osama H. Arif and Omar Eidous. Fourth-order kernel method for simple linear degradation model. Communications in Statistics: Simulation and Computation, 47(1): 16–29, 2018. CODEN CSSCDB. ISSN 0361-0918.

Abuzaid:2017:RRM

[AEQEH17] A. H. Abuzaid, M. K. Abu El-Qumsan, and A. M. El-Habil. On the robustness of right and middle censoring schemes in parametric survival models. Communications in Statistics: Simulation and Computation, 46(3):1771–1780, 2017. CODEN CSSCDB. ISSN 0361-0918.

Afify:2014:SLS

[Afi14] W. M. Afify. Separation and lack of separation of subpopulation in the mixed distributions. Communications in Statistics: Simulation and Computation, 43(2):417–427, 2014. CODEN CSSCDB. ISSN 0361-0918.

Aliev:2012:PLC

[AFL^{+}12] Alexander Aliev, Valerii Fedorov, Sergei Leonov, Brian McHugh, and Mindy Magee. PkStaMp library for constructing optimal population designs for PK/PD studies. Communications in Statistics: Simulation and Computation, 41(6): 717–729, 2012. CODEN CSSCDB. ISSN 0361-0918.

Ahmadabadi:2012:ADE

[AFM12] M. Nili Ahmadabadi, Y. Farjami, and M. B. Moghadam. Approximating distributions by extended generalized Lambda distribution (XGLD). Communications in Statistics: Simulation and Computation, 41(1):1–23, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Atil:2014:BSU

[AFS14] Lynda Atil, Hocine Fellag, and Ana Garcia Sipols. Behavior of the size in the unit root testing under contamination. Communications in Statistics: Simulation and Computation, 43(4):790–803, 2014. CODEN CSSCDB. ISSN 0361-0918.

Asimit:2016:SIN

[AFV16] Alexandru V. Asimit, Edward Furman, and Raluca Vernic. Statistical inference for a new class of multivariate Pareto distributions. Communications in Statistics: Simulation and Computation, 45(2):456–471, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ahmed:2015:ESI

[AFvRvR15] S. Ejaz Ahmed, Saber Fallahpour, Dietrich von Rosen, and Tatjana von Rosen. Estimation of several intraclass correlation coefficients. Communications in Statistics: Simulation and Computation, 44(9):2315–2328, 2015. CODEN CSSCDB. ISSN 0361-0918.

Asar:2016:NSP

[AG16] Yasin Asar and Asir Genç. New shrinkage parameters for the Liu-type logistic estimators. Communications in Statistics: Simulation and Computation, 45(3):1094–1103, 2016. CODEN CSSCDB. ISSN 0361-0918.

Asar:2017:TPR

[AG17] Yasin Asar and Asir Genç. Two-parameter ridge estimator in the binary logistic regression. Communications in Statistics: Simulation and Computation, 46(9):7088–7099, 2017. CODEN CSSCDB. ISSN 0361-0918.

Altunkaynak:2019:BCI

[AG19] Bulent Altunkaynak and Hamza Gamgam. Bootstrap confidence intervals for the coefficient of quartile variation. Communications in Statistics: Simulation and Computation, 48 (7):2138–2146, 2019. CODEN CSSCDB. ISSN 0361-0918.

Asghari:2017:LPI

[AGAB17] S. Asghari, B. Sadeghpour Gildeh, J. Ahmadi, and G. Mohtashami Borzadaran. Lifetime performance index based on ranked set sampling. Communications in Statistics: Simulation and Computation, 46(9):7405–7422, 2017. CODEN CSSCDB. ISSN 0361-0918.

Aghamohammadi:2018:BMC

[Agh18] A. Aghamohammadi. A Bayesian multiple comparison procedure for simply ordered treatment medians. Communications in Statistics: Simulation and Computation, 47(4): 939–953, 2018. CODEN CSSCDB. ISSN 0361-0918.

Aggarwal:2015:CCD

[AGJ15] Manohar Aggarwal, Seema Gupta, and Mithilesh Kumar Jha. Construction of cross-over designs for comparing test treatments with control using terraces. Communications in Statistics: Simulation and Computation, 44(3):580–590, 2015. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2017:DMA

[AGKJ17] Muhammad Aslam, Wenhao Gui, Nasrullah Khan, and Chi-Hyuck Jun. Double moving average-EWMA control chart for exponentially distributed quality. Communications in Statistics: Simulation and Computation, 46(9):7351–7364, 2017. CODEN CSSCDB. ISSN 0361-0918.

Abravesh:2019:CBE

[AGM19] Akbar Abravesh, Masoud Ganji, and Behdad Mostafaiy. Classical and Bayesian estimation of stress-strength reliability in type II censored Pareto distributions. Communications in Statistics: Simulation and Computation, 48(8): 2333–2358, 2019. CODEN CSSCDB. ISSN 0361-0918.

Andres:2013:OMR

[AH13a] A. Martín Andrés and M. Álvarez Hernández. Optimal method for realizing two-sided inferences about a linear combination of two proportions. Communications in Statistics: Simulation and Computation, 42(2):327–343, 2013. CODEN CSSCDB. ISSN 0361-0918.

Araki:2013:ERP

[AH13b] Yuko Araki and Satoshi Hattori. Efficient regularization parameter selection via information criteria. Communications in Statistics: Simulation and Computation, 42(2):280–293, 2013. CODEN CSSCDB. ISSN 0361-0918.

Abdel-Hamid:2015:IOD

[AHAH15] Alaa H. Abdel-Hamid and Essam K. Al-Hussaini. Inference and optimal design based on step-partially accelerated life tests for the generalized Pareto distribution under progressive Type-I censoring. Communications in Statistics: Simulation and Computation, 44(7):1750–1769, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ahrari:2019:ETB

[AHB19] V. Ahrari, A. Habibirad, and S. Baratpour. Exponentiality test based on alpha-divergence and gamma-divergence. Communications in Statistics: Simulation and Computation, 48(4):1138–1152, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ariza-Hernandez:2016:BEL

[AHGJRC16] Francisco J. Ariza-Hernandez, Flaviano Godínez-Jaimes, and Ramón Reyes-Carreto. Bayesian estimation of the log-linear exponential regression model with censorship and collinearity. Communications in Statistics: Simulation and Computation, 45(1):152–164, 2016. CODEN CSSCDB. ISSN 0361-0918.

Aly:2017:ODA

[AHM17] Aya A. Aly, Ramadan M. Hamed, and Mahmoud A. Mahmoud. Optimal design of the adaptive exponentially weighted moving average control chart over a range of mean shifts. Communications in Statistics: Simulation and Computation, 46(2):890–902, 2017. CODEN CSSCDB. ISSN 0361-0918.

Abuzaid:2012:SNT

[AHRM12] A. H. Abuzaid, A. G. Hussin, A. Rambli, and I. Mohamed. Statistics for a new test of discordance in circular data. Communications in Statistics: Simulation and Computation, 41 (10):1882–1890, 2012. CODEN CSSCDB. ISSN 0361-0918.

Asar:2017:EBC

[AID17] Özgür Asar, Ozlem Ilk, and Osman Dag. Estimating Box– Cox power transformation parameter via goodness-of-fit tests. Communications in Statistics: Simulation and Computation, 46(1):91–105, 2017. CODEN CSSCDB. ISSN 0361-0918.

Al-Jararha:2013:CST

[AJ13] Jehad Al-Jararha. A class of sampling two units with probability proportional to size. Communications in Statistics: Simulation and Computation, 42(8):1906–1916, 2013. CO-DEN CSSCDB. ISSN 0361-0918.

Aminnejad:2017:BOD

[AJ17] Minoo Aminnejad and Habib Jafari. Bayesian A- and Doptimal designs for gamma regression model with inverse link function. Communications in Statistics: Simulation and Computation, 46(10):8166–8189, 2017. CODEN CSSCDB. ISSN 0361-0918.

Abdollahnezhad:2018:TEQ

[AJ18] Kamel Abdollahnezhad and Ali Akbar Jafari. Testing the equality of quantiles for several normal populations. Communications in Statistics: Simulation and Computation, 47 (7):1890–1898, 2018. CODEN CSSCDB. ISSN 0361-0918.

Amini:2011:ADG

[AJB11] M. Amini, H. Jabbari, and G. R. Mohtashami Borzadaran. Aspects of dependence in generalized Farlie-Gumbel-Morgenstern distributions. Communications in Statistics: Simulation and Computation, 40(8):1192–1205, 2011. CO-DEN CSSCDB. ISSN 0361-0918.

Ahmad:2010:EDF

[AJM10] K. E. Ahmad, Z. F. Jaheen, and A. A. Modhesh. Estimation of a discriminant function based on small sample size from a mixture of two Gumbel distributions. Communications in Statistics: Simulation and Computation, 39(4):713–725, 2010. CODEN CSSCDB. ISSN 0361-0918.

Armstrong:2011:CHT

[AK11] Ronald D. Armstrong and Mabel T. Kung. CUSUM hypothesis tests and alternative response probabilities for finite Poisson trials. Communications in Statistics: Simulation and Computation, 40(7):1057–1073, 2011. CODEN CSSCDB. ISSN 0361-0918.

Arakelian:2014:CDM

[AK14] Veni Arakelian and Dimitris Karlis. Clustering dependencies via mixtures of copulas. Communications in Statistics: Simulation and Computation, 43(7):1644–1661, 2014. CODEN CSSCDB. ISSN 0361-0918.

Abdel-Karim:2015:CSC

[AK15] Amany Hassan Abdel-Karim. Construction of simultaneous confidence intervals for ratios of means of lognormal distributions. Communications in Statistics: Simulation and Computation, 44(2):271–283, 2015. CODEN CSSCDB. ISSN 0361-0918.

Abdel-Karim:2017:EZO

[AK17] Amany Hassan Abdel-Karim. Extended zero-one inflated beta and adjusted three-part regression models for proportional data analysis. Communications in Statistics: Simulation and Computation, 46(8):6155–6172, 2017. CODEN CSSCDB. ISSN 0361-0918.

Atalay:2019:PRL

[AK19] Sevcan Demir Atalay and Gözde Kus. Profust reliability of linear and circular type F and G systems having two failure criteria under Markov dependency. Communications in Statistics: Simulation and Computation, 48(1):219–239, 2019. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2017:DCC

[AKJ17] Muhammad Aslam, Nasrullah Khan, and Chi-Hyuck Jun. Designing of a control chart using belief statistic for exponential distribution. Communications in Statistics: Simulation and Computation, 46(5):3781–3793, 2017. CODEN CSSCDB. ISSN 0361-0918.

Androulakis:2011:TPE

[AKM11] E. Androulakis, C. Koukouvinos, and K. Mylona. Tuning parameter estimation in penalized least squares methodology. Communications in Statistics: Simulation and Computation, 40(9):1444–1457, 2011. CODEN CSSCDB. ISSN 0361-0918.

Angelopoulos:2013:CEU

[AKS13a] P. Angelopoulos, C. Koukouvinos, and A. Skountzou. Clustering effects in unreplicated factorial experiments. Communications in Statistics: Simulation and Computation, 42(9): 1998–2007, 2013. CODEN CSSCDB. ISSN 0361-0918.

Angelopoulos:2013:CNT

[AKS13b] P. Angelopoulos, C. Koukouvinos, and A. Skountzou. Construction of new three-level response surface designs. Communications in Statistics: Simulation and Computation, 42 (7):1587–1595, 2013. CODEN CSSCDB. ISSN 0361-0918.

Androulakis:2016:NMA

[AKV16a] E. Androulakis, C. Koukouvinos, and F. Vonta. A new method for the analysis of supersaturated designs with discrete data. Communications in Statistics: Simulation and Computation, 45(6):1971–1990, 2016. CODEN CSSCDB. ISSN 0361-0918.

Androulakis:2016:TPS

[AKV16b] E. Androulakis, C. Koukouvinos, and F. Vonta. Tuning parameter selection in penalized frailty models. Communications in Statistics: Simulation and Computation, 45(5): 1538–1553, 2016. CODEN CSSCDB. ISSN 0361-0918.

Albatineh:2017:ESD

[AKZK17] Ahmed N. Albatineh, Hafiz M. R. Khan, Bashar Zogheib, and Golam B. M. Kibria. Effects of some design factors on the distribution of similarity indices in cluster analysis. Communications in Statistics: Simulation and Computation, 46(5):4018–4034, 2017. CODEN CSSCDB. ISSN 0361-0918.

AlMutairi:2015:VTS

[AL15] Alya O. Al Mutairi and Heng Chin Low. p-value for two-sample linear rank tests using permutation simulations and the normal approximation method. Communications in Statistics: Simulation and Computation, 44(3):819–826, 2015. CODEN CSSCDB. ISSN 0361-0918.

AlMutairi:2016:IMS

[AL16] Alya O. Al Mutairi and Heng Chin Low. Improved measures of the spread of data for some unknown complex distributions using saddlepoint approximations. Communications in Statistics: Simulation and Computation, 45(1):33–47, 2016. CODEN CSSCDB. ISSN 0361-0918.

Angali:2014:BEB

[ALH14] Kambiz Ahmadi Angali, S. Mahmoud Latifi, and David D. Hanagal. Bayesian estimation of bivariate exponential distributions based on LINEX and quadratic loss functions: A survival approach with censored samples. Communications in Statistics: Simulation and Computation, 43(1):31– 44, 2014. CODEN CSSCDB. ISSN 0361-0918.

Alhamzawi:2016:BEN

[Alh16] Rahim Alhamzawi. Bayesian elastic net tobit quantile regression. Communications in Statistics: Simulation and Computation, 45(7):2409–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Alkan:2017:ACD

[Alk17] Nesrin Alkan. Assessing convergence diagnostic tests for Bayesian Cox regression. Communications in Statistics: Simulation and Computation, 46(4):3201–3212, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Almasri:2010:TTS

[Alm10] Abdullah Almasri. Tests for trend: A simulation study. Communications in Statistics: Simulation and Computation, 39(3):598–611, 2010. CODEN CSSCDB. ISSN 0361-0918.

Ala-Luhtala:2016:GSM

[ALP16] Juha Ala-Luhtala and Robert Piché. Gaussian scale mixture models for robust linear multivariate regression with missing data. Communications in Statistics: Simulation and Computation, 45(3):791–813, 2016. CODEN CSSCDB. ISSN 0361-0918.

Amiri:2013:SND

[AM13] Amirhossein Amiri and Elnaz Moein. Some notes on diagnostic procedures in simple and multiple linear regression profiles monitoring. Communications in Statistics: Simulation and Computation, 42(5):981–1002, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Ajdari:2014:AEE

[AM14] Ali Ajdari and Hashem Mahlooji. An adaptive explorationexploitation algorithm for constructing metamodels in random simulation using a novel sequential experimental design. Communications in Statistics: Simulation and Computation, 43(5):947–968, 2014. CODEN CSSCDB. ISSN 0361-0918.

Aghamohammadi:2017:GSM

[AM17a] Ali Aghamohammadi and M. R. Meshkani. Gibbs sampling method for the Bayesian adaptive elastic net. Communications in Statistics: Simulation and Computation, 46(6): 4264–4273, 2017. CODEN CSSCDB. ISSN 0361-0918.

Albis:2017:BAC

[AM17b] Manuel Leonard F. Albis and Dennis S. Mapa. Bayesian averaging of classical estimates in asymmetric vector autoregressive models. Communications in Statistics: Simulation and Computation, 46(3):1760–1770, 2017. CODEN CSSCDB. ISSN 0361-0918.

Amini:2017:NEE

[AM17c] Morteza Amini and Mahdi Mahdizadeh. Nonparametric estimation of the entropy using a ranked set sample. Communications in Statistics: Simulation and Computation, 46(9): 6719–6737, 2017. CODEN CSSCDB. ISSN 0361-0918.

Abdelfatah:2019:ICT

[AM19] Sally Abdelfatah and Reda Mazloum. An improved cluster two-stage randomized response model. Communications in Statistics: Simulation and Computation, 48(1):58–72, 2019. CODEN CSSCDB. ISSN 0361-0918.

Aguilera-Morillo:2015:PSE

[AMA15] M. Carmen Aguilera-Morillo and Ana M. Aguilera. P-spline estimation of functional classification methods for improving the quality in the food industry. Communications in Statistics: Simulation and Computation, 44(10):2513–2534, 2015. CODEN CSSCDB. ISSN 0361-0918.

Al-Marshadi:2017:EMG

[AMAA17] Ali Hussein Al-Marshadi, Abdullah Hamoud Alharby, and Muhammad Aslam. Evaluating modified generalized information criterion in presence of multicollinearity. Communications in Statistics: Simulation and Computation, 46(8): 6298–6307, 2017. CODEN CSSCDB. ISSN 0361-0918.

Amin:2019:BID

[Ami19a] Ayman A. Amin. Bayesian identification of double seasonal autoregressive time series models. Communications in Statistics: Simulation and Computation, 48(8):2501–2511, 2019. CODEN CSSCDB. ISSN 0361-0918.

Amin:2019:KLD

[Ami19b] Ayman A. Amin. Kullback–Leibler divergence to evaluate posterior sensitivity to different priors for autoregressive time series models. Communications in Statistics: Simulation and Computation, 48(5):1277–1291, 2019. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2017:NCU

[AMJ17] Muhammad Aslam, Muhammad Mohsin, and Chi-Hyuck Jun. A new t-chart using process capability index. Communications in Statistics: Simulation and Computation, 46 (7):5141–5150, 2017. CODEN CSSCDB. ISSN 0361-0918.

Auda:2019:MCS

[AMKS19] Hend A. Auda, Joseph W. McKean, John D. Kloke, and Mahmoud Sadek. A Monte Carlo study of REML and robust rank-based analyses for the random intercept mixed model. Communications in Statistics: Simulation and Computation, 48(3):837–860, 2019. CODEN CSSCDB. ISSN 0361-0918.

Almasi:2017:BLU

[AMM17] Isaac Almasi, Adel Mohammadpour, and Mohammad Mohammadi. Best linear unbiased interpolation of order statistics. Communications in Statistics: Simulation and Computation, 46(5):4161–4171, 2017. CODEN CSSCDB. ISSN 0361-0918.

Almasi:2018:BUP

[AMM18] Isaac Almasi, Adel Mohammadpour, and Mohammad Mohammadi. Best unbiased prediction of order statistics in stable distributions. Communications in Statistics: Simulation and Computation, 47(8):2424–2435, 2018. CODEN CSSCDB. ISSN 0361-0918.

Acosta-Mejia:2010:RSC

[AMP10] Cesar Acosta-Mejia and Joseph J. Pignatiello, Jr. The run sum R chart with fast initial response. Communications in Statistics: Simulation and Computation, 39(5):921–932, 2010. CODEN CSSCDB. ISSN 0361-0918.

Al-Moisheer:2016:EDF

[AMS16] A. S. Al-Moisheer and K. S. Sultan. Estimation of a discriminant function from a mixture of two Burr Type III distributions. Communications in Statistics: Simulation and Computation, 45(10):3760–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Aly:2015:CPP

[AMW15] Aya A. Aly, Mahmoud A. Mahmoud, and William H. Woodall. A comparison of the performance of Phase II simple linear profile control charts when parameters are estimated. Communications in Statistics: Simulation and Computation, 44(6):1432–1440, 2015. CODEN CSSCDB. ISSN 0361-0918.

Abebe:2011:SNA

[AN11] Asheber Abebe and Sai V. Nudurupati. Smooth nonparametric allocation of classification. Communications in Statistics: Simulation and Computation, 40(5):694–709, 2011. CODEN CSSCDB. ISSN 0361-0918.

Akhundov:2012:SDS

[AN12] Ilham Akhundov and Valery Nevzorov. From Student’s t_{2} distribution to Student’s t_{3} distribution through characterizations. Communications in Statistics: Simulation and Computation, 41(6):710–716, 2012. CODEN CSSCDB. ISSN 0361-0918.

Ahsanullah:2016:SNR

[AN16] M. Ahsanullah and Prof. V. B. Nevzorov. On some new record schemes. Communications in Statistics: Simulation and Computation, 45(5):1467–1476, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Amiri:2014:NAV

[ANA14] Amirhossein Amiri, Ali Nedaie, and Mahdi Alikhani. A new adaptive variable sample size approach in EWMA control chart. Communications in Statistics: Simulation and Computation, 43(4):804–812, 2014. CODEN CSSCDB. ISSN 0361-0918.

Auda:2019:CNG

[ANB19] Hend Auda and Magdalena Niewiadomska-Bugaj. Comparing a new Gini test with other symmetry tests when median is known. Communications in Statistics: Simulation and Computation, 48(8):2401–2412, 2019. CODEN CSSCDB. ISSN 0361-0918.

Alizadeh:2017:CBO

[AND^{+}17] M. Alizadeh, S. Nadarajah, M. Doostparast, A. Parchami, and M. Mashinchi. On confidence bounds for one-parameter exponential families. Communications in Statistics: Simulation and Computation, 46(2):1569–1582, 2017. CODEN CSSCDB. ISSN 0361-0918.

Angus:2013:VCC

[Ang13] John E. Angus. A variation on the coupon collecting problem. Communications in Statistics: Simulation and Computation, 42(10):2197–2202, 2013. CODEN CSSCDB. ISSN 0361-0918.

Anisimov:2016:PHM

[Ani16] Vladimir V. Anisimov. Predictive hierarchic modeling of operational characteristics in clinical trials. Communications in Statistics: Simulation and Computation, 45(5): 1477–1488, 2016. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2011:C

[Ano11a] Anonymous. Corrigendum. Communications in Statistics: Simulation and Computation, 40(9):1524, 2011. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2011:EBE

[Ano11b] Anonymous. Editorial board EOV. Communications in Statistics: Simulation and Computation, 40(10):ebi–ebii, 2011. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2012:EBE

[Ano12] Anonymous. Editorial board EOV. Communications in Statistics: Simulation and Computation, 41(10):ebi–ebii, 2012. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2014:C

[Ano14a] Anonymous. Corrigendum. Communications in Statistics: Simulation and Computation, 43(2):441, 2014. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2014:E

[Ano14b] Anonymous. Erratum. Communications in Statistics: Simulation and Computation, 43(2):442, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Anonymous:2015:Ea

[Ano15a] Anonymous. Erratum. Communications in Statistics: Simulation and Computation, 44(1):270, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Anonymous:2015:Eb

[Ano15b] Anonymous. Erratum. Communications in Statistics: Simulation and Computation, 44(6):1657, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Anonymous:2017:EBE

[Ano17a] Anonymous. Editorial Board:EOV. Communications in Statistics: Simulation and Computation, 46(10):ebi, 2017. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2017:E

[Ano17b] Anonymous. Erratum. Communications in Statistics: Simulation and Computation, 46(6):5021, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Anonymous:2018:C

[Ano18a] Anonymous. Corrigendum. Communications in Statistics: Simulation and Computation, 47(2):629, 2018. CODEN CSSCDB. ISSN 0361-0918.

Anonymous:2018:EBE

[Ano18b] Anonymous. Editorial Board: Eov. Communications in Statistics: Simulation and Computation, 47(10):ebi, 2018. CODEN CSSCDB. ISSN 0361-0918.

Alih:2017:RCB

[AO17] Ekele Alih and Hong Choon Ong. Robust cluster-based multivariate outlier diagnostics and parameter estimation in regression analysis. Communications in Statistics: Simulation and Computation, 46(1):244–260, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ahmed:2012:SPK

[AOJ12] S. Ejaz Ahmed, M. Hafidz Omar, and Anwar H. Joarder. Stabilizing the performance of kurtosis estimator of multivariate data. Communications in Statistics: Simulation and Computation, 41(10):1860–1871, 2012. CODEN CSSCDB. ISSN 0361-0918.

Akahira:2013:HOA

[AOK13] M. Akahira, N. Ohyauchi, and S. Kawai. A higher order approximation to a percentage point of the distribution of a noncentral t-statistic without the normality assumption. Communications in Statistics: Simulation and Computation, 42(9):2086–2105, 2013. CODEN CSSCDB. ISSN 0361-

Aoki:2015:MMB

[Aok15] Satoshi Aoki. Minimal Markov basis for tests of main effect models for 2^{p−1} fractional factorial designs of resolution p. Communications in Statistics: Simulation and Computation, 44(9):2371–2386, 2015. CODEN CSSCDB. ISSN 0361-0918.

Alam:2017:TET

[AP17] Khurshid Alam and Sudhir Paul. Testing equality of two beta binomial proportions in the presence of unequal extradispersion parameters. Communications in Statistics: Simulation and Computation, 46(4):2784–2799, 2017. CODEN CSSCDB. ISSN 0361-0918.

Alharbi:2018:EVF

[AP18] Yousef F. Alharbi and Prakash N. Patili. Error variance function estimation in nonparametric regression models. Communications in Statistics: Simulation and Computation, 47(5):1479–1491, 2018. CODEN CSSCDB. ISSN 0361-0918.

Asmussen:2010:NSR

[AR10] Søren Asmussen and Tobias Rydén. A note on skewness in regenerative simulation. Communications in Statistics: Simulation and Computation, 40(1):45–57, 2010. CODEN CSSCDB. ISSN 0361-0918.

Almomani:2012:SGS

[AR12] Mohammad H. Almomani and Rosmanjawati Abdul Rahman. Selecting a good stochastic system for the large number of alternatives. Communications in Statistics: Simulation and Computation, 41(2):222–237, 2012. CODEN CSSCDB. ISSN 0361-0918.

Andrade:2017:LRD

[AR17a] P. Andrade and L. Rifo. Long-range dependence and approximate Bayesian computation. Communications in Statistics: Simulation and Computation, 46(2):1219–1237, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Asha:2017:EFB

[AR17b] G. Asha and A. Vincent Raja. An extension of the Freund’s bivariate distribution to model cascading failures. Communications in Statistics: Simulation and Computation, 46(7): 5516–5530, 2017. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2013:EER

[ARA13] Muhammad Aslam, Tahira Riaz, and Saima Altaf. Efficient estimation and robust inference of linear regression models in the presence of heteroscedastic errors and high leverage points. Communications in Statistics: Simulation and Computation, 42(10):2223–2238, 2013. CODEN CSSCDB. ISSN 0361-0918.

Awwad:2015:SIB

[ARAM15] Raed R. Abu Awwad, Mohammad Z. Raqab, and Intesar M. Al-Mudahakha. Statistical inference based on progressively Type II censored data from Weibull model. Communications in Statistics: Simulation and Computation, 44(10): 2654–2670, 2015. CODEN CSSCDB. ISSN 0361-0918.

AmorimAmaral:2013:GNM

[ARC13] Getulio Jose Amorim Amaral, Olga Patricia Reyes Floréz, and Francisco José A. Cysneiros. Graphical and numerical methods for detecting influential observations in complex Bingham data. Communications in Statistics: Simulation and Computation, 42(8):1801–1814, 2013. CODEN CSSCDB. ISSN 0361-0918.

Ahmad:2017:DBG

[ARK17] Murad A. Ahmad, Mohammad Z. Raqab, and Debasis Kundu. Discriminating between the generalized Rayleigh and Weibull distributions: Some comparative studies. Communications in Statistics: Simulation and Computation, 46 (6):4880–4895, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ali:2013:SMS

[ARKB13] Irfan Ali, Yashpal Singh Raghav, M. Faisal Khan, and Abdul Bari. Selective maintenance in system reliability with random costs of repairing and replacing the components. Communications in Statistics: Simulation and Computation, 42 (9):2026–2039, 2013. CODEN CSSCDB. ISSN 0361-0918.

Alizadeh:2019:NUF

[ARM^{+}19] Morad Alizadeh, Thiago G. Ramires, S. M. T. K. Mir-Mostafaee, Maryam Samizadeh, and Edwin M. M. Ortega.

A new useful four-parameter extension of the Gumbel distribution: Properties, regression model and applications using the GAMLSS framework. Communications in Statistics: Simulation and Computation, 48(6):1746–1767, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Azuero:2012:SDF

[ART^{+}12] Andres Azuero, David T. Redden, Hemant K. Tiwari, Senait G. Asmellash, and Chandrika J. Piyathilake. A simple distribution-free algorithm for generating simulated highdimensional correlated data with an autoregressive structure. Communications in Statistics: Simulation and Computation, 41(1):89–98, 2012. CODEN CSSCDB. ISSN 0361-0918.

Arnold:2016:IDA

[ARW16] Matthias Arnold, Nils Raabe, and Dominik Wied. Identifying different areas of inhomogeneous mineral subsoil: Spatial fluctuation approaches. Communications in Statistics: Simulation and Computation, 45(1):252–263, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ahmadi:2018:PTI

[ARY18] K. Ahmadi, M. Rezaei, and F. Yousefzadeh. Progressively Type-II censored competing risks data for exponential distributions based on sequential order statistics. Communications in Statistics: Simulation and Computation, 47(5): 1276–1296, 2018. CODEN CSSCDB. ISSN 0361-0918.

Adler:2012:RRO

[AS12a] André Adler and George Skountrianos. Resampling ratios of order statistics (RROS). Communications in Statistics: Simulation and Computation, 41(10):1891–1894, 2012. CO-DEN CSSCDB. ISSN 0361-0918.

Alkahtani:2012:DME

[AS12b] Saad Alkahtani and Jay Schaffer. A double multivariate exponentially weighted moving average (dMEWMA) control chart for a process location monitoring. Communications in Statistics: Simulation and Computation, 41(2):238–252, 2012. CODEN CSSCDB. ISSN 0361-0918.

Andronov:2012:SEM

[AS12c] Alexander Andronov and Diana Santalova. Statistical estimates for modified gravitation model by aggregated data. Communications in Statistics: Simulation and Computation, 41(6):730–745, 2012. CODEN CSSCDB. ISSN 0361-0918.

Alzubaidi:2014:IST

[AS14a] Hasan Alzubaidi and Tony Shardlow. Improved simulation techniques for first exit time of neural diffusion models. Communications in Statistics: Simulation and Computation, 43(10):2508–2520, 2014. CODEN CSSCDB. ISSN 0361-0918.

Awan:2014:ORE

[AS14b] Wajid Hussain Awan and Javid Shabbir. On optimum regression estimator for population mean using two auxiliary variables in simple random sampling. Communications in Statistics: Simulation and Computation, 43(6):1508–1522, 2014. CODEN CSSCDB. ISSN 0361-0918.

Aytacoglu:2015:RCC

[AS15] Burcu Aytacoglu and Hakan S. Sazak. A robust control chart for monitoring the mean of an autocorrelated process. Communications in Statistics: Simulation and Computation, 44 (7):1787–1800, 2015. CODEN CSSCDB. ISSN 0361-0918.

Amro:2017:MPP

[AS17] Lubna Amro and Monjed H. Samuh. More powerful permutation test based on multistage ranked set sampling. Communications in Statistics: Simulation and Computation, 46 (7):5271–5284, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ahmed:2013:MND

[ASA13] Rashid Ahmed, Farrukh Shehzad, and Munir Akhtar. Minimal neighbor designs in linear blocks. Communications in Statistics: Simulation and Computation, 42(9):2135–2139, 2013. CODEN CSSCDB. ISSN 0361-0918.

Asar:2017:SNM

[Asa17] Yasin Asar. Some new methods to solve multicollinearity in logistic regression. Communications in Statistics: Simulation and Computation, 46(4):2576–2586, 2017. CODEN CSSCDB. ISSN 0361-0918.

Asikgil:2014:NAE

[Asi14] Baris Asikgil. A novel approach for estimating seemingly unrelated regressions with high-order autoregressive disturbances. Communications in Statistics: Simulation and Computation, 43(9):2061–2080, 2014. CODEN CSSCDB. ISSN 0361-0918.

Akat:2019:EAM

[ASKT19] Fulya Akat, A. Sevtap Selcuk-Kestel, and Fatih Tank. The estimation of adopted mortality and morbidity rates using model and the phase type law: the Turkish case. Communications in Statistics: Simulation and Computation, 48(9): 2552–2565, 2019. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2014:PKM

[Asl14a] Muhammad Aslam. Performance of Kibria’s method for the heteroscedastic ridge regression model: Some Monte Carlo evidence. Communications in Statistics: Simulation and Computation, 43(4):673–686, 2014. CODEN CSSCDB. ISSN 0361-0918.

Aslam:2014:UHC

[Asl14b] Muhammad Aslam. Using heteroscedasticity-consistent standard errors for the linear regression model with correlated regressors. Communications in Statistics: Simulation and Computation, 43(10):2353–2373, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Aslam:2019:TTA

[Asl19] Muhammad Aslam. Time truncated attribute control chart for the Weibull distribution using multiple dependent state sampling. Communications in Statistics: Simulation and Computation, 48(4):1219–1228, 2019. CODEN CSSCDB. ISSN 0361-0918.

Amo-Salas:2012:ODT

[ASLFLR12] M. Amo-Salas, J. López-Fidalgo, and V. I. López-Ríos. Optimal designs for two nested pharmacokinetic models with correlated observations. Communications in Statistics: Simulation and Computation, 41(7):944–963, 2012. CODEN CSSCDB. ISSN 0361-0918.

Aregay:2015:CAM

[ASM15] Mehreteab Aregay, Ziv Shkedy, and Geert Molenberghs. Comparison of additive and multiplicative Bayesian models for longitudinal count data with overdispersion parameters: A simulation study. Communications in Statistics: Simulation and Computation, 44(2):454–473, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Ambati:2018:ECD

[ASS18] Sai Kiran Ambati, Stephen A. Sedory, and Sarjinder Singh. Estimation of coefficient of dispersion using auxiliary information. Communications in Statistics: Simulation and Computation, 47(7):1996–2009, 2018. CODEN CSSCDB. ISSN 0361-0918.

Aggarwal:2010:NOO

[ASSH10] M. L. Aggarwal, Poonam Singh, Vandana Sarin, and Bushra Husain. Nearly optimal orthogonally blocked designs for four mixture components based on F -squares. Communications in Statistics: Simulation and Computation, 40(1):165–183, 2010. CODEN CSSCDB. ISSN 0361-0918.

Abbas:2016:OBA

[AT16] Kamran Abbas and Yincai Tang. Objective Bayesian analysis for log-logistic distribution. Communications in Statistics: Simulation and Computation, 45(8):2782–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Arashi:2014:SRE

[ATB14] M. Arashi, S. M. M. Tabatabaey, and M. Hassanzadeh Bashtian. Shrinkage ridge estimators in linear regression. Communications in Statistics: Simulation and Computation, 43 (4):871–904, 2014. CODEN CSSCDB. ISSN 0361-0918.

Alkan:2017:UIP

[ATC17] Nesrin Alkan, Yuksel Terzi, and M. Ali Cengiz. Using informative priors for handling missing data problem in Cox regression. Communications in Statistics: Simulation and Computation, 46(10):7614–7623, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ateya:2017:EUI

[Ate17] Saieed F. Ateya. Estimation under inverse Weibull distribution based on Balakrishnan’s unified hybrid censored scheme. Communications in Statistics: Simulation and Computation, 46(5):3645–3666, 2017. CODEN CSSCDB. ISSN 0361-0918.

Atilgan:2016:RCA

[Ati16] Yasemin Kayhan Atilgan. Robust coplot analysis. Communications in Statistics: Simulation and Computation, 45(5): 1763–1775, 2016. CODEN CSSCDB. ISSN 0361-0918.

Abebe:2014:CPB

[ATV^{+}14] Haftom T. Abebe, Frans E. S. Tan, Gerard J. P. Van Breukelen, Jan Serroyen, and Martijn P. F. Berger. On the choice of a prior for Bayesian D-optimal designs for the logistic regression model with a single predictor. Communications in Statistics: Simulation and Computation, 43(7):1811–1824, 2014. CODEN CSSCDB. ISSN 0361-0918.

Asghar:2017:ASL

[AU17] Zahid Asghar and Amena Urooj. Analysis of seasonal level shift (SLS) detection in SARIMA models. Communications in Statistics: Simulation and Computation, 46(9): 7264–7318, 2017. CODEN CSSCDB. ISSN 0361-0918.

Austin:2010:DGP

[Aus10] Peter C. Austin. A data-generation process for data with specified risk differences or numbers needed to treat. Communications in Statistics: Simulation and Computation, 39 (3):563–577, 2010. CODEN CSSCDB. ISSN 0361-0918.

Asgharzadeh:2017:ETP

[AVR17] A. Asgharzadeh, R. Valiollahi, and M. Z. Raqab. Estimation of Pr(Y < X) for the two-parameter generalized exponential records. Communications in Statistics: Simulation and Computation, 46(1):379–394, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ahmadi:2015:EPG

[AY15] K. Ahmadi and F. Yousefzadeh. Estimation for the parameters of generalized half-normal distribution based on progressive Type-I interval censoring. Communications in Statistics: Simulation and Computation, 44(10):2671–2695, 2015. CODEN CSSCDB. ISSN 0361-0918.

Akay:2018:CMP

[AY18a] Özlem Akay and Güzin Yüksel. Clustering the mixed panel dataset using Gower’s distance and k-prototypes algorithms. Communications in Statistics: Simulation and Computation, 47(10):3031–3041, 2018. CODEN CSSCDB. ISSN 0361-0918.

Aydin:2018:MSR

[AY18b] Dursun Aydin and Ersin Yilmaz. Modified spline regression based on randomly right-censored data: a comparative study. Communications in Statistics: Simulation and Computation, 47(9):2587–2611, 2018. CODEN CSSCDB. ISSN 0361-0918.

Ahmadi:2017:PTI

[AYR17] K. Ahmadi, F. Yousefzadeh, and M. Rezaei. Progressively Type-I interval censored competing risks data for the proportional hazards family. Communications in Statistics: Simulation and Computation, 46(8):5924–5950, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Amiri:2010:INB

[AZ10] Saeid Amiri and Silvelyn Zwanzig. An improvement of the nonparametric bootstrap test for the comparison of the coefficient of variations. Communications in Statistics: Simulation and Computation, 39(9):1726–1734, 2010. CODEN CSSCDB. ISSN 0361-0918.

Al-Zoubi:2010:AIM

[AZCS10] Loai Mahmoud Al-Zou’bi, Robert Graham Clark, and David G. Steel. Adaptive inference for multi-stage survey data. Communications in Statistics: Simulation and Computation, 39(7):1334–1350, 2010. CODEN CSSCDB. ISSN 0361-0918.

Asadzadeh:2011:CEB

[AZSA11] Shervin Asadzadeh, Yaser Zerehsaz, Abbas Saghaei, and Abdollah Aghaie. Compound-estimator based cause-selecting control chart for monitoring multistage processes. Communications in Statistics: Simulation and Computation, 40(3): 322–344, 2011. CODEN CSSCDB. ISSN 0361-0918.

Beyaztas:2013:JAB

[BA13] Ufuk Beyaztas and Aylin Alin. Jackknife-after-bootstrap method for detection of influential observations in linear regression models. Communications in Statistics: Simulation and Computation, 42(6):1256–1267, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Beyaztas:2014:JAB

[BA14] Ufuk Beyaztas and Aylin Alin. Jackknife-after-bootstrap as logistic regression diagnostic tool. Communications in Statistics: Simulation and Computation, 43(9):2047–2060, 2014. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2019:MLE

[BA19] N. Balakrishnan and Farouq Mohammad A. Alam. Maximum likelihood estimation of the parameters of a multiple step-stress model from the Birnbaum–Saunders distribution under time-constraint: a comparative study. Communications in Statistics: Simulation and Computation, 48 (5):1535–1559, 2019. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2016:AVI

[BAA16] S. Balamurali, Muhammad Azam, and Muhammad Aslam. Attribute-variable inspection policy for lots using resampling based on EWMA. Communications in Statistics: Simulation and Computation, 45(8):3014–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bahcecitapar:2018:SFA

[Bah18] Melike Kaya Bahçecitapar. Some factors affecting statistical power of approximate tests in the linear mixed model for longitudinal data. Communications in Statistics: Simulation and Computation, 47(1):294–314, 2018. CODEN CSSCDB. ISSN 0361-0918.

Baklizi:2016:CIE

[Bak16] Ayman Baklizi. Comparison of interval estimators of Pr(X < Y) in the two-parameter exponential distribution. Communications in Statistics: Simulation and Computation, 45(8):2937–??, 2016. CODEN CSSCDB. ISSN 0361-

Balamurali:2017:NMC

[Bal17] S. Balamurali. A new mixed chain sampling plan based on the process capability index for product acceptance. Communications in Statistics: Simulation and Computation, 46 (7):5423–5439, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bansal:2011:THG

[BAM11] Pooja Bansal, Sangeeta Arora, and Kalpana K. Mahajan. Testing homogeneity of Gini indices against simple-ordered alternative. Communications in Statistics: Simulation and Computation, 40(2):185–198, 2011. CODEN CSSCDB. ISSN 0361-0918.

Bidram:2015:EEW

[BAN15] H. Bidram, M. H. Alamatsaz, and V. Nekoukhou. On an extension of the exponentiated Weibull distribution. Communications in Statistics: Simulation and Computation, 44 (6):1389–1404, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bagheri:2016:EEPa

[BAN16] S. F. Bagheri, M. Alizadeh, and S. Nadarajah. Efficient estimation of the PDF and the CDF of the exponentiated Gumbel distribution. Communications in Statistics: Simulation and Computation, 45(1):339–361, 2016. CODEN CSSCDB. ISSN 0361-0918.

Banik:2017:SAB

[Ban17] A. D. Banik. Stationary analysis of a BMAP/R/1 queue with R-type multiple working vacations. Communications in Statistics: Simulation and Computation, 46(2):1035–1061, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bagheri:2016:EEPb

[BAND16] S. F. Bagheri, M. Alizadeh, S. Nadarajah, and E. Deiri. Efficient estimation of the PDF and the CDF of the Weibull extension model. Communications in Statistics: Simulation and Computation, 45(6):2191–2207, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Borgatto:2015:CAE

[BAPA15] Adriano Ferreti Borgatto, Caio Azevedo, Aluisio Pinheiro, and Dalton Andrade. Comparison of ability estimation methods using IRT for tests with different degrees of difficulty. Communications in Statistics: Simulation and Computation, 44(2):474–488, 2015. CODEN CSSCDB. ISSN 0361-0918.

Barbiero:2011:CIR

[Bar11] Alessandro Barbiero. Confidence intervals for reliability of stress-strength models in the normal case. Communications in Statistics: Simulation and Computation, 40(6):907–925, 2011. CODEN CSSCDB. ISSN 0361-0918.

Barbiero:2017:LSM

[Bar17a] Alessandro Barbiero. Least-squares and minimum chi-square estimation in a discrete Weibull model. Communications in Statistics: Simulation and Computation, 46(10):8028–8048, 2017. CODEN CSSCDB. ISSN 0361-0918.

Barghout:2017:PES

[Bar17b] May Barghout. On parameter estimation of software reliability models. Communications in Statistics: Simulation and Computation, 46(2):910–932, 2017. CODEN CSSCDB. ISSN 0361-0918.

Barnabani:2017:ACG

[Bar17c] Marco Barnabani. An approximation to the convolution of gamma distributions. Communications in Statistics: Simulation and Computation, 46(1):331–343, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Bayoud:2011:BEB

[Bay11] Husam Awni Bayoud. Bayes and empirical Bayes estimation of the parameter n in a binomial distribution. Communications in Statistics: Simulation and Computation, 40(9): 1422–1433, 2011. CODEN CSSCDB. ISSN 0361-0918.

Bazyari:2017:BAT

[Baz17] Abouzar Bazyari. Bootstrap approach to test the homogeneity of order restricted mean vectors when the covariance matrices are unknown. Communications in Statistics: Simulation and Computation, 46(9):7194–7209, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bastero:2011:RES

[BB11] Rowena F. Bastero and Erniel B. Barrios. Robust estimation of a spatiotemporal model with structural change. Communications in Statistics: Simulation and Computation, 40(3): 448–468, 2011. CODEN CSSCDB. ISSN 0361-0918.

Bayramoglu:2014:CBR

[BB14] Konul Bayramoglu and Ismihan Bayramoglu (Bairamov). On censored bivariate random variables: Copula, characterization, and estimation. Communications in Statistics: Simulation and Computation, 43(10):2173–2185, 2014. CODEN CSSCDB. ISSN 0361-0918.

Bakhtiarifar:2017:PCM

[BB17a] Mohammad Hasan Bakhtiarifar and Mahdi Bashiri. A probabilistic clustering method for data elements with normal distributed attributes. Communications in Statistics: Simulation and Computation, 46(4):2563–2575, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balabdaoui:2017:LEC

[BB17b] Fadoua Balabdaoui and Saonli Basu. Letter to the editor: Comments on Groparu-Cojocaru and Doray (2013). Communications in Statistics: Simulation and Computation, 46 (5):3833–3840, 2017. CODEN CSSCDB. ISSN 0361-0918. See [GCD13].

Banerjee:2017:TER

[BB17c] Buddhananda Banerjee and Atanu Biswas. True endpoint reduction by surrogate endpoints. Communications in Statistics: Simulation and Computation, 46(8):6645–6653, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2017:PIV

[BB17d] Mohamed Bentarzi and Wissam Bentarzi. Periodic integervalued GARCH(1, 1) model. Communications in Statistics: Simulation and Computation, 46(2):1167–1188, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Bonner:2017:EPS

[BB17e] Ashley J. Bonner and Joseph Beyene. Evaluating the performance of sparse principal component analysis methods in high-dimensional data scenarios. Communications in Statistics: Simulation and Computation, 46(5):3794–3811, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bakhtiarifar:2017:OPM

[BBA17] Mohammad Hasan Bakhtiarifar, Mahdi Bashiri, and Amirhossein Amiri. Optimization of problems with multivariate multiple functional responses: a case study in air quality. Communications in Statistics: Simulation and Computation, 46(10):8049–8063, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bansal:2013:BEI

[BBF13] Naveen K. Bansal, Madhusudan Bhandary, and Koji Fujiwara. Bayes estimation of intraclass correlation coefficients under unequal family sizes. Communications in Statistics: Simulation and Computation, 42(2):294–302, 2013. CODEN CSSCDB. ISSN 0361-0918.

Boubaker:2017:WEL

[BBK17] H. Boubaker, M. Boutahar, and R. Khalfaoui. Wavelets and estimation of long memory in nonstationary models: Does anything beat the exact local Whittle estimator? Communications in Statistics: Simulation and Computation, 46(2): 1189–1218, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2017:GFT

[BC17] Narayanaswamy Balakrishnan and Ekaterina Chimitova. Goodness-of-fit tests for one-shot device accelerated life testing data. Communications in Statistics: Simulation and Computation, 46(5):3723–3734, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2013:TGF

[BCGV13] N. Balakrishnan, E. Chimitova, N. Galanova, and M. Vedernikova. Testing goodness of fit of parametric AFT and PH models with residuals. Communications in Statistics: Simulation and Computation, 42(6):1352–1367, 2013. CODEN CSSCDB. ISSN 0361-0918.

Bayer:2017:MSC

[BCN17] Fábio M. Bayer and Francisco Cribari-Neto. Model selection criteria in beta regression with varying dispersion. Communications in Statistics: Simulation and Computation, 46(1): 729–746, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2015:EAS

[BCV15] N. Balakrishnan, E. Chimitova, and M. Vedernikova. An empirical analysis of some nonparametric goodness-of-fit tests for censored data. Communications in Statistics: Simulation and Computation, 44(4):1101–1115, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2014:AEP

[BD14a] M. Bentarzi and L. Djeddou. Adaptive estimation of periodic first-order threshold autoregressive model. Communications in Statistics: Simulation and Computation, 43(7): 1611–1630, 2014. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2014:MPV

[BD14b] M. Bentarzi and L. Djeddou. On mixture periodic vector autoregressive models. Communications in Statistics: Simulation and Computation, 43(10):2325–2352, 2014. CODEN CSSCDB. ISSN 0361-0918.

Baratnia:2017:MLS

[BD17a] Mohammad Baratnia and Mahdi Doostparast. Modeling lifetime of sequential r-out-of-n systems with independent and heterogeneous components. Communications in Statistics: Simulation and Computation, 46(9):7365–7375, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bhuyan:2017:RCU

[BD17b] Prajamitra Bhuyan and Anup Dewanji. Reliability computation under dynamic stress-strength modeling with cumulative stress and strength degradation. Communications in Statistics: Simulation and Computation, 46(4):2701–2713, 2017. CODEN CSSCDB. ISSN 0361-0918.

Braschel:2017:SGF

[BDK17] Melissa C. Braschel, Gerarda A. Darlington, and Neil Klar. A study of grouped failure time data and the misapplication of recurrent event modeling. Communications in Statistics: Simulation and Computation, 46(7):5313–5322, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2011:CTG

[BDKM11] N. Balakrishnan, Katherine F. Davies, Jerome P. Keating, and Robert L. Mason. Correlation-type goodness of fit test for extreme value distribution based on simultaneous closeness. Communications in Statistics: Simulation and Computation, 40(7):1074–1095, 2011. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2017:SLD

[BDL17] N. Balakrishnan, Qingzhou Dai, and Kai Liu. A skew logistic distribution as an alternative to the model of van Staden and King. Communications in Statistics: Simulation and Computation, 46(5):4082–4097, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bowman:2010:EGB

[BDS10] Alan Bowman, Necip Doganaksoy, and Josef Schmee. Efficient gradient-based optimization of process capability with multiple, potentially nonnormal outputs. Communications in Statistics: Simulation and Computation, 39(4):807–827, 2010. CODEN CSSCDB. ISSN 0361-0918.

Bee:2015:ELP

[Bee15] Marco Bee. Estimation of the lognormal-Pareto distribution using probability weighted moments and maximum likelihood. Communications in Statistics: Simulation and Computation, 44(8):2040–2060, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bee:2017:DAV

[Bee17] M. Bee. Density approximations and VaR computation for compound Poisson-lognormal distributions. Communications in Statistics: Simulation and Computation, 46(3): 1825–1841, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bang:2017:CKQ

[BEJC17] Sungwan Bang, Soo-Heang Eo, Myoungshic Jhun, and Hyung Jun Cho. Composite kernel quantile regression. Communications in Statistics: Simulation and Computation, 46 (3):2228–2240, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bensic:2014:FDD

[Ben14] Mirta Bensić. Fitting distribution to data by a generalized nonlinear least squares method. Communications in Statistics: Simulation and Computation, 43(4):687–705, 2014. CODEN CSSCDB. ISSN 0361-0918.

Benkhelifa:2017:MOE

[Ben17] Lazhar Benkhelifa. The Marshall–Olkin extended generalized Lindley distribution: Properties and applications. Communications in Statistics: Simulation and Computation, 46 (10):8306–8330, 2017. CODEN CSSCDB. ISSN 0361-0918.

Barbiero:2017:RPS

[BF17] Alessandro Barbiero and Pier Alda Ferrari. An R package for the simulation of correlated discrete variables. Communications in Statistics: Simulation and Computation, 46(7): 5123–5140, 2017. CODEN CSSCDB. ISSN 0361-0918.

Beyaztas:2017:NBB

[BFB17] Beste H. Beyaztas, Esin Firuzan, and Ufuk Beyaztas. New block bootstrap methods: Sufficient and/or ordered. Communications in Statistics: Simulation and Computation, 46 (5):3942–3951, 2017. CODEN CSSCDB. ISSN 0361-0918.

Baghfalaki:2012:EEA

[BG12] T. Baghfalaki and M. Ganjali. An ECM estimation approach for analyzing multivariate skew-normal data with dropout. Communications in Statistics: Simulation and Computation, 41(10):1970–1988, 2012. CODEN CSSCDB. ISSN 0361-0918.

Bargagliotti:2015:CSI

[BG15] Anna E. Bargagliotti and Raymond N. Greenwell. Combinatorics and statistical issues related to the Kruskal-Wallis statistic. Communications in Statistics: Simulation and Computation, 44(2):533–550, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bulinskaya:2016:OCS

[BG16] Ekaterina Bulinskaya and Julia Gusak. Optimal control and sensitivity analysis for two risk models. Communications in Statistics: Simulation and Computation, 45(5):1451–1466, 2016. CODEN CSSCDB. ISSN 0361-0918.

Baghfalaki:2017:RWG

[BG17] Taban Baghfalaki and Mojtaba Ganjali. Robust weighted generalized estimating equations based on statistical depth.

Communications in Statistics: Simulation and Computation, 46(10):8283–8305, 2017. CODEN CSSCDB. ISSN 0361-0918.

Barbosa:2013:RCC

[BGM13] Emanuel Pimentel Barbosa, Mario Antonio Gneri, and Ariane Meneguetti. Range control charts revisited: Simpler Tippett-like formulae, its practical implementation, and the study of false alarm. Communications in Statistics: Simulation and Computation, 42(2):247–262, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Bagnato:2014:SDN

[BGP14] Luca Bagnato, Francesca Greselin, and Antonio Punzo. On the spectral decomposition in normal discriminant analysis. Communications in Statistics: Simulation and Computation, 43(6):1471–1489, 2014. CODEN CSSCDB. ISSN 0361-0918.

Bairamov:2015:MRS

[BGU15] Ismihan Bairamov (Bayramoglu), Selma Gurler, and Burcu Ucer. On the mean remaining strength of the k-out-of-n: F system with exchangeable components. Communications in Statistics: Simulation and Computation, 44(1):1–13, 2015. CODEN CSSCDB. ISSN 0361-0918.

Beauducel:2016:CCF

[BH16] André Beauducel and Norbert Hilger. On the correlation of common factors with variance not accounted for by the factor model. Communications in Statistics: Simulation and Computation, 45(6):2145–2157, 2016. CODEN CSSCDB. ISSN 0361-0918.

Beauducel:2017:BFS

[BH17] André Beauducel and Norbert Hilger. On the bias of factor score determinacy coefficients based on different estimation methods of the exploratory factor model. Communications in Statistics: Simulation and Computation, 46(8): 6144–6154, 2017. CODEN CSSCDB. ISSN 0361-0918.

Baffoun:2017:ESE

[BHF17] Hatem Baffoun, Mekki Hajlaoui, and Abdeljelil Farhat. Equation-solving estimator based on Metropolis–Hastings algorithm with delayed rejection. Communications in Statistics: Simulation and Computation, 46(4):3103–3111, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bagheri:2012:NCI

[BHI12] Arezoo Bagheri, M. Habshah, and R. H. M. R. Imon. A novel collinearity-influential observation diagnostic measure based on a group deletion approach. Communications in Statistics: Simulation and Computation, 41(8):1379–1396, 2012. CODEN CSSCDB. ISSN 0361-0918.

Bialek:2014:SSO

[Bia14] Jacek Bialek. Simulation study of an original price index formula. Communications in Statistics: Simulation and Computation, 43(2):285–297, 2014. CODEN CSSCDB. ISSN 0361-0918.

Bialek:2017:AFPb

[Bia17a] Jacek Bialek. Approximation of the Fisher price index by using Lowe, Young, and AG Mean indices. Communications in Statistics: Simulation and Computation, 46(8):6454–6467, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bialek:2017:AFPa

[Bia17b] Jacek Bialek. Approximation of the Fisher price index by using the Lloyd–Moulton index: Simulation study. Communications in Statistics: Simulation and Computation, 46(5): 3588–3598, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bidram:2012:BEG

[Bid12] Hamid Bidram. The beta exponential-geometric distribution. Communications in Statistics: Simulation and Computation, 41(9):1606–1622, 2012. CODEN CSSCDB. ISSN 0361-0918.

Bang:2014:WSV

[BJ14a] Sungwan Bang and Myoungshic Jhun. Weighted support vector machine using k-means clustering. Communications in Statistics: Simulation and Computation, 43(10):2307– 2324, 2014. CODEN CSSCDB. ISSN 0361-0918.

Bouveyron:2014:AMR

[BJ14b] C. Bouveyron and J. Jacques. Adaptive mixtures of regressions: Improving predictive inference when population has changed. Communications in Statistics: Simulation and Computation, 43(10):2570–2592, 2014. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2018:ODQ

[BJU18] S. Balamurali, P. Jeyadurga, and M. Usha. Optimal designing of quick switching sampling system assuring percentile life under Weibull distribution. Communications in Statistics: Simulation and Computation, 47(5):1550–1563, 2018. CODEN CSSCDB. ISSN 0361-0918.

Banik:2010:CSP

[BK10] Shipra Banik and B. M. Golam Kibria. Comparison of some parametric and nonparametric type one sample confidence intervals for estimating the mean of a positively skewed distribution. Communications in Statistics: Simulation and Computation, 39(2):361–389, 2010. CODEN CSSCDB. ISSN 0361-0918.

Banik:2011:EPC

[BK11] Shipra Banik and B. M. Golam Kibria. Estimating the population coefficient of variation by confidence intervals. Communications in Statistics: Simulation and Computation, 40 (8):1236–1261, 2011. CODEN CSSCDB. ISSN 0361-0918.

Bayoud:2016:TET

[BK16] Husam A. Bayoud and Omar A. Kittaneh. Testing the equality of two exponential distributions. Communications in Statistics: Simulation and Computation, 45(7):2249–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Barakat:2017:TEF

[BK17] H. M. Barakat and O. M. Khaled. Toward the establishment of a family of distributions that may fit any dataset. Communications in Statistics: Simulation and Computation, 46 (8):6129–6143, 2017. CODEN CSSCDB. ISSN 0361-0918.

Babayi:2018:ISS

[BK18] Salman Babayi and Esmaile Khorram. Inference of stressstrength for the Type-II generalized logistic distribution under progressively Type-II censored samples. Communications in Statistics: Simulation and Computation, 47(7): 1975–1995, 2018. CODEN CSSCDB. ISSN 0361-0918.

Bhaumik:2015:SST

[BKB^{+}15] Dulal K. Bhaumik, Kush Kapur, Narayanaswamy Balakrishnan, Jerome P. Keating, and Robert D. Gibbons. Small sample tests for shape parameters of gamma distributions. Communications in Statistics: Simulation and Computation, 44 (5):1339–1363, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bang:2010:SVM

[BKJ10] Sungwan Bang, Ja-Yong Koo, and Myoungshic Jhun. Support vector machine using k-spatial medians clustering and recovery process. Communications in Statistics: Simulation and Computation, 39(7):1422–1434, 2010. CODEN CSS-CDB. ISSN 0361-0918.

Benson-Karhi:2018:MFG

[BKSM18] Diamanta Benson-Karhi, Haim Shore, and Maya Malamud. Modeling fetal-growth biometry with Response Modeling Methodology (RMM) and comparison to current models. Communications in Statistics: Simulation and Computation, 47(1):129–142, 2018. CODEN CSSCDB. ISSN 0361-0918.

Buzkova:2015:TDA

[BL15] Petra Buzková and Thomas Lumley. Time to diagnosis: Accounting for differential endpoint follow-up in multi-cohort studies. Communications in Statistics: Simulation and Computation, 44(1):247–256, 2015. CODEN CSSCDB. ISSN 0361-0918.

Blacher:2014:PRN

[Bla14] René Blacher. Proved random numbers obtained from hardware devices. Communications in Statistics: Simulation and Computation, 43(5):1020–1035, 2014. CODEN CSSCDB. ISSN 0361-0918.

Borlongan:2019:RSC

[BLB19] Martin Augustine B. Borlongan, Joseph Ryan G. Lansangan, and Erniel B. Barrios. Robust simultaneous confidence interval estimation of principal component loadings. Communications in Statistics: Simulation and Computation, 48 (2):431–446, 2019. CODEN CSSCDB. ISSN 0361-0918.

Bar-Lev:2016:NSS

[BLBK16] Shaul K. Bar-Lev, Benzion Boukai, and Igor Kleiner. A numerical study of small parameter behavior of some families of distributions. Communications in Statistics: Simulation and Computation, 45(10):3534–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Brumback:2017:UBW

[BLC17] Babette A. Brumback, Li Li, and Zhuangyu Cai. On the use of between-within models to adjust for confounding due to unmeasured cluster-level covariates. Communications in Statistics: Simulation and Computation, 46(5):3841–3854, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bagdonavicius:2012:MTP

[BLN12] V. Bagdonavicius, R. Levuliene, and M. Nikulin. Modelling and testing of presence of hazard rates crossing under censoring. Communications in Statistics: Simulation and Computation, 41(7):980–991, 2012. CODEN CSSCDB. ISSN 0361-0918.

Bentarzi:2010:ATP

[BM10] M. Bentarzi and M. Merzougui. Adaptive test for periodicity in autoregressive conditional heteroskedastic processes. Communications in Statistics: Simulation and Computation, 39(9):1735–1753, 2010. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2014:SFI

[BM14] N. Balakrishnan and Debanjan Mitra. Some further issues concerning likelihood inference for left truncated and right censored lognormal data. Communications in Statistics: Simulation and Computation, 43(2):400–416, 2014. CODEN CSSCDB. ISSN 0361-0918.

Balakrishna:2017:EFM

[BM17a] N. Balakrishna and P. Anvar Muhammed. Estimating function method for product autoregressive models. Communications in Statistics: Simulation and Computation, 46(5): 3962–3979, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2017:DNS

[BM17b] Saminathan Balamurali and Usha Mahalingam. Developing of a new system of skip-lot reinspection schemes. Communications in Statistics: Simulation and Computation, 46(2): 1481–1491, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bibi:2019:MME

[BM19] Abdelouahab Bibi and Fateh Merahi. Moment method estimation of first-order continuous-time bilinear processes. Communications in Statistics: Simulation and Computation, 48(4):1070–1087, 2019. CODEN CSSCDB. ISSN 0361-0918.

Barco:2017:IPL

[BMJ17] Kelly Vanessa Parede Barco, Josmar Mazucheli, and Vanderly Janeiro. The inverse power Lindley distribution. Communications in Statistics: Simulation and Computation, 46 (8):6308–6323, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bagdonavicius:2012:HTR

[BMN12] Vilijandas Bagdonavicius, Sergey V. Malov, and Mikhail Nikulin. Homogeneity tests for related samples under censoring. Communications in Statistics: Simulation and Computation, 41(6):764–775, 2012. CODEN CSSCDB. ISSN 0361-0918.

Batsidis:2013:NPD

[BMPZ13] Apostolos Batsidis, Nirian Martin, Leandro Pardo, and Konstantinos Zografos. A necessary power divergence type family tests of multivariate normality. Communications in Statistics: Simulation and Computation, 42(10):2253–2271, 2013. CODEN CSSCDB. ISSN 0361-0918.

Barakat:2015:BMO

[BNKM15] H. M. Barakat, E. M. Nigm, O. M. Khaled, and F. A. Momenkhan. Bootstrap method for order statistics and modeling study of the air pollution. Communications in Statistics: Simulation and Computation, 44(6):1477–1491, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Bonnini:2016:MAC

[Bon16] Stefano Bonnini. Multivariate approach for comparative evaluations of customer satisfaction with application to transport services. Communications in Statistics: Simulation and Computation, 45(5):1554–1568, 2016. CODEN CSSCDB. ISSN 0361-0918.

Baghdali-Ourbih:2017:IRD

[BOOTD17] Latifa Baghdali-Ourbih, Megdouda Ourbih-Tari, and Abdelnasser Dahmani. Implementing refined descriptive sampling into three-phase discrete-event simulation models. Communications in Statistics: Simulation and Computation, 46(5): 4035–4049, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bozkurt:2013:MFG

[Boz13] Olcay Bozkurt. Mixtures of Farlie-Gumbel-Morgenstern copulas. Communications in Statistics: Simulation and Computation, 42(1):171–182, 2013. CODEN CSSCDB. ISSN 0361-0918.

Bahari:2017:DRG

[BPG17] Fayyaz Bahari, Safar Parsi, and Mojtaba Ganjali. A doubly robust goodness-of-fit test in general linear models with missing covariates. Communications in Statistics: Simulation and Computation, 46(10):7909–7923, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bekrizadeh:2017:NAC

[BPJ17] H. Bekrizadeh, G. A. Parham, and B. Jamshidi. A new asymmetric class of bivariate copulas for modeling dependence. Communications in Statistics: Simulation and Computation, 46(7):5594–5609, 2017. CODEN CSSCDB. ISSN 0361-0918.

Bertoluzzo:2013:MSP

[BPS13] F. Bertoluzzo, F. Pesarin, and L. Salmaso. On multi-sided permutation tests. Communications in Statistics: Simulation and Computation, 42(6):1380–1390, 2013. CODEN CSSCDB. ISSN 0361-0918.

Bole:2013:BHT

[BR13] Velimir Bole and Peter Rebec. Bootstrapping the Hausman test in panel data models. Communications in Statistics: Simulation and Computation, 42(3):650–670, 2013. CODEN CSSCDB. ISSN 0361-0918.

Balakrishna:2014:IGD

[BR14] N. Balakrishna and T. Rahul. Inverse Gaussian distribution for modeling conditional durations in finance. Communications in Statistics: Simulation and Computation, 43(3): 476–486, 2014. CODEN CSSCDB. ISSN 0361-0918.

Baratpour:2016:ETB

[BR16a] S. Baratpour and A. Habibi Rad. Exponentiality test based on the progressive Type II censoring via cumulative entropy. Communications in Statistics: Simulation and Computation, 45(7):2625–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bulinski:2016:SAA

[BR16b] Alexander V. Bulinski and Alexander S. Rakitko. Simulation and analytical approach to the identification of significant factors. Communications in Statistics: Simulation and Computation, 45(5):1430–1450, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bhat:2017:CGR

[BR17] Satish Bhat and Vidya Raju. A class of generalized ridge estimators. Communications in Statistics: Simulation and Computation, 46(7):5105–5112, 2017. CODEN CSSCDB. ISSN 0361-0918.

Behzad:2018:SAC

[BR18] Ramin Behzad and Mohammad Reza Salehi Rad. Simultaneous arrival of customers to two different queues and modeling dependence via copula approach. Communications in Statistics: Simulation and Computation, 47(10):3118–3131, 2018. CODEN CSSCDB. ISSN 0361-0918.

Bianchi:2017:TSO

[BRF17] Michele Leonardo Bianchi, Svetlozar T. Rachev, and Frank J. Fabozzi. Tempered stable Ornstein–Uhlenbeck processes: a practical view. Communications in Statistics: Simulation and Computation, 46(1):423–445, 2017. CODEN CSSCDB. ISSN 0361-0918.

BarrosoDaSilva:2018:CBR

[BRFD18] Eveliny Barroso Da Silva, Carlos Alberto Ribeiro Diniz, Jalmar Manuel Farfan Carrasco, and Mário De Castro. A class of beta regression models with multiplicative log-normal measurement errors. Communications in Statistics: Simulation and Computation, 47(1):229–248, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Best:2016:CNT

[BRT^{+}16] D. J. Best, J. C. W. Rayner, O. Thas, Jan De Neve, and David Allingham. Comparing nonparametric tests of equality of means for randomized block designs. Communications in Statistics: Simulation and Computation, 45(5): 1718–1730, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bandyopadhyay:2010:BMR

[BS10] Nibedita Bandyopadhyay and Ananda Sen. Bayesian modeling of recurrent event data with dependent censoring. Communications in Statistics: Simulation and Computation, 39 (3):641–654, 2010. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2015:ELIa

[BS15] N. Balakrishnan and Feng Su. Exact likelihood inference for k exponential populations under joint Type-II censoring. Communications in Statistics: Simulation and Computation, 44(3):591–613, 2015. CODEN CSSCDB. ISSN 0361-0918.

Brandner:2016:SIL

[BS16a] Reinhard Brandner and Ernst Stadlober. Samples of iid lognormals: Approximations for characteristics of minima. Communications in Statistics: Simulation and Computation, 45(2):504–518, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bright:2016:EPS

[BS16b] Brianna C. Bright and Julia N. Soulakova. Evaluating the performance of simultaneous stepwise confidence intervals for the difference between two Poisson rates. Communications in Statistics: Simulation and Computation, 45(7): 2338–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2017:RRS

[BS17a] N. Balakrishnan and H. M. Saleh. Recurrence relations for single and product moments of progressively Type-II censored order statistics from generalized logistic distribution with applications to inference. Communications in Statistics: Simulation and Computation, 46(6):4559–4577, 2017. CODEN CSSCDB. ISSN 0361-0918.

Batool:2017:GRD

[BS17b] Fatima Batool and Javid Shabbir. A generic randomization device for quantitative response variables. Communications in Statistics: Simulation and Computation, 46(8): 6399–6413, 2017. CODEN CSSCDB. ISSN 0361-0918.

Berchtold:2017:IRO

[BS17c] André Berchtold and Joan-Carles Surís. Imputation of repeatedly observed multinomial variables in longitudinal surveys. Communications in Statistics: Simulation and Computation, 46(4):3267–3283, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2018:GFT

[BS18] Kirushanthini Balakrishnan and M. R. Sooriyarachchi. A goodness of fit test for multilevel survival data. Communications in Statistics: Simulation and Computation, 47(1): 30–47, 2018. CODEN CSSCDB. ISSN 0361-0918.

Brombin:2010:MAP

[BSFG10] Chiara Brombin, Luigi Salmaso, Giuseppe Ferronato, and Pier-Francesco Galzignato. Multi-aspect procedures for paired data with application to biometric morphing. Communications in Statistics: Simulation and Computation, 40 (1):1–12, 2010. CODEN CSSCDB. ISSN 0361-0918.

Batool:2017:IBR

[BSH17] Fatima Batool, Javid Shabbir, and Zawar Hussain. An improved binary randomized response model using six decks of cards. Communications in Statistics: Simulation and Computation, 46(4):2548–2562, 2017. CODEN CSSCDB. ISSN 0361-0918.

Brusco:2019:RAU

[BSK19] Michael J. Brusco, Douglas Steinley, and Hans-Friedrich Köhn. Residual analysis for unidimensional scaling in the L_{2}-norm. Communications in Statistics: Simulation and Computation, 48(7):2210–2221, 2019. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2015:ELIb

[BSL15] N. Balakrishnan, Feng Su, and Kin-Yat Liu. Exact likelihood inference for k exponential populations under joint progressive Type-II censoring. Communications in Statistics: Simulation and Computation, 44(4):902–923, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Bai:2019:SIM

[BSLF19] Xuchao Bai, Yimin Shi, Bin Liu, and Qianrao Fu. Statistical inference of Marshall–Olkin bivariate Weibull distribution with three shocks based on progressive interval censored data. Communications in Statistics: Simulation and Computation, 48(3):637–654, 2019. CODEN CSSCDB. ISSN 0361-0918.

Baran:2013:ERD

[BSV13] Sándor Baran, Kinga Sikolya, and Lajos Veress. Estimating the risk of a Down’s syndrome term pregnancy using age and serum markers: Comparison of various methods. Communications in Statistics: Simulation and Computation, 42 (7):1654–1672, 2013. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2016:BMT

[BSZ16] N. Balakrishnan, H. Y. So, and X. J. Zhu. On Box–Muller transformation and simulation of normal record data. Communications in Statistics: Simulation and Computation, 45 (10):3670–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bakoyannis:2017:IDL

[BT17] Giorgos Bakoyannis and Giota Touloumi. Impact of dependent left truncation in semiparametric competing risks methods: A simulation study. Communications in Statistics: Simulation and Computation, 46(3):2025–2042, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balamurali:2015:ODV

[BU15] S. Balamurali and M. Usha. Optimal designing of variables sampling plan for resubmitted lots. Communications in Statistics: Simulation and Computation, 44(5):1210–1224, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bush:2015:SSD

[Bus15] Stephen Bush. Sample size determination for logistic regression: A simulation study. Communications in Statistics: Simulation and Computation, 44(2):360–373, 2015. CODEN CSSCDB. ISSN 0361-0918.

Brusco:2019:ILD

[BVC^{+}19] Michael J. Brusco, Clay M. Voorhees, Roger J. Calantone, Michael K. Brady, and Douglas Steinley. Integrating linear discriminant analysis, polynomial basis expansion, and genetic search for two-group classification. Communications in Statistics: Simulation and Computation, 48(6):1623–1636, 2019. CODEN CSSCDB. ISSN 0361-0918.

Bhat:2016:MLE

[BVP16] Satish Bhat, R. Vidya, and V. Pandit Parameshwar. Maximum likelihood estimation of parameters in a mixture model. Communications in Statistics: Simulation and Computation, 45(5):1776–1784, 2016. CODEN CSSCDB. ISSN 0361-0918.

Bizuneh:2019:CDC

[BW19] Berihun Bizuneh and Fu-Kwun Wang. Comparison of different control charts for a Weibull process with type-I censoring. Communications in Statistics: Simulation and Computation, 48(4):1088–1100, 2019. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2017:LIT

[BYAJ17] N. Balakrishnan, Y. Ye, and R. A. Al-Jarallah. Likelihood inference for Type I bivariate Pólya–Aeppli distribution. Communications in Statistics: Simulation and Computation, 46(8):6436–6453, 2017. CODEN CSSCDB. ISSN 0361-0918.

Balakrishnan:2015:IBS

[BZ15] N. Balakrishnan and Xiaojun Zhu. Inference for the Birnbaum–Saunders lifetime regression model with applications. Communications in Statistics: Simulation and Computation, 44(8):2073–2100, 2015. CODEN CSSCDB. ISSN 0361-0918.

Bouadoumou:2015:JEL

[BZL15] Maxime Bouadoumou, Yichuan Zhao, and Yinghua Lu. Jackknife empirical likelihood for the accelerated failure time model with censored data. Communications in Statistics: Simulation and Computation, 44(7):1818–1832, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Bi:2013:MDR

[BZW13] Tao Bi, Bo Zhang, and Huishan Wu. Measuring downside risk using high-frequency data: Realized downside risk measure. Communications in Statistics: Simulation and Computation, 42(4):741–754, 2013. CODEN CSSCDB. ISSN 0361-0918.

Bi:2011:DIS

[BZX11] Tao Bi, Bo Zhang, and Rong Xu. Dynamics of intraday serial correlation in China’s stock market. Communications in Statistics: Simulation and Computation, 40(10):1637–1650, 2011. CODEN CSSCDB. ISSN 0361-0918.

Cuervo:2010:HNR

[CA10] Edilberto Cepeda Cuervo and Jorge Alberto Achcar. Heteroscedastic nonlinear regression models. Communications in Statistics: Simulation and Computation, 39(2):405–419, 2010. CODEN CSSCDB. ISSN 0361-0918.

Cohen:2017:ENM

[CA17] Achraf Cohen and Raid Amin. The effects of normal mixtures and autocorrelation on the fraction non-conforming. Communications in Statistics: Simulation and Computation, 46(10):8105–8117, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cahoy:2013:EML

[Cah13] Dexter O. Cahoy. Estimation of Mittag-Leffler parameters. Communications in Statistics: Simulation and Computation, 42(2):303–315, 2013. CODEN CSSCDB. ISSN 0361-0918.

Calzolari:2019:EER

[Cal19] Giorgio Calzolari. Econometrics exams and round numbers: Use or misuse of indirect estimation methods? Communications in Statistics: Simulation and Computation, 48(1): 1–14, 2019. CODEN CSSCDB. ISSN 0361-0918.

Caron:2019:MAP

[Car19] P.-O. Caron. Minimum average partial correlation and parallel analysis: The influence of oblique structures. Communications in Statistics: Simulation and Computation, 48(7): 2110–2117, 2019. CODEN CSSCDB. ISSN 0361-0918.

Cruz:2014:EPM

[CB14] Rolan Paul K. Veron Cruz and Erniel B. Barrios. Estimation procedure for a multiple time series model. Communications in Statistics: Simulation and Computation, 43(10): 2415–2431, 2014. CODEN CSSCDB. ISSN 0361-0918.

Cirillo:2017:EOG

[CB17] Marcelo Ângelo Cirillo and Lúcia Pereira Barroso. Effect of outliers on the GFI quality adjustment index in structural equation model and proposal of alternative indices. Communications in Statistics: Simulation and Computation, 46 (3):1895–1905, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cambon:2017:PAC

[CBB^{+}17a] A. C. Cambon, K. B. Baumgartner, G. N. Brock, N. G. F. Cooper, D. Wu, and S. N. Rai. Properties of adaptive clinical trial signature design in the presence of gene and genetreatment interaction. Communications in Statistics: Simulation and Computation, 46(10):8233–8250, 2017. CODEN CSSCDB. ISSN 0361-0918.

Canary:2017:CHL

[CBB^{+}17b] Jana D. Canary, Leigh Blizzard, Ronald P. Barry, David W. Hosmer, and Stephen J. Quinn. A comparison of the Hosmer–Lemeshow, Pigeon–Heyse, and Tsiatis goodness-of-fit tests for binary logistic regression under two grouping methods. Communications in Statistics: Simulation and Computation, 46(3):1871–1894, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chandran:2016:UIM

[CBDD16] Avinash Chandran, Derek Brown, Jerome Danoff, and Loretta DiPietro. Using the inverse maximum ratio-Λ as a technique to quantify surface uniformity. Communications in Statistics: Simulation and Computation, 45(3): 1129–1142, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chesneau:2019:NCP

[CBH19] Christophe Chesneau, Hassan S. Bakouch, and Tassaddaq Hussain. A new class of probability distributions via cosine and sine functions with applications. Communications in Statistics: Simulation and Computation, 48(8):2287–2300, 2019. CODEN CSSCDB. ISSN 0361-0918.

Cepeda-Cuervo:2010:GER

[CC10] Edilberto Cepeda-Cuervo. Generalized EGARCH random effect models application to financial time series. Communications in Statistics: Simulation and Computation, 39(8): 1517–1529, 2010. CODEN CSSCDB. ISSN 0361-0918.

Chakrabarty:2019:CIW

[CC19a] Jimut Bahan Chakrabarty and Shovan Chowdhury. Compounded inverse Weibull distributions: Properties, inference and applications. Communications in Statistics: Simulation and Computation, 48(7):2012–2033, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Chen:2019:FTG

[CC19b] Haiqing Chen and Weihu Cheng. Fitting type I generalized logistic distribution by modified method based on percentiles. Communications in Statistics: Simulation and Computation, 48(7):2222–2227, 2019. CODEN CSSCDB. ISSN 0361-0918.

Crevits:2019:REL

[CC19c] Ruben Crevits and Christophe Croux. Robust estimation of linear state space models. Communications in Statistics: Simulation and Computation, 48(6):1694–1705, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Chen:2010:HBB

[CCC10] Henian Chen, Patricia Cohen, and Sophie Chen. How big is a big odds ratio? Interpreting the magnitudes of odds ratios in epidemiological studies. Communications in Statistics: Simulation and Computation, 39(4):860–864, 2010. CODEN CSSCDB. ISSN 0361-0918.

Chang:2016:BVS

[CCC16a] Sheng-Mao Chang, Ray-Bing Chen, and Yunchan Chi. Bayesian variable selections for probit models with componentwise Gibbs samplers. Communications in Statistics: Simulation and Computation, 45(8):2752–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chen:2016:IAN

[CCC16b] Hsin-Hung Chen, Mingchih Chen, and Chun-Cheng Chiu. The integration of artificial neural networks and text mining to forecast gold futures prices. Communications in Statistics: Simulation and Computation, 45(4):1213–1225, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ciampi:2012:MBC

[CCD^{+}12] A. Ciampi, H. Campbell, A. Dyachenko, B. Rich, J. Mc-Cusker, and M. G. Cole. Model-based clustering of longitudinal data: Application to modeling disease course and gene expression trajectories. Communications in Statistics: Simulation and Computation, 41(7):992–1005, 2012. CODEN CSSCDB. ISSN 0361-0918.

Chen:2018:RAC

[CCH18] J. H. Chen, Y. C. Chang, and W. L. Hung. A robust automatic clustering algorithm for probability density functions with application to categorizing color images. Communications in Statistics: Simulation and Computation, 47(7): 2152–2168, 2018. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2010:NAM

[CCL10] Conghua Cheng, Jinyuan Chen, and Zehui Li. A new algorithm for maximum likelihood estimation with progressive Type-I interval censored data. Communications in Statistics: Simulation and Computation, 39(4):750–766, 2010. CODEN CSSCDB. ISSN 0361-0918.

Chen:2014:AMI

[CCLT14] Man-Hua Chen, Li-Ching Chen, Kuen-Hung Lin, and Xingwei Tong. Analysis of multivariate interval censoring by diabetic retinopathy study. Communications in Statistics: Simulation and Computation, 43(7):1825–1835, 2014. CODEN CSSCDB. ISSN 0361-0918.

Chalise:2013:PPV

[CCM13] Prabhakar Chalise, Eric Chicken, and Daniel McGee. Performance and prediction for varying survival time scales. Communications in Statistics: Simulation and Computation, 42 (3):636–649, 2013. CODEN CSSCDB. ISSN 0361-0918.

Carrozzo:2016:NAR

[CCM^{+}16] Eleonora Carrozzo, Livio Corain, Remigio Musci, Luigi Salmaso, and Luca Spadoni. A new approach to rank several multivariate normal populations with application to life cycle assessment. Communications in Statistics: Simulation and Computation, 45(5):1583–1599, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Cao:2018:RRH

[CCRX18] Chunzheng Cao, Mengqian Chen, Yuqian Ren, and Yue Xu. Robust replicated heteroscedastic measurement error model using heavy-tailed distribution. Communications in Statistics: Simulation and Computation, 47(6):1771–1784, 2018. CODEN CSSCDB. ISSN 0361-0918.

Cepeda-Cuervo:2012:DGS

[CCUR12] Edilberto Cepeda-Cuervo, B. Piedad Urdinola, and Diana Rodríguez. Double generalized spatial econometric models. Communications in Statistics: Simulation and Computation, 41(5):671–685, 2012. CODEN CSSCDB. ISSN 0361-0918.

Cui:2017:QLB

[CCY17] Xiaohua Cui, Xia Chen, and Li Yan. Quasi-likelihood bridge estimators for high-dimensional generalized linear models. Communications in Statistics: Simulation and Computation, 46(10):8190–8204, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chang:2013:GCI

[cCyYnH13] Feng chen Chang, Shean ya Yeh, and Hsin neng Hsieh. Generalized confidence interval estimation for the difference in paired areas under the ROC curves in the absence of a gold standard. Communications in Statistics: Simulation and Computation, 42(9):2056–2072, 2013. CODEN CSSCDB. ISSN 0361-0918.

Corkum:2019:NCI

[CCZ19] A. Corkum, P. Cabilio, and Y. Zhang. Nonparametric confidence intervals for location in time series data. Communications in Statistics: Simulation and Computation, 48(5): 1292–1311, 2019. CODEN CSSCDB. ISSN 0361-0918.

Chen:2017:PEG

[CCZZ17] Haiqing Chen, Weihu Cheng, Jing Zhao, and Xu Zhao. Parameter estimation for generalized Pareto distribution by generalized probability weighted moment-equations. Communications in Statistics: Simulation and Computation, 46 (10):7761–7776, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cai:2012:SBM

[CD12] Yuzhi Cai and Neville Davies. A simple bootstrap method for time series. Communications in Statistics: Simulation and Computation, 41(5):621–631, 2012. CODEN CSSCDB. ISSN 0361-0918.

Cui:2014:PAP

[CD14] Lirong Cui and Tadashi Dohi. Preface to The 5th Asia– Pacific International Symposium on Advanced Reliability and Maintenance Modeling (APARM 2012). Communications in Statistics: Simulation and Computation, 43(8):1837, 2014. CODEN CSSCDB. ISSN 0361-0918.

Cramer:2018:ROP

[CD18] E. Cramer and K. Davies. Restricted optimal progressive censoring. Communications in Statistics: Simulation and Computation, 47(4):1216–1239, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chatrabgoun:2019:OMI

[CDE19] O. Chatrabgoun, A. Daneshkhah, and M. Esmaeilbeigi. Optimizing minimum information pair-copula using genetic algorithm to select optimal basis functions. Communications in Statistics: Simulation and Computation, 48(2):494–505, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ciampi:2012:TWC

[CDGML12] Antonio Ciampi, Alina Dyachenko, Ana Gonzalez-Marcos, and Yves Lechevallier. Two-way classification of a data table with non negative entries: The role of the χ^{2} distance and correspondence analysis. Communications in Statistics: Simulation and Computation, 41(7):1006–1022, 2012. CO-DEN CSSCDB. ISSN 0361-0918.

Cai:2014:CFS

[CDkP14] Yuzhi Cai, Neville Davies, and Wan kai Pang. Combining forecasts via simulations. Communications in Statistics: Simulation and Computation, 43(6):1318–1332, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Chaubey:2012:AKE

[CDL12] Yogendra P. Chaubey, Isha Dewan, and Jun Li. An asymmetric kernel estimator of density function for stationary associated sequences. Communications in Statistics: Simulation and Computation, 41(4):554–572, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Coudret:2015:CKD

[CDS15] Raphaël Coudret, Gilles Durrieu, and Jérôme Saracco. Comparison of kernel density estimators with assumption on number of modes. Communications in Statistics: Simulation and Computation, 44(1):196–216, 2015. CODEN CSSCDB. ISSN 0361-0918.

Christen:2017:ODG

[CFSC17] J. Andrés Christen, Colin Fox, and Mario Santana-Cibrian. Optimal direction Gibbs sampler for truncated multivariate normal distributions. Communications in Statistics: Simulation and Computation, 46(4):2587–2600, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chen:2017:SCS

[CGST17] Huifen Chen, David Goldsman, Bruce W. Schmeiser, and Kwok-Leung Tsui. Symmetric X̄-charts: Sensitivity to nonnormality and control-limit estimation. Communications in Statistics: Simulation and Computation, 46(1):358–378, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cai:2010:MSS

[CH10] Yuzhi Cai and Dominic Hames. Minimum sample size determination for generalized extreme value distribution. Communications in Statistics: Simulation and Computation, 40 (1):87–98, 2010. CODEN CSSCDB. ISSN 0361-0918.

Chen:2011:LRT

[CH11] Shun-Yi Chen and Ching-Feng Hsu. Likelihood ratio tests for the non equivalence of means. Communications in Statistics: Simulation and Computation, 40(6):818–838, 2011. CODEN CSSCDB. ISSN 0361-0918.

Chen:2017:SPS

[CH17] Ding-Geng (Din) Chen and Shuyen Ho. From statistical power to statistical assurance: It’s time for a paradigm change in clinical trial design. Communications in Statistics: Simulation and Computation, 46(10):7957–7971, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chang:2012:MSC

[Cha12] Youngjae Chang. Multi-step classification trees. Communications in Statistics: Simulation and Computation, 41(9): 1728–1744, 2012. CODEN CSSCDB. ISSN 0361-0918.

Chang:2013:VSR

[Cha13] Youngjae Chang. Variable selection via regression trees in the presence of irrelevant variables. Communications in Statistics: Simulation and Computation, 42(8):1703–1726, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chang:2018:PPC

[Cha18] Xinfeng Chang. Performance of the principal component two-parameter estimator in misspecified linear regression model. Communications in Statistics: Simulation and Computation, 47(7):2070–2082, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chakraborty:2017:GWM

[CHB17] N. Chakraborty, S. W. Human, and N. Balakrishnan. A generally weighted moving average chart for time between events. Communications in Statistics: Simulation and Computation, 46(10):7790–7817, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cheong:2011:UMV

[Che11] Chin Wen Cheong. Univariate and multivariate value-at-risk: Application and implication in energy markets. Communications in Statistics: Simulation and Computation, 40 (7):957–977, 2011. CODEN CSSCDB. ISSN 0361-0918.

Cheikh:2014:CSR

[Che14] Malika Cheikh. Comparative study of robust estimators based on a sensitivity coefficient in principal component analysis. Communications in Statistics: Simulation and Computation, 43(10):2639–2648, 2014. CODEN CSSCDB. ISSN 0361-0918.

Chen:2015:MCP

[Che15] Zhanshou Chen. Monitoring change in persistence against the null of difference-stationarity in infinite variance observations. Communications in Statistics: Simulation and Computation, 44(1):71–87, 2015. CODEN CSSCDB. ISSN 0361-

Chen:2017:NAD

[Che17] Zhongxue Chen. A nonparametric approach to detecting the difference of survival medians. Communications in Statistics: Simulation and Computation, 46(1):395–403, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chiu:2015:ESD

[Chi15] Wen-Chih Chiu. Economic-statistical design of EWMA control charts based on Taguchi’s loss function. Communications in Statistics: Simulation and Computation, 44(1):137– 153, 2015. CODEN CSSCDB. ISSN 0361-0918.

Cao:2017:LBB

[CHJL17] Yongxiu Cao, Jian Huang, Yuling Jiao, and Yanyan Liu. A lower bound based smoothed quasi-Newton algorithm for group bridge penalized regression. Communications in Statistics: Simulation and Computation, 46(6):4694–4707, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chvostekova:2017:CBP

[Chv17] Martina Chvosteková. Confidence band for percentile line in multiple linear regression. Communications in Statistics: Simulation and Computation, 46(5):3933–3941, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Casanova:2010:BSA

[CIB10] Maria Paz Casanova, Pilar Iglesias, and Heleno Bolfarine. A Bayesian semiparametric approach for solving the discrete calibration problem. Communications in Statistics: Simulation and Computation, 39(2):347–360, 2010. CODEN CSS-CDB. ISSN 0361-0918.

Chang:2017:MCC

[cIC17] Yuan chin Ivan Chang. Multiple-class classification: Ordinal and categorical labels. Communications in Statistics: Simulation and Computation, 46(10):7561–7581, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chiang:2017:CCG

[CJB^{+}17] Jyun-You Chiang, Nan Jiang, Trenton N. Brown, Tzong-Ru Tsai, and Y. L. Lio. Control charts for generalized exponential distribution percentiles. Communications in Statistics: Simulation and Computation, 46(10):7827–7843, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Campillo:2018:PIS

[CJLV18] Fabien Campillo, Marc Joannides, and Irène Larramendy-Valverde. Parameter identification for a stochastic logistic growth model with extinction. Communications in Statistics: Simulation and Computation, 47(3):721–737, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chiang:2018:IBX

[CJTL18] Jyun-You Chiang, Nan Jiang, Tzong-Ru Tsai, and Y. L. Lio. Inference of δ = P (X < Y) for Burr XII distributions with record samples. Communications in Statistics: Simulation and Computation, 47(3):822–838, 2018. CODEN CSSCDB. ISSN 0361-0918.

Crabbe:2013:CTS

[CJV13] Marjolein Crabbe, Bradley Jones, and Martina Vandebroek. Comparing two-stage segmentation methods for choice data with a one-stage latent class choice analysis. Communications in Statistics: Simulation and Computation, 42(5): 1188–1212, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chu:2014:TML

[CK14] Man-Kee M. Chu and John J. Koval. Trajectory modeling of longitudinal binary data: Application of the EM algorithm for mixture models. Communications in Statistics: Simulation and Computation, 43(3):495–519, 2014. CODEN CSSCDB. ISSN 0361-0918.

Cuny:2016:MWF

[CK16] Henri E. Cuny and Tristan Senga Kiessé. On modeling wood formation using parametric and semiparametric regressions for count data. Communications in Statistics: Simulation and Computation, 45(5):1748–1762, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Casement:2018:GPE

[CK18] Christopher J. Casement and David J. Kahle. Graphical prior elicitation in univariate models. Communications in Statistics: Simulation and Computation, 47(10):2906–2924, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chaturvedi:2019:RBA

[CK19] Ajit Chaturvedi and Taruna Kumari. Robust Bayesian analysis of generalized inverted family of distributions. Communications in Statistics: Simulation and Computation, 48(8): 2244–2268, 2019. CODEN CSSCDB. ISSN 0361-0918.

Choi:2014:VSN

[CKC^{+}14] Byoung-Jeong Choi, Kwang-Rae Kim, Kyu-Dong Cho, Changyi Park, and Ja-Yong Koo. Variable selection for naive Bayes semisupervised learning. Communications in Statistics: Simulation and Computation, 43(10):2702–2713, 2014. CODEN CSSCDB. ISSN 0361-0918.

Choi:2017:RSV

[CKKP17] Hosik Choi, Yongdai Kim, Sunghoon Kwon, and Changyi Park. A robust support vector machine for labeling errors. Communications in Statistics: Simulation and Computation, 46(8):6061–6073, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chatterjee:2017:AUT

[CKP17] K. Chatterjee, C. Koukouvinos, and C. Parpoula. On the analysis of unbalanced two-level supersaturated designs via generalized linear models. Communications in Statistics: Simulation and Computation, 46(5):3383–3395, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Craciun:2013:SMB

[CKPR13] Gheorghe Craciun, Jaejik Kim, Casian Pantea, and Grzegorz A. Rempala. Statistical model for biochemical network inference. Communications in Statistics: Simulation and Computation, 42(1):121–137, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chen:2015:EBD

[CL15] Yuh-Ing Chen and Wen-Ming Lin. Estimation of the benchmark dose for right-censored lifetime data. Communications in Statistics: Simulation and Computation, 44(6): 1514–1531, 2015. CODEN CSSCDB. ISSN 0361-0918.

Chen:2018:BHJ

[CL18] Geng Chen and Sheng Luo. Bayesian hierarchical joint modeling using skew-normal/independent distributions. Communications in Statistics: Simulation and Computation, 47 (5):1420–1438, 2018. CODEN CSSCDB. ISSN 0361-0918.

Castellares:2022:CEG

[CL22] Fredy Castellares and Artur J. Lemonte. Comments on “An extended Gompertz–Makeham distribution with application to lifetime data”. Communications in Statistics: Simulation and Computation, 51(5):2769–2777, 2022. CODEN CSS-CDB. ISSN 0361-0918. See [Abd18].

Chantarangsi:2018:NPP

[CLB^{+}18] W. Chantarangsi, W. Liu, F. Bretz, S. Kiatsupaibul, and A. J. Hayter. Normal probability plots with confidence for the residuals in linear regression. Communications in Statistics: Simulation and Computation, 47(2):367–379, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chen:2013:LCL

[CLJ13] Ding-Geng (Din) Chen, Y. L. Lio, and Nan Jiang. Lower confidence limits on the generalized exponential distribution percentiles under progressive Type-I interval censoring. Communications in Statistics: Simulation and Computation, 42(9):2106–2117, 2013. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2016:MSR

[CLLH16] Sen Lin Cheng, Buqing Liu, Lei Li, and Jiansong Huang. MLE solution research for a certain kind of two-dimensional lognormal distribution function. Communications in Statistics: Simulation and Computation, 45(7):2459–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Cancho:2013:PSC

[CLO13] Vicente G. Cancho, Francisco Louzada, and Edwin M. Ortega. The power series cure rate model: An application to a cutaneous melanoma data. Communications in Statistics: Simulation and Computation, 42(3):586–602, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chin:2017:HMH

[CLT17] Wen Cheong Chin, Min Cherng Lee, and Pei Pei Tan. Heterogeneous market hypothesis evaluation using multipower variation volatility. Communications in Statistics: Simulation and Computation, 46(8):6574–6587, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chang:2012:ILI

[CLW12] F. Chang, T. Lei, and A. C. M. Wong. Improved likelihood inference on testing the difference of non centrality parameters of two independent non central t distributions with identical degrees of freedom. Communications in Statistics: Simulation and Computation, 41(3):342–354, 2012. CODEN CSSCDB. ISSN 0361-0918.

Cui:2019:PSE

[CLY19] Lirong Cui, Zhongping Li, and He Yi. Partial self-exciting point processes and their parameter estimations. Communications in Statistics: Simulation and Computation, 48(10): 2913–2935, 2019. CODEN CSSCDB. ISSN 0361-0918.

Chi:2013:TSH

[CM13] Yueh-Yun Chi and Keith E. Muller. Two-step hypothesis testing when the number of variables exceeds the sample size. Communications in Statistics: Simulation and Computation, 42(5):1113–1125, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chukova:2015:PAO

[CM15] Stefanka Chukova and Leda D. Minkova. Pólya–Aeppli of order k risk model. Communications in Statistics: Simulation and Computation, 44(3):551–564, 2015. CODEN CSSCDB. ISSN 0361-0918.

Corlu:2016:EPG

[CM16] Canan G. Corlu and Melike Meterelliyoz. Estimating the parameters of the generalized Lambda distribution: Which method performs best? Communications in Statistics: Simulation and Computation, 45(7):2276–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chukova:2019:NHP

[CM19] Stefanka Chukova and Leda D. Minkova. Non-homogeneous Pólya–Aeppli process. Communications in Statistics: Simulation and Computation, 48(10):2955–2967, 2019. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2011:VSI

[CMB11] Zhi-Qiang Cheng, Yi-Zhong Ma, and Jing Bu. Variance shifts identification model of bivariate process based on LS-SVM pattern recognizer. Communications in Statistics:

Simulation and Computation, 40(2):274–284, 2011. CODEN CSSCDB. ISSN 0361-0918.

Cordeiro:2018:KNL

[CMBS18] Gauss M. Cordeiro, Elizabete C. Machado, Denise A. Botter, and Mônica C. Sandoval. The Kumaraswamy normal linear regression model with applications. Communications in Statistics: Simulation and Computation, 47(10):3062–3082, 2018. CODEN CSSCDB. ISSN 0361-0918.

Cassarly:2017:ATE

[CMC^{+}17] Christy Cassarly, Renee’ H. Martin, Marc Chimowitz, Edsel A. Peña, Viswanathan Ramakrishnan, and Yuko Y. Palesch. Assessing Type I error and power of multistate Markov models for panel data — a simulation study. Communications in Statistics: Simulation and Computation, 46 (9):7040–7061, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chakraborti:2012:SEL

[CMM12] S. Chakraborti, G. Michaelson, and A. K. McCracken. Shortest expected length confidence interval for the power of the t-test. Communications in Statistics: Simulation and Computation, 41(8):1336–1345, 2012. CODEN CSSCDB. ISSN 0361-0918.

Caudill:2014:NDE

[CMM14] Steven B. Caudill, Franklin G. Mixon,, Jr., and C. Paul Mixon. NFL draftnikology: Euclidean metrics and other approaches to scoring ranking predictions. Communications in Statistics: Simulation and Computation, 43(2):237–248, 2014. CODEN CSSCDB. ISSN 0361-0918.

Chowdhury:2017:CGD

[CMN17] Shovan Chowdhury, Amitava Mukherjee, and Asok K. Nanda. On compounded geometric distributions and their applications. Communications in Statistics: Simulation and Computation, 46(3):1715–1734, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cuddington:2011:CAS

[CN11] John T. Cuddington and William Navidi. A critical assessment of simulated critical values. Communications in Statistics: Simulation and Computation, 40(5):719–727, 2011. CODEN CSSCDB. ISSN 0361-0918.

Chen:2019:IAL

[CN19] Ying-Ju Chen and Wei Ning. Information approach for a lifetime change-point model based on the exponential– logarithmic distribution. Communications in Statistics: Simulation and Computation, 48(7):1996–2003, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Chan:2016:RPV

[CNA16] Stephen Chan, Saralees Nadarajah, and Emmanuel Afuecheta. An R package for value at risk and expected shortfall. Communications in Statistics: Simulation and Computation, 45 (9):3416–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Choi:2017:HBM

[CNK17] Seongmi Choi, Balgobin Nandram, and Dalho Kim. A hierarchical Bayesian model for binary data incorporating selection bias. Communications in Statistics: Simulation and Computation, 46(6):4767–4782, 2017. CODEN CSSCDB. ISSN 0361-0918.

Costantini:2015:SSI

[CNPW15] Mauro Costantini, Paresh Narayan, Stephan Popp, and Joakim Westerlund. Small-sample improved seasonal unit root tests for trending and breaking series. Communications in Statistics: Simulation and Computation, 44(4):868–877, 2015. CODEN CSSCDB. ISSN 0361-0918.

Chen:2011:DON

[CP11] Sheau-Chiann Chen and Jeh-Nan Pan. Determining optimal number of samples for constructing multivariate control charts. Communications in Statistics: Simulation and Computation, 40(2):216–228, 2011. CODEN CSSCDB. ISSN 0361-0918.

Chatrabgoun:2016:CDE

[CP16] O. Chatrabgoun and G. Parham. Copula density estimation using multiwavelets based on the multiresolution analysis. Communications in Statistics: Simulation and Computation, 45(9):3350–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chang:2010:NCS

[CPL10] Ching-Hui Chang, Nabendu Pal, and Jyh-Jiuan Lin. A note on comparing several Poisson means. Communications in Statistics: Simulation and Computation, 39(8):1605–1627, 2010. CODEN CSSCDB. ISSN 0361-0918.

Chang:2017:RTE

[CPL17] Ching-Hui Chang, Nabendu Pal, and Jyh-Jiuan Lin. A revisit to test the equality of variances of several populations. Communications in Statistics: Simulation and Computation, 46(8):6360–6384, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cruz:2017:BET

[CQH17] F. R. B. Cruz, R. C. Quinino, and L. L. Ho. Bayesian estimation of traffic intensity based on queue length in a multiserver M/M/s queue. Communications in Statistics: Simulation and Computation, 46(9):7319–7331, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chachi:2017:FRR

[CR17] Jalal Chachi and Mahdi Roozbeh. A fuzzy robust regression approach applied to bedload transport data. Communications in Statistics: Simulation and Computation, 46(3): 1703–1714, 2017. CODEN CSSCDB. ISSN 0361-0918.

Corain:2013:PAS

[CRS13] Livio Corain, Susanna Ragazzi, and Luigi Salmaso. A permutation approach to split-plot experiments. Communications in Statistics: Simulation and Computation, 42(6): 1391–1408, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chandra:2012:SAE

[CS12] Hukum Chandra and U. C. Sud. Small area estimation for zero-inflated data. Communications in Statistics: Simulation and Computation, 41(5):632–643, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Crudu:2017:FSP

[CS17a] Federico Crudu and Zsolt Sándor. On the finite-sample properties of conditional empirical likelihood estimators. Communications in Statistics: Simulation and Computation, 46 (2):1520–1545, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cugnata:2017:CAI

[CS17b] Federica Cugnata and Silvia Salini. Comparison of alternative imputation methods for ordinal data. Communications in Statistics: Simulation and Computation, 46(1):315–330, 2017. CODEN CSSCDB. ISSN 0361-0918.

Celik:2019:TWA

[CS19] Nuri Celik and Birdal Senoglu. Two-way ANOVA when the distribution of the error terms is skew t. Communications in Statistics: Simulation and Computation, 48(1):287–301, 2019. CODEN CSSCDB. ISSN 0361-0918.

Chouia:2017:MCS

[CSA17] Sana Chouia and Nacira Seddik-Ameur. A modified chisquare test for Bertholon model with censored data. Communications in Statistics: Simulation and Computation, 46 (1):593–602, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chandra:2015:SAE

[CSG15] Hukum Chandra, U. C. Sud, and Yogita Gharde. Small area estimation using estimated population level auxiliary data. Communications in Statistics: Simulation and Computation, 44(5):1197–1209, 2015. CODEN CSSCDB. ISSN 0361-0918.

Chauhan:2016:CSV

[CSK16] Rajvir Singh Chauhan, Parminder Singh, and Narinder Kumar. Comparing several variances with a control in direction-mixed families of hypotheses. Communications in Statistics: Simulation and Computation, 45(7):2297–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chen:2017:BEH

[CSL17] Chen-Yueh Chen, Khalil Shafie, and Yen-Kuang Lin. Bayesian estimation of the Hurst parameter of fractional Brownian motion. Communications in Statistics: Simulation and Computation, 46(6):4760–4766, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chaubey:2013:STS

[CSS13] Yogendra P. Chaubey, Murari Singh, and Debaraj Sen. On symmetrizing transformation of the sample coefficient of variation from a normal population. Communications in Statistics: Simulation and Computation, 42(9):2118–2134, 2013. CODEN CSSCDB. ISSN 0361-0918.

Cai:2017:ALT

[CSY17] Jing Cai, Yimin Shi, and Hongbo Yue. Accelerated life tests for log-normal series system with dependent masked data under Type-I progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 46(2): 1628–1646, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chen:2012:PEP

[CT12] Sy-Mien Chen and Ta-Jung Tsai. Properties of estimators of the process capability index. Communications in Statistics: Simulation and Computation, 41(8):1444–1462, 2012. CODEN CSSCDB. ISSN 0361-0918.

Cortes-Toto:2017:TSA

[CTGR17] Daniela Cortés-Toto, Víctor M. Guerrero, and Hortensia J. Reyes. Trend smoothness achieved by penalized least squares with the smoothing parameter chosen by optimality criteria. Communications in Statistics: Simulation and Computation, 46(2):1492–1507, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chen:2017:IPN

[CTH17] Ding-Geng (Din) Chen, Naitee Ting, and Shuyen Ho. Informative priors or noninformative priors? A Bayesian reanalysis of binary data from Macugen phase III clinical trials. Communications in Statistics: Simulation and Computation, 46(6):4535–4546, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chi:2016:CSC

[CTTT16] Yunchan Chi, Tsung-Hsien Tsai, Yi-Hsuan Tu, and Wen-Yan Tsai. Comparison of several confidence intervals for median residual lifetime with left-truncated and right-censored data. Communications in Statistics: Simulation and Computation, 45(2):701–716, 2016. CODEN CSSCDB. ISSN 0361-0918.

Cai:2017:WAR

[CTXW17] Xia Cai, Yubin Tian, Houbao Xu, and Jun Wang. WCF approach of reliability assessment for solid state power controller with accelerate degradation data. Communications in Statistics: Simulation and Computation, 46(1):458–468, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chigbu:2019:EOC

[CU19] Polycarp E. Chigbu and Eugene C. Ukaegbu. The efficiency and optimality characterization of certain six-treatment incomplete-block designs emanating from some quasi-semi-Latin squares. Communications in Statistics: Simulation and Computation, 48(5):1406–1428, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Cheng:2012:EBX

[CW12a] Yung-Fu Cheng and Fu-Kwun Wang. Estimating the Burr XII parameters in constant-stress partially accelerated life tests under multiple censored data. Communications in Statistics: Simulation and Computation, 41(9):1711–1727, 2012. CODEN CSSCDB. ISSN 0361-0918.

Chiang:2012:AOD

[CW12b] Min-Hsien Chiang and Li-Min Wang. Additive outlier detection and estimation for the logarithmic autoregressive conditional duration model. Communications in Statistics: Simulation and Computation, 41(3):287–301, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Cui:2014:PTA

[CW14] Yunwei Cui and Rongning Wu. A portmanteau test for ARMA processes with infinite variance. Communications in Statistics: Simulation and Computation, 43(3):597–614, 2014. CODEN CSSCDB. ISSN 0361-0918.

Choi:2015:PLM

[CW15] Taeryon Choi and Yoonsung Woo. A partially linear model using a Gaussian process prior. Communications in Statistics: Simulation and Computation, 44(7):1770–1786, 2015. CODEN CSSCDB. ISSN 0361-0918.

Chen:2016:TME

[CWSW16] Hubert J. Chen, Shu-Fei Wu, Pei-Chi Shao, and Wen-Gine Wang. On testing the mean equivalence of treatments from correlated normal populations. Communications in Statistics: Simulation and Computation, 45(6):2104–2121, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chen:2012:NMA

[CWYF12] Jianwei Chen, Xiaofei Wang, Dai Yang, and Boming Fan. Nonparametric modeling auxiliary covariates in random coefficient models. Communications in Statistics: Simulation and Computation, 41(8):1271–1281, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Cheng:2017:PEC

[CX17] Maolin Cheng and Minyin Xiang. Parameter estimation of a composite production function model based on improved artificial fish swarm algorithm and model application. Communications in Statistics: Simulation and Computation, 46 (10):8218–8232, 2017. CODEN CSSCDB. ISSN 0361-0918.

Chan:2018:SMU

[CXC18] Priscilla H. Chan, Ronghui Xu, and Christina D. Chambers. A study of R^{2} measure under the accelerated failure time models. Communications in Statistics: Simulation and Computation, 47(2):380–391, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chen:2016:MML

[CXW16] Wangxue Chen, Minyu Xie, and Ming Wu. Modified maximum likelihood estimator of scale parameter using moving extremes ranked set sampling. Communications in Statistics: Simulation and Computation, 45(6):2232–2240, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chang:2012:NNT

[CY12] Chia-Hao Chang and Ching-Ho Yen. A new nonparametric test for the umbrella alternative problem. Communications in Statistics: Simulation and Computation, 41(8): 1478–1488, 2012. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2016:ASS

[CY16] Yu-Ting Cheng and Chih-Ching Yang. An approach of stocks substitution strategy using fuzzy interval correlation coefficient. Communications in Statistics: Simulation and Computation, 45(4):1187–1196, 2016. CODEN CSSCDB. ISSN 0361-0918.

Chen:2017:CLT

[CY17] Man-Hua Chen and Hsin-Yu Yang. A class of linear transformation models for interval censored data with a cured proportion. Communications in Statistics: Simulation and Computation, 46(10):8251–8263, 2017. CODEN CSSCDB. ISSN 0361-0918.

Cheng:2013:OVS

[CYW13] Yu-Ting Cheng, Chih-Ching Yang, and Sou-Chein Wu. Optimal variable sample size and sampling interval MSE chart based on improved square root transformation. Communications in Statistics: Simulation and Computation, 42(3): 534–551, 2013. CODEN CSSCDB. ISSN 0361-0918.

Chen:2018:NMC

[CYY18] Jiaqi Chen, Hualong Yang, and Jianfeng Yao. A new multivariate CUSUM chart using principal components with a revision of Crosier’s chart. Communications in Statistics: Simulation and Computation, 47(2):464–476, 2018. CODEN CSSCDB. ISSN 0361-0918.

Chaji:2019:ESE

[CZ19] Alireza Chaji and Konstantinos Zografos. An estimator of Shannon entropy of beta-generated distributions and a goodness-of-fit test. Communications in Statistics: Simulation and Computation, 48(2):396–415, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Chen:2016:ESC

[CZL16] Henian Chen, Nanhua Zhang, and Xiaosun Lu. Is early smoking a causal risk factor for later cognitive impairment? A20-year prospective study with time-varying propensity score matching based on random intercept and slope. Communications in Statistics: Simulation and Computation, 45 (8):2733–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dogru:2018:RMR

[DA18] Fatma Zehra Dogru and Olcay Arslan. Robust mixture regression modeling using the least trimmed squares (LTS)- estimation method. Communications in Statistics: Simulation and Computation, 47(7):2184–2196, 2018. CODEN CSSCDB. ISSN 0361-0918.

Doostparast:2013:BEB

[DAA13] Mahdi Doostparast, Mohammad Vali Ahmadi, and Jafar Ahmadi. Bayes estimation based on joint progressive Type II censored data under LINEX loss function. Communications in Statistics: Simulation and Computation, 42(8): 1865–1886, 2013. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2014:BRP

[DAD14] Hakan Demirtas, Anup Amatya, and Beyza Doganay. Bin-Nor: An r package for concurrent generation of binary and normal data. Communications in Statistics: Simulation and Computation, 43(3):569–579, 2014. CODEN CSSCDB. ISSN 0361-0918.

Dag:2014:MIB

[DAI14] Osman Dag, Ozgur Asar, and Ozlem Ilk. A methodology to implement Box–Cox transformation when no covariate is available. Communications in Statistics: Simulation and Computation, 43(7):1740–1759, 2014. CODEN CSSCDB. ISSN 0361-0918.

deAndrade:2017:TGM

[dALA17] Breno Silveira de Andrade, Jacek Leskow, and Marinho G. Andrade. Transformed GARMA model: Properties and simulations. Communications in Statistics: Simulation and Computation, 46(9):7166–7179, 2017. CODEN CSSCDB. ISSN 0361-0918.

Das:2015:RDG

[DAM15] I. Das, M. Aggarwal, and S. Mukhopadhyay. Robust designs in generalized linear models: A quantile dispersion graphs approach. Communications in Statistics: Simulation and Computation, 44(9):2348–2370, 2015. CODEN CSSCDB. ISSN 0361-0918.

Darkhovskiy:2012:NAM

[Dar12] Boris Darkhovskiy. Non asymptotic minimax estimation of functionals with noisy observations. Communications in Statistics: Simulation and Computation, 41(6):787–803, 2012. CODEN CSSCDB. ISSN 0361-0918.

Dadashova:2017:SBM

[DARMM^{+}17] Bahar Dadashova, Blanca Arenas-Ramírez, José Mira-Mcwilliams, Camino González-Fernández, and Francisco Aparicio-Izquierdo. Simulation-based model comparison methodology with application to road accident models. Communications in Statistics: Simulation and Computation, 46(7):5340–5366, 2017. CODEN CSSCDB. ISSN 0361-

Dasgupta:2017:GTC

[Das17] Ratan Dasgupta. Growth of tuber crops and almost sure band for quantiles. Communications in Statistics: Simulation and Computation, 46(2):1062–1085, 2017. CODEN CSSCDB. ISSN 0361-0918.

Dazi:2019:RBS

[DAY19] G. Dazi, M. Azimmohseni, and F. Yaghmaie. Rank-based statistics for testing the whiteness hypothesis of time series. Communications in Statistics: Simulation and Computation, 48(8):2452–2466, 2019. CODEN CSSCDB. ISSN 0361-0918.

Denhere:2016:RPC

[DB16] Melody Denhere and Nedret Billor. Robust principal component functional logistic regression. Communications in Statistics: Simulation and Computation, 45(1):264–281, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dutta:2017:EQE

[DB17] Santanu Dutta and Suparna Biswas. Extreme quantile estimation based on financial time series. Communications in Statistics: Simulation and Computation, 46(6):4226–4243, 2017. CODEN CSSCDB. ISSN 0361-0918.

Dutta:2018:NEE

[DB18] Santanu Dutta and Suparna Biswas. Nonparametric estimation of 100(1 − p) % expected shortfall: p → 0 as sample size is increased. Communications in Statistics: Simulation and Computation, 47(2):338–352, 2018. CODEN CSSCDB. ISSN 0361-0918.

Dalla:2017:CFB

[DBM17] Violetta Dalla, Yannis Bassiakos, and Simos G. Meintanis. Characteristic function-based inference for GARCH models with heavy-tailed innovations. Communications in Statistics: Simulation and Computation, 46(4):2733–2755, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mattos:2016:MSO

[dBMdSF16] Thalita do Bem Mattos and Clécio da Silva Ferreira. The mean-shift outlier model under skew normal distribution.

Communications in Statistics: Simulation and Computation, 45(6):1905–1917, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dovoedo:2015:BBO

[DC15] Y. H. Dovoedo and S. Chakraborti. Boxplot-based outlier detection for the location-scale family. Communications in Statistics: Simulation and Computation, 44(6):1492–1513, 2015. CODEN CSSCDB. ISSN 0361-0918.

DePinho:2017:CVF

[DC17] Frank M. De Pinho and Ricardo F. Couto. Comparing volatility forecasting models during the global financial crisis. Communications in Statistics: Simulation and Computation, 46(7):5257–5270, 2017. CODEN CSSCDB. ISSN 0361-0918.

Dumont:2016:TSA

[DCM16] Cyrielle Dumont, Marylore Chenel, and France Mentré. Two-stage adaptive designs in nonlinear mixed effects models: Application to pharmacokinetics in children. Communications in Statistics: Simulation and Computation, 45(5): 1511–1525, 2016. CODEN CSSCDB. ISSN 0361-0918.

Drignei:2016:UAF

[DD16] Dorin Drignei and Mihaela C. Drignei. Uncertainty analysis for functional computer output: A simulation study in inverse problems. Communications in Statistics: Simulation and Computation, 45(9):3281–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dolgun:2017:PNM

[DD17] Anil Dolgun and Haydar Demirhan. Performance of nonparametric multiple comparison tests under heteroscedasticity, dependency, and skewed error distribution. Communications in Statistics: Simulation and Computation, 46(7): 5166–5183, 2017. CODEN CSSCDB. ISSN 0361-0918.

Diop:2016:SBI

[DDD16] Aba Diop, Aliou Diop, and Jean-François Dupuy. Simulationbased inference in a zero-inflated Bernoulli regression model. Communications in Statistics: Simulation and Computation, 45(10):3597–??, 2016. CODEN CSSCDB. ISSN 0361-

Dey:2016:PTI

[DDK16] Tanujit Dey, Sanku Dey, and Debasis Kundu. On progressively Type-II censored two-parameter Rayleigh distribution. Communications in Statistics: Simulation and Computation, 45(2):438–455, 2016. CODEN CSSCDB. ISSN 0361-0918.

DeMartini:2011:RCS

[De 11] Daniele De Martini. Robustness and corrections for sample size adaptation strategies based on effect size estimation. Communications in Statistics: Simulation and Computation, 40(9):1263–1277, 2011. CODEN CSSCDB. ISSN 0361-0918.

DeCapitani:2014:IES

[De 14] Lucio De Capitani. Interval estimation for the Sortino ratio and the Omega ratio. Communications in Statistics: Simulation and Computation, 43(6):1385–1429, 2014. CODEN CSSCDB. ISSN 0361-0918.

Deiri:2012:CEP

[Dei12] Einolah Deiri. Classical estimator of parameters of the gamma case with presence of two outliers. Communications in Statistics: Simulation and Computation, 41(5):590–597, 2012. CODEN CSSCDB. ISSN 0361-0918.

Delzell:2017:IQD

[Del17] Darcie A. P. Delzell. An investigation of a quadratic design criterion for estimation of the hemodynamic response function. Communications in Statistics: Simulation and Computation, 46(4):2945–2956, 2017. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2017:CGB

[Dem17a] Hakan Demirtas. Concurrent generation of binary and nonnormal continuous data through fifth-order power polynomials. Communications in Statistics: Simulation and Computation, 46(1):344–357, 2017. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2017:APG

[Dem17b] Hakan Demirtas. On accurate and precise generation of generalized Poisson variates. Communications in Statistics:

Simulation and Computation, 46(1):489–499, 2017. CODEN CSSCDB. ISSN 0361-0918.

Demirhan:2018:GGD

[Dem18] Haydar Demirhan. A generalized Gumbel distribution and its parameter estimation. Communications in Statistics: Simulation and Computation, 47(10):2829–2848, 2018. CO-DEN CSSCDB. ISSN 0361-0918.

Deng:2011:NFM

[Den11] Yihao Deng. A note on the factorization of a matrix inverse. Communications in Statistics: Simulation and Computation, 40(8):1122–1129, 2011. CODEN CSSCDB. ISSN 0361-0918.

deFreitas:2013:SEC

[dFR13] Luiz Antonio de Freitas and Josemar Rodrigues. Standard exponential cure rate model with informative censoring. Communications in Statistics: Simulation and Computation, 42(1):8–23, 2013. CODEN CSSCDB. ISSN 0361-0918.

Eustaquio:2017:NHT

[dGEB17] John de Guzman Eustaquio and Erniel B. Barrios. Nonparametric hypothesis testing in clustered survival model. Communications in Statistics: Simulation and Computation, 46 (9):7485–7500, 2017. CODEN CSSCDB. ISSN 0361-0918.

Debusho:2019:ACO

[DGH19] Legesse K. Debusho, Dibaba B. Gemechu, and Linda M. Haines. Algorithmic construction of optimal block designs for two-colour cDNA microarray experiments using the linear mixed effects model. Communications in Statistics: Simulation and Computation, 48(7):1948–1963, 2019. CODEN CSSCDB. ISSN 0361-0918.

Dunder:2018:VSL

[DGMC18] Emre Dünder, Serpil Gümüstekin, Naci Murat, and Mehmet Ali Cengiz. Variable selection in linear regression analysis with alternative Bayesian information criteria using differential evaluation algorithm. Communications in Statistics: Simulation and Computation, 47(2):605–614, 2018. CODEN CSSCDB. ISSN 0361-0918.

Dedecius:2012:AMP

[DH12] Kamil Dedecius and Radek Hofman. Autoregressive model with partial forgetting within Rao-blackwellized particle filter. Communications in Statistics: Simulation and Computation, 41(5):582–589, 2012. CODEN CSSCDB. ISSN 0361-0918.

Demirtas:2016:CPB

[DH16] Hakan Demirtas and Donald Hedeker. Computing the pointbiserial correlation under any underlying continuous distribution. Communications in Statistics: Simulation and Computation, 45(8):2744–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dag:2017:AEB

[DI17] Osman Dag and Ozlem Ilk. An algorithm for estimating Box–Cox transformation parameter in ANOVA. Communications in Statistics: Simulation and Computation, 46(8): 6424–6435, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ding:2018:BES

[Din18] Jianhua Ding. Bayesian estimation on semiparametric models with shape restricted B-splines. Communications in Statistics: Simulation and Computation, 47(5):1315–1325, 2018. CODEN CSSCDB. ISSN 0361-0918.

deJong:2016:MIP

[dJvBS16] Roel de Jong, Stef van Buuren, and Martin Spiess. Multiple imputation of predictor variables using generalized additive models. Communications in Statistics: Simulation and Computation, 45(3):968–985, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dogu:2011:ECP

[DK11] Eralp Doğu and Ipek Deveci Kocakoç. Estimation of change point in generalized variance control chart. Communications in Statistics: Simulation and Computation, 40(3):345–363, 2011. CODEN CSSCDB. ISSN 0361-0918.

Dogu:2013:MCP

[DK13] Eralp Dogu and Ipek Deveci Kocakoç. A multivariate change point detection procedure for monitoring mean and covariance simultaneously. Communications in Statistics: Simulation and Computation, 42(6):1235–1255, 2013. CODEN CSSCDB. ISSN 0361-0918.

Dawoud:2017:EPP

[DK17a] Issam Dawoud and Selahattin Kaçiranlar. Evaluation of the predictive performance of the Liu type estimator. Communications in Statistics: Simulation and Computation, 46(4): 2800–2820, 2017. CODEN CSSCDB. ISSN 0361-0918.

Dawoud:2017:OTMa

[DK17b] Issam Dawoud and Selahattin Kaçiranlar. An optimal k of k th MA-ARIMA models under AR(p) models. Communications in Statistics: Simulation and Computation, 46(4): 2842–2864, 2017. CODEN CSSCDB. ISSN 0361-0918.

Dawoud:2017:OTMb

[DK17c] Issam Dawoud and Selahattin Kaçiranlar. An optimal k of k th MA-ARIMA models under MA(q) models. Communications in Statistics: Simulation and Computation, 46(6): 4185–4198, 2017. CODEN CSSCDB. ISSN 0361-0918.

Drosou:2019:SIS

[DK19] K. Drosou and C. Koukouvinos. Sure independence screening for analyzing supersaturated designs. Communications in Statistics: Simulation and Computation, 48(7):1979– 1995, 2019. CODEN CSSCDB. ISSN 0361-0918.

Djerroud:2018:MNP

[DKA18] Lamia Djerroud, Tristan Senga Kiessé, and Smail Adjabi. A multivariate non-parametric kernel estimator for global sensitivity analysis. Communications in Statistics: Simulation and Computation, 47(6):1606–1622, 2018. CODEN CSSCDB. ISSN 0361-0918.

Drosou:2018:MAS

[DKL18] K. Drosou, C. Koukouvinos, and A. Lappa. A method for analyzing supersaturated designs inspired by control charts. Communications in Statistics: Simulation and Computation, 47(4):1134–1145, 2018. CODEN CSSCDB. ISSN 0361-0918.

Dey:2017:ECD

[DKRL17] Sanku Dey, Devendra Kumar, Pedro L. Ramos, and Francisco Louzada. Exponentiated Chen distribution: Properties and estimation. Communications in Statistics: Simulation and Computation, 46(10):8118–8139, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Dufour:2015:FSR

[DKV15] Jean-Marie Dufour, Lynda Khalaf, and Marcel Voia. Finitesample resampling-based combined hypothesis tests, with applications to serial correlation and predictability. Communications in Statistics: Simulation and Computation, 44 (9):2329–2347, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ding:2017:CRT

[DL17] Juan Ding and Hao Li. Comparison of robust tests for genetic association analysis incorporating uncertain genotype. Communications in Statistics: Simulation and Computation, 46(5):3436–3443, 2017. CODEN CSSCDB. ISSN 0361-0918.

delaHorra:2017:BID

[dlHMRB17] Julián de la Horra, J. Miguel Marín, and María Teresa Rodríguez-Bernal. Bayesian inference and data cloning in the calibration of population projection matrices. Communications in Statistics: Simulation and Computation, 46(3): 1669–1681, 2017. CODEN CSSCDB. ISSN 0361-0918.

deLara:2019:GLT

[dLHT19] Idemauro Antonio Rodrigues de Lara, John Hinde, and Cesar Augusto Taconeli. Global and local tests to assess stationarity of Markov transition models. Communications in Statistics: Simulation and Computation, 48(4):1019–1039, 2019. CODEN CSSCDB. ISSN 0361-0918.

Dubeau:2014:EGM

[DM14] F. Dubeau and Y. Mir. Exponential growth model: From horizontal to linear asymptote. Communications in Statistics: Simulation and Computation, 43(10):2186–2204, 2014. CODEN CSSCDB. ISSN 0361-0918.

Das:2019:RDM

[DM19] Ishapathik Das and Siuli Mukhopadhyay. Robust designs for multinomial models. Communications in Statistics: Simulation and Computation, 48(10):2998–3021, 2019. CODEN CSSCDB. ISSN 0361-0918.

Brito:2017:OAA

[dMBSFB17] José André de Moura Brito, Gustavo Silva Semaan, Augusto Cesar Fadel, and Luciana Roque Brito. An optimization approach applied to the optimal stratification problem. Communications in Statistics: Simulation and Computation, 46(6):4419–4451, 2017. CODEN CSSCDB. ISSN 0361-0918.

deMorais:2017:IDC

[dMdCQSH17] Ricardo Saldanha de Morais, Roberto da Costa Quinino, Emílio Suyama, and Linda Lee Ho. Inferences in dichotomous classifications with misclassifications based on sequential repetitive classifications. Communications in Statistics: Simulation and Computation, 46(10):8264–8282, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Donneau:2015:SSC

[DMMA15] A. F. Donneau, M. Mauer, G. Molenberghs, and A. Albert. A simulation study comparing multiple imputation methods for incomplete longitudinal ordinal data. Communications in Statistics: Simulation and Computation, 44(5):1311–1338, 2015. CODEN CSSCDB. ISSN 0361-0918.

Diongue:2016:FGP

[DN16] Abdou Kâ Diongue and Mor Ndongo. The k-factor GARMA process with infinite variance innovations. Communications in Statistics: Simulation and Computation, 45(2):420–437, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dauxois:2018:SEP

[DN18] J. Y. Dauxois and A. Nucit. Some estimation problems in epidemic modeling. Communications in Statistics: Simulation and Computation, 47(4):1066–1087, 2018. CODEN CSSCDB. ISSN 0361-0918.

Dogan:2017:EDD

[DÖ17] Ilkay Dogan and Kazim Özdamar. The effect of different data structures, sample sizes on model fit measures. Communications in Statistics: Simulation and Computation, 46 (9):7525–7533, 2017. CODEN CSSCDB. ISSN 0361-0918.

deOliveira:2019:CBR

[dODD19] Willian Luís de Oliveira, Carlos Alberto Ribeiro Diniz, and Maria Durbán. A class of bivariate regression models for discrete and/or continuous responses. Communications in Statistics: Simulation and Computation, 48(8):2359–2383, 2019. CODEN CSSCDB. ISSN 0361-0918.

Dong:2010:BMD

[Don10] Fanglong Dong. Bayesian method to detect outliers for ordinal data. Communications in Statistics: Simulation and Computation, 39(7):1470–1484, 2010. CODEN CSSCDB. ISSN 0361-0918.

Deng:2017:SEM

[DOZ17] Wenli Deng, Fei Ouyang, and Jiajia Zhang. Semiparametric estimation method for accelerated failure time model with dependent censoring. Communications in Statistics: Simulation and Computation, 46(9):6947–6958, 2017. CODEN CSSCDB. ISSN 0361-0918.

Daneshkhah:2016:AMD

[DPCJ16] A. Daneshkhah, G. Parham, O. Chatrabgoun, and M. Jokar. Approximation multivariate distribution with pair copula using the orthonormal polynomial and Legendre multiwavelets basis functions. Communications in Statistics: Simulation and Computation, 45(2):389–419, 2016. CODEN CSSCDB. ISSN 0361-0918.

Du:2017:RSR

[DPS17] Wenyu Du, Aleksey S. Polunchenko, and Grigory Sokolov. On robustness of the Shiryaev–Roberts change-point detection procedure under parameter misspecification in the postchange distribution. Communications in Statistics: Simulation and Computation, 46(3):2185–2206, 2017. CODEN CSSCDB. ISSN 0361-0918.

deSouzaChaves:2011:SEC

[dR11] Josenildo de Souza Chaves and Josemar Rodrigues. Standard exponential cure rate model with noninformative or informative uniform-exponential censoring. Communications in Statistics: Simulation and Computation, 40(3):364–382, 2011. CODEN CSSCDB. ISSN 0361-0918.

Dhhan:2018:SIS

[DRAM18] Waleed Dhhan, Sohel Rana, Taha Alshaybawee, and Habshah Midi. The single-index support vector regression model to address the problem of high dimensionality. Communications in Statistics: Simulation and Computation, 47(9): 2792–2799, 2018. CODEN CSSCDB. ISSN 0361-0918.

Duarte:2013:ODA

[DS13] Belmiro P. M. Duarte and Pedro M. Saraiva. Optimal design of acceptance sampling plans by variables for nonconforming proportions when the standard deviation is unknown. Communications in Statistics: Simulation and Computation, 42 (6):1318–1342, 2013. CODEN CSSCDB. ISSN 0361-0918.

Davalos:2014:EIC

[DS14] Angel Dávalos and Naijun Sha. Estimating intraclass correlation coefficient and identifying influential observations under one-way random effects model. Communications in Statistics: Simulation and Computation, 43(10):2374–2389, 2014. CODEN CSSCDB. ISSN 0361-0918.

daSilva:2019:SAC

[dSBHM19] Marcelo A. da Silva, Jorge L. Bazán, and Anne Corinne Huggins-Manley. Sensitivity analysis and choosing between alternative polytomous IRT models using Bayesian model comparison criteria. Communications in Statistics: Simulation and Computation, 48(2):601–620, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Das:2019:ORT

[DSD19] Sudipta Das, Debasis Sengupta, and Anup Dewanji. Optimum release time of a software under periodic debugging schedule. Communications in Statistics: Simulation and Computation, 48(5):1516–1534, 2019. CODEN CSSCDB. ISSN 0361-0918.

Dey:2018:BCI

[DSMJ18] Sanku Dey, Mahendra Saha, Sudhansu S. Maiti, and Chi-Hyuck Jun. Bootstrap confidence intervals of generalized process capability index C_{pyk} for Lindley and power Lindley distributions. Communications in Statistics: Simulation and Computation, 47(1):249–262, 2018. CODEN CSSCDB. ISSN 0361-0918.

Da-Silva:2015:BDD

[DSR15] Cibele Queiroz Da-Silva and Guilherme Souza Rodrigues. Bayesian dynamic Dirichlet models. Communications in Statistics: Simulation and Computation, 44(3):787–818, 2015. CODEN CSSCDB. ISSN 0361-0918.

Deng:2012:PEL

[DTL12] Wenli Deng, Yong Tian, and Qiuping Lv. Parametric estimator of linear model with interval-censored data. Communications in Statistics: Simulation and Computation, 41 (10):1794–1804, 2012. CODEN CSSCDB. ISSN 0361-0918.

Drezner:2010:MKS

[DTZ10] Zvi Drezner, Ofir Turel, and Dawit Zerom. A modified Kolmogorov–Smirnov test for normality. Communications in Statistics: Simulation and Computation, 39(4):693–704, 2010. CODEN CSSCDB. ISSN 0361-0918.

Draxler:2016:CML

[DTZG16] Clemens Draxler, Gerhard Tutz, Katharina Zink, and Can Gürer. Comparison of maximum likelihood with conditional pairwise likelihood estimation of person parameters in the Rasch model. Communications in Statistics: Simulation and Computation, 45(6):2007–2017, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dutta:2014:LSU

[Dut14] Santanu Dutta. Local smoothing using the bootstrap. Communications in Statistics: Simulation and Computation, 43 (2):378–389, 2014. CODEN CSSCDB. ISSN 0361-0918.

Dutta:2015:LSK

[Dut15] Santanu Dutta. Local smoothing for kernel distribution function estimation. Communications in Statistics: Simulation and Computation, 44(4):878–891, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Dutta:2016:CVR

[Dut16] Santanu Dutta. Cross-validation revisited. Communications in Statistics: Simulation and Computation, 45(2):472–490, 2016. CODEN CSSCDB. ISSN 0361-0918.

Dong:2016:VIL

[DW16] Zijian Dong and Zhongming Wang. Variational inference of linear regression with nonzero prior means. Communications in Statistics: Simulation and Computation, 45(7):2241–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

deWinter:2016:CFA

[dWD16] Joost C. F. de Winter and Dimitra Dodou. Common factor analysis versus principal component analysis: A comparison of loadings by means of simulations. Communications in Statistics: Simulation and Computation, 45(1):299–321, 2016. CODEN CSSCDB. ISSN 0361-0918.

Diawara:2019:SLM

[DWKL19] Norou Diawara, Lance Waller, Robert King, and Jennifer Lorio. Simulations of local Moran’s index in a spatiotemporal setting. Communications in Statistics: Simulation and Computation, 48(6):1849–1859, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Duan:2018:IGP

[DWW18] Fengjun Duan, Guanjun Wang, and Huan Wang. Inverse Gaussian process models for bivariate degradation analysis: a Bayesian perspective. Communications in Statistics: Simulation and Computation, 47(1):166–186, 2018. CODEN CSSCDB. ISSN 0361-0918.

Ding:2016:NEP

[DX16] Juan Ding and Wenjun Xiong. A new estimator for a population proportion using group testing. Communications in Statistics: Simulation and Computation, 45(1):101–114, 2016. CODEN CSSCDB. ISSN 0361-0918.

Drezner:2016:SED

[DZ16] Zvi Drezner and Dawit Zerom. A simple and effective discretization of a continuous random variable. Communications in Statistics: Simulation and Computation, 45(10): 3798–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ding:2017:EPL

[DZ17] Jianhua Ding and Zhongzhan Zhang. M-estimation in the partially linear model with Bernstein polynomials under shape constrains. Communications in Statistics: Simulation and Computation, 46(1):779–794, 2017. CODEN CSSCDB. ISSN 0361-0918.

Dharmasena:2010:EPC

[DZC10] L. Sandamali Dharmasena, P. Zeephongsekul, and P. Castagliola. Estimating process capability index C_{P M} using a bootstrap sequential sampling procedure. Communications in Statistics: Simulation and Computation, 39(6):1097–1110, 2010. CODEN CSSCDB. ISSN 0361-0918.

Duan:2019:SBA

[DZT19] Xing-De Duan, Yuan-Ying Zhao, and An-Min Tang. A semiparametric Bayesian approach to simplex regression model with heterogeneous dispersion. Communications in Statistics: Simulation and Computation, 48(8):2487–2500, 2019. CODEN CSSCDB. ISSN 0361-0918.

Du:2018:ECP

[DZX18] Jiang Du, Zhongzhan Zhang, and Dengke Xu. Estimation for the censored partially linear quantile regression models. Communications in Statistics: Simulation and Computation, 47(8):2393–2408, 2018. CODEN CSSCDB. ISSN 0361-0918.

Erdugan:2012:CMJ

[EA12] Funda Erdugan and Fikri Akdeniz. Computational method for jackknifed generalized ridge tuning parameter based on generalized maximum entropy. Communications in Statistics: Simulation and Computation, 41(8):1411–1429, 2012. CODEN CSSCDB. ISSN 0361-0918.

Erees:2017:IMA

[EA17] Suay Erees and Aylin Alin. Influences of misspecification on asymptotic relative efficiency of coefficients of determination: Application to agriculture. Communications in Statistics: Simulation and Computation, 46(3):1842–1857, 2017. CODEN CSSCDB. ISSN 0361-0918.

El-Aroui:2018:POG

[EAG18] Mhamed-Ali El-Aroui and Olivier Gaudoin. Prequential omnibus goodness-of-fit tests for stochastic processes: a numerical study. Communications in Statistics: Simulation and Computation, 47(3):747–761, 2018. CODEN CSSCDB. ISSN 0361-0918.

Edwards:2014:AFP

[EB14] David J. Edwards and J. Paul Brooks. Alternative foldover plans for two-level nonregular designs. Communications in Statistics: Simulation and Computation, 43(1):209–224, 2014. CODEN CSSCDB. ISSN 0361-0918.

Esmenda:2018:REM

[EB18] Mary Jane Esmenda and Erniel B. Barrios. Robust estimation of a multilevel model with structural change. Communications in Statistics: Simulation and Computation, 47(4): 1014–1027, 2018. CODEN CSSCDB. ISSN 0361-0918.

El-Din:2017:SIG

[EDAAAM17] M. M. Mohie El-Din, Y. Abdel-Aty, and M. H. Abu-Moussa. Statistical inference for the Gompertz distribution based on Type-II progressively hybrid censored data. Communications in Statistics: Simulation and Computation, 46(8): 6242–6260, 2017. CODEN CSSCDB. ISSN 0361-0918.

El-Din:2018:PIO

[EDE18] M. M. Mohie El-Din and W. S. Emam. Prediction intervals for ordinary and dual generalized order statistics from two independent sequences. Communications in Statistics: Simulation and Computation, 47(9):2703–2723, 2018. CODEN CSSCDB. ISSN 0361-0918.

Erisoglu:2010:MHS

[EE10] Ülkü Erisoglu and Hamza Erol. Modeling heterogeneous survival data using mixture of extended exponential-geometric distributions. Communications in Statistics: Simulation and Computation, 39(10):1939–1952, 2010. CODEN CSSCDB. ISSN 0361-0918.

ElHefnawy:2014:CNP

[EF14] Ali El Hefnawy and Aya Farag. A combined nonlinear programming model and Kibria method for choosing ridge parameter regression. Communications in Statistics: Simulation and Computation, 43(6):1442–1470, 2014. CODEN CSSCDB. ISSN 0361-0918.

Escudeiro:2017:ACI

[EFA17] Sara Escudeiro, Adelaide Freitas, and Vera Afreixo. Approximate confidence intervals for a linear combination of binomial proportions: a new variant. Communications in Statistics: Simulation and Computation, 46(9):7501–7524, 2017. CODEN CSSCDB. ISSN 0361-0918.

Evans:2016:DBM

[EG16] Dafydd Evans and Jonathan Gillard. Difference-based methods for truncating the singular value decomposition. Communications in Statistics: Simulation and Computation, 45 (3):863–879, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ellingson:2017:NBS

[EGO^{+}17] Leif Ellingson, David Groisser, Daniel Osborne, Vic Patrangenaru, and Armin Schwartzman. Nonparametric bootstrap of sample means of positive-definite matrices with an application to diffusion-tensor-imaging data analysis. Communications in Statistics: Simulation and Computation, 46 (6):4851–4879, 2017. CODEN CSSCDB. ISSN 0361-0918.

El-Horbaty:2018:TRE

[EH18] Yahia S. El-Horbaty. On testing random effects in the nested-error regression model. Communications in Statistics: Simulation and Computation, 47(6):1670–1676, 2018. CODEN CSSCDB. ISSN 0361-0918.

Eryilmaz:2010:DRE

[EK10] Serkan Eryilmaz and Cihangir Kan. Dynamic reliability evaluation of consecutive-k-within-m-out-of-n: F system. Communications in Statistics: Simulation and Computation, 40(1):58–71, 2010. CODEN CSSCDB. ISSN 0361-0918.

Egorov:2012:SAP

[EK12] Vladimir Egorov and Nourlan Kondybaev. On some applications of Poisson point processes. Communications in Statistics: Simulation and Computation, 41(6):804–819, 2012. CODEN CSSCDB. ISSN 0361-0918.

Ene:2016:NEM

[EK16] Nehir Ene and Mehmet Karahasan. A new estimation method based on order statistics in the families of symmetric location-scale distributions. Communications in Statistics: Simulation and Computation, 45(5):1802–1825, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Ertas:2017:RLT

[EKG17] Hasan Ertas, Selahattin Kaçiranlar, and Hüseyin Güler. Robust Liu-type estimator for regression based on Mestimator. Communications in Statistics: Simulation and Computation, 46(5):3907–3932, 2017. CODEN CSSCDB. ISSN 0361-0918.

Eftekhar:2019:TES

[EKK19] Sana Eftekhar and Mahmood Kharrati-Kopaei. Testing the equality of several inverse Gaussian means under heterogeneity. Communications in Statistics: Simulation and Computation, 48(9):2757–2774, 2019. CODEN CSSCDB. ISSN 0361-0918.

Emura:2018:CRC

[EL18] Takeshi Emura and Yu-Ting Liao. Critical review and comparison of continuity correction methods: The normal approximation to the binomial distribution. Communications in Statistics: Simulation and Computation, 47(8): 2266–2285, 2018. CODEN CSSCDB. ISSN 0361-0918.

Elhassanein:2017:CDF

[Elh17] A. Elhassanein. On the complex dynamics of functionalcoefficients nonlinear autoregressive time series models. Communications in Statistics: Simulation and Computation, 46(3):2406–2417, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ellis:2011:AUQ

[Ell11] Steven P. Ellis. An algorithm for unconstrained quadratically penalized convex optimization. Communications in Statistics: Simulation and Computation, 40(7):1006–1029, 2011. CODEN CSSCDB. ISSN 0361-0918.

Emura:2017:RRP

[ELS17] Takeshi Emura, Ting-Hsuan Long, and Li-Hsien Sun. R routines for performing estimation and statistical process control under copula-based time series models. Communications in Statistics: Simulation and Computation, 46(4): 3067–3087, 2017. CODEN CSSCDB. ISSN 0361-0918.

Elsawah:2019:COR

[Els19] A. M. Elsawah. Constructing optimal router bit life sequential experimental designs: New results with a case study. Communications in Statistics: Simulation and Computation, 48(3):723–752, 2019. CODEN CSSCDB. ISSN 0361-

Emadi:2013:MMV

[Ema13] M. Emadi. Mixture models in view of evidential analysis. Communications in Statistics: Simulation and Computation, 42(8):1824–1835, 2013. CODEN CSSCDB. ISSN 0361-0918.

ElMatouat:2015:ROE

[EMH15] Abdelaziz El Matouat, Freedath Djibril Moussa, and Hassania Hamzaoui. Resampling for order estimation of autoregressive models with missing data. Communications in Statistics: Simulation and Computation, 44(5):1187–1196, 2015. CODEN CSSCDB. ISSN 0361-0918.

Eskandari:2016:FMG

[EO16] Farzad Eskandari and Ehsan Ormoz. Finite mixture of generalized semiparametric models: Variable selection via penalized estimation. Communications in Statistics: Simulation and Computation, 45(10):3744–??, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Ertas:2018:MRE

[Ert18] Hasan Ertas. A modified ridge m-estimator for linear regression model with multicollinearity and outliers. Communications in Statistics: Simulation and Computation, 47(4): 1240–1250, 2018. CODEN CSSCDB. ISSN 0361-0918.

Eryilmaz:2010:NWC

[Ery10] Serkan Eryilmaz. Number of working components in consecutive k-out-of-n system while it is working. Communications in Statistics: Simulation and Computation, 39(4):683–692, 2010. CODEN CSSCDB. ISSN 0361-0918.

Eryilmaz:2014:SSP

[Ery14] Serkan Eryilmaz. On signatures of series and parallel systems consisting of modules with arbitrary structures. Communications in Statistics: Simulation and Computation, 43 (5):1202–1211, 2014. CODEN CSSCDB. ISSN 0361-0918.

Eryilmaz:2015:MTF

[Ery15] Serkan Eryilmaz. Mean time to failure of weighted k-out-of-n: G systems. Communications in Statistics: Simulation and Computation, 44(10):2705–2713, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Eryilmaz:2017:RAS

[Ery17] Serkan Eryilmaz. Reliability analysis of systems with components having two dependent subcomponents. Communications in Statistics: Simulation and Computation, 46(10): 8005–8013, 2017. CODEN CSSCDB. ISSN 0361-0918.

Eryilmaz:2018:PTS

[Ery18] Serkan Eryilmaz. Phase type stress-strength models with reliability applications. Communications in Statistics: Simulation and Computation, 47(4):954–963, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Eryilmaz:2019:MST

[Ery19] Serkan Eryilmaz. Modeling systems with two dependent components under bivariate shock models. Communications in Statistics: Simulation and Computation, 48(6): 1714–1728, 2019. CODEN CSSCDB. ISSN 0361-0918.

Economou:2015:SST

[ES15] Polychronis Economou and Milan Stehlík. On small samples testing for frailty through homogeneity test. Communications in Statistics: Simulation and Computation, 44(1): 40–65, 2015. CODEN CSSCDB. ISSN 0361-0918.

Emura:2016:EMS

[ES16] Takeshi Emura and Shau-Kai Shiu. Estimation and model selection for left-truncated and right-censored lifetime data with application to electric power transformers analysis. Communications in Statistics: Simulation and Computation, 45(9):3171–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Elghafghuf:2017:RPL

[ES17a] Adel Elghafghuf and Henrik Stryhn. Robust Poisson likelihood estimation for frailty Cox models: A simulation study. Communications in Statistics: Simulation and Computation, 46(4):2907–2923, 2017. CODEN CSSCDB. ISSN 0361-0918.

Epasinghe:2017:GFT

[ES17b] Nirodha Epasinghe and Roshini Sooriyarachchi. A goodness of fit test for the multilevel proportional odds model. Communications in Statistics: Simulation and Computation, 46 (7):5610–5626, 2017. CODEN CSSCDB. ISSN 0361-0918.

Esmaeeli:2017:MTS

[ESAD17] Hamid Esmaeeli, Ahmad Sadegheih, Amirhossein Amiri, and Mohammad Hadi Doroudyan. Monitoring two-stage processes with a profile at the second stage. Communications in Statistics: Simulation and Computation, 46(3): 2457–2473, 2017. CODEN CSSCDB. ISSN 0361-0918.

ElAdlouni:2018:RBQ

[ESSH18] Salaheddine El Adlouni, Garba Salaou, and André St-Hilaire. Regularized Bayesian quantile regression. Communications in Statistics: Simulation and Computation, 47(1): 277–293, 2018. CODEN CSSCDB. ISSN 0361-0918.

Elbayoumi:2016:WEF

[ET16] Tamer M. Elbayoumi and Jeff Terpstra. Weighted L_{1}- estimates for the first-order bifurcating autoregressive model. Communications in Statistics: Simulation and Computation, 45(8):2991–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Endharta:2017:CBM

[EY17] Alfonsus Julanto Endharta and Won Young Yun. Conditionbased maintenance policy for linear consecutive-k-out-of-n: F system. Communications in Statistics: Simulation and Computation, 46(4):3088–3102, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ehlers:2015:BEP

[EZ15] Ricardo Ehlers and M. Zevallos. Bayesian estimation and prediction of stochastic volatility models via INLA. Communications in Statistics: Simulation and Computation, 44 (3):683–693, 2015. CODEN CSSCDB. ISSN 0361-0918.

Esmailzadeh:2017:UOA

[EZ17] Nabaz Esmailzadeh and Zahra Zandi. Using orthogonal array for constructing three-level search designs. Communications in Statistics: Simulation and Computation, 46(3): 1906–1917, 2017. CODEN CSSCDB. ISSN 0361-0918.

Feroze:2017:SSP

[FA17] Navid Feroze and Muhammad Aslam. On selection of a suitable prior for the posterior analysis of censored mixture of Topp Leone distribution. Communications in Statistics: Simulation and Computation, 46(7):5184–5211, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Fan:2018:PLA

[Fan18] Chunpeng Fan. A pseudo likelihood approach to analyze rate difference of binary response data in longitudinal factorial studies. Communications in Statistics: Simulation and Computation, 47(7):2169–2183, 2018. CODEN CSSCDB. ISSN 0361-0918.

Farne:2016:ASV

[Far16] Matteo Farné. An algorithm to simulate VMA processes having a spectrum with fixed condition number. Communications in Statistics: Simulation and Computation, 45(5): 1664–1675, 2016. CODEN CSSCDB. ISSN 0361-0918.

Feng:2018:TSV

[FCC18] Mu Feng, Zhao Chen, and Ximing Cheng. Two-step variable selection in partially linear additive models with time series data. Communications in Statistics: Simulation and Computation, 47(3):661–671, 2018. CODEN CSSCDB. ISSN 0361-0918.

Feddag:2012:CML

[Fed12] M.-L. Feddag. Composite marginal likelihoods to the normal Bradley-Terry model. Communications in Statistics: Simulation and Computation, 41(3):279–286, 2012. CODEN CSSCDB. ISSN 0361-0918.

Feng:2016:ABS

[Fen16] Xinwei Feng. Anticipated backward stochastic differential equation with reflection. Communications in Statistics: Simulation and Computation, 45(5):1676–1688, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Feng:2017:DDS

[Fen17] Guangchao C. Feng. Do difference scores make a difference on the third-person effect? Communications in Statistics: Simulation and Computation, 46(7):5085–5104, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Frobish:2016:SEC

[FEP16] Daniel Frobish, Nader Ebrahimi, and Dung Pham. Semiparametric estimation of a change-point for recurrent events data. Communications in Statistics: Simulation and Computation, 45(9):3339–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ferreira:2017:EMM

[FFNM17] Dário Ferreira, Sandra S. Ferreira, Célia Nunes, and João T. Mexia. Estimation in mixed models through three step minimization. Communications in Statistics: Simulation and Computation, 46(2):1156–1166, 2017. CODEN CSSCDB. ISSN 0361-0918.

Figueiredo:2015:CVB

[FG15] Adelaide Figueiredo and Paulo Gomes. Clustering of variables based on Watson distribution on hypersphere: A comparison of algorithms. Communications in Statistics: Simulation and Computation, 44(10):2622–2635, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ferrara:2010:TFO

[FGL10] Laurent Ferrara, Dominique Guegan, and Zhiping Lu. Testing fractional order of long memory processes: A Monte Carlo study. Communications in Statistics: Simulation and Computation, 39(4):795–806, 2010. CODEN CSSCDB. ISSN 0361-0918.

Fuentes-Garcia:2010:NBN

[FGMW10] R. Fuentes-García, R. H. Mena, and S. G. Walker. A new Bayesian nonparametric mixture model. Communications in Statistics: Simulation and Computation, 39(4):669–682, 2010. CODEN CSSCDB. ISSN 0361-0918.

Farahani:2014:BSI

[FK14] Zahra Sadat Meshkani Farahani and Esmaile Khorram. Bayesian statistical inference for weighted exponential distribution. Communications in Statistics: Simulation and Computation, 43(6):1362–1384, 2014. CODEN CSSCDB. ISSN 0361-0918.

Fischer:2015:PMG

[FK15] T. Fischer and U. Kamps. Power maps in goodness-of-fit testing based on censored samples. Communications in Statistics: Simulation and Computation, 44(7):1931–1974, 2015. CODEN CSSCDB. ISSN 0361-0918.

Fika:2017:SET

[FK17] Paraskevi Fika and Christos Koukouvinos. Stochastic estimates for the trace of functions of matrices via Hadamard matrices. Communications in Statistics: Simulation and Computation, 46(5):3491–3503, 2017. CODEN CSSCDB. ISSN 0361-0918.

Filzmoser:2018:RLR

[FK18] Peter Filzmoser and Fatma Sevinç Kurnaz. A robust Liu regression estimator. Communications in Statistics: Simulation and Computation, 47(2):432–443, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Firinguetti:2017:SPL

[FKA17] Luis Firinguetti, Golam Kibria, and Rodrigo Araya. Study of partial least squares and ridge regression methods. Communications in Statistics: Simulation and Computation, 46 (8):6631–6644, 2017. CODEN CSSCDB. ISSN 0361-0918.

Farahani:2019:CUW

[FKG19] Zahra Sadat Meshkani Farahani, Esmaile Khorram, and Mojtaba Ganjali. A comparison of using weighted distribution and joint modeling for analyzing non-ignorable missing responses. Communications in Statistics: Simulation and Computation, 48(3):704–722, 2019. CODEN CSSCDB. ISSN 0361-0918.

Fang:2013:ABD

[FL13] Liang Fang and Thomas M. Loughin. Analyzing binomial data in a split-plot design: Classical approach or modern techniques? Communications in Statistics: Simulation and Computation, 42(4):727–740, 2013. CODEN CSSCDB. ISSN 0361-0918.

Ferreira:2019:MLE

[FL19] Paulo H. Ferreira and Francisco Louzada. Maximum likelihood estimation for bivariate SUR tobit modeling in presence of two right-censored dependent variables. Communications in Statistics: Simulation and Computation, 48(1): 150–168, 2019. CODEN CSSCDB. ISSN 0361-0918.

Fierro:2015:SIS

[FLB15] Raúl Fierro, Víctor Leiva, and N. Balakrishnan. Statistical inference on a stochastic epidemic model. Communications in Statistics: Simulation and Computation, 44(9): 2297–2314, 2015. CODEN CSSCDB. ISSN 0361-0918.

Fang:2019:OMA

[FLW19] Fang Fang, Jialiang Li, and Jingli Wang. Optimal model averaging estimation for correlation structure in generalized estimating equations. Communications in Statistics: Simulation and Computation, 48(5):1574–1593, 2019. CODEN CSSCDB. ISSN 0361-0918.

Fox:2016:CLV

[FMMV16] Jean-Paul Fox, Maarten Marsman, Joris Mulder, and Josine Verhagen. Complex latent variable modeling in educational assessment. Communications in Statistics: Simulation and Computation, 45(5):1499–1510, 2016. CODEN CSSCDB. ISSN 0361-0918.

Fallahnezhad:2017:MPM

[FN17] M. S. Fallahnezhad and E. Najafian. A model of preventive maintenance for parallel, series, and single-item replacement systems based on statistical analysis. Communications in Statistics: Simulation and Computation, 46(7):5846–5859, 2017. CODEN CSSCDB. ISSN 0361-0918.

Fattah:2017:ETW

[FNA17] Ahmed A. Fattah, Saralees Nadarajah, and A-Hadi N. Ahmed. The exponentiated transmuted Weibull geometric distribution with application in survival analysis. Communications in Statistics: Simulation and Computation, 46(6): 4244–4263, 2017. CODEN CSSCDB. ISSN 0361-0918.

Filipiak:2014:PSI

[FP14] Katarzyna Filipiak and Simo Puntanen. Preface to the special issue on advances on linear models and inference: Computational aspects. Communications in Statistics: Simulation and Computation, 43(9):2045–2046, 2014. CODEN CSSCDB. ISSN 0361-0918.

Fosdick:2016:VSC

[FP16] Bailey K. Fosdick and Michael D. Perlman. Variancestabilizing and confidence-stabilizing transformations for the normal correlation coefficient with known variances. Communications in Statistics: Simulation and Computation, 45 (6):1918–1935, 2016. CODEN CSSCDB. ISSN 0361-0918.

Feng:2018:CME

[FP18] Cindy X. Feng and Andrea Piesse. On correcting measurement error in the persistence rate estimator. Communications in Statistics: Simulation and Computation, 47(4):997– 1013, 2018. CODEN CSSCDB. ISSN 0361-0918.

Faria:2012:CES

[FS12] Susana Faria and Gilda Soromenho. Comparison of EM and SEM algorithms in Poisson regression models: A simulation study. Communications in Statistics: Simulation and Computation, 41(4):497–509, 2012. CODEN CSSCDB. ISSN 0361-0918.

Fu:2017:CPA

[FS17] Qianrao Fu and Yimin Shi. Constant-partially accelerated life tests for Marshall–Olkin exponential series system with dependent masked data. Communications in Statistics: Simulation and Computation, 46(10):7879–7893, 2017. CODEN CSSCDB. ISSN 0361-0918.

Feng:2012:SLP

[FSJ12] Jing Feng, Quan Sun, and Tongdan Jin. Storage life prediction for a high-performance capacitor using multi-Phase Wiener degradation model. Communications in Statistics: Simulation and Computation, 41(8):1317–1335, 2012. CO-DEN CSSCDB. ISSN 0361-0918.

Faghihi:2016:OIB

[FT16] M. R. Faghihi and S. Tat. On Φ_{p}-optimality of incomplete block designs: An algorithm. Communications in Statistics: Simulation and Computation, 45(2):758–769, 2016. CODEN CSSCDB. ISSN 0361-0918.

Fan:2017:SAP

[FT17] Ai-Xia Fan and Nian-Sheng Tang. Sensitivity analysis of partially linear models with response missing at random. Communications in Statistics: Simulation and Computation, 46(7):5323–5339, 2017. CODEN CSSCDB. ISSN 0361-

Franko:2017:RWS

[FTE17] C. Franko, G. Y. Tütüncü, and S. Eryilmaz. Reliability of weighted k-out-of-n: G systems consisting of two types of components and a cold standby component. Communications in Statistics: Simulation and Computation, 46(5): 4067–4081, 2017. CODEN CSSCDB. ISSN 0361-0918.

Fan:2014:NAB

[FWY14] Tsai-Hung Fan, Guo-Tzau Wang, and Jenn-Hwa Yu. A new algorithm in Bayesian model averaging in regression models. Communications in Statistics: Simulation and Computation, 43(2):315–328, 2014. CODEN CSSCDB. ISSN 0361-0918.

Fu:2012:GVS

[FX12] Guang-Hui Fu and Qing-Song Xu. Grouping variable selection by weight fused elastic net for multi-collinear data. Communications in Statistics: Simulation and Computation, 41(2):205–221, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yu:2014:DDT

[fYyP14] Hong fwu Yu and Chien yu Peng. Designing a degradation test with a two-parameter exponential lifetime distribution. Communications in Statistics: Simulation and Computation, 43(8):1938–1958, 2014. CODEN CSSCDB. ISSN 0361-0918.

Fan:2017:PSS

[FZ17] Chunpeng Fan and Donghui Zhang. On power and sample size of the ANOVA-type rank test. Communications in Statistics: Simulation and Computation, 46(4):3224–3241, 2017. CODEN CSSCDB. ISSN 0361-0918.

Fu:2014:GVS

[FZDF14] Guang-Hui Fu, Wen-Ming Zhang, Lin Dai, and Ying-Zi Fu. Group variable selection with oracle property by weightfused adaptive elastic net model for strongly correlated data. Communications in Statistics: Simulation and Computation, 43(10):2468–2481, 2014. CODEN CSSCDB. ISSN 0361-0918.

Gong:2012:IRR

[GA12] Yankun Gong and Asheber Abebe. On the iteratively reweighted rank regression estimator. Communications in Statistics: Simulation and Computation, 41(2):155–166, 2012. CODEN CSSCDB. ISSN 0361-0918.

Gupta:2015:P

[GA15] Sat Gupta and Manohar Aggarwal. Preface. Communications in Statistics: Simulation and Computation, 44(9):2209, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gui:2017:ASP

[GA17] Wenhao Gui and Muhammad Aslam. Acceptance sampling plans based on truncated life tests for weighted exponential distribution. Communications in Statistics: Simulation and Computation, 46(3):2138–2151, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2019:IAC

[GA19] M. P. Gadre and S. B. Adnaik. Independent attributes control chart for Markov-dependent processes. Communications in Statistics: Simulation and Computation, 48(4):985–997, 2019. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2014:MGR

[Gad14] M. P. Gadre. A multivariate group runs control chart for process dispersion. Communications in Statistics: Simulation and Computation, 43(4):813–837, 2014. CODEN CSSCDB. ISSN 0361-0918.

Gao:2016:PSE

[GAK16] Hongjiang Gao, Inmaculada Aban, and Charles R. Katholi. Pool screening: An example of independent nonidentical Bernoulli trial. Communications in Statistics: Simulation and Computation, 45(9):3307–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ghashghaei:2019:NCC

[GAK19] Reza Ghashghaei, Amirhossein Amiri, and Peyman Khosravi. New control charts for simultaneous monitoring of the mean vector and covariance matrix of multivariate multiple linear profiles. Communications in Statistics: Simulation and Computation, 48(5):1382–1405, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Ghitany:2015:ERS

[GAMA15] M. E. Ghitany, D. K. Al-Mutairi, and S. M. Aboukhamseen. Estimation of the reliability of a stress-strength system from power Lindley distributions. Communications in Statistics: Simulation and Computation, 44(1):118–136, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gomez:2010:ATF

[GAPSÁ10] Víctor Gómez, Félix Aparicio-Pérez, and Ángel Sánchez-Ávila. An alternative to transfer function forecasting based on subspace methods. Communications in Statistics: Simulation and Computation, 39(10):1855–1867, 2010. CODEN CSSCDB. ISSN 0361-0918.

Guarte:2013:NHT

[GB13] Jacqueline Guarte and Erniel Barrios. Nonparametric hypothesis testing in a spatial-temporal model: A simulation study. Communications in Statistics: Simulation and Computation, 42(1):153–170, 2013. CODEN CSSCDB. ISSN 0361-0918.

Gauran:2017:NMC

[GB17a] Iris Ivy M. Gauran and Erniel B. Barrios. Nonparametric modeling of clustered customer survival data. Communications in Statistics: Simulation and Computation, 46(1): 603–618, 2017. CODEN CSSCDB. ISSN 0361-0918.

Guilatco:2017:NES

[GB17b] Ruffy S. Guilatco and Erniel B. Barrios. Nonparametric estimation of a switching regression model. Communications in Statistics: Simulation and Computation, 46(2):840–854, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gomaa:2019:SNM

[GB19] Abdel-Salam Gomaa and Jeffrey B. Birch. A semiparametric nonlinear mixed model approach to phase I profile monitoring. Communications in Statistics: Simulation and Computation, 48(6):1677–1693, 2019. CODEN CSSCDB. ISSN 0361-0918.

Giancristofaro:2012:ACT

[GBB^{+}12] Rosa Arboretti Giancristofaro, Mario Bolzan, Stefano Bonnini, Livio Corain, and Francesca Solmi. Advantages of the closed testing method in multiple comparisons procedures. Communications in Statistics: Simulation and Computation, 41(6):746–763, 2012. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2014:IAN

[GBO14] João Gomes, Isabel Barão, and Tiago Oliveira. Is it always necessary to take sample selection into account? Communications in Statistics: Simulation and Computation, 43(10): 2264–2274, 2014. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2016:NRB

[GBP16] M. Ivette Gomes, M. Fátima Brilhante, and Dinis Pestana. New reduced-bias estimators of a positive extreme value index. Communications in Statistics: Simulation and Computation, 45(3):833–862, 2016. CODEN CSSCDB. ISSN 0361-0918.

Gallardo:2016:PTC

[GBPDL16] Diego I. Gallardo, Heleno Bolfarine, and Antonio C. Pedroso-De-Lima. Promotion time cure rate model with bivariate random effects. Communications in Statistics: Simulation and Computation, 45(2):603–624, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Giancristofaro:2016:SME

[GBTV16] R. Arboretti Giancristofaro, V. Boatto, T. Tempesta, and D. Vecchiato. Statistical modeling for the evaluation of consumers’ preferences. Communications in Statistics: Simulation and Computation, 45(5):1569–1582, 2016. CODEN CSSCDB. ISSN 0361-0918.

Grabchak:2017:PTF

[GC17] Michael Grabchak and Victor Cosme. On the performance of Turing’s formula: A simulation study. Communications in Statistics: Simulation and Computation, 46(6):4199–4209, 2017. CODEN CSSCDB. ISSN 0361-0918.

Groparu-Cojocaru:2013:IGN

[GCD13] Ionica Groparu-Cojocaru and Louis G. Doray. Inference for the generalized normal Laplace distribution. Communications in Statistics: Simulation and Computation, 42(9): 1989–1997, 2013. CODEN CSSCDB. ISSN 0361-0918. See comments [BB17b].

Giancristofaro:2012:CAC

[GCR12] Rosa Arboretti Giancristofaro, Livio Corain, and Susanna Ragazzi. A comparison among combination-based permutation statistics for randomized complete block design. Communications in Statistics: Simulation and Computation, 41 (7):964–979, 2012. CODEN CSSCDB. ISSN 0361-0918.

Gallego:2015:CVM

[GD15] José L. Gallego and Carlos Díaz. Cointegrated VARIMA models: Specification and simulation. Communications in Statistics: Simulation and Computation, 44(1):66–70, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gomez-Deniz:2015:PEN

[GDCO15] E. Gómez-Déniz and E. Calderín-Ojeda. Parameters estimation for a new generalized geometric distribution. Communications in Statistics: Simulation and Computation, 44(8): 2023–2039, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gomez-Deniz:2016:PMI

[GDGG16] Emilio Gómez-Déniz, M. E. Ghitany, and Ramesh C. Gupta. Poisson-mixed inverse Gaussian regression model and its application. Communications in Statistics: Simulation and Computation, 45(8):2767–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Gomez-Deniz:2019:GAD

[GDSCO19] E. Gómez-Déniz, J. M. Sarabia, and E. Calderín-Ojeda. The geometric ArcTan distribution with applications to model demand for health services. Communications in Statistics: Simulation and Computation, 48(4):1101–1120, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Gott:2013:NEW

[GE13] Aimee N. Gott and Idris A. Eckley. A note on the effect of wavelet choice on the estimation of the evolutionary wavelet spectrum. Communications in Statistics: Simulation and Computation, 42(2):393–406, 2013. CODEN CSSCDB. ISSN 0361-0918.

Grogan:2017:SBM

[GE17] Tristan R. Grogan and David A. Elashoff. A simulation based method for assessing the statistical significance of logistic regression models after common variable selection procedures. Communications in Statistics: Simulation and Computation, 46(9):7180–7193, 2017. CODEN CSSCDB. ISSN 0361-0918.

Genc:2017:TIG

[Gen17] Ali I. Genç. Truncated inverted generalized exponential distribution and its properties. Communications in Statistics: Simulation and Computation, 46(6):4654–4670, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Gerhard:2016:SSS

[Ger16] Daniel Gerhard. Simultaneous small sample inference for linear combinations of generalized linear model parameters. Communications in Statistics: Simulation and Computation, 45(8):2678–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Estrada:2010:GFT

[GEVA10] Elizabeth González-Estrada and José A. Villaseñor-Alva. A goodness-of-fit test for location-scale max-stable distributions. Communications in Statistics: Simulation and Computation, 39(3):557–562, 2010. CODEN CSSCDB. ISSN 0361-0918.

Gokpinar:2017:TEV

[GG17a] Esra Gokpinar and Fikri Gokpinar. Testing equality of variances for several normal populations. Communications in Statistics: Simulation and Computation, 46(1):38–52, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gokpinar:2017:TES

[GG17b] Fikri Gokpinar and Esra Gokpinar. Testing the equality of several log-normal means based on a computational approach. Communications in Statistics: Simulation and Computation, 46(3):1998–2010, 2017. CODEN CSSCDB. ISSN 0361-0918.

Godwin:2019:ABC

[GG19] Ryan T. Godwin and David E. Giles. Analytic bias correction for maximum likelihood estimators when the bias function is non-constant. Communications in Statistics: Simulation and Computation, 48(1):15–26, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Gokpinar:2017:PAO

[GGB17] Fikri Gökpinar, Esra Gökpinar, and Hülya Bayrak. Permutation approach for ordinal preference data. Communications in Statistics: Simulation and Computation, 46(3): 2321–2332, 2017. CODEN CSSCDB. ISSN 0361-0918.

Guler:2017:CAB

[GGK17] Hüseyin Güler, Berrin Gültay, and Selahattin Kaçiranlar. Comparisons of the alternative biased estimators for the distributed lag models. Communications in Statistics: Simulation and Computation, 46(4):3306–3318, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gill:2017:SIO

[GGM17] A. N. Gill, Anju Goyal, and Vishal Maurya. Simultaneous inferences for ordered exponential location parameters under unbalanced data and heteroscedasticity of scale parameters. Communications in Statistics: Simulation and Computation, 46(4):3129–3139, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ghahroodi:2010:AMA

[GGNB10] Z. Rezaei Ghahroodi, M. Ganjali, H. Navvabpour, and D. Berridge. An appraisal of methods for the analysis of longitudinal ordinal response data with random dropout using a nonhomogeneous Markov model. Communications in Statistics: Simulation and Computation, 39(5):1027–1048, 2010. CODEN CSSCDB. ISSN 0361-0918.

Garcia:2019:NPP

[GGP19] Eunice Campirán García and Eduardo Gutiérrez-Peña. Nonparametric product partition models for multiple changepoints analysis. Communications in Statistics: Simulation and Computation, 48(7):1922–1947, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Gao:2017:RWE

[GGZG17] Bingbing Gao, Shesheng Gao, Yongmin Zhong, and Chengfan Gu. Random weighting estimation of sampling distributions via importance resampling. Communications in Statistics: Simulation and Computation, 46(1):640–654, 2017. CODEN CSSCDB. ISSN 0361-0918.

Guerbyenne:2015:BPS

[GH15] Hafida Guerbyenne and Fayçal Hamdi. Bootstrapping periodic state-space models. Communications in Statistics: Simulation and Computation, 44(2):374–401, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ghosh:2016:FEC

[GH16] Subir Ghosh and Mike (Fu Ze) Huang. Full estimation capacity hierarchical fractional factorial designs for a class of models. Communications in Statistics: Simulation and Computation, 45(5):1613–1620, 2016. CODEN CSSCDB. ISSN 0361-0918.

Gomes:2011:RBL

[GHRM11] M. Ivette Gomes, Lígia Henriques-Rodrigues, and M. Cristina Miranda. Reduced-bias location-invariant extreme value index estimation: A simulation study. Communications in Statistics: Simulation and Computation, 40(3):424–447, 2011. CODEN CSSCDB. ISSN 0361-0918.

Gosho:2014:SWC

[GHY14] Masahiko Gosho, Chikuma Hamada, and Isao Yoshimura. Selection of working correlation structure in weighted generalized estimating equation method for incomplete longitudinal data. Communications in Statistics: Simulation and Computation, 43(1):62–81, 2014. CODEN CSSCDB. ISSN 0361-0918.

Garren:2014:EED

[GK14a] Steven T. Garren and Phillip S. Kott. Evaluating the effective degrees of freedom of the delete-a-group jackknife. Communications in Statistics: Simulation and Computation, 43 (10):2649–2660, 2014. CODEN CSSCDB. ISSN 0361-0918.

Grover:2014:GCR

[GK14b] Lovleen Kumar Grover and Parmdeep Kaur. A generalized class of ratio type exponential estimators of population mean under linear transformation of auxiliary variable. Communications in Statistics: Simulation and Computation, 43(7): 1552–1574, 2014. CODEN CSSCDB. ISSN 0361-0918.

Gorecki:2015:RMC

[GK15] T. Górecki and M. Krzyśko. Regression methods for combining multiple classifiers. Communications in Statistics:

Simulation and Computation, 44(3):739–755, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gadre:2018:TGI

[GK18] M. P. Gadre and V. C. Kakade. Two group inspection-based control charts for dispersion matrix. Communications in Statistics: Simulation and Computation, 47(6):1652–1669, 2018. CODEN CSSCDB. ISSN 0361-0918.

Ghufran:2012:OMS

[GKA12] Shazia Ghufran, Saman Khowaja, and M. J. Ahsan. Optimum multivariate stratified sampling designs with travel cost: A multiobjective integer nonlinear programming approach. Communications in Statistics: Simulation and Computation, 41(5):598–610, 2012. CODEN CSSCDB. ISSN 0361-0918.

Gokpinar:2019:AFM

[GKA19] F. Gökpinar, T. A. Khaniyev, and R. T. Aliyev. Approximation formulas for the moments of the boundary functional of a Gaussian random walk with positive drift by using Siegmund’s formula. Communications in Statistics: Simulation and Computation, 48(9):2679–2688, 2019. CODEN CSSCDB. ISSN 0361-0918.

Gupta:2017:REF

[GKS17] Sat Gupta, Geeta Kalucha, and Javid Shabbir. A regression estimator for finite population mean of a sensitive variable using an optional randomized response model. Communications in Statistics: Simulation and Computation, 46(3): 2393–2405, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gorecki:2010:SMR

[GL10] Tomasz Górecki and Maciej Luczak. Some methods of replacing the nearest neighbor method. Communications in Statistics: Simulation and Computation, 39(2):262–276, 2010. CODEN CSSCDB. ISSN 0361-0918.

Gao:2011:LRR

[GL11] Feng Gao and Xu-Qing Liu. Linearized ridge regression estimator under the mean squared error criterion in a linear regression model. Communications in Statistics: Simulation and Computation, 40(9):1434–1443, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Gorecki:2014:FSD

[GL14] Tomasz Górecki and Maciej Luczak. First and second derivatives in time series classification using DTW. Communications in Statistics: Simulation and Computation, 43(9): 2081–2092, 2014. CODEN CSSCDB. ISSN 0361-0918.

Gob:2015:LSD

[GL15] Rainer Göb and Kristina Lurz. Lot sampling distributions for bounded number of nonconformities. Communications in Statistics: Simulation and Computation, 44(10):2562–2591, 2015. CODEN CSSCDB. ISSN 0361-0918.

Guo:2016:PBO

[GL16] Guangbao Guo and Lu Lin. Parallel bootstrap and optimal subsample lengths in smooth function models. Communications in Statistics: Simulation and Computation, 45(6): 2208–2231, 2016. CODEN CSSCDB. ISSN 0361-0918.

Gonzalez-Lopez:2017:BES

[GLGG17] V. A. Gonzalez-Lopez, Ramin Gholizadeh, and Christian E. Galarza. E-Bayesian estimation for system reliability and availability analysis based on exponential distribution. Communications in Statistics: Simulation and Computation, 46 (8):6221–6241, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gouet:2012:LTC

[GLS12] Raúl Gouet, F. Javier López, and Gerardo Sanz. Limit theorems for the counting process of near-records. Communications in Statistics: Simulation and Computation, 41(6): 820–832, 2012. CODEN CSSCDB. ISSN 0361-0918.

Goyal:2013:MTD

[GMG13] Anju Goyal, Vishal Maurya, and Amar Nath Gill. A multiple three-decision procedures for comparing several treatments with a control under heteroscedasticity. Communications in Statistics: Simulation and Computation, 42(8):1763–1784, 2013. CODEN CSSCDB. ISSN 0361-0918.

Gharibzadeh:2018:CDP

[GMR^{+}18] Safoora Gharibzadeh, Mohammad Ali Mansournia, Abbas Rahimiforoushani, Ahad Alizadeh, Atieh Amouzegar, Kamran Mehrabani-Zeinabad, and Kazem Mohammad. Comparing different propensity score estimation methods for estimating the marginal causal effect through standardization to propensity scores. Communications in Statistics: Simulation and Computation, 47(4):964–976, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Gomes:2015:BOR

[GN15] D. Prata Gomes and M. Manuela Neves. Bootstrap and other resampling methodologies in statistics of extremes. Communications in Statistics: Simulation and Computation, 44(10):2592–2607, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gupta:2017:PPL

[GNG17] Arjun K. Gupta, Daya K. Nagar, and Armando Gómez. Percentage points of LRC for testing equality of covariance matrices under intraclass correlation structure. Communications in Statistics: Simulation and Computation, 46(5): 3855–3869, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gertsbakh:2015:MCE

[GNV15] Ilya Gertsbakh, Eyal Neuman, and Radislav Vaisman. Monte Carlo for estimating exponential convolution. Communications in Statistics: Simulation and Computation, 44 (10):2696–2704, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gomez:2013:SCC

[Góm13] Víctor Gómez. A strongly consistent criterion to decide between I(1) and I(0) processes based on different convergence rates. Communications in Statistics: Simulation and Computation, 42(8):1848–1864, 2013. CODEN CSSCDB. ISSN 0361-0918.

Gorecki:2018:CTS

[Gór18] Tomasz Górecki. Classification of time series using combination of DTW and LCSS dissimilarity measures. Communications in Statistics: Simulation and Computation, 47(1): 263–276, 2018. CODEN CSSCDB. ISSN 0361-0918.

Greselin:2015:PVN

[GP15] Francesca Greselin and Leo Pasquazzi. Parametric versus nonparametric inference on Zenga index of inequality: Issues and evidence from survey data. Communications in Statistics: Simulation and Computation, 44(7):1702–1719, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gauthier:2016:AIO

[GP16] Bertrand Gauthier and Luc Pronzato. Approximation of IMSE-optimal designs via quadrature rules and spectral decomposition. Communications in Statistics: Simulation and Computation, 45(5):1600–1612, 2016. CODEN CSSCDB. ISSN 0361-0918.

Golparvar:2017:SBE

[GP17] Leila Golparvar and Ahmad Parsian. Selecting the best exponential population under Type-II progressive censoring scheme via empirical Bayes approach. Communications in Statistics: Simulation and Computation, 46(1):404–422, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gopinath:2018:TDB

[GPM18] Pratheesh P. Gopinath, Rajender Parsad, and B. N. Mandal. Two-dimensional balanced sampling plans excluding adjacent units under sharing a border and island adjacency schemes. Communications in Statistics: Simulation and Computation, 47(3):712–720, 2018. CODEN CSSCDB. ISSN 0361-0918.

Gao:2018:ACN

[GQYY18] Yue Gao, Zong-Feng Qi, Jian-Feng Yang, and Jin-Yu Yang. Algorithmic construction of nearly column-orthogonal designs. Communications in Statistics: Simulation and Computation, 47(3):924–937, 2018. CODEN CSSCDB. ISSN 0361-0918.

Ghadge:2010:RTT

[GR10] Chanchala Ghadge and T. V. Ramanathan. Rank test for testing randomness of the technology parameters in a stochastic frontier regression model. Communications in Statistics: Simulation and Computation, 40(1):99–112, 2010. CODEN CSSCDB. ISSN 0361-0918.

Gunel:2017:FET

[GR17] Erdogan Gunel and Kenneth Joseph Ryan. Fisher’s exact test from a Bayes perspective. Communications in Statistics: Simulation and Computation, 46(9):7393–7404, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ghapani:2017:MSI

[GRB17] F. Ghapani, A. R. Rasekh, and B. Babadi. Mean shift and influence measures in linear measurement error models with stochastic linear restrictions. Communications in Statistics: Simulation and Computation, 46(6):4499–4512, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Grcar:2017:RTM

[Grc17] Joseph F. Grcar. Research trends in MSC 65 numerical analysis. Communications in Statistics: Simulation and Computation, 46(4):2505–2512, 2017. CODEN CSSCDB. ISSN 0361-0918.

Greco:2017:WLB

[Gre17] Luca Greco. Weighted likelihood based inference for P (X < Y). Communications in Statistics: Simulation and Computation, 46(10):7777–7789, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gallardo:2017:SEP

[GRM17] Diego I. Gallardo, Jose S. Romeo, and Renate Meyer. A simplified estimation procedure based on the EM algorithm for the power series cure rate model. Communications in Statistics: Simulation and Computation, 46(8):6342–6359, 2017. CODEN CSSCDB. ISSN 0361-0918.

Goldoust:2019:LDM

[GRSN19] M. Goldoust, S. Rezaei, Y. Si, and S. Nadarajah. Lifetime distributions motivated by series and parallel structures. Communications in Statistics: Simulation and Computation, 48(2):556–579, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ghafouri:2017:TSP

[GRY17] S. Ghafouri, A. Habibi Rad, and F. Yousefzadeh. Twosample prediction for progressively Type-II censored Weibull lifetimes. Communications in Statistics: Simulation and Computation, 46(2):1381–1400, 2017. CODEN CSSCDB. ISSN 0361-0918.

Grothe:2011:KR

[GS11] Oliver Grothe and Friedrich Schmid. Kendall’s W reconsidered. Communications in Statistics: Simulation and Computation, 40(2):285–305, 2011. CODEN CSSCDB. ISSN 0361-0918.

Gawarammana:2017:CMA

[GS17] M. B. M. B. K. Gawarammana and M. R. Sooriyarachchi. Comparison of methods for analyzing binary repeated measures data: A simulation-based study (comparison of methods for binary repeated measures). Communications in Statistics: Simulation and Computation, 46(3):2103–2120, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gurung:2017:AMF

[GSP^{+}17] Bishal Gurung, K. N. Singh, Ranjit Kumar Paul, Sanjeev Panwar, Biwash Gurung, and Lawrence Lepcha. An alternative method for forecasting price volatility by combining models. Communications in Statistics: Simulation and Computation, 46(6):4627–4636, 2017. CODEN CSSCDB. ISSN 0361-0918.

Gupta:2016:IET

[GSSCR16] Sat Gupta, Javid Shabbir, Rita Sousa, and Pedro Corte-Real. Improved exponential type estimators of the mean of a sensitive variable in the presence of nonsensitive auxiliary information. Communications in Statistics: Simulation and Computation, 45(9):3317–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Guan:2014:OMC

[GTFX14] Qiang Guan, Yincai Tang, Jiayu Fu, and Ancha Xu. Optimal multiple constant-stress accelerated life tests for generalized exponential distribution. Communications in Statistics: Simulation and Computation, 43(8):1852–1865, 2014. CODEN CSSCDB. ISSN 0361-0918.

Giusti:2014:ROQ

[GTPS14] C. Giusti, N. Tzavidis, M. Pratesi, and N. Salvati. Resistance to outliers of M-quantile and robust random effects small area models. Communications in Statistics: Simulation and Computation, 43(3):549–568, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Gui:2017:EHL

[Gui17] Wenhao Gui. Exponentiated half logistic distribution: Different estimation methods and joint confidence regions.

Communications in Statistics: Simulation and Computation, 46(6):4600–4617, 2017. CODEN CSSCDB. ISSN 0361-0918.

Guillaume:2018:CQP

[Gui18] Tristan Guillaume. Computation of the quadrivariate and pentavariate normal cumulative distribution functions. Communications in Statistics: Simulation and Computation, 47(3):839–851, 2018. CODEN CSSCDB. ISSN 0361-0918.

Gunasekera:2017:IFP

[Gun17] Sumith Gunasekera. Inferences on functions of Pareto parameters with application to income inequality measures. Communications in Statistics: Simulation and Computation, 46(2):933–947, 2017. CODEN CSSCDB. ISSN 0361-0918.

Guris:2019:NNU

[Gür19] Burak Güris. A new nonlinear unit root test with Fourier function. Communications in Statistics: Simulation and Computation, 48(10):3056–3062, 2019. CODEN CSSCDB. ISSN 0361-0918.

Gutierrez:2015:UWV

[Gut15] Andres Gutierrez. The use of working variables in the Bayesian modeling of mean and dispersion parameters in generalized nonlinear models with random effects. Communications in Statistics: Simulation and Computation, 44(1): 168–195, 2015. CODEN CSSCDB. ISSN 0361-0918.

Gurevich:2010:RCP

[GV10] Gregory Gurevich and Albert Vexler. Retrospective change point detection: From parametric to distribution free policies. Communications in Statistics: Simulation and Computation, 39(5):899–920, 2010. CODEN CSSCDB. ISSN 0361-0918.

Gauchi:2010:NPI

[GVC10] Jean-Pierre Gauchi, Jean-Pierre Vila, and Louis Coroller. New prediction interval and band in the nonlinear regression model: Application to predictive modeling in foods. Communications in Statistics: Simulation and Computation, 39 (2):322–334, 2010. CODEN CSSCDB. ISSN 0361-0918.

Garza-Venegas:2017:ECP

[GVTGF^{+}17] Jorge Garza-Venegas, Victor Tercero-Gomez, Alvaro Cordero Franco, María Temblador-Pérez, and Mario Beruvides. An evaluation of change-point estimators for a sequence of normal observations with unknown parameters. Communications in Statistics: Simulation and Computation, 46(6): 4297–4317, 2017. CODEN CSSCDB. ISSN 0361-0918.

Guo:2019:IPR

[GY19] Jianfeng Guo and Jiameng Yang. An iterative procedure for robust circle fitting. Communications in Statistics: Simulation and Computation, 48(6):1872–1879, 2019. CODEN CSSCDB. ISSN 0361-0918.

Gao:2010:NMV

[GZ10] Chao Gao and Caiya Zhang. A novel method for visualization of clustering results. Communications in Statistics: Simulation and Computation, 39(5):1049–1056, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Hosseinifard:2012:PCA

[HA12] S. Z. Hosseinifard and B. Abbasi. Process capability analysis in non normal linear regression profiles. Communications in Statistics: Simulation and Computation, 41(10):1761–1784, 2012. CODEN CSSCDB. ISSN 0361-0918.

Harandi:2017:CGL

[HA17a] S. Shams Harandi and M. H. Alamatsaz. A complementary generalized linear failure rate-geometric distribution. Communications in Statistics: Simulation and Computation, 46 (1):679–703, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hernandez:2017:NAI

[HA17b] M. Álvarez Hernández and A. Martín Andrés. New asymptotic inferences about the difference, ratio, and linear combination of two independent proportions. Communications in Statistics: Simulation and Computation, 46(2):1557–1568, 2017. CODEN CSSCDB. ISSN 0361-0918.

Habibi:2018:PBE

[HA18] M. Habibi and A. Asgharzadeh. Power binomial exponential distribution: Modeling, simulation and application. Communications in Statistics: Simulation and Computation, 47 (10):3042–3061, 2018. CODEN CSSCDB. ISSN 0361-0918.

Huszti:2011:CSM

[HAAQ11] Ella Huszti, Michal Abrahamowicz, Ahmadou Alioum, and Catherine Quantin. Comparison of selected methods for modeling of multi-state disease progression processes: A simulation study. Communications in Statistics: Simulation and Computation, 40(9):1402–1421, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Haseen:2016:MSM

[HAB16] Sanam Haseen, Irfan Ali, and Abdul Bari. Multiobjective stochastic multivariate stratified sampling in presence of nonresponse. Communications in Statistics: Simulation and Computation, 45(8):2810–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Han:2019:BEM

[Han19] Ming Han. E-Bayesian estimation and its E-MSE under the scaled squared error loss function, for exponential distribution as example. Communications in Statistics: Simulation and Computation, 48(6):1880–1890, 2019. CODEN CSSCDB. ISSN 0361-0918.

Haq:2019:NNS

[Haq19] Abdul Haq. A new nonparametric synthetic EWMA control chart for monitoring process mean. Communications in Statistics: Simulation and Computation, 48(6):1665–1676, 2019. CODEN CSSCDB. ISSN 0361-0918.

Harlamov:2012:SMG

[Har12] B. P. Harlamov. Stochastic model of gas capillary chromatography. Communications in Statistics: Simulation and Computation, 41(7):1023–1031, 2012. CODEN CSSCDB. ISSN 0361-0918.

Han:2014:SRM

[HAT14] Seungbong Han, Adin-Cristian Andrei, and Kam-Wah Tsui. A semiparametric regression method for interval-censored data. Communications in Statistics: Simulation and Computation, 43(1):18–30, 2014. CODEN CSSCDB. ISSN 0361-

Han:2015:EST

[HAT15] Seungbong Han, Adin-Cristian Andrei, and Kam-Wah Tsui. Efficient stratified testing procedure for a false discovery rate. Communications in Statistics: Simulation and Computation, 44(5):1117–1125, 2015. CODEN CSSCDB. ISSN 0361-0918.

Hayes:2014:FSB

[Hay14] Kevin Hayes. Finite-sample bias-correction factors for the median absolute deviation. Communications in Statistics: Simulation and Computation, 43(10):2205–2212, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Hussain:2016:UNB

[HAZSK16] Zawar Hussain, Bander Al-Zahrani, Javid Shabbir, and Manzoor Khan. On using negative binomial distribution as a randomization device in sensitive surveys. Communications in Statistics: Simulation and Computation, 45(10): 3584–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Huda:2010:RSC

[HB10] S. Huda and L. Benkherouf. Rotatability is a sufficient condition for A- and D-rotatability. Communications in Statistics: Simulation and Computation, 39(6):1174–1182, 2010. CODEN CSSCDB. ISSN 0361-0918.

Han:2015:BBC

[HB15] L. Spectrum Han and W. John Braun. Block bootstrap calibration with application to the fire weather index. Communications in Statistics: Simulation and Computation, 44(3): 647–665, 2015. CODEN CSSCDB. ISSN 0361-0918.

Huda:2016:STT

[HB16] S. Huda and L. Benkherouf. On some two- and threedimensional d-minimax designs for estimating slopes of a third-order response surface. Communications in Statistics: Simulation and Computation, 45(6):1885–1895, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Helian:2017:SCC

[HBC17] Shanjun Helian, Babette A. Brumback, and Robert L. Cook. Sparse canonical correlation analysis between an alcohol biomarker and self-reported alcohol consumption. Communications in Statistics: Simulation and Computation, 46(10): 7924–7941, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hsieh:2017:TST

[HC17a] Jin-Jian Hsieh and Hsin-Yu Chen. A testing strategy for two crossing survival curves. Communications in Statistics: Simulation and Computation, 46(8):6685–6696, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Hung:2017:SSM

[HC17b] Ying-Chao Hung and Wei-Cheng Chen. Simulation of some multivariate distributions related to the Dirichlet distribution with application to Monte Carlo simulations. Communications in Statistics: Simulation and Computation, 46(6): 4281–4296, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hendrych:2018:SWR

[HC18] Radek Hendrych and Tomás Cipra. Self-weighted recursive estimation of GARCH models. Communications in Statistics: Simulation and Computation, 47(2):315–328, 2018. CODEN CSSCDB. ISSN 0361-0918.

Haghighi:2019:CCM

[HC19] Firoozeh Haghighi and Philippe Castagliola. Control chart for monitoring the Weibull shape parameter under two competing risks. Communications in Statistics: Simulation and Computation, 48(7):2125–2137, 2019. CODEN CSSCDB. ISSN 0361-0918.

Hsiao:2017:SFS

[HCW17] Tien-Mu Hsiao, Huifen Chen, and Chien-Hua Wu. Statistical frameworks for setting cutoff points of metabolic syndrome criteria. Communications in Statistics: Simulation and Computation, 46(7):5666–5681, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Hanagal:2013:BEP

[HD13a] David D. Hanagal and Alok D. Dabade. Bayesian estimation of parameters and comparison of shared gamma frailty models. Communications in Statistics: Simulation and Computation, 42(4):910–931, 2013. CODEN CSSCDB. ISSN 0361-

Huang:2013:CME

[HD13b] Yangxin Huang and Getachew Dagne. Comparison of mixedeffects models for skew-normal responses with an application to AIDS data: A Bayesian approach. Communications in Statistics: Simulation and Computation, 42(6):1268–1287, 2013. CODEN CSSCDB. ISSN 0361-0918.

Huang:2017:SEP

[HD17] Zhensheng Huang and Haiying Ding. Statistical estimation for partially linear error-in-variable models with error-prone covariates. Communications in Statistics: Simulation and Computation, 46(8):6559–6573, 2017. CODEN CSSCDB. ISSN 0361-0918.

Haggstrom:2010:EPE

[HdL10] Jenny Häggström and Xavier de Luna. Estimating prediction error: Cross-validation vs. accumulated prediction error. Communications in Statistics: Simulation and Computation, 39(5):880–898, 2010. CODEN CSSCDB. ISSN 0361-0918.

Hartmann:2017:BIG

[HE17] Marcelo Hartmann and Ricardo S. Ehlers. Bayesian inference for generalized extreme value distributions via Hamiltonian Monte Carlo. Communications in Statistics: Simulation and Computation, 46(7):5285–5302, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hamed:2013:SRP

[HEF13] Ramadan Hamed, Ali El Hefnawy, and Aya Farag. Selection of the ridge parameter using mathematical programming. Communications in Statistics: Simulation and Computation, 42(6):1409–1432, 2013. CODEN CSSCDB. ISSN 0361-0918.

Hanusz:2018:MCC

[HESK18] Zofia Hanusz, Rie Enomoto, Takashi Seo, and Kazuyuki Koizumi. A Monte Carlo comparison of Jarque–Bera type tests and Henze–Zirkler test of multivariate normality. Communications in Statistics: Simulation and Computation, 47 (5):1439–1452, 2018. CODEN CSSCDB. ISSN 0361-0918.

Hajrajabi:2018:NSA

[HF18] Arezo Hajrajabi and Afshin Fallah. Nonlinear semiparametric AR(1) model with skew-symmetric innovations. Communications in Statistics: Simulation and Computation, 47(5): 1453–1462, 2018. CODEN CSSCDB. ISSN 0361-0918.

Herschtal:2016:CBM

[HFKM16] Alan Herschtal, Farshad Foroudi, Tomas Kron, and Kerrie Mengersen. A comparison of Bayesian models of heteroscedasticity in nested normal data. Communications in Statistics: Simulation and Computation, 45(8):2947–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Hossain:2016:BAN

[HHH16] Shahadut Hossain, Zahirul Hoque, and A. H. M. Saidul Hasan. A better alternative to non-parametric approaches for adjusting for covariate measurement errors in logistic regression. Communications in Statistics: Simulation and Computation, 45(8):2659–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Hua:2010:MFT

[HHHZ10] Steven Y. Hua, Doyle L. Hawkins, Chien-Pai Han, and Yingqi Zhao. Multivariate failure time analysis when only the time to first failure is observed. Communications in Statistics: Simulation and Computation, 40(1):113–128, 2010. CODEN CSSCDB. ISSN 0361-0918.

Hlavka:2016:BPO

[HHKM16] Z. Hlávka, M. Husková, C. Kirch, and S. G. Meintanis. Bootstrap procedures for online monitoring of changes in autoregressive models. Communications in Statistics: Simulation and Computation, 45(7):2471–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Herbert:2011:IED

[HHMW11] Robert D. Herbert, Andrew Hayen, Petra Macaskill, and Stephen D. Walter. Interval estimation for the difference of two independent variances. Communications in Statistics: Simulation and Computation, 40(5):744–758, 2011. CODEN CSSCDB. ISSN 0361-0918.

Hasan:2016:CSO

[HHS16] M. Tariqul Hasan, Shahariar Huda, and Gary Sneddon. A comparative study of observation- and parameter-driven zero-inflated Poisson models for longitudinal count data. Communications in Statistics: Simulation and Computation, 45(10):3643–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Hilafu:2017:RSI

[Hil17] Haileab Hilafu. Random sliced inverse regression. Communications in Statistics: Simulation and Computation, 46(5): 3516–3526, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hirano:2017:MLP

[Hir17] Toshihiro Hirano. Modified linear projection for large spatial datasets. Communications in Statistics: Simulation and Computation, 46(2):870–889, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hou:2018:PEU

[HJ18] Jian Hou and Bo Jiang. Predictions and estimations under a group of linear models with random coefficients. Communications in Statistics: Simulation and Computation, 47(2): 510–525, 2018. CODEN CSSCDB. ISSN 0361-0918.

Han:2015:CLB

[HJST15] Sung Won Han, Wei Jiang, Lianjie Shu, and Kwok-Leung Tsui. A comparison of likelihood-based spatiotemporal monitoring methods under non-homogeneous population size. Communications in Statistics: Simulation and Computation, 44(1):14–39, 2015. CODEN CSSCDB. ISSN 0361-0918.

Hong:2010:SPG

[HK10] Chong Sun Hong and Jong-Min Kim. The SSR plot: A graphical representation for regression. Communications in Statistics: Simulation and Computation, 39(4):726–735, 2010. CODEN CSSCDB. ISSN 0361-0918.

Hemmati:2013:SAL

[HK13] Fariba Hemmati and Esmaile Khorram. Statistical analysis of the log-normal distribution under Type-II progressive hybrid censoring schemes. Communications in Statistics:

Simulation and Computation, 42(1):52–75, 2013. CODEN CSSCDB. ISSN 0361-0918.

Hemmati:2017:APT

[HK17] Fariba Hemmati and Esmaile Khorram. On adaptive progressively Type-II censored competing risks data. Communications in Statistics: Simulation and Computation, 46(6): 4671–4693, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hajjar:2018:BNS

[HK18a] S. Bohlouri Hajjar and S. Khazaei. Bayesian nonparametric survival analysis using mixture of Burr XII distributions. Communications in Statistics: Simulation and Computation, 47(9):2724–2738, 2018. CODEN CSSCDB. ISSN 0361-0918.

Hoxhaj:2018:SGS

[HK18b] Valmira Hoxhaj and Ravindra Khattree. A study of geometric and statistical curvatures for the skew-normal family of distributions. Communications in Statistics: Simulation and Computation, 47(7):2010–2022, 2018. CODEN CSSCDB. ISSN 0361-0918.

Hoffman:2010:EPJ

[HKC10] Lorrie L. Hoffman, Greg Knofczynski, and Steve Clark. Estimation of parameters in the joinpoint two-regime regression model. Communications in Statistics: Simulation and Computation, 39(8):1562–1576, 2010. CODEN CSSCDB. ISSN 0361-0918.

Hussein:2017:PRA

[HKNC17] A. Hussein, A. Kasem, S. Nkurunziza, and S. Campostrini. Performance of risk-adjusted cumulative sum charts when some assumptions are not met. Communications in Statistics: Simulation and Computation, 46(2):823–830, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hossain:2018:NWL

[HKP18] Md. Mobarak Hossain, Tomasz J. Kozubowski, and Krzysztof Podgórski. A novel weighted likelihood estimation with empirical Bayes flavor. Communications in Statistics: Simulation and Computation, 47(2):392–412, 2018. CODEN CSSCDB. ISSN 0361-0918.

He:2014:CCC

[HLH14] Shuguang He, Shijie Li, and Zhen He. A combination of CUSUM charts for monitoring a zero-inflated Poisson process. Communications in Statistics: Simulation and Computation, 43(10):2482–2497, 2014. CODEN CSSCDB. ISSN 0361-0918.

Heuchenne:2014:LBI

[HLLV14] Cédric Heuchenne, Stéphane Laurent, Catherine Legrand, and Ingrid Van Keilegom. Likelihood-based inference for semi-competing risks. Communications in Statistics: Simulation and Computation, 43(5):1112–1132, 2014. CODEN CSSCDB. ISSN 0361-0918.

Han:2017:REE

[HLWH17] Zhong-Cheng Han, Jin-Guan Lin, Hong-Xia Wang, and Xing-Fang Huang. A robust and efficient estimation method for nonparametric models with jump points. Communications in Statistics: Simulation and Computation, 46(8): 6283–6297, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hlynka:2017:MFD

[Hly17] Myron Hlynka. MCMC and the Fibonacci distribution. Communications in Statistics: Simulation and Computation, 46(5):3375–3382, 2017. CODEN CSSCDB. ISSN 0361-0918.

Han:2015:CSS

[HLZK15] Sung Won Han, Kyu Jong Lee, Hua Zhong, and Seoung Bum Kim. A comparison of spatiotemporal surveillance methods for nonhomogeneous change size. Communications in Statistics: Simulation and Computation, 44(10):2714–2730, 2015. CODEN CSSCDB. ISSN 0361-0918.

He:2012:AIS

[HM12a] Jianghua He and Matthew S. Mayo. Adjusted interim survival analysis under nonproportional hazards. Communications in Statistics: Simulation and Computation, 41(1): 111–124, 2012. CODEN CSSCDB. ISSN 0361-0918.

Huskova:2012:PDC

[HM12b] Marie Husková and Miriam Marusiaková. M-procedures for detection of changes for dependent observations. Communications in Statistics: Simulation and Computation, 41(7): 1032–1050, 2012. CODEN CSSCDB. ISSN 0361-0918.

Huskova:2012:TSE

[HM12c] Marie Husková and Simos G. Meintanis. Tests for symmetric error distribution in linear and nonparametric regression models. Communications in Statistics: Simulation and Computation, 41(6):833–851, 2012. CODEN CSSCDB. ISSN 0361-0918.

Holgersson:2013:ANH

[HM13] H. E. T. Holgersson and Rashid Mansoor. Assessing normality of high-dimensional data. Communications in Statistics: Simulation and Computation, 42(2):360–369, 2013. CODEN CSSCDB. ISSN 0361-0918.

Heinzl:2017:AHT

[HM17] Harald Heinzl and Martina Mittlboeck. Assessing a hypothesis test for the difference between two quantiles from independent populations. Communications in Statistics: Simulation and Computation, 46(5):3540–3552, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hussain:2010:SME

[HMH^{+}10] Shakir Hussain, Mohammed A. Mohammed, M. Sayeed Haque, Roger Holder, John Macleod, and Richard Hobbs. A simple method to ensure plausible multiple imputation for continuous multivariate data. Communications in Statistics: Simulation and Computation, 39(9):1779–1784, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Holgersson:2016:TPU

[HMS16] H. E. T. Holgersson, K. Månsson, and G. Shukur. Testing for panel unit roots under general cross-sectional dependence. Communications in Statistics: Simulation and Computation, 45(5):1785–1801, 2016. CODEN CSSCDB. ISSN 0361-0918.

Haghbin:2013:LBI

[HN13] H. Haghbin and A. R. Nematollahi. Likelihood-based inference in autoregressive models with scaled t-distributed innovations by means of EM-based algorithms. Communications in Statistics: Simulation and Computation, 42(10): 2239–2252, 2013. CODEN CSSCDB. ISSN 0361-0918.

Howe:2015:GAW

[HN15] Eylem Deniz Howe and Orietta Nicolis. Genetic algorithm in the wavelet domain for large p small n regression. Communications in Statistics: Simulation and Computation, 44 (5):1144–1157, 2015. CODEN CSSCDB. ISSN 0361-0918.

Hermans:2018:FCF

[HNM^{+}18] Lisa Hermans, Vahid Nassiri, Geert Molenberghs, Michael G. Kenward, Wim Van der Elst, Marc Aerts, and Geert Verbeke. Fast, closed-form, and efficient estimators for hierarchical models with AR(1) covariance and unequal cluster sizes. Communications in Statistics: Simulation and Computation, 47(5):1492–1505, 2018. CODEN CSSCDB. ISSN 0361-0918.

Hosseini-Nasab:2012:PPB

[HNV12] S. Mohammad E. Hosseini-Nasab and Masoumeh Vakili. On properties of percentile bootstrap confidence intervals for period using periodogram. Communications in Statistics: Simulation and Computation, 41(3):316–326, 2012. CODEN CSSCDB. ISSN 0361-0918.

Hatjispyros:2017:BND

[HNW17] Spyridon J. Hatjispyros, Theodoros Nicoleris, and Stephen G. Walker. Bayesian nonparametric density estimation under length bias. Communications in Statistics: Simulation and Computation, 46(10):8064–8076, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hodoshima:2013:EEU

[Hod13] Jiro Hodoshima. The effect of estimation of unknown degrees of freedom on estimation of remaining parameters in Spanos’ conditional t heteroskedastic model. Communications in Statistics: Simulation and Computation, 42(8):1727–1749, 2013. CODEN CSSCDB. ISSN 0361-0918.

Holgersson:2010:MSM

[Hol10] H. E. T. Holgersson. A modified skewness measure for testing asymmetry. Communications in Statistics: Simulation and Computation, 39(2):335–346, 2010. CODEN CSSCDB. ISSN 0361-0918.

Hou:2016:MAG

[Hou16] Fujun Hou. A multiplicative Alo-group based hierarchical decision model and application. Communications in Statistics: Simulation and Computation, 45(8):2846–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Hanagal:2017:CGF

[HPG17] David D. Hanagal, Arvind Pandey, and Ayon Ganguly. Correlated gamma frailty models for bivariate survival data. Communications in Statistics: Simulation and Computation, 46(5):3627–3644, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hanagal:2017:SFM

[HPS17] David D. Hanagal, Arvind Pandey, and P. G. Sankaran. Shared frailty model based on reversed hazard rate for left censored data. Communications in Statistics: Simulation and Computation, 46(1):230–243, 2017. CODEN CSSCDB. ISSN 0361-0918.

Henry:2014:NEH

[HRS14] Guillermo Henry, Daniela Rodriguez, and Mariela Sued. A note on the estimation of the hölder constant. Communications in Statistics: Simulation and Computation, 43(4): 905–912, 2014. CODEN CSSCDB. ISSN 0361-0918.

Hashemiparast:2012:TDA

[HS12] S. M. Hashemiparast and M. Sabzevari. The twodimensional asymmetric Gambler’s ruin problem with n players and simulations. Communications in Statistics: Simulation and Computation, 41(10):1895–1911, 2012. CODEN CSSCDB. ISSN 0361-0918.

Hor:2017:MVU

[HS17] Parimal Hor and Babulal Seal. Minimum variance unbiased estimation of stress-strength reliability under bivariate normal and its comparisons. Communications in Statistics: Simulation and Computation, 46(3):2447–2456, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Hayakawa:2019:IVE

[HS19a] Kazuhiko Hayakawa and Qi Sun. Instrumental variable estimation of factor models with possibly many variables. Communications in Statistics: Simulation and Computation, 48 (6):1729–1745, 2019. CODEN CSSCDB. ISSN 0361-0918.

Hor:2019:SR

[HS19b] Parimal Hor and Babulal Seal. SAVE: Robust or not? Communications in Statistics: Simulation and Computation, 48 (10):2855–2865, 2019. CODEN CSSCDB. ISSN 0361-0918.

Hussain:2017:GGD

[HSP^{+}17] Zawar Hussain, Javid Shabbir, Zahid Pervez, Said Farooq Shah, and Manzoor Khan. Generalized geometric distribution of order k: A flexible choice to randomize the response. Communications in Statistics: Simulation and Computation, 46(6):4708–4721, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hussain:2010:BEU

[HSR10] Zawar Hussain, Javid Shabbir, and Muhammad Riaz. Bayesian estimation using Warner’s randomized response model through simple and mixture prior distributions. Communications in Statistics: Simulation and Computation, 40 (1):147–164, 2010. CODEN CSSCDB. ISSN 0361-0918.

Hwang:2017:CNM

[HST17] Yi-Ting Hwang, Yu-Hui Su, and Harn-Jing Terng. Comparisons of normalization methods for relative quantitative in real-time polymerase chain reaction. Communications in Statistics: Simulation and Computation, 46(2):1336–1359, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hanusz:2014:SSI

[HT14] Zofia Hanusz and Joanna Tarasińska. Simulation study on improved Shapiro–Wilk tests for normality. Communications in Statistics: Simulation and Computation, 43(9): 2093–2105, 2014. CODEN CSSCDB. ISSN 0361-0918.

Hayashi:2017:FSA

[HT17] Kenichi Hayashi and Keiji Takai. Finite-sample analysis of impacts of unlabeled data and their labeling mechanisms in linear discriminant analysis. Communications in Statistics: Simulation and Computation, 46(1):184–203, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hyodo:2014:MCA

[HTN14] Masashi Hyodo, Sho Takahashi, and Takahiro Nishiyama. Multiple comparisons among mean vectors when the dimension is larger than the total sample size. Communications in Statistics: Simulation and Computation, 43(10):2283–2306, 2014. CODEN CSSCDB. ISSN 0361-0918.

Huang:2014:TAE

[Hua14] Zhensheng Huang. Two adjusted empirical-likelihood-based methods in generalized varying-coefficient partially linear model. Communications in Statistics: Simulation and Computation, 43(10):2252–2263, 2014. CODEN CSSCDB. ISSN 0361-0918.

Huang:2016:ELB

[Hua16] Zhensheng Huang. Empirical-likelihood-based test for partially linear single-index models with error-prone linear covariates. Communications in Statistics: Simulation and Computation, 45(7):2638–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Huang:2017:AWS

[Hua17] Li-Fei Huang. Adjusted Wilcoxon signed rank test tables for ratio of percentiles. Communications in Statistics: Simulation and Computation, 46(7):5763–5771, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hinloopen:2014:EFS

[HW14] Jeroen Hinloopen and Rien J. L. M. Wagenvoort. On the exact finite sample distribution of the L_{1}-FCvM test statistic. Communications in Statistics: Simulation and Computation, 43(6):1304–1317, 2014. CODEN CSSCDB. ISSN 0361-0918.

Hepworth:2017:RRE

[HW17a] Graham Hepworth and Ray Watson. Revisiting retesting in the estimation of proportions by group testing. Communications in Statistics: Simulation and Computation, 46(1): 261–274, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hsu:2017:SAF

[HW17b] Yuting Hsu and Qi Wang. Statistical analysis of firm interdependence using duration data-parametric approach. Communications in Statistics: Simulation and Computation, 46 (2):1292–1301, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hu:2018:NOT

[HW18] Sigui Hu and Honglei Wang. Nearly optimal truncated group sequential test on binomial proportions. Communications in Statistics: Simulation and Computation, 47(8):2332–2342, 2018. CODEN CSSCDB. ISSN 0361-0918.

Hwang:2017:CSQ

[Hwa17] Wook-Yeon Hwang. Chi-square quantile-based multivariate variance monitoring for individual observations. Communications in Statistics: Simulation and Computation, 46(7): 5392–5409, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hossain:2013:IMW

[HWB13] Ahmed Hossain, Andrew R. Willan, and Joseph Beyene. An improved method on Wilcoxon rank sum test for gene selection from microarray experiments. Communications in Statistics: Simulation and Computation, 42(7):1563–1577, 2013. CODEN CSSCDB. ISSN 0361-0918.

Hauenstein:2018:CAB

[HWD18] Severin Hauenstein, Simon N. Wood, and Carsten F. Dormann. Computing AIC for black-box models using generalized degrees of freedom: a comparison with cross-validation. Communications in Statistics: Simulation and Computation, 47(5):1382–1396, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wei:2013:REP

[hWjJsH13] Chuan hua Wei, Xu jie Jia, and Hong sheng Hu. Restricted estimation in partially linear errors-in-variables models. Communications in Statistics: Simulation and Computation, 42(8):1836–1847, 2013. CODEN CSSCDB. ISSN 0361-0918.

Huang:2018:VSC

[HWW18] Jun Huang, Haibo Wang, and Wei Wang. Variable selection in classification model via quadratic programming. Communications in Statistics: Simulation and Computation, 47(7): 1922–1939, 2018. CODEN CSSCDB. ISSN 0361-0918.

He:2016:CTP

[HWX16] Daojiang He, Yan Wu, and Kai Xu. A class of s − K type principal components estimators in the linear model. Communications in Statistics: Simulation and Computation, 45 (8):2709–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Hamid:2018:IPG

[HWXH18] Hamzah Abdul Hamid, Yap Bee Wah, Xian-Jin Xie, and Ong Seng Huat. Investigating the power of goodness-of-fit tests for multinomial logistic regression. Communications in Statistics: Simulation and Computation, 47(4):1039–1055, 2018. CODEN CSSCDB. ISSN 0361-0918.

Hao:2018:GAB

[HWZ18] Wu Hao, Wang Weifeng, and Guo Zhongkai. Generalized anticipated backward stochastic differential equations driven by Brownian motion and continuous increasing process. Communications in Statistics: Simulation and Computation, 47(3):809–821, 2018. CODEN CSSCDB. ISSN 0361-0918.

Huang:2013:CMC

[HXB13] Xuan Huang, Nuo Xu, and Søren Bisgaard. A class of Markov chain models for average run length computations for autocorrelated processes. Communications in Statistics: Simulation and Computation, 42(7):1495–1513, 2013. CO-DEN CSSCDB. ISSN 0361-0918.

He:2015:RWE

[HXX15] Daojiang He, Kai Xu, and Xingzhong Xu. Random weighting empirical distribution function and its applications to goodness-of-fit testing. Communications in Statistics: Simulation and Computation, 44(6):1441–1452, 2015. CODEN CSSCDB. ISSN 0361-0918.

Huang:2018:MIE

[HYS18] Yisong Huang, Jingjing Yin, and Hani Samawi. Methods improving the estimate of diagnostic odds ratio. Communications in Statistics: Simulation and Computation, 47(2): 353–366, 2018. CODEN CSSCDB. ISSN 0361-0918.

Han:2019:NMM

[HYZ19] Liyan Han, Hanhuan Yan, and Chengli Zheng. Normal mixture method for stock daily returns over different subperiods. Communications in Statistics: Simulation and Computation, 48(2):447–457, 2019. CODEN CSSCDB. ISSN 0361-0918.

Huang:2019:IVB

[HZ19] Jiting Huang and Peixin Zhao. Instrumental variable based variable selection for generalized linear models with endogenous covariates. Communications in Statistics: Simulation and Computation, 48(6):1891–1900, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Han:2017:EBA

[HZHK17] Shengtong Han, Hongmei Zhang, Ramin Homayouni, and Wilfried Karmaus. An efficient Bayesian approach for Gaussian Bayesian network structure learning. Communications in Statistics: Simulation and Computation, 46(7):5070– 5084, 2017. CODEN CSSCDB. ISSN 0361-0918.

Hendry:2017:EMA

[HZNN17] Gillian M. Hendry, Temesgen Zewotir, Rajen N. Naidoo, and Delia North. The effect of the mechanism and amount of missingness on subset correspondence analysis. Communications in Statistics: Simulation and Computation, 46(9): 7100–7115, 2017. CODEN CSSCDB. ISSN 0361-0918.

Han:2015:EOC

[HZP15] Sung Won Han, Hua Zhong, and Mary Putt. An efficient operator for the change point estimation in partial spline model. Communications in Statistics: Simulation and Computation, 44(5):1171–1186, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2014:QRE

[hZxNL14] Ying hui Zhou, Zhong xin Ni, and Yong Li. Quantile regression via the EM algorithm. Communications in Statistics: Simulation and Computation, 43(10):2162–2172, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Iannario:2016:TOC

[Ian16] Maria Iannario. Testing overdispersion in CUBE models. Communications in Statistics: Simulation and Computation, 45(5):1621–1635, 2016. CODEN CSSCDB. ISSN 0361-

Inan:2013:LTL

[IE13] Deniz Inan and Birsen E. Erdogan. Liu-type logistic estimator. Communications in Statistics: Simulation and Computation, 42(7):1578–1586, 2013. CODEN CSSCDB. ISSN 0361-0918.

Imada:2011:MTS

[Ima11] T. Imada. Multivariate two-sided tests for normal mean vectors based on approximations of likelihood ratio test. Communications in Statistics: Simulation and Computation, 40 (6):894–906, 2011. CODEN CSSCDB. ISSN 0361-0918.

Imada:2013:MTS

[Ima13] T. Imada. Multivariate two-sided tests for normal mean vectors with unknown covariance matrix. Communications in Statistics: Simulation and Computation, 42(1):183–201, 2013. CODEN CSSCDB. ISSN 0361-0918.

Ikeda:2014:TSR

[IMS14] Shu Ikeda, Shun Matsuura, and Hideo Suzuki. Two-step residual-based estimation of error variances for generalized least squares in split-plot experiments. Communications in Statistics: Simulation and Computation, 43(2):342–358, 2014. CODEN CSSCDB. ISSN 0361-0918.

Iida:2018:BEM

[IMS18] Masanari Iida, Yoichi Miyata, and Takayuki Shiohama. Bootstrap estimation and model selection for multivariate normal mixtures using parallel computing with graphics processing units. Communications in Statistics: Simulation and Computation, 47(5):1326–1342, 2018. CODEN CSSCDB. ISSN 0361-0918.

Idjis:2017:VRM

[IOTBO17] Khelidja Idjis, Megdouda Ourbih-Tari, and Latifa Baghdali-Ourbih. Variance reduction in M/M/1 retrial queues using refined descriptive sampling. Communications in Statistics: Simulation and Computation, 46(6):5002–5020, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Iqbal:2011:WLE

[Iqb11] Farhat Iqbal. A weighted linear estimator of multivariate ARCH parameters. Communications in Statistics: Simulation and Computation, 40(4):544–560, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Ibrahim:2013:ODC

[IRHM13] S. Ibrahim, A. Rambli, A. G. Hussin, and I. Mohamed. Outlier detection in a circular regression model using COVRA-TIO statistic. Communications in Statistics: Simulation and Computation, 42(10):2272–2280, 2013. CODEN CSSCDB. ISSN 0361-0918.

Islam:2011:TPS

[Isl11] M. Shahidul Islam. Testing periodicity in short series and application to gene expression data. Communications in Statistics: Simulation and Computation, 40(4):561–573, 2011. CODEN CSSCDB. ISSN 0361-0918.

Intarapak:2017:CEI

[ISS17] Sukanya Intarapak, Rawee Suwandechochai, and Thidaporn Supapakorn. Comparison of the estimators of the intracluster correlation for the nested error regression model. Communications in Statistics: Simulation and Computation, 46(3):2057–2070, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ivanescu:2018:FFR

[Iva18] Andrada E. Ivanescu. Function-on-function regression for two-dimensional functional data. Communications in Statistics: Simulation and Computation, 47(9):2656–2669, 2018. CODEN CSSCDB. ISSN 0361-0918.

Jaheen:2011:BEE

[JA11] Zeinhum F. Jaheen and Mashail M. Al Harbi. Bayesian estimation for the exponentiated Weibull model via Markov chain Monte Carlo simulation. Communications in Statistics: Simulation and Computation, 40(4):532–543, 2011. CODEN CSSCDB. ISSN 0361-0918.

Jafari:2015:IMT

[JA15] A. A. Jafari and K. Abdollahnezhad. Inferences on the means of two log-normal distributions: A computational approach test. Communications in Statistics: Simulation and Computation, 44(7):1659–1672, 2015. CODEN CSSCDB. ISSN 0361-0918.

Jafari:2017:TEM

[JA17] Ali Akbar Jafari and Kamel Abdollahnezhad. Testing the equality means of several log-normal distributions. Communications in Statistics: Simulation and Computation, 46(3): 2311–2320, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jung:2018:MRM

[JCL18] Byoung Cheol Jung, Sooyoung Cheon, and Hwa Kyung Lim. Mixtures of regression models with incomplete and noisy data. Communications in Statistics: Simulation and Computation, 47(2):444–463, 2018. CODEN CSSCDB. ISSN 0361-0918.

Jia:2017:RGE

[JG17a] Chao Jia and Yonghui Ge. A resisting gross errors capability study of robust estimation of unary linear regression method. Communications in Statistics: Simulation and Computation, 46(2):815–822, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jia:2017:EIE

[JG17b] Xiang Jia and Bo Guo. Exact inference for exponential distribution with multiply Type-I censored data. Communications in Statistics: Simulation and Computation, 46(9): 7210–7220, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jianjun:2014:DFA

[JH14] Wu Jianjun and Sun Huijun. Detrended fluctuation analysis of mixed traffic flow time series with a blockage. Communications in Statistics: Simulation and Computation, 43(3): 487–494, 2014. CODEN CSSCDB. ISSN 0361-0918.

Jin:2011:AFT

[Jin11] Man Jin. Analysis of failure time data with mixed-effects accelerated failure time model. Communications in Statistics: Simulation and Computation, 40(4):614–619, 2011. CODEN CSSCDB. ISSN 0361-0918.

Joshi:2017:ECP

[JJB17] Savitri Joshi, K. K. Jose, and Deepesh Bhati. Estimation of a change point in the hazard rate of Lindley model under right censoring. Communications in Statistics: Simulation and Computation, 46(5):3563–3574, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jadhav:2016:RLR

[JK16] N. H. Jadhav and D. N. Kashid. Robust linearized ridge Mestimator for linear regression model. Communications in Statistics: Simulation and Computation, 45(3):1001–1024, 2016. CODEN CSSCDB. ISSN 0361-0918.

Jafari:2017:CAT

[JK17a] A. A. Jafari and M. R. Kazemi. Computational approach test for inference about several correlation coefficients: Equality and common. Communications in Statistics: Simulation and Computation, 46(3):2043–2056, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jana:2017:CTN

[JK17b] Nabakumar Jana and Somesh Kumar. Classification into two normal populations with a common mean and unequal variances. Communications in Statistics: Simulation and Computation, 46(1):546–558, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jin:2017:CSG

[JK17c] Renhao Jin and Gabrielle E. Kelly. A comparison of sampling grids, cut-off distance and type of residuals in parametric variogram estimation. Communications in Statistics: Simulation and Computation, 46(3):1781–1795, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Junior:2016:MSP

[JKBR16] Hélcio Vieira Junior, Karl Heinz Kienitz, and Mischel Carmen Neyra Belder-Rain. Multinomial selection in the presence of infinite alternatives. Communications in Statistics: Simulation and Computation, 45(5):1731–1747, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Jones:2016:ECN

[JKKR16] Daniel Jones, Michael Kohler, Adam Krzyzak, and Alexander Richter. Empirical comparison of nonparametric regression estimates on real data. Communications in Statistics: Simulation and Computation, 45(7):2309–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Jose:2012:CIP

[JL12] K. K. Jose and Jane A. Luke. On confidence intervals for process capability indices in a one-way random model. Communications in Statistics: Simulation and Computation, 41 (10):1805–1815, 2012. CODEN CSSCDB. ISSN 0361-0918.

Jimenez:2017:EET

[JLGD17] Sara Jimenez, Dejian Lai, K. Lance Gould, and Barry R. Davis. Evaluating the effects of treatment switching with randomization as an instrumental variable in a randomized controlled trial. Communications in Statistics: Simulation and Computation, 46(6):4966–4990, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Javed:2013:GTM

[JM13] Farrukh Javed and Panagiotis Mantalos. GARCH-type models and performance of information criteria. Communications in Statistics: Simulation and Computation, 42(8): 1917–1933, 2013. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2016:PMC

[JMCT16] Wenyu Jiang, Nathalie C. Moon, Bingshu E. Chen, and Dongsheng Tu. Permutation methods for comparing the accuracy of nested prediction models in survival analysis. Communications in Statistics: Simulation and Computation, 45(8):2691–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Jayalath:2017:ITH

[JNMR17] Kalanka P. Jayalath, Hon Keung Tony Ng, Ananda B. Manage, and Kent E. Riggs. Improved tests for homogeneity of variances. Communications in Statistics: Simulation and Computation, 46(9):7423–7446, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jovanovic:2017:EGE

[Jov17] Milan Jovanović. Estimation of P {X < Y } for geometric– exponential model based on complete and censored samples. Communications in Statistics: Simulation and Computation, 46(4):3050–3066, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jaggi:2018:NCN

[JPV^{+}18] Seema Jaggi, Dinesh Kumar Pateria, Cini Varghese, Eldho Varghese, and Arpan Bhowmik. A note on circular neighbor balanced designs. Communications in Statistics: Simulation and Computation, 47(10):2896–2905, 2018. CODEN CSSCDB. ISSN 0361-0918.

Jensen:2012:VRT

[JR12] Donald R. Jensen and Donald E. Ramirez. Variations on ridge traces in regression. Communications in Statistics: Simulation and Computation, 41(2):265–278, 2012. CODEN CSSCDB. ISSN 0361-0918.

Jankowski:2012:ISC

[JS12] Hanna Jankowski and Larissa Stanberry. Identifying skeleton curves in noisy data. Communications in Statistics: Simulation and Computation, 41(6):852–864, 2012. CODEN CSSCDB. ISSN 0361-0918.

Jayasekara:2014:SBS

[JS14] D. N. Jayasekara and M. R. Sooriyarachchi. A simulation based study for comparing tests associated with receiver operating characteristic (ROC) curves. Communications in Statistics: Simulation and Computation, 43(10):2444–2467, 2014. CODEN CSSCDB. ISSN 0361-0918.

Jayakumar:2019:DLW

[JS19a] K. Jayakumar and K. K. Sankaran. Discrete Linnik Weibull distribution. Communications in Statistics: Simulation and Computation, 48(10):3092–3117, 2019. CODEN CSSCDB. ISSN 0361-0918.

Jlassi:2019:VIA

[JS19b] Ines Jlassi and Jérôme Saracco. Variable importance assessment in sliced inverse regression for variable selection. Communications in Statistics: Simulation and Computation, 48 (1):169–199, 2019. CODEN CSSCDB. ISSN 0361-0918.

Jaki:2018:EBM

[JSKV18] Thomas Jaki, Ting-Li Su, Minjung Kim, and M. Lee Van Horn. An evaluation of the bootstrap for model validation in mixture models. Communications in Statistics: Simulation and Computation, 47(4):1028–1038, 2018. CODEN CSSCDB. ISSN 0361-0918.

Jammalamadaka:2019:SPE

[JT19] S. Rao Jammalamadaka and Emanuele Taufer. Semiparametric estimation of the autoregressive parameter in non-Gaussian Ornstein–Uhlenbeck processes. Communications in Statistics: Simulation and Computation, 48(9): 2791–2811, 2019. CODEN CSSCDB. ISSN 0361-0918.

Jin:2010:EMC

[JTY10] Hao Jin, Zheng Tian, and Yunfeng Yang. Estimation mean change-point in ARCH models with heavy-tailed innovations. Communications in Statistics: Simulation and Computation, 39(2):390–404, 2010. CODEN CSSCDB. ISSN 0361-0918.

Jihn:2019:PDM

[JU19] Chang-Ho Jihn and M. Mujiya Ulkhaq. Parallel discretization of the Markov chain approximation for the autoregressive moving average chart. Communications in Statistics: Simulation and Computation, 48(9):2660–2678, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Jhajj:2012:GDR

[JW12] H. S. Jhajj and Gurjeet Singh Walia. A generalized difference-cum-ratio type estimator for the population mean in double sampling. Communications in Statistics: Simulation and Computation, 41(1):58–64, 2012. CODEN CSS-CDB. ISSN 0361-0918.

John:2018:CIF

[JW18] Majnu John and Yihren Wu. Confidence intervals for finite difference solutions. Communications in Statistics: Simulation and Computation, 47(7):2102–2118, 2018. CODEN CSSCDB. ISSN 0361-0918.

Jia:2014:RRD

[JWW14] Xujie Jia, Liying Wang, and Chuanhua Wei. Reliability research of dependent failure systems using copula. Communications in Statistics: Simulation and Computation, 43(8): 1838–1851, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wang:2013:BAT

[jWzM13] Jian jun Wang and Yi zhong Ma. Bayesian analysis of two-level fractional factorial experiments with non-normal responses. Communications in Statistics: Simulation and Computation, 42(9):1970–1988, 2013. CODEN CSSCDB. ISSN 0361-0918.

Jeong:2010:EPE

[JY10] Jinook Jeong and Byung Ho Yoon. The effect of pseudoexogenous instrumental variables on Hausman test. Communications in Statistics: Simulation and Computation, 39 (2):315–321, 2010. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2014:PFA

[JZ14] Wenhua Jiang and Cun-Hui Zhang. Paths following algorithm for penalized logistic regression using SCAD and MCP. Communications in Statistics: Simulation and Computation, 43(5):1064–1077, 2014. CODEN CSSCDB. ISSN 0361-0918.

Jinnah:2017:ELI

[JZ17] Ali Jinnah and Yichuan Zhao. Empirical likelihood inference for the bivariate survival function under univariate censoring. Communications in Statistics: Simulation and Computation, 46(6):4348–4355, 2017. CODEN CSSCDB. ISSN 0361-0918.

Jin:2018:SRB

[JZ18] Hao Jin and Si Zhang. Spurious regression between long memory series due to mis-specified structural breaks. Communications in Statistics: Simulation and Computation, 47 (3):692–711, 2018. CODEN CSSCDB. ISSN 0361-0918.

Jiang:2019:UJC

[JZ19] Fen Jiang and Jin Zhang. Using jackknife to correct bias of MLE for the truncated Pareto distribution. Communications in Statistics: Simulation and Computation, 48(6): 1706–1713, 2019. CODEN CSSCDB. ISSN 0361-0918.

Jin:2017:ESR

[JZZZ17] Hao Jin, Danshi Zhang, Jinsuo Zhang, and Si Zhang. Evidence of spurious regression driven by heavy-tailed observations with structural changes. Communications in Statistics: Simulation and Computation, 46(2):1086–1101, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Karadag:2019:GMS

[KA19a] Özge Karadag and Serpil Aktas. A generalized, multi-stage adjusted, latent class linear mixed model for testing genetic association. Communications in Statistics: Simulation and Computation, 48(8):2301–2312, 2019. CODEN CSSCDB. ISSN 0361-0918.

Khosravi:2019:SSC

[KA19b] Peyman Khosravi and Amirhossein Amiri. Self-starting control charts for monitoring logistic regression profiles. Communications in Statistics: Simulation and Computation, 48 (6):1860–1871, 2019. CODEN CSSCDB. ISSN 0361-0918.

Khan:2019:MDS

[KAAJ19] Nasrullah Khan, Muhammad Aslam, Liaquat Ahmad, and Chi-Hyuck Jun. Multiple dependent state repetitive sampling plans with or without auxiliary variable. Communications in Statistics: Simulation and Computation, 48(4): 1055–1069, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kleyn:2017:PTC

[KABM17] J. Kleyn, M. Arashi, A. Bekker, and S. Millard. Preliminary testing of the Cobb–Douglas production function and related inferential issues. Communications in Statistics: Simulation and Computation, 46(1):469–488, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kang:2016:OFD

[Kan16] Moonsu Kang. Optimal false discovery rate control with kernel density estimation in a microarray experiment. Communications in Statistics: Simulation and Computation, 45 (3):771–780, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kamranrad:2017:PIM

[KAN17] Reza Kamranrad, Amirhossein Amiri, and Seyed Taghi Akhavan Niaki. Phase-II monitoring and diagnosing of multivariate categorical processes using generalized linear test-based control charts. Communications in Statistics: Simulation and Computation, 46(8):5951–5980, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Kao:2017:CCB

[Kao17] Shih-Chou Kao. X̄ and R control charts based on weighted variance with left-right tail-weighted ratio for skewed distributions. Communications in Statistics: Simulation and Computation, 46(4):2714–2732, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kao:2018:CCB

[Kao18] Shih-Chou Kao. A control chart based on weighted bootstrap with strata. Communications in Statistics: Simulation and Computation, 47(2):556–581, 2018. CODEN CSSCDB. ISSN 0361-0918.

Karabatsos:2018:MML

[Kar18] George Karabatsos. Marginal maximum likelihood estimation methods for the tuning parameters of ridge, power ridge, and generalized ridge regression. Communications in Statistics: Simulation and Computation, 47(6):1632–1651, 2018. CODEN CSSCDB. ISSN 0361-0918.

Khan:2019:CCP

[KAS^{+}19] Abid Khan, Rashid Ahmed, Farrukh Shehzad, M. H. Tahir, and Syed Shakir Ali Ghazali. Construction of circular partially balanced-repeated measurement designs using cyclic shifts. Communications in Statistics: Simulation and Computation, 48(2):506–515, 2019. CODEN CSSCDB. ISSN 0361-0918.

Klakattawi:2011:BNB

[KBAD11] Hadeel S. Klakattawi, Lamya A. Baharith, and Gannat R. Al-Dayian. Bayesian and non Bayesian estimations on the exponentiated modified Weibull distribution for progressive censored sample. Communications in Statistics: Simulation and Computation, 40(9):1291–1309, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Kang:2014:SCI

[KBJ14] Jongkyeong Kang, Sungwan Bang, and Myoungshic Jhun. Simultaneous confidence intervals for the ratios of marginal means of a multivariate Poisson distribution. Communications in Statistics: Simulation and Computation, 43(7): 1783–1796, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kong:2017:AMI

[KC17] Dejing Kong and Lirong Cui. The analysis of mixed intervalcensored and complete data. Communications in Statistics: Simulation and Computation, 46(1):145–163, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kao:2016:EMD

[KCC16] Lie-Jane Kao, Tai-Yuan Chen, and Kuang-Chao Chang. Efficient mixture design fitting quadratic surface with quantile responses using first-degree polynomial. Communications in Statistics: Simulation and Computation, 45(4):1365–1380, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kamranfar:2017:DOG

[KCM17] Hoda Kamranfar, Rahim Chinipardaz, and Behzad Mansouri. Detecting outliers in GARCH(p,q) models. Communications in Statistics: Simulation and Computation, 46(10): 7844–7854, 2017. CODEN CSSCDB. ISSN 0361-0918.

Klimenok:2012:BPQ

[KD12] Valentina Klimenok and Alexander Dudin. A BMAP/PH/N queue with negative customers and partial protection of service. Communications in Statistics: Simulation and Computation, 41(7):1062–1082, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kaciranlar:2018:PPN

[KD18] Selahattin Kaçiranlar and Issam Dawoud. On the performance of the Poisson and the negative binomial ridge predictors. Communications in Statistics: Simulation and Computation, 47(6):1751–1770, 2018. CODEN CSSCDB. ISSN 0361-0918.

Karahasan:2017:NEB

[KE17] Mehmet Karahasan and Nehir Ene. New estimators based on order statistics in some families of scale distributions. Communications in Statistics: Simulation and Computation, 46 (4):2880–2906, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kittaneh:2017:EET

[KEB17] Omar A. Kittaneh and Mohamed A. El-Beltagy. Efficiency estimation of Type-I censored sample from the Weibull distribution based on sup-entropy. Communications in Statistics: Simulation and Computation, 46(4):2678–2688, 2017. CODEN CSSCDB. ISSN 0361-0918.

Khribi:2016:PME

[KFM16] Lotfi Khribi, Marc Fredette, and Brenda MacGibbon. The Poisson maximum entropy model for homogeneous Poisson processes. Communications in Statistics: Simulation and Computation, 45(9):3435–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kan:2014:RSC

[KFO14] Cihangir Kan, Ceki Franko, and C. Murat Ozkut. Reliability and some characteristics of circular consecutive- k-within-m-out-of-n: F system. Communications in Statistics: Simulation and Computation, 43(5):1078–1093, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kibria:2014:MTP

[KG14] B. M. Golam Kibria and Florence George. Methods for testing population signal-to-noise ratio. Communications in Statistics: Simulation and Computation, 43(3):443–461, 2014. CODEN CSSCDB. ISSN 0361-0918.

Korkmaz:2017:NGT

[KG17] Mustafa Ç. Korkmaz and Ali I. Genç. A new generalized two-sided class of distributions with an emphasis on twosided generalized normal distribution. Communications in Statistics: Simulation and Computation, 46(2):1441–1460, 2017. CODEN CSSCDB. ISSN 0361-0918.

Khowaja:2011:EPM

[KGA11] Saman Khowaja, Shazia Ghufran, and M. J. Ahsan. Estimation of population means in multivariate stratified random sampling. Communications in Statistics: Simulation and Computation, 40(5):710–718, 2011. CODEN CSSCDB. ISSN 0361-0918.

Khowaja:2013:PCA

[KGA13] Saman Khowaja, Shazia Ghufran, and M. J. Ahsan. On the problem of compromise allocation in multi-response stratified sample surveys. Communications in Statistics: Simulation and Computation, 42(4):790–799, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Kim:2014:HBA

[KGK14] Dal Ho Kim, Gyuhyeong Goh, and Yongku Kim. Hierarchical Bayesian approach to a multi-site hidden Markov model. Communications in Statistics: Simulation and Computation, 43(6):1241–1252, 2014. CODEN CSSCDB. ISSN 0361-0918.

Koyuncu:2014:ETE

[KGS14] Nursel Koyuncu, Sat Gupta, and Rita Sousa. Exponentialtype estimators of the mean of a sensitive variable in the presence of nonsensitive auxiliary information. Communications in Statistics: Simulation and Computation, 43(7): 1583–1594, 2014. CODEN CSSCDB. ISSN 0361-0918.

Krit:2016:SLB

[KGXR16] Meryam Krit, Olivier Gaudoin, Min Xie, and Emmanuel Remy. Simplified likelihood based goodness-of-fit tests for the Weibull distribution. Communications in Statistics: Simulation and Computation, 45(3):920–951, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kim:2017:DDG

[KH17] Jong-Min Kim and S. Y. Hwang. Directional dependence via Gaussian copula beta regression model with asymmetric GARCH marginals. Communications in Statistics: Simulation and Computation, 46(10):7639–7653, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kuo:2018:CPM

[KH18] Tzu-Chun Kuo and Todd C. Headrick. A characterization of power method transformations through the method of percentiles. Communications in Statistics: Simulation and Computation, 47(8):2197–2215, 2018. CODEN CSSCDB. ISSN 0361-0918.

Khashei:2019:CSS

[KH19a] Mehdi Khashei and Zahra Hajirahimi. A comparative study of series ARIMA/MLP hybrid models for stock price forecasting. Communications in Statistics: Simulation and Computation, 48(9):2625–2640, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kim:2019:CDD

[KH19b] Jong-Min Kim and S. Y. Hwang. The copula directional dependence by stochastic volatility models. Communications in Statistics: Simulation and Computation, 48(4): 1153–1175, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kennes:2012:CRS

[KHH12] Lieven Nils Kennes, Ralf-Dieter Hilgers, and Nicole Heussen. Choice of the reference set in a randomization test based on linear ranks in the presence of missing values. Communications in Statistics: Simulation and Computation, 41(7): 1051–1061, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kim:2014:RAB

[Kim14] Yang-Jin Kim. Regression analysis of bivariate current status data using a multistate model. Communications in Statistics: Simulation and Computation, 43(3):462–475, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kalema:2016:CMT

[KIM16a] George Kalema, Samuel Iddi, and Geert Molenberghs. The combined model: A tool for simulating correlated counts with overdispersion. Communications in Statistics: Simulation and Computation, 45(7):2491–??, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Kim:2016:CSS

[Kim16b] Young-Ju Kim. A comparative study of semiparametric estimation in partially linear single-index models. Communications in Statistics: Simulation and Computation, 45(7): 2577–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kim:2017:CRM

[Kim17] Yang-Jin Kim. Cure rate model with bivariate interval censored data. Communications in Statistics: Simulation and Computation, 46(9):7116–7124, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kizilaslan:2019:BEP

[Kiz19] Fatih Kizilaslan. E-Bayesian estimation for the proportional hazard rate model based on record values. Communications in Statistics: Simulation and Computation, 48(2):350–371, 2019. CODEN CSSCDB. ISSN 0361-0918.

Khan:2013:CJG

[KJ13] N. Mamode Khan and V. Jowaheer. Comparing joint GQL estimation and GMM adaptive estimation in COM-Poisson longitudinal regression model. Communications in Statistics: Simulation and Computation, 42(4):755–770, 2013. CODEN CSSCDB. ISSN 0361-0918.

Kim:2017:NBA

[KJK17] Dal Ho Kim, Aejung Jo, and Yongku Kim. Nonparametric Bayesian analysis for multi-site hidden Markov model. Communications in Statistics: Simulation and Computation, 46 (6):4896–4907, 2017. CODEN CSSCDB. ISSN 0361-0918.

Krishna:2013:AMO

[KJR13] E. Krishna, K. K. Jose, and Miroslav M. Ristić. Applications of Marshall–Olkin Fréchet distribution. Communications in Statistics: Simulation and Computation, 42(1):76–89, 2013. CODEN CSSCDB. ISSN 0361-0918.

Kim:2010:OCC

[KJS10] Seoung Bum Kim, Weerawat Jitpitaklert, and Thuntee Sukchotrat. One-class classification-based control charts for monitoring autocorrelated multivariate processes. Communications in Statistics: Simulation and Computation, 39(3): 461–474, 2010. CODEN CSSCDB. ISSN 0361-0918.

Kayano:2010:SFP

[KK10a] Mitsunori Kayano and Sadanori Konishi. Sparse functional principal component analysis via regularized basis expansions and its application. Communications in Statistics: Simulation and Computation, 39(7):1318–1333, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Kossler:2010:ATT

[KK10b] W. Kössler and N. Kumar. An adaptive test for the twosample scale problem based on U-statistics. Communications in Statistics: Simulation and Computation, 39(9): 1785–1802, 2010. CODEN CSSCDB. ISSN 0361-0918.

Kallioras:2014:PEI

[KK14a] Athanasios G. Kallioras and Ioannis A. Koutrouvelis. Percentile estimation in inverse Gaussian distributions. Communications in Statistics: Simulation and Computation, 43 (2):269–284, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kaveh:2014:ETG

[KK14b] Marjan Kaveh and Majid Jafari Khaledi. Estimation in truncated GLG model for ordered categorical spatial data using the SAEM algorithm. Communications in Statistics: Simulation and Computation, 43(3):528–537, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Koyuncu:2016:CWS

[KK16] Nursel Koyuncu and Cem Kadilar. Calibration weighting in stratified random sampling. Communications in Statistics: Simulation and Computation, 45(7):2267–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kanika:2017:BEL

[KK17a] Kanika and Somesh Kumar. Bayes estimation of a Langevin mean direction. Communications in Statistics: Simulation and Computation, 46(4):2769–2783, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Kim:2017:BFC

[KK17b] Jaehee Kim and Minjung Kyung. Bayesian Fourier clustering of gene expression data. Communications in Statistics: Simulation and Computation, 46(8):6475–6494, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

PhD:2019:FEM

[KKGK19] Eda Karaismailoglu, Naime Meric Konar, Dincer Goksuluk, and Ahmet Ergun Karaagaoglu. Factors effecting the model performance measures area under the ROC curve, net reclassification improvement and integrated discrimination improvement. Communications in Statistics: Simulation and Computation, 48(9):2586–2598, 2019. CODEN CSSCDB. ISSN 0361-0918.

Khan:2017:TGE

[KKH17] Muhammad Shuaib Khan, Robert King, and Irene Lena Hudson. Transmuted generalized exponential distribution: A generalization of the exponential distribution with applications to survival data. Communications in Statistics: Simulation and Computation, 46(6):4377–4398, 2017. CODEN CSSCDB. ISSN 0361-0918.

Khaw:2017:MCV

[KKYW17] Khai Wah Khaw, Michael B. C. Khoo, Wai Chung Yeong, and Zhang Wu. Monitoring the coefficient of variation using a variable sample size and sampling interval control chart. Communications in Statistics: Simulation and Computation, 46(7):5772–5794, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kuang:2014:MDP

[KL14] Nenghui Kuang and Bingquan Liu. Moderate deviation for parameter estimation in the Rayleigh diffusion process. Communications in Statistics: Simulation and Computation, 43(7):1631–1643, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kim:2017:VBB

[KL17a] Jongphil Kim and Ji-Hyun Lee. The validation of a betabinomial model for overdispersed binomial data. Communications in Statistics: Simulation and Computation, 46(2): 807–814, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kong:2017:SSB

[KL17b] Taewoon Kong and Kichun Lee. Semi-supervised Bayesian adaptive multiresolution shrinkage for wavelet-based denoising. Communications in Statistics: Simulation and Computation, 46(6):4399–4418, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kang:2018:TLR

[KL18] Xiaojuan Kang and Tizheng Li. Testing a linear relationship in varying coefficient spatial autoregressive models. Communications in Statistics: Simulation and Computation, 47(1): 187–205, 2018. CODEN CSSCDB. ISSN 0361-0918.

Kim:2019:MPP

[KLK19] Yongku Kim, Woo Dong Lee, and Sang Gil Kang. A matching prior for the product of normal means based on the modified profile likelihood. Communications in Statistics: Simulation and Computation, 48(5):1312–1329, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kaczynski:2012:BNR

[KLLM12a] W. Kaczynski, L. Leemis, N. Loehr, and J. McQueston. Bivariate nonparametric random variate generation using a piecewise-linear cumulative distribution function. Communications in Statistics: Simulation and Computation, 41(4): 469–496, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kaczynski:2012:NRV

[KLLM12b] W. Kaczynski, L. Leemis, N. Loehr, and J. McQueston. Nonparametric random variate generation using a piecewiselinear cumulative distribution function. Communications in Statistics: Simulation and Computation, 41(4):449–468, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kachouie:2017:FCD

[KLS17] Nezamoddin N. Kachouie, Xihong Lin, and Armin Schwartzman. FDR control of detected regions by multiscale matched filtering. Communications in Statistics: Simulation and Computation, 46(1):127–144, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kawai:2012:IVT

[KM12] Reiichiro Kawai and Hiroki Masuda. Infinite variation tempered stable Ornstein–Uhlenbeck processes with discrete observations. Communications in Statistics: Simulation and Computation, 41(1):125–139, 2012. CODEN CSSCDB. ISSN 0361-0918.

Khoshkenar:2013:NTR

[KM13] Amin Khoshkenar and Hashem Mahlooji. A new test of randomness for Lehmer generators based on the Manhattan distance between pairs of consecutive random numbers. Communications in Statistics: Simulation and Computation, 42 (1):202–214, 2013. CODEN CSSCDB. ISSN 0361-0918.

Kayal:2019:SDG

[KM19a] Suchandan Kayal and Rajesh Moharana. A shift-dependent generalized cumulative entropy of order n. Communications in Statistics: Simulation and Computation, 48(6): 1768–1783, 2019. CODEN CSSCDB. ISSN 0361-0918.

Khadem:2019:ESD

[KM19b] Y. Khadem and M. Bameni Moghadam. Economic statistical design of X̄-control charts: Modified version of Rahim and Banerjee (1993) model. Communications in Statistics: Simulation and Computation, 48(3):684–703, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kharazmi:2015:GWE

[KMF15] Omid Kharazmi, Abbas Mahdavi, and Malek Fathizadeh. Generalized weighted exponential distribution. Communications in Statistics: Simulation and Computation, 44(6): 1557–1569, 2015. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2011:VSM

[KMS11] C. Koukouvinos, K. Mylona, and A. Skountzou. A variable selection method for analyzing supersaturated designs. Communications in Statistics: Simulation and Computation, 40(4):484–496, 2011. CODEN CSSCDB. ISSN 0361-0918.

Kibria:2015:SSS

[KMS15] B. M. Golam Kibria, Kristofer Månsson, and Ghazi Shukur. A simulation study of some biasing parameters for the ridge type estimation of Poisson regression. Communications in Statistics: Simulation and Computation, 44(4):943–957, 2015. CODEN CSSCDB. ISSN 0361-0918.

Koukouvinos:2013:GCM

[KMSG13] C. Koukouvinos, K. Mylona, A. Skountzou, and P. Goos. A general construction method for five-level second-order rotatable designs. Communications in Statistics: Simulation and Computation, 42(9):1961–1969, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Kelkar:2015:SAU

[KMSK15] V. A. Kelkar, R. V. Mohan, H. Shiferaw, and A. D. Kelkar. Statistical analysis of uncertainties in deterministic computational modeling — application to composite process resin infusion flow model. Communications in Statistics: Simulation and Computation, 44(9):2251–2263, 2015. CODEN CSSCDB. ISSN 0361-0918.

Karagrigoriou:2019:CCG

[KMV19] A. Karagrigoriou, A. Makrides, and I. Vonta. On a control chart for the Gini index with simulations. Communications in Statistics: Simulation and Computation, 48(4): 1121–1137, 2019. CODEN CSSCDB. ISSN 0361-0918.

Khedmati:2017:PMG

[KN17] Majid Khedmati and Seyed Taghi Akhavan Niaki. Phase-I monitoring of general linear profiles in multistage processes.

Communications in Statistics: Simulation and Computation, 46(6):4465–4489, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kazemzadeh:2015:CPE

[KNA15] Reza Baradaran Kazemzadeh, Rassoul Noorossana, and Mona Ayoubi. Change point estimation of multivariate linear profiles under linear drift. Communications in Statistics: Simulation and Computation, 44(6):1570–1599, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Knaeble:2019:ATC

[Kna19] Brian Knaeble. Adjustment with three continuous variables. Communications in Statistics: Simulation and Computation, 48(2):627–633, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kim:2016:EPB

[KNJ16] Seong W. Kim, Hon Keung Tony Ng, and Hakjin Jang. Estimation of parameters in a bivariate generalized exponential distribution based on Type-II censored samples. Communications in Statistics: Simulation and Computation, 45(10): 3776–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kim:2019:CSN

[KNLK19] Seong W. Kim, Hon Keung Tony Ng, Jung-Dong Lee, and Jinheum Kim. A comparative study on nonparametric estimation procedures for survival quantiles. Communications in Statistics: Simulation and Computation, 48(10):2968– 2984, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kardiyen:2016:CTG

[KO16] Filiz Kardiyen and Hülya Olmus. A comparison of two group classification approaches to fat-tailed and skewed data. Communications in Statistics: Simulation and Computation, 45(1):17–32, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kurtoglu:2019:RRE

[KÖ19] Fikriye Kurtoglu and M. Revan Özkale. Restricted ridge estimator in generalized linear models: Monte Carlo simulation studies on Poisson and binomial distributed responses. Communications in Statistics: Simulation and Computation, 48 (4):1191–1218, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kwon:2016:RBC

[KOC16] Amy M. Kwon, Ming Ouyang, and Andrew Cheng. Resampling-based classification using depth for functional curves. Communications in Statistics: Simulation and Computation, 45(9):3329–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kocak:2017:MAU

[Koc17] Mehmet Kocak. Meta-analysis of univariate P -values. Communications in Statistics: Simulation and Computation, 46 (2):1257–1265, 2017. CODEN CSSCDB. ISSN 0361-0918.

Komori:2015:SAA

[Kom15] Yoshio Komori. Suitable algorithm associated with a parameterization for the three-parameter log-normal distribution. Communications in Statistics: Simulation and Computation, 44(1):239–246, 2015. CODEN CSSCDB. ISSN 0361-0918.

Korobeynikov:2012:CPM

[Kor12] Anton Korobeynikov. Consistency of parametric MLE under mixed case interval censoring. Communications in Statistics: Simulation and Computation, 41(7):1083–1092, 2012. CO-DEN CSSCDB. ISSN 0361-0918.

Korkmaz:2017:GSS

[Kor17] Mustafa Ç. Korkmaz. A generalized skew slash distribution via gamma-normal distribution. Communications in Statistics: Simulation and Computation, 46(2):1647–1660, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kozak:2014:CRS

[Koz14] Marcin Kozak. Comparison of random search method and genetic algorithm for stratification. Communications in Statistics: Simulation and Computation, 43(2):249–253, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kulkarni:2010:CSA

[KP10] H. V. Kulkarni and V. V. Patil. A comparative study of analyses of 2^{n} factorial experiments with a Poisson distributed response variable. Communications in Statistics: Simulation and Computation, 39(8):1530–1547, 2010. CODEN CSS-CDB. ISSN 0361-0918.

Koukouvinos:2012:ASD

[KP12] C. Koukouvinos and C. Parpoula. Analyzing supersaturated designs by means of an information based criterion. Communications in Statistics: Simulation and Computation, 41 (1):44–57, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kim:2017:VED

[KP17] Gwansu Kim and Mingue Park. Variance estimation of DEE and regression estimator when a first-phase cluster sample is restratified. Communications in Statistics: Simulation and Computation, 46(4):3020–3033, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kpanzou:2015:IFQ

[Kpa15] Tchilabalo Abozou Kpanzou. On the influence function of the quintile share ratio. Communications in Statistics: Simulation and Computation, 44(10):2492–2499, 2015. CODEN CSSCDB. ISSN 0361-0918.

Karim:2017:ASL

[KPG^{+}17] Mohammad Ehsanul Karim, John Petkau, Paul Gustafson, Helen Tremlett, and The Beams Study Group. On the application of statistical learning approaches to construct inverse probability weights in marginal structural Cox models: Hedging against weight-model misspecification. Communications in Statistics: Simulation and Computation, 46(10): 7668–7697, 2017. CODEN CSSCDB. ISSN 0361-0918.

Khan:2016:CBN

[KPIY16] Shahjahan Khan, Budi Pratikno, Adriana I. N. Ibrahim, and Rossita M. Yunus. The correlated bivariate noncentral F distribution and its application. Communications in Statistics: Simulation and Computation, 45(10):3491–??, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Kollo:2017:TDS

[KPV17] Tõnu Kollo, Gaida Pettere, and Marju Valge. Tail dependence of skew t-copulas. Communications in Statistics: Simulation and Computation, 46(2):1024–1034, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kaur:2012:RRE

[KR12] Inderdeep Kaur and M. B. Rajarshi. Ridge regression for estimation of transition probabilities from aggregate data.

Communications in Statistics: Simulation and Computation, 41(4):524–530, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2017:SAG

[KR17] C. Satheesh Kumar and A. Riyaz. On some aspects of a generalized alternative zero-inflated logarithmic series distribution. Communications in Statistics: Simulation and Computation, 46(4):2689–2700, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kotb:2018:BIP

[KR18] Mohammed S. Kotb and Mohammad Z. Raqab. Bayesian inference and prediction of the Rayleigh distribution based on ordered ranked set sampling. Communications in Statistics: Simulation and Computation, 47(3):905–923, 2018. CODEN CSSCDB. ISSN 0361-0918.

Khan:2015:DOS

[KRAA15] M. G. M. Khan, Dinesh Rao, A. H. Ansari, and M. J. Ahsan. Determining optimum strata boundaries and sample sizes for skewed population with log-normal distribution. Communications in Statistics: Simulation and Computation, 44(5): 1364–1387, 2015. CODEN CSSCDB. ISSN 0361-0918.

Krishnamoorthy:2016:MNB

[Kri16] K. Krishnamoorthy. Modified normal-based approximation to the percentiles of linear combination of independent random variables with applications. Communications in Statistics: Simulation and Computation, 45(7):2428–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Kohansal:2015:PET

[KRK15] A. Kohansal, S. Rezakhah, and E. Khorram. Parameter estimation of Type-II hybrid censored weighted exponential distribution. Communications in Statistics: Simulation and Computation, 44(5):1273–1299, 2015. CODEN CSSCDB. ISSN 0361-0918.

Kim:2018:NBM

[KRK18] Dal Ho Kim, Hyunnam Ryu, and Yongku Kim. Nonparametric Bayesian modeling for monotonicity in catch ratio. Communications in Statistics: Simulation and Computation, 47(4):1056–1065, 2018. CODEN CSSCDB. ISSN 0361-

Kawasaki:2015:TST

[KS15] Tamae Kawasaki and Takashi Seo. A two sample test for mean vectors with unequal covariance matrices. Communications in Statistics: Simulation and Computation, 44(7): 1850–1866, 2015. CODEN CSSCDB. ISSN 0361-0918.

Kumar:2016:SAC

[KS16] Deepak Kumar and S. B. Singh. Stochastic analysis of complex repairable system with deliberate failure emphasizing reboot delay. Communications in Statistics: Simulation and Computation, 45(2):583–602, 2016. CODEN CSSCDB. ISSN 0361-0918.

Khan:2015:GSS

[KSG15] Zaheen Khan, Javid Shabbir, and Sat Gupta. Generalized systematic sampling. Communications in Statistics: Simulation and Computation, 44(9):2240–2250, 2015. CODEN CSSCDB. ISSN 0361-0918.

Khan:2018:CVS

[KSG18] Zaheen Khan, Javid Shabbir, and S. N. Gupta. Circular versions of systematic sampling in the presence of linear trend. Communications in Statistics: Simulation and Computation, 47(7):1899–1921, 2018. CODEN CSSCDB. ISSN 0361-0918.

Kundu:2013:BAD

[KSGM13] D. Kundu, D. Samanta, A. Ganguly, and S. Mitra. Bayesian analysis of different hybrid and progressive life tests. Communications in Statistics: Simulation and Computation, 42 (9):2160–2173, 2013. CODEN CSSCDB. ISSN 0361-0918.

Keown-Stoneman:2019:WMS

[KSHD19] Charles D. G. Keown-Stoneman, Julie Horrocks, and Gerarda A. Darlington. Weibull multi-state models with misclassification. Communications in Statistics: Simulation and Computation, 48(1):39–57, 2019. CODEN CSSCDB. ISSN 0361-0918.

Kim:2012:ASD

[KSJ^{+}12] Jinseog Kim, Insuk Sohn, Sin-Ho Jung, Sujong Kim, and Changyi Park. Analysis of survival data with group lasso.

Communications in Statistics: Simulation and Computation, 41(9):1593–1605, 2012. CODEN CSSCDB. ISSN 0361-0918.

Karcher:2013:SID

[KSS13] Wolfgang Karcher, Hans-Peter Scheffler, and Evgeny Spodarev. Simulation of infinitely divisible random fields. Communications in Statistics: Simulation and Computation, 42 (1):215–246, 2013. CODEN CSSCDB. ISSN 0361-0918.

Kaushik:2017:BIP

[KSS17] Arun Kaushik, Umesh Singh, and Sanjay Kumar Singh. Bayesian inference for the parameters of Weibull distribution under progressive Type-I interval censored data with beta-binomial removals. Communications in Statistics: Simulation and Computation, 46(4):3140–3158, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kara:2017:SIS

[KTA17] Mahmut Kara, Özlem Türksen, and Halil Aydogdu. Statistical inference for α-series process with the inverse Gaussian distribution. Communications in Statistics: Simulation and Computation, 46(6):4938–4950, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kayal:2017:IBX

[KTRA17] T. Kayal, Yogesh Mani Tripathi, M. K. Rastogi, and A. Asgharzadeh. Inference for Burr XII distribution under Type I progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 46(9):7447–7465, 2017. CODEN CSSCDB. ISSN 0361-0918.

Khokhar:2014:BSF

[KU14] Ankit Khokhar and S. K. Upadhyay. A Bayes study on the feasibility of shape and threshold parameters based on the assumption of exponential and Weibull models. Communications in Statistics: Simulation and Computation, 43(4): 851–870, 2014. CODEN CSSCDB. ISSN 0361-0918.

Khalil:2019:GEP

[KuAH19] Sadia Khalil, Muhammad Noor ul Amin, and Muhammad Hanif. Generalized estimator of population mean by using conventional and non-conventional measures in the presence of measurement errors. Communications in Statistics: Simulation and Computation, 48(2):516–529, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Kurita:2013:EIM

[Kur13] Takamitsu Kurita. Exploring the impact of multivariate GARCH innovations on hypothesis testing for cointegrating vectors. Communications in Statistics: Simulation and Computation, 42(8):1785–1800, 2013. CODEN CSSCDB. ISSN 0361-0918.

Kuruwita:2016:NIE

[Kur16] C. N. Kuruwita. Non-iterative estimation and variable selection in the single-index quantile regression model. Communications in Statistics: Simulation and Computation, 45 (10):3615–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Krzysko:2014:CKD

[KWW14] Mirosław Krzyśko, Łukasz Waszak, and Waldemar Wołyński. Comparison of kernel discriminant coordinates algorithms. Communications in Statistics: Simulation and Computation, 43(9):2138–2148, 2014. CODEN CSSCDB. ISSN 0361-0918.

Kuang:2017:ABW

[KX17] Nenghui Kuang and Huantian Xie. Asymptotic behavior of weighted cubic variation of sub-fractional Brownian motion. Communications in Statistics: Simulation and Computation, 46(1):215–229, 2017. CODEN CSSCDB. ISSN 0361-0918.

Kantar:2015:REP

[KY15] Yeliz Mert Kantar and Vural Yildirim. Robust estimation for parameters of the extended Burr Type III distribution. Communications in Statistics: Simulation and Computation, 44 (7):1901–1930, 2015. CODEN CSSCDB. ISSN 0361-0918.

Khawsithiwong:2011:AKF

[KYP11] P. Khawsithiwong, N. Yatawara, and V. Pongsapukdee. Adaptive Kalman filtering with multivariate generalized Laplace system noise. Communications in Statistics: Simulation and Computation, 40(9):1278–1290, 2011. CODEN CSSCDB. ISSN 0361-0918.

Kyung:2016:CBM

[Kyu16] Minjung Kyung. A computational Bayesian method for generalized semiparametric regression models. Communications in Statistics: Simulation and Computation, 45(3): 1104–1128, 2016. CODEN CSSCDB. ISSN 0361-0918.

LopezQuintero:2017:SAV

[Lan17] Freddy Omar López Quintero. Sensitivity analysis for variance parameters in Bayesian simplex mixed models for proportional data. Communications in Statistics: Simulation and Computation, 46(7):5212–5228, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Loehle:2013:RVM

[LA13] Craig Loehle and Nasser Arghami. Reduced-variance methods for detectability correction of population abundance. Communications in Statistics: Simulation and Computation, 42(6):1343–1351, 2013. CODEN CSSCDB. ISSN 0361-0918.

Lak:2017:EBU

[LAG17] Fazlolah Lak, Mahmood Afshari, and Behzad Gholizadeh. An efficiency Bayesian unit root test in unobserved-ARCH models. Communications in Statistics: Simulation and Computation, 46(6):4841–4850, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lagos-Alvarez:2011:BCT

[LÁJGAF11] B. Lagos-Álvarez, M. D. Jiménez-Gamero, and V. Alba-Fernández. Bias correction in the Type I generalized logistic distribution. Communications in Statistics: Simulation and Computation, 40(4):511–531, 2011. CODEN CSSCDB. ISSN 0361-0918.

Lak:2011:BUR

[Lak11] Fazlollah Lak. Bayesian unit root testing in unobserved-ARCH models. Communications in Statistics: Simulation and Computation, 40(2):208–215, 2011. CODEN CSSCDB. ISSN 0361-0918.

LaMotte:2014:IPM

[LaM14] Lynn R. LaMotte. On inverse prediction in mixed linear models. Communications in Statistics: Simulation and Computation, 43(9):2106–2116, 2014. CODEN CSSCDB. ISSN 0361-0918.

Landes:2010:HBC

[Lan10] Reid D. Landes. Hierarchical Bayesian calibration of untested devices. Communications in Statistics: Simulation and Computation, 39(7):1351–1364, 2010. CODEN CSS-CDB. ISSN 0361-0918.

Lavery:2019:NPM

[LASX19] Matthew Ryan Lavery, Parul Acharya, Stephen A. Sivo, and Lihua Xu. Number of predictors and multicollinearity: What are their effects on error and bias in regression? Communications in Statistics: Simulation and Computation, 48(1): 27–38, 2019. CODEN CSSCDB. ISSN 0361-0918.

Loose:2017:BBC

[LBP17] Laís H. Loose, Fábio M. Bayer, and Tarciana L. Pereira. Bootstrap Bartlett correction in inflated beta regression. Communications in Statistics: Simulation and Computation, 46(4):2865–2879, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lin:2011:PLT

[LBW11] Chien-Tai Lin, N. Balakrishnan, and Sam J. S. Wu. Planning life tests based on progressively Type-I grouped censored data from the Weibull distribution. Communications in Statistics: Simulation and Computation, 40(4):574–595, 2011. CODEN CSSCDB. ISSN 0361-0918.

L:2012:MDB

[LC12] Liliana Garrido L. and Edilberto Cepeda. Mixture of distributions in the biparametric exponential family: A Bayesian approach. Communications in Statistics: Simulation and Computation, 41(3):355–375, 2012. CODEN CSSCDB. ISSN 0361-0918.

Liu:2016:DAD

[LC16] Fangyu Liu and Lirong Cui. A design of attributes double sampling plans for three-class products. Communications in Statistics: Simulation and Computation, 45(3):1054–1071, 2016. CODEN CSSCDB. ISSN 0361-0918.

Lu:2018:CTM

[LCD18] Yan Lu, Zhongxue Chen, and Danielle Durán. Comparison of two methods in estimating standard error of the method of simulated moments estimators for generalized linear mixed models. Communications in Statistics: Simulation and Computation, 47(10):2886–2895, 2018. CODEN CSSCDB. ISSN 0361-0918.

Lima:2019:PML

[LCN19] Verônica M. C. Lima and Francisco Cribari-Neto. Penalized maximum likelihood estimation in the modified extended Weibull distribution. Communications in Statistics: Simulation and Computation, 48(2):334–349, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Lee:2017:PAS

[LDK17] Shun Fu Lee, Allan Donner, and Neil Klar. Power of the adjusted Q statistic to evaluate heterogeneity in meta-analyses of cluster randomized trials. Communications in Statistics: Simulation and Computation, 46(9):7062–7073, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Laud:2010:SST

[LDS10] Purushottam W. Laud, Paul Damien, and Thomas S. Shively. Sampling some truncated distributions via rejection algorithms. Communications in Statistics: Simulation and Computation, 39(6):1111–1121, 2010. CODEN CSSCDB. ISSN 0361-0918.

Li:2014:MCP

[LDW14] Zhonghua Li, Yi Dai, and Zhaojun Wang. Multivariate change point control chart based on data depth for Phase I analysis. Communications in Statistics: Simulation and Computation, 43(6):1490–1507, 2014. CODEN CSSCDB. ISSN 0361-0918.

Lee:2010:EDC

[Lee10a] M. H. Lee. Economic design of cumulative sum control charts for monitoring a process with correlated samples. Communications in Statistics: Simulation and Computation, 39(10):1909–1922, 2010. CODEN CSSCDB. ISSN 0361-0918.

Lee:2010:MEC

[Lee10b] M. H. Lee. Multivariate EWMA control chart with adaptive sample sizes. Communications in Statistics: Simulation and Computation, 39(8):1548–1561, 2010. CODEN CSSCDB. ISSN 0361-0918.

Lee:2012:DMS

[Lee12] Ming Ha Lee. The design of the multivariate synthetic exponentially weighted moving average control chart. Communications in Statistics: Simulation and Computation, 41(10): 1785–1793, 2012. CODEN CSSCDB. ISSN 0361-0918.

Lu:2011:ETV

[LG11] Zhiping Lu and Dominique Guegan. Estimation of timevarying long memory parameter using wavelet method. Communications in Statistics: Simulation and Computation, 40(4):596–613, 2011. CODEN CSSCDB. ISSN 0361-0918.

Lukyanenko:2012:ABP

[LGM12] Andrey Lukyanenko, Andrei Gurtov, and Evsey Morozov. An adaptive backoff protocol with Markovian contention window control. Communications in Statistics: Simulation and Computation, 41(7):1093–1106, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Liu:2013:THI

[LH13] Xuhua Liu and Daojiang He. Testing homogeneity of inverse Gaussian scale parameters based on generalized likelihood ratio. Communications in Statistics: Simulation and Computation, 42(2):382–392, 2013. CODEN CSSCDB. ISSN 0361-0918.

Lin:2014:SRC

[LH14] Yi-Kuei Lin and Cheng-Fu Huang. Spare reliability for capacitated computer networks under tolerable error rate and latency considerations. Communications in Statistics: Simulation and Computation, 43(8):1879–1899, 2014. CODEN CSSCDB. ISSN 0361-0918.

Laurini:2017:GGE

[LH17] Márcio Poletti Laurini and Luiz Koodi Hotta. GMC/ GEL estimation of stochastic volatility models. Communications in Statistics: Simulation and Computation, 46(9): 6828–6844, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lin:2008:EBV

[LHB08] Chien-Tai Lin, Yen-Lung Huang, and N. Balakrishnan. Exact Bayesian variable sampling plans for the exponential distribution based on Type-I and Type-II hybrid censored samples. Communications in Statistics: Simulation and Computation, 37(6):1101–1116, 2008. CODEN CSSCDB. ISSN 0361-0918. See corrections [LHB10].

Lin:2010:CEB

[LHB10] Chien-Tai Lin, Yen-Lung Huang, and N. Balakrishnan. Corrections on “Exact Bayesian Variable Sampling Plans for the Exponential Distribution Based on Type-I and Type-II Hybrid Censored Samples”. Communications in Statistics: Simulation and Computation, 39(7):1499–1505, 2010. CO-DEN CSSCDB. ISSN 0361-0918. See [LHB08].

Li:2010:SNB

[Li10a] Chin-Shang Li. Semiparametric negative binomial regression models. Communications in Statistics: Simulation and Computation, 39(3):475–486, 2010. CODEN CSSCDB. ISSN 0361-0918.

Li:2010:BIW

[Li10b] Longhai Li. Are Bayesian inferences weak for Wasserman’s example? Communications in Statistics: Simulation and Computation, 39(3):655–667, 2010. CODEN CSSCDB. ISSN 0361-0918.

Li:2011:TLF

[Li11] Chin-Shang Li. Testing the lack-of-fit of zero-inflated Poisson regression models. Communications in Statistics: Simulation and Computation, 40(4):497–510, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Li:2012:ABR

[Li12] Jialiang Li. Applications of the bootstrap in ROC analysis. Communications in Statistics: Simulation and Computation, 41(6):865–877, 2012. CODEN CSSCDB. ISSN 0361-0918.

Li:2013:PLM

[Li13] Chin-Shang Li. A partially linear model with its applications. Communications in Statistics: Simulation and Computation, 42(7):1673–1680, 2013. CODEN CSSCDB. ISSN 0361-0918.

Li:2014:CPT

[Li14] Liang Li. A class of permutation tests for the equality of two marginal survival functions using paired censored data. Communications in Statistics: Simulation and Computation, 43(10):2498–2507, 2014. CODEN CSSCDB. ISSN 0361-0918.

Lin:2015:MSI

[Lin15] Ting Hsiang Lin. Model selection information criteria in latent class models with missing data and contingency question. Communications in Statistics: Simulation and Computation, 44(2):319–331, 2015. CODEN CSSCDB. ISSN 0361-0918.

Lin:2017:BSA

[Lin17] Xiaoyan Lin. A Bayesian semiparametric accelerated failure time model for arbitrarily censored data with covariates subject to measurement error. Communications in Statistics: Simulation and Computation, 46(1):747–756, 2017. CODEN CSSCDB. ISSN 0361-0918.

Liu:2013:UWR

[Liu13] Xiaofeng Steven Liu. A unified way to represent and compute the cumulative distribution function for noncentral t, F, and χ^{2}. Communications in Statistics: Simulation and Computation, 42(6):1433–1436, 2013. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:APD

[Liu17a] Xiaohui Liu. Approximating the projection depth median of dimensions p ≥ 3. Communications in Statistics: Simulation and Computation, 46(5):3756–3768, 2017. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:ELR

[Liu17b] Xiaohui Liu. Empirical likelihood for the response mean of generalized linear models with missing at random responses.

Communications in Statistics: Simulation and Computation, 46(1):164–173, 2017. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:BMA

[Liu17c] Zhe Liu. Bayesian model-averaged regularization for Gaussian graphical models. Communications in Statistics: Simulation and Computation, 46(4):3213–3223, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2012:PMS

[LJ12] Sang-Ho Lee and Chi-Hyuck Jun. A process monitoring scheme controlling false discovery rate. Communications in Statistics: Simulation and Computation, 41(10):1912–1920, 2012. CODEN CSSCDB. ISSN 0361-0918.

Law:2017:RPL

[LJ17] Martin Law and Dan Jackson. Residual plots for linear regression models with censored outcome data: a refined method for visualizing residual uncertainty. Communications in Statistics: Simulation and Computation, 46(4): 3159–3171, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2014:SMS

[LK14] Ming Ha Lee and Michael B. C. Khoo. The synthetic mean square error control chart. Communications in Statistics: Simulation and Computation, 43(6):1523–1542, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Lee:2015:MSC

[LK15a] Ming Ha Lee and Michael B. C. Khoo. Multivariate synthetic —s— control chart with variable sampling interval. Communications in Statistics: Simulation and Computation, 44 (4):924–942, 2015. CODEN CSSCDB. ISSN 0361-0918.

Lee:2015:VSI

[LK15b] Ming Ha Lee and Michael B. C. Khoo. Variable sampling interval cumulative count of conforming chart with runs rules. Communications in Statistics: Simulation and Computation, 44(9):2410–2430, 2015. CODEN CSSCDB. ISSN 0361-0918.

Lee:2017:CSC

[LK17a] Ming Ha Lee and Michael B. C. Khoo. Combined synthetic and |S| chart for monitoring process dispersion. Communications in Statistics: Simulation and Computation, 46(7): 5698–5711, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2017:DSM

[LK17b] Ming Ha Lee and Michael B. C. Khoo. Double sampling max chart. Communications in Statistics: Simulation and Computation, 46(10):7855–7878, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2018:DSC

[LK18a] Ming Ha Lee and Michael B. C. Khoo. Double sampling |S| control chart with variable sample size and variable sampling interval. Communications in Statistics: Simulation and Computation, 47(2):615–628, 2018. CODEN CSSCDB. ISSN 0361-0918.

Lee:2018:ESD

[LK18b] Ming Ha Lee and Michael B. C. Khoo. Economic-statistical design of control chart with runs rules for correlation within sample. Communications in Statistics: Simulation and Computation, 47(10):2849–2864, 2018. CODEN CSSCDB. ISSN 0361-0918.

Lee:2019:EES

[LK19] Ming Ha Lee and Michael B. C. Khoo. The economic and economic statistical designs of synthetic double sampling X̄ chart. Communications in Statistics: Simulation and Computation, 48(8):2313–2332, 2019. CODEN CSSCDB. ISSN 0361-0918.

Liu:2011:BCA

[LKHH11] Tzu-Hsin Liu, Jau-Chuan Ke, Yu-Lun Hsu, and Ying-Lin Hsu. Bootstrapping computation of availability for a repairable system with standby subject to imperfect switching. Communications in Statistics: Simulation and Computation, 40(4):469–483, 2011. CODEN CSSCDB. ISSN 0361-0918.

Low:2012:RRR

[LKTW12] Chun Kit Low, Michael B. C. Khoo, Wei Lin Teoh, and Zhang Wu. The revised m-of-k runs rule based on median run length. Communications in Statistics: Simulation and Computation, 41(8):1463–1477, 2012. CODEN CSSCDB. ISSN 0361-0918.

Kuo:2013:FLR

[lKW13] Kun lin Kuo and Yuchung J. Wang. A fresh look at the running time analysis for the Gibbs sampler. Communications in Statistics: Simulation and Computation, 42(8): 1815–1823, 2013. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:SIA

[LKYH17] Tzu-Hsin Liu, Jau-Chuan Ke, Meng-Feng Yen, and Ying-Lin Hsu. Statistical inference for the availability of repairable system with general repair, reboot delay, and switching failure. Communications in Statistics: Simulation and Computation, 46(1):579–592, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lui:2014:NTH

[LL14] Kung-Jong Lui and Chii-Dean Lin. Notes on test homogeneity of the odds ratio in matched pairs under stratified sampling: A Monte Carlo evaluation. Communications in Statistics: Simulation and Computation, 43(10):2403–2414, 2014. CODEN CSSCDB. ISSN 0361-0918.

Lin:2017:TSA

[LL17a] Chien-Tai Lin and Chung-Tao Lu. A two-stage algorithm for computation of the exact distribution of sums of spacings. Communications in Statistics: Simulation and Computation, 46(10):8205–8217, 2017. CODEN CSSCDB. ISSN 0361-0918.

Luo:2017:DHP

[LL17b] Ronghua Luo and Wei Lan. Detecting homogeneous predictors in high-dimensional panel model with an MCMC algorithm. Communications in Statistics: Simulation and Computation, 46(9):7376–7392, 2017. CODEN CSSCDB. ISSN 0361-0918.

Li:2018:BAH

[LL18] Xuerui Li and Luqin Liu. Bayesian analysis for heteroscedastic normal-Pareto mixture model with application. Communications in Statistics: Simulation and Computation, 47(9): 2775–2791, 2018. CODEN CSSCDB. ISSN 0361-0918.

Lee:2019:CTT

[LL19] Youngmi Lee and Sangyeol Lee. On causality test for time series of counts based on Poisson INGARCH models with application to crime and temperature data. Communications in Statistics: Simulation and Computation, 48(6): 1901–1911, 2019. CODEN CSSCDB. ISSN 0361-0918.

Lin:2019:CDL

[LLB19] Chien-Tai Lin, Bo-Yi Lee, and N. Balakrishnan. On the computation of distributions of linear combinations of Laplace order statistics and their applications. Communications in Statistics: Simulation and Computation, 48(10):3031–3055, 2019. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:NGC

[LLH17] Xuhua Liu, Na Li, and Yuqin Hu. A new generalized confidence interval for the among-group variance in the heteroscedastic one-way random effects model. Communications in Statistics: Simulation and Computation, 46(3): 2299–2310, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lim:2012:RMA

[LLJ^{+}12] Johan Lim, Woojoo Lee, Sin-Ho Jung, Kyeong Eun Lee, and Sung-Cheol Yun. A regression model for the AUC of clustered ordinal test results and working independent optimal weights. Communications in Statistics: Simulation and Computation, 41(8):1397–1410, 2012. CODEN CSSCDB. ISSN 0361-0918.

Lloyd:2017:CEO

[Llo17] Chris J. Lloyd. Computing exact one-sided confidence limits for treatment effect in clinical trials. Communications in Statistics: Simulation and Computation, 46(1):795–805, 2017. CODEN CSSCDB. ISSN 0361-0918.

Liu:2016:NFS

[LLS16] Miao Liu, Xiaoling Lu, and Jie Song. A new feature selection method for text categorization of customer reviews. Communications in Statistics: Simulation and Computation, 45 (4):1397–1409, 2016. CODEN CSSCDB. ISSN 0361-0918.

Liao:2017:RCI

[LLWY17] Mou-Yuan Liao, Chien-Hua Lin, Chien-Wei Wu, and Chien-Hui Yang. Reliable confidence intervals for assessing normal process incapability. Communications in Statistics: Simulation and Computation, 46(1):446–457, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Liu:2016:ELA

[LLX16] Wanrong Liu, Xuewen Lu, and Changchun Xie. Empirical likelihood for the additive hazards model with current status data. Communications in Statistics: Simulation and Computation, 45(8):2720–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Luo:2016:RSR

[LLZ16] Ji Luo, Haifen Li, and Jiajia Zhang. Robust smoothed rank estimation methods for accelerated failure time model allowing clusters. Communications in Statistics: Simulation and Computation, 45(6):1865–1884, 2016. CODEN CSSCDB. ISSN 0361-0918.

Liu:2018:ADT

[LLZ18] Yang Liu, Ji Li, and Xu Zhang. Analysis of dependently truncated data in Cox framework. Communications in Statistics: Simulation and Computation, 47(6):1677–1695, 2018. CODEN CSSCDB. ISSN 0361-0918.

Li:2010:TLR

[LM10] Tizheng Li and Changlin Mei. Testing for linear regression relationships in randomly right-censored varying-coefficient models. Communications in Statistics: Simulation and Computation, 39(6):1122–1146, 2010. CODEN CSSCDB. ISSN 0361-0918.

Latif:2016:CCT

[LM16a] Sumaia A. Latif and Pedro A. Morettin. Curve of correlation for time series. Communications in Statistics: Simulation and Computation, 45(8):2792–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Li:2016:SIP

[LM16b] Tizheng Li and Changlin Mei. Statistical inference on the parametric component in partially linear spatial autoregressive models. Communications in Statistics: Simulation and Computation, 45(6):1991–2006, 2016. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:MMS

[LM17] Haigang Liu and Reza Modarres. Modeling a minimal spanning tree. Communications in Statistics: Simulation and Computation, 46(7):5246–5256, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2011:HDN

[LML11] Hyejin Lee, Ming Meng, and Junsoo Lee. How do nonlinear unit root tests perform with non normal errors? Communications in Statistics: Simulation and Computation, 40(8): 1182–1191, 2011. CODEN CSSCDB. ISSN 0361-0918.

Locking:2013:PSR

[LMS13] Håkan Locking, Kristofer Månsson, and Ghazi Shukur. Performance of some ridge parameters for probit regression: With application to Swedish job search data. Communications in Statistics: Simulation and Computation, 42(3): 698–710, 2013. CODEN CSSCDB. ISSN 0361-0918.

Livingston:2017:BIS

[LN17] Glen Livingston, Jr. and Darfiana Nur. Bayesian inference for smooth transition autoregressive (STAR) model: A prior sensitivity analysis. Communications in Statistics: Simulation and Computation, 46(7):5440–5461, 2017. CODEN CSSCDB. ISSN 0361-0918.

Li:2011:SSA

[LNL11] Yong Li, Zhong-Xin Ni, and Jin-Guan Lin. A stochastic simulation approach to model selection for stochastic volatility models. Communications in Statistics: Simulation and Computation, 40(7):1043–1056, 2011. CODEN CSSCDB. ISSN 0361-0918.

Liang:2017:SRN

[LNT17] Jiajuan Liang, Kai Wang Ng, and Guoliang Tian. A stochastic representation for the l_{p}-norm symmetric distribution and its applications. Communications in Statistics: Simulation and Computation, 46(8):6011–6018, 2017. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:DCA

[LNUT17] Yan Liu, Hideaki Nagahata, Hirotaka Uchiyama, and Masanobu Taniguchi. Discriminant and cluster analysis of possibly high-dimensional time series data by a class of disparities. Communications in Statistics: Simulation and Computation, 46(10):8014–8027, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2016:DIO

[LO16] Hyejin Lee and Dong-Yop Oh. Dealing with the initial observation in the LM unit root test. Communications in Statistics: Simulation and Computation, 45(10):3660–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Lin:2016:PIF

[LOL16] Yu-Jau Lin, H. M. Okasha, and Y. L. Lio. Prediction intervals for future order statistics from a generalized modified Weibull. Communications in Statistics: Simulation and Computation, 45(10):3508–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Longford:2014:SSC

[Lon14] Nicholas T. Longford. Sample size calculation for comparing two normal random samples using equilibrium priors. Communications in Statistics: Simulation and Computation, 43 (10):2149–2161, 2014. CODEN CSSCDB. ISSN 0361-0918.

Li:2010:MDT

[LP10] Yan Li and Xiaolong Pu. A method for designing threehypothesis test problems and sequential schemes. Communications in Statistics: Simulation and Computation, 39(9): 1690–1708, 2010. CODEN CSSCDB. ISSN 0361-0918.

Lecoutre:2016:NRC

[LP16] Bruno Lecoutre and Jacques Poitevineau. New results for computing Blaker’s exact confidence interval for one parameter discrete distributions. Communications in Statistics: Simulation and Computation, 45(3):1041–1053, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Luo:2017:GTF

[LPC17] Lingling Luo, Sattayatham Pairote, and Ratthachat Chatpatanasiri. GARCH-type forecasting models for volatility of stock market and MCS test. Communications in Statistics: Simulation and Computation, 46(7):5303–5312, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lopes:2013:MDP

[LPZ13] Sílvia R. C. Lopes, Guilherme Pumi, and Karine Zaniol. Mallows distance in VARFIMA(0, d, 0) processes. Communications in Statistics: Simulation and Computation, 42(1): 24–51, 2013. CODEN CSSCDB. ISSN 0361-0918.

Liu:2011:TCC

[LqQlX11] Miao Liu, Xiu quan Qiao, and Wang li Xu. Three categories customer churn prediction based on the adjusted Real Adaboost. Communications in Statistics: Simulation and Computation, 40(10):1548–1562, 2011. CODEN CSSCDB. ISSN 0361-0918.

Li:2016:GVS

[LQW^{+}16] Yang Li, Yichen Qin, Limin Wang, Jiaxu Chen, and Shuangge Ma. Grouped variable selection using area under the ROC with imbalanced data. Communications in Statistics: Simulation and Computation, 45(4):1268–1280, 2016. CODEN CSSCDB. ISSN 0361-0918.

Livacic-Rojas:2010:ATE

[LRVF10] Pablo Livacic-Rojas, Guillermo Vallejo, and Paula Fernández. Analysis of Type I error rates of univariate and multivariate procedures in repeated measures designs. Communications in Statistics: Simulation and Computation, 39(3):624–640, 2010. CODEN CSSCDB. ISSN 0361-0918.

Li:2010:TUR

[LS10] Yushu Li and Ghazi Shukur. Testing for unit root against LSTAR model: Wavelet improvement under GARCH distortion. Communications in Statistics: Simulation and Computation, 39(2):277–286, 2010. CODEN CSSCDB. ISSN 0361-0918.

Loganathan:2014:DSS

[LS14] A. Loganathan and K. Shalini. Determination of single sampling plans by attributes under the conditions of zeroinflated Poisson distribution. Communications in Statistics: Simulation and Computation, 43(3):538–548, 2014. CODEN CSSCDB. ISSN 0361-0918.

Lamichhane:2016:EMU

[LS16] Rajan Lamichhane and Sarjinder Singh. Estimation of mode using two-phase sampling. Communications in Statistics: Simulation and Computation, 45(7):2586–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Lamichhane:2017:IEP

[LSD17] R. Lamichhane, S. Singh, and N. Diawara. Improved estimation of population mean using known median of auxiliary variable. Communications in Statistics: Simulation and Computation, 46(4):2821–2828, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lu:2011:NTG

[LSLW11] Xiaoling Lu, Jie Song, Miao Liu, and Xizhi Wu. A new two-group-forecast-and-selection method for direct marketing. Communications in Statistics: Simulation and Computation, 40(10):1627–1636, 2011. CODEN CSSCDB. ISSN 0361-0918.

Lee:2013:SMS

[LSS13] Cheon-Sig Lee, Stephen A. Sedory, and Sarjinder Singh. Simulated minimum sample size requirements in various randomized response models. Communications in Statistics: Simulation and Computation, 42(4):771–789, 2013. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:EMT

[LSTM17] Xiaobin Liu, Guogen Shan, Lili Tian, and Chang-Xing Ma. Exact methods for testing homogeneity of proportions for multiple groups of paired binary data. Communications in Statistics: Simulation and Computation, 46(8):6074–6082, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lien:2016:CVA

[LSY16] Donald Lien, Christopher Stroud, and Keying Ye. Comparing VaR approximation methods that use the first four moments as inputs. Communications in Statistics: Simulation and Computation, 45(2):491–503, 2016. CODEN CSSCDB. ISSN 0361-0918.

Lee:2017:DTA

[LSY17] Keunbaik Lee, Hyejin Song, and Jae Keun Yoo. Dimension test approach of heteroscedasticity in the linear model. Communications in Statistics: Simulation and Computation, 46 (6):4356–4366, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lee:2015:NMT

[LT15] Pei-Hsi Lee and Chau-Chen Torng. A note on “Modified Tukey’s Control Chart”. Communications in Statistics: Simulation and Computation, 44(1):233–238, 2015. CODEN CSSCDB. ISSN 0361-0918. See [TGRGCF^{+}12].

Lio:2014:CCM

[LTAJ14] Y. L. Lio, Tzong-Ru Tsai, M. Aslam, and Nan Jiang. Control charts for monitoring Burr Type-X percentiles. Communications in Statistics: Simulation and Computation, 43 (4):761–776, 2014. CODEN CSSCDB. ISSN 0361-0918.

Tang:2013:CIA

[lTHT13] Man lai Tang, Xinjie He, and Guoliang Tian. A confidence interval approach for comparative studies involving binary outcomes in paired organs. Communications in Statistics: Simulation and Computation, 42(2):425–453, 2013. CODEN CSSCDB. ISSN 0361-0918.

Li:2011:CID

[LTPT11] Hui-Qiong Li, Man-Lai Tang, Wai-Yin Poon, and Nian-Sheng Tang. Confidence intervals for difference between two Poisson rates. Communications in Statistics: Simulation and Computation, 40(9):1478–1493, 2011. CODEN CSSCDB. ISSN 0361-0918.

Li:2015:SCM

[LTQ15] Fuxiao Li, Zheng Tian, and Peiyan Qi. Structural change monitoring for random coefficient autoregressive time series. Communications in Statistics: Simulation and Computation, 44(4):996–1009, 2015. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:POM

[LTZ17] Jian Liu, Chao Tan, and Chao Zeng. Parameter optimization for a modified MEWMA control chart based on a PSO algorithm. Communications in Statistics: Simulation and Computation, 46(5):4140–4153, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lu:2013:SEM

[Lu13] Yan Lu. Standard error of the method of simulated moment estimator for generalized linear mixed models. Communications in Statistics: Simulation and Computation, 42(1):1–7, 2013. CODEN CSSCDB. ISSN 0361-0918.

Lu:2014:RCE

[Lu14] Yan Lu. Regression coefficient estimation in dual frame surveys. Communications in Statistics: Simulation and Computation, 43(7):1675–1684, 2014. CODEN CSSCDB. ISSN 0361-0918.

Luepsen:2017:ART

[Lue17] Haiko Luepsen. The aligned rank transform and discrete variables: a warning. Communications in Statistics: Simulation and Computation, 46(9):6923–6936, 2017. CODEN CSSCDB. ISSN 0361-0918.

Luepsen:2018:CNA

[Lue18] Haiko Luepsen. Comparison of nonparametric analysis of variance methods: a vote for van der Waerden. Communications in Statistics: Simulation and Computation, 47(9): 2547–2576, 2018. CODEN CSSCDB. ISSN 0361-0918.

Lui:2016:NIE

[Lui16] Kung-Jong Lui. Notes on interval estimation of odds ratio in matched pairs under stratified sampling. Communications in Statistics: Simulation and Computation, 45(7):2562–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Lui:2017:NEP

[Lui17a] Kung-Jong Lui. Notes on estimation of the proportion ratio in the presence of carryover effects under the AB/BA crossover trial. Communications in Statistics: Simulation and Computation, 46(1):757–771, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lui:2017:NTC

[Lui17b] Kung-Jong Lui. Notes on testing carry-over effects in continuous data under an incomplete block crossover design. Communications in Statistics: Simulation and Computation, 46 (7):5229–5239, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lui:2019:NTS

[Lui19] Kung-Jong Lui. Notes on the two-stage procedure under an incomplete block two-period crossover design. Communications in Statistics: Simulation and Computation, 48(8): 2384–2400, 2019. CODEN CSSCDB. ISSN 0361-0918.

Li:2011:SMT

[LW11a] Xinmin Li and Juan Wang. A simple method for testing variance components in unbalanced nested model. Communications in Statistics: Simulation and Computation, 40(9): 1310–1323, 2011. CODEN CSSCDB. ISSN 0361-0918.

Lin:2011:BPO

[LW11b] Xiaoyan Lin and Lianming Wang. Bayesian proportional odds models for analyzing current status data: Univariate, clustered, and multivariate. Communications in Statistics: Simulation and Computation, 40(8):1171–1181, 2011. CO-DEN CSSCDB. ISSN 0361-0918.

Liu:2014:DCV

[LW14] Shih-Wen Liu and Chien-Wei Wu. Design and construction of a variables repetitive group sampling plan for unilateral specification limit. Communications in Statistics: Simulation and Computation, 43(8):1866–1878, 2014. CODEN CSSCDB. ISSN 0361-0918.

Lu:2016:IDD

[LW16] Tao Lu and Min Wang. Investigate data dependency for dynamic gene regulatory network identification through highdimensional differential equation approach. Communications in Statistics: Simulation and Computation, 45(7): 2377–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

LaMotte:2017:IPH

[LW17a] Lynn R. LaMotte and Jeffrey D. Wells. Inverse prediction for heteroscedastic response using mixed models software. Communications in Statistics: Simulation and Computation, 46 (6):4490–4498, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ledwina:2017:CPT

[LW17b] Teresa Ledwina and Grzegorz Wylupek. On Charlier polynomials in testing Poissonity. Communications in Statistics:

Simulation and Computation, 46(3):1918–1932, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Li:2015:TSC

[LWH^{+}15] Jianbo Li, Yuanyuan Wang, Zhensheng Huang, Riquan Zhang, and Lin Liu. Testing serial correlation in single index models. Communications in Statistics: Simulation and Computation, 44(4):1020–1035, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ling:2017:EHS

[LWT17] M. H. Ling, S. Y. Wong, and K. L. Tsui. Efficient heterogeneous sampling for stochastic simulation with an illustration in health care applications. Communications in Statistics: Simulation and Computation, 46(1):631–639, 2017. CODEN CSSCDB. ISSN 0361-0918.

Leung:2010:SSE

[LWZH10] Bartholomew Leung, Heung Wong, Riquan Zhang, and Shengtong Han. Smoothing spline estimation for partially linear single-index models. Communications in Statistics: Simulation and Computation, 39(10):1953–1961, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Luo:2013:YIA

[LX13] Jingqin Luo and Chengjie Xiong. Youden index and associated cut-points for three ordinal diagnostic groups. Communications in Statistics: Simulation and Computation, 42 (6):1213–1234, 2013. CODEN CSSCDB. ISSN 0361-0918.

Liang:2017:TLA

[LXPL17] Rubing Liang, Qiang Xia, Jiazhu Pan, and Jinshan Liu. Testing a linear ARMA model against threshold-ARMA models: A Bayesian approach. Communications in Statistics: Simulation and Computation, 46(2):1302–1317, 2017. CODEN CSSCDB. ISSN 0361-0918.

Li:2019:SIP

[LXS19] Xu Li, Juxia Xiao, and Jianhong Shi. Statistical inference in the partial linear models with the inverse Gaussian kernel. Communications in Statistics: Simulation and Computation, 48(1):240–263, 2019. CODEN CSSCDB. ISSN 0361-0918.

Li:2013:WSR

[LY13] Wenxue Li and Hu Yang. Weighted stochastic restricted estimation in linear measurement error models. Communications in Statistics: Simulation and Computation, 42(4): 932–968, 2013. CODEN CSSCDB. ISSN 0361-0918.

Li:2016:SSD

[LY16] Xiaomao Li and Jianxin Yin. Sparse sufficient dimension reduction for Markov blanket discovery. Communications in Statistics: Simulation and Computation, 45(4):1355–1364, 2016. CODEN CSSCDB. ISSN 0361-0918.

Liu:2017:EVS

[LY17] Huilan Liu and Hu Yang. Estimation and variable selection in single-index composite quantile regression. Communications in Statistics: Simulation and Computation, 46(9): 7022–7039, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lv:2017:VSP

[LYG17] Jing Lv, Hu Yang, and Chaohui Guo. Variable selection in partially linear additive models for modal regression. Communications in Statistics: Simulation and Computation, 46 (7):5646–5665, 2017. CODEN CSSCDB. ISSN 0361-0918.

Lien:2014:AAV

[LYY14] Donald Lien, Xiaobin Yang, and Keying Ye. Alternative approximations to value-at-risk: A comparison. Communications in Statistics: Simulation and Computation, 43(10): 2225–2240, 2014. CODEN CSSCDB. ISSN 0361-0918.

Li:2010:PSV

[LZ10] Jianbo Li and Riquan Zhang. Penalized spline varyingcoefficient single-index model. Communications in Statistics: Simulation and Computation, 39(2):221–239, 2010. CODEN CSSCDB. ISSN 0361-0918.

Liu:2014:CHD

[LZ14] Xiaohui Liu and Yijun Zuo. Computing halfspace depth and regression depth. Communications in Statistics: Simulation and Computation, 43(5):969–985, 2014. CODEN CSSCDB. ISSN 0361-0918.

Luo:2016:SDE

[LZ16] Chun Luo and Yingshan Zhang. Symmetrical design of experiment in global sensitivity analysis based on ANOVA high-dimensional model representation. Communications in Statistics: Simulation and Computation, 45(1):48–72, 2016. CODEN CSSCDB. ISSN 0361-0918.

Li:2018:SIH

[LZ18] Rui Li and Haibing Zhao. Sparsity identification for highdimensional partially linear model with measurement error. Communications in Statistics: Simulation and Computation, 47(8):2378–2392, 2018. CODEN CSSCDB. ISSN 0361-0918.

Liu:2010:LLR

[LZS10] Songhua Liu, Junying Zhang, and Keguo Sun. Learning lowrank kernel matrices with column-based methods. Communications in Statistics: Simulation and Computation, 39(7): 1485–1498, 2010. CODEN CSSCDB. ISSN 0361-0918.

Liu:2015:DNC

[LZZ15] Liu Liu, Jian Zhang, and Xuemin Zi. Dual nonparametric CUSUM control chart based on ranks. Communications in Statistics: Simulation and Computation, 44(3):756–772, 2015. CODEN CSSCDB. ISSN 0361-0918.

Li:2016:FAA

[LZZ16] Shilong Li, Xia Zhao, and Jingxiao Zhang. Fractional age assumption based on cubic polynomial interpolation. Communications in Statistics: Simulation and Computation, 45 (4):1173–1186, 2016. CODEN CSSCDB. ISSN 0361-0918.

Liu:2013:SRB

[LZZW13] Liu Liu, Xuemin Zi, Jian Zhang, and Zhaojun Wang. A sequential rank-based nonparametric adaptive EWMA control chart. Communications in Statistics: Simulation and Computation, 42(4):841–859, 2013. CODEN CSSCDB. ISSN 0361-0918.

Mahdizadeh:2012:QEU

[MA12] M. Mahdizadeh and N. R. Arghami. Quantile estimation using ranked set samples from a population with known mean. Communications in Statistics: Simulation and Computation, 41(10):1872–1881, 2012. CODEN CSSCDB. ISSN 0361-0918.

Mansouri:2016:DMF

[MA16] Behzad Mansouri and Mitra Abasnezhad. Discrimination measures for fractional integrated models based on wavelets. Communications in Statistics: Simulation and Computation, 45(9):3476–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Macci:2011:LDR

[Mac11a] Claudio Macci. Large deviation results for wave governed random motions driven by semi-Markov processes. Communications in Statistics: Simulation and Computation, 40(9): 1342–1363, 2011. CODEN CSSCDB. ISSN 0361-0918.

Marcon:2011:GKI

[MAC11b] G. Marcon, G. Adelfio, and M. Chiodi. Gamma kernel intensity estimation in temporal point processes. Communications in Statistics: Simulation and Computation, 40(8): 1146–1162, 2011. CODEN CSSCDB. ISSN 0361-0918.

Macedo:2017:RRG

[Mac17] Pedro Macedo. Ridge Regression and Generalized Maximum Entropy: An improved version of the Ridge–GME parameter estimator. Communications in Statistics: Simulation and Computation, 46(5):3527–3539, 2017. CODEN CSSCDB. ISSN 0361-0918.

Milani:2017:GNA

[MAD17] Eder Milani, Marinho Andrade, and Carlos Diniz. Generalized normal ARMA model applied to the areas of economy, hydrology, and public policy. Communications in Statistics: Simulation and Computation, 46(7):5819–5835, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Mahmoud:2012:PPI

[Mah12] Mahmoud A. Mahmoud. The performance of Phase II simple linear profile approaches when parameters are estimated. Communications in Statistics: Simulation and Computation, 41(10):1816–1833, 2012. CODEN CSSCDB. ISSN 0361-0918.

Mahdizadeh:2015:EBT

[Mah15] M. Mahdizadeh. On entropy-based test of exponentiality in ranked set sampling. Communications in Statistics: Simulation and Computation, 44(4):979–995, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Mahdizadeh:2017:TUU

[Mah17] M. Mahdizadeh. On testing uniformity using an informationtheoretic measure. Communications in Statistics: Simulation and Computation, 46(8):6173–6196, 2017. CODEN CSSCDB. ISSN 0361-0918.

Maiti:2017:LST

[Mai17] Saran Ishika Maiti. Large sample test criteria on Toeplitz covariance structure. Communications in Statistics: Simulation and Computation, 46(1):1–15, 2017. CODEN CSSCDB. ISSN 0361-0918.

Maki:2011:PEC

[Mak11] Daiki Maki. Pitfalls in estimating cointegrating vector when cointegration relationship has nonlinear adjustment. Communications in Statistics: Simulation and Computation, 40 (8):1111–1121, 2011. CODEN CSSCDB. ISSN 0361-0918.

Moeini:2011:CDI

[MAM11] A. Moeini, B. Abbasi, and H. Mahlooji. Conditional distribution inverse method in generating uniform random vectors over a simplex. Communications in Statistics: Simulation and Computation, 40(5):685–693, 2011. CODEN CSSCDB. ISSN 0361-0918.

Moreira:2017:EBD

[MAMM17] Ana Moreira, Artur Araújo, and Luís Meira-Machado. Estimation of the bivariate distribution function for censored gap times. Communications in Statistics: Simulation and Computation, 46(1):275–300, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mandal:2015:CEF

[Man15a] B. N. Mandal. Combinatorial equivalence of fractional factorial designs. Communications in Statistics: Simulation and Computation, 44(4):1010–1019, 2015. CODEN CSSCDB. ISSN 0361-0918.

Mandal:2015:LIP

[Man15b] B. N. Mandal. Linear integer programming approach to construction of balanced incomplete block designs. Communications in Statistics: Simulation and Computation, 44(6): 1405–1411, 2015. CODEN CSSCDB. ISSN 0361-0918.

Mansoor:2017:UPC

[Man17] Rashid Mansoor. Using principal components to test normality of high-dimensional data. Communications in Statistics: Simulation and Computation, 46(5):3396–3405, 2017. CODEN CSSCDB. ISSN 0361-0918.

Munir:2011:PED

[MAR11] Shahzad Munir, Zahid Asghar, and Muhammad Riaz. Performance evaluation of different tests for location parameters. Communications in Statistics: Simulation and Computation, 40(6):839–853, 2011. CODEN CSSCDB. ISSN 0361-0918.

Marozzi:2012:DFT

[Mar12] Marco Marozzi. A distribution free test for the equality of scales. Communications in Statistics: Simulation and Computation, 41(6):878–889, 2012. CODEN CSSCDB. ISSN 0361-0918.

Marozzi:2013:NST

[Mar13a] Marco Marozzi. Nonparametric simultaneous tests for location and scale testing: A comparison of several methods. Communications in Statistics: Simulation and Computation, 42(6):1298–1317, 2013. CODEN CSSCDB. ISSN 0361-0918.

Martin:2013:ABM

[Mar13b] Ryan Martin. An approximate Bayesian marginal likelihood approach for estimating finite mixtures. Communications in Statistics: Simulation and Computation, 42(7):1533–1548, 2013. CODEN CSSCDB. ISSN 0361-0918.

Martin:2015:VDS

[Mar15] D. E. K. Martin. p-values for the discrete scan statistic through slack variables. Communications in Statistics: Simulation and Computation, 44(9):2223–2239, 2015. CODEN CSSCDB. ISSN 0361-0918.

Martin:2019:MAM

[Mar19] Donald E. K. Martin. Minimal auxiliary Markov chains through sequential elimination of states. Communications in Statistics: Simulation and Computation, 48(4):1040–1054, 2019. CODEN CSSCDB. ISSN 0361-0918.

Maleki:2017:PIM

[MAT17] Mohammad Reza Maleki, Amirhossein Amiri, and Ali Reza Taheriyoun. Phase II monitoring of binary profiles in the presence of within-profile autocorrelation based on Markov model. Communications in Statistics: Simulation and Computation, 46(10):7710–7732, 2017. CODEN CSSCDB. ISSN 0361-0918.

Matsui:2019:SGL

[Mat19] Hidetoshi Matsui. Sparse group lasso for multiclass functional logistic regression models. Communications in Statistics: Simulation and Computation, 48(6):1784–1797, 2019. CODEN CSSCDB. ISSN 0361-0918.

Mazumder:2017:SSM

[Maz17] Satyaki Mazumder. Single-step and multiple-step forecasting in one-dimensional single chirp signal using MCMCbased Bayesian analysis. Communications in Statistics: Simulation and Computation, 46(4):2529–2547, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Manalaysay:2017:SPC

[MB17] Kristina Celene M. Manalaysay and Erniel B. Barrios. Semiparametric principal component Poisson regression on clustered data. Communications in Statistics: Simulation and Computation, 46(2):1546–1556, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mehrdoust:2017:EMC

[MBF17] F. Mehrdoust, S. Babaei, and S. Fallah. Efficient Monte Carlo option pricing under CEV model. Communications in Statistics: Simulation and Computation, 46(3):2254–2266, 2017. CODEN CSSCDB. ISSN 0361-0918.

Martinez:2012:HMR

[MBL12] Zoraida C. Martínez, Daniel Barráez, and Carenne C. Ludeñez. A hybrid method for random pattern sequence classification. Communications in Statistics: Simulation and Computation, 41(1):24–31, 2012. CODEN CSSCDB. ISSN 0361-0918.

Mara:2012:PST

[MC12] Constance A. Mara and Robert A. Cribbie. Paired-samples tests of equivalence. Communications in Statistics: Simulation and Computation, 41(10):1928–1943, 2012. CODEN CSSCDB. ISSN 0361-0918.

Mohammadinia:2019:CBM

[MC19] Farideh Mohammadinia and Rahim Chinipardaz. Combining Bayesian method and Kalman smoother for detection additive outlier patches in autoregressive time series. Communications in Statistics: Simulation and Computation, 48 (7):2191–2209, 2019. CODEN CSSCDB. ISSN 0361-0918.

Maleki:2019:EPE

[MCAK19] M. R. Maleki, P. Castagliola, A. Amiri, and Michael B. C. Khoo. The effect of parameter estimation on phase II monitoring of Poisson regression profiles. Communications in Statistics: Simulation and Computation, 48(7):1964–1978, 2019. CODEN CSSCDB. ISSN 0361-0918.

Modak:2018:CGR

[MCC18] Soumita Modak, Asis Kumar Chattopadhyay, and Tanuka Chattopadhyay. Clustering of gamma-ray bursts through kernel principal component analysis. Communications in Statistics: Simulation and Computation, 47(4):1088–1102, 2018. CODEN CSSCDB. ISSN 0361-0918.

McIntosh:2014:IPA

[McI14] James McIntosh. Inference problems in the analysis of the relationship between alcohol consumption and coronary heart disease. Communications in Statistics: Simulation and Computation, 43(10):2275–2282, 2014. CODEN CSSCDB. ISSN 0361-0918.

McIntosh:2018:NRI

[McI18] James McIntosh. Normalization rules for item response models: Simulation results. Communications in Statistics: Simulation and Computation, 47(10):3095–3105, 2018. CODEN CSSCDB. ISSN 0361-0918.

Ma:2016:MAP

[MCM16] Yan Ma, Haitao Chu, and Madhu Mazumdar. Meta-analysis of proportions of rare events — a comparison of exact likelihood methods with robust variance estimation. Communications in Statistics: Simulation and Computation, 45(8): 3036–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Mostajabi:2013:PPS

[MDD13] Farida Mostajabi, Somnath Datta, and Susmita Datta. Predicting patient survival from proteomic profile using mass spectrometry data: An empirical study. Communications in Statistics: Simulation and Computation, 42(3):485–498, 2013. CODEN CSSCDB. ISSN 0361-0918.

Mu:2017:SDR

[MDX17] Rongji Mu, Lixiang Dai, and Jin Xu. Sequential design for response surface model fit in computer experiments using derivative information. Communications in Statistics: Simulation and Computation, 46(2):1148–1155, 2017. CODEN CSSCDB. ISSN 0361-0918.

Melas:2012:Pa

[Mel12a] Viatcheslav B. Melas. Preface. Communications in Statistics: Simulation and Computation, 41(6):687, 2012. CODEN CSSCDB. ISSN 0361-0918.

Melas:2012:Pb

[Mel12b] Viatcheslav B. Melas. Preface. Communications in Statistics: Simulation and Computation, 41(7):943, 2012. CODEN CSSCDB. ISSN 0361-0918.

Merzougui:2018:EPR

[Mer18] M. Merzougui. Estimation in periodic restricted EXPAR(1) models. Communications in Statistics: Simulation and Computation, 47(10):2819–2828, 2018. CODEN CSSCDB. ISSN 0361-0918.

Mokarram:2018:CPS

[MERN18] Reza Mokarram, Mahdi Emadi, Arezou Habibi Rad, and Mahdi Jabbari Nooghabi. A comparison of parametric and semi-parametric survival models with artificial neural networks. Communications in Statistics: Simulation and Computation, 47(3):738–746, 2018. CODEN CSSCDB. ISSN 0361-0918.

Meyer:2013:SNA

[Mey13] Mary C. Meyer. A simple new algorithm for quadratic programming with applications in statistics. Communications in Statistics: Simulation and Computation, 42(5): 1126–1139, 2013. CODEN CSSCDB. ISSN 0361-0918.

Malekzadeh:2012:SAG

[MG12] P. Malekzadeh and S. D. Gore. A stochastic approach to general nonlinear programming problems. Communications in Statistics: Simulation and Computation, 41(10):1989–1997, 2012. CODEN CSSCDB. ISSN 0361-0918.

Moradzadeh:2017:WKF

[MGB17] N. Moradzadeh, M. Ganjali, and T. Baghfalaki. Weighted kappa as a function of unweighted kappas. Communications in Statistics: Simulation and Computation, 46(5): 3769–3780, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mukherjee:2013:DFE

[MGC13] A. Mukherjee, M. A. Graham, and S. Chakraborti. Distribution-free exceedance CUSUM control charts for location. Communications in Statistics: Simulation and Computation, 42(5):1153–1187, 2013. CODEN CSSCDB. ISSN 0361-0918.

Maurya:2011:MCM

[MGG11] Vishal Maurya, Anju Goyal, and Amar Nath Gill. Multiple comparisons with more than one control for exponential location parameters under heteroscedasticity. Communications in Statistics: Simulation and Computation, 40(5):621– 644, 2011. CODEN CSSCDB. ISSN 0361-0918.

Mazucheli:2017:CTE

[MGL17] J. Mazucheli, M. E. Ghitany, and F. Louzada. Comparisons of ten estimation methods for the parameters of Marshall–Olkin extended exponential distribution. Communications in Statistics: Simulation and Computation, 46(7): 5627–5645, 2017. CODEN CSSCDB. ISSN 0361-0918.

Martinez:2014:AEA

[MH14] Marie-José Martinez and Emma Holian. An alternative estimation approach for the heterogeneity linear mixed model.

Communications in Statistics: Simulation and Computation, 43(10):2628–2638, 2014. CODEN CSSCDB. ISSN 0361-0918.

Makinen:2017:MPD

[MH17a] Terhi Mäkinen and Lasse Holmström. Modeling probability density through ultraspherical polynomial transformations. Communications in Statistics: Simulation and Computation, 46(8):5879–5900, 2017. CODEN CSSCDB. ISSN 0361-0918.

McNeish:2017:CDS

[MH17b] Daniel M. McNeish and Jeffery R. Harring. Clustered data with small sample sizes: Comparing the performance of model-based and design-based approaches. Communications in Statistics: Simulation and Computation, 46(2):855–869, 2017. CODEN CSSCDB. ISSN 0361-0918.

Milani:2019:BAM

[MHAD19] Eder Angelo Milani, Marcelo Hartmann, Marinho G. Andrade, and Carlos Alberto Ribeiro Diniz. A Bernoulli autoregressive moving average model applied to rainfall occurrence. Communications in Statistics: Simulation and Computation, 48(9):2743–2756, 2019. CODEN CSSCDB. ISSN 0361-0918.

Mielke:2012:ODP

[Mie12] Tobias Mielke. D-optimal designs for paired observations in quadratic regression with mixed effects. Communications in Statistics: Simulation and Computation, 41(7):1107–1119, 2012. CODEN CSSCDB. ISSN 0361-0918.

Mahmoudi:2017:CCL

[MJ17] Eisa Mahmoudi and A. A. Jafari. The compound class of linear failure rate-power series distributions: Model, properties, and applications. Communications in Statistics: Simulation and Computation, 46(2):1414–1440, 2017. CODEN CSSCDB. ISSN 0361-0918.

Moghadam:2018:ENN

[MKFP18] M. Bameni Moghadam, Y. Khadem, S. Fani, and M. A. Pasha. Effects of non-normality on economic and economic statistical designs of -control charts with multiple assignable causes and Weibull in-control times. Communications in Statistics: Simulation and Computation, 47(7):2055–2069, 2018. CODEN CSSCDB. ISSN 0361-0918.

Mannseth:2018:AIS

[MKS18] Janne Mannseth, Tore S. Kleppe, and Hans J. Skaug. On the application of improved symplectic integrators in Hamiltonian Monte Carlo. Communications in Statistics: Simulation and Computation, 47(2):500–509, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Monfared:2017:BEC

[ML17] M. Esmail Dehghan Monfared and Fazlollah Lak. Bayesian estimation of the change point in a gamma process using control chart. Communications in Statistics: Simulation and Computation, 46(3):2333–2345, 2017. CODEN CSSCDB. ISSN 0361-0918.

Martino:2019:ARS

[ML19a] L. Martino and F. Louzada. Adaptive rejection sampling with fixed number of nodes. Communications in Statistics: Simulation and Computation, 48(3):655–665, 2019. CODEN CSSCDB. ISSN 0361-0918.

Martino:2019:EEA

[ML19b] Luca Martino and David Luengo. Extremely efficient acceptance-rejection method for simulating uncorrelated Nakagami fading channels. Communications in Statistics: Simulation and Computation, 48(6):1798–1814, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Miao:2016:SNL

[MLC16] Daniel Wei-Chung Miao, Hsiu-Chun Lee, and Hung Chen. A standardized normal-Laplace mixture distribution fitted to symmetric implied volatility smiles. Communications in Statistics: Simulation and Computation, 45(4):1249–1267, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ma:2015:ABT

[MLZ15] Xiao-Bing Ma, Feng-Chun Lin, and Yu Zhao. An adjustment to the Bartlett’s test for small sample size. Communications in Statistics: Simulation and Computation, 44(1):257–269, 2015. CODEN CSSCDB. ISSN 0361-0918.

Mahmoud:2010:PMC

[MM10] M. A. Mahmoud and P. E. Maravelakis. The performance of the MEWMA control chart when parameters are estimated. Communications in Statistics: Simulation and Computation, 39(9):1803–1817, 2010. CODEN CSSCDB. ISSN 0361-0918.

Minh:2015:UHA

[MM15] David D. L. Minh and Do Le (Paul) Minh. Understanding the Hastings algorithm. Communications in Statistics: Simulation and Computation, 44(2):332–349, 2015. CODEN CSSCDB. ISSN 0361-0918.

Michael:2016:SCM

[MM16] Semhar Michael and Volodymyr Melnykov. Studying complexity of model-based clustering. Communications in Statistics: Simulation and Computation, 45(6):2051–2069, 2016. CODEN CSSCDB. ISSN 0361-0918.

Moghaddam:2017:NMB

[MM17] Samira Moghaddam and Hashem Mahlooji. A new metamodel-based method for solving semi-expensive simulation optimization problems. Communications in Statistics: Simulation and Computation, 46(6):4795–4811, 2017. CODEN CSSCDB. ISSN 0361-0918.

Maiti:2018:EPC

[MM18a] Sudhansu S. Maiti and Indrani Mukherjee. On estimation of the PDF and CDF of the Lindley distribution. Communications in Statistics: Simulation and Computation, 47(5): 1370–1381, 2018. CODEN CSSCDB. ISSN 0361-0918.

Masmoudi:2018:SGG

[MM18b] Khalil Masmoudi and Afif Masmoudi. Singular Gaussian graphical models: Structure learning. Communications in Statistics: Simulation and Computation, 47(10):3106–3117, 2018. CODEN CSSCDB. ISSN 0361-0918.

Mitra:2017:SEN

[MMB17] Sharmishtha Mitra, Amit Mitra, and Sanket Bose. Simultaneous estimation of number of signals and signal parameters of superimposed sinusoidal model: A robust sequential bivariate m-periodogram approach. Communications in Statistics: Simulation and Computation, 46(4):2654–2677, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mohsenkhani:2019:AMB

[MMB19] Zohreh Fallah Mohsenkhani, Mohsen Mohhamadzadeh, and Taban Baghfalaki. Augmented mixed beta regression models with skew-normal independent distributions: Bayesian analysis of labor force data. Communications in Statistics: Simulation and Computation, 48(7):2147–2164, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Mazucheli:2019:BCM

[MMD19] Josmar Mazucheli, André Felipe Berdusco Menezes, and Sanku Dey. Bias-corrected maximum likelihood estimators of the parameters of the inverse Weibull distribution. Communications in Statistics: Simulation and Computation, 48 (7):2046–2055, 2019. CODEN CSSCDB. ISSN 0361-0918.

Meira-Machado:2014:SSC

[MMF14] Luís Meira-Machado and Susana Faria. A simulation study comparing modeling approaches in an illness-death multistate model. Communications in Statistics: Simulation and Computation, 43(5):929–946, 2014. CODEN CSSCDB. ISSN 0361-0918.

Mantalos:2010:IDI

[MMK10] P. Mantalos, K. Mattheou, and A. Karagrigoriou. An improved divergence information criterion for the determination of the order of an AR process. Communications in Statistics: Simulation and Computation, 39(5):865–879, 2010. CODEN CSSCDB. ISSN 0361-0918.

Madbuly:2013:EMH

[MMM13] Doaa F. Madbuly, Petros E. Maravelakis, and Mahmoud A. Mahmoud. The effect of methods for handling missing values on the performance of the MEWMA control chart. Communications in Statistics: Simulation and Computation, 42(6): 1437–1454, 2013. CODEN CSSCDB. ISSN 0361-0918.

Martelli:2016:BEM

[MMM16] Irene Martelli, Mariagiulia Matteucci, and Stefania Mignani. Bayesian estimation of a multidimensional additive graded response model for correlated traits. Communications in Statistics: Simulation and Computation, 45(5):1636–1654, 2016. CODEN CSSCDB. ISSN 0361-0918.

Minh:2012:RMC

[MMN12] Do Le (Paul) Minh, David D. L. Minh, and Andrew L. Nguyen. Regenerative Markov chain Monte Carlo for any distribution. Communications in Statistics: Simulation and Computation, 41(9):1745–1760, 2012. CODEN CSSCDB. ISSN 0361-0918.

Minh:2016:LS

[MMN16] David D. L. Minh, Do Le (Paul) Minh, and Andrew L. Nguyen. Layer sampling. Communications in Statistics: Simulation and Computation, 45(1):73–100, 2016. CODEN CSSCDB. ISSN 0361-0918.

Menon:2013:COC

[MMP^{+}13] Sandeep Menon, Joseph Massaro, Michael J. Pencina, Jerry Lewis, and Yong Cheng Wang. Comparison of operating characteristics of commonly used sample size re-estimation procedures in a two-stage design. Communications in Statistics: Simulation and Computation, 42(5):1140–1152, 2013. CODEN CSSCDB. ISSN 0361-0918.

Malik:2016:CBT

[MMP16] W. A. Malik, J. Möhring, and H. P. Piepho. A clusteringbased test for nonadditivity in an unreplicated two-way layout. Communications in Statistics: Simulation and Computation, 45(2):660–670, 2016. CODEN CSSCDB. ISSN 0361-0918.

Monari:2016:P

[MMS16] Paola Monari, Viatcheslav B. Melas, and Luigi Salmaso. Preface. Communications in Statistics: Simulation and Computation, 45(5):1429, 2016. CODEN CSSCDB. ISSN 0361-0918.

Mameli:2018:BAS

[MMV18] V. Mameli, M. Musio, and L. Ventura. Bootstrap adjustments of signed scoring rule root statistics. Communications in Statistics: Simulation and Computation, 47(4): 1204–1215, 2018. CODEN CSSCDB. ISSN 0361-0918.

McDonald:2015:HDA

[MN15] James B. McDonald and Hieu Nguyen. Heteroscedasticity and distributional assumptions in the censored regression model. Communications in Statistics: Simulation and Computation, 44(8):2151–2168, 2015. CODEN CSSCDB. ISSN 0361-0918.

Morimoto:2017:REH

[MN17a] Takayuki Morimoto and Shuichi Nagata. Robust estimation of a high-dimensional integrated covariance matrix. Communications in Statistics: Simulation and Computation, 46 (2):1102–1112, 2017. CODEN CSSCDB. ISSN 0361-0918.

Morina:2017:CRS

[MN17b] D. Moriña and A. Navarro. Competing risks simulation with the survsim R package. Communications in Statistics: Simulation and Computation, 46(7):5712–5722, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mirniam:2018:MLE

[MN18] A. S. Mirniam and A. R. Nematollahi. Maximum likelihood estimation in vector autoregressive models with multivariate scaled t-distributed innovations using EM-based algorithms. Communications in Statistics: Simulation and Computation, 47(3):890–904, 2018. CODEN CSSCDB. ISSN 0361-0918.

Moheghi:2017:EIN

[MNBB17] Hamidreza Moheghi, Seyed Taghi Akhavan Niaki, Behrang Bootaki, and Davood Bakhshesh. On the effect of inducted negative correlation rate for beta acceptance-rejection algorithms. Communications in Statistics: Simulation and Computation, 46(3):2152–2167, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mkhadri:2010:MAC

[MNH10] Abdallah Mkhadri, Assi N’Guessan, and Bezza Hafidi. An MM algorithm for constrained estimation in a road safety measure modeling. Communications in Statistics: Simulation and Computation, 39(5):1057–1071, 2010. CODEN CSSCDB. ISSN 0361-0918.

Mahmoudvand:2013:OPR

[MNZ13] Rahim Mahmoudvand, Nader Najari, and Mohammad Zokaei. On the optimal parameters for reconstruction and forecasting in singular spectrum analysis. Communications in Statistics: Simulation and Computation, 42(4):860–870, 2013. CODEN CSSCDB. ISSN 0361-0918.

Miljkovic:2017:ERC

[MO17] Tatjana Miljkovic and Megan Orr. An evaluation of the reconstructed coefficient of determination and potential adjustments. Communications in Statistics: Simulation and Computation, 46(9):6705–6718, 2017. CODEN CSSCDB. ISSN 0361-0918.

Madukaife:2018:PAI

[MO18] Mbanefo S. Madukaife and Fabian C. Okafor. A powerful affine invariant test for multivariate normality based on interpoint distances of principal components. Communications in Statistics: Simulation and Computation, 47(5): 1264–1275, 2018. CODEN CSSCDB. ISSN 0361-0918.

Madukaife:2019:NLS

[MO19] Mbanefo S. Madukaife and Fabian C. Okafor. A new large sample goodness of fit test for multivariate normality based on chi squared probability plots. Communications in Statistics: Simulation and Computation, 48(6):1651–1664, 2019. CODEN CSSCDB. ISSN 0361-0918.

Moradi:2019:UAL

[Mor19] M. Moradi. Using adaptive line-transect sampling in airborne geophysics studies. Communications in Statistics: Simulation and Computation, 48(3):675–683, 2019. CODEN CSSCDB. ISSN 0361-0918.

Md:2018:NQE

[MOS18] Augustus Jayaraj Md, Oluseun Odumade, and Sarjinder Singh. A new quasi empirical Bayes estimate in randomized response sampling. Communications in Statistics: Simulation and Computation, 47(7):1879–1889, 2018. CODEN CSSCDB. ISSN 0361-0918.

Mandal:2013:ODO

[MP13a] N. K. Mandal and Manisha Pal. Optimum designs for optimum mixtures in multiresponse experiments. Communications in Statistics: Simulation and Computation, 42(5): 1104–1112, 2013. CODEN CSSCDB. ISSN 0361-0918.

Mandal:2013:OMD

[MP13b] Nripes Kumar Mandal and Manisha Pal. Optimum mixture design for the detection of synergistic effects. Communications in Statistics: Simulation and Computation, 42(3): 711–726, 2013. CODEN CSSCDB. ISSN 0361-0918.

Macci:2017:ABS

[MP17a] Claudio Macci and Barbara Pacchiarotti. Asymptotic behavior of some hitting probabilities for sums of IID Gaussian random sets. Communications in Statistics: Simulation and Computation, 46(6):4318–4332, 2017. CODEN CSSCDB. ISSN 0361-0918.

Min:2017:BVS

[MP17b] S. Min and T. Park. Bayesian variable selection in Poisson change-point regression analysis. Communications in Statistics: Simulation and Computation, 46(3):2267–2282, 2017. CODEN CSSCDB. ISSN 0361-0918.

Majakwara:2019:SII

[MP19a] Jacob Majakwara and Suvra Pal. On some inferential issues for the destructive COM–Poisson-generalized gamma regression cure rate model. Communications in Statistics: Simulation and Computation, 48(10):3118–3142, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Mingoti:2019:SDA

[MP19b] Sueli A. Mingoti and Letícia P. Pinto. Step-down diagnostic analysis for monitoring the covariance matrix of bivariate normal processes. Communications in Statistics: Simulation and Computation, 48(9):2615–2624, 2019. CODEN CSSCDB. ISSN 0361-0918.

Munoz-Pichardo:2011:IAD

[MPEGMGMR11] J. M. Muñoz-Pichardo, A. Enguix-González, J. Muñoz-García, and J. L. Moreno-Rebollo. Influence analysis on discriminant coordinates. Communications in Statistics: Simulation and Computation, 40(6):793–807, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Mohsin:2016:EDM

[MPG16] Muhammad Mohsin, Jürgen Pilz, and Albrecht Gebhardt. An explicit distribution to model the proportion of heating degree day and cooling degree day. Communications in Statistics: Simulation and Computation, 45(7):2617–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Maruotti:2014:BCZ

[MR14] Antonello Maruotti and Valentina Raponi. On baseline conditions for zero-inflated longitudinal count data. Communications in Statistics: Simulation and Computation, 43(4): 743–760, 2014. CODEN CSSCDB. ISSN 0361-0918.

Marozzi:2018:PSA

[MR18] Marco Marozzi and Jenő Reiczigel. A progressive shift alternative to evaluate nonparametric tests for skewed data. Communications in Statistics: Simulation and Computation, 47(10):3083–3094, 2018. CODEN CSSCDB. ISSN 0361-0918.

Munir:2018:EPD

[MRA18] Shahzad Munir, Muhammad Riaz, and Zahid Asghar. On evaluating the performance of different forecasters. Communications in Statistics: Simulation and Computation, 47(2): 542–555, 2018. CODEN CSSCDB. ISSN 0361-0918.

Marra:2017:RSB

[MRF17] Giampiero Marra, Rosalba Radice, and Panagiota Filippou. Regression spline bivariate probit models: A practical approach to testing for exogeneity. Communications in Statistics: Simulation and Computation, 46(3):2283–2298, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mohamed:2016:NDT

[MRKI16] I. B. Mohamed, A. Rambli, N. Khaliddin, and A. I. N. Ibrahim. A new discordancy test in circular data using spacings theory. Communications in Statistics: Simulation and Computation, 45(8):2904–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Moustafa:2010:NDF

[MRM10] H. M. Moustafa, S. G. Ramadan, and S. A. Mohammed. Nonlinear discriminant functions for mixed random walk models. Communications in Statistics: Simulation and Computation, 39(10):1923–1938, 2010. CODEN CSSCDB. ISSN 0361-0918.

Mombeni:2017:LDB

[MRN17] H. Mombeni, S. Rezaei, and S. Nadarajah. Linex discrepancy for bandwidth selection. Communications in Statistics: Simulation and Computation, 46(7):5054–5069, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Mukhopadhyay:2010:PIT

[MS10] Purna Mukhopadhyay and V. A. Samaranayake. Prediction intervals for time series: A modified sieve bootstrap approach. Communications in Statistics: Simulation and Computation, 39(3):517–538, 2010. CODEN CSSCDB. ISSN 0361-0918.

Muthukumarana:2014:BAO

[MS14] Saman Muthukumarana and Tim B. Swartz. Bayesian analysis of ordinal survey data using the Dirichlet process to account for respondent personality traits. Communications in Statistics: Simulation and Computation, 43(1):82–98, 2014. CODEN CSSCDB. ISSN 0361-0918.

Morrell:2011:CSA

[MSB11] Christopher H. Morrell, Shan L. Sheng, and Larry J. Brant. A comparative study of approaches for predicting prostate cancer from longitudinal data. Communications in Statistics: Simulation and Computation, 40(9):1494–1513, 2011. CODEN CSSCDB. ISSN 0361-0918.

Moslemi:2018:RSE

[MSEN18] Amir Moslemi, Mirmehdi Seyyed-Esfahani, and Seyed Taghi Akhavan Niaki. Robust surface estimation in multiresponse multistage statistical optimization problems. Communications in Statistics: Simulation and Computation, 47 (3):762–782, 2018. CODEN CSSCDB. ISSN 0361-0918.

Maansson:2010:SSS

[MSK10] Kristofer Månsson, Ghazi Shukur, and B. M. Golam Kibria. A simulation study of some ridge regression estimators under different distributional assumptions. Communications in Statistics: Simulation and Computation, 39(8):1639–1670, 2010. CODEN CSSCDB. ISSN 0361-0918.

Mita:2017:MCT

[MSK^{+}17] Norihiro Mita, Hiroshi Sasaki, Kazutaka Kani, Akio Tabuchi, and Heihachiro Hara. Multiple comparisons in two-factor testing of psychophysical thresholds by the delta test method. Communications in Statistics: Simulation and Computation, 46(1):669–678, 2017. CODEN CSSCDB. ISSN 0361-0918.

Miretskiy:2012:EMS

[MSM12] D. I. Miretskiy, W. R. W. Scheinhardt, and M. R. H. Mandjes. On efficiency of multilevel splitting. Communications in Statistics: Simulation and Computation, 41(6):890–904, 2012. CODEN CSSCDB. ISSN 0361-0918.

Mohammadi:2012:BIP

[MSR12] A. Mohammadi and M. R. Salehi-Rad. Bayesian inference and prediction in an M/G/1 with optional second service. Communications in Statistics: Simulation and Computation, 41(3):419–435, 2012. CODEN CSSCDB. ISSN 0361-0918.

Macedo:2010:GCE

[MSS10a] Pedro Macedo, Manuel Scotto, and Elvira Silva. A general class of estimators for the linear regression model affected by collinearity and outliers. Communications in Statistics: Simulation and Computation, 39(5):981–993, 2010. CODEN CSSCDB. ISSN 0361-0918.

Macedo:2010:CRP

[MSS10b] Pedro Macedo, Manuel Scotto, and Elvira Silva. On the choice of the ridge parameter: A maximum entropy approach. Communications in Statistics: Simulation and Computation, 39(8):1628–1638, 2010. CODEN CSSCDB. ISSN 0361-0918.

Macedo:2016:RME

[MSS16] Pedro Macedo, Manuel Scotto, and Elvira Silva. Regularization with maximum entropy and quantum electrodynamics: The merg(e) estimators. Communications in Statistics: Simulation and Computation, 45(3):1143–1157, 2016. CODEN CSSCDB. ISSN 0361-0918.

Mohamed:2017:IUH

[MSS17] Choukri Mohamed, Stephen A. Sedory, and Sarjinder Singh. Imputation using higher order moments of an auxiliary variable. Communications in Statistics: Simulation and Computation, 46(8):6588–6617, 2017. CODEN CSSCDB. ISSN 0361-0918.

Maboudou-Tchao:2018:KMC

[MT18] Edgard M. Maboudou-Tchao. Kernel methods for changes detection in covariance matrices. Communications in Statistics: Simulation and Computation, 47(6):1704–1721, 2018. CODEN CSSCDB. ISSN 0361-0918.

Magel:2010:NTM

[MTCP10] Rhonda Magel, Jeff Terpstra, Katrina Canonizado, and Ja In Park. Nonparametric tests for mixed designs. Communications in Statistics: Simulation and Computation, 39(6): 1228–1250, 2010. CODEN CSSCDB. ISSN 0361-0918.

Malefaki:2010:ESV

[MTL10] Sonia Malefaki, Samis Trevezas, and Nikolaos Limnios. An EM and a stochastic version of the EM algorithm for nonparametric hidden semi-Markov models. Communications in Statistics: Simulation and Computation, 39(2):240–261, 2010. CODEN CSSCDB. ISSN 0361-0918.

Murakami:2015:PCM

[Mur15] Hidetoshi Murakami. Power comparison of multivariate Wilcoxon-type tests based on the Jurecková-Kalina’s ranks of distances. Communications in Statistics: Simulation and Computation, 44(8):2176–2194, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ma:2010:FDT

[MV10] Jielai Ma and Wim P. M. Vijverberg. Five diagnostic tests for unobserved cluster effects. Communications in Statistics: Simulation and Computation, 39(6):1212–1227, 2010. CODEN CSSCDB. ISSN 0361-0918.

Marella:2019:TIB

[MV19] Daniela Marella and Paola Vicard. Toward an integrated Bayesian network approach to measurement error detection and correction. Communications in Statistics: Simulation and Computation, 48(2):544–555, 2019. CODEN CSSCDB. ISSN 0361-0918.

Mahmoud:2010:EDE

[MW10] Mahmoud A. Mahmoud and William H. Woodall. An evaluation of the double exponentially weighted moving average control chart. Communications in Statistics: Simulation and Computation, 39(5):933–949, 2010. CODEN CSSCDB. ISSN 0361-0918.

Miecznikowski:2010:BME

[MWH10] Jeffrey C. Miecznikowski, Dongliang Wang, and Alan Hutson. Bootstrap MISE estimators to obtain bandwidth for kernel density estimation. Communications in Statistics: Simulation and Computation, 39(7):1455–1469, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Men:2016:SBI

[MWK16] Zhongxian Men, Tony S. Wirjanto, and Adam W. Kolkiewicz. Sampling-based inference of time deformation models with heavy tail distributions. Communications in Statistics: Simulation and Computation, 45(9):3128–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Mao:2017:SMC

[MWL17] Guangcai Mao, Yi Wei, and Yanyan Liu. SIMEX method for censored quantile regression with measurement error. Communications in Statistics: Simulation and Computation, 46 (10):7552–7560, 2017. CODEN CSSCDB. ISSN 0361-0918.

Mojiri:2018:CPK

[MWSB18] A. Mojiri, Y. Waghei, H. R. Nili Sani, and G. R. Mohtashami Borzadaran. Comparison of predictions by kriging and spatial autoregressive models. Communications in Statistics: Simulation and Computation, 47(6):1785–1795, 2018. CO-DEN CSSCDB. ISSN 0361-0918.

Mu:2017:ACM

[MX17a] Weiyan Mu and Shifeng Xiong. On algorithmic construction of maximin distance designs. Communications in Statistics: Simulation and Computation, 46(10):7972–7985, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Mu:2017:RGC

[MX17b] Weiyan Mu and Shifeng Xiong. Robust generalized confidence intervals. Communications in Statistics: Simulation and Computation, 46(8):6049–6060, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Ma:2016:BAE

[MYS^{+}16] Junsheng Ma, Xiaoying Yu, Elaine Symanski, Rachelle Doody, and Wenyaw Chan. A Bayesian approach in estimating transition probabilities of a discrete-time Markov chain for ignorable intermittent missing data. Communications in Statistics: Simulation and Computation, 45(7):2598–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Manisera:2017:ENC

[MZ17] Marica Manisera and Paola Zuccolotto. Estimation of nonlinear CUB models via numerical optimization and EM algorithm. Communications in Statistics: Simulation and Computation, 46(7):5723–5739, 2017. CODEN CSSCDB. ISSN 0361-0918.

Noorossana:2012:PMU

[NA12] R. Noorossana and M. Ayoubi. Profile monitoring using nonparametric bootstrap T^{2} control chart. Communications in Statistics: Simulation and Computation, 41(3):302–315, 2012. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2013:GFT

[NA13] Hadi Alizadeh Noughabi and Naser Reza Arghami. Goodnessof-fit tests based on correcting moments of entropy estimators. Communications in Statistics: Simulation and Computation, 42(3):499–513, 2013. CODEN CSSCDB. ISSN 0361-0918.

Najarzadeh:2017:TEG

[Naj17] Dariush Najarzadeh. Testing equality of generalized variances of k multivariate normal populations. Communications in Statistics: Simulation and Computation, 46(8): 6414–6423, 2017. CODEN CSSCDB. ISSN 0361-0918.

Naderi:2018:MNM

[NAJ18] Mehrdad Naderi, Alireza Arabpour, and Ahad Jamalizadeh. Multivariate normal mean-variance mixture distribution based on Lindley distribution. Communications in Statistics: Simulation and Computation, 47(4):1179–1192, 2018. CODEN CSSCDB. ISSN 0361-0918.

Najarzadeh:2019:CIP

[Naj19] Dariush Najarzadeh. Confidence intervals for product of powers of the generalized variances of k multivariate normal populations. Communications in Statistics: Simulation and Computation, 48(4):1264–1276, 2019. CODEN CSSCDB. ISSN 0361-0918.

Najarian:2013:SSS

[NAK13] S. Najarian, M. Arashi, and B. M. Golam Kibria. A simulation study on some restricted ridge regression estimators. Communications in Statistics: Simulation and Computation, 42(4):871–890, 2013. CODEN CSSCDB. ISSN 0361-0918.

Namba:2017:BTS

[Nam17] Akio Namba. Bootstrap test for a structural break under possible heteroscedasticity. Communications in Statistics: Simulation and Computation, 46(5):4127–4139, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Nagatsuka:2015:EMP

[NB15] Hideki Nagatsuka and N. Balakrishnan. An efficient method of parameter and quantile estimation for the threeparameter Weibull distribution based on statistics invariant to unknown location parameter. Communications in Statistics: Simulation and Computation, 44(2):295–318, 2015. CODEN CSSCDB. ISSN 0361-0918.

Nezhad:2017:RAL

[NBOM17] Mohammad Saber Fallah Nezhad, Abolghasem Yousefi Babadi, Mohammad Saleh Owlia, and Ali Mostafaeipour. A recursive approach for lot sentencing problem in the presence of inspection errors. Communications in Statistics: Simulation and Computation, 46(3):2376–2392, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ng:2019:GGL

[NC19] Victoria K. Y. Ng and Robert A. Cribbie. The gamma generalized linear model, log transformation, and the robust Yuen–Welch test for analyzing group means with skewed and heteroscedastic data. Communications in Statistics: Simulation and Computation, 48(8):2269–2286, 2019. CODEN CSSCDB. ISSN 0361-0918.

Navarro:2015:AWE

[NCF15] Fabien Navarro, Christophe Chesneau, and Jalal Fadili. On adaptive wavelet estimation of a class of weighted densities. Communications in Statistics: Simulation and Computation, 44(8):2137–2150, 2015. CODEN CSSCDB. ISSN 0361-0918.

Norton:2018:SHH

[NCF18] Richard A. Norton, J. Andrés Christen, and Colin Fox. Sampling hyperparameters in hierarchical models: Improving on Gibbs for high-dimensional latent fields and large datasets. Communications in Statistics: Simulation and Computation, 47(9):2639–2655, 2018. CODEN CSSCDB. ISSN 0361-0918.

Nakamura:2010:DAG

[ND10] Tomohiro Nakamura and Hideyuki Douke. Development of adaptive group sequential procedure for changing sample size. Communications in Statistics: Simulation and Computation, 39(5):950–970, 2010. CODEN CSSCDB. ISSN 0361-0918.

Nie:2017:ICP

[ND17] Bin Nie and Mengying Du. Identifying change-point in polynomial profiles based on data-segmentation. Communications in Statistics: Simulation and Computation, 46(4): 2513–2528, 2017. CODEN CSSCDB. ISSN 0361-0918.

Nduka:2018:EBA

[Ndu18] Uchenna Chinedu Nduka. EM-based algorithms for autoregressive models with t-distributed innovations. Communications in Statistics: Simulation and Computation, 47(1): 206–228, 2018. CODEN CSSCDB. ISSN 0361-0918.

Neyens:2019:INL

[NFM19] Thomas Neyens, Christel Faes, and Geert Molenberghs. Integrated nested Laplace approximation for the analysis of count data via the combined model: a simulation study.

Communications in Statistics: Simulation and Computation, 48(3):819–836, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ng:2013:PSG

[Ng13] C. K. Ng. Procedures for selecting good exponential populations. Communications in Statistics: Simulation and Computation, 42(7):1681–1692, 2013. CODEN CSSCDB. ISSN 0361-0918.

Ng:2017:NPS

[Ng17] Allen C. K. Ng. New procedures for selecting the k-best exponential populations. Communications in Statistics: Simulation and Computation, 46(3):2176–2184, 2017. CODEN CSSCDB. ISSN 0361-0918.

Niaki:2014:PTG

[NGT14] Seyed Taghi Akhavan Niaki, Fazlollah Masoumi Gazaneh, and Moslem Toosheghanian. A parameter-tuned genetic algorithm for economic-statistical design of variable sampling interval X-bar control charts for non-normal correlated samples. Communications in Statistics: Simulation and Computation, 43(5):1212–1240, 2014. CODEN CSSCDB. ISSN 0361-0918.

Nguyen:2017:RMC

[Ngu17] Andrew L. Nguyen. Regenerative Markov chain importance sampling. Communications in Statistics: Simulation and Computation, 46(5):3892–3906, 2017. CODEN CSSCDB. ISSN 0361-0918.

Nomura:2018:EAC

[NH18] Takanobu Nomura and Satoshi Hattori. Estimation of the average causal effect via multiple propensity score stratification. Communications in Statistics: Simulation and Computation, 47(1):48–62, 2018. CODEN CSSCDB. ISSN 0361-0918.

Nandy:2010:ODB

[NHLL10] Karabi Nandy, Sami Helle, Antti Liski, and Erkki Liski. Optimal designs for binary logistic regression with a qualitative classifier with independent levels. Communications in Statistics: Simulation and Computation, 39(10):1962–1977, 2010. CODEN CSSCDB. ISSN 0361-0918.

Nicolau:2010:TDS

[Nic10] João Nicolau. Transition density and simulated likelihood estimation for time-inhomogeneous diffusions. Communications in Statistics: Simulation and Computation, 39(7): 1435–1454, 2010. CODEN CSSCDB. ISSN 0361-0918.

Nam:2017:MLD

[NK17] Jin Hyun Nam and Donguk Kim. Modified linear discriminant analysis using block covariance matrix in highdimensional data. Communications in Statistics: Simulation and Computation, 46(3):1796–1807, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Ng:2017:OEP

[NKKC17] Hon Keung Tony Ng, Ismail Kinaci, Coskun Kus, and Ping Shing Chan. Optimal experimental plan for multilevel stress testing with Weibull regression under progressive Type-II extremal censoring. Communications in Statistics: Simulation and Computation, 46(4):2611–2637, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Nadarajah:2017:EDM

[NL17] Saralees Nadarajah and Rui Li. Exact distribution of a modified Behrens–Fisher statistic. Communications in Statistics: Simulation and Computation, 46(9):6845–6864, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Najafi:2018:PAP

[NMS18] A. R. Najafi, Farshid Mehrdoust, and Shima Shirinpour. Pricing American put option on zero-coupon bond under fractional CIR model with transaction cost. Communications in Statistics: Simulation and Computation, 47(3):864– 870, 2018. CODEN CSSCDB. ISSN 0361-0918.

Noorossana:2015:PIM

[NN15] R. Noorossana and M. Nikoo. Phase II monitoring of geometric profiles. Communications in Statistics: Simulation and Computation, 44(4):1036–1049, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Navruz:2018:QEC

[NÖ18] Gözde Navruz and A. Firat Özdemir. Quantile estimation and comparing two independent groups with an approach based on percentile bootstrap. Communications in Statistics: Simulation and Computation, 47(7):2119–2138, 2018. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2017:EPE

[Nou17a] H. Alizadeh Noughabi. An extensive power evaluation of some tests for the inverse Gaussian distribution. Communications in Statistics: Simulation and Computation, 46(7): 5410–5422, 2017. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2017:EBT

[Nou17b] Hadi Alizadeh Noughabi. Entropy-based tests of uniformity: A Monte Carlo power comparison. Communications in Statistics: Simulation and Computation, 46(2):1266–1279, 2017. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2017:TEB

[Nou17c] Hadi Alizadeh Noughabi. Testing exponentiality based on Kullback–Leibler information for progressively Type II censored data. Communications in Statistics: Simulation and Computation, 46(10):7624–7638, 2017. CODEN CSSCDB. ISSN 0361-0918.

Noughabi:2017:TVL

[Nou17d] Hadi Alizadeh Noughabi. Testing the validity of the logistic model based on the empirical distribution function. Communications in Statistics: Simulation and Computation, 46 (7):5531–5540, 2017. CODEN CSSCDB. ISSN 0361-0918.

Nasir:2017:SBB

[NP17] Ehab Nasir and Rong Pan. Simulation-based Bayesian optimal design for multi-factor accelerated life tests. Communications in Statistics: Simulation and Computation, 46(2): 980–993, 2017. CODEN CSSCDB. ISSN 0361-0918.

Najafabadi:2015:GAE

[NQ15] Amir T. Payandeh Najafabadi and Marjan Qazvini. A GLM approach to estimating copula models. Communications in Statistics: Simulation and Computation, 44(6):1641–1656, 2015. CODEN CSSCDB. ISSN 0361-0918.

Nayeban:2013:BKB

[NRB13] S. Nayeban, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. Bhattacharyya and Kshirsagar bounds in generalized gamma distribution. Communications in Statistics: Simulation and Computation, 42(5):969–980, 2013. CODEN CSSCDB. ISSN 0361-0918.

Nezhad:2018:DEO

[NS18] Mohammad Saber Fallah Nezhad and Faeze Zahmatkesh Saredorahi. Designing an economically optimal repetitive group-sampling plan based on loss functions. Communications in Statistics: Simulation and Computation, 47(3): 783–799, 2018. CODEN CSSCDB. ISSN 0361-0918.

Nenes:2010:ECC

[NT10] George Nenes and George Tagaras. Evaluation of CUSUM charts for finite-horizon processes. Communications in Statistics: Simulation and Computation, 39(3):578–597, 2010. CODEN CSSCDB. ISSN 0361-0918.

Nezhad:2019:MPE

[NWB^{+}19] Emad Ashtari Nezhad, Yadollah Waghei, G. R. Mohtashami Borzadaran, H. R. Nilli Sani, and Hadi Alizadeh Noughabi. The modified permutation entropy-based independence test of time series. Communications in Statistics: Simulation and Computation, 48(10):2877–2897, 2019. CODEN CSSCDB. ISSN 0361-0918.

Neuhauser:2017:STC

[NWR17] Markus Neuhäuser, Anke Welz, and Graeme D. Ruxton. Statistical tests for the comparison of two samples: The general alternative. Communications in Statistics: Simulation and Computation, 46(2):903–909, 2017. CODEN CSSCDB. ISSN 0361-0918.

Nakamura:2016:SHT

[NYD16] Tomohiro Nakamura, Yoshiro Yamamoto, and Hideyuki Douke. Sequential hypothesis tests for identifying the minimum dose with a threshold effect. Communications in Statistics: Simulation and Computation, 45(6):1950–1970, 2016. CODEN CSSCDB. ISSN 0361-0918.

Nieto:2013:URJ

[NZL13] Fabio H. Nieto, Hanwen Zhang, and Wen Li. Using the reversible jump MCMC procedure for identifying and estimating univariate TAR models. Communications in Statistics:

Simulation and Computation, 42(4):814–840, 2013. CODEN CSSCDB. ISSN 0361-0918.

Ozdemir:2019:AAE

[ÖA19] Senay Özdemir and Olcay Arslan. An alternative algorithm of the empirical likelihood estimation for the parameter of a linear regression model. Communications in Statistics: Simulation and Computation, 48(7):1913–1921, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ozonur:2017:CTD

[OAB17] Deniz Ozonur, Hatice Tul Kubra Akdur, and Hulya Bayrak. Comparisons of tests of distributional assumption in Poisson regression model. Communications in Statistics: Simulation and Computation, 46(8):6197–6207, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ozel:2017:OLL

[OAC^{+}17] Gamze Ozel, Morad Alizadeh, Selen Cakmakyapan, G. G. Hamedani, Edwin M. M. Ortega, and Vicente G. Cancho. The odd log-logistic Lindley Poisson model for lifetime data. Communications in Statistics: Simulation and Computation, 46(8):6513–6537, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ossima:2016:UAC

[OBD16] Arnaud D. Nze Ossima, Mohamed C. Belkacemi, and Jean-Pierre Daurès. Using adjacent-category logits procedure for estimating receiver operating characteristic surface. Communications in Statistics: Simulation and Computation, 45 (3):902–919, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ortega:2011:LLR

[OCH11] Edwin M. M. Ortega, Gauss M. Cordeiro, and Elizabeth M. Hashimoto. A log-linear regression model for the beta-Weibull distribution. Communications in Statistics: Simulation and Computation, 40(8):1206–1235, 2011. CODEN CSSCDB. ISSN 0361-0918.

Ozturk:2016:EPV

[OD16] Omer Ozturk and Neslihan Demirel. Estimation of population variance from multi-ranker ranked set sampling designs. Communications in Statistics: Simulation and Computation, 45(10):3568–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ozdemir:2017:TSB

[ÖEG17] Yaprak Arzu Özdemir, Meral Ebegil, and Fikri Gökpinar. A test statistic based on ranked set sampling for two normal means. Communications in Statistics: Simulation and Computation, 46(10):8077–8085, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ortega:2014:CBM

[OG14] Francisco J. Ortega and Jose M. Gavilan. A comparison between maximum likelihood and Bayesian estimation of stochastic frontier production models. Communications in Statistics: Simulation and Computation, 43(7):1714–1725, 2014. CODEN CSSCDB. ISSN 0361-0918.

Oh:2015:URT

[OK15] Yujin Oh and Yu-Seop Kim. Unit root tests in the presence of multi-variance break and level shifts that have power against the piecewise stationary alternative. Communications in Statistics: Simulation and Computation, 44(6): 1465–1476, 2015. CODEN CSSCDB. ISSN 0361-0918.

Olkin:2016:LDB

[Olk16] Ingram Olkin. Life distributions: A brief discussion. Communications in Statistics: Simulation and Computation, 45 (5):1489–1498, 2016. CODEN CSSCDB. ISSN 0361-0918.

Orenti:2017:PCB

[OM17] Annalisa Orenti and Ettore Marubini. A performance counterexample of Billor–Chatterjee–Hadi procedure and an improvement proposal for robust regression. Communications in Statistics: Simulation and Computation, 46(5): 3980–3989, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ong:2013:BGN

[ON13] S. H. Ong and C. M. Ng. A bivariate generalization of the noncentral negative binomial distribution. Communications in Statistics: Simulation and Computation, 42(3):570–585, 2013. CODEN CSSCDB. ISSN 0361-0918.

Orbe:2013:CIR

[ONa13] Jesus Orbe and Vicente Núñez-antón. Confidence intervals on regression models with censored data. Communications in Statistics: Simulation and Computation, 42(9): 2140–2159, 2013. CODEN CSSCDB. ISSN 0361-0918.

Onder:2016:CSG

[Önd16] Hasan Önder. Comparative study of generalized estimating equations and logistic regressions on different sample sizes and correlation levels. Communications in Statistics: Simulation and Computation, 45(10):3528–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Onozawa:2016:TSP

[ONS16] Mizuki Onozawa, Takahiro Nishiyama, and Takashi Seo. On test statistics in profile analysis with high-dimensional data. Communications in Statistics: Simulation and Computation, 45(10):3716–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

ODriscoll:2011:GVM

[OR11] Diarmuid O’Driscoll and Donald E. Ramirez. Geometric view of measurement errors. Communications in Statistics: Simulation and Computation, 40(9):1373–1382, 2011. CO-DEN CSSCDB. ISSN 0361-0918.

Orzeszko:2017:NTS

[Orz17] Witold Orzeszko. Nonparametric testing for serial independence using the NRL statistic. Communications in Statistics: Simulation and Computation, 46(7):5151–5165, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ozkan:2017:MRS

[ÖSY17] Harun Özkan, Thanasis Stengos, and Ege Yazgan. Markov regime switching in mean and in fractional integration parameter. Communications in Statistics: Simulation and Computation, 46(9):6974–6981, 2017. CODEN CSSCDB. ISSN 0361-0918.

Otto:2017:NES

[Ott17] Philipp Otto. A note on efficient simulation of multidimensional spatial autoregressive processes. Communications in Statistics: Simulation and Computation, 46(6):4547–4558, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ourbih-Tari:2017:VRS

[OTZA17] Megdouda Ourbih-Tari, Arezki Zioui, and Abdelouhab Aloui. Variance reduction in the simulation of a stationary M/G/1 queuing system using refined descriptive sampling. Communications in Statistics: Simulation and Computation, 46(5):3504–3515, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ouysse:2013:FIB

[Ouy13] Rachida Ouysse. A fast iterated bootstrap procedure for approximating the small-sample bias. Communications in Statistics: Simulation and Computation, 42(7):1472–1494, 2013. CODEN CSSCDB. ISSN 0361-0918.

Ozdemir:2012:NRS

[ÖW12] A. Firat Özdemir and Rand Wilcox. New results on the small-sample properties of some robust univariate estimators of location. Communications in Statistics: Simulation and Computation, 41(9):1544–1556, 2012. CODEN CSSCDB. ISSN 0361-0918.

Okamura:2014:VBP

[OWD14] H. Okamura, R. Watanabe, and T. Dohi. Variational Bayes for phase-type distribution. Communications in Statistics: Simulation and Computation, 43(8):2031–2044, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Ozdemir:2013:CML

[OWY13] A. F. Ozdemir, Rand R. Wilcox, and E. Yildiztepe. Comparing measures of location: Some small-sample results when distributions differ in skewness and kurtosis under heterogeneity of variances. Communications in Statistics: Simulation and Computation, 42(2):407–424, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Ozdemir:2018:CIG

[OWY18] A. Firat Ozdemir, Rand Wilcox, and Engin Yildiztepe. Comparing J independent groups with a method based on trimmed means. Communications in Statistics: Simulation and Computation, 47(3):852–863, 2018. CODEN CSSCDB. ISSN 0361-0918.

Ogura:2016:PTT

[OY16] Toru Ogura and Takemi Yanagimoto. Powerful test of two proportions by assuming a registered prior density. Communications in Statistics: Simulation and Computation, 45(6): 1936–1949, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ozkale:2014:MCS

[Özk14] M. Revan Özkale. Monte Carlo simulation study of biased estimators in the linear regression models with correlated or heteroscedastic errors. Communications in Statistics: Simulation and Computation, 43(5):1143–1186, 2014. CODEN CSSCDB. ISSN 0361-0918.

Pakdaman:2018:PES

[PA18a] Zohreh Pakdaman and Jafar Ahmadi. Point estimation of the stress-strength reliability parameter for parallel system with independent and non-identical components. Communications in Statistics: Simulation and Computation, 47(4): 1193–1203, 2018. CODEN CSSCDB. ISSN 0361-0918.

Pereira:2018:CMS

[PA18b] Marcos Antonio Alves Pereira and Betsabé G. Blas Achic. Calibration model with skew scale mixtures of normal distributions. Communications in Statistics: Simulation and Computation, 47(5):1463–1478, 2018. CODEN CSSCDB. ISSN 0361-0918.

Palanisamy:2017:SDB

[Pal17] T. Palanisamy. Smoothing the difference-based estimates of variance using variational mode decomposition. Communications in Statistics: Simulation and Computation, 46(6): 4991–5001, 2017. CODEN CSSCDB. ISSN 0361-0918.

Pereira:2015:OFT

[PAM15] Dulce G. Pereira, Anabela Afonso, and Fátima Melo Medeiros. Overview of Friedman’s test and post-hoc analysis. Communications in Statistics: Simulation and Computation, 44(10):2636–2653, 2015. CODEN CSSCDB. ISSN 0361-0918.

Papastamoulis:2014:HLS

[Pap14] Panagiotis Papastamoulis. Handling the label switching problem in latent class models via the ECR algorithm. Communications in Statistics: Simulation and Computation, 43 (4):913–927, 2014. CODEN CSSCDB. ISSN 0361-0918.

Park:2011:NTP

[Par11] Hyo-Il Park. A nonparametric test procedure based on a group of quantile tests. Communications in Statistics: Simulation and Computation, 40(5):759–783, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Park:2015:NTP

[Par15] Hyo-Il Park. A nonparametric test procedure based on quantiles for the right censored data. Communications in Statistics: Simulation and Computation, 44(7):1735–1749, 2015. CODEN CSSCDB. ISSN 0361-0918.

Park:2017:GFT

[Par17] Sangun Park. On the goodness-of-fit test based on the ratio of cumulative hazard functions with the Type II censored data. Communications in Statistics: Simulation and Computation, 46(4):2935–2944, 2017. CODEN CSSCDB. ISSN 0361-0918.

Peerajit:2019:EAS

[PAS19] Wilasinee Peerajit, Yupaporn Areepong, and Saowanit Sukparungsee. Explicit analytical solutions for ARL of CUSUM chart for a long-memory SARFIMA model. Communications in Statistics: Simulation and Computation, 48(4): 1176–1190, 2019. CODEN CSSCDB. ISSN 0361-0918.

Pan:2010:MSS

[PB10] Zhengqiang Pan and N. Balakrishnan. Multiple-steps stepstress accelerated degradation modeling based on Wiener and gamma processes. Communications in Statistics: Simulation and Computation, 39(7):1384–1402, 2010. CODEN CSSCDB. ISSN 0361-0918.

Poblador:2016:CCM

[PB16] Ma. Sofia Criselda A. Poblador and Erniel B. Barrios. Control chart for monitoring autocorrelated process with multiple exogenous inputs. Communications in Statistics: Simulation and Computation, 45(9):3373–??, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Pannu:2017:RGL

[PB17a] Jasdeep Pannu and Nedret Billor. Robust group-lasso for functional regression model. Communications in Statistics: Simulation and Computation, 46(5):3356–3374, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Premaratne:2017:ATS

[PB17b] Gamini Premaratne and Anil K. Bera. Adjusting the tests for skewness and kurtosis for distributional misspecifications. Communications in Statistics: Simulation and Computation, 46(5):3599–3613, 2017. CODEN CSSCDB. ISSN 0361-0918.

Pal:2018:LIB

[PB18a] Suvra Pal and Narayanaswamy Balakrishnan. Likelihood inference based on EM algorithm for the destructive lengthbiased Poisson cure rate model with Weibull lifetime. Communications in Statistics: Simulation and Computation, 47 (3):644–660, 2018. CODEN CSSCDB. ISSN 0361-0918.

Palm:2018:BBI

[PB18b] Bruna Gregory Palm and Fábio M. Bayer. Bootstrap-based inferential improvements in beta autoregressive moving average model. Communications in Statistics: Simulation and Computation, 47(4):977–996, 2018. CODEN CSSCDB. ISSN 0361-0918.

Pan:2011:BCS

[PBS11] Zhengqiang Pan, N. Balakrishnan, and Quan Sun. Bivariate constant-stress accelerated degradation model and inference. Communications in Statistics: Simulation and Computation, 40(2):247–257, 2011. CODEN CSSCDB. ISSN 0361-0918.

Pereira:2013:ACA

[PC13] Luis Nobre Pereira and Pedro Simões Coelho. An applied comparison of area-level linear mixed models in small area estimation. Communications in Statistics: Simulation and Computation, 42(3):671–685, 2013. CODEN CSSCDB. ISSN 0361-0918.

Pereira:2014:DMM

[PCN14] Tarciana L. Pereira and Francisco Cribari-Neto. Detecting model misspecification in inflated beta regressions. Communications in Statistics: Simulation and Computation, 43(3): 631–656, 2014. CODEN CSSCDB. ISSN 0361-0918.

Pekmezci:2016:CPW

[PD16] Aytaç Pekmezci and Mustafa Dilek. The comparison of performances of widely used cointegration tests. Communications in Statistics: Simulation and Computation, 45(6): 2070–2080, 2016. CODEN CSSCDB. ISSN 0361-0918.

Pan:2019:SIG

[PDL19] Yingli Pan, Jieli Ding, and Yanyan Liu. Statistical inference for generalized case-cohort design under the proportional hazards model with parameter constraints. Communications in Statistics: Simulation and Computation, 48(8): 2467–2486, 2019. CODEN CSSCDB. ISSN 0361-0918.

Pereira:2019:QRB

[Per19] Gustavo H. A. Pereira. On quantile residuals in beta regression. Communications in Statistics: Simulation and Computation, 48(1):302–316, 2019. CODEN CSSCDB. ISSN 0361-0918.

Patricio:2017:CPN

[PFOC17] Miguel Patrício, Fábio Ferreira, Bárbara Oliveiros, and Francisco Caramelo. Comparing the performance of normality tests with ROC analysis and confidence intervals. Communications in Statistics: Simulation and Computation, 46 (10):7535–7551, 2017. CODEN CSSCDB. ISSN 0361-0918.

Peguin-Feissolle:2013:TGN

[PFST13] Anne Péguin-Feissolle, Birgit Strikholm, and Timo Teräsvirta. Testing the Granger noncausality hypothesis in stationary nonlinear models of unknown functional form. Communications in Statistics: Simulation and Computation, 42(5): 1063–1087, 2013. CODEN CSSCDB. ISSN 0361-0918.

Paredes-Garcia:2017:DEA

[PGCT17] Wilfrido J. Paredes-García and Eduardo Castaño-Tostado. Differential evolutionary algorithm in the construction process of optimal experimental designs. Communications in Statistics: Simulation and Computation, 46(10):7733–7743, 2017. CODEN CSSCDB. ISSN 0361-0918.

Palacios-Gonzalez:2014:MMB

[PGGF14] F. Palacios-González and R. M. García-Fernández. Mixtures of mixtures based on multiresolution analysis theory. Communications in Statistics: Simulation and Computation, 43 (4):723–742, 2014. CODEN CSSCDB. ISSN 0361-0918.

Payne:2018:EAD

[PGH^{+}18] Elizabeth H. Payne, Mulugeta Gebregziabher, James W. Hardin, Viswanathan Ramakrishnan, and Leonard E. Egede. An empirical approach to determine a threshold for assessing overdispersion in Poisson and negative binomial models for count data. Communications in Statistics: Simulation and Computation, 47(6):1722–1738, 2018. CODEN CSSCDB. ISSN 0361-0918.

Panteleeva:2015:ICE

[PGHA15] Olga Vladimirovna Panteleeva, Eduardo Gutiérrez González, Humberto Vaquera Huerta, and José A. Villaseñor Alva. Identifiability and comparison of estimation methods on Weibull mixture models. Communications in Statistics: Simulation and Computation, 44(7):1879–1900, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Pant:2017:CBT

[PH17] Mohan D. Pant and Todd C. Headrick. A characterization of the Burr Type III and Type XII distributions through the method of percentiles and the Spearman correlation. Communications in Statistics: Simulation and Computation, 46 (2):1611–1627, 2017. CODEN CSSCDB. ISSN 0361-0918.

Phillips:2012:CML

[Phi12] Robert F. Phillips. On computing maximum-likelihood estimates of the unbalanced two-way random-effects model. Communications in Statistics: Simulation and Computation, 41(10):1921–1927, 2012. CODEN CSSCDB. ISSN 0361-0918.

Picinbono:2013:APP

[Pic13] B. Picinbono. Algorithms for point processes analysis. Communications in Statistics: Simulation and Computation, 42 (1):90–106, 2013. CODEN CSSCDB. ISSN 0361-0918.

Picinbono:2014:SSD

[Pic14] B. Picinbono. Simulations of some doubly stochastic Poisson point processes. Communications in Statistics: Simulation and Computation, 43(7):1700–1713, 2014. CODEN CSSCDB. ISSN 0361-0918.

Picinbono:2016:IRD

[Pic16] Bernard Picinbono. Inversion of the renewal density with dead-time. Communications in Statistics: Simulation and Computation, 45(3):1083–1093, 2016. CODEN CSSCDB. ISSN 0361-0918.

Picchini:2019:LFS

[Pic19] Umberto Picchini. Likelihood-free stochastic approximation EM for inference in complex models. Communications in Statistics: Simulation and Computation, 48(3):861–881, 2019. CODEN CSSCDB. ISSN 0361-0918.

Pitselis:2014:REO

[Pit14] Georgios Pitselis. Robust eligible own funds and value at risk under solvency II system. Communications in Statistics: Simulation and Computation, 43(1):161–182, 2014. CODEN CSSCDB. ISSN 0361-0918.

Pedeli:2013:EBP

[PK13a] Xanthi Pedeli and Dimitris Karlis. On estimation of the bivariate Poisson INAR process. Communications in Statistics: Simulation and Computation, 42(3):514–533, 2013. CODEN CSSCDB. ISSN 0361-0918.

Pobiruchin:2013:SSC

[PK13b] Monika Pobiruchin and Meinhard Kieser. Sample size calculation and blinded sample size recalculation in clinical trials where the treatment effect is measured by the relative risk. Communications in Statistics: Simulation and Computation, 42(7):1643–1653, 2013. CODEN CSSCDB. ISSN 0361-0918.

Park:2016:RCD

[PK16] H. Park and S. Konishi. Robust coordinate descent algorithm robust solution path for high-dimensional sparse regression modeling. Communications in Statistics: Simulation and Computation, 45(1):115–129, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Phaladiganon:2011:BBM

[PKC^{+}11] Poovich Phaladiganon, Seoung Bum Kim, Victoria C. P. Chen, Jun-Geol Baek, and Sun-Kyoung Park. Bootstrapbased T^{2} multivariate control charts. Communications in Statistics: Simulation and Computation, 40(5):645–662, 2011. CODEN CSSCDB. ISSN 0361-0918.

Park:2012:RRU

[PKK12] Yousung Park, Daeyoung Kim, and Seongyong Kim. Robust regression using data partitioning and M-estimation. Communications in Statistics: Simulation and Computation, 41 (8):1282–1300, 2012. CODEN CSSCDB. ISSN 0361-0918.

Paik:2013:FRH

[PL13] Minhui Paik and Michael D. Larsen. Fractional regression hot deck imputation weight adjustment. Communications in Statistics: Simulation and Computation, 42(7):1514–1532, 2013. CODEN CSSCDB. ISSN 0361-0918.

Prass:2016:MBE

[PLA16] Taiane S. Prass, Sílvia R. C. Lopes, and Jorge A. Achcar. MCMC Bayesian estimation in FIEGARCH models. Communications in Statistics: Simulation and Computation, 45 (9):3238–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Pan:2016:NPJ

[PLLF16] Dong-Dong Pan, Zheng-Bang Li, Qi-Zhai Li, and Wing Kam Fung. A novel powerful joint analysis with data fusion in two-stage case-control genome-wide association studies. Communications in Statistics: Simulation and Computation, 45(7):2362–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Pan:2018:RJM

[PLT18] Jia-Chiun Pan, Mei-Hsien Lee, and Miao-Yu Tsai. Reversible jump Markov chain Monte Carlo algorithms for Bayesian variable selection in logistic mixed models. Communications in Statistics: Simulation and Computation, 47 (8):2234–2247, 2018. CODEN CSSCDB. ISSN 0361-0918.

Pan:2014:TSB

[PLW^{+}14] Haitao Pan, Chuanjuan Li, Ling Wang, Zhiwei Jiang, and Jielai Xia. A three-stage Bayesian adaptive Phase I/II design and simulation studies. Communications in Statistics: Simulation and Computation, 43(2):254–268, 2014. CODEN CSSCDB. ISSN 0361-0918.

Penghui:2015:RTS

[PLZ15] Guo Penghui, Liu Lihu, and Qian Zhengming. Robust test for spatial error model: Considering changes of spatial layouts and distribution misspecification. Communications in Statistics: Simulation and Computation, 44(2):402–416, 2015. CODEN CSSCDB. ISSN 0361-0918.

Pal:2012:ODO

[PM12] Manisha Pal and Nripes Kumar Mandal. D-optimum designs for optimum mixture in a quadratic log contrast model. Communications in Statistics: Simulation and Computation, 41(2):253–264, 2012. CODEN CSSCDB. ISSN 0361-0918.

Park:2016:PCG

[PM16] Taeyoung Park and Seunghyun Min. Partially collapsed Gibbs sampling for linear mixed-effects models. Communications in Statistics: Simulation and Computation, 45(1): 165–180, 2016. CODEN CSSCDB. ISSN 0361-0918.

Pal:2017:ODP

[PM17] Manisha Pal and Nripes Kumar Mandal. Optimum designs for parameter estimation in a mixture experiment with two correlated responses. Communications in Statistics: Simulation and Computation, 46(10):7698–7709, 2017. CODEN CSSCDB. ISSN 0361-0918.

Pal:2018:OMD

[PM18] Manisha Pal and N. K. Mandal. Optimum mixture designs for the log-logistic dose-response model with mixture of two similar compounds. Communications in Statistics: Simulation and Computation, 47(3):800–808, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Pasha:2018:GMM

[PMNF18] M. A. Pasha, M. Bameni Moghadam, N. Nematollahi, and S. Fani. A generalized model for multiplicity-cause economic and economic statistical design of X̄-control charts with proportional hazards shock model. Communications in Statistics: Simulation and Computation, 47(6):1866–1878, 2018. CODEN CSSCDB. ISSN 0361-0918.

Pakyari:2017:TGF

[PN17] Reza Pakyari and K. Resalati Nia. Testing goodness-of-fit for some lifetime distributions with conventional Type-I censoring. Communications in Statistics: Simulation and Computation, 46(4):2998–3009, 2017. CODEN CSSCDB. ISSN 0361-0918.

Prataviera:2019:HOL

[POCdSB19] Fábio Prataviera, Edwin M. M. Ortega, Gauss M. Cordeiro, and Altemir da Silva Braga. The heteroscedastic odd loglogistic generalized gamma regression model for censored data. Communications in Statistics: Simulation and Computation, 48(6):1815–1839, 2019. CODEN CSSCDB. ISSN 0361-0918.

Polajnar:2017:URC

[Pol17] Emil Polajnar. Using restricted CCA for cross-language information retrieval. Communications in Statistics: Simulation and Computation, 46(6):4618–4626, 2017. CODEN CSSCDB. ISSN 0361-0918.

Polajnar:2018:ULR

[Pol18] Emil Polajnar. Using lasso RCCA for cross-language information retrieval. Communications in Statistics: Simulation and Computation, 47(9):2739–2748, 2018. CODEN CSSCDB. ISSN 0361-0918.

Palanisamy:2014:EVF

[PR14] Thangavel Palanisamy and Joghee Ravichandran. Estimation of variance function in heteroscedastic regression models by generalized coiflets. Communications in Statistics: Simulation and Computation, 43(10):2213–2224, 2014. CODEN CSSCDB. ISSN 0361-0918.

Palanisamy:2015:VEH

[PR15a] T. Palanisamy and J. Ravichandran. Variance estimation in heteroscedastic models by undecimated Haar transform. Communications in Statistics: Simulation and Computation, 44(6):1532–1544, 2015. CODEN CSSCDB. ISSN 0361-0918.

Puccetti:2015:CSB

[PR15b] Giovanni Puccetti and Ludger Rüschendorf. Computation of sharp bounds on the expected value of a supermodular function of risks with given marginals. Communications in Statistics: Simulation and Computation, 44(3):705–718, 2015. CODEN CSSCDB. ISSN 0361-0918.

Procter:2019:IMQ

[PR19a] Marion Procter and Chris Robertson. Imputing missing quality of life data as covariate in survival analysis of the International Breast Cancer Study Group Trials VI and VII. Communications in Statistics: Simulation and Computation, 48(2):580–590, 2019. CODEN CSSCDB. ISSN 0361-0918.

Procter:2019:PSI

[PR19b] Marion Procter and Chris Robertson. Performance of standard imputation methods for missing quality of life data as covariate in survival analysis based on simulations from the International Breast Cancer Study Group Trials VI and VII*. Communications in Statistics: Simulation and Computation, 48(10):3063–3077, 2019. CODEN CSSCDB. ISSN 0361-0918.

Pradhan:2014:ICS

[PRA14] Biswabrata Pradhan, Soumya Roy, and Md. Zafar Anis. Inference for the component and system lifetime distribution of a k-unit parallel system based on system data. Communications in Statistics: Simulation and Computation, 43(1): 45–61, 2014. CODEN CSSCDB. ISSN 0361-0918.

Prajitha:2015:MDH

[Pra15] P. Prajitha. Modeling of discrete HNBUE class with hypothesis testing application. Communications in Statistics: Simulation and Computation, 44(7):1867–1878, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Pawar:2010:NST

[PS10] V. Y. Pawar and D. T. Shirke. A nonparametric Shewharttype synthetic control chart. Communications in Statistics: Simulation and Computation, 39(8):1493–1505, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Pesarin:2013:WCP

[PS13] F. Pesarin and L. Salmaso. On the weak consistency of permutation tests. Communications in Statistics: Simulation and Computation, 42(6):1368–1379, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Pan:2014:ODS

[PS14] Zhengqiang Pan and Quan Sun. Optimal design for stepstress accelerated degradation test with multiple performance characteristics based on gamma processes. Communications in Statistics: Simulation and Computation, 43(2): 298–314, 2014. CODEN CSSCDB. ISSN 0361-0918.

Pal:2018:EFP

[PS18a] Surya K. Pal and Housila P. Singh. Estimation of finite population mean using auxiliary information in presence of nonresponse. Communications in Statistics: Simulation and Computation, 47(1):143–165, 2018. CODEN CSSCDB. ISSN 0361-0918.

Pwasong:2018:FCU

[PS18b] Augustine Pwasong and Saratha Sathasivam. Forecasting comparisons using a hybrid ARFIMA and LRNN models. Communications in Statistics: Simulation and Computation, 47(8):2286–2303, 2018. CODEN CSSCDB. ISSN 0361-0918.

Pakhteev:2019:SNR

[PS19a] A. Pakhteev and A. Stepanov. On simulation of normal records. Communications in Statistics: Simulation and Computation, 48(8):2413–2424, 2019. CODEN CSSCDB. ISSN 0361-0918.

Pawar:2019:NTM

[PS19b] Somanath D. Pawar and Digambar T. Shirke. Nonparametric tests for multivariate locations based on data depth. Communications in Statistics: Simulation and Computation, 48(3):753–776, 2019. CODEN CSSCDB. ISSN 0361-0918.

Pervez:2017:ANB

[PSHP17] Zahid Pervez, Said Farooq Shah, Zawar Hussain, and Andreas N. Philippou. An application of negative binomial distribution of order k in sensitive surveys. Communications in Statistics: Simulation and Computation, 46(8):6654–6660, 2017. CODEN CSSCDB. ISSN 0361-0918.

Phang:2013:SAI

[PSO13] Y. N. Phang, S. Z. Sim, and S. H. Ong. Statistical analysis for the inverse trinomial distribution. Communications in Statistics: Simulation and Computation, 42(9):2073–2085, 2013. CODEN CSSCDB. ISSN 0361-0918.

Parmet:2010:FAR

[PSS10] Yisrael Parmet, Edna Schechtman, and Michael Sherman. Factor analysis revisited — how many factors are there? Communications in Statistics: Simulation and Computation, 39(10):1893–1908, 2010. CODEN CSSCDB. ISSN 0361-0918.

Perera:2016:GFT

[PSW16] A. A. P. N. M. Perera, M. R. Sooriyarachchi, and S. L. Wickramasuriya. A goodness of fit test for the multilevel logistic model. Communications in Statistics: Simulation and Computation, 45(2):643–659, 2016. CODEN CSSCDB. ISSN 0361-0918.

Panaretos:2019:IRO

[PTVP19] Dimitris Panaretos, George Tzavelas, Malvina Vamvakari, and Demosthenes Panagiotakos. Investigating the role of orthogonal and non-orthogonal rotation in multivariate factor analysis, in regard to the repeatability of the extracted factors: a simulation study. Communications in Statistics: Simulation and Computation, 48(7):2165–2176, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Pepler:2016:CSM

[PUN16] P. T. Pepler, D. W. Uys, and D. G. Nel. A comparison of some methods for the selection of a common eigenvector model for the covariance matrices of two groups. Communications in Statistics: Simulation and Computation, 45(8): 2917–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Pepler:2017:DAU

[PUN17] P. T. Pepler, D. W. Uys, and D. G. Nel. Discriminant analysis under the common principal components model. Communications in Statistics: Simulation and Computation, 46 (6):4812–4827, 2017. CODEN CSSCDB. ISSN 0361-0918.

Pepler:2018:RCM

[PUN18] P. T. Pepler, D. W. Uys, and D. G. Nel. Regularized covariance matrix estimation under the common principal components model. Communications in Statistics: Simulation and Computation, 47(3):631–643, 2018. CODEN CSSCDB. ISSN 0361-0918.

Paul:2017:BPN

[PV17] Chinmoy Paul and Gajendra K. Vishwakarma. Back propagation neural networks and multiple regressions in the case of heteroskedasticity. Communications in Statistics: Simulation and Computation, 46(9):6772–6789, 2017. CODEN CSSCDB. ISSN 0361-0918.

Pingel:2017:CEP

[PW17] Ronnie Pingel and Ingeborg Waernbaum. Correlation and efficiency of propensity score-based estimators for average causal effects. Communications in Statistics: Simulation and Computation, 46(5):3458–3478, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Pan:2011:RAA

[PWT11] Rui Pan, Hansheng Wang, and Chih-Ling Tsai. Regression analysis of asymmetric pairs in large-scale network data. Communications in Statistics: Simulation and Computation, 40(10):1540–1547, 2011. CODEN CSSCDB. ISSN 0361-0918.

Paige:2018:SSE

[PZ18] Robert L. Paige and Renren Zhao. Small sample equivalence tests for exponentiality. Communications in Statistics: Simulation and Computation, 47(6):1696–1703, 2018. CODEN CSSCDB. ISSN 0361-0918.

Qi:2018:MMV

[QDT18] Peiyan Qi, Xifa Duan, and Zheng Tian. MOSUM monitoring for variance change in nonparametric regression models. Communications in Statistics: Simulation and Computation, 47(9):2532–2546, 2018. CODEN CSSCDB. ISSN 0361-0918.

Qi:2016:SMC

[QDTL16] Peiyan Qi, Xifa Duan, Zheng Tian, and Fuxiao Li. Sequential monitoring for changes in models with a polynomial trend. Communications in Statistics: Simulation and Computation, 45(1):222–239, 2016. CODEN CSSCDB. ISSN 0361-0918.

Qaffou:2017:BPI

[QEK17] Abdelaziz Qaffou, Hamid El Maroufy, and Tewfik Kernane. Bayesian parameter inference for stochastic SIR epidemic model with hyperbolic diffusion. Communications in Statistics: Simulation and Computation, 46(9):6907–6922, 2017. CODEN CSSCDB. ISSN 0361-0918.

Qu:2019:TSE

[QG19] Hao Qu and Wei Gao. A two-stage estimation for panel data models with grouped fixed effects. Communications in Statistics: Simulation and Computation, 48(9):2539–2551, 2019. CODEN CSSCDB. ISSN 0361-0918.

Qihong:2014:PES

[QJ14] Duan Qihong and Liu Junrong. Parameter estimation for a k-out-of-n: F system. Communications in Statistics: Simulation and Computation, 43(1):99–114, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Qin:2017:ITI

[QLAJB17] Yangyang Qin, Kai Liu, Reem Al-Jarallah, and N. Balakrishnan. Inference for Type II bivariate Pólya–Aeppli distribution. Communications in Statistics: Simulation and Computation, 46(8):5981–5990, 2017. CODEN CSSCDB. ISSN 0361-0918.

Qin:2016:LRB

[QMLS16] Lei Qin, Shuangge Ma, Jung-Chen Lin, and Ben-Chang Shia. Lasso regression based on empirical mode decomposition. Communications in Statistics: Simulation and Computation, 45(4):1281–1294, 2016. CODEN CSSCDB. ISSN 0361-0918.

Qi:2017:TSV

[QQYZ17] Ai-Jun Qi, Zong-Feng Qi, Jian-Feng Yang, and Qiao-Zhen Zhang. A three-stage variable selection method for supersaturated designs. Communications in Statistics: Simulation and Computation, 46(4):2601–2610, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Qin:2016:MLO

[QWM^{+}16] Li Qin, Lisa Weissfeld, Marsha D. Marcus, Michele D. Levine, and Feng Dai. Modeling longitudinal obesity data with intermittent missingness using a new latent variable model. Communications in Statistics: Simulation and Computation, 45(6):2018–2031, 2016. CODEN CSSCDB. ISSN 0361-0918.

Qi:2019:RPL

[QZF19] Zong-Feng Qi, Yong-Dao Zhou, and Kai-Tai Fang. Representative points for location-biased datasets. Communications in Statistics: Simulation and Computation, 48(2):458–471, 2019. CODEN CSSCDB. ISSN 0361-0918.

Roozbeh:2014:FRE

[RA14] M. Roozbeh and M. Arashi. Feasible ridge estimator in seemingly unrelated semiparametric models. Communications in Statistics: Simulation and Computation, 43(10):2593–2613, 2014. CODEN CSSCDB. ISSN 0361-0918.

Roozbeh:2016:NRR

[RA16] Mahdi Roozbeh and Mohammad Arashi. New ridge regression estimator in semiparametric regression models. Communications in Statistics: Simulation and Computation, 45 (10):3683–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rao:2017:RLS

[RA17] G. Srinivasa Rao and Muhammad Aslam. Resubmitted lots with single sampling plans by attributes under the conditions of zero-inflated Poisson distribution. Communications in Statistics: Simulation and Computation, 46(3): 1814–1824, 2017. CODEN CSSCDB. ISSN 0361-0918.

Raqab:2018:DAW

[RAAK18] Mohammad Z. Raqab, Shafiqah A. Al-Awadhi, and Debasis Kundu. Discriminating among Weibull, log-normal, and loglogistic distributions. Communications in Statistics: Simulation and Computation, 47(5):1397–1419, 2018. CODEN CSSCDB. ISSN 0361-0918.

Rahman:2017:SAH

[Rah17] Azizur Rahman. Small area housing stress estimation in Australia: Calculating confidence intervals for a spatial microsimulation model. Communications in Statistics: Simulation and Computation, 46(9):7466–7484, 2017. CODEN CSSCDB. ISSN 0361-0918.

Raimova:2016:PAS

[Rai16] Gulnora Raimova. Probabilistic approach to solution of the Neumann problem for some nonlinear equation. Communications in Statistics: Simulation and Computation, 45(8): 2981–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rakitzis:2017:PMR

[Rak17] Athanasios C. Rakitzis. On the performance of modified runs rules charts with estimated parameters. Communications in Statistics: Simulation and Computation, 46(2):1360–1380, 2017. CODEN CSSCDB. ISSN 0361-0918.

Roozbeh:2010:SRR

[RAN10] M. Roozbeh, M. Arashi, and H. A. Niroumand. Semiparametric ridge regression approach in partially linear models. Communications in Statistics: Simulation and Computation, 39(3):449–460, 2010. CODEN CSSCDB. ISSN 0361-0918.

Rosychuk:2015:BAT

[RAN15] Rhonda J. Rosychuk, Christina Alloway, and Amanda S. Newton. A bootstrap approach for testing marginal independence between two categorical variables when subjects have repeated responses. Communications in Statistics: Simulation and Computation, 44(2):350–359, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Raschke:2013:PET

[Ras13] Mathias Raschke. Parameter estimation for the tail distribution of a random sequence. Communications in Statistics: Simulation and Computation, 42(5):1013–1043, 2013. CO-DEN CSSCDB. ISSN 0361-0918.

Raschke:2017:OMA

[Ras17] Mathias Raschke. Opportunities of the minimum Anderson– Darling estimator as a variant of the maximum likelihood method. Communications in Statistics: Simulation and Computation, 46(9):6879–6888, 2017. CODEN CSSCDB. ISSN 0361-0918.

Rutkowska:2016:CTL

[RB16] Agnieszka Rutkowska and Kazimierz Banasik. The Cucconi test for location-scale alternatives in application to asymmetric hydrological variables. Communications in Statistics: Simulation and Computation, 45(3):986–1000, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Resende:2017:PIO

[RBC17] Mariana Resende, Carla Regina Guimarães Brighenti, and Marcelo Ângelo Cirillo. Procedure to identify outliers through cumulative distribution of extremes in a Gamma response model. Communications in Statistics: Simulation and Computation, 46(9):6937–6946, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Reeve:2015:BMB

[RBX^{+}15] Russell Reeve, Seth Berry, Wei Xiao, Bradley Ferguson, Marcel Thürk, and Rüdiger Goetz. Benefits of model-based drug development: A rigorous, planned case study. Communications in Statistics: Simulation and Computation, 44(9): 2210–2222, 2015. CODEN CSSCDB. ISSN 0361-0918.

Rekabdar:2017:ADQ

[RCM17] G. Rekabdar, R. Chinipardaz, and B. Mansouri. On approximating distribution of the quadratic discriminant function. Communications in Statistics: Simulation and Computation, 46(5):3614–3626, 2017. CODEN CSSCDB. ISSN 0361-0918.

Rhein:2014:RMN

[RCR14] Beate Rhein, Tanja Clees, and Margot Ruschitzka. Robustness measures and numerical approximation of the cumulative density function of response surfaces. Communications in Statistics: Simulation and Computation, 43(1):1–17, 2014. CODEN CSSCDB. ISSN 0361-0918.

Rubinstein:2012:SMD

[RDV12] Reuven Rubinstein, Andrey Dolgin, and Radislav Vaisman. The splitting method for decision making. Communications in Statistics: Simulation and Computation, 41(6):905–921, 2012. CODEN CSSCDB. ISSN 0361-0918.

Reybod:2019:PMS

[REM19] Arman Reybod, Javad Etminan, and Adel Mohammadpour. A Pitman measure of similarity in k-means for clustering heavy-tailed data. Communications in Statistics: Simulation and Computation, 48(6):1595–1605, 2019. CODEN CSSCDB. ISSN 0361-0918.

Reschenhofer:2019:HRE

[Res19] E. Reschenhofer. Heteroscedasticity-robust estimation of autocorrelation. Communications in Statistics: Simulation and Computation, 48(4):1251–1263, 2019. CODEN CSS-CDB. ISSN 0361-0918.

RezendeDosSantos:2010:IHS

[RF10] Thiago Rezende Dos Santos and Glaura C. Franco. Inference for the hyperparameters of structural models under classical and Bayesian perspectives: A comparison study. Communications in Statistics: Simulation and Computation, 39(9): 1671–1689, 2010. CODEN CSSCDB. ISSN 0361-0918.

Ram:2018:BDS

[RG18] Mangey Ram and Nupur Goyal. Bi-directional system analysis under copula-coverage approach. Communications in Statistics: Simulation and Computation, 47(6):1831–1844, 2018. CODEN CSSCDB. ISSN 0361-0918.

Roy:2017:IBP

[RGP17] Soumya Roy, E. V. Gijo, and Biswabrata Pradhan. Inference based on progressive Type I interval censored data from log-normal distribution. Communications in Statistics: Simulation and Computation, 46(8):6495–6512, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ruwet:2011:ICT

[RH11] C. Ruwet and G. Haesbroeck. Impact of contamination on training and test error rates in statistical clustering. Communications in Statistics: Simulation and Computation, 40 (3):394–411, 2011. CODEN CSSCDB. ISSN 0361-0918.

Roozbeh:2017:FRE

[RH17] Mahdi Roozbeh and Nor Aishah Hamzah. Feasible robust estimator in restricted semiparametric regression models based on the LTS approach. Communications in Statistics: Simulation and Computation, 46(9):7332–7350, 2017. CODEN CSSCDB. ISSN 0361-0918.

Riffi:2019:EEG

[Rif19] Mohamed I. Riffi. Explicit expressions for gamma spacings and their moments. Communications in Statistics: Simulation and Computation, 48(2):621–626, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Rivaz:2017:MCA

[RK17] Firoozeh Rivaz and Majid Jafari Khaledi. Monte Carlo approximation of likelihood function in spatial GLMMs through an empirical Bayes method. Communications in Statistics: Simulation and Computation, 46(2):1322–1335, 2017. CODEN CSSCDB. ISSN 0361-0918.

Rodrigues:2016:P

[RL16] Paulo C. Rodrigues and Eric Laber. Preface. Communications in Statistics: Simulation and Computation, 45(7): 2657–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rong:2017:OND

[RL17] Jian-Ying Rong and Xu-Qing Liu. Optimal non-diagonal-type estimators in linear regression under the prediction error sum of squares criterion. Communications in Statistics: Simulation and Computation, 46(3):2418–2429, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Rocha:2011:BCA

[RLF11] Gustavo H. M. A. Rocha, Rosangela H. Loschi, and Glaura C. Franco. Bayesian and classical approaches for hypotheses testing in trisomies. Communications in Statistics: Simulation and Computation, 40(3):307–321, 2011. CODEN CSSCDB. ISSN 0361-0918.

Rahman:2016:EPT

[RLTJ16] Mohammad A. Rahman, Karl K. Lin, Ram C. Tiwari, and Matthew Jackson. Exact poly-K test. Communications in Statistics: Simulation and Computation, 45(7):2257–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ram:2016:PEM

[RM16a] Mangey Ram and Monika Manglik. Performance evaluation of a multi-state system covering imperfect fault coverage. Communications in Statistics: Simulation and Computation, 45(9):3259–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Riazoshams:2016:PRM

[RM16b] Hossein Riazoshams and Habshah Bt. Midi. The performance of a robust multistage estimator in nonlinear regression with heteroscedastic errors. Communications in Statistics: Simulation and Computation, 45(9):3394–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Roli:2016:MBA

[RM16c] Giulia Roli and Paola Monari. A model-based approach to measure school achievements in latent groups of students: A simulation study. Communications in Statistics: Simulation and Computation, 45(5):1655–1663, 2016. CODEN CSSCDB. ISSN 0361-0918.

Richter:2017:PTS

[RM17] Scott J. Richter and Melinda H. McCann. Permutation tests of scale using deviances. Communications in Statistics: Simulation and Computation, 46(7):5553–5565, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ramos:2015:SPE

[RMM15] Luís P. Ramos, Pedro P. Mota, and João T. Mexia. Sample partitioning estimation for ergodic diffusions. Communications in Statistics: Simulation and Computation, 44(1): 105–117, 2015. CODEN CSSCDB. ISSN 0361-0918.

Riazoshams:2010:PRT

[RMS10] Hossein Riazoshams, Habshah Bt. Midi, and Olimjon Sh. Sharipov. The performance of robust two-stage estimator in nonlinear regression with autocorrelated error. Communications in Statistics: Simulation and Computation, 39(6): 1251–1268, 2010. CODEN CSSCDB. ISSN 0361-0918.

Ranganai:2017:PLS

[RN17] Edmore Ranganai and Saralees Nadarajah. A predictive leverage statistic for quantile regression with measurement errors. Communications in Statistics: Simulation and Computation, 46(8):6385–6398, 2017. CODEN CSSCDB. ISSN 0361-0918.

Roozegar:2016:PIN

[RNJ16] Roohollah Roozegar, Alireza Nematollahi, and Ahad Jamalizadeh. Properties and inference for a new class of skew-t distributions. Communications in Statistics: Simulation and Computation, 45(9):3217–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ramires:2018:NRM

[ROC^{+}18] Thiago G. Ramires, Edwin M. M. Ortega, Gauss M. Cordeiro, Gilberto A. Paula, and Niel Hens. New regression model with four regression structures and computational aspects. Communications in Statistics: Simulation and Computation, 47(7):1940–1962, 2018. CODEN CSSCDB. ISSN 0361-0918.

Rodriguez:2011:LLI

[Rod11] Abel Rodriguez. On-line learning for the infinite hidden Markov model. Communications in Statistics: Simulation and Computation, 40(6):879–893, 2011. CODEN CSSCDB. ISSN 0361-0918.

Rodriguez:2016:CNA

[Rod16] Gabriel Rodríguez. A comparative note about estimation of the fractional parameter under additive outliers. Communications in Statistics: Simulation and Computation, 45(1): 207–221, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rong:2010:ALT

[Ron10] Jian-Ying Rong. Adjustive Liu-type estimators in linear regression models. Communications in Statistics: Simulation and Computation, 39(6):1162–1173, 2010. CODEN CSS-CDB. ISSN 0361-0918.

Roozegar:2017:RWA

[Roo17] Rasool Roozegar. Randomly weighted average with Dirichlet component proportions. Communications in Statistics: Simulation and Computation, 46(3):1808–1813, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Rosenthal:2018:SSS

[Ros18] Jeffrey S. Rosenthal. Stochastic simulation of sequential game-theory voting models. Communications in Statistics: Simulation and Computation, 47(7):2040–2054, 2018. CO-DEN CSSCDB. ISSN 0361-0918.

Ross:2019:IDR

[Ros19] Joseph Ross. Incremental-decremental rolling window double exponential smoothing. Communications in Statistics: Simulation and Computation, 48(10):3078–3091, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Roth:2012:SDD

[Rot12] Katrin Roth. Sequential designs for dose escalation studies in oncology. Communications in Statistics: Simulation and Computation, 41(7):1131–1141, 2012. CODEN CSSCDB. ISSN 0361-0918.

Roy:2012:ARM

[Roy12] Surupa Roy. Accounting for response misclassification and covariate measurement error using a random effects logit model. Communications in Statistics: Simulation and Computation, 41(9):1623–1636, 2012. CODEN CSSCDB. ISSN 0361-0918.

Rodriguez-Picon:2018:DMB

[RPRPMG^{+}18] Luis Alberto Rodríguez-Picón, Anna Patricia Rodríguez-Picón, Luis Carlos Méndez-González, Manuel I. Rodríguez-Borbón, and Alejandro Alvarado-Iniesta. Degradation modeling based on gamma process models with random effects. Communications in Statistics: Simulation and Computation, 47(6):1796–1810, 2018. CODEN CSSCDB. ISSN 0361-0918.

Rohan:2012:AVM

[RR12] Neelabh Rohan and T. V. Ramanathan. Asymmetric volatility models with structural breaks. Communications in Statistics: Simulation and Computation, 41(9):1519–1543, 2012. CODEN CSSCDB. ISSN 0361-0918.

Roy:2013:DAT

[RR13] Dilip Roy and Rishideep Roy. Distribution of the activity time in network analysis: A critical revisit with a gamma alternative. Communications in Statistics: Simulation and Computation, 42(6):1288–1297, 2013. CODEN CSSCDB. ISSN 0361-0918.

Raji:2019:RDC

[RRA19a] Ishaq Adeyanju Raji, Muhammad Riaz, and Nasir Abbas. Robust dual-CUSUM control charts for contaminated processes. Communications in Statistics: Simulation and Computation, 48(7):2177–2190, 2019. CODEN CSSCDB. ISSN 0361-0918.

Rasheed:2019:CBR

[RRA^{+}19b] H. M. Kashif Rasheed, Mujahid Rasul, Rashid Ahmed, Maria Batool, M. H. Tahir, and Farrukh Shehzad. Circular balanced repeated measurement designs in periods of three different sizes. Communications in Statistics: Simulation and Computation, 48(10):3022–3030, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Rattihalli:2019:SCB

[RRR19] R. N. Rattihalli, K. S. Madhava Rao, and M. Raghunath. Some count-based nonparametric tests for circular symmetry of a bivariate distribution. Communications in Statistics: Simulation and Computation, 48(3):922–943, 2019. CODEN CSSCDB. ISSN 0361-0918.

Rezaei:2018:ISS

[RSBZ18] Amir Rezaei, Maryam Sharafi, Javad Behboodian, and Atefeh Zamani. Inferences on stress-strength parameter based on GLD5 distribution. Communications in Statistics: Simulation and Computation, 47(5):1251–1263, 2018. CODEN CSSCDB. ISSN 0361-0918.

Raman:2016:CAB

[RSC16] Rohan Kumar Raman, U. C. Sud, and Hukum Chandra. Calibration approach-based product estimator of finite population total with subsampling of nonrespondents under single and two-phase sampling. Communications in Statistics: Simulation and Computation, 45(8):2965–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rao:2017:LAM

[RSR17] A. Vasudeva Rao, P. Sitaramacharyulu, and M. Chenchu Ramaiah. Linear approximate ML estimation in scaled Type-I generalized logistic distribution based on Type-II censored samples. Communications in Statistics: Simulation and Computation, 46(3):1682–1702, 2017. CODEN CSSCDB. ISSN 0361-0918.

Rathgeber:2017:FGH

[RSS17] A. W. Rathgeber, J. Stadler, and S. Stöckl. Fitting generalized hyperbolic processes — new insights for generating initial values. Communications in Statistics: Simulation and Computation, 46(7):5752–5762, 2017. CODEN CSSCDB. ISSN 0361-0918.

Rasch:2012:MES

[RSW12] D. Rasch, B. Spangl, and M. Wang. Minimal experimental size in the three way ANOVA cross classification model with approximate F -tests. Communications in Statistics: Simulation and Computation, 41(7):1120–1130, 2012. CODEN CSSCDB. ISSN 0361-0918.

Ramakrishnan:2012:PAF

[RT12] Viswanathan Ramakrishnan and Leroy R. Thacker. Population attributable fraction as a measure of heritability in dichotomous twin data. Communications in Statistics: Simulation and Computation, 41(3):405–418, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Roozegar:2017:MGD

[RTJ17] Rasool Roozegar, Saeid Tahmasebi, and Ali Akbar Jafari. The McDonald Gompertz distribution: Properties and applications. Communications in Statistics: Simulation and Computation, 46(5):3341–3355, 2017. CODEN CSSCDB. ISSN 0361-0918.

Ramdhani:2018:UST

[RTKB18] Satesh Ramdhani, Ram Tripathi, Jerome Keating, and Narayanaswamy Balakrishnan. Underground storage tanks (UST): a closer investigation statistical implications to changing the shape of a UST. Communications in Statistics: Simulation and Computation, 47(9):2612–2623, 2018. CODEN CSSCDB. ISSN 0361-0918.

Rasch:2016:CAB

[RTV16] D. Rasch, F. Teuscher, and L. R. Verdooren. A conjecture about BIBDs. Communications in Statistics: Simulation and Computation, 45(5):1526–1537, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rutkowska:2015:PCS

[Rut15] Agnieszka Rutkowska. Properties of the Cox-Stuart test for trend in application to hydrological series: The simulation study. Communications in Statistics: Simulation and Computation, 44(3):565–579, 2015. CODEN CSSCDB. ISSN 0361-0918.

Remenyi:2015:WSD

[RV15] Norbert Reményi and Brani Vidakovic. Wavelet shrinkage with double Weibull prior. Communications in Statistics: Simulation and Computation, 44(1):88–104, 2015. CODEN CSSCDB. ISSN 0361-0918.

Ruli:2016:HOB

[RV16] Erlis Ruli and Laura Ventura. Higher-order Bayesian approximations for pseudo-posterior distributions. Communications in Statistics: Simulation and Computation, 45(8): 2863–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rigdon:2017:UPD

[RW17] Steven E. Rigdon and William H. Woodall. Using the predictive distribution to determine control limits for the Bayesian MEWMA chart. Communications in Statistics: Simulation and Computation, 46(10):7818–7826, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Rakhmawati:2016:PCA

[RYW16] Dwi Yuli Rakhmawati, Chao-Lung Yang, and Chien-Wei Wu. Process capability assessment for asymmetric tolerances with consideration of gauge measurement errors. Communications in Statistics: Simulation and Computation, 45(2): 519–547, 2016. CODEN CSSCDB. ISSN 0361-0918.

Rahardja:2010:BIO

[RZZ10] Dewi Rahardja, Yan D. Zhao, and Hao Helen Zhang. Bayesian inference of odds ratios in misclassified binary data with a validation substudy. Communications in Statistics: Simulation and Computation, 39(10):1845–1854, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Shemehsavar:2016:FIO

[SA16] S. Shemehsavar and Morteza Amini. Failure inference and optimization for step stress model based on bivariate Wiener model. Communications in Statistics: Simulation and Computation, 45(1):130–151, 2016. CODEN CSSCDB. ISSN 0361-0918.

Siddiqui:2017:ASC

[SA17a] Fatima Siddiqui and Qazi M. Ali. An analytical study of the classification of highly skewed data. Communications in Statistics: Simulation and Computation, 46(10):7582–7601, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sogandi:2017:MCP

[SA17b] Fatemeh Sogandi and Amirhossein Amiri. Monotonic change point estimation of generalized linear model-based regression profiles. Communications in Statistics: Simulation and Computation, 46(3):2207–2227, 2017. CODEN CSSCDB. ISSN 0361-0918.

Saghir:2019:ECC

[SAAJ19] Aamir Saghir, Liaquat Ahmad, Muhammad Aslam, and Chi-Hyuck Jun. A EWMA control chart based on an auxiliary variable and repetitive sampling for monitoring process location. Communications in Statistics: Simulation and Computation, 48(7):2034–2045, 2019. CODEN CSSCDB. ISSN 0361-0918.

Sabzevari:2018:VAP

[Sab18] M. Sabzevari. Variance of the asymmetric n -player gambler’s ruin problem with ties allowed. Communications in Statistics: Simulation and Computation, 47(5):1540–1549, 2018. CODEN CSSCDB. ISSN 0361-0918.

Safariyan:2019:IES

[SAB19] A. Safariyan, M. Arashi, and R. Arabi Belaghi. Improved estimators for stress-strength reliability using record ranked set sampling scheme. Communications in Statistics: Simulation and Computation, 48(9):2708–2726, 2019. CODEN CSSCDB. ISSN 0361-0918.

Stephenson:2017:SAA

[SAD17] Matthew Stephenson, R. Ayesha Ali, and Gerarda A. Darlington. A simple approach to analyzing clustered longitudinal data. Communications in Statistics: Simulation and Computation, 46(5):3553–3562, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sadooghi-alvandi:2016:CSR

[SaJMF16] S. M. Sadooghi-alvandi, A. A. Jafari, and H. A. Mardani-Fard. Comparing several regression models with unequal variances. Communications in Statistics: Simulation and Computation, 45(9):3190–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Sangnawakij:2017:CIC

[SaN17] Patarawan Sangnawakij and Sa aat Niwitpong. Confidence intervals for coefficients of variation in two-parameter exponential distributions. Communications in Statistics: Simulation and Computation, 46(8):6618–6630, 2017. CODEN CSSCDB. ISSN 0361-0918.

Satman:2012:GAB

[Sat12] M. Hakan Satman. A genetic algorithm based modification on the LTS algorithm for large data sets. Communications in Statistics: Simulation and Computation, 41(5):644–652, 2012. CODEN CSSCDB. ISSN 0361-0918.

Santos:2012:NDT

[SB12a] Edward Santos and Erniel Barrios. Nonparametric decomposition of time series data with inputs. Communications in Statistics: Simulation and Computation, 41(9):1693–1710, 2012. CODEN CSSCDB. ISSN 0361-0918.

Shafay:2012:OTS

[SB12b] A. R. Shafay and N. Balakrishnan. One- and two-sample Bayesian prediction intervals based on Type-I hybrid censored data. Communications in Statistics: Simulation and Computation, 41(1):65–88, 2012. CODEN CSSCDB. ISSN 0361-0918.

Santos:2017:IPA

[SB17] Kevin Carl P. Santos and Erniel B. Barrios. Improving predictive accuracy of logistic regression model using ranked set samples. Communications in Statistics: Simulation and Computation, 46(1):78–90, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sadeghpour:2018:ETB

[SBH18] M. Sadeghpour, S. Baratpour, and A. Habibirad. Exponentiality test based on Renyi distance between equilibrium distributions. Communications in Statistics: Simulation and Computation, 47(10):2960–2967, 2018. CODEN CSSCDB. ISSN 0361-0918.

Said:2018:KSD

[SBNG18] Khamis K. Said, Doaa Basalamah, Wei Ning, and Arjun Gupta. The Kumaraswamy skew-t distribution and its related properties. Communications in Statistics: Simulation and Computation, 47(8):2409–2423, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Sharma:2017:EMD

[SBS17] Vikas Kumar Sharma, Hassan S. Bakouch, and Khushboo Suthar. An extended Maxwell distribution: Properties and applications. Communications in Statistics: Simulation and Computation, 46(9):6982–7007, 2017. CODEN CSSCDB. ISSN 0361-0918.

Saulo:2019:SSE

[SBZB19] Helton Saulo, Marcelo Bourguignon, Xiaojun Zhu, and N. Balakrishnan. Some simple estimators for the twoparameter gamma distribution. Communications in Statistics: Simulation and Computation, 48(8):2425–2437, 2019. CODEN CSSCDB. ISSN 0361-0918.

Stogiannis:2013:IFI

[SC13] D. Stogiannis and C. Caroni. Issues in fitting inverse Gaussian first hitting time regression models for lifetime data. Communications in Statistics: Simulation and Computation, 42(9):1948–1960, 2013. CODEN CSSCDB. ISSN 0361-0918.

Shan:2018:OIS

[SC18a] Guogen Shan and John J. Chen. Optimal inference for Simon’s two-stage design with over or under enrollment at the second stage. Communications in Statistics: Simulation and Computation, 47(4):1157–1167, 2018. CODEN CSSCDB. ISSN 0361-0918.

Su:2018:PSS

[SC18b] Pei-Fang Su and Sheng-Mao Chang. Power and sample size calculation for paired right-censored data based on survival copula models. Communications in Statistics: Simulation and Computation, 47(6):1565–1582, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Sun:2014:NAI

[SCH14] Bruce Qiang Sun, Xinfu Chen, and Ting Ting Huang. A new approach to importance sampling in Taylor’s stochastic volatility model. Communications in Statistics: Simulation and Computation, 43(3):580–596, 2014. CODEN CSSCDB. ISSN 0361-0918.

Santos:2017:HTL

[SCN17] Jéssica Santos and Francisco Cribari-Neto. Hypothesis testing in log-Birnbaum–Saunders regressions. Communications in Statistics: Simulation and Computation, 46(5): 3990–4003, 2017. CODEN CSSCDB. ISSN 0361-0918.

Shvartsman:2012:FMC

[SCpL12] Ilya Shvartsman, Vernon M. Chinchilli, and Hsiao pin Liu. A frailty model in crossover studies with time-to-event response variable. Communications in Statistics: Simulation and Computation, 41(10):1960–1969, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Schaarschmidt:2016:SCI

[SD16] Frank Schaarschmidt and Gemechis D. Djira. Simultaneous confidence intervals for ratios of fixed effect parameters in linear mixed models. Communications in Statistics: Simulation and Computation, 45(5):1704–1717, 2016. CODEN CSSCDB. ISSN 0361-0918.

Shukur:2017:CEF

[SDD17] Ghazi Shukur, Mariusz Doszyń, and Krzysztof Dmytrów. Comparison of the effectiveness of forecasts obtained by means of selected probability functions with respect to forecast error distributions. Communications in Statistics: Simulation and Computation, 46(5):3667–3679, 2017. CODEN CSSCDB. ISSN 0361-0918.

Staniswalis:2017:LPD

[SDS17] Joan G. Staniswalis, Christopher Dodoo, and Anu Sharma. Local principal differential analysis: Graphical methods for functional data with covariates. Communications in Statistics: Simulation and Computation, 46(3):2346–2359, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sharma:2018:LAB

[SDSM18] Vikas Kumar Sharma, Sanku Dey, Sanjay Kumar Singh, and Uzma Manzoor. On length and area-biased Maxwell distributions. Communications in Statistics: Simulation and Computation, 47(5):1506–1528, 2018. CODEN CSSCDB. ISSN 0361-0918.

Shan:2015:FSS

[SDZL15] Rui Shan, Haibo Dai, Jingyi Zhao, and Wen Liu. Forecasting study of Shanghai’s and Shenzhen’s stock markets using a hybrid forecast method. Communications in Statistics: Simulation and Computation, 44(4):1066–1077, 2015. CODEN CSSCDB. ISSN 0361-0918.

Sousa:2012:MLV

[SEG12] J. Beleza Sousa, M. L. Esquível, and R. M. Gaspar. Machine learning Vasicek model calibration with Gaussian processes. Communications in Statistics: Simulation and Computation, 41(6):776–786, 2012. CODEN CSSCDB. ISSN 0361-0918.

Sousa:2015:BBO

[SEG15] J. Beleza Sousa, M. L. Esquível, and R. M. Gaspar. Brownian bridge and other path-dependent Gaussian processes vectorial simulation. Communications in Statistics: Simulation and Computation, 44(10):2608–2621, 2015. CODEN CSSCDB. ISSN 0361-0918.

Seif:2019:NMC

[Sei19] Asghar Seif. A new Markov chain approach for the economic statistical design of the VSS T^{2} control chart. Communications in Statistics: Simulation and Computation, 48(1): 200–218, 2019. CODEN CSSCDB. ISSN 0361-0918.

Sen:2018:SUR

[Sen18] Amit Sen. A simple unit root testing methodology that does not require knowledge regarding the presence of a break. Communications in Statistics: Simulation and Computation, 47(3):871–889, 2018. CODEN CSSCDB. ISSN 0361-0918.

Su:2012:NMC

[SF12] Zheng Su and Liang Fang. A novel method to calculate mean survival time for time-to-event data. Communications in Statistics: Simulation and Computation, 41(5):611–620, 2012. CODEN CSSCDB. ISSN 0361-0918.

Santos:2019:MAO

[SF19] James D. Santos and Nelson L. Souza Filho. A Metropolis algorithm to obtain co-sufficient samples with applications in conditional tests. Communications in Statistics: Simulation and Computation, 48(9):2655–2659, 2019. CODEN CSSCDB. ISSN 0361-0918.

Santos:2019:CIB

[SFC19] T. R. Santos, G. C. Franco, and T. B. Ceccotti. Confidence intervals based on the deviance statistic for the hyperparameters in state space models. Communications in Statistics: Simulation and Computation, 48(4):998–1018, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Shahriari:2015:VSM

[SFG15] Shirin Shahriari, Susana Faria, and A. Manuela Gonçalves. Variable selection methods in high-dimensional regression — a simulation study. Communications in Statistics: Simulation and Computation, 44(10):2548–2561, 2015. CODEN CSSCDB. ISSN 0361-0918.

Salomon:2014:ACL

[SFLC14] Luis A. Salomón, Jean-Claude Fort, and Li-Vang Lozada-Chang. Average competitive learning vector quantization. Communications in Statistics: Simulation and Computation, 43(6):1288–1303, 2014. CODEN CSSCDB. ISSN 0361-0918.

Stehlik:2012:SSR

[SFS12] Milan Stehlík, Zdenek Fabián, and Lubos Strelec. Small sample robust testing for normality against Pareto tails. Communications in Statistics: Simulation and Computation, 41(7):1167–1194, 2012. CODEN CSSCDB. ISSN 0361-0918.

Song:2012:NLB

[SG12] Myung Soon Song and Leon J. Gleser. A numerical likelihood-based approach to combining correlation matrices. Communications in Statistics: Simulation and Computation, 41(9):1679–1692, 2012. CODEN CSSCDB. ISSN 0361-0918.

Shabbir:2014:IGD

[SG14] Javid Shabbir and Sat Gupta. An improved generalized difference-cum-ratio-type estimator for the population variance in two-phase sampling using two auxiliary variables.

Communications in Statistics: Simulation and Computation, 43(10):2540–2550, 2014. CODEN CSSCDB. ISSN 0361-0918.

Schubert:2015:CPM

[SG15a] Christine M. Schubert and Tobias Guennel. Comparing performance of multiclass classification systems with ROC manifolds: When volume and correct classification fails. Communications in Statistics: Simulation and Computation, 44 (3):719–738, 2015. CODEN CSSCDB. ISSN 0361-0918.

Sihm:2015:TSB

[SG15b] Jeong Sep Sihm and Sat Gupta. A two-stage binary optional randomized response model. Communications in Statistics: Simulation and Computation, 44(9):2278–2296, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Shabbir:2017:EFP

[SG17] Javid Shabbir and Sat Gupta. Estimation of finite population mean using two auxiliary variables in stratified twophase sampling. Communications in Statistics: Simulation and Computation, 46(2):1238–1256, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Sysoev:2016:BCI

[SGB16] Oleg Sysoev, Anders Grimvall, and Oleg Burdakov. Bootstrap confidence intervals for large-scale multivariate monotonic regression problems. Communications in Statistics: Simulation and Computation, 45(3):1025–1040, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Samanta:2017:ORB

[SGKM17] D. Samanta, A. Ganguly, D. Kundu, and S. Mitra. Order restricted Bayesian inference for exponential simple step-stress model. Communications in Statistics: Simulation and Computation, 46(2):1113–1135, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sturludottir:2017:DCM

[SGNS17] Erla Sturludottir, Helga Gunnlaugsdottir, Olafur K. Nielsen, and Gunnar Stefansson. Detection of a changepoint, a meanshift accompanied with a trend change, in short time-series with autocorrelation. Communications in Statistics: Simulation and Computation, 46(7):5808–5818, 2017. CODEN CSSCDB. ISSN 0361-0918.

Schiffl:2012:ODT

[SH12] Katharina Schiffl and Ralf-Dieter Hilgers. A-optimal designs for two-color microarray experiments for treatmentcontrol and all-to-next contrasts. Communications in Statistics: Simulation and Computation, 41(7):1142–1151, 2012. CODEN CSSCDB. ISSN 0361-0918.

Shang:2015:RTE

[Sha15] Han Lin Shang. Resampling techniques for estimating the distribution of descriptive statistics of functional data. Communications in Statistics: Simulation and Computation, 44 (3):614–635, 2015. CODEN CSSCDB. ISSN 0361-0918.

Shafay:2016:EIS

[Sha16] A. R. Shafay. Exact inference for a simple step-stress model with generalized Type-I hybrid censored data from the exponential distribution. Communications in Statistics: Simulation and Computation, 45(1):181–206, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Shen:2012:EBC

[She12] Pao-Sheng Shen. Estimation of the bivariate cause-specific distribution function with left-truncated competing risks data. Communications in Statistics: Simulation and Computation, 41(1):99–110, 2012. CODEN CSSCDB. ISSN 0361-0918.

Shen:2013:AME

[She13a] Pao-Sheng Shen. Additive mixed effect model for clustered doubly censored data. Communications in Statistics: Simulation and Computation, 42(10):2295–2307, 2013. CODEN CSSCDB. ISSN 0361-0918.

Shen:2013:REO

[She13b] Pao-Sheng Shen. Regression M-estimators with one modified form of doubly censored data. Communications in Statistics: Simulation and Computation, 42(3):603–612, 2013. CODEN CSSCDB. ISSN 0361-0918.

Shen:2014:AAR

[She14] Pao-Sheng Shen. Aalen’s additive risk model for lefttruncated and right-censored data. Communications in Statistics: Simulation and Computation, 43(5):1006–1019, 2014. CODEN CSSCDB. ISSN 0361-0918.

Shen:2015:ATM

[She15a] Pao-Sheng Shen. Additive transformation models for clustered doubly censored data. Communications in Statistics: Simulation and Computation, 44(8):1988–2002, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Shen:2015:PER

[She15b] Pao-Sheng Shen. Parameter estimation in regression for long-term survival rate from left-truncated and rightcensored data. Communications in Statistics: Simulation and Computation, 44(4):958–978, 2015. CODEN CSSCDB. ISSN 0361-0918.

Shen:2017:EJS

[She17a] Pao-Sheng Shen. Estimation of the joint survival function for successive duration times. Communications in Statistics: Simulation and Computation, 46(8):6672–6684, 2017. CODEN CSSCDB. ISSN 0361-0918.

Shen:2017:SAT

[She17b] Pao-Sheng Shen. Semiparametric analysis of transformation models with dependently left-truncated and right-censored data. Communications in Statistics: Simulation and Computation, 46(3):2474–2487, 2017. CODEN CSSCDB. ISSN 0361-0918.

Shen:2017:SRC

[She17c] Pao-Sheng Shen. A semiparametric regression cure model for interval-censored and left-truncated data. Communications in Statistics: Simulation and Computation, 46(5): 3690–3700, 2017. CODEN CSSCDB. ISSN 0361-0918.

Simino:2012:CPH

[SHM12] J. Simino, M. Hollander, and D. McGee. Calibration of proportional hazards and accelerated failure time models. Communications in Statistics: Simulation and Computation, 41 (6):922–941, 2012. CODEN CSSCDB. ISSN 0361-0918.

Strandberg:2013:NTD

[SI13] Alicia Graziosi Strandberg and Boris Iglewicz. A nonparametric test for deviation from randomness with applications to stock market index data. Communications in Statistics: Simulation and Computation, 42(3):686–697, 2013. CODEN CSSCDB. ISSN 0361-0918.

Strandberg:2014:DRR

[SI14] Alicia Graziosi Strandberg and Boris Iglewicz. Detecting randomness: A review of existing tests with new comparisons. Communications in Statistics: Simulation and Computation, 43(5):1187–1201, 2014. CODEN CSSCDB. ISSN 0361-0918.

Suliadi:2013:SEP

[SID13] Suliadi Suliadi, Noor Akma Ibrahim, and Isa Daud. Semiparametric estimation with profile algorithm for longitudinal binary data. Communications in Statistics: Simulation and Computation, 42(1):138–152, 2013. CODEN CSSCDB. ISSN 0361-0918.

Siman:2011:ECS

[Sim11] Miroslav Siman. On exact computation of some statistics based on projection pursuit in a general regression context. Communications in Statistics: Simulation and Computation, 40(6):948–956, 2011. CODEN CSSCDB. ISSN 0361-0918.

Sadikon:2019:NTD

[SIMS19] Nurul Hidayah Sadikon, Adriana Irawati Nur Ibrahim, Ibrahim Mohamed, and Kunio Shimizu. A new test of discordancy in cylindrical data. Communications in Statistics: Simulation and Computation, 48(8):2512–2522, 2019. CO-DEN CSSCDB. ISSN 0361-0918.

Singh:2014:CCV

[Sin14] Parminder Singh. Calculation of critical values for Haibing’s procedure for successive comparisons between ordered treatment effects. Communications in Statistics: Simulation and Computation, 43(1):151–160, 2014. CODEN CSSCDB. ISSN 0361-0918.

Samart:2018:EBC

[SJC18] Klairung Samart, Naratip Jansakul, and Mitchai Chongcheawchamnan. Exact bootstrap confidence intervals for regression coefficients in small samples. Communications in Statistics: Simulation and Computation, 47(10):2953–2959, 2018. CO-DEN CSSCDB. ISSN 0361-0918.

Schutzenmeister:2012:CNH

[SJP12] A. Schützenmeister, U. Jensen, and H.-P. Piepho. Checking normality and homoscedasticity in the general linear model using diagnostic plots. Communications in Statistics: Simulation and Computation, 41(2):141–154, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Spainhour:2019:SPG

[SJR19] John Christian G. Spainhour, Michael G. Janech, and Viswanathan Ramakrishnan. A study of the properties of Gaussian mixture model for stable isotope standard quantification in MALDI-TOF MS. Communications in Statistics: Simulation and Computation, 48(6):1637–1650, 2019. CODEN CSSCDB. ISSN 0361-0918.

Shen:2019:SAB

[SJT19] Xiaohui Shen, Long Jiang, and Dejian Tian. L^{p} solutions of anticipated BSDEs with weak monotonicity and general growth generators. Communications in Statistics: Simulation and Computation, 48(1):73–90, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Seo:2017:ITP

[SK17a] Jung-In Seo and Suk-Bok Kang. Inference for the twoparameter half-logistic distribution using pivotal quantities under progressively Type-II censoring schemes. Communications in Statistics: Simulation and Computation, 46(7): 5462–5478, 2017. CODEN CSSCDB. ISSN 0361-0918.

Seo:2017:RBEb

[SK17b] Jung In Seo and Yongku Kim. Robust Bayesian estimation of a two-parameter exponential distribution under generalized Type-I progressive hybrid censoring. Communications in Statistics: Simulation and Computation, 46(7):5795–5807, 2017. CODEN CSSCDB. ISSN 0361-0918.

Song:2017:BER

[SK17c] Joon Jin Song and Jong-Min Kim. Bayesian estimation of rare sensitive attribute. Communications in Statistics: Simulation and Computation, 46(5):4154–4160, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sukparungsee:2018:BCH

[SKBA18] Saowanit Sukparungsee, Sasigarn Kuvattana, Piyapatr Busababodhin, and Yupaporn Areepong. Bivariate copulas on the Hotelling’s T^{2} control chart. Communications in Statistics: Simulation and Computation, 47(2):413–419, 2018. CODEN CSSCDB. ISSN 0361-0918.

Srimaneekarn:2015:EDS

[SKHL15] N. Srimaneekarn, S. Kiatsupaibul, A. J. Hayter, and W. Liu. Estimating drug shelf-life with unknown lot-to-lot variability. Communications in Statistics: Simulation and Computation, 44(8):2195–2207, 2015. CODEN CSSCDB. ISSN 0361-0918.

Sunecher:2017:EPB

[SKJ17] Yuvraj Sunecher, Naushad Mamode Khan, and Vandna Jowaheer. Estimating the parameters of a BINMA Poisson model for a non-stationary bivariate time series. Communications in Statistics: Simulation and Computation, 46(9): 6803–6827, 2017. CODEN CSSCDB. ISSN 0361-0918.

Singh:2011:TMS

[SKJH11] Sarjinder Singh, Jong-Min Kim, Hakjin Jang, and Stephen Horn. Truncated Midzuno–Sen sampling schemes for estimating distribution functions. Communications in Statistics: Simulation and Computation, 40(7):1096–1110, 2011. CODEN CSSCDB. ISSN 0361-0918.

Seo:2017:RBEa

[SKK17] Jung In Seo, Suk Bok Kang, and Yongku Kim. Robust Bayesian estimation of a bathtub-shaped distribution under progressive Type-II censoring. Communications in Statistics: Simulation and Computation, 46(2):1008–1023, 2017. CODEN CSSCDB. ISSN 0361-0918.

Saha:2018:SSM

[SKLC18] Sajal Saha, Michael B. C. Khoo, Ming Ha Lee, and Philippe Castagliola. A side-sensitive modified group runs double sampling (SSMGRDS) control chart for detecting mean shifts. Communications in Statistics: Simulation and Computation, 47(5):1353–1369, 2018. CODEN CSSCDB. ISSN 0361-0918.

Son:2019:GAN

[SKLK19] W. Son, Y. Kim, J. Lim, and H.-C. Kuo. A general algorithm for non-parametric maximum likelihood estimator of stochastically ordered survival functions from case 2 intervalcensored data. Communications in Statistics: Simulation and Computation, 48(3):807–818, 2019. CODEN CSSCDB. ISSN 0361-0918.

Shutoh:2010:TEM

[SKM^{+}10] Nobumichi Shutoh, Makiko Kusumi, Wataru Morinaga, Shunichi Yamada, and Takashi Seo. Testing equality of mean vectors in two sample problem with missing data. Communications in Statistics: Simulation and Computation, 39(3): 487–500, 2010. CODEN CSSCDB. ISSN 0361-0918.

Skoulakis:2019:SPN

[Sko19] Georgios Skoulakis. Simulating from polynomial-normal distributions. Communications in Statistics: Simulation and Computation, 48(2):472–477, 2019. CODEN CSSCDB. ISSN 0361-0918.

Shin:2016:AIE

[SKS16] Dong Wan Shin, Hyo Jin Kim, and Jinwook Seo. SUR approach for IV estimation of canonical contagion models. Communications in Statistics: Simulation and Computation, 45(1):378–387, 2016. CODEN CSSCDB. ISSN 0361-0918.

Sarikavanij:2014:ELS

[SKST14] Sukuman Sarikavanij, Subramanyam Kasala, Bimal K. Sinha, and Montip Tiensuwan. Estimation of location and scale parameters in two-parameter exponential distribution based on ranked set sample. Communications in Statistics: Simulation and Computation, 43(1):132–141, 2014. CODEN CSSCDB. ISSN 0361-0918.

Seals:2016:EUS

[SKZA16] Samantha R. Seals, Charles R. Katholi, Jing Zhang, and Inmaculada B. Aban. Evaluating the use of spatial covariance structures in the analysis of cardiovascular imaging data. Communications in Statistics: Simulation and Computation, 45(9):3149–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Salari:2018:CDE

[SKZD18] Maryam Salari, Anoshirvan Kazemnejad, Farid Zayeri, and Ramin Daneshvar. Comparison of different estimation methods in growth curve models for categorical data: a simulation study. Communications in Statistics: Simulation and Computation, 47(6):1811–1830, 2018. CODEN CSSCDB. ISSN 0361-0918.

Saghir:2014:SRC

[SL14] Aamir Saghir and Zhengyan Lin. A study on the robustness of G-chart to non-normality. Communications in Statistics: Simulation and Computation, 43(10):2241–2251, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Saghir:2015:DGC

[SL15] Aamir Saghir and Zhengyan Lin. Designing of Gini-chart for exponential, t, logistic and Laplace distributions. Communications in Statistics: Simulation and Computation, 44 (9):2387–2409, 2015. CODEN CSSCDB. ISSN 0361-0918.

Su:2017:ACL

[SL17] Xiao Su and Sheng Luo. Analysis of censored longitudinal data with skewness and a terminal event. Communications in Statistics: Simulation and Computation, 46(7): 5378–5391, 2017. CODEN CSSCDB. ISSN 0361-0918.

Steel:2011:VSK

[SLB11] S. J. Steel, N. Louw, and S. Bierman. Variable selection for kernel classification. Communications in Statistics: Simulation and Computation, 40(2):229–246, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Song:2011:SNL

[SLLW11] Jie Song, Xiaoling Lu, Miao Liu, and Xizhi Wu. Stratified normalization LogitBoost for two-class unbalanced data classification. Communications in Statistics: Simulation and Computation, 40(10):1587–1593, 2011. CODEN CSSCDB. ISSN 0361-0918.

Strandberg:2014:EME

[SLX14] Erika Strandberg, Xinyi Lin, and Ronghui Xu. Estimation of main effect when covariates have non-proportional hazards. Communications in Statistics: Simulation and Computation, 43(7):1760–1770, 2014. CODEN CSSCDB. ISSN 0361-0918.

Samani:2010:PSE

[SM10] Afshin Samani and Pascal Madeleine. Permuted sample entropy. Communications in Statistics: Simulation and Computation, 39(8):1506–1516, 2010. CODEN CSSCDB. ISSN 0361-0918.

Saleh:2017:APS

[SM17] Nesma A. Saleh and Mahmoud A. Mahmoud. Accounting for phase I sampling variability in the performance of the MEWMA control chart with estimated parameters. Communications in Statistics: Simulation and Computation, 46 (6):4333–4347, 2017. CODEN CSSCDB. ISSN 0361-0918.

Smaga:2017:BMM

[Sma17] Lukasz Smaga. Bootstrap methods for multivariate hypothesis testing. Communications in Statistics: Simulation and Computation, 46(10):7654–7667, 2017. CODEN CSSCDB. ISSN 0361-0918.

Schimek:2012:IIA

[SMB12] Michael G. Schimek, Alena Mysicková, and Eva Budinská. An inference and integration approach for the consolidation of ranked lists. Communications in Statistics: Simulation and Computation, 41(7):1152–1166, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Santos:2015:KMU

[SMF15] R. Santos, J. P. Martins, and M. Felgueiras. Known mean, unknown maxima? Testing the maximum knowing only the mean. Communications in Statistics: Simulation and Computation, 44(10):2473–2491, 2015. CODEN CSSCDB. ISSN 0361-0918.

Salehi:2019:MTS

[SMMA19] Mohsen Salehi, Adel Mohammadpour, Mohammad Mohammadi, and Mina Aminghafari. A modified two-sample t -test based on permutation method for large-scale data. Communications in Statistics: Simulation and Computation, 48(2): 372–384, 2019. CODEN CSSCDB. ISSN 0361-0918.

Sjolander:2017:NNA

[SMS17a] Pär Sjölander, Kristofer Månsson, and Ghazi Shukur. A new nonlinear asymmetric cointegration approach using error correction models. Communications in Statistics: Simulation and Computation, 46(2):1661–1668, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sjolander:2017:TPC

[SMS17b] Pär Sjölander, Kristofer Månsson, and Ghazi Shukur. Testing for panel cointegration in an error-correction framework with an application to the Fisher hypothesis. Communications in Statistics: Simulation and Computation, 46(3): 1735–1745, 2017. CODEN CSSCDB. ISSN 0361-0918.

Shi:2014:NDT

[SMWR14] Jian Shi, Yixuan Meng, Shaoping Wang, and Yongmei Ruan. A novel double-time-scale evaluation method for Ethernet performability. Communications in Statistics: Simulation and Computation, 43(8):1913–1937, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Steane:2012:MBC

[SMY12] Michelle A. Steane, Paul D. McNicholas, and Rickey Y. Yada. Model-based classification via mixtures of multivariate t-factor analyzers. Communications in Statistics: Simulation and Computation, 41(4):510–523, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Salazar:2013:NTC

[SN13a] Yuri Salazar and Wing Lon Ng. Nonparametric tail copula estimation: An application to stock and volatility index returns. Communications in Statistics: Simulation and Computation, 42(3):613–635, 2013. CODEN CSSCDB. ISSN 0361-0918.

Simsek:2013:MCG

[SN13b] Gülhayat Gölbasi Simsek and Fatma Noyan. McDonald’s ω_{t}, Cronbach’s α, and generalized θ for composite reliability of common factors structures. Communications in Statistics: Simulation and Computation, 42(9):2008–2025, 2013. CODEN CSSCDB. ISSN 0361-0918.

Sankaran:2016:NQF

[SNM16] P. G. Sankaran, N. Unnikrishnan Nair, and N. N. Midhu. A new quantile function with applications to reliability analysis. Communications in Statistics: Simulation and Computation, 45(2):566–582, 2016. CODEN CSSCDB. ISSN 0361-0918.

Sharifdoust:2016:PGF

[SNO16] Maryam Sharifdoust, Choung Min Ng, and Seng-Huat Ong. Probability generating function based Jeffrey’s divergence for statistical inference. Communications in Statistics: Simulation and Computation, 45(7):2445–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Said:2017:LPT

[SNT17] Khamis K. Said, Wei Ning, and Yubin Tian. Likelihood procedure for testing changes in skew normal model with applications to stock returns. Communications in Statistics: Simulation and Computation, 46(9):6790–6802, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sim:2010:PED

[SO10] S. Z. Sim and S. H. Ong. Parameter estimation for discrete distributions by generalized Hellinger-type divergence based on probability generating function. Communications in Statistics: Simulation and Computation, 39(2):305–314, 2010. CODEN CSSCDB. ISSN 0361-0918.

Sayyareh:2010:ECB

[SOBH10] Abdolreza Sayyareh, Raouf Obeidi, and Avner Bar-Hen. Empirical comparison between some model selection criteria. Communications in Statistics: Simulation and Computation, 40(1):72–86, 2010. CODEN CSSCDB. ISSN 0361-0918.

Solberger:2016:LRT

[Sol16] Martin Solberger. A likelihood ratio test for idiosyncratic unit roots in the exact factor model with integrated factors. Communications in Statistics: Simulation and Computation, 45(6):2032–2050, 2016. CODEN CSSCDB. ISSN 0361-0918.

Solmi:2014:EPS

[SOSB14a] Francesca Solmi, Patrick Onghena, Luigi Salmaso, and Isis Bulté. Extensions of permutation solutions to test for treatment effects in replicated single-case alternation experiments with multivariate response. Communications in Statistics: Simulation and Computation, 43(5):1036–1051, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Solmi:2014:PST

[SOSB14b] Francesca Solmi, Patrick Onghena, Luigi Salmaso, and Isis Bulté. A permutation solution to test for treatment effects in alternation design single-case experiments. Communications in Statistics: Simulation and Computation, 43(5): 1094–1111, 2014. CODEN CSSCDB. ISSN 0361-0918.

Sezer:2017:CCI

[SOY17] Ahmet Sezer, Evren Ozkip, and Berna Yazici. Comparison of confidence intervals for the Behrens–Fisher problem. Communications in Statistics: Simulation and Computation, 46 (4):3242–3266, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sparks:2014:DMO

[SP14] Ross Sparks and Ellis Patrick. Detection of multiple outbreaks using spatio-temporal EWMA-ordered statistics. Communications in Statistics: Simulation and Computation, 43(10):2678–2701, 2014. CODEN CSSCDB. ISSN 0361-0918.

Saleh:2016:CEE

[SP16] Moein Saleh and Rong Pan. Constructing efficient experimental designs for generalized linear models. Communications in Statistics: Simulation and Computation, 45(8): 2827–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Simkova:2017:CMM

[SP17a] Tereza Simková and Jan Picek. A comparison of L-, LQ- , T L-moment and maximum likelihood high quantile estimates of the GPD and GEV distribution. Communications in Statistics: Simulation and Computation, 46(8): 5991–6010, 2017. CODEN CSSCDB. ISSN 0361-0918.

Singh:2017:EPM

[SP17b] Housila P. Singh and Surya K. Pal. Estimation of population mean in two-occasion successive sampling. Communications in Statistics: Simulation and Computation, 46(6): 4908–4937, 2017. CODEN CSSCDB. ISSN 0361-0918.

Sankaran:2018:ARR

[SP18a] P. G. Sankaran and S. Prasad. Additive risks regression model for middle censored exponentiated-exponential lifetime data. Communications in Statistics: Simulation and Computation, 47(7):1963–1974, 2018. CODEN CSSCDB. ISSN 0361-0918.

Sciandra:2018:GMS

[SP18b] M. Sciandra and A. Plaia. A graphical model selection tool for mixed models. Communications in Statistics: Simulation and Computation, 47(9):2624–2638, 2018. CODEN CSSCDB. ISSN 0361-0918.

Sparks:2012:SCO

[Spa12] Ross Sparks. Spatially clustered outbreak detection using the EWMA SCAN statistics with multiple sized Windows. Communications in Statistics: Simulation and Computation, 41(9):1637–1653, 2012. CODEN CSSCDB. ISSN 0361-0918.

Souza:2016:TII

[SPCNL16] Tatiene C. Souza, Tarciana L. Pereira, Francisco Cribari-Neto, and Veônica M. C. Lima. Testing inference in inflated beta regressions under model misspecification. Communications in Statistics: Simulation and Computation, 45(2): 625–642, 2016. CODEN CSSCDB. ISSN 0361-0918.

Shahzad:2019:NCR

[SPH19] Usman Shahzad, Pier Francesco Perri, and Muhammad Hanif. A new class of ratio-type estimators for improving mean estimation of nonsensitive and sensitive variables by using supplementary information. Communications in Statistics: Simulation and Computation, 48(9):2566–2585, 2019. CODEN CSSCDB. ISSN 0361-0918.

Soulakova:2012:PAC

[SR12] Julia N. Soulakova and Ananya Roy. On power approximations and comparison of several asymptotic tests to detect a specified difference between two proportions. Communications in Statistics: Simulation and Computation, 41(3): 376–390, 2012. CODEN CSSCDB. ISSN 0361-0918.

Sparks:2017:ISM

[SRP^{+}17] Ross S. Sparks, Bella Robinson, Robert Power, Mark Cameron, and Sam Woolford. An investigation into social media syndromic monitoring. Communications in Statistics: Simulation and Computation, 46(8):5901–5923, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Shutoh:2010:AED

[SS10] Nobumichi Shutoh and Takashi Seo. Asymptotic expansion of the distribution of the Studentized linear discriminant function based on two-step monotone missing samples. Communications in Statistics: Simulation and Computation, 39 (7):1365–1383, 2010. CODEN CSSCDB. ISSN 0361-0918.

Sengupta:2012:EMR

[SS12a] Raghu Nandan Sengupta and Sachin Srivastava. Estimation for the multiple regression setup using balanced loss function. Communications in Statistics: Simulation and Computation, 41(5):653–670, 2012. CODEN CSSCDB. ISSN 0361-0918.

Sengupta:2012:MRE

[SS12b] Raghu Nandan Sengupta and Sachin Srivastava. Minimum risk estimation of scalar means under convex combination of loss functions. Communications in Statistics: Simulation and Computation, 41(8):1346–1371, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Shen:2012:ELR

[sS12c] Pao sheng Shen. Empirical likelihood ratio for linear transformation models with doubly censored data. Communications in Statistics: Simulation and Computation, 41(4): 531–543, 2012. CODEN CSSCDB. ISSN 0361-0918.

Sedory:2014:EMU

[SS14] Stephen A. Sedory and Sarjinder Singh. Estimation of mode using auxiliary information. Communications in Statistics: Simulation and Computation, 43(10):2390–2402, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Sun:2016:EBS

[SS16] Tianyu Sun and Yimin Shi. Estimation for Birnbaum– Saunders distribution in simple step stress-accelerated life test with Type-II censoring. Communications in Statistics: Simulation and Computation, 45(3):880–901, 2016. CODEN CSSCDB. ISSN 0361-0918.

Sarabia:2017:BEI

[SS17a] José María Sarabia and Golnaz Shahtahmassebi. Bayesian estimation of incomplete data using conditionally specified priors. Communications in Statistics: Simulation and Computation, 46(5):3419–3435, 2017. CODEN CSSCDB. ISSN 0361-0918.

Slama:2017:BAC

[SS17b] Abdeldjalil Slama and Hafida Saggou. A Bayesian analysis of a change in the parameters of autoregressive time series. Communications in Statistics: Simulation and Computation, 46(9):7008–7021, 2017. CODEN CSSCDB. ISSN 0361-0918.

Srivastava:2017:TSI

[SS17c] Muni S. Srivastava and Martin Singull. Testing sphericity and intraclass covariance structures under a growth curve model in high dimension. Communications in Statistics: Simulation and Computation, 46(7):5740–5751, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Shen:2019:IEJ

[sS19a] Pao sheng Shen. The IPW estimator for the joint distribution function of the gap times from recurrent event data. Communications in Statistics: Simulation and Computation, 48(4):957–967, 2019. CODEN CSSCDB. ISSN 0361-0918.

Shin:2019:VEC

[SS19b] Ji Won Shin and Dong Wan Shin. Vector error correction heterogeneous autoregressive forecast model of realized volatility and implied volatility. Communications in Statistics: Simulation and Computation, 48(5):1503–1515, 2019. CODEN CSSCDB. ISSN 0361-0918.

Singh:2015:EFP

[SSA15] Sarjinder Singh, Stephen A. Sedory, and Raghunath Arnab. Estimation of finite population variance using scrambled responses in the presence of auxiliary information. Communications in Statistics: Simulation and Computation, 44(4): 1050–1065, 2015. CODEN CSSCDB. ISSN 0361-0918.

Shokrizadeh:2018:ODV

[SSA18] Reza Shokrizadeh, Abbas Saghaei, and Vahid Amirzadeh. Optimal design of the variable sampling size and sampling interval variable dimension T^{2} control chart for monitoring the mean vector of a multivariate normal process. Communications in Statistics: Simulation and Computation, 47(2): 329–337, 2018. CODEN CSSCDB. ISSN 0361-0918.

Singh:2016:BEP

[SSK16] Sanjay Kumar Singh, Umesh Singh, and Manoj Kumar. Bayesian estimation for Poisson-exponential model under progressive Type-II censoring data with binomial removal and its application to ovarian cancer data. Communications in Statistics: Simulation and Computation, 45(9):3457–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Singh:2014:ELE

[SSmK14] Sarjinder Singh, Stephen A. Sedory, and Jong min Kim. An empirical likelihood estimate of the finite population correlation coefficient. Communications in Statistics: Simulation and Computation, 43(6):1430–1441, 2014. CODEN CSSCDB. ISSN 0361-0918.

Saggou:2019:AUQ

[SSOTB19] Hafida Saggou, Ines Sadeg, Megdouda Ourbih-Tari, and El-Bay Bourennane. The analysis of unreliable M^{[X]}/G/1 queuing system with loss, vacation and two delays of verification. Communications in Statistics: Simulation and Computation, 48(5):1366–1381, 2019. CODEN CSSCDB. ISSN 0361-0918.

Sanso:2014:DRU

[SSS14] Andreu Sansó, Clara Simón, and Ana E. Sipols. A detrended range unit root (DRUR) test. Communications in Statistics: Simulation and Computation, 43(6):1253–1264, 2014. CO-DEN CSSCDB. ISSN 0361-0918.

Stock:2015:BIE

[SSY15] Eileen M. Stock, James D. Stamey, and Dean M. Young. Bayesian interval estimation for the difference in TPRs and FPRs of two diagnostic tests with unverified negatives. Communications in Statistics: Simulation and Computation, 44 (2):505–524, 2015. CODEN CSSCDB. ISSN 0361-0918.

Song:2018:DWN

[SSY18] Hyejin Song, Dong Wan Shin, and Jae Keun Yoo. Do we need the constant term in the heterogeneous autoregressive model for forecasting realized volatilities? Communications in Statistics: Simulation and Computation, 47(1): 63–73, 2018. CODEN CSSCDB. ISSN 0361-0918.

Shao:2019:PIV

[SSYS19] Hui Shao, Charles Stoecker, Shuang Yang, and Lizheng Shi. The pitfall of instrumental variables in big data: What the rule of thumb can’t give you. Communications in Statistics: Simulation and Computation, 48(7):2118–2124, 2019. CODEN CSSCDB. ISSN 0361-0918.

StLaurent:2013:EMP

[ST13] Roy St. Laurent and Philip Turk. The effects of misconceptions on the properties of Friedman’s test. Communications in Statistics: Simulation and Computation, 42(7): 1596–1615, 2013. CODEN CSSCDB. ISSN 0361-0918.

Shalabh:2017:RPM

[ST17] Shalabh and Jia-Ren Tsai. Ratio and product methods of estimation of population mean in the presence of correlated measurement errors. Communications in Statistics: Simulation and Computation, 46(7):5566–5593, 2017. CODEN CSSCDB. ISSN 0361-0918.

Stapleton:2012:ECW

[Sta12] Laura Stapleton. Evaluation of conditional weight approximations for two-level models. Communications in Statistics: Simulation and Computation, 41(2):182–204, 2012. CODEN CSSCDB. ISSN 0361-0918.

Staggs:2017:CNK

[Sta17] Vincent S. Staggs. Comparison of naïve, Kenward–Roger, and parametric bootstrap interval approaches to smallsample inference in linear mixed models. Communications in Statistics: Simulation and Computation, 46(3):1933–1943, 2017. CODEN CSSCDB. ISSN 0361-0918.

Steland:2012:SDA

[Ste12] Ansgar Steland. Sequential data-adaptive bandwidth selection by cross-validation for nonparametric prediction. Communications in Statistics: Simulation and Computation, 41 (7):1195–1219, 2012. CODEN CSSCDB. ISSN 0361-0918.

Stepanov:2019:SWR

[Ste19] A. Stepanov. On simulation of weak records. Communications in Statistics: Simulation and Computation, 48(3): 797–806, 2019. CODEN CSSCDB. ISSN 0361-0918.

Singh:2016:OSS

[STHS16] Warsha Singh, Gudmundur Thordarson, Sisira Haputhantri, and Gunnar Stefansson. Optimized sampling strategies for identifying modes in length-frequency distributions. Communications in Statistics: Simulation and Computation, 45 (8):2874–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Siray:2015:RLE

[STK15] Gülesen Üstündag Siray, Selma Toker, and Selahattin Kaçiranlar. On the restricted Liu estimator in the logistic regression model. Communications in Statistics: Simulation and Computation, 44(1):217–232, 2015. CODEN CSSCDB. ISSN 0361-0918.

Song:2017:DMV

[STLH17] Lianlian Song, Geoffrey Tso, Hing-Po Lo, and Zhongsheng Hua. A dynamic multiple-variety choice adaption model. Communications in Statistics: Simulation and Computation, 46(1):515–529, 2017. CODEN CSSCDB. ISSN 0361-0918.

Su:2011:OPD

[Su11a] Jen-Je Su. On the oversized problem of Dickey–Fuller-type tests with GARCH errors. Communications in Statistics: Simulation and Computation, 40(9):1364–1372, 2011. CO-DEN CSSCDB. ISSN 0361-0918.

Su:2011:NIR

[Su11b] Zheng Su. A note on importance resampling for multidimensional statistics. Communications in Statistics: Simulation and Computation, 40(8):1163–1170, 2011. CODEN CSSCDB. ISSN 0361-0918.

Sharma:2018:HBA

[SU18] Reema Sharma and S. K. Upadhyay. A hierarchical Bayes analysis for one-shot device testing experiment under the assumption of exponentiality. Communications in Statistics: Simulation and Computation, 47(5):1297–1314, 2018. CO-DEN CSSCDB. ISSN 0361-0918.

Sun:2016:ELI

[Sun16] Zhimeng Sun. Empirical likelihood inference of the partial linear isotonic errors-in-variables regression models with missing data. Communications in Statistics: Simulation and Computation, 45(2):671–688, 2016. CODEN CSSCDB. ISSN 0361-0918.

Sun:2017:TAR

[Sun17] Huihui Sun. Testing for autocorrelation and random-effects in nonlinear mixed effects models based on M-estimation. Communications in Statistics: Simulation and Computation, 46(9):7125–7137, 2017. CODEN CSSCDB. ISSN 0361-0918.

Samawi:2013:MEA

[SV13] Hani M. Samawi and Robert Vogel. More efficient approximation of multiple integrals using steady state ranked simulated sampling. Communications in Statistics: Simulation and Computation, 42(2):370–381, 2013. CODEN CSSCDB. ISSN 0361-0918.

Shan:2010:SEE

[SVWH10] Guogen Shan, Albert Vexler, Gregory E. Wilding, and Alan D. Hutson. Simple and exact empirical likelihood ratio tests for normality based on moment relations. Communications in Statistics: Simulation and Computation, 40(1): 129–146, 2010. CODEN CSSCDB. ISSN 0361-0918.

Shen:2011:ACB

[SW11a] Chuanhe Shen and Xiangrong Wang. Analysis of convertible bond value based on integration of support vector machine and copula function. Communications in Statistics: Simulation and Computation, 40(10):1563–1575, 2011. CODEN CSSCDB. ISSN 0361-0918.

Stehlik:2011:ELR

[SW11b] Milan Stehlík and Helga Wagner. Exact likelihood ratio testing for homogeneity of the exponential distribution. Communications in Statistics: Simulation and Computation, 40 (5):663–684, 2011. CODEN CSSCDB. ISSN 0361-0918.

Swift:2013:MBA

[SW13] Andrew W. Swift and Beichen Wang. Moment-based approximations of discrete probability distributions using rational functions. Communications in Statistics: Simulation and Computation, 42(10):2203–2222, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Shieh:2016:MCD

[SW16] Jiunn-I Shieh and Hsin-Hung Wu. Measures of consistency for DEMATEL method. Communications in Statistics: Simulation and Computation, 45(3):781–790, 2016. CODEN CSSCDB. ISSN 0361-0918.

Sun:2018:SEC

[SWGX18] Lirong Sun, Bing Xing Wang, Baocai Guo, and Min Xie. Synthetic exponential control charts with unknown parameter. Communications in Statistics: Simulation and Computation, 47(8):2360–2377, 2018. CODEN CSSCDB. ISSN 0361-0918.

Shieh:2014:ATV

[SWL14] Jiunn-I Shieh, Hsin-Hung Wu, and Hsiang-Chuan Liu. Analysis of the threshold values of semantic structure analysis in identifying causal relationships. Communications in Statistics: Simulation and Computation, 43(7):1543–1551, 2014. CODEN CSSCDB. ISSN 0361-0918.

Shojaei:2017:SCS

[SWM17] S. Reza H. Shojaei, Yadollah Waghei, and Mohsen Mohammadzadeh. A simulation comparison on spatial Poisson mixed models. Communications in Statistics: Simulation and Computation, 46(8):6208–6220, 2017. CODEN CSSCDB. ISSN 0361-0918.

Song:2017:ICS

[SXYF17] Huaming Song, Qian Xu, Hui Yang, and Jun Fang. Interpreting out-of-control signals using instance-based Bayesian classifier in multivariate statistical process control. Communications in Statistics: Simulation and Computation, 46(1): 53–77, 2017. CODEN CSSCDB. ISSN 0361-0918.

Salmasi:2017:GOS

[SY17a] Mehdi Rajaei Salmasi and Gholamhosein Yari. On generalized order statistics from generalized Pareto distribution. Communications in Statistics: Simulation and Computation, 46(7):5682–5697, 2017. CODEN CSSCDB. ISSN 0361-0918.

Schroder:2017:RSS

[SY17b] Carsten Schröder and Shlomo Yitzhaki. Reasonable sample sizes for convergence to normality. Communications in Statistics: Simulation and Computation, 46(9):7074–7087, 2017. CODEN CSSCDB. ISSN 0361-0918.

Shia:2016:P

[SZ16] Ben-Chang Shia and Bo Zhang. Preface. Communications in Statistics: Simulation and Computation, 45(4):1159, 2016. CODEN CSSCDB. ISSN 0361-0918.

Shia:2011:FCD

[SZFM11] Ben-Chang Shia, Jianping Zhu, Kuangnan Fang, and Shuangge Ma. Fuzzy canonical discriminant analysis: Theory and practice. Communications in Statistics: Simulation and Computation, 40(10):1526–1539, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Shadman:2017:CPM

[SZMY17] Alireza Shadman, Changliang Zou, Hashem Mahlooji, and Arthur B. Yeh. A change point method for Phase II monitoring of generalized linear profiles. Communications in Statistics: Simulation and Computation, 46(1):559–578, 2017. CODEN CSSCDB. ISSN 0361-0918.

Strand:2010:MNR

[SZS10] Matthew Strand, Yu Zhang, and Bruce J. Swihart. Monotone nonparametric regression and confidence intervals.

Communications in Statistics: Simulation and Computation, 39(4):828–845, 2010. CODEN CSSCDB. ISSN 0361-0918.

Shahid:2017:SIC

[SZSA17] Muhammad Rizwan Shahid, Muhammad Zakria, Farrukh Shehzad, and Rashid Ahmed. Some important classes of generalized neighbor designs for linear blocks. Communications in Statistics: Simulation and Computation, 46(3): 1991–1997, 2017. CODEN CSSCDB. ISSN 0361-0918.

Tovar:2013:DBT

[TA13] José Rafael Tovar and Jorge Alberto Achcar. Dependence between two diagnostic tests with copula function approach: A simulation study. Communications in Statistics: Simulation and Computation, 42(2):454–475, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Thanoon:2017:MCL

[TA17] Thanoon Y. Thanoon and Robiah Adnan. Model comparison of linear and nonlinear Bayesian structural equation models with dichotomous data. Communications in Statistics: Simulation and Computation, 46(6):4578–4599, 2017. CODEN CSSCDB. ISSN 0361-0918.

Tang:2016:NKM

[Tan16] Yongqiang Tang. Notes on Kolassa’s method for estimating the power of Wilcoxon–Mann–Whitney test. Communications in Statistics: Simulation and Computation, 45(1): 240–251, 2016. CODEN CSSCDB. ISSN 0361-0918.

Tastan:2017:SBE

[Tas17] Hüseyin Tastan. Simulation-based estimation of threshold moving average models with contemporaneous shock asymmetry and an application to Turkish business cycles. Communications in Statistics: Simulation and Computation, 46 (5):3870–3891, 2017. CODEN CSSCDB. ISSN 0361-0918.

Tasdan:2018:SRT

[Tas18] Feridun Tasdan. Smoothed ranks for two or multi-sample location problems. Communications in Statistics: Simulation and Computation, 47(2):526–541, 2018. CODEN CSSCDB. ISSN 0361-0918.

Tabass:2017:GMT

[TB17] M. Sanei Tabass and G. R. Mohtashami Borzadaran. The generalized maximum Tsallis entropy estimators and applications to the Portland cement dataset. Communications in Statistics: Simulation and Computation, 46(4):3284–3293, 2017. CODEN CSSCDB. ISSN 0361-0918.

Temsamani:2014:RBF

[TC14] Jamal Temsamani and Juan A. Carrasco. Reliability bounds for fault-tolerant systems with deferred repair using bounding split regenerative randomization. Communications in Statistics: Simulation and Computation, 43(1):183–208, 2014. CODEN CSSCDB. ISSN 0361-0918.

Tahir:2016:NWP

[TCA^{+}16] M. H. Tahir, Gauss M. Cordeiro, Ayman Alzaatreh, M. Mansoor, and M. Zubair. A new Weibull–Pareto distribution: Properties and applications. Communications in Statistics: Simulation and Computation, 45(10):3548–??, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Tenreiro:2017:NTM

[Ten17] C. Tenreiro. A new test for multivariate normality by combining extreme and nonextreme BHEP tests. Communications in Statistics: Simulation and Computation, 46(3): 1746–1759, 2017. CODEN CSSCDB. ISSN 0361-0918.

Tewfik:2017:OTT

[Tew17] Lounis Tewfik. Optimal tests in AR(m) time series model. Communications in Statistics: Simulation and Computation, 46(2):1583–1610, 2017. CODEN CSSCDB. ISSN 0361-0918.

Tamura:2010:PML

[TG10] Karin Ayumi Tamura and Viviana Giampaoli. Prediction in multilevel logistic regression. Communications in Statistics: Simulation and Computation, 39(6):1083–1096, 2010. CODEN CSSCDB. ISSN 0361-0918.

Tercero-Gomez:2012:MTC

[TGRGCF^{+}12] Victor Tercero-Gomez, Jose Ramirez-Galindo, Alavarado Cordero-Franco, Milton Smith, and Mario Beruvides. Modified Tukey’s control chart. Communications in Statistics:

Simulation and Computation, 41(9):1566–1579, 2012. CO-DEN CSSCDB. ISSN 0361-0918. See note [LT15].

Tuac:2018:RPE

[TGSA18] Y. Tuaç, Y. Güney, B. Senoglu, and O. Arslan. Robust parameter estimation of regression model with AR(p) error terms. Communications in Statistics: Simulation and Computation, 47(8):2343–2359, 2018. CODEN CSSCDB. ISSN 0361-0918.

Turner:2017:EMC

[TJ17] Rolf Turner and Celeste Jeffs. An extension of Monte Carlo hypothesis tests. Communications in Statistics: Simulation and Computation, 46(8):6545–6558, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Tadayon:2015:BAS

[TK15a] Vahid Tadayon and Majid Jafari Khaledi. Bayesian analysis of skew Gaussian spatial models based on censored data. Communications in Statistics: Simulation and Computation, 44(9):2431–2441, 2015. CODEN CSSCDB. ISSN 0361-0918.

Tuerhong:2015:CNS

[TK15b] Gulanbaier Tuerhong and Seoung Bum Kim. Comparison of novelty score-based multivariate control charts. Communications in Statistics: Simulation and Computation, 44(5): 1126–1143, 2015. CODEN CSSCDB. ISSN 0361-0918.

Tu:2016:UIP

[TK16] Chunhao Tu and Woon Yuen Koh. Using inverse probability weighting estimators to evaluate various propensity scores when treatment switching exists. Communications in Statistics: Simulation and Computation, 45(6):2182–2190, 2016. CODEN CSSCDB. ISSN 0361-0918.

Tu:2017:CBS

[TK17] Chunhao Tu and Woon Yuen Koh. A comparison of balancing scores for estimating rate ratios of count data in observational studies. Communications in Statistics: Simulation and Computation, 46(1):772–778, 2017. CODEN CSSCDB. ISSN 0361-0918.

Teoh:2016:ERL

[TKCL16] W. L. Teoh, Michael B. C. Khoo, Philippe Castagliola, and M. H. Lee. The exact run length distribution and design of the Shewhart chart with estimated parameters based on median run length. Communications in Statistics: Simulation and Computation, 45(6):2081–2103, 2016. CODEN CSSCDB. ISSN 0361-0918.

Tuerhong:2014:MCC

[TKKC14] Gulanbaier Tuerhong, Seoung Bum Kim, Pilsung Kang, and Sungzoon Cho. Multivariate control charts based on hybrid novelty scores. Communications in Statistics: Simulation and Computation, 43(1):115–131, 2014. CODEN CSSCDB. ISSN 0361-0918.

Tang:2013:SSM

[TKX13] Fei Tang, B. M. Golam Kibria, and Huiliang Xie. Some statistical methods in assessing linear relational agreement: An empirical comparative study. Communications in Statistics: Simulation and Computation, 42(2):263–279, 2013. CODEN CSSCDB. ISSN 0361-0918.

Tian:2015:NAN

[TL15] Ye Tian and Heng Lian. A note on application of Nesterov’s method in solving lasso-type problems. Communications in Statistics: Simulation and Computation, 44(7):1673–1682, 2015. CODEN CSSCDB. ISSN 0361-0918.

Lin:2014:TMO

[tLB14] Chien tai Lin and N. Balakrishnan. Tests for multiple outliers in an exponential sample. Communications in Statistics: Simulation and Computation, 43(4):706–722, 2014. CODEN CSSCDB. ISSN 0361-0918.

Torabi:2014:LUD

[TM14] Hamzeh Torabi and Narges H. Montazeri. The logisticuniform distribution and its applications. Communications in Statistics: Simulation and Computation, 43(10):2551– 2569, 2014. CODEN CSSCDB. ISSN 0361-0918.

Tamandi:2016:EPK

[TN16] Mostafa Tamandi and Saralees Nadarajah. On the estimation of parameters of Kumaraswamy-G distributions. Communications in Statistics: Simulation and Computation, 45 (10):3811–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Teimouri:2017:STS

[TN17] Mahdi Teimouri and Saralees Nadarajah. On simulating truncated skewed Cauchy random variables. Communications in Statistics: Simulation and Computation, 46(2): 1318–1321, 2017. CODEN CSSCDB. ISSN 0361-0918.

Tarvirdizade:2019:NFH

[TN19] Bahman Tarvirdizade and Nader Nematollahi. A new flexible hazard rate distribution: Application and estimation of its parameters. Communications in Statistics: Simulation and Computation, 48(3):882–899, 2019. CODEN CSSCDB. ISSN 0361-0918.

Tay:2019:PEM

[TNO19] S. Y. Tay, C. M. Ng, and S. H. Ong. Parameter estimation by minimizing a probability generating function-based power divergence. Communications in Statistics: Simulation and Computation, 48(10):2898–2912, 2019. CODEN CSSCDB. ISSN 0361-0918.

Takahashi:2012:STI

[TNS12] Sho Takahashi, Takahiro Nishiyama, and Takashi Seo. Simultaneous tests for independence among components of random vector by step-down multiple comparison procedure. Communications in Statistics: Simulation and Computation, 41(10):1998–2005, 2012. CODEN CSSCDB. ISSN 0361-0918.

Tabakan:2017:LEP

[TÖ17] Gülin Tabakan and M. Revan Özkale. Liu estimator in partly linear regression models with correlated errors. Communications in Statistics: Simulation and Computation, 46(3): 1958–1973, 2017. CODEN CSSCDB. ISSN 0361-0918.

Turkan:2018:JEN

[TÖ18] Semra Türkan and Gamze Özel. A jackknifed estimators for the negative binomial regression model. Communications in Statistics: Simulation and Computation, 47(6):1845–1865, 2018. CODEN CSSCDB. ISSN 0361-0918.

Tommasi:2012:ODR

[Tom12] C. Tommasi. Optimal design robust to a misspecified model. Communications in Statistics: Simulation and Computation, 41(7):1220–1231, 2012. CODEN CSSCDB. ISSN 0361-0918.

Torabi:2015:LIS

[Tor15] Mahmoud Torabi. Likelihood inference for spatial generalized linear mixed models. Communications in Statistics: Simulation and Computation, 44(7):1692–1701, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Touboul:2011:PPT

[Tou11] Jacques Touboul. Projection pursuit through relative entropy minimization. Communications in Statistics: Simulation and Computation, 40(6):854–878, 2011. CODEN CSS-CDB. ISSN 0361-0918.

Tripathy:2015:NED

[TP15] Manas Ranjan Tripathy and Nabendu Pal. A note on estimation of a distribution function in a nonparametric setup using Stein’s shrinkage estimation technique. Communications in Statistics: Simulation and Computation, 44(10): 2731–2741, 2015. CODEN CSSCDB. ISSN 0361-0918.

Trivedi:2017:EPM

[TP17] B. S. Trivedi and M. N. Patel. Estimation of parameters of Misclassified Size Biased Discrete Lindley Distribution. Communications in Statistics: Simulation and Computation, 46(7):5541–5552, 2017. CODEN CSSCDB. ISSN 0361-0918.

Teel:2015:EEF

[TPS15] Chen Teel, Taeyoung Park, and Allan R. Sampson. EM estimation for finite mixture models with known mixture component size. Communications in Statistics: Simulation and Computation, 44(6):1545–1556, 2015. CODEN CSSCDB. ISSN 0361-0918.

Tikkiwal:2013:PGR

[TRG13] G. C. Tikkiwal, Piyush Kant Rai, and Alka Ghiya. On the performance of generalized regression estimator for small domains. Communications in Statistics: Simulation and Computation, 42(4):891–909, 2013. CODEN CSSCDB. ISSN 0361-0918.

Triantafyllou:2018:NCC

[Tri18a] Ioannis S. Triantafyllou. Nonparametric control charts based on order statistics: Some advances. Communications in Statistics: Simulation and Computation, 47(9):2684–2702, 2018. CODEN CSSCDB. ISSN 0361-0918.

Tripathy:2018:IEC

[Tri18b] Manas Ranjan Tripathy. Improved estimation of common location of two exponential populations with order restricted scale parameters using censored samples. Communications in Statistics: Simulation and Computation, 47(9): 2800–2818, 2018. CODEN CSSCDB. ISSN 0361-0918.

Teimouri:2018:EAS

[TRM18] Mahdi Teimouri, Saeid Rezakhah, and Adel Mohammadpour. EM algorithm for symmetric stable mixture model. Communications in Statistics: Simulation and Computation, 47(2):582–604, 2018. CODEN CSSCDB. ISSN 0361-0918.

Tseng:2011:EES

[TS11] Yi-Kuan Tseng and Ken-Ning Shu. Efficient estimation for a semiparametric extended hazards model. Communications in Statistics: Simulation and Computation, 40(2):258–273, 2011. CODEN CSSCDB. ISSN 0361-0918.

Teles:2017:ETA

[TS17] Paulo Teles and Paulo S. A. Sousa. The effect of temporal aggregation on the estimation accuracy of time series models. Communications in Statistics: Simulation and Computation, 46(9):6738–6759, 2017. CODEN CSSCDB. ISSN 0361-0918.

Teles:2018:ETA

[TS18] Paulo Teles and Paulo S. A. Sousa. The effect of temporal aggregation on the estimation accuracy of ARMA models. Communications in Statistics: Simulation and Computation, 47(10):2865–2885, 2018. CODEN CSSCDB. ISSN 0361-0918.

Tian:2019:BLR

[TSL^{+}19] Yuzhu Tian, Silian Shen, Ge Lu, Manlai Tang, and Maozai Tian. Bayesian LASSO-regularized quantile regression for linear regression models with autoregressive errors. Communications in Statistics: Simulation and Computation, 48 (3):777–796, 2019. CODEN CSSCDB. ISSN 0361-0918.

Tudor:2014:EMW

[TT14] Ciprian A. Tudor and Cristiana Tudor. EGARCH model with weighted liquidity. Communications in Statistics: Simulation and Computation, 43(5):1133–1142, 2014. CODEN CSSCDB. ISSN 0361-0918.

Turkmen:2015:ORE

[TT15] Asuman S. Turkmen and Gulin Tabakan. Outlier resistant estimation in difference-based semiparametric partially linear models. Communications in Statistics: Simulation and Computation, 44(2):417–432, 2015. CODEN CSSCDB. ISSN 0361-0918.

Tung:2019:BAL

[TTC19] Dao Thanh Tung, Minh-Ngoc Tran, and Tran Manh Cuong. Bayesian adaptive lasso with variational Bayes for variable selection in high-dimensional generalized linear mixed models. Communications in Statistics: Simulation and Computation, 48(2):530–543, 2019. CODEN CSSCDB. ISSN 0361-0918.

Tian:2014:TLE

[TTZ14] Yuzhu Tian, Maozai Tian, and Qianqian Zhu. Transmuted linear exponential distribution: A new generalization of the linear exponential distribution. Communications in Statistics: Simulation and Computation, 43(10):2661–2677, 2014. CODEN CSSCDB. ISSN 0361-0918.

Tarima:2013:RME

[TVS13] Sergey S. Tarima, Albert Vexler, and Siddhartha Singh. Robust mean estimation under a possibly incorrect lognormality assumption. Communications in Statistics: Simulation and Computation, 42(2):316–326, 2013. CODEN CSS-CDB. ISSN 0361-0918.

Tsai:2011:MSC

[TW11] Shing Chih Tsai and Jung Wei. Multinomial selection for comparison with a standard. Communications in Statistics: Simulation and Computation, 40(6):926–947, 2011. CODEN CSSCDB. ISSN 0361-0918.

Tan:2016:LMB

[TW16] Lixiang Tan and Lihong Wang. The LASSO method for bilinear time series models. Communications in Statistics: Simulation and Computation, 45(3):1072–1082, 2016. CO-DEN CSSCDB. ISSN 0361-0918.

Tian:2015:STG

[TX15] Ruiqin Tian and Liugen Xue. Smooth-threshold GEE variable selection in high-dimensional partially linear models with longitudinal data. Communications in Statistics: Simulation and Computation, 44(7):1720–1734, 2015. CODEN CSSCDB. ISSN 0361-0918.

Tseng:2016:KSA

[TY16] Y. K. Tseng and Y. F. Yang. A kernel smooth approach for joint modeling of accelerated failure time and longitudinal data. Communications in Statistics: Simulation and Computation, 45(4):1240–1248, 2016. CODEN CSSCDB. ISSN 0361-0918.

Tang:2014:BAN

[TZ14] Nian-Sheng Tang and Hui Zhao. Bayesian analysis of nonlinear reproductive dispersion mixed models for longitudinal data with nonignorable missing covariates. Communications in Statistics: Simulation and Computation, 43(6): 1265–1287, 2014. CODEN CSSCDB. ISSN 0361-0918.

Tong:2013:LRC

[TZD^{+}13] Qiaoling Tong, Xuecheng Zou, Qian Ding, Tianzhen Liu, and Hengqing Tong. Latent regression with constrained parameters to determine the weight coefficients in summary index model. Communications in Statistics: Simulation and Computation, 42(7):1628–1642, 2013. CODEN CSSCDB. ISSN 0361-0918.

Urooj:2017:APT

[UA17] Amena Urooj and Zahid Asghar. Analysis of the performance of test statistics for detection of outliers (additive, innovative, transient, and level shift) in AR(1) processes. Communications in Statistics: Simulation and Computation, 46(2):948–979, 2017. CODEN CSSCDB. ISSN 0361-0918.

Umali:2014:NPC

[UB14] Jennifer Umali and Erniel Barrios. Nonparametric principal components regression. Communications in Statistics: Simulation and Computation, 43(7):1797–1810, 2014. CODEN CSSCDB. ISSN 0361-0918.

Uddin:2017:TET

[UH17] Nizam Uddin and M. S. Hasan. Testing equality of two normal means using combined samples of paired and unpaired data. Communications in Statistics: Simulation and Computation, 46(3):2430–2446, 2017. CODEN CSSCDB. ISSN 0361-0918.

Uhm:2011:ARM

[UHP11] Daiho Uhm, Fred W. Huffer, and Cheolyong Park. Additive risk model using piecewise constant hazard function. Communications in Statistics: Simulation and Computation, 40 (9):1458–1477, 2011. CODEN CSSCDB. ISSN 0361-0918.

UlIslam:2017:SBR

[Ul 17] Tanweer Ul Islam. Stringency-based ranking of normality tests. Communications in Statistics: Simulation and Computation, 46(1):655–668, 2017. CODEN CSSCDB. ISSN 0361-0918.

Uzuke:2017:PKK

[UMN17] C. A. Uzuke, J. I. Mbegbu, and C. R. Nwosu. Performance of Kibria, Khalaf, and Shurkur’s methods when the eigenvalues are skewed. Communications in Statistics: Simulation and Computation, 46(3):2071–2102, 2017. CODEN CSSCDB. ISSN 0361-0918.

Uraibi:2017:SOF

[UMR17] Hassan S. Uraibi, Habshah Midi, and Sohel Rana. Selective overview of forward selection in terms of robust correlations. Communications in Statistics: Simulation and Computation, 46(7):5479–5503, 2017. CODEN CSSCDB. ISSN 0361-0918.

Uys:2010:UTS

[US10] D. W. Uys and S. J. Steel. Using a truncated C_{p} statistic for variable selection in multiple linear regression. Communications in Statistics: Simulation and Computation, 39(2): 420–432, 2010. CODEN CSSCDB. ISSN 0361-0918.

Usami:2017:BLP

[Usa17] Satoshi Usami. Bayesian longitudinal paired comparison model and its application to sports data using weighted likelihood bootstrap. Communications in Statistics: Simulation and Computation, 46(3):1974–1990, 2017. CODEN CSSCDB. ISSN 0361-0918.

Veerakumari:2017:ESS

[VA17] K. Pradeepa Veerakumari and S. Azarudheen. Evaluation of single sampling plan under the conditions of intervened Poisson distribution. Communications in Statistics: Simulation and Computation, 46(8):6106–6114, 2017. CODEN CSSCDB. ISSN 0361-0918.

Villasenor-Alva:2015:CTN

[VAGE15] José A. Villaseñor-Alva and Elizabeth González-Estrada. A correlation test for normality based on the lévy characterization. Communications in Statistics: Simulation and Computation, 44(5):1225–1238, 2015. CODEN CSSCDB. ISSN 0361-0918.

VanZyl:2012:MRM

[Van12] J. Martin Van Zyl. A median regression model to estimate the parameters of the three-parameter generalized Pareto distribution. Communications in Statistics: Simulation and Computation, 41(4):544–553, 2012. CODEN CSSCDB. ISSN 0361-0918.

VanZyl:2015:ALA

[Van15] J. Martin Van Zyl. Applying least absolute deviation regression to regression-type estimation of the index of a stable distribution using the characteristic function. Communications in Statistics: Simulation and Computation, 44(9): 2442–2462, 2015. CODEN CSSCDB. ISSN 0361-0918.

Valiollahi:2019:PFF

[VAR19] R. Valiollahi, A. Asgharzadeh, and M. Z. Raqab. Prediction of future failures times based on Type-I hybrid censored samples of random sample sizes. Communications in Statistics: Simulation and Computation, 48(1):109–125, 2019. CODEN CSSCDB. ISSN 0361-0918.

Vijayaraghavan:2016:BDS

[VB16] R. Vijayaraghavan and S. Banumathi. Bayesian design of special type of double sampling plans for compliance testing. Communications in Statistics: Simulation and Computation, 45(1):362–377, 2016. CODEN CSSCDB. ISSN 0361-0918.

Vajargah:2013:OHS

[VC13] Behrouz Fathi Vajargah and Asghar Eskandari Chechaglou. Optimal Halton sequence via inversive scrambling. Communications in Statistics: Simulation and Computation, 42(2): 476–484, 2013. CODEN CSSCDB. ISSN 0361-0918.

Vajargah:2015:NMS

[VC15] Behrouz Fathi Vajargah and Asghar Eskandari Chechaglou. New modified scrambled Faure sequences. Communications in Statistics: Simulation and Computation, 44(3):666–682, 2015. CODEN CSSCDB. ISSN 0361-0918.

Volterman:2014:FRO

[VCDB14] W. Volterman, E. Cramer, K. F. Davies, and N. Balakrishnan. Further results on order statistics generated by two simulation methods. Communications in Statistics: Simulation and Computation, 43(10):2732–2743, 2014. CODEN CSSCDB. ISSN 0361-0918.

vandenHeuvel:2010:CEM

[vdH10] Edwin R. van den Heuvel. A comparison of estimation methods on the coverage probability of Satterthwaite confidence intervals for assay precision with unbalanced data. Communications in Statistics: Simulation and Computation, 39(4): 777–794, 2010. CODEN CSSCDB. ISSN 0361-0918.

Vexler:2010:DBE

[VG10] Albert Vexler and Gregory Gurevich. Density-based empirical likelihood ratio change point detection policies. Communications in Statistics: Simulation and Computation, 39 (9):1709–1725, 2010. CODEN CSSCDB. ISSN 0361-0918.

Vajargah:2015:RPF

[VG15] Behrouz Fathi Vajargah and Maryam Gharehdaghi. Reducing periodicity of fuzzy Markov chains based on simulation using Halton sequence. Communications in Statistics: Simulation and Computation, 44(8):2169–2175, 2015. CODEN CSSCDB. ISSN 0361-0918.

Villagran:2016:NPS

[VHV^{+}16] Alejandro Villagran, Gabriel Huerta, Marina Vannucci, Charles S. Jackson, and Alvaro Nosedal. Non-parametric sampling approximation via Voronoi tessellations. Communications in Statistics: Simulation and Computation, 45(2): 717–736, 2016. CODEN CSSCDB. ISSN 0361-0918.

Viljoen:2017:CSC

[Vil17] Helena Viljoen. A comparison of stepwise common singular spectrum analysis and horizontal multi-channel singular spectrum analysis. Communications in Statistics: Simulation and Computation, 46(9):6865–6878, 2017. CODEN CSSCDB. ISSN 0361-0918.

Vinod:2017:GCK

[Vin17] Hrishikesh D. Vinod. Generalized correlation and kernel causality with applications in development economics. Communications in Statistics: Simulation and Computation, 46 (6):4513–4534, 2017. CODEN CSSCDB. ISSN 0361-0918.

Vieira:2010:MSP

[VKB10] Hélcio Vieira, Jr., Karl Heinz Kienitz, and Mischel Carmen Neyra Belderrain. Multinomial selection problem: A study of BEM and AVC algorithms. Communications in Statistics: Simulation and Computation, 39(5):971–980, 2010. CODEN CSSCDB. ISSN 0361-0918.

Valle:2010:NMM

[VL10] Luciana Dalla Valle and Fabrizio Leisen. A new multinomial model and a zero variance estimation. Communications in Statistics: Simulation and Computation, 39(4):846–859, 2010. CODEN CSSCDB. ISSN 0361-0918.

Volaufova:2013:SCA

[VL13] Julia Volaufova and Lynn Roy Lamotte. A simulation comparison of approximate tests for fixed effects in random coefficients growth curve models. Communications in Statistics:

Simulation and Computation, 42(2):344–359, 2013. CODEN CSSCDB. ISSN 0361-0918.

VanOirbeek:2016:ECE

[VL16] Robin Van Oirbeek and Emmanuel Lesaffre. Exploring the clustering effect of the frailty survival model by means of the Brier score. Communications in Statistics: Simulation and Computation, 45(9):3294–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Vahl:2011:WPE

[VM11] Christopher I. Vahl and George A. Milliken. Whole-plot exchange algorithms for constructing D-optimal multistratum designs. Communications in Statistics: Simulation and Computation, 40(7):1030–1042, 2011. CODEN CSSCDB. ISSN 0361-0918.

vanNiekerk:2017:GMC

[vNBA17] Janet van Niekerk, Andriëtte Bekker, and Mohammad Arashi. A gamma-mixture class of distributions with Bayesian application. Communications in Statistics: Simulation and Computation, 46(10):8152–8165, 2017. CODEN CSSCDB. ISSN 0361-0918.

Verma:2019:BBP

[VND19] Vivek Verma, Dilip C. Nath, and Radhakanta Das. Bayesian bounds for population proportion under ranked set sampling. Communications in Statistics: Simulation and Computation, 48(2):478–493, 2019. CODEN CSSCDB. ISSN 0361-0918.

Volaufova:2014:ASS

[Vol14] Julia Volaufova. Approximate small-sample tests of fixed effects in nonlinear mixed models. Communications in Statistics: Simulation and Computation, 43(9):2117–2137, 2014. CODEN CSSCDB. ISSN 0361-0918.

Vidal-Puig:2014:CSD

[VPF14] S. Vidal-Puig and A. Ferrer. A comparative study of different methodologies for fault diagnosis in multivariate quality control. Communications in Statistics: Simulation and Computation, 43(5):986–1005, 2014. CODEN CSSCDB. ISSN 0361-0918.

Voinov:2013:GFT

[VPSV13] Vassilly Voinov, Natalya Pya, Niyaz Shapakov, and Yevgeniy Voinov. Goodness-of-fit tests for the power-generalized Weibull probability distribution. Communications in Statistics: Simulation and Computation, 42(5):1003–1012, 2013. CODEN CSSCDB. ISSN 0361-0918.

Vrbik:2015:MAK

[Vrb15] Jan Vrbik. Moments of AR(k) parameter estimators. Communications in Statistics: Simulation and Computation, 44 (5):1239–1252, 2015. CODEN CSSCDB. ISSN 0361-0918.

Vaysi:2018:SAR

[VRN18] Mohammad Salehi Vaysi, Sadegh Rezaei, and Saralees Nadarajah. Sensitivity analysis of reliability functions of the exponential power series lifetime distribution. Communications in Statistics: Simulation and Computation, 47(10): 2938–2952, 2018. CODEN CSSCDB. ISSN 0361-0918.

Villa-Vialaneix:2016:CWE

[VVHP^{+}16] Nathalie Villa-Vialaneix, Noslen Hernández, Alain Paris, Céline Domange, Nathalie Priymenko, and Philippe Besse. On combining wavelets expansion and sparse linear models for regression on metabolomic data and biomarker selection. Communications in Statistics: Simulation and Computation, 45(1):282–298, 2016. CODEN CSSCDB. ISSN 0361-0918.

vanZyl:2011:AKS

[vZ11] J. Martin van Zyl. Application of the Kolmogorov–Smirnov test to estimate the threshold when estimating the extreme value index. Communications in Statistics: Simulation and Computation, 40(2):199–207, 2011. CODEN CSSCDB. ISSN 0361-0918.

vanZyl:2018:PUG

[vZ18] J. Martin van Zyl. The performance of univariate goodnessof-fit tests for normality based on the empirical characteristic function in large samples. Communications in Statistics: Simulation and Computation, 47(4):1146–1156, 2018. CO-DEN CSSCDB. ISSN 0361-0918.

vanZyl:2012:PET

[vZS12] J. Martin van Zyl and Robert Schall. Parameter estimation through weighted least-squares rank regression with specific reference to the Weibull and Gumbel distributions. Communications in Statistics: Simulation and Computation, 41 (9):1654–1666, 2012. CODEN CSSCDB. ISSN 0361-0918.

Wang:2014:ABC

[WA14] Jing Wang and Yves F. Atchadé. Approximate Bayesian computation for exponential random graph models for large social networks. Communications in Statistics: Simulation and Computation, 43(2):359–377, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wu:2017:MRL

[WA17] Jibo Wu and Yasin Asar. More on the restricted Liu estimator in the logistic regression model. Communications in Statistics: Simulation and Computation, 46(5):3680–3689, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wu:2018:PAU

[WA18] Jibo Wu and Yasin Asar. Performance of the almost unbiased ridge-type principal component estimator in logistic regression model. Communications in Statistics: Simulation and Computation, 47(10):2925–2937, 2018. CODEN CSSCDB. ISSN 0361-0918.

Walker:2011:PSW

[Wal11] Stephen G. Walker. Posterior sampling when the normalizing constant is unknown. Communications in Statistics: Simulation and Computation, 40(5):784–792, 2011. CODEN CSSCDB. ISSN 0361-0918.

Walker:2014:CML

[Wal14] Stephen G. Walker. Computing marginal likelihoods via posterior sampling. Communications in Statistics: Simulation and Computation, 43(3):520–527, 2014. CODEN CSSCDB. ISSN 0361-0918.

Walker:2017:IAS

[Wal17] Stephen G. Walker. An iterative algorithm for solving sparse linear equations. Communications in Statistics: Simulation and Computation, 46(7):5113–5122, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Wang:2010:DPN

[Wan10a] Jing Wang. Dirichlet processes in nonlinear mixed effects models. Communications in Statistics: Simulation and Computation, 39(3):539–556, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wang:2010:SEC

[Wan10b] Qin Wang. A shrinkage estimation of central subspace in sufficient dimension reduction. Communications in Statistics: Simulation and Computation, 39(10):1868–1876, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wang:2012:CTA

[Wan12] Yuchung J. Wang. Comparisons of three approaches for discrete conditional models. Communications in Statistics: Simulation and Computation, 41(1):32–43, 2012. CODEN CSSCDB. ISSN 0361-0918.

Wang:2014:UBA

[Wan14] Fu-Kwun Wang. Using BBPSO algorithm to estimate the Weibull parameters with censored data. Communications in Statistics: Simulation and Computation, 43(10):2614–2627, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wang:2017:APG

[Wan17] Li Wang. Adaptive procedure for generalized familywise error rate control. Communications in Statistics: Simulation and Computation, 46(10):8140–8151, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Wang:2013:CNG

[WANB13] Jung-Chao Wang, Hend Auda, and Magdalena Niewiadomska-Bugaj. Comparing a new Gini test with other symmetry tests when median is unknown. Communications in Statistics: Simulation and Computation, 42(7):1616–1627, 2013. CODEN CSSCDB. ISSN 0361-0918.

Warrens:2017:SKF

[War17] Matthijs J. Warrens. Symmetric kappa as a function of unweighted kappas. Communications in Statistics: Simulation and Computation, 46(7):5240–5245, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wolters:2018:ESC

[WB18] Mark A. Wolters and W. John Braun. Enforcing shape constraints on a probability density estimate using an additive adjustment curve. Communications in Statistics: Simulation and Computation, 47(3):672–691, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Wang:2016:MSA

[WBG16a] X. Wang, N. Balakrishnan, and B. Guo. Mis-specification analyses of nonlinear Wiener process-based degradation models. Communications in Statistics: Simulation and Computation, 45(3):814–832, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wang:2016:RLE

[WBG16b] Xiaolin Wang, Narayanaswamy Balakrishnan, and Bo Guo. Residual life estimation based on a generalized Wiener process with skew-normal random effects. Communications in Statistics: Simulation and Computation, 45(6):2158–2181, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2015:HGT

[WC15] Rand R. Wilcox and Florence Clark. Heteroscedastic global tests that the regression parameters for two or more independent groups are identical. Communications in Statistics: Simulation and Computation, 44(3):773–786, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wu:2016:OCR

[WC16] Shu-Fei Wu and Hubert J. Chen. An optimal confidence region for the largest and the smallest means from a multivariate normal distribution. Communications in Statistics: Simulation and Computation, 45(3):952–967, 2016. CODEN CSSCDB. ISSN 0361-0918.

Weng:2017:CRD

[WC17] Li-Jen Weng and Chung-Ping Cheng. Is categorization of random data necessary for parallel analysis on Likert-type data? Communications in Statistics: Simulation and Computation, 46(7):5367–5377, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wu:2016:SIL

[WCC16] Chin-Chuan Wu, Li-Ching Chen, and Yi-Jie Chen. Statistical inferences for the lifetime performance index of the products with the Gompertz distribution under censored samples. Communications in Statistics: Simulation and Computation, 45(4):1318–1336, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wu:2018:HTP

[WCC^{+}18] Shu-Fei Wu, Tzu-Chin Chen, Wen-Jui Chang, Chia-Wei Chang, and Chen Lin. A hypothesis testing procedure for the evaluation on the lifetime performance index of products with Burr XII distribution under progressive type I interval censoring. Communications in Statistics: Simulation and Computation, 47(9):2670–2683, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wang:2012:PAL

[WCL12] F. K. Wang, Y. F. Cheng, and W. L. Lu. Partially accelerated life tests for the Weibull distribution under multiply censored data. Communications in Statistics: Simulation and Computation, 41(9):1667–1678, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Wen:2017:MWS

[WCSL17] Yanqing Wen, Lirong Cui, Shubin Si, and Baoliang Liu. A multiple warm standby δ-shock system with a repairman having multiple vacations. Communications in Statistics: Simulation and Computation, 46(4):3172–3186, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Waymyers:2018:NGG

[WDC18] Sophia D. Waymyers, Sanku Dey, and Hrishikesh Chakraborty. A new generalization of the gamma distribution with application to negatively skewed survival data. Communications in Statistics: Simulation and Computation, 47(7): 2083–2101, 2018. CODEN CSSCDB. ISSN 0361-0918.

Weiss:2010:IMO

[Wei10] Christian H. Weiß. The INARCH(1) model for overdispersed time series of counts. Communications in Statistics: Simulation and Computation, 39(6):1269–1291, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wei:2011:EVC

[Wei11] Chuan-Hua Wei. Estimation in varying-coefficient errors-in-variables models with missing response variables. Communications in Statistics: Simulation and Computation, 40(3): 383–393, 2011. CODEN CSSCDB. ISSN 0361-0918.

Wei:2019:SSS

[Wei19] Fengrong Wei. Sparse structure selection and estimation. Communications in Statistics: Simulation and Computation, 48(4):1229–1238, 2019. CODEN CSSCDB. ISSN 0361-0918.

Westgate:2014:CUT

[Wes14] Philip M. Westgate. A comparison of utilized and theoretical covariance weighting matrices on the estimation performance of quadratic inference functions. Communications in Statistics: Simulation and Computation, 43(10):2432–2443, 2014. CODEN CSSCDB. ISSN 0361-0918.

West:2017:PSC

[Wes17] P. W. West. Population structure and correlation between auxiliary and target variables may affect precision of estimates in forest inventory. Communications in Statistics: Simulation and Computation, 46(6):4951–4965, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2017:PWT

[WF17] Yuh-Jenn Wu and Wei-Quan Fang. Performance of Waldtype estimator for parametric component in partial linear regression with a mixture of Berkson and classical error models. Communications in Statistics: Simulation and Computation, 46(5):3444–3457, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wu:2010:EST

[WGS10] Chih-Chieh Wu, Roger C. Grimson, and Sanjay Shete. Exact statistical tests for heterogeneity of frequencies based on extreme values. Communications in Statistics: Simulation and Computation, 39(3):612–623, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wang:2018:RSA

[WGX18] Ronghua Wang, Beiqing Gu, and Xiaoling Xu. Reliability statistical analysis about products of Birnbaum–Saunders fatigue life distribution for life test and progressive stress accelerated life test. Communications in Statistics: Simulation and Computation, 47(8):2304–2331, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wei:2017:DGE

[WH17] Yonghua Wei and Gabriel Huerta. Dynamic generalized extreme value modeling via particle filters. Communications in Statistics: Simulation and Computation, 46(8):6324–6341, 2017. CODEN CSSCDB. ISSN 0361-0918.

Whitfield:2018:CSE

[Whi18] Paul H. Whitfield. Clustering of seasonal events: a simulation study using circular methods. Communications in Statistics: Simulation and Computation, 47(10):3008–3030, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wiwatwattana:2017:WPC

[WHK17] N. Wiwatwattana, A. J. Hayter, and S. Kiatsupaibul. Winprobabilities for comparing two binary outcomes. Communications in Statistics: Simulation and Computation, 46(1): 204–214, 2017. CODEN CSSCDB. ISSN 0361-0918.

Williamson:2012:UAD

[Wil12] Patricia Pepple Williamson. Usefulness of asymptotic distributions in the classical occupancy problem. Communications in Statistics: Simulation and Computation, 41(8): 1501–1517, 2012. CODEN CSSCDB. ISSN 0361-0918.

Williamson:2015:EPS

[Wil15] Patricia Pepple Williamson. Estimation of the population size in the classical occupancy problem. Communications in Statistics: Simulation and Computation, 44(4):851–867, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wilcox:2017:GCM

[Wil17] Rand R. Wilcox. Global comparisons of medians and other quantiles in a one-way design when there are tied values. Communications in Statistics: Simulation and Computation, 46(4):3010–3019, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wagh:2017:CME

[WK17] Y. S. Wagh and K. K. Kamalja. Comparison of methods of estimation for parameters of generalized Poisson distribution through simulation study. Communications in Statistics: Simulation and Computation, 46(5):4098–4112, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wagh:2018:ZIM

[WK18] Yogita S. Wagh and Kirtee K. Kamalja. Zero-inflated models and estimation in zero-inflated Poisson distribution. Communications in Statistics: Simulation and Computation, 47 (8):2248–2265, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wang:2010:DDP

[WL10] Liuxia Wang and Yulin Li. Detecting the dimensionality for principal components model. Communications in Statistics: Simulation and Computation, 39(6):1073–1082, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wang:2011:BIP

[WL11] Yan-Ping Wang and Zhen-Zhou Lu. Bayesian inference and prediction analysis of the power law process based on a gamma prior distribution. Communications in Statistics: Simulation and Computation, 40(9):1383–1401, 2011. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2014:NDT

[WL14] Yan-Dong Wu and Xu-Qing Liu. Non-diagonal-type estimator in linear regression. Communications in Statistics: Simulation and Computation, 43(1):142–150, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wu:2015:OAC

[WL15a] Shu-Fei Wu and Ying-Po Lin. The optimal allocation combination for the one-sided sequential screening procedure based on the individual misclassification error. Communications in Statistics: Simulation and Computation, 44(4): 833–850, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wu:2015:CSM

[WL15b] Xionghua Wilson Wu and Dejian Lai. Comparison of statistical methods for pretest-posttest designs in terms of Type I error probability and statistical power. Communications in Statistics: Simulation and Computation, 44(2):284–294, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wang:2019:SOD

[WLF19] Qianru Wang, Junxia Liao, and Shengjun Fan. L^{p}(1 < p ≤ 2) solutions of one-dimensional BSDEs whose generator is weakly monotonic in y and non-Lipschitz in z. Communications in Statistics: Simulation and Computation, 48(7): 2099–2109, 2019. CODEN CSSCDB. ISSN 0361-0918.

Wu:2010:MOS

[WLL10] Shu-Fei Wu, Ying-Po Lin, and Huei-Jiuan Lin. A modified one-sided sequential screening procedure based on individual misclassification error. Communications in Statistics: Simulation and Computation, 39(9):1754–1778, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wu:2016:CFD

[WLWT16] Berlin Wu, Wentsung Lai, Ching-Ling Wu, and Tsung-Kuo Tienliu. Correlation with fuzzy data and its applications in the 12-year compulsory education in Taiwan. Communications in Statistics: Simulation and Computation, 45(4): 1337–1354, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wang:2018:VBS

[WLX18] Pan Wang, Zhenzhou Lu, and Sinan Xiao. Variance-based sensitivity analysis with the uncertainties of the input variables and their distribution parameters. Communications in Statistics: Simulation and Computation, 47(4):1103–1125, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wang:2019:MCD

[WLZL19] Tao Wang, Qun Li, Qingpei Zang, and Zhonghua Li. A multiple-case deletion approach for detecting influential points in high-dimensional regression. Communications in Statistics: Simulation and Computation, 48(7):2065–2082, 2019. CODEN CSSCDB. ISSN 0361-0918.

Watcharotone:2017:MCS

[WMH17] Kuanwong Watcharotone, Joseph W. McKean, and Bradley E. Huitema. A Monte Carlo study of traditional and rank-based picked-points analyses. Communications in Statistics: Simulation and Computation, 46(5):4050–4066, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wu:2014:BVS

[WMK14] Jin Chu Wu, Alvin F. Martin, and Raghu N. Kacker. Bootstrap variability studies in ROC analysis on large datasets. Communications in Statistics: Simulation and Computation, 43(1):225–236, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wu:2016:VNT

[WMK16] Jin Chu Wu, Alvin F. Martin, and Raghu N. Kacker. Validation of nonparametric two-sample bootstrap in ROC analysis on large datasets. Communications in Statistics: Simulation and Computation, 45(5):1689–1703, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wu:2019:MCS

[WMK19] Jin Chu Wu, Alvin F. Martin, and Raghu N. Kacker. Monte Carlo studies of bootstrap variability in ROC analysis with data dependency. Communications in Statistics: Simulation and Computation, 48(2):317–333, 2019. CODEN CSSCDB. ISSN 0361-0918.

Withers:2015:CIN

[WN15a] Christopher S. Withers and Saralees Nadarajah. Confidence intervals for the normal multiple correlation. Communications in Statistics: Simulation and Computation, 44(8): 2003–2022, 2015. CODEN CSSCDB. ISSN 0361-0918.

Withers:2015:IEC

[WN15b] Christopher S. Withers and Saralees Nadarajah. Improved estimates and confidence intervals for regenerative simulation. Communications in Statistics: Simulation and Computation, 44(7):1801–1817, 2015. CODEN CSSCDB. ISSN 0361-0918.

Withers:2017:NCR

[WN17] Christopher S. Withers and Saralees Nadarajah. A new confidence region for the multinomial distribution. Communications in Statistics: Simulation and Computation, 46(5): 4113–4126, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wong:2017:PCM

[WODY17] George Y. C. Wong, Michael P. Osborne, Qinggang Diao, and Qiqing Yu. Piecewise Cox models with right-censored data. Communications in Statistics: Simulation and Computation, 46(10):7894–7908, 2017. CODEN CSSCDB. ISSN 0361-0918.

Weiss:2013:SPA

[WP13] Christian H. Weiß and Miia Peltola. Sequential pattern analysis: A statistical investigation of sequence length and support. Communications in Statistics: Simulation and Computation, 42(5):1044–1062, 2013. CODEN CSSCDB. ISSN 0361-0918.

Wang:2015:STS

[WPLL15] Lei Wang, Xiaolong Pu, Yan Li, and Yukun Liu. Sequential two-stage D-optimality sensitivity test for binary response data. Communications in Statistics: Simulation and Computation, 44(7):1833–1849, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wu:2017:LCB

[WPTL17] Chia Huang Wu, Wen Lea Pearn, Yu Ting Tai, and Pi Chuan Lin. Lower confidence bounds as precision measure for truncated processes. Communications in Statistics: Simulation and Computation, 46(2):1461–1480, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Wan:2018:SSA

[WSCW18] Shu-Mei Wan, Tien-Yu Sun, Kai-Hui Chen, and Chien-Hua Wu. Sample size algorithm for randomized response model with dichotomized response. Communications in Statistics: Simulation and Computation, 47(2):477–499, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wang:2015:SHG

[WSQ15] Shuai Wang, Yang Shen, and Linyi Qian. Static hedging of geometric average Asian options with standard options. Communications in Statistics: Simulation and Computation, 44(8):2101–2116, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wang:2016:QAS

[WSW16] Mingqiu Wang, Lixin Song, and Xiaoguang Wang. Quadratic approximation via the SCAD penalty with a diverging number of parameters. Communications in Statistics: Simulation and Computation, 45(1):1–16, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wang:2018:UPM

[WSX18] Junjie Wang, Qin Su, and Min Xie. A univariate procedure for monitoring location and dispersion with ordered categorical data. Communications in Statistics: Simulation and Computation, 47(1):115–128, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wu:2017:SAD

[WSZ17] Min Wu, Yimin Shi, and Chunfang Zhang. Statistical analysis of dependent competing risks model in accelerated life testing under progressively hybrid censoring using copula function. Communications in Statistics: Simulation and Computation, 46(5):4004–4017, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wang:2010:DEG

[WTCZ10] Xiaodi Wang, Yincai Tang, Xueping Chen, and Yingshan Zhang. Design of experiment in global sensitivity analysis based on ANOVA high-dimensional model representation. Communications in Statistics: Simulation and Computation, 39(6):1183–1195, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wu:2016:SAS

[WTLWL16] Berlin Wu, Tsung-Kuo Tien-Liu, Ching-Ling Wu, and Wentsung Lai. Statistical analysis with soft computation for fuzzy answering in sampling survey. Communications in Statistics: Simulation and Computation, 45(4):1295–1306, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wang:2011:OAE

[WTZ11] Xiaodi Wang, Yincai Tang, and Yingshan Zhang. Orthogonal arrays for the estimation of global sensitivity indices based on ANOVA high-dimensional model representation. Communications in Statistics: Simulation and Computation, 40(9):1324–1341, 2011. CODEN CSSCDB. ISSN 0361-0918.

Wu:2010:LLS

[Wu10] Shaowen Wu. Lag length selection in DF-GLS unit root tests. Communications in Statistics: Simulation and Computation, 39(8):1590–1604, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wu:2014:VSJ

[Wu14] Liu-Cang Wu. Variable selection in joint location and scale models of the skew-t-normal distribution. Communications in Statistics: Simulation and Computation, 43(3):615–630, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wu:2016:MRA

[Wu16] Jibo Wu. Modified restricted almost unbiased Liu estimator in linear regression model. Communications in Statistics: Simulation and Computation, 45(2):689–700, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wu:2017:PTL

[Wu17a] Jibo Wu. Preliminary test Liu-type estimators based on W, LR, and LM test statistics in a regression model. Communications in Statistics: Simulation and Computation, 46(9): 6760–6771, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wu:2017:MCE

[Wu17b] Shu-Fei Wu. Multiple comparisons of exponential location parameters with a control based on doubly censored sample under heteroscedasticity. Communications in Statistics: Simulation and Computation, 46(3):1858–1870, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2018:MCA

[Wu18a] Shu-Fei Wu. Multiple comparisons with the average for exponential location parameters based on doubly censored sample under heteroscedasticity. Communications in Statistics: Simulation and Computation, 47(5):1529–1539, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wu:2018:OSM

[Wu18b] Shu-Fei Wu. One stage multiple comparisons of k − 1 treatment mean lifetimes with the control for exponential distributions under heteroscedasticity. Communications in Statistics: Simulation and Computation, 47(10):2968–2978, 2018. CODEN CSSCDB. ISSN 0361-0918.

Wu:2019:MMC

[Wu19] Shu-Fei Wu. A modified multiple comparisons with a control for exponential location parameters based on doubly censored sample under heteroscedasticity. Communications in Statistics: Simulation and Computation, 48(7):2056–2064, 2019. CODEN CSSCDB. ISSN 0361-0918.

Wang:2010:WEU

[WW10] Xue Wang and Andrew T. A. Wood. Wavelet estimation of an unknown function observed with correlated noise. Communications in Statistics: Simulation and Computation, 39 (2):287–304, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wang:2017:BCM

[WW17] Min Wang and Wentao Wang. Bias-corrected maximum likelihood estimation of the parameters of the weighted Lindley distribution. Communications in Statistics: Simulation and Computation, 46(1):530–545, 2017. CODEN CSSCDB. ISSN 0361-0918.

Wu:2014:CIE

[WWJ14] Jianrong Wu, Augustine Wong, and Guoyong Jiang. Confidence intervals of effect size for randomized comparative parallel-group studies with unequal variances. Communications in Statistics: Simulation and Computation, 43(7): 1662–1674, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wei:2015:EEH

[WWL15] Chuan-Hua Wei, Li-Jie Wan, and Chun-Ling Liu. Efficient estimation in heteroscedastic partially linear varying coefficient models. Communications in Statistics: Simulation and Computation, 44(4):892–901, 2015. CODEN CSSCDB. ISSN 0361-0918.

Wei:2019:IMC

[WWWS19] Yuan Wei, Huiwen Wang, Shanshan Wang, and Gilbert Saporta. Incremental modelling for compositional data streams. Communications in Statistics: Simulation and Computation, 48(8):2229–2243, 2019. CODEN CSSCDB. ISSN 0361-0918.

Wu:2019:OES

[WX19] Jian Wu and Liugen Xue. Oracally efficient spline-backfitted kernel smoothing of additive partial linear measurement error model. Communications in Statistics: Simulation and Computation, 48(10):2985–2997, 2019. CODEN CSSCDB. ISSN 0361-0918.

Wong:2012:EUL

[WY12] George Y. C. Wong and Qiqing Yu. Estimation under the Lehmann regression model with interval-censored data. Communications in Statistics: Simulation and Computation, 41(8):1489–1500, 2012. CODEN CSSCDB. ISSN 0361-0918.

Wu:2014:SRA

[WY14] Jibo Wu and Hu Yang. On the stochastic restricted almost unbiased estimators in linear regression model. Communications in Statistics: Simulation and Computation, 43(2): 428–440, 2014. CODEN CSSCDB. ISSN 0361-0918.

Wu:2015:PCL

[WY15a] Jibo Wu and Hu Yang. On the principal component Liu-type estimator in linear regression. Communications in Statistics: Simulation and Computation, 44(8):2061–2072, 2015. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2015:CVN

[WY15b] Shaowen Wu and Yong Yin. Combining p-values in nonstationary panels. Communications in Statistics: Simulation and Computation, 44(6):1412–1431, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Wu:2017:RMM

[WY17] Qiang Wu and Weixin Yao. Relabel mixture models via modal clustering. Communications in Statistics: Simulation and Computation, 46(5):3406–3418, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Wang:2019:WTL

[WY19] Leyang Wang and Hang Yu. Weighted total least-squares joint adjustment with weight correction factors. Communications in Statistics: Simulation and Computation, 48(9): 2689–2707, 2019. CODEN CSSCDB. ISSN 0361-0918.

Wang:2010:GPS

[WZ10] Dehui Wang and Haixiang Zhang. Generalized RCINAR(p) process with signed thinning operator. Communications in Statistics: Simulation and Computation, 40(1):13–44, 2010. CODEN CSSCDB. ISSN 0361-0918.

Wang:2016:REL

[WZCJ16] Liying Wang, Jie Zhang, Wei Chen, and Xujie Jia. Reliability evaluation of a load-sharing parallel system with failure dependence. Communications in Statistics: Simulation and Computation, 45(9):3094–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Wang:2015:IER

[WZG15] Guan-Wei Wang, Chun-Xia Zhang, and Gao Guo. Investigating the effect of randomly selected feature subsets on bagging and boosting. Communications in Statistics: Simulation and Computation, 44(3):636–646, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Wang:2017:TSM

[WZT17] Xiaodi Wang, Yingshan Zhang, and Yincai Tang. A twostage method for the estimation of global sensitivity indices of non-parametric models. Communications in Statistics: Simulation and Computation, 46(4):2957–2975, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Wu:2016:RVS

[WZTX16] Liu-Cang Wu, Zhong-Zhan Zhang, Guo-Liang Tian, and Deng-Ke Xu. A robust variable selection to t-type joint generalized linear models via penalized t-type pseudo-likelihood. Communications in Statistics: Simulation and Computation, 45(7):2320–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Xu:2015:TAO

[XC15] Chen Xu and Jiahua Chen. A thresholding algorithm for order selection in finite mixture models. Communications in Statistics: Simulation and Computation, 44(2):433–453, 2015. CODEN CSSCDB. ISSN 0361-0918.

Xiao:2018:EIV

[XC18] Yanting Xiao and Zhanshou Chen. Estimation and inference for varying coefficient partially nonlinear errors-in-variables models. Communications in Statistics: Simulation and Computation, 47(7):2023–2039, 2018. CODEN CSS-CDB. ISSN 0361-0918.

Xu:2018:SLQ

[XCJ^{+}18] Qifa Xu, Chao Cai, Cuixia Jiang, Fang Sun, and Xue Huang. Sampling Lasso quantile regression for large-scale data. Communications in Statistics: Simulation and Computation, 47(1):92–114, 2018. CODEN CSSCDB. ISSN 0361-0918.

Xiong:2015:FNT

[XD15] Wenjun Xiong and Juan Ding. A flexible nonparametric test for ordered and umbrella alternatives. Communications in Statistics: Simulation and Computation, 44(5):1257–1263, 2015. CODEN CSSCDB. ISSN 0361-0918.

Xiong:2017:CIP

[XD17] Wenjun Xiong and Juan Ding. Confidence intervals for a population proportion using group testing in the presence of misclassification. Communications in Statistics: Simulation and Computation, 46(1):174–183, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xampeny:2018:SSE

[XGTM18] Rafel Xampeny, Pere Grima, and Xavier Tort-Martorell. Selecting significant effects in factorial designs: Lenth’s method versus using negligible interactions. Communications in Statistics: Simulation and Computation, 47(5): 1343–1352, 2018. CODEN CSSCDB. ISSN 0361-0918.

Xin:2018:WMT

[XHD18] Xin Xin, Julie Horrocks, and Gerarda A. Darlington. Weighting methods for ties between event times and covariate change times. Communications in Statistics: Simulation and Computation, 47(1):1–15, 2018. CODEN CSSCDB. ISSN 0361-0918.

Xiang:2015:BDF

[Xia15] Jim X. Xiang. A bivariate distribution family with specified correlation coefficient and given marginals. Communications in Statistics: Simulation and Computation, 44(9): 2463–2472, 2015. CODEN CSSCDB. ISSN 0361-0918.

Xu:2018:FVA

[XL18] Jing Xu and Xinmin Li. A fiducial p -value approach for comparing heteroscedastic regression models. Communications in Statistics: Simulation and Computation, 47(2):420–431, 2018. CODEN CSSCDB. ISSN 0361-0918.

Xu:2014:ABO

[XLC14] Ying Xu, Chun Fan Lee, and Yin Bun Cheung. Analyzing binary outcome data with small clusters: A simulation study. Communications in Statistics: Simulation and Computation, 43(7):1771–1782, 2014. CODEN CSSCDB. ISSN 0361-0918.

Xie:2013:WMT

[XLPC13] Changchun Xie, Xuewen Lu, Janice Pogue, and Ding-Geng (Din) Chen. Weighted multiple testing correction for correlated binary endpoints. Communications in Statistics: Simulation and Computation, 42(8):1693–1702, 2013. CO-DEN CSSCDB. ISSN 0361-0918.

Xie:2017:STH

[XLW17] Feng-Chang Xie, Jin-Guan Lin, and Bo-Cheng Wei. Score test for homogeneity of dispersion in generalized Poisson mixed models with excess zeros. Communications in Statistics: Simulation and Computation, 46(1):301–314, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xu:2019:RPM

[XLX^{+}19] Cong Xu, Zheng Li, Yuan Xue, Lijun Zhang, and Ming Wang. An R package for model fitting, model selection and the simulation for longitudinal data with dropout missingness. Communications in Statistics: Simulation and Computation, 48(9):2812–2829, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xu:2019:PBP

[XP19] Jiaqing Xu and Cheng Peng. Parametric bootstrap process capability index control charts for both mean and dispersion. Communications in Statistics: Simulation and Computation, 48(10):2936–2954, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xu:2016:PBT

[XQW^{+}16] Liwen Xu, Kaiyi Qu, Mixia Wu, Bo Mei, and Ranran Chen. Parametric bootstrap tests for unbalanced threefactor nested designs under heteroscedasticity. Communications in Statistics: Simulation and Computation, 45(1): 322–338, 2016. CODEN CSSCDB. ISSN 0361-0918.

Xue:2011:AHA

[XS11] Wei Xue and Peng Shen. Application of HITS algorithm in Web mining in literature evaluation and empirical study. Communications in Statistics: Simulation and Computation, 40(10):1576–1586, 2011. CODEN CSSCDB. ISSN 0361-0918.

Xu:2017:MFT

[XSA^{+}17] Jianbo Xu, Guogen Shan, Amei Amei, Jiwei Zhao, Daniel Young, and Sheila Clark. A modified Friedman test for randomized complete block designs. Communications in Statistics: Simulation and Computation, 46(2):1508–1519, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xiang:2011:NTZ

[XT11] Liming Xiang and Guo Shou Teo. A note on tests for zero-inflation in correlated count data. Communications in Statistics: Simulation and Computation, 40(7):992–1005, 2011. CODEN CSSCDB. ISSN 0361-0918.

Xu:2019:BAL

[XT19] Dengke Xu and Niansheng Tang. Bayesian adaptive lasso for quantile regression models with nonignorably missing response data. Communications in Statistics: Simulation and Computation, 48(9):2727–2742, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xu:2014:BAM

[XTG14] Ancha Xu, Yincai Tang, and Qiang Guan. Bayesian analysis of masked data in step-stress accelerated life testing. Communications in Statistics: Simulation and Computation, 43 (8):2016–2030, 2014. CODEN CSSCDB. ISSN 0361-0918.

Xu:2010:EES

[Xu10] Dinghai Xu. An efficient estimation for switching regression models: A Monte Carlo study. Communications in Statistics: Simulation and Computation, 39(7):1403–1421, 2010. CODEN CSSCDB. ISSN 0361-0918.

Xu:2014:AKD

[Xu14] Sun Xu. Asymmetric kernel density estimation based on grouped data with applications to loss model. Communications in Statistics: Simulation and Computation, 43(3): 657–672, 2014. CODEN CSSCDB. ISSN 0361-0918.

Xuemei:2017:RER

[Xue17] Hu Xuemei. Robust estimators and robust tests for the slightly contaminated stochastic logistic population models. Communications in Statistics: Simulation and Computation, 46(4):2756–2768, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xu:2017:PBI

[XW17] Liwen Xu and Dengkui Wang. Parametric bootstrap inferences for unbalanced panel data models. Communications in Statistics: Simulation and Computation, 46(10):7602–7613, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xiong:2016:CIS

[XX16] Cui Xiong and Jin Xu. Confidence intervals from stochastic approximation. Communications in Statistics: Simulation and Computation, 45(6):1827–1837, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Xu:2012:PTB

[XY12] Jianwen Xu and Hu Yang. On the preliminary test backfitting and Speckman estimators in partially linear models and numerical comparisons. Communications in Statistics: Simulation and Computation, 41(3):327–341, 2012. CODEN CSSCDB. ISSN 0361-0918.

Xu:2016:PBS

[XY16] Liwen Xu and Lingli Yuan. A parametric bootstrap solution to two-way MANOVA without interaction under heteroscedasticity: Fixed and mixed models. Communications in Statistics: Simulation and Computation, 45(8):2888–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Xing:2017:USC

[XY17a] Guo-Dong Xing and Shan-Chao Yang. Uniformly strong consistency of frequency polygons for negatively associated samples. Communications in Statistics: Simulation and Computation, 46(3):2168–2175, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xing:2017:LTA

[XY17b] Yu Xing and Xiaoping Yang. Laplace transform approach to option pricing for time-changed Brownian models. Communications in Statistics: Simulation and Computation, 46 (3):2121–2137, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xu:2015:PBT

[XYCY15] Liwen Xu, Fangqin Yang, Ranran Chen, and Shenghua Yu. A parametric bootstrap test for two-way ANOVA model without interaction under heteroscedasticity. Communications in Statistics: Simulation and Computation, 44(5): 1264–1272, 2015. CODEN CSSCDB. ISSN 0361-0918.

Xing:2019:USC

[XYL19] Guo-Dong Xing, Shan-Chao Yang, and Xiaohu Li. Uniformly strong consistency and Berry–Esseen bound of frequency polygons for α-mixing samples. Communications in Statistics: Simulation and Computation, 48(2):416–430, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xu:2017:PSM

[XYZ17] Jianwen Xu, Qianlin Yang, and Jiamao Zhang. The performances of several modified CIC criteria for working intracluster correlation structure selection in GEE analysis. Communications in Statistics: Simulation and Computation, 46 (8):6034–6048, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xue:2016:DID

[XZ16] Wei Xue and Jing Zhang. Dealing with imbalanced dataset: A re-sampling method based on the improved SMOTE algorithm. Communications in Statistics: Simulation and Computation, 45(4):1160–1172, 2016. CODEN CSSCDB. ISSN 0361-0918.

Xu:2017:BIJ

[XZ17] Dengke Xu and Zhongzhan Zhang. Bayesian inference for joint location and scale nonlinear models with skew-normal errors. Communications in Statistics: Simulation and Computation, 46(1):619–630, 2017. CODEN CSSCDB. ISSN 0361-0918.

Xu:2019:MAA

[XZ19] Liwen Xu and Jiabin Zhou. A model-averaging approach for smoothing spline regression. Communications in Statistics: Simulation and Computation, 48(8):2438–2451, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xu:2019:VSG

[XZF19] Jianwen Xu, Jiamao Zhang, and Liya Fu. Variable selection in generalized estimating equations via empirical likelihood and Gaussian pseudo-likelihood. Communications in Statistics: Simulation and Computation, 48(4):1239–1250, 2019. CODEN CSSCDB. ISSN 0361-0918.

Xiao:2015:PID

[XZZ15] Weilin Xiao, Weiguo Zhang, and Xili Zhang. Parameter identification for drift fractional Brownian motions with application to the Chinese stock markets. Communications in Statistics: Simulation and Computation, 44(8):2117–2136, 2015. CODEN CSSCDB. ISSN 0361-0918.

Youssef:2017:AGE

[YA17] Ahmed H. Youssef and Mohamed R. Abonazel. Alternative GMM estimators for first-order autoregressive panel model: An improving efficiency approach. Communications in Statistics: Simulation and Computation, 46(4): 3112–3128, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yuzbasi:2019:DSS

[YA19] B. Yüzbasi and M. Arashi. Double shrunken selection operator. Communications in Statistics: Simulation and Computation, 48(3):666–674, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yasmin:2015:CBI

[YAA15] Fariha Yasmin, Rashid Ahmed, and Munir Akhtar. Construction of balanced incomplete block designs using cyclic shifts. Communications in Statistics: Simulation and Computation, 44(2):525–532, 2015. CODEN CSSCDB. ISSN 0361-0918.

Yamamura:2016:EPA

[Yam16] Kohji Yamamura. Estimation of the predictive ability of ecological models. Communications in Statistics: Simulation and Computation, 45(6):2122–2144, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Yang:2013:USA

[Yan13] Su-Fen Yang. Using a single average loss control chart to monitor process mean and variability. Communications in Statistics: Simulation and Computation, 42(7):1549–1562, 2013. CODEN CSSCDB. ISSN 0361-0918.

Yang:2016:IDF

[Yan16] Su-Fen Yang. An improved distribution-free EWMA mean chart. Communications in Statistics: Simulation and Computation, 45(4):1410–1427, 2016. CODEN CSSCDB. ISSN 0361-0918.

Yao:2013:SSB

[Yao13] Weixin Yao. A simple solution to Bayesian mixture labeling. Communications in Statistics: Simulation and Computation, 42(4):800–813, 2013. CODEN CSSCDB. ISSN 0361-0918.

Yao:2018:CME

[Yao18] Lihua Yao. Comparing methods for estimating the abilities for the multidimensional models of mixed item types. Communications in Statistics: Simulation and Computation, 47 (1):74–91, 2018. CODEN CSSCDB. ISSN 0361-0918.

Yarovaya:2012:SBW

[Yar12] Elena Yarovaya. Symmetric branching walks in homogeneous and non homogeneous random environments. Communications in Statistics: Simulation and Computation, 41 (7):1232–1249, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yousof:2019:NLM

[YAR^{+}19] Haitham M. Yousof, Emrah Altun, Mahdi Rasekhi, Morad Alizadeh, G. G. Hamedani, and M. Masoom Ali. A new lifetime model with regression models, characterizations and applications. Communications in Statistics: Simulation and Computation, 48(1):264–286, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yang:2019:LPB

[YAT19] Jun Yang, Fady Alajaji, and Glen Takahara. Linear programming bounds on the union probability. Communications in Statistics: Simulation and Computation, 48(9): 2845–2854, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yoo:2012:GPD

[YBW^{+}12] Changwon Yoo, Erik M. Brilz, Meredith Wilcox, Mark A. Pershouse, and Elizabeth A. Putnam. Gene pathways discovery in asbestos-related diseases using local causal discovery algorithm. Communications in Statistics: Simulation and Computation, 41(10):1840–1859, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Yeh:2010:EBR

[YC10] Li Lun Yeh and Fei Long Chen. An extension of Banerjee and Rahim’s model for an economic X-bar control chart for non-normally distributed data under gamma failure models. Communications in Statistics: Simulation and Computation, 39(5):994–1015, 2010. CODEN CSSCDB. ISSN 0361-0918.

Yue:2011:EMS

[YC11] Jack C. Yue and Murray K. Clayton. The effect of migration on a similarity index. Communications in Statistics: Simulation and Computation, 40(3):412–423, 2011. CODEN CSSCDB. ISSN 0361-0918.

Yue:2016:MCO

[YC16] Jack C. Yue and Murray K. Clayton. A multiple-community overlap measure based on a probabilistic approach. Communications in Statistics: Simulation and Computation, 45(4): 1307–1317, 2016. CODEN CSSCDB. ISSN 0361-0918.

Yan:2016:MEM

[YCH16] Chunning Yan, Ren Chen, and Yangxin Huang. Mixedeffects models with skewed distributions for time-varying decay rate in HIV dynamics. Communications in Statistics: Simulation and Computation, 45(2):737–757, 2016. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:OBV

[YCL17] Ming-Chung Yang, Lee-Shen Chen, and Tachen Liang. Optimal Bayesian variable sampling plans for exponential distributions based on modified type-II hybrid censored samples. Communications in Statistics: Simulation and Computation, 46(6):4722–4744, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yeh:2012:IMO

[YCS12] Hung-Wen Yeh, Wenyaw Chan, and Elaine Symanski. Intermittent missing observations in discrete-time hidden Markov models. Communications in Statistics: Simulation and Computation, 41(2):167–181, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:VRE

[YCS17] Xinxia Yang, Ratthachat Chatpatanasiri, and Pairote Sattayatham. Value at risk estimation under stochastic volatility models using adaptive PMCMC methods. Communications in Statistics: Simulation and Computation, 46(9): 7221–7237, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yeh:2010:ETP

[YCSD10] Hung-Wen Yeh, Wenyaw Chan, Elaine Symanski, and Barry R. Davis. Estimating transition probabilities for ignorable intermittent missing data in a discrete-time Markov chain. Communications in Statistics: Simulation and Computation, 39(2):433–448, 2010. CODEN CSSCDB. ISSN 0361-0918.

Yu:2019:JMD

[YCSP19] Xiaoying Yu, Wenyaw Chan, Michael D. Swartz, and Linda B. Piller. Joint models of dynamics of mothers’ stress and children’s disease. Communications in Statistics: Simulation and Computation, 48(9):2775–2790, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yuksel:2019:FCO

[YDA19] Selcen Yüksel, Pervin Demir, and Afra Alkan. Factors causing occurrence of artificial dif: a simulation study for dichotomous data. Communications in Statistics: Simulation and Computation, 48(7):2004–2011, 2019. CODEN CSS-CDB. ISSN 0361-0918.

Yang:2017:BAG

[YE17] Szu-Peng Yang and Takeshi Emura. A Bayesian approach with generalized ridge estimation for high-dimensional regression and testing. Communications in Statistics: Simulation and Computation, 46(8):6083–6105, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yeh:2017:FGM

[Yeh17] Hsiaw-Chan Yeh. Four general multivariate stationary extremal Markovian processes. Communications in Statistics: Simulation and Computation, 46(1):16–37, 2017. CODEN CSSCDB. ISSN 0361-0918.

Young:2012:CLE

[YGL12] Linda J. Young, Carol A. Gotway, and Kenneth K. Lopiano. A closer look at errors in variables methods for use in regression models with spatially misaligned data. Communications in Statistics: Simulation and Computation, 41(7): 1250–1269, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yang:2016:TSC

[YGW16] Jin Yang, Shuang Guo, and Chuanhua Wei. Testing serial correlation in partially linear additive errors-in-variables models. Communications in Statistics: Simulation and Computation, 45(9):3114–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Yu:2019:OCC

[YH19] Doyeong Yu and Wei-Ting Hwang. Optimal cutoffs for continuous biomarkers for survival data under competing risks. Communications in Statistics: Simulation and Computation, 48(5):1330–1345, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ye:2012:IAG

[YHL12] Ren-Dao Ye, Yu-Qin Hu, and Kun Luo. Inferences on the among-group variance component in unbalanced heteroscedastic one-fold nested design. Communications in Statistics: Simulation and Computation, 41(3):391–404, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yildiz:2017:WML

[Yil17] Nilgun Yildiz. On the weighted mixed Liu-type estimator under unbiased stochastic restrictions. Communications in Statistics: Simulation and Computation, 46(9):7238–7248, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yildiz:2019:NSR

[Yil19] Nilgün Yildiz. A new stochastic restricted Liu-type estimator in linear regression model. Communications in Statistics: Simulation and Computation, 48(1):91–108, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yeong:2017:DPM

[YKLL17] Wai Chung Yeong, Michael B. C. Khoo, Sok Li Lim, and Ming Ha Lee. A direct procedure for monitoring the coefficient of variation using a variable sample size scheme. Communications in Statistics: Simulation and Computation, 46 (6):4210–4225, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yeong:2014:EOD

[YKLR14] Wai Chung Yeong, Michael B. C. Khoo, M. H. Lee, and M. A. Rahim. Economically optimal design of a multivariate synthetic T^{2} chart. Communications in Statistics: Simulation and Computation, 43(6):1333–1361, 2014. CODEN CSSCDB. ISSN 0361-0918.

Yeong:2016:SAE

[YKSC16] Wai Chung Yeong, Michael B. C. Khoo, M. Shamsuzzaman, and Philippe Castagliola. Sensitivity analyses of the economic and economic statistical designs of the multivariate synthetic T^{2} chart. Communications in Statistics: Simulation and Computation, 45(7):2541–??, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Yang:2010:UVL

[YKY10] Su-Fen Yang, Chih-Ying Ko, and Jin-Tyan Yeh. Using VSI loss control charts to monitor a process with incorrect adjustment. Communications in Statistics: Simulation and Computation, 39(4):736–749, 2010. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:SQI

[YL17] Weiming Yang and Shu Liao. A study of quadratic inference functions with alternative weighting matrices. Communications in Statistics: Simulation and Computation, 46 (2):994–1007, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yan:2017:DMS

[YLA17] Aijun Yan, Sanyang Liu, and Muhammad Azam. Designing a multiple state repetitive group sampling plan based on the coefficient of variation. Communications in Statistics: Simulation and Computation, 46(9):7154–7165, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2016:PLR

[YLG16] Hu Yang, Jing Lv, and Chaohui Guo. Penalized LAD regression for single-index models. Communications in Statistics: Simulation and Computation, 45(7):2392–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Yoo:2011:JMC

[YLK11] Hanna Yoo, Jaewon Lee, and Yang-Jin Kim. Joint model of clustered failure time data with informative cluster size. Communications in Statistics: Simulation and Computation, 40(6):808–817, 2011. CODEN CSSCDB. ISSN 0361-0918.

Yoon:2017:ALL

[YLL17] Young Joo Yoon, Sooyong Lee, and Taewook Lee. Adaptive LASSO for linear regression models with ARMA-GARCH errors. Communications in Statistics: Simulation and Computation, 46(5):3479–3490, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yu:2017:ACC

[YLM17] Tao Yu, Jialiang Li, and Shuangge Ma. Accounting for clinical covariates and interactions in ranking genomic markers using ROC. Communications in Statistics: Simulation and Computation, 46(5):3735–3755, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2015:OFD

[YM15] Jie Yang and Abhyuday Mandal. D-optimal factorial designs under generalized linear models. Communications in Statistics: Simulation and Computation, 44(9):2264–2277, 2015. CODEN CSSCDB. ISSN 0361-0918.

Yamashita:2017:EPP

[YMS17] Haruka Yamashita, Shun Matsuura, and Hideo Suzuki. Estimation of principal points for a multivariate binary distribution using a log-linear model. Communications in Statistics: Simulation and Computation, 46(2):1136–1147, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Yoshimori:2015:NCB

[Yos15] Masayo Yoshimori. Numerical comparison between different empirical prediction intervals under the Fay–Herriot model. Communications in Statistics: Simulation and Computation, 44(5):1158–1170, 2015. CODEN CSSCDB. ISSN 0361-

Young:2013:RTI

[You13] Derek S. Young. Regression tolerance intervals. Communications in Statistics: Simulation and Computation, 42(9): 2040–2055, 2013. CODEN CSSCDB. ISSN 0361-0918.

Yun:2014:ORM

[YPH14] Won Young Yun, Goeun Park, and Young Jin Han. An optimal reliability and maintainability design of a searching system. Communications in Statistics: Simulation and Computation, 43(8):1959–1978, 2014. CODEN CSSCDB. ISSN 0361-0918.

Yuan:2011:CIS

[YR11] Xiaobin Yuan and Shesh N. Rai. Confidence intervals for survival probabilities: A comparison study. Communications in Statistics: Simulation and Computation, 40(7):978–991, 2011. CODEN CSSCDB. ISSN 0361-0918.

Yari:2012:UWM

[YS12] Gh. Yari and A. Saghafi. Unbiased Weibull modulus estimation using differential cumulative entropy. Communications in Statistics: Simulation and Computation, 41(8): 1372–1378, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yang:2019:PMS

[YSC19] Jingyun Yang, Mingzhi Shao, and Guoliang Cai. On the performance of MixTVEM: a simulation study. Communications in Statistics: Simulation and Computation, 48(9): 2830–2844, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yagi:2019:IST

[YSH19] Ayaka Yagi, Takashi Seo, and Zofia Hanusz. Improved simplified T^{2} test statistics for a mean vector with monotone missing data. Communications in Statistics: Simulation and Computation, 48(6):1606–1622, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yue:2017:PBA

[YSS17] Li-Li Yue, Jian-Hong Shi, and Wei-Xing Song. A parametric bootstrap approach for two-way error component regression models. Communications in Statistics: Simulation and Computation, 46(5):3952–3961, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2013:NPF

[YSWM13] Dong-Fang Yang, Fu-Chun Sun, Shi-Cheng Wang, and Hai-Bo Min. A novel particle filter aided by interval analysis approach. Communications in Statistics: Simulation and Computation, 42(8):1887–1905, 2013. CODEN CSSCDB. ISSN 0361-0918.

Yang:2019:GSM

[YSY19] Fengkai Yang, Ang Shan, and Haijing Yuan. Gibbs sampling for mixture quantile regression based on asymmetric Laplace distribution. Communications in Statistics: Simulation and Computation, 48(5):1560–1573, 2019. CODEN CSSCDB. ISSN 0361-0918.

Yu:2012:CAI

[YT12] Ching-Yun Yu and Jia-Ren Tsai. A comparison of approaches to interval estimation of mediated effects in linear regression models. Communications in Statistics: Simulation and Computation, 41(8):1301–1316, 2012. CODEN CSSCDB. ISSN 0361-0918.

Yari:2017:EEP

[YT17] Gholamhossein Yari and Zahra Tondpour. Estimation of the exponential Pareto II distribution parameters. Communications in Statistics: Simulation and Computation, 46(9): 6889–6906, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2016:BHK

[YW16] Hongxia Yang and Jun Wang. Bayesian hierarchical kernelized probabilistic matrix factorization. Communications in Statistics: Simulation and Computation, 45(7):2528–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Ye:2017:TSL

[YW17a] Qian Ye and Lang Wu. Two-step and likelihood methods for joint models of longitudinal and survival data. Communications in Statistics: Simulation and Computation, 46(8): 6019–6033, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yin:2017:ECR

[YW17b] Li Yin and Xiaoqin Wang. Estimating confidence regions of common measures of the baseline and treatment effect on dichotomous outcome of a population. Communications in Statistics: Simulation and Computation, 46(4):3034–3049, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yin:2017:MLE

[YWY17] Li Yin, Xiaoqin Wang, and Weimin Ye. Maximum-likelihood estimation and presentation for the interaction between treatments in observational studies with a dichotomous outcome. Communications in Statistics: Simulation and Computation, 46(9):7138–7153, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2018:GFB

[YWZ18] Xiaona Yang, Zhaojun Wang, and Xuemin Zi. Goodnessof-fit-based outlier detection for Phase I monitoring. Communications in Statistics: Simulation and Computation, 47 (10):2979–2991, 2018. CODEN CSSCDB. ISSN 0361-0918.

Yang:2010:VSP

[YXC10] Yiping Yang, Liugen Xue, and Weihu Cheng. Variable selection for partially linear models with randomly censored data. Communications in Statistics: Simulation and Computation, 39(8):1577–1589, 2010. CODEN CSSCDB. ISSN 0361-0918.

Yang:2014:LEL

[YXS14] Yujiao Yang, Song Xu, and Qiongxia Song. A logspline estimation for a linear regression model with an intervalcensored continuous covariate. Communications in Statistics: Simulation and Computation, 43(10):2521–2539, 2014. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:RFM

[YXY17] Li Yang, Sijia Xiang, and Weixin Yao. Robust fitting of mixtures of factor analyzers using the trimmed likelihood estimator. Communications in Statistics: Simulation and Computation, 46(2):1280–1291, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:NIPa

[YY17a] Fengkai Yang and Haijing Yuan. A non-iterative posterior sampling algorithm for Laplace linear regression model. Communications in Statistics: Simulation and Computation, 46(3):2488–2503, 2017. CODEN CSSCDB. ISSN 0361-

Yang:2017:NIPb

[YY17b] Fengkai Yang and Haijing Yuan. A non-iterative posterior sampling algorithm for linear quantile regression model. Communications in Statistics: Simulation and Computation, 46(8):5861–5878, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:QRV

[YY17c] Jing Yang and Hu Yang. Quantile regression and variable selection for single-index varying-coefficient models. Communications in Statistics: Simulation and Computation, 46 (6):4637–4653, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yang:2017:RME

[YY17d] Jing Yang and Hu Yang. Robust modal estimation and variable selection for single-index varying-coefficient models. Communications in Statistics: Simulation and Computation, 46(4):2976–2997, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yu:2017:RLR

[YY17e] Chun Yu and Weixin Yao. Robust linear regression: A review and comparison. Communications in Statistics: Simulation and Computation, 46(8):6261–6282, 2017. CODEN CSSCDB. ISSN 0361-0918.

Yun:2014:SBI

[yYjHwK14] Won young Yun, Young jin Han, and Ha won Kim. Simulation-based inspection policies for a one-shot system in storage over a finite time span. Communications in Statistics: Simulation and Computation, 43(8):1979–2003, 2014. CODEN CSSCDB. ISSN 0361-0918.

Yang:2016:CDS

[YYY16] Hu Yang, Danhui Yi, and Chenqun Yu. Cluster data streams with noisy variables. Communications in Statistics: Simulation and Computation, 45(4):1381–1396, 2016. CODEN CSSCDB. ISSN 0361-0918.

Yu:2011:BAM

[YZ11] Chao Yu and Jingxiao Zhang. Bayesian approach to Markov switching stochastic volatility model with jumps. Communications in Statistics: Simulation and Computation, 40(10): 1613–1626, 2011. CODEN CSSCDB. ISSN 0361-0918.

Yu:2015:GEB

[YZCW15] Ching-Ray Yu, Kelly H. Zou, Martin O. Carlsson, and Samaradasa Weerahandi. Generalized estimation of the BLUP in mixed-effects models: A comparison with ML and REML. Communications in Statistics: Simulation and Computation, 44(3):694–704, 2015. CODEN CSSCDB. ISSN 0361-0918.

Yu:2019:BQR

[YZW19] Yang Yu, Zhihong Zou, and Shanshan Wang. Bayesian quantile regression and variable selection for partial linear singleindex model: Using free knot spline. Communications in Statistics: Simulation and Computation, 48(5):1429–1449, 2019. CODEN CSSCDB. ISSN 0361-0918.

Ye:2017:COO

[YZZ17a] Jane J. Ye, Julie Zhou, and Wenjie Zhou. Computing Aoptimal and E-optimal designs for regression models via semidefinite programming. Communications in Statistics: Simulation and Computation, 46(3):2011–2024, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Yu:2017:NEJ

[YZZ17b] Chao Yu, Xujie Zhao, and Bo Zhang. Nonparametric estimation of jump characteristics under market microstructure noise. Communications in Statistics: Simulation and Computation, 46(5):3575–3587, 2017. CODEN CSSCDB. ISSN 0361-0918.

Z:2018:AMC

[Z.18] Rezaei Ghahroodi Z. Analysis of mixed correlated semicontinuous and β-inflated Poisson responses. Communications in Statistics: Simulation and Computation, 47(10): 2992–3007, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zghoul:2010:NTB

[ZA10] Ahmad A. Zghoul and Adnan M. Awad. Normality test based on a truncated mean characterization. Communications in Statistics: Simulation and Computation, 39(9): 1818–1843, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zadlo:2017:PPS

[Za̧d17] Tomasz Za̧dło. On prediction of population and subpopulation characteristics for future periods. Communications in Statistics: Simulation and Computation, 46(10):8086–8104, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zougab:2014:BAN

[ZAK14] Nabil Zougab, Smail Adjabi, and Célestin C. Kokonendji. Bayesian approach in nonparametric count regression with binomial kernel. Communications in Statistics: Simulation and Computation, 43(5):1052–1063, 2014. CODEN CSS-CDB. ISSN 0361-0918.

Zardasht:2019:NTR

[Zar19] V. Zardasht. On nonparametric test for the residual probability order. Communications in Statistics: Simulation and Computation, 48(6):1840–1848, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2016:ISC

[ZB16] Bo Zhang and Tao Bi. Intraday serial correlation,volatility, and jump: Evidence from China’s stock market. Communications in Statistics: Simulation and Computation, 45(4): 1226–1239, 2016. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2011:MMF

[ZC11] Zhengcheng Zhang and Yingxiang Chang. MRL and MPL functions of parallel systems with INID components under double monitoring. Communications in Statistics: Simulation and Computation, 40(9):1514–1523, 2011. CODEN CSSCDB. ISSN 0361-0918.

Zhong:2012:NTT

[ZCJ12] Zihong Zhong, Weiquan Chen, and Hua Jin. A new test for testing non inferiority in matched-pairs design. Communications in Statistics: Simulation and Computation, 41(9): 1557–1565, 2012. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2010:CBC

[ZCR10] P. G. Zhang, D. G. Chen, and T. Roe. Choice of baselines in clinical trials: A simulation study from statistical power perspective. Communications in Statistics: Simulation and Computation, 39(7):1305–1317, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2019:ODS

[ZCY^{+}19] Shiliang Zhang, Hui Cao, Zonglin Ye, Yanbin Zhang, and Xiali Hei. An outlier detection scheme for dynamical sequential datasets. Communications in Statistics: Simulation and Computation, 48(5):1450–1502, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2014:NER

[ZDX14] Liangyong Zhang, Xiaofang Dong, and Xingzhong Xu. Nonparametric estimation for random censored data based on ranking set sampling. Communications in Statistics: Simulation and Computation, 43(8):2004–2015, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zeng:2017:CSC

[Zen17] Guoping Zeng. A comparison study of computational methods of Kolmogorov–Smirnov statistic in credit scoring. Communications in Statistics: Simulation and Computation, 46 (10):7744–7760, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2016:MCS

[ZF16] Jianping Zhu and Kuangnan Fang. A multistage chi-square test for measurement of the degree of association in twoway table. Communications in Statistics: Simulation and Computation, 45(4):1197–1212, 2016. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:FCU

[ZF17] Haibing Zhao and Wing Kam Fung. FDR control under block dependence. Communications in Statistics: Simulation and Computation, 46(8):6468–6474, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2016:ACB

[ZGC16] Guoyi Zhang, Maozhen Gong, and Yang Cheng. Adjusted confidence bands for complex survey data. Communications in Statistics: Simulation and Computation, 45(6): 1896–1904, 2016. CODEN CSSCDB. ISSN 0361-0918.

Zevallos:2017:RML

[ZGE17] Mauricio Zevallos, Loretta Gasco, and Ricardo Ehlers. Riemann manifold Langevin methods on stochastic volatility estimation. Communications in Statistics: Simulation and Computation, 46(10):7942–7956, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zghoul:2010:GFT

[Zgh10] Ahmad A. Zghoul. A goodness of fit test for normality based on the empirical moment generating function. Communications in Statistics: Simulation and Computation, 39(6): 1292–1304, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zitouni:2017:SGC

[ZGR^{+}17] Djamel Zitouni, Benjamin C. Guinhouya, Pierre Ravaux, Christian Vilhelm, Bruno Sarrazin, Mohamed Lemdani, and Hossein Mehdaoui. Si-GARCH: Construction and validation of a new method for the detection of breaking points in models. Communications in Statistics: Simulation and Computation, 46(4):3294–3305, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2019:BSN

[ZH19] Haiming Zhou and Xianzheng Huang. Bandwidth selection for nonparametric modal regression. Communications in Statistics: Simulation and Computation, 48(4):968–984, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2011:RCI

[Zha11a] Qiang Zhang. Repeated confidence intervals under fractional Brownian motion in long-term clinical trials. Communications in Statistics: Simulation and Computation, 40(8): 1130–1145, 2011. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2011:DFN

[Zha11b] Zhong-Yuan Zhang. Divergence functions of non negative matrix factorization: A comparison study. Communications in Statistics: Simulation and Computation, 40(10): 1594–1612, 2011. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2015:PBA

[Zha15a] Guoyi Zhang. A parametric bootstrap approach for oneway ANOVA under unequal variances with unbalanced data. Communications in Statistics: Simulation and Computation, 44(4):827–832, 2015. CODEN CSSCDB. ISSN 0361-

Zhang:2015:NII

[Zha15b] Xu Zhang. Nonparametric inference for an inverseprobability-weighted estimator with doubly truncated data. Communications in Statistics: Simulation and Computation, 44(2):489–504, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2015:GSB

[Zha15c] Junlong Zhao. General sparse boosting: Improving feature selection of L_{2} boosting by correlation-based penalty family. Communications in Statistics: Simulation and Computation, 44(6):1612–1640, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:EFD

[Zha17a] Haibing Zhao. Estimating false discovery proportion in multiple comparison under dependency. Communications in Statistics: Simulation and Computation, 46(9):6697–6704, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:NTM

[Zha17b] Junguang Zhao. A new test for the mean vector in large dimension and small samples. Communications in Statistics: Simulation and Computation, 46(8):6115–6128, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2018:FAL

[Zha18] Yanyan Zhao. Feasible algorithm for linear mixed model for massive data. Communications in Statistics: Simulation and Computation, 47(4):1126–1133, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zheng:2017:PFB

[ZHKR17] Wei Zheng, Juan Han, Weijie Kong, and Dewang Ren. The particle filter based on random number searching algorithm for parameter estimation. Communications in Statistics: Simulation and Computation, 46(2):1401–1413, 2017. CO-DEN CSSCDB. ISSN 0361-0918.

Zhang:2014:SAA

[ZHL14] Jiajia Zhang, Wenqing He, and Haifen Li. A semiparametric approach for accelerated failure time models with covariates subject to measurement error. Communications in Statistics: Simulation and Computation, 43(2):329–341, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:LRT

[ZHLW17] Jiujun Zhang, Chuan He, Zhonghua Li, and Zhaojun Wang. Likelihood ratio test-based chart for monitoring the process variability. Communications in Statistics: Simulation and Computation, 46(1):704–728, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zimmer:2017:TLU

[Zim17] Zachary Zimmer. Tolerance limits under Poisson regression based on small-sample asymptotic methodology. Communications in Statistics: Simulation and Computation, 46(7): 5836–5845, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2013:EEM

[ZJNG13] Hanhua Zhang, Hongfei Jin, Wei Ning, and Arjun K. Gupta. Edgeworth expansion of the moment-based test for homogeneity in an NEF-QVF mixture model. Communications in Statistics: Simulation and Computation, 42(10):2281–2294, 2013. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2018:ACA

[ZK18] Yaqian Zhu and John Kolassa. Assessing and comparing the accuracy of various bootstrap methods. Communications in Statistics: Simulation and Computation, 47(8):2436–2453, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2019:CIM

[ZKL^{+}19] Yunxi Zhang, Soeun Kim, Ye Lin, George Baum, Karen M. Basen-Engquist, and Michael D. Swartz. Comparisons of imputation methods with application to assess factors associated with self efficacy of physical activity in breast cancer survivors. Communications in Statistics: Simulation and Computation, 48(8):2523–2537, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2010:OCG

[ZL10] Qiang Zhang and Dejian Lai. Operating characteristics for group sequential trials monitored under fractional Brownian motion. Communications in Statistics: Simulation and Computation, 39(4):705–712, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:BCE

[ZL17] Guoyi Zhang and Rong Liu. Bias-corrected estimators of scalar skew normal. Communications in Statistics: Simulation and Computation, 46(2):831–839, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Zahra:2017:MAP

[ZLA17] Nighat Zahra, Supranee Lisawadi, and S. Ejaz Ahmed. Meta-analysis, pretest, and shrinkage estimation of kurtosis parameters. Communications in Statistics: Simulation and Computation, 46(10):7986–8004, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2017:MIM

[ZLC^{+}17] Xiao Zhang, Quanlin Li, Karen Cropsey, Xiaowei Yang, Kui Zhang, and Thomas Belin. A multiple imputation method for incomplete correlated ordinal data using multivariate probit models. Communications in Statistics: Simulation and Computation, 46(3):2360–2375, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Zou:2018:DTF

[ZLGS18] Qingrong Zou, Qi Li, Hao Guo, and Jian Shi. A discrete-time and finite-state Markov chain model for association football matches. Communications in Statistics: Simulation and Computation, 47(8):2476–2485, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2015:MCC

[ZLGZ15] Maoyuan Zhou, Liu Liu, Wei Geng, and Jie Zhou. Multivariate control chart based on multivariate Smirnov test. Communications in Statistics: Simulation and Computation, 44 (6):1600–1611, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2019:IAO

[ZLH^{+}19] Shaonan Zhang, Shanshan Li, Jiaqiao Hu, Haipeng Xing, and Wei Zhu. An iterative algorithm for optimal variable weighting in K-means clustering. Communications in Statistics: Simulation and Computation, 48(5):1346–1365, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2019:IWE

[ZLHW19] Yan-Yong Zhao, Jin-Guan Lin, Xing-Fang Huang, and Hong-Xia Wang. Iterative weighted estimation based on variance modelling in linear regression models. Communications in Statistics: Simulation and Computation, 48(9): 2599–2614, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:IPS

[ZLL17] Xu Zhang, Ji Li, and Yang Liu. Inference for probability of selection with dependently truncated data using a Cox model. Communications in Statistics: Simulation and Computation, 46(3):1944–1957, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2014:DCB

[ZLS14] Boan Zhang, Yunzhi Lin, and Zheng Su. Dichotomizing a continuous baseline covariate in randomized clinical trials with binary outcomes. Communications in Statistics: Simulation and Computation, 43(2):390–399, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zhai:2019:NCV

[ZLTX19] Min Zhai, Guolin Liu, Qiuxiang Tao, and Mingzhen Xin. A novel characteristic value correction iteration method. Communications in Statistics: Simulation and Computation, 48 (2):591–600, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:RBE

[ZLW17] Yan-Yong Zhao, Jin-Guan Lin, and Hong-Xia Wang. Robust bootstrap estimates in heteroscedastic semi-varying coefficient models and applications in analyzing Australia CPI data. Communications in Statistics: Simulation and Computation, 46(4):2638–2653, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:RMP

[ZLZ17] Li Zhang, Cunjie Lin, and Yong Zhou. A resampling method by perturbing the estimating functions for quantile regression with missing data. Communications in Statistics: Simulation and Computation, 46(8):6661–6671, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2016:ASR

[ZLZW16] Jiujun Zhang, Zhonghua Li, Qin Zhou, and Zhaojun Wang. An adaptive Shiryaev–Roberts procedure for signalling varying location shifts. Communications in Statistics: Simulation and Computation, 45(7):2511–??, 2016. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2012:SIV

[ZM12] Hui-Guo Zhang and Chang-Lin Mei. SiZer inference for varying coefficient models. Communications in Statistics: Simulation and Computation, 41(10):1944–1959, 2012. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2013:SSA

[ZM13] Xiaojian Zhou and Yizhong Ma. A study on SMO algorithm for solving ϵ-SVR with non-PSD kernels. Communications in Statistics: Simulation and Computation, 42(10):2175–2196, 2013. CODEN CSSCDB. ISSN 0361-0918.

Zhong:2017:IMB

[ZM17] Jianlan Zhong and Yizhong Ma. An integrated model based on statistical process control and maintenance for two-stage dependent processes. Communications in Statistics: Simulation and Computation, 46(1):106–126, 2017. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2018:PTD

[ZN18] Y. Zhang and S. Nadarajah. On partially truncated distributions. Communications in Statistics: Simulation and Computation, 47(6):1623–1631, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zargar:2016:TIB

[ZNA16] M. Zargar, H. Jabbari Nooghabi, and M. Amini. Test of independence for Baker’s bivariate distributions. Communications in Statistics: Simulation and Computation, 45(9): 3074–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Zavorka:2014:MSS

[ZP14] Shannon Zavorka and Jamis J. Perrett. Minimum sample size considerations for two-group linear and quadratic discriminant analysis with rare populations. Communications in Statistics: Simulation and Computation, 43(7): 1726–1739, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2014:SST

[ZQL14] Wei Zhang, Lianfen Qian, and Yunxia Li. Semiparametric sequential testing for multiple change points in piecewise constant hazard functions with long-term survivors. Communications in Statistics: Simulation and Computation, 43 (7):1685–1699, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2011:P

[ZS11] Bo Zhang and Benchang Shia. Preface. Communications in Statistics: Simulation and Computation, 40(10):1525, 2011. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:VSS

[ZS17] Xuejing Zhao and Jinxia Su. Variable selection for semiparametric proportional hazards model under progressive Type-II censoring. Communications in Statistics: Simulation and Computation, 46(6):4367–4376, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zawieja:2017:SSB

[ZSÁ17] Bogna Zawieja and Anna Szczepańska-Álvarez. A simulation study of the Bennett test and Miller test for coefficients of variation. Communications in Statistics: Simulation and Computation, 46(5):3701–3711, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:MPC

[ZSGW17] Yang Zhang, Yanfen Shang, Nini Gao, and Qing Wang. Monitoring prespecified changes in linear profiles using control charts with supplementary runs rules. Communications in Statistics: Simulation and Computation, 46(9): 7249–7263, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:CUW

[ZSY17] Hua Zhang, Guogen Shan, and Daniel Young. Comparison of unweighted and weighted rank based tests for an ordered alternative in randomized complete block designs. Communications in Statistics: Simulation and Computation, 46(6): 4452–4464, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:CEA

[ZSZ17] Yongjin Zhang, Youchao Sun, and Ming Zhao. A combinatorial estimation approach for storage reliability with initial failures based on periodic testing data. Communications in Statistics: Simulation and Computation, 46(4):3319–3340, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2018:ZCC

[ZSZW18] Yang Zhang, Yanfen Shang, Min Zhang, and Xiaofang Wu. Zone control charts with estimated parameters for detecting prespecified changes in linear profiles. Communications in Statistics: Simulation and Computation, 47(9):2749–2774, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zhu:2010:THT

[ZTT10] Liting Zhu, Xianming Tan, and Dongsheng Tu. Testing the homogeneity of two survival functions against a mixture alternative based on censored data. Communications in Statistics: Simulation and Computation, 39(4):767–776, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zuo:2019:NAC

[Zuo19] Yijun Zuo. A new approach for the computation of halfspace depth in high dimensions. Communications in Statistics: Simulation and Computation, 48(3):900–921, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:SSP

[ZVHC17] Yang Zhao, Albert Vexler, Alan Hutson, and Xiwei Chen. A statistical software procedure for exact parametric and nonparametric likelihood-ratio tests for two-sample comparisons. Communications in Statistics: Simulation and Computation, 46(4):2829–2841, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2014:RAU

[ZW14] Nan Zhang and Jinting Wang. Reliability analysis of unrepairable systems with k-out-of-m: G subsystems subject to suspended animation. Communications in Statistics: Simulation and Computation, 43(8):1900–1912, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zeng:2015:NML

[ZW15a] Michael Ruochen Zeng and Tony Siu Tung Wong. A note on the modified likelihood ratio test for homogeneity in beta mixture models. Communications in Statistics: Simulation and Computation, 44(10):2500–2512, 2015. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2015:IRC

[ZW15b] Haixiang Zhang and Dehui Wang. Inference for random coefficient INAR(1) process based on frequency domain analysis. Communications in Statistics: Simulation and Computation, 44(4):1078–1100, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2017:SDO

[ZW17] Zhenzhong Zhang and Wenlong Wang. The stationary distribution of Ornstein–Uhlenbeck process with a two-state Markov switching. Communications in Statistics: Simulation and Computation, 46(6):4783–4794, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2012:NLR

[ZWCC12] Lingyun Zhang, Dong Q. Wang, Min Chen, and Gemai Chen. A note on the likelihood ratio test for equality of k normal populations. Communications in Statistics: Simulation and Computation, 41(5):573–581, 2012. CODEN CSS-CDB. ISSN 0361-0918.

Zhang:2010:ESS

[ZWJ10] Lei Zhang, Jixiang Wu, and William D. Johnson. Empirical study of six tests for equality of populations with zeroinflated continuous distributions. Communications in Statistics: Simulation and Computation, 39(6):1196–1211, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2014:INI

[ZWZ^{+}14a] Chun-Xia Zhang, Guan-Wei Wang, Jiang-She Zhang, Gao Guo, and Qing-Yan Ying. IRUSRT: A novel imbalanced learning technique by combining inverse random under sampling and random tree. Communications in Statistics: Simulation and Computation, 43(10):2714–2731, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2014:VSB

[ZWZ14b] Xinxing Zhou, Kaibo Wang, and Lei Zhao. Variableselection-based epidemic disease diagnosis. Communications in Statistics: Simulation and Computation, 43(7): 1595–1610, 2014. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2012:EPT

[ZWZC12] Hualing Zhao, Xiaoxia Wu, Hong Zhang, and Hanfeng Chen. Estimating the proportion of true null hypotheses in nonparametric exponential mixture model with appication to the leukemia gene expression data. Communications in Statistics: Simulation and Computation, 41(9):1580–1592, 2012. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2010:CIP

[ZX10] Qian Zhang and Xingzhong Xu. Confidence intervals of performance measures for an M/G/1 queueing system. Communications in Statistics: Simulation and Computation, 39 (3):501–516, 2010. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2013:ELN

[ZX13] Peixin Zhao and Liugen Xue. Empirical likelihood for nonparametric components in additive partially linear models. Communications in Statistics: Simulation and Computation, 42(9):1935–1947, 2013. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:VOS

[ZX17] Guimei Zhao and Xingzhong Xu. The p-values for one-sided hypothesis testing in univariate linear calibration. Communications in Statistics: Simulation and Computation, 46(6): 4828–4840, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2015:PIE

[ZXJ15] Lei Zhang, Wenting Xie, and William D. Johnson. Performance of interval estimators for the inverse hypergeometric distribution. Communications in Statistics: Simulation and Computation, 44(5):1300–1310, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2016:EPL

[ZYHZ16] Yousheng Zhou, Yiping Yang, Jian Han, and Peixin Zhao. Estimation for partially linear single-index instrumental variables models. Communications in Statistics: Simulation and Computation, 45(10):3629–??, 2016. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2013:DML

[ZYM13] Ying Zhang, Hao Yu, and A. Ian McLeod. Developments in maximum likelihood unit root tests. Communications in Statistics: Simulation and Computation, 42(5):1088–1103, 2013. CODEN CSSCDB. ISSN 0361-0918.

Zhuang:2019:CEM

[ZYS19] Xinchen Zhuang, Tianxiang Yu, and Linjie Shen. On capacity evaluation for multi-state weighted k-out-of-n system. Communications in Statistics: Simulation and Computation, 48(7):2083–2098, 2019. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2017:SEP

[ZYW17] Xueqin Zhou, Lei Yang, and Xianyi Wu. On several estimates to the precision parameter of Dirichlet process prior. Communications in Statistics: Simulation and Computation, 46(4):3187–3200, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2018:HCR

[ZYX18] Li Zhao, Liang Yan, and Xingzhong Xu. High-correlated residuals improved estimation in the high-dimensional SUR model. Communications in Statistics: Simulation and Computation, 47(6):1583–1605, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2013:CPM

[ZZ13] Xin Zhang and Jiujun Zhang. CUSUM procedures for monitoring process mean and variability. Communications in Statistics: Simulation and Computation, 42(7):1455–1471, 2013. CODEN CSSCDB. ISSN 0361-0918.

Zhao:2017:MTI

[ZZ17] Xiaobing Zhao and Xian Zhou. Multi-type insurance claim processes with high-dimensional covariates. Communications in Statistics: Simulation and Computation, 46(1):500– 514, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2018:NEM

[ZZ18] Xiaoying Zhou and Feipeng Zhang. A new estimation method for continuous threshold expectile model. Communications in Statistics: Simulation and Computation, 47(8): 2486–2498, 2018. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2015:DFM

[ZZGL15] Maoyuan Zhou, Xuemin Zi, Wei Geng, and Zhonghua Li. A distribution-free multivariate change-point model for statistical process control. Communications in Statistics: Simulation and Computation, 44(8):1975–1987, 2015. CODEN CSSCDB. ISSN 0361-0918.

Zhou:2017:LFD

[ZZS17] Yingchun Zhou, Liping Zhu, and Nell Sedransk. Local functional data model for characterizing shape variation of similar curves. Communications in Statistics: Simulation and Computation, 46(6):4745–4759, 2017. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2011:NCD

[ZZW11] Jiujun Zhang, Changliang Zou, and Zhaojun Wang. A new chart for detecting the process mean and variability. Communications in Statistics: Simulation and Computation, 40 (5):728–743, 2011. CODEN CSSCDB. ISSN 0361-0918.

Zhang:2010:BEO

[ZZZ10] Gai-Ying Zhang, Chun-Xia Zhang, and Jiang-She Zhang. Out-of-bag estimation of the optimal hyperparameter in SubBag ensemble method. Communications in Statistics: Simulation and Computation, 39(10):1877–1892, 2010. CO-DEN CSSCDB. ISSN 0361-0918.

Zhang:2014:MSC

[ZZZJ14] Jieting Zhang, Minqiang Zhang, Wenyi Zhang, and Can Jiao. Model selection for complex multilevel latent class model. Communications in Statistics: Simulation and Computation, 43(4):838–850, 2014. CODEN CSSCDB. ISSN 0361-0918.